A Bibliography of Publications in Aquatic Ecology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

18 April 2025 Version 1.20

Title word cross-reference

< 200 [DDV91]. ^{+} [Raa89]. ^{14} [DB85]. ^{206/207} [WMM94]. ^{32} [JC83]. _{1} [Nwa95]. _{2} [APS81, CBP22, EAK11, FM10, LAC99, LWX19, RPS20, WYW^{+}24, Zev82]. 3 [MB81]. _{4} [BVM15, MB81]. : [PMZMJ16]. δ^{13} [BBM09, Fra00, TKR10]. δ^{15} [BBM09, TKR10]. γ [HYH03, VV93b]. ω [GSM07]. × [CB00a, GGT14].

-fixing [Zev82]. -HCH [VV93b]. -Proteobacteria [HYH03].

1-1-1971 [Ano70b]. 1-1-1973 [Ano72c]. 1-4020-1512-7 [Cad05e]. 1-4020-1804-5 [Ano05b]. 110 [Ano05b]. 16S [FO98]. 175/Hydrobiologia [Ano05b]. 1757 [AIS02]. 1960s [LSO10]. 1969 [Gel70]. 1975 [Van76]. 1990s [LSO10]. 1998 [vdV99].

2 [Gel70]. 20-year [Bra21]. 20th [Flo01]. 25^{o} [dSLM94]. 25th [Van76].

3D [CN94, HEZ03]. 3D-numerical [CN94].

44p [Har73a]. 48^{o} [dSLM94]. 4E [TJB98].

6-month [ZJN^{+}23]. 6-month-old [ACB12]. 62.00 [Cad05e].

90-m [KPS09].

A. [ACL13, Bak94, GAAT^{+}24, Har94, dS68]. Aan [Sch72b, Ano81a]. aangroei [Per69]. abandoned [dBPG00]. abbreviatum [Coe97, HP06]. Aberrant [MFS95]. abilities [MF07]. ability [CSA23, LMBM12]. Abiotic [MR11, CF09, FSN21, GDB19, Hei88, Mag00, PPK^{+}23, ŻKŻ10]. Abonnees [Ano72c, Ano70b]. Aboveground [HC09, LAM06]. Abramis [Caz82, Geu84, Raa89]. abrupt [BBD17]. absence [HPM20, Irv89]. absorption [PMSS18]. abstract [Beu74, DV74, dL74]. Abundance [KiNS05, KRATA20, LM92, MBM^{+}22, SAB05, UKS05, BSA24, BBM09, BZL22, BZG^{+}22, BHPT20, CMV91, Cra91, DRN09, DGGW20, dSDMD20, ECGC25, FWS^{+}23, FRP08, FCDA09, GJW20, GDB19, GD91, HJ23, HNiN03, KGM13, KGM22, KP24, Lee84, LNHN08, LMD11, MAA22, OSO09, PPH16, RFW17, SL07, SBLdAM23, The02, TGV00, VHN00, VH08, WB10, YYS^{+}23, ZTG10, ZYL22, dSCT23]. abundance-frequency [Lee84]. abundance/biomass [Cra91]. Abundancy [Lin78]. Abundant [GBJM22, BBP08, IRT06, MBM^{+}22, YL11]. Academic [Ano05b, Cad05e]. Acanthodiaptomus [DK80]. Acari [MSW10, VS96b]. Acartia [BTV88]. accentuated [GB07]. Acclimation [DZG17, TPS09]. according [Dar73, DB85]. account [Har94]. Accumulatie [KvdM70]. accumulating [Cad78]. accumulation [DBO11, GC08, Lij86, Pus94, SMPK09, KvdM70]. Acentria [JGH97]. acetic [EAK11]. Acid [Gos99, EAK11, GSK08, HFS08, IM06, KMS82a, KMS82b, KCR92, MMM11, MKG03, NAT08, Pay10, ST09, ŠNZ21, TBA03, TSZ10]. acidic [MD92, Gos99]. acidification [ACB12, Bra01, HXR09, SMR08, SL24, SG08, TK19b, TK19c, Van96, vDSB80]. acidifying [BPD15, Wor90]. acidophilic [ČN10]. acids [KKI04, MCvE09, MKA79, SGK07, VK12, WLB22, WBB00, ZWY^{+}24]. Acipensor [WXW11]. Acknowledgements [Ano91]. acoustic [CJS^{+}23, GCN^{+}23, HHP12, PJH06]. Acoustical [GJ06]. AcousticIA [GCN^{+}23]. acquisition [FSL^{+}23, MLG20]. Acroloxus [HVS87]. Acropora [NYT09b, SS24]. across [AC16, BGM08, BSW^{+}24, CGC19, CGL21, HLK22, MCV08, PF07, PCGW20, RBN20, RQ93, RAA98, TLC93, VEK20, WMV09]. act [KJP14]. Actinocyclus [RSK93]. action [EMN04]. actions [SdAdC^{+}25]. Active [LWM89, OI89, DCG19, GMVCLGGP24, HWK11, MSK93, TJB98, vdB70]. activities [AAA^{+}24, BGRM^{+}24, EPL^{+}23, Gul89, TB92, Vos82, VM25, nWdLjZ21]. activity [BNV03, BLP16, CLRdSR23, CB00c, EHM91, GZG02, GX11, Her83, HPPL00, HPT95, KS94, Koo76, KMG^{+}23, LAC99, MAM13, Nau00, NN05a, NN05b, ÖE04, RCSF93, SWB13, SvFN15, ŠV09, TPT22, TW03]. actual [Dew80]. aculeatus [BM98, HBL13, SDRM16, VRR17]. acuminata [XWC15]. acuta [BA14, HD21, ZKN09a]. Acute [GFSN04, DDZ^{+}24, QLW^{+}23]. adaptability [GYW17]. adaptation [GRL20, HBR12, TJZ^{+}24, VM25]. adaptations [DZL21, NZS16, Sch97b, ZBA19]. adaptative [ABF21]. adapted [ZLT22]. adaptive [Vij91, VRR17]. addition [BVI16, BVV80, RKS14, TZD21]. additional [DK80]. additions [BBM09]. additive [MKV15]. address [Gin80]. adenosine [Koo76]. adequacy [SRF21, dSSR^{+}23]. Adige [SB08]. adjacent [SHV96, vdHD17]. adjust [ZCA20]. Adriatic [Lov74]. adsorbents [DHL16]. adsorption [TW03]. Adsorptive [MD92]. adult [HF97, KAN20, NNN22, PC14, Pen00, DAO24, RMH21, SS92]. ADV [PJH06]. Advanced [LZH^{+}25]. advantage [BWS19, CYW24, DF11, OLM09]. advantages [SJD22]. adverse [LAZ21]. adversely [PKR12]. Aedes [RHO09]. Aegagropila [BI09, SYW09]. Aegean [Bal09]. Aegiceras [QLZ20, QZL^{+}24]. aegypti [RHO09]. aerial [Del92]. Aerobic [OBC82, Sep79]. aeruginosa [BHB01, DZT13, dSFFA03, JJP24, KBS05, MF07, MHG16, NR97, Nan00, PMZMJ16, YP17, YZY09, ZWY20]. Aeshna [RIT04]. aestuarius [FvM03]. aethiopica [PDD07, PCD11]. afdeling [Web71a]. affect [BCP10, BKG16, CLRdSR23, DCC20, GCE11, LWY^{+}24, MKT23, PKR12, SAB10, WZL18, ZWF11, ZJN^{+}23]. affected [Bol04, TDBW95]. affecting [BH22, MR11, Mit78, PdE^{+}24, SAACR09, WMV09]. affects [ACB12, dSDMD20, GKM22, KÇA22, LMD11, NMdAM21, RFW17, SZH20, VRR17]. affinis [GD91, Leh04, NCT91]. afforestation [CPA15]. afforested [LWO01]. Africa [TAN14, AANO25, CDW92]. African [ABF01, BPA01, BBR01, FAK01, Flo01, FDP01, FC93, KEL09, KKO08, KHF01, PBR01, PJF01, RFP01, REP01, RFF01]. afrotropical [CDW92]. after [Doo82, FdLY^{+}24, FWR^{+}23a, FWR^{+}23b, Gul89, MLG95, NNB^{+}24, PdCF^{+}25, RCSF93, SMPK09, SSV21, SDN25, Vaa79, Ver80a, WPvB17, YF18]. afval [Ros69b]. Afvalwater [De 77b]. against [BPD15, Gib86, GH04, MNS05a, MNS05b, RIT04, Van97, VvGV12]. agardhi [VM79]. agardhii [Ber75, Bij75, VZM75]. Agassiz [Bak80]. Age [DLG20, DF83, OTJ19, Dar73, DDL19, GX11, KW12, KHF01, SH16]. age-0 [KW12]. age-related [SH16]. ages [LOZ24]. aggregate [KS94]. aggregates [ZWY20]. aggregation [WXZ12]. aggressive [JCF17]. agitation [RVK96]. Aglaodiaptomus [DM04]. agonistic [HKB19]. agricultural [EAO20, JGUF23, KGM13, KGM22, dCNGF25, OTJ19, SLL09]. agriculture [DN23, LRMdF19]. aid [DL89, Jan75]. aim [CBdO^{+}24]. airborne [Alt98]. Aires [SSI21]. akinete [PLdFC22]. akinetes [KIG06]. Al [AAMS24, TKM03, UT16, WvRvdV03b, XCL10]. Al-Namil [AAMS24]. Alagoas [NCDS^{+}25]. Alalay [AAV07]. alarm [THS17a, THS17b]. Alaska [FM99]. Alaskan [CS13]. alba [PB84]. Albemarle [BB23]. Albert [DVZ94].

Alberta [IJ06]. Alcedo [NBT21]. Aldehydes [Jüt05]. alder [VK12]. alewife [PHD13]. alga [Bij75, DNG22, DCL19, HL95, KA97, LL77, MMV05, SMGAGG00, SYW09, WBB00]. Algae [WNS08, AMSN07, Ano83b, BHM03, Bor73, Coe77, dNFdNM21, Gon79, Hil73, HK76, HV79, HW13, Kap76, Kla80, KM17, LB10, MKA79, MMV09, Nie73, PS06, PIM95, PB04, PDvdV06a, PDvdV06b, RM81, TKR10, TNC08, VKK73, WKT05, WTC09, YZ08, ZD21, lCdB70, vD97]. Algae-based [WNS08]. Algal [dSDMD20, Los80, Som83, AZD02, AC16, Bir78, BB23, CPA15, FGF13, GFSN04, Har73b, JBS05, MNK04, Mun94, NS07, PR84, Pol75, SAN99, SAGN07, SB04, SHP21, ŠV09, TBA03, TWC13, TBS18, VPP16, VPM82, VP77, WP12]. algal-bacterial [AZD02]. Algeciras [SMGAGG00]. algen [lCdB70]. Algeria [KND^{+}22]. algological [ÁBZ08]. algorithms [WSvA98]. alien [BKG16, GBPR23, MTC^{+}24, Pin75, SAR21, WHS13, dPBCP22]. alkaline [DDV91, Nau00, NN05a, NN05b]. alkalinity [GPS^{+}23, PKR12]. Allegheny [PKE22a, PKE22b]. allelochemical [TPT22]. Allelopathic [DCG19, LRSN22, MMV05, BV14, JJP24, MCB22, XWC15]. allelopathy [GXY14, MMV09]. alleni [AP00]. alleviates [WCD17]. allis [BFS21]. allmanni [Spa71]. allocation [GOB24, KA97, SXY20]. allochthonous [KTJ10]. Allometric [ABF21, PHM21]. almost [WKH82, WA84]. aloides [MMV05, MMV09, RIT04]. Alonella [NSK08]. Along [Gys72a, AT24, ABA^{+}24, BJS22, BTP24, CPM09, Dav93, Den94, HvAV83, IUK22, KND^{+}22, KFS04, KRATA20, LKSK15, MHE93, MDH93, MJJ21, NvdVdlM01, NEP05a, NEP05b, PFT12, RBPF11, RSM24, SCN^{+}24, VvdMP12, VIM17, WZL18, WiNDF22, YS98, YMB93, dW71b]. alpine [BBD17, Bla82, Bra99, FWB05, GA10, GDT08, MSW10, ZBA19, GSB13]. alpinus [AK09, Bra99]. Alps [MB06, MB06]. alter [BBF17, KMR20, SH16, SdAdC^{+}25]. Alteration [ÓP04, TW03]. alterations [COC18]. altered [OFCP22]. Alternanthera [LWY^{+}24, SJD22, YF18]. alternating [MOPCP08]. Alternative [Sch89]. Altiplano [JM08]. altitude [BN24, WMD07]. altitudinal [VTE19]. Aluminium [BZL22, DDV91, VdJ94, BZG^{+}22]. Alvarado [RRRRA07]. Alvarez [Art99]. Alvarez-Cobelas [Art99]. Amarasinghe [Gul09b]. amarus [DBO11]. Amazon [DMD19, NMdAM21, TBSZ22, dRFC21]. Amazonian [AFC20, DEP22, PBM21, SBLdAM23]. amazonicum [PNC22]. ambient [VRR17]. Ambystoma [GK07]. Ameiurus [Pad13]. Ameland [Hol70, Hol70]. America [SC09, Van78b, BZ97, Pad13, SMS^{+}24, Cad01]. American [DH22, EK93, FvM03, Hey92, RJ11, SM12, ZLC12, dLFTB22]. americanus [ARG20]. amid [KÇA22]. amino [KKI04, KMS82a, KMS82b, MKA79]. amino-acid [KMS82a]. ammonia [HWS22, JDJ^{+}25, LSY20, RCB95, YYS^{+}23]. ammonia-oxidizing [HWS22, YYS^{+}23]. ammonium [CVG03, FV95a, GC23, PMSS18]. amoeba [Pay10]. amoebae [ASLT15, VvGV12]. among [BDF20, BHG18a, BHG18b, CMH14, Dew80, FKB92, Fra00, FGM20, GSH17, INJ20, IIMRC21, LB10, PU15, Pri03, SRF21, SJ00b, TCH18, TH16, Vaa75, WZL19]. amount [SJD22]. Amphiagrion [HP06]. Amphibia [OR11]. amphibian [Bri99, DF11, GRDPP09, MM96, OTJ19, QCF19, TNtK00]. amphibians [NGSOAD20, PVH17]. Amphichaeta [Mas88]. Amphidinium [RP22]. amphidromous [JBS05]. amphimelas [MOM24]. amphipod [BLP16, CB16, DRM18, Den74, IDP19, JTL09, KJP14, dSLML21, Pin75, PP95, PDvdV06a, PDvdV06b, Wil07]. Amphipoda [BHG18a, BHG18b, FvM03, DGD06, Dor74, Dor77, Kle84, KKM09b, Pin75, YTD02, ŽG09]. amphipods [BG15, FvM03, KML02, Leh04, LA19, MSB10, PK08, TET23, TPS09]. amplitude [Rat14]. Ampullariidae [GD09]. Amsterdam [Ano83b, Gel70]. Amur [RPN15]. Anabaena [KIG06, TK88a]. anaerobic [VV90]. analog [GCT^{+}23]. analoga [FCPCMB^{+}25]. analyses [BPA01, KSY19, KTJ10, MMM11, MDP08, Moe71, Rat17, dCdCP19, vIR92]. analysing [SCC19]. Analysis [BTP24, DBdF02, Fli85, GGD07, Nwa95, PRW08, ABA^{+}24, AMTSJ15, BPD15, CCS09, DDF93, DGH^{+}24, DH09, DL89, DV74, FDB97, GM93, HK82, HF08, HH80, IÓG04, JBB07, Jan75, KGS02, LTH22, MNJ21, NBT21, PFLO14, PV95, PMD10, RHA00, RB82, RL05, ST09, SMR08, SCS12, TRHHAR10, UOJGRL21, VP12, VB98, VM25, VGV94, WWS21, WRS05, vdTS97]. analytical [HCS02]. analyzed [MKV15]. analyzes [Moe71]. Analyzing [DGBL15, SJ00b, DT23b]. Anax [Cru10]. anchialine [CCH11]. anchovy [IT09]. ancient [BVM15]. Ancylidae [Van91]. Ancylus [HVS87]. Andean [AAV07, BMR07, CMB07, IAHB18]. Andhra [BJS22]. Angewandte [Gos98b]. angiosperms [JC80, PSZ79]. anglica [LvBR06]. Anguilla [ACL13, DTC19]. angularis [GFSN04]. anhydrobiotic [RC98]. Aniene [TBF09]. animal [CRR22, NS07]. animals [BT92, GC05a, GC05b, HDF88, RC98, Vaa75, Wol00]. annandale [SC00]. annandalei [BJS22]. Announcement [Ano95a, Ano99a, Ano02]. Announcements [Ano96, Ano68c, Ano69a, Ano72d]. Annual [Ker81, KK89, KBO74, SMC21, BLB^{+}23, HDF88, LFR75, MMBV24, SLG09a, SEE22, Pod74a, Pod74b, Pod74c]. annuel [Pod74a, Pod74b, Pod74c]. annular [CLH06, NLC06]. Anomopoda [MJEC05]. Anomura [FBNF08]. anostracan [YP17]. anoxia [LMBM12]. anoxic [JTGS24, RZG09]. Antarctic [HH13, WHP76]. Antarctica [APS81, WNZ24]. antennatus [KK09]. Anthology [vdV99]. Anthozoa [SPB^{+}24]. Anthropogenic [MCC21, SdAdC^{+}25, AAA^{+}24, CBF19, EPL^{+}23, JVCJ25, LPT21, MOM24, TK95, ZPD06]. anthropogenically [QSA09]. anti [KK11]. anti-predator [KK11]. antibiotic [LMP02, MRR10]. antibiotic-resistant [MRR10]. Antibiotics [ZWF11]. anticyclonic [MML^{+}22, MML^{+}23]. antillarum [SHP21]. antioxidant [xHCjX^{+}23, LSY20, LYM21, QLW^{+}23, VM25]. antioxidative [TPS09]. antipodarum [HKT16, HBR12, LDK20]. Antipredator [Lin07, SG20]. Anuraeopsis [SAGN07]. Anzali [MKG20, MNJ21]. Apatani [SD09].

Apennine [DCC20]. Aperiodic [GKK95]. Apophallus [RVB08]. apparatus [Goo79, HH80]. apparent [BRG02, CB13]. Appendages [RR03]. apple [DD17, GD09, GFW24, MV14, SMM23, SM12]. applesnails [YB08]. Applicability [Cra91, PMSS18, SRH08]. Application [AVD07, AR09, De 80b, DT23a, KSY19, Los80, SD86, SKV02, SCC19, BVM15, DT23b, DHL16, HVB97, Hos80, LCS01, LXH21]. Applications [GDD07, GV89, Her83, LZL22, Goe07, SNKN^{+}24]. Applied [Meu91, DB85, RBAB24, Rav01, Roi88b, Van82a, WMM94]. apply [SA95]. Applying [JJP24, DDPT24]. appraisal [Gul07b]. Approach [HG05, Kal86, AdAPC20, Bro98, DdNM22, DDT24, ES09, GPV19, JGM^{+}23, KW12, KS94, KCG05, MGVC91, PPCH22, ST09, SNO98, SJ00b, VPP16, dSSB24]. approaches [FDG97, KT15, MTJ10]. appropriate [SKRB09]. Aquaculture [Sar05e, Sar05f, QLZ20, QZL^{+}24, Shi98]. aquae [KIG06]. aquarium [PGM^{+}24]. Aquatic [Ano97b, Ano97c, Ano97d, Ano98e, BV96, CH83, FCG16, Flo01, Gro22, JVCJ25, MK05b, PCD11, TB22, ŻD06, Adm90, AdAPC20, Ara01, BVI16, BSMVP24, BCN12, Bel87, BA97, Bes76, Bes81, Beu74, BT92, Blo70, Bol04, BvdH90, BZL22, BZG^{+}22, Bro98, CAPGA08, CGO19, CZC14, CKM19, CBE20, Coe77, Coe97, CH12, CRR22, CMBG20, DWB10, De 79c, EAK11, FDG97, FWR^{+}23a, FWR^{+}23b, GDB19, GPV19, Gra12, Gri89, HCS02, HBB13, HS85, HS86, Her83, Hig80, HSY18, JC80, JC83, JTGS24, JAM15, KSY19, KK12b, KMR20, Ker97, Kir22, KPJP23, KH92, KÇA22, LCS01, LMBM12, LZL22, LZY21, MLTH15, Mos97, MMV05, NNN22, NLF10, NZ98, NRS19, OTJ19, PPY^{+}22, PME08, PDE08, PFT12, PMM10, Pip87, PB04, PSZ79, PKR12, RMR21, RASA13, Rin76, Rin93, RU87, RAN24]. aquatic [SKW21, SSV21, SKRB09, SJ00a, SSN24, SLL09, SdAdC^{+}25, Sur01, TK89, TMM21, TG19, UET12, VNSFG16, VSA97, VBK96, WBR17, WRH98, Wet93, WvLA14, WRS05, Wol97, YAC21, ZWY^{+}22, Gul09b, Gul07b]. aquaticus [Ara01, HJV08]. aquatisghe [Blo70]. aqueous [SGM22, ZWY^{+}24]. aquifer [MBM15, NHF96]. Arabian [AAMS24, DGN21]. arable [SEB11]. archaea [FO98]. archaeological [Jan75]. archeological [BH82]. archipelago [KBN20]. Architecture [Roo79, FGF13, LMD11]. Arctic [AK09, GA10, Gud04, KJS15, SMJ03, SSK14]. Arctodiaptomus [Bra99, TSZ10, Tol02, ZG02]. Area [GWF92, WB10, BB78, Bol04, CGL21, Coe78, CB94, CDD20, DV92, HSCR10, KK07, Kla88, KCR92, KÁ04, MO92, NB22, OSR88, Pee74, PEREBM^{+}23, Rin81c, SWW02, SEE22, SKH19, SDW^{+}24, Vaa75, Van89, VBD96, WMTR07, Zim78, dHC76, vG84, vdVP76]. areas [CV95, Dro84, GCS93, MM95, Mor99, PDD07, RP92, RPSSDS23, RCB95, SMC21, SBLdAM23, SLC10, TNtK00, TNtK03, YRR09]. Arembergergracht [Sch68a]. Arenicola [Van82b, SHK92]. Arequipa [FCPCMB^{+}25]. Argentina [SC09, ARP14, CTO01, CB00c, GBZ21, GL01, SSI21]. Argia [HP06]. arginine [GMVCLGGP24]. arguinensis [DGGW20]. Ariake [IT09, ISY10a, ISY10b]. arid [BKD^{+}24, FdLY^{+}24, dCMMB15, MdMdLB21, RBN20, RTW17]. aridland [RAD21]. Aristeus [KK09]. aromatic [OMNC24]. arriving [RRK15]. arsenic [LZZ^{+}25]. Artemisia [xHCjX^{+}23]. artesian [RFW17]. artifacts [Bre85, ZD21]. Artificial [Bol04, VIH16, ZYS14, ASA06, CvDG07, DT23b, GGD07, GFZ15, GDD07, HB75, HV79, JBB07, LWM89, Lin78, MS08, MBMV20, NC20, PT14, SGM99, SB88, STH21, TXW^{+}24, XWJ22, dPBCP22, SOW01]. artisanal [MOM24]. Arts [Gul99]. Arunachal [BN24, SD09]. arundinacea [CDL14]. Ascertaining [RSM24]. ascidian [KHG20]. Asellus [Ara01, HJV08]. asexual [RAP22, YYR20]. Asia [PVC96]. Asian [Gul09b, DTH20, PHM21]. Asiatic [ZWJ19]. Asimple [DB81]. aspecten [Dee69]. Aspects [CY95, DWV81, Mit74, BHV82, Cap75, Dee69, FBNF08, Par78, Roi88b, RB78, Sha78, TD87a, VS96b, WWS20, Wil80]. Asplanchna [HLSN24, STHN03]. assays [Rat17]. Assemblage [RR03, ATB07, BDG10, DSG^{+}24, DSL12, Fli85, GTI99, IRT06, KND^{+}22, MKV15, RB08, SC09, VN22, VHB09]. Assemblages [RRRRA07, BD74, Bal09, BR07, Den94, DMD19, dSDMD20, ERA17, ED95, GB07, Hig09, JTL09, KK07, LHV07, dSLML21, Mag00, MS08, MCC21, MKT23, MB14, MTS11, NWL11, NHF96, dCNGF25, NMP12, PU15, RSM24, RAD21, RHH17, SVM08, SVK94, SVW95, SLL09, SBS12, TWC13, TBSZ22, TH16, ŽAL24, ZPW15, dCAP23]. assemblies [De 84b]. assembly [ASLT15, MKS22, RRL09]. assess [ASLT15, ABF21, CTO01, CSV06, PLM^{+}22, SVB98]. assessed [MNJ21, NBT21, vGVV11]. Assessing [Bro98, CDD20, EBM18, FBS19, Gib86, Soi05a, Soi05b, WXW11, DT23a, HS85, HS86, SEP09]. Assessment [BDR97, BSF11, DGH^{+}24, DDPT24, STT01, TPT22, WWG^{+}25, AdAPC20, Ano05b, BKC11, Bra01, CGC19, CS13, DH01, DTG09, GL01, Hig74, KWO11, LCS01, Lah98, Lee84, LTH22, MAS14, MD05, Mol80, MBFD24, PGM^{+}24, PDD07, Sch74, Sch83, SB88, SDW^{+}24, TG88, Van74, VDB81, WDA15, ZD21, dL74]. Assessments [AOS^{+}24, BKC11, LH93]. Assignment [MD05]. Assimilation [SFB04, GC05a, GC05b, LW11, dCAP23]. Associated [RRRRA07, AAR^{+}24, BM03, BJ99, BSA24, BR07, CBdO^{+}24, GJW20, IAHB18, JTL09, JGH97, KH92, Kor98, Kor14, dSLML21, LRMdF19, MLB16, PGSL02, PEREBM^{+}23, PDD07, PCD11, PHD13, RSM24, Rod10, SC00, VNKdOR15, VNSFG16, WKT05, dSCT23]. Association [KK07, MLTH15, DGN21, GUT21, WGK95]. associations [DGGW20, HD79a, HvdGvS10, PAAJM17]. assortative [Wil07]. assumptions [BKC11]. Astacus [Gee75]. astaxanthin [DCL19]. astaxanthin-producing [DCL19]. Asterionella [BR87]. Asymmetrical [HXR09]. asymmetry [KAN20]. Atherina [vdVP76]. Atlantic [RMS11, ABF21, Baa80, Cad08, GA10, GBJM22, HdC24, LLL22, MFTM20, PFSPdS24, RAO^{+}24, VNSFG16, VSM19a, dNdSBJdCeS21]. Atlas [Gul98]. atoll [dSDLdA22]. ATP [Koo76]. atrazine [HAB98]. attached [WZL19, YZY09]. attempt [YO02]. attention [KP97, Sim94]. attenuation [BP94, Bui95, De 00, GEK97]. atthis [NBT21]. attitudes [VBA22]. attractant [RHA00]. attributed [Kae20]. attributes [NMP12]. attu [IRT06]. Atyaephyra [VV85]. Atyidae [CCH11]. Aubl. [HHEM19]. Aude [GL92]. August [Van76]. Aulacoseira [JGG09, MF07]. Aurantiochytrium [SNKN^{+}24]. Aurelia [RAP22]. auricularia [LWX19]. auriculata [HHEM19, PLdFC22]. aurita [Baa85]. auritus [TE04]. AUSRIVAS [STT01]. Australia [DWB10, WB10]. Australian [RTW17, RFW17, STT01]. australis [AH82, BZD81, BZD82, CBE20, GYW17, HHY21]. Austria [DH09, GDT08]. Autecological [Baa79, Baa81, Baa85, Baa88a, Baa88b]. Autecology [SK09]. Author [Ano03a]. Authors [Ano98a, Ano04c]. autochthonous [dCAP23]. automated [Sch78]. automatic [CJS^{+}23]. autonomous [Ald79]. autostabilization [Deg10]. autotrophic [GCE11, HNiN03, VMS09]. autotrophs [CMB07]. autotrophy [BWH02, DWA11]. autumn [CVM21, MS02, TGD16]. autumnal [BdSCH09, NSK08]. Auvergne [DK80, Dor76a]. availabilities [KPV16]. availability [CB00a, DWV81, DYS86, HEPH09, HH13, INJ20, LCPM15, Rat17, WXL16, WMD07]. available [GEK97]. Aveiro [CM95b, MQC93, PDM95, AA95, ACA02]. avert [WWS20]. Aves [TE04]. avoid [KLG17]. Avoidance [DK80, PK08, Sch97a]. avoiding [FAC16]. avoids [MMBV24]. awatschensis [nWdLjZ21]. Axarus [Fer92b, MF92]. Axios [DDD07]. axis [VvdMP12]. Azolla [SK84, TK88a, TK95]. Azolla-Anabaena [TK88a]. Azores [WNS08]. azote [Pod74a, Pod74b, Pod74c]. Azul [MT09]. azurea [ATB07].

B [Cad01, KA97, KT09, Nie98, SMC08, Van97, War97]. baby [Sch88]. Bacillariophyceae [MNT16, RSK93, Ste94, WP91]. back [BKO92]. backbarrier [FN94]. background [Gib86, NRS19]. backscatter [BP94]. Backwater [Wel74]. Bacteria [VV05, BVM15, BKW94, BT01, BASJ94, Cap75, CVM21, GBJM22, GC23, HPS11, LKM01, LMP02, MRR10, NYT09b, RZG09, RDE93, Sch77, WZL19, YZY09, dWvG88, vG72]. bacteria-sized [GBJM22]. bacterial [AZD02, BSA14, BEZ91, CZZ^{+}25, GGT14, HHY21, Joi73, KIJ11, KD82, KKR02, LTL09, LDL11, MGL22, Nau00, NMB09, NN05a, NN05b, OMV14, RKS14, RBN20, SL07, WBR17, ZWY20]. Bacterien [vG72]. bacteriennes [Joi73]. bacteriological [ÁBZ08]. bacterioplankton [ACA02, BLS88, BLG99, CDA03, FRP08, MML^{+}22, MML^{+}23, PAP24]. bacterium [DZT13]. bacterivory [KKR02, NMS21]. bad [MCvE09]. Baden [Mar01]. Baetis [IGS10]. bag [BHV82]. bags [TFH^{+}24]. Bahamian [SZH21]. baía [dSPdS10]. baicalensis [JGG09]. Baikal [JGG09, KiNS05, UKS05]. Baird [NSK08]. Bakker [NL91]. Balance [NMB09, Pod74a, Pod74b, Pod74c, VML93]. balances [Kal86]. Balancing [AL24]. Balaton [PPV13, VMS09]. Balbiani [FKB92]. Balgzand [Cad78, Van78a, Zim78]. Balili [SDN25]. ballast [HVB97]. balls [BI09]. balthica [dW71a]. Baltic [Leh04, Ole97, Alb04, BHM03, BSB97, Cad05c, EMT00, EMN04, GTPH08, KIJ11, KSBR94, KSBŁ95, KML19, Kow94, KP97, KHK17, LMV09, Nau00, NEP05a, NEP05b, Pus94, SN03, Sch97b, Sör82, VET20, Zet97b]. baltica [EMT00]. Banff [HP06]. Bangladesh [HAS22]. bank [BEQP18, Bel87, DHPN^{+}24, dSSSES21, Sto76, VSB10]. Banks [Gys72a, HEB00, NMdAM21, PVB20]. Banyoles [PR92]. barb [TP09]. barbel [HPT95]. barbs [SNO98]. Barbus [HPT95]. Barguzin [KiNS05, UKS05]. barnacle [MSB15]. barrages [Sha78]. barreimiae [TP09]. Barrett [Ver97]. barrier [BKC11]. Barriers [BKC11]. base [ODI08]. Based [HG05, BCN12, CCM21, ECB09, HM84, HHC21, PPV13, RJ11, Roe96, RGD10, SOW01, SEB11, SR88, TSV16, VHC92, VN22, VGV94, WNS08, ZWY^{+}22]. Baseline [FMdCFdCV24, dSDLdA22, PJF01]. Basics [Sta06]. Basin [GB82, KEH09, Ole97, PDD07, ŠDT^{+}24a, ŠDT^{+}24b, WiNDF22, BFB15, DSL12, Joi73, MCC21, OMM05a, OMM05b, PdCF^{+}25, Pod74a, Pod74b, Pod74c, Rat14, SGRMP20, SSI21, WFFM10, YKU01, YBZ20, ZYL22, NMdAM21, TJZ^{+}24]. Basin-scale [KEH09]. basing [BvdH90]. basins [AvdV90, KHD99, PPV13, TH16, VHN00, Osk72]. basis [DBdF02, MVI16, SMJ03]. bass [GS11, SXC^{+}25]. bassin [Joi73, Pod74a, Pod74b, Pod74c]. bat [RHO99]. batch [JB83, PPH16]. bath [Sch88]. bath-water [Sch88]. Batrachochytrium [NGSOAD20, dNdSBJdCeS21]. Bay [AC86, BDR97, FM99, IT09, ISY10a, ISY10b, KiNS05, KHG20, LKM01, LKE94, NSB24, PHD13, SMGAGG00, SHV96, STH21, TGV00, TBF99, UKS05, YO02, Bar21, dSLML21, NvdVdlM01, WCZ11, XCL10, CP10, CAFA04, FR09, GBJM22, GW73, GW74, Haa74, Lin72, NudWB^{+}25, RFO94, RDE93, TYN02, YHM05, ZSD94]. Bayesian [MKV15]. bays [PIM95]. BC [Ano81a]. bdelloid [Dev09]. be [LLX11, MRR10, MCSV21]. beach [Nie73]. beaches [BGL13, FCPCMB^{+}25, FA99]. beam [BP94, FWS^{+}23]. beaming [DTG09]. bearing [Roi88a]. bears [ZBA19]. beaver [WKK24]. Beavers [Bas20]. become [Zev82]. bed [ACA02, DR96, FD94, KKS11, NBD^{+}25]. beds [Bak80, CM95b, HSCR10, SHV96, WWG^{+}25, WAW92]. beetle [MB14]. beetles [DB02, PMM10, VNSFG16]. before [DR96, Doo82, SDN25, Vaa79, Ver80a, ŽG09]. behavior [BDP21b, MLTH15, Pen00, PK08, RHS18, SM12, TNC08, ZWJ19, dCdCP19, dG71]. Behavioral [HLSN24]. Behaviour [GCS93, VV93b, ÅDB00, AD89, BT92, Cra85, CBNF19, De 76a, ECB09, FTH22, GBPR23, Irv86, KK11, MEB93, MFP01, NPH03, PdE^{+}24, RR98, WBS22, WBS24]. Behavioural [BJ99, NZS16, TET23, HS85, HS86, SHK92]. behaviours [GBA20]. Behoud [Gys72a]. Beij. [DCG19]. being [CH83]. Beira [DGH^{+}24]. Bekannten [De 79a]. Belgie [Dum69, dB69]. Belgique [Joi73]. Belgium [Joi73, PN94, TKD01, DH01, Den94, Dum69, GGD07, MDV03, VDB81, dB69]. Belize [CFRNT21]. Bellamya [OLM09]. belong [AYM09]. belowground [LAM06]. belt [NGSOAD20]. Beltschutsloot [Hog69]. benasi [DT23a].

Beneath [BB23]. benedeni [HGR11]. beneficial [TRP22]. benefit [SDRM16]. Benefits [vdHD17, SB97]. Bengal [CP10, FR09, NudWB^{+}25, NSB24, PPCH22, RAN24]. Benthic [AC16, CV77, DG06, MLB16, NP09, Ole97, RB08, SC09, TGE04, VS96a, WFH93, Wol74, ABT21, ACE10, BM98, BBF17, CTO01, CCS09, CMBI25, CAJ11, CS13, DDF93, DVK81, DLG20, DMD19, EÖ04, FLP08, FCG16, GAA24, HH08, HKO14, HK76, HDF88, IAHB18, JM84, JJG^{+}25, JvDS06, KF95, Kla80, KM17, KWO11, LH93, Leh04, LKSK15, LTH22, MGT17, MS02, MSB15, MBMV20, MvAV20, Mun94, NWL11, NČŠ12, NW24, Nie73, NRB21, NHL18, ODI08, Ola92, OMNC24, ÖE04, PV95, PS06, PU15, PGT20b, RFF01, RP22, SWB13, SOW01, SLM17, SMJ95, SYW09, SK04, Soi05a, Soi05b, TWC13, TPT22, TVD91, Van82b, VN22, VSW00, WWS21, WTC09, YMB93, ŽAL24, vdTS97]. Benthivorous [WPB08]. Benthos [SMJ03, De 80c, SGK07]. benzene [VV90]. Bergum [De 79b, Wan79]. Bergumermeer [Lee84, Wan79]. berilloni [SDRM16]. Bertness [Cad08]. Berula [RPS20]. besmetting [Swe71]. bestrijdingsmiddelen [lCdB70]. Bestuur [Ano72d, Ano68c, Ano69a]. beta [BDF20, CCS20, SLM17, TRP22]. betekenis [Spa71]. better [CKM19]. Between [BJJ12, AAPdG10, AC16, BV14, BCN12, BT23, Bri99, Bro84, CGC19, CZC14, CW04, CFRNT21, CDA03, Dan84, DS75, EMF14, EMN04, FBL13, GUT21, Gie71, GCE11, HH08, HWK11, HKT16, HÅB08, JB87, KW12, KMR20, KTH14, KKM09a, KPJP23, KJS15, KJP14, KHK17, LKM01, LLF99, MKA79, MBFD24, MOdSCP07, MMV09, Nau00, NMB09, NB22, OSM13, Pad13, PF07, PE78, PU15, PN94, PSD97, Rav01, RM00, Rin93, Rin97, Roi81a, Rus92, SAN99, SAGN07, SLS15, ST20, SHV96, SBH17, TXW^{+}24, TBBS17, TKH20, UET12, VNSFG16, VvOT09, Wol79, Wor90, WZL19, YZY09, ZPD06, ZTNW21, ZCZ14, dPNF91, dWvG88, vDB78]. Between-lake [BJJ12]. between-site [MOdSCP07]. Bibliography [Zet97a, Tol84]. bicarbonate [PSZ79]. bicolor [ACL13]. Bidasoa [Sol97]. Biddulphia [Baa85, JB83]. Biesbosch [HvB85, Osk72, Par68, Pee74]. Biesbosch-Spaarbekkens [Osk72]. bight [Mom73, WGK95]. bij [Ker70, Web71b, vdE70]. bilan [Pod74a, Pod74b, Pod74c]. binding [DE93, GMVCLGGP24]. Binnenpolder [Bel76a]. bio [HK82, JTGS24, KBN20]. bio-diverse [KBN20]. bio-indicator [JTGS24]. bio-organic [HK82]. Bioaccumulation [SGA06]. bioassay [MD05, PHB09]. bioassessment [ERA17, MSP13]. biocenoses [Waa80]. biochemical [ADS25, BGC00, DSN09, FAL07, JDJ^{+}25]. biocide [MvAV20]. biocoenotic [Mar01]. biodegradation [HZZ22, RKS14]. Biodeposition [FD94, MS08]. Biodiversity [MK05b, VVK96, Bar21, BSW^{+}24, CGL21, EAB08, FGA^{+}24, Flo01, GFF17, HCS02, JVCJ25, NWL11, PMB^{+}24, PGA^{+}22, PMM10, PKE22a, PKE22b, PVC96, RAO^{+}24, RFGLSB20, SSB21, WDA15]. biofilm [BLB^{+}23, LTL09, Rao10]. biofilms [ÁBZ08]. Biofoul [Hod98]. Biofoul-ing [Hod98]. biogas [TK89]. biogenic [ÓP04, Sta06, dPBCP22]. biogeochemical [DGN21]. biogeochemistry [BZV93]. biogeographic [SSS98]. biogeographical [Baa79]. Biogeography [Gul98, CGL21, DHEG21a, DHEG21b, NB22, PVC96, vD97]. Bioindication [SOW01, JB83]. bioindicators [CMBI25, Sur01, TBF09]. Biological [FC93, MFS03, Sch74, Sch83, WDA15, Bra01, CTO01, Coe75, DBG02b, DZG10, Dor87, DV74, HEZ03, Hig74, JVCJ25, JvDS06, KSBŁ95, Koo76, M.75, Mol80, OI89, PLM18, RGF21, RASA13, Rin73, STT01, TG88, TCH18, Van74, VDB81]. biological-ecological [VDB81]. biologically [Sta06]. biologie [Web71a]. Biologisch [De 77b]. Biology [BBP08, Cad05c, Gos98a, LMM05, vdV01, BSB97, CY95, CM05, Lee75, vdV99]. Bioluminescence [RMR21, VKK02, VPF21]. biomanipulated [JVG95]. Biomanipulation [GV89, Hos89, Ric86, TSV16, VKG95, Gul89, MRD89, MLG95, VDV86]. biomanipulations [HBO03]. Biomass [SMPK09, TK95, Ver79, BGM08, Bes87, BHPT20, CMB07, Cha87, Cra91, DTG09, EHM91, EAB08, GC08, GZN20, Gon79, GCE11, HKH^{+}23, HXR09, HH91, KD82, Koo76, Kou83, MG91a, MSD08, Nie82, NDW96, NHL18, PFMD00, SXY20, SKM05a, SKM05b, TK88a, VBD96, WF05a, WF05b, YHM05, vdTS97]. biomasses [HB03]. biomes [VNSFG16]. Biometry [BJS22]. Biomonitoring [JGUF23, KF95, MSK93, OB13, PDD07, Ros92a, SdAdC^{+}25, WD00]. bionomics [GSG24]. bioreactors [TBS18]. biosedimentary [RQ93]. Biosphere [War97]. Biosystematics [dHC76]. Biota [Hei88, CPM09, DJD92]. biotelemetry [MMC14]. Biotic [DH01, MGVC91, CHF09, HHC21, LLF99, Mag00, PB04, Rav01, SAR21]. biotopes [Df82, NvdVdlM01, WNS08]. bioturbation [CSA23, Dav93, HZC95]. biotypology [Cla87]. bird [GUT21, MŚSJ16]. birds [Mor99, Pra78]. Biscay [RFO94]. Bison [Lin72]. bisphenol [HZZ22]. bispinosus [BLC12]. bit [OSM13]. biting [MSP13]. bitterling [DBO11]. Bivalve [DP97, DBG02a, LB14, PCB10, PSD97, SPB97, ZSD94, dW71a]. bivalves [BBF17, HPM20, JJG^{+}25, PSP91]. Bivalvia [CB00c, DG84, Ken11, MSDDD22, PCB10]. Biwa [WiNDF22, YKU01]. Bizonbaai [Lin72]. bjoerkna [Geu84]. Black [GKM17, GCE11, KKS11, MML^{+}22, MML^{+}23, SAL16]. blackfish [BLC12]. bleached [SN03]. Bleiswijkse [MRD89]. blend [HH13]. Blicca [Geu84]. Blood [AIS02, Dor77, Dor79, KBO74]. Bloom [Los80, BCP10, CBC99, DCL19, DKI05, HLL14, MJM16, MFL09, SFV09, STH21, VPP16, VvGV12, ZQH16, Los80]. bloom-forming [BCP10, DCL19, VvGV12]. blooming [DG02]. blooms [BH16, BMJ16, HSP16, IFV16a, IBV16, IFV16b, MS11, SK16, TSV16, VIH16, ZGZ20, ZWY20, vGVV11]. Blue [Kap76, WB10, Bij75, HdC24, dCMMB15, Rei99, SZH21]. Blue-green [Kap76, Bij75, dCMMB15]. bluegill [EBM18, KW12]. Board [Ano68c, Ano69a, Ano72d]. Bode [CB00a, CB00b]. bodemhapper [Beu71b]. bodied [SOC12]. bodies [HVS87, Hig74, Ric86, SKV02]. Body [CBF19, Cru10, VV07, VALRdF20, ACE10, Jar08, MLTH15, Nat05, NSB24, Pen00, RM00, SBLdAM23, VRR17]. Boeken [Dor70a, dH72, hvRB69]. bog [ČN10, NČŠ12, RD77]. bogs [Hig71]. Bohemia [Dvo69]. boil [VFSH12]. Bojorquez [LGCS09]. Bolivia [AAV07]. Bolivian [JM08]. Bolshaya [GZN20]. boltoni [Sæt92]. Bolu [Kül05]. Boois [War97]. Book [Ano81a, Ano97a, Ano98b, Ano98c, Ano98d, Ano99b, Ano00a, Ano00b, Ano00c, Ano04a, Ano05a, Ano05b, Art99, Cad04, Cad99, Cad01, Cad05c, Cad05d, Cad05e, Cad05a, Cad05b, Cad08, Ess06, Gos98a, Gos98b, Gos99, Gul98, Gul99, Gul09a, Gul09b, Har73a, Hod98, Laa97, Moo98, Nie98, Sar05d, Sar05e, Sar05f, Sar05a, Sar05b, Sar05c, Shi98, Ver97, VDD77, War97, vD97, vdV99, vdV01]. Books [hvRB69]. border [Har73b]. bordering [Ber75]. Borderlakes [NvZvG16]. boreal [Das07, KOA00, RKS14, TGZ09, TRP22]. Boreoheptagyia [Mou92]. Bory [HL95]. bostoniensis [OLTV19]. both [LME10]. Bothnia [Sör82]. Bothnian [Haa74]. Botryococcus [KZV02]. Botshol [REH92, vIR92, SOD92, dR92]. bottlenose [PBDJ20]. bottles [Ber84]. Bottom [Cha71, Hig81, MBMV20, Alb04, BBM21, Cla87, DdNM22, DGBL15, FBS19, HKO14, KEH09, KIG06, KMK09, MQC93, iNTT05a, iNTT05b, PJH06, PT14, Ros92b, TET23, ŻKŻ10]. Bottom-up [MBMV20, BBM21, DdNM22, DGBL15, FBS19, PT14]. bottoms [HHV93, HWG77]. botulism [Haa73]. boundary [KEH09, PV95, RAA98]. bouquet [FTH22]. Bouwhuis [Gee76]. box [JBB07]. boyeri [vdVP76]. Bozcaada [Bal09]. Brabant [Van79a]. Brachionus [GFSN04, LSY20, LYM21, LRSN22, SAN99, SAGN07, YN08b, YZ08, YYR20, ZXZ09]. Brachyura [LMM05]. Brackish [Bak72, Har74, BD74, Bak77, Bak78, Bak79, DWA11, De 74, DD84, DDV84, Dor74, KS94, KKR02, KKD02, KRZ03, Kre82, KBO74, LMV09, Lov74, MR81, MGC19, Mur69, Mur74b, Muu74, Nie79b, Nor98, Pol75, Pos74, RM81, SSR13, SC09, SR88, The02]. Brackish-water [Har74, KKR02, Nor98]. Brady [Hol70, Hol70]. brak [Mur69]. Brakwatermeer [Bak72]. brama [Caz82, Geu84, Raa89]. Brancelj [Gos98a]. Branchinella [YP17]. Branchiopoda [VSB10]. branchiopods [MMBP21]. Branchipodopsis [VSB10]. branickii [NH92]. Brantas [STT01]. brasiliensis [dCKL23]. Brazil [ATB07, BSMVP24, BdSCH09, CSV06, FGL19, FAB18, FBNF08, GFZ15, GDB19, MSDDD22, dSLM94, NNB^{+}24, NCDS^{+}25, NP09, PMVdC^{+}23, PNC22, RFGLSB20, SdAdC^{+}25, dSPdS10, dSSR^{+}23]. Brazilian [AL24, BA07, DMD19, FdLY^{+}24, FNS^{+}24a, FNS^{+}24b, LSdS^{+}24, MCV08, SDD^{+}23, dCKL23, dPBCP22]. breadth [SAACR09]. breakdown [BHV82, CYW24, DCC20, PB84]. Bream [SGA06, Caz82, Geu84, Lam89, Raa89]. Bréb. [Ste76]. Brebisson [Baa85]. Breeding [Bre80, LMM05, GRDPP09, KLG17]. Brehm [MJEC05]. brengen [Ros69b]. brevis [RVB08]. bridgesii [AR05]. brief [vdB70]. Brielle [Kap80]. Brielse [Kap76]. Brightwell [Baa88a, Baa88b]. Britain [LH93]. British [CF96, Cor78a, Mit74, Mit78]. broad [Cla82]. Brokopondo [Van78b]. Brong. [SBS12]. brook [Sch92]. brooks [Dor76a]. brown [AK09, GS04, HKH^{+}23, HFS08, dCMMB15, SH16]. Bruce [Gul99]. Brunswick [JLG06]. Brussel [Pol73]. Brussels [Pol73, TKD01]. Bubbler [LMM05]. Budaun [KP24]. budget [JLS09, MNK04]. budgets [Bro81, Van77b]. Budzyńskie [Nat05]. Buenos [SSI21]. Bufflehead [HF08]. Bufo [ZWJ19]. Bugach [MKG03]. bugensis [ISJ10, OTS05, OMM05a, OMM05b]. bugs [Ara01, NCDS^{+}25]. building [ÓP04]. Buleji [GSG24]. Bulinus [KC22]. Bulk [LKE94]. bull [RMH21, WF05a, WF05b]. bullhead [BKG16, GA10, Pad13]. Bullseye [CBdO^{+}24]. buoyant [MOCPG09]. burbot [BR05]. Burchell [Nwa95]. burden [PFLM15]. burial [CDL14]. burrowing [PdE^{+}24, RL05]. burrows [MKL10]. butt [WPB21]. Bysmatrum [RP22]. byssus [RVK96]. Bythotrephes [KG94, NAT08, Vis88].

C [Art99, Gos98b, Gul99, Shi98, War97, APS81, BBM09, BPD15, DB85, Fra00, GCL99, MB81, TKR10]. Côte [KKO08, AGB94]. CA [LME10]. Cab [ECGC25]. Cab-ilan [ECGC25]. caddis [Hig80, MM96]. caddisflies [Hof99, MO92]. caddisfly [TNtK00]. Cadmium [LWY^{+}24, DDZ^{+}24, GFSN07, HLSN24, PMSS18]. Caenogastropoda [GOB24]. Caffrey [Ver97]. cages [KM17]. Caicara [dSPdS10]. Calabar [NE85]. calamity [VSA97]. calanoid [BV89, DM04, NBM98, NSB24, PM00]. Calanoida [FG91, NCT91, NSB24, TSZ10, Tol02]. calanoids [LRSN22, MJJ21]. calcein [TFT10]. calcifying [HEF13]. calcium [BA14, GD09, JCL09]. Calculating [CC94]. Calculation [RCL94, Ald78, Ald79, Bui95]. Calculations [DB85, TV85]. caldera [MT09]. California [Bau00, BC23, PEREBM^{+}23]. call [NRS19]. Callinectes [HdC24]. Calmano [Gos98b]. Calopteryx [IS06]. calyciflorus [LSY20, LYM21, SAN99, YN08b, YYR20, ZXZ09]. cameras [CJS^{+}23, GCN^{+}23]. Cameroon [TAN14]. camp [Bel76a]. Campotosto [SB88]. Can [BKG16, MRR10, MCSV21, RTW17, SCG09, TPS09, VPS21, XWJ22, DMM97, HLL^{+}25, LLX11, Par69, PD10, Ros69b, WHW14]. Canada [GB82, EPL^{+}23, HP06, HF08, IJ06, JLG06, Pip87, VH08]. canadensis [Bes76]. Canadian [BTB09, KCG05, PFDM99, PFMD00]. Canal [KK20, CM95b, GE80b]. canaliculata [SMM23, SM12]. canals [Bel79, RU87, ZPW15]. Cananéia [dSLM94]. Canary [RMS11]. Cancun [LGCS09]. Candia [RG07]. cannibalism [Cru10, IS06, IDP19, KKM09a]. cannot [DMM97]. canoe [SPA12]. canoe-mediated [SPA12]. canopy [CvNB02, LB10, WKT05]. can’t [OR06]. Cantabria [BD09]. canyon [VEK20]. canyon-shaped [VEK20]. capability [QJD02]. capacity [BDR97, DP97, FDB97, HFS08, PBM21, SPB97, SGR86]. Cape [AANO25, Bar21, HHN03]. captivity [SSV21]. captured [TRHHAR10]. Carassius [RRK15]. carbo [VBA22]. Carbohydrate [BWM86]. carbohydrates [Bes79]. Carbon [KT09, BVM15, BT01, Cad80, CH91b, DDV91, DGB17, FVW03, GC05a, GC05b, HFS08, HWS22, HH80, KTJ10, KA97, KRJ09, LTL09, LSR08, LZN14, MV83, Nie79a, Nie79b, PPL16, PFDM99, PFT12, RKS14, RHS18, RBN20, Roe96, SKW21, Sch78, THS17a, THS17b, Van77b, Bes79]. carbon-to-nutrient [PPL16]. Carcharhinus [RMH21]. Carcinus [FTH22, RMH22]. Caretta [NPH03]. cari [ŽG09]. Caribbean [ARG20, FBM22, HSCR10]. Carijoa [dPBCP22]. carinata [GX11]. Carlin [HHP12]. Carlingford [BDR97, FDB97]. carnivores [CMH14]. carnivorous [CBE20]. carnivory [GANA00]. Carolina [PDP02, AP12, BB23]. carp [DD99, JM84, KG03, Raa89, RRK15, SD09, Van75b, ZYL22]. carpio [Raa89, SD09]. carried [Bel76a]. carrying [BDR97, DP97, SPB97]. carterae [RP22]. Cascade [SR92]. cascades [CH12]. Cascading [FGB20, SSB21]. case [AKF^{+}23, AAA^{+}24, Ald78, ABDPPP^{+}24, AAV07, BVM15, BPD15, DDPT24, EBM18, FAC16, FA18, GR97, GAA24, HBO03, HLL^{+}25, JGUF23, Lee82, MHYL23, Nie78, NRB21, vOCRM19, PMM10, RKLB22, RKLB23, SWW02, SEB11, SKV02, dJG07]. case-studies [dJG07]. Caspian [BKG16, BHG18a, BHG18b, MAS14, OMM05a, OMM05b, PDvdV06a, PDvdV06b, VPS13]. CASSARINA [ABF01, BPA01, BBR01, FAK01, Flo01, FDP01, PBR01, PJF01, RFP01, REP01, RFF01]. Castellammare [MS08]. Castelo [CB00a, CB00b]. Castelo-do-Bode [CB00a, CB00b]. Castillo [Art99]. Catalogue [Ash92]. catastrophic [RAD21]. catch [Gud04]. catchment [LWO01, PRW08, SLL09]. catchments [JGUF23, MB06]. catesbeiana [STV06]. Catfish [PGM^{+}24, BFS21, Pad13, SBLdAM23, WBS22, WBS24]. cations [LDK20]. Catla [SMC08]. Catta [SDRM16]. Caulerpa [CCS09, DNG22, KS79, KMS82a, KMS82b, PVC96]. causaal [Gie71]. causal [Gie71]. cause [CDL13, FN94, FASdPN^{+}25, Ros69b, SASA^{+}24, Swe71]. caused [ABDPPP^{+}24, Pee74, WCD17, ZD21]. Causes [Lam89, Nie94, PPY^{+}22, BKW94, BJS12, HGM22]. caution [NRS19]. cave [FTP08, GFF17]. cavernosa [STG18]. Cd [RAP22, WA84]. cell [FO98, NHiN15, Ste94]. Cellana [FCS15]. cells [CDL13]. cellular [TDL23]. cellulose [BLB^{+}23]. centennial [Gud04]. Central [BB87, BST89, Bra21, DN23, HGM22, KND^{+}22, Mas85, SB88, TAN14, TBF09, AKL15, GBZ21, Har04, Hig09, HC09, PFLO14, PFLM15, RJD07, SNG21, TWC13, VD94, DCC20, GTPH08, Grz92, Pip87, PSSKŽ24a, PSSKŽ24b, SKM05a, SKM05b, TDHH23]. Central-Baltic [GTPH08]. centrales [Ste94]. century [Flo01]. Cerastoderma [PSP91]. Ceratophyllum [Bes76, Bes79, DCG19, MB81]. Cercopagidae [NAT08]. Ceriantharia [CBdO^{+}24, SPB^{+}24]. Ceriodaphnia [AMSN07, GTI99]. Cerrado [PMVdC^{+}23, VNSFG16, dRFC21]. Cerro [MT09]. certain [TG04]. cervula [HP06]. Cesium [RCSF93]. cestode [WCT00]. cf [PAAJM17, TPT22, KP97]. cf. [BRB97, BSB97, PGV95, Sch97b, SVB97, Zet97b]. CH [BVM15]. Ch. [SHK92]. chaetoceras [Coe97]. Chaetoceros [Bak94]. Chaetognatha [FNW94]. chain [Beu71a, HMB88, SGK07, WXL16]. chalk [Rus92]. Challenges [MTJ10]. chamber [TFT10, TGE04]. Champsocephalus [WNZ24]. Change [Flo01, BGRM^{+}24, CPA15, DTH20, FDP01, Gib86, HCS02, HHEM19, HPT95, LMA10, MHL05, MRR10, PJF01, RPS20, SMS^{+}24, TPS09, Vaa79, VFSH12, Wey09]. Changes [ASA06, Bes87, CPM09, CC99, Doo82, EK98, FdLY^{+}24, HB85, IUK22, LSO10, LSY20, MSD08, Pee74, PHB09, PFT12, PP95, PHD13, QLZ20, SMJ95, VBD96, YSY^{+}22, ZJN^{+}23, vDB78, ARP14, AL09, AV92, ABF21, BPA01, CSK10, CM05, Coe78, CV96, dSDLdA22, DDF93, DBV96, EÖ04, FSN21, GH06, GOB24, Har04, HNiN03, JM84, JBS05, KK12a, KC82, KBO74, LAM06, LBKV18, MMM11, Mar01, MJJ21, Nie78, Nie82, NDW96, NHL18, PBR01, QLW^{+}23, Rei92, RP22, SH16, TŠB17, VvdMP12, Ver76a, Wel74, ZGZ20, ZWY^{+}24]. changing [MVI16, Nie94, QSA09, RMH22]. channel [BWH02, MG06, SMC21, DDF93]. Chaoboridae [Par69]. Chaoborus [BNPC07, Par69, PAY03]. Chara [BA97]. Characean [SOD92]. characid [PdCF^{+}25]. characterisation [WGK95]. characteristic [Roi88a]. Characteristics [SAR21, AAA^{+}24, Bel76b, BP94, BTP24, BCP10, BAG94, BRP15, Hal76, HWS22, JvDS06, KJF00, KPV16, LKSK15, MB81, MBFD24, MD92, PCGW20, SOW01, SHK92, SH11, VKK02, WBR17, YTD02, YZ08, ZGZ20]. Characterization [Bla82, OMNC24, AVD07, CGO19, LKE94, RFP01, SLH87, VH89, VSM19b, ŽRP15]. Charophyta [Har73a]. charophyte [HLK22]. charr [AK09, Bra99, Gud04, KJS15, SSK14]. chasse [Joi73, Pod74a, Pod74b, Pod74c]. check [SS98]. Checklist [WZ92, NE85, VMS94]. Chelicorophium [BHG18a, BHG18b]. Chemical [ABV72, BH22, GEÓ04b, Gys72b, JB83, KGS02, Żur06a, ÅDB00, AOS^{+}24, BD09, Bel76b, BG15, Bra01, Bro98, CBC99, CM95a, DE93, De 80b, DZG10, Gro22, Hog69, Kir22, KPJP23, KSBŁ95, Lac91, LKE94, LHV87, MIMRRZ08, MLZ10, MJ93, REH92, STTS18a, STTS18b, SOW01, SHP21, TB22, VvGV12, WV02, Wey09, WWŻ06, Zad03, ZL07, Moe71]. chemically [KM83]. chemicals [Her83, RHA00, ZG02]. chemie [Dum69]. Chemische [ABV72, Gys72b, Moe71, Hog69]. Chemistry [Gol71, KÁ04, AIS02, Dum69, Nwa87, QPP95, SSB21, Van96, vDSB80]. chemo [VGP^{+}23]. chemo-ecology [VGP^{+}23]. chemoautotrophic [GFF17]. chemocline [KBR10]. Chemoecological [DNG22]. cheni [KKS11]. Chennai [NSB24]. Chernobyl [RCSF93]. Chesapeake [YHM05]. Cheshire [BM03]. Chile [PFLO14, PFLM15]. Chilean [BSW^{+}24]. Chilomonas [SFA05]. China [LSO10, CZC14, CCK10, CDL13, CZZ^{+}25, DZT13, HWS22, MHYL23, PLM^{+}22, SLC10, STH21, TWC13, WZ92, WXW11, WZL18, WLB22, WWG^{+}25, WTC09, XCL10, YL08, YBZ20, ZZL16, ZPW13, ZPW15]. chinensis [OLM09]. Chinese [DGBL15, HHC21, MSF21, OLM09, ŠDT^{+}24a, ŠDT^{+}24b, WXW11, WXL16]. Chinook [KAN20, DAO24]. Chionoecetes [YSY^{+}22]. Chironomid [KKV92b, BP92, Dvô70, FKB92, GEÓ04a, GL92, KK12a, KKV92a, KH92, LR92, MSW10, Ola92, ÓP04, Ros92b, Rus92, TD87a, Tok92, WB10, HD93]. Chironomidae [Mou92, Ash92, BH92, BW92, CDW92, Del92, Dvô70, FLA09, GWF92, Grz92, IÓG04, Jac92, KSW92, KML19, Kre79, Kuk92, Lar92, LM92, MCC21, Mic92, MF92, MSP13, NH92, PGV95, PR92, RFF01, Ros92a, Sæt92, SB88, SR92, SKK92, SHK92, SS92, SKM05a, SKM05b, WZ92, WH92, WW84, WAW92]. Chironomids [BB87, Fer92a, BST89, BBD17, CZC14, MO92, RP92, RHO99, Sch92]. Chironomus [BW92, GPA98, KSW92, KML19, MKL10, PGV95, SHK92]. Chiroptera [RHO99]. chitala [PPK^{+}23]. Chlidonias [GKM17]. Chlorella [AND25, DCG19, GFSN04, HZZ22]. chloride [LKSK15]. chlorin [Kow94]. chlorinity [DDV84]. Chlorococcales [IM06, Mur69]. chloroform [VV90, VV93a]. Chlorophyceae [MSK93, Nie74, Ste76, WBB00, dHC76]. Chlorophyll [AAPdG10, CSM08, PPC08, BS95, CH91b, GEG91, HB85, MAS14, NO85, PFMD00, Rat17, Roi81a, SGM99, ZZL16]. Chlorophyll-a [AAPdG10, MAS14, Rat17, Roi81a, ZZL16, BS95, NO85]. chlorophylland [Bes79]. Chlorophyta [AND25, BI09, IM06, KML02, MR11, PVC96, RJD07, Sch95]. chlorpyrifos [TJB98]. choice [GBA20, Pri03]. choices [HWK11]. choreus [Lar92]. chorus [KBN20]. chromium [VRH^{+}24]. chromophoric [GM16]. chromosomes [FKB92]. chronic [GFSN04, JDJ^{+}25]. chroöcoccal [RCV91]. Chrysaora [BJS22]. Chrysophyceae [Roi81b, Roi88a]. chydorid [NSK08]. Chydoridae [NSK08]. Chytridiomycota [dNdSBJdCeS21]. cichlid [TBA03]. Cichliformes [JDJ^{+}25]. Cigarette [WPB21]. ciliate [CLL10, GAAT^{+}24, LDL11, SFA05]. ciliates [KK20, KBR10]. Cinachyrella [STG18]. circadian [dSSR^{+}23]. Circeo [BST89]. circulation [KSF95, KHM04]. Cirripedia [CY95]. citarum [CMBI25]. citizen [MTC^{+}24, MTdS^{+}24, PFSPdS24]. Cladocera [BM03, BV84, DHEG21a, DHEG21b, EÖ04, GFSN07, Gie71, Irv86, Ker91, LPT21, LXH21, NMS21, NSK08, NZS16, NRVV76, ÖE04, RFF01, SAB05, SSS98, Vij91, Gos98a]. Cladoceran [JTGS24, AVdNM19, DHPN^{+}24, KG94, LAZ21, MR17, NR97, Sir17, Zad03]. cladocerans [AMSN07, dFBdGGMLT21, BA07, DSD87, FA18, dSFFA03, FAS05a, FAS05b, FAL07, MNS05a, MNS05b, Nan00, NEGS07, SNGV06, WFFM10]. Cladopelma [PR92]. Cladophoraceae [Pol76]. Cladophorales [Nie74, dHC76]. Cladophorophyceae [BI09]. clam [PHM21]. clamitans [CB13, STV06]. Clarias [Nwa95]. Clariidae [Nwa95]. clarki [VK12]. clarkii [AFR14, MLG20, RHS18]. class [HJ23, LMV09]. Classical [BMS20]. Classification [De 74, Aag92, CNN^{+}24, ES09, Har74, LH93, RRS^{+}22, RRS^{+}23, VHC92]. classifications [AR09]. Classifying [PME08]. clay [Den94, HEPH09, TW03]. clay-turbid [HEPH09]. clean [YRR09]. clear [LJL05, MOdSCP07, MOPCP08, NvZvG16, Soi05a, Soi05b]. Clearance [Dev09, HPM20]. clearwater [MDV03]. Cleve [Baa88b]. Clibanarius [TK19b, TK19c]. Climate [BJ04, RP08, VFSH12, BGRM^{+}24, GSB13, HHEM19, Kae20, KOA00, MHL05, RPS20, SMS^{+}24, ZXH15]. climatic [BTP24, BJ04, LPT21, TŠB17]. clonal [LW11, qLlZhY16, MLL21, RW98]. clones [BDP21a, RR98]. cloning [FO98]. close [TBBS17]. closed [BV78, CSV19, DB85, Nie78]. closely [KML19, VvdMP12, YN08a]. closest [GCT^{+}23]. closing [SLG09a]. closure [Par78, Ver80a]. clousure [Doo82]. club [GL88]. clustering [AVD07]. Cnidaria [MTdS^{+}24, NNB^{+}24, SPB^{+}24, SC00]. Co [WMV09, HHY21, KKM09b, LZZ^{+}25, PMA22, SKH19, VHN00, YYS^{+}23, CBP22, FM10, LAC99, LWX19, RPS20, WYW^{+}24]. co-culture [YYS^{+}23]. co-existence [KKM09b]. co-existing [VHN00]. co-occurrence [HHY21]. co-occurring [PMA22, SKH19]. co-polluted [LZZ^{+}25]. coal [DSG^{+}24]. coarse [TRP22]. Coast [KBN20, ABA^{+}24, BJS22, BTP24, CY95, DNG22, FBNF08, LWM89, MAA22, NSB24, PFLM15, RNJ97, SDD^{+}23, SRI20, TRHHAR10, WZL18, dW71b, BASJ94, GBJM22, zEGB13, CP10]. Coastal [Bou07, Cad05d, Nie98, BV89, Bal09, BBM09, CBC99, CPP22, CL11, CP10, Clo99, CVG03, CKK99, CCS20, CVM21, CSV19, DP97, De 74, DPL03, Den94, DGG15, DGB17, DGGW20, DBdF02, EFD99, FBV02, FO98, GB82, GKK95, HKO14, HEZ03, HMB88, Kat92, KSBR94, LLF99, LGCS09, MHYL23, MAS14, MB14, MBM^{+}22, MM95, MTSJA19, iNTT05a, iNTT05b, NudWB^{+}25, NP09, PP95, RVJ98, RAO^{+}24, SB78, Sch97b, SOU14, TC15, The02, WSvA98, Zet97b, KND^{+}22, Cad01, Nie98]. coasts [RVJ06]. Cobelas [Art99]. Cochin [Wel74]. cod [NZ75]. coded [VMS94]. coefficient [BP94, GEK97]. coefficients [AZD02]. Coenagrionidae [HP06]. Coexistence [Deg10, dSSR^{+}23, GA10]. Coexisting [dSGD17, BDG06]. cohabiting [WHS13]. Cohesive [FCW94, CN94, FSC06, ÓP04, TFP06, TDP06, VML93]. cohort [DF11]. cohorts [IÓG04]. cold [DCC20]. coldwater [GMW09]. Coleoptera [HH08, CAPGA08, DB02, Gie83, MB14]. coleoptrata [DS75]. collapse [CMM05, HGM22, TDHH23]. collected [NBT21, WKH82, WA84]. collecting [Kor98, Kor14]. collection [KZV02, Kle84, PGT20a]. collections [Rus92]. Colombia [JDF23, RPSSDS23]. Colombian [TBSZ22]. colonial [dSFFA03, KHG20, YZY09]. colonies [LRG21, LZN14, TKR10, WZL19, YSYZ23, dSCT23]. Colonisation [HF97, AANO25, CF96, TGD16]. Colonization [WAW92, BCP10, DD84, DLG20, GAA24, HEB00, KMR20, KTH14, KM17, LMA10, MQC93, PT14, RGF21, UET12]. colonization-related [KTH14]. colonizing [ERA17]. colony [BHB01, MMV05, WZL19]. colony-attached [WZL19]. color [NNN22]. Colorado [VSW00]. colorimetric [Sch78]. colour [WSvA98]. colourless [GEG91]. Columbia [HBR12]. column [BZV93, BT01, KFS04, KPS09, KIJ11, KIG06, Rin81a, ZV93]. combat [De 80b, Los80, SK16]. Combating [FAC16, JCF17]. combinations [Dor74]. Combined [BJS12, SAGN07, WMTR07, BVM15, FCW94, KT15, LSY20, MRBQ^{+}25, PMSS18, QLW^{+}23, RAP22]. combining [NMP12]. Combretaceae [WZL18]. come [BRJ97]. commensal [Pen00]. Comments [De 80a, TKM03, Beu74, WvRvdV03b]. commercial [ABF21, CC99]. common [DTH20, HPT95, SD09, TK89, VSM19a, ZNP13]. commonly [Kae20]. communication [Kir22]. communis [AND25]. communites [Pip87]. communities [Aag92, AVD07, AT24, AFC20, BGRM^{+}24, BB00, BLB^{+}23, BKD^{+}24, BD09, BM98, BCN12, BA97, BDF20, Bra21, CPA15, CAJ11, CDW92, Cra91, DDF93, DD99, Df82, EMN04, EK98, EPL^{+}23, FdLY^{+}24, FAK01, FLP08, FSN21, FNS^{+}24a, FNS^{+}24b, FWR^{+}23a, FWR^{+}23b, FNW94, GGD07, GGT14, GZN20, Gru11, HGM22, IUK22, JCS01, JVCJ25, KK12a, KJF00, KIJ11, KPJP23, KRJ09, KCG05, KMK09, Kuk92, LPT22, LH93, LTL09, LLF99, LPB07, LBY17, MGT17, MTA15, MDH93, MLG95, MBMV20, MvAV20, MAD10, MB06, MNT16, NSK08, NZ98, NMdAM21, PV95, PRW08, PD10, PVV07, PCGW20, RNJ97, REP01, RHL02, SWF17, SEP09, SLM17, Soi05a, Soi05b, SA95, STH21, SZH20, TŠB17, TRP22, The02, TLH18, TKH20, TK19b, TK19c, VS96a, WRH98]. communities [WvLA14, WRS05, WMTR07, ZWY20, dSPdS10]. Community [CM95b, FGM20, GPGSMM^{+}23a, GCT^{+}23, HHA92, MKS22, UKS05, WGL21, ARP14, AZD02, AP12, AR97, ASLT15, AOS^{+}24, AC16, BR05, Bol07, BB23, BV96, CGC19, dSCdMM09, CCK10, CZZ^{+}25, dMCFS^{+}23, CAFA04, Dav93, DÖD21, FR09, FGF13, FM99, FBL13, GFF17, GMW09, GFZ15, Gra12, GCE11, Gul76b, Gul89, HLK22, HRP91, HAS22, HHY21, Irv86, JAM15, JJG^{+}25, Kim99, Kir22, LSO10, LLL^{+}23, LRMdF19, MGL22, MBM15, MSB15, dSLM94, MSP13, MSF21, MNK04, NP09, NČŠ12, NAM19, NHL18, OLJB20, PDE08, PEREBM^{+}23, PS06, PFT12, PDP02, PIM95, PHD13, QPP95, Rav01, RRL09, RU87, RAA98, SGRMP20, SB08, SSI21, SN03, SL07, SCS12, SB88, SOD92, SMJ95, SK04, SKM05a, SKM05b, TFH^{+}24, TGW06, TK19a, Tok92, VDV86, VL18, VSW00, WWS21, WBR17, WZM^{+}23, WT14]. community [WZL19, YYS^{+}23, YHM05, ZWY^{+}22, ZPW13, dSSB24, dRFC21, GPGSMM^{+}23b]. community-level [SGRMP20]. Como [Rav01]. Comoé [KKO08, KKO08]. comoensis [KKO08]. Comparative [ÁBZ08, Bak77, Dor74, DMD95, Lee84, Rao10, Rat17, SK04, TKD01, VM25, vdV76, CSA23, Gul09b, PJH06, GB69]. compared [BWS19, CH91b, OLJB20, SHK92]. Comparing [Fra00, HF08, YB08, dBPG00, CAPGA08]. Comparison [BD74, DPL03, HD79a, VDB81, VK12, WFFM10, WZL19, AMTSJ15, BDR97, Cra91, DWB10, DCL19, FED95, GB82, GDT08, JvDS06, KNH07, LDL11, MKG03, MF92, MSD08, PU15, PR84, PN94, RGF21, Rat14, Rav01, RM00, RDG86, Roi81a, UET12, YRR09, ZCZ14]. Comparisons [DBG13, HSY18, NJ00]. compartment [DB85, TZ05]. compartments [Kal86]. Compensatory [LLG10]. compete [PMA22]. Competition [HÅB08, SAN99, SLS15, SBH17, CLRdSR23, Den74, FM99, FLA09, KK17, MLG20, NEGS07, NS21, Pin75, SAGN07, SZY19, Som83, WCT00, Wor90, dPNF91]. Competitive [CW04, MCB22, DH22, GA10, KML19, MF07, NC20, RG07, RTL17, TPT22]. competitor [ZCA20]. compilation [MDP08]. complex [ARG20, BDP21a, Bol04, DW14, HK82, SB97, TK88a]. complexes [TW03]. complexity [DTA22, dSDMD20, FGF13, HWK11, KKM09b, MSB15, Rin73]. component [DGH^{+}24]. components [DZG10, ED95, OLTV19, TRP22, TDL23]. Composition [AP12, GWF92, HHY21, STH21, WLB22, AC16, Bel87, BR07, Bes81, Bes87, BJJ12, BB23, BGC00, CH91b, CCK10, Coe78, CS13, DHPN^{+}24, DA99, DG84, DSL12, DÖD21, ECGC25, EDLC24, GMW09, GBZ21, GEÓ04b, GSM07, GDB19, GZN20, GDT08, GR91, GC23, HRP91, IUK22, IAHB18, JBS05, KP24, KBS05, LBY17, MGL22, MKG03, MB14, MDH93, MSF21, Nor98, OSO09, PFT12, PD10, PR84, PKE22a, PKE22b, RMW91, RAA98, SL07, SKM05a, SKM05b, ŠNZ21, Str88, SZH20, SGGO06, TŠB17, WZL19]. compositions [KZV02, NHiN15, ZKN09a]. compounds [DE93, DF90, GEG91, GM16, MJM16, NN05a, NN05b, STTS18a, STTS18b, Sep79, dH71a]. comprehensive [SL24]. comprising [DSN09, KBN20]. compromise [CvNB02]. computer [BGK02]. Concentration [Ver76a, CLH06, Hos80, KK07, Ker78a, NO85, PGV95, PFMD00, RPS20, SAN99, ZJN^{+}23]. concentrations [BHL09, CMG95, CBN94, DV92, GV99, HB85, HFS08, KMS82a, LAZ21, PDP02, PMZMJ16, SHK92, TK95, VVW95, WYZ13, Gul98]. concept [LCS01, Par80]. Concepts [DVK78]. concern [Cra85, FASdPN^{+}25, SASA^{+}24]. concinnus [Baa88a, WP91]. Concluding [vL86]. concomitant [HMB88]. concordance [FGM20]. condition [AL09, CBF19, LCPM15, Leh04, LTH22, SBLdAM23, VRR17]. conditioned [PA98]. conditioning [CKM19]. Conditions [DG06, AP00, AKF^{+}23, BSA14, BZD82, BTB09, BHL09, BHP05, CBC99, CM95a, CGD19, CSM08, CC94, DTH20, DE93, Dvo69, FSN21, FLA09, GBM10, GLGU22, HD93, HD21, KBS05, LCPM15, LZD16, MS07, MV14, MAD10, MD92, Nie74, NMP12, OLJB20, OMV14, vOCRM19, PLC22, Rei92, REH92, RZG09, RFW17, SVW95, SB14, ŠNZ21, TŠB17, TMM21, Wey09, WDP15, YSY^{+}22, YO02, ZG02]. conductivity [GANA00, LDK20]. configuration [GSB13]. confirm [WVM18]. conflict [MŚSJ16]. confluences [HZC95, NWL11]. congener [PPH16]. congeneric [DTA22]. Congress [Ano77a, Ano97e]. Congressen [Ano70a, Ano71a, Ano72a]. Congresses [Ano72a, Ano78a, Ano80, Ano81b, Ano82, Ano89a, SC83, Wet81]. congruence [TKH20]. Conidae [GSG24]. conifer [VK12]. conjugated [MKA79]. connectivity [BKC11, RMH21, SPA12]. consequence [RBAB24, TDHH23, TLH18]. Consequences [Blo70, BVV80, CHF09, KM83, Lam89, NZ75, VGV82, XYH11]. conservation [AL24, CvNB02, LH93, Mit78, NCDS^{+}25, PPCH22, RASA13, VBA22]. consideration [SB78]. considerations [Ald78, KPV16]. conspecific [Zad03]. Constance [WMTR07]. constant [SH11, SBH17, WPvB17]. Constantine [KND^{+}22]. constituents [KBO74]. constraints [BTB09, SJ00a]. constricta [LLL^{+}23]. constructed [vGVV11]. construction [SMJ95, Ver76b]. consume [OLM09]. consumer [PB04, SAB10, SB04, YP17]. consumers [CCS09, YL11]. consumption [FA18, HBL13, LA19, dSLM94, RB83, SRG18, WVM18]. consumptive [PMA22, SGD22]. containing [LLL^{+}23]. contaminant [PJF01]. contaminants [Pus94]. contaminated [GPA98]. contamination [HPT95, LCF91, MvAV20, MFP01]. content [Bes76, DPT04, HF08, KFS04, KTJ10, MSD08, NC20, SPB^{+}24, SSR13, SRG18, TBA03, TRHHAR10, UOJGRL21, VALRdF20, WvRvdV03a, WYW^{+}24, dCdCP19]. Contents [WA84, Ano98i, Ano03c, Bes79, MMM11, NAT08, Ano68a]. context [GAAT^{+}24]. Continental [Nie98, GPV19, SDD^{+}23]. contingency [Gra12]. continuous [Ald78, Boe86, Deg10, GBM75, Som83]. Contrasting [HPS11, MBM15, MK05a, Pad13, VV05, BGL13, FBL13, KJS15, MAD10, PAP24, ŠNZ21, TK19b, TK19c]. Contrasts [Mos97]. contribute [VPS21]. contributing [CGL21]. Contribution [BB12, BDF20, Del92, JB83, Lam97, MTdS^{+}24, PMM10, RAO^{+}24, UOJGRL21, zEGB13]. contributions [DBV96, FDG97]. Control [Har82a, Hod98, HM86, AVdNM19, Bel87, BLG99, HSP16, IFV16a, IBV16, IFV16b, Lah98, LWdSD16, MVI16, MBMV20, Osk72, SON11, SK16, Van75b, Ver76b, VIH16, WT14]. Controlling [BMJ16, CCS99, KP97, SB08, YP17]. controls [CKK99, CSV19, DGBL15, FED95, FTH22, GC23, HAS22, OMC22]. Conus [GSG24]. convection [KPS09]. convective [VS09]. convenient [SVB98]. conventional [KT15]. convergent [RTW17]. conversion [HS85, HS86]. Convolutional [RRS^{+}22, RRS^{+}23]. cooling [Kel73, Koo73, Ver73, Wil79]. copepod [BV89, BNPC07, BGC00, CGD19, DM04, GIK21, MBM15, NS07, NBM98, OCS20, PM00, RK02, RM00, WCT00]. Copepoda [Bra99, FG91, LRSN22, NCT91, NSB24, NudWB^{+}25, TSZ10, Tol02]. Copepods [MJJ21, BTV88, Bre80, CDD20, HJ23, KRATA20, NSB24, SZH21]. copepoo [GEG91]. Copidodiaptomus [CB00b]. copper [Ade71, MEB93, RMW91]. Coral [TDHH23, GLGU22, JCS01, NvdVdlM01, NYT09a, NYT09b, RRS^{+}22, RRS^{+}23, TKR10, Vie20, dSCT23]. coral-reef [NYT09a]. corals [WLN12]. Corbicula [HPM20, PHM21]. cordgrass [CF09]. Cordillera [DN23]. Coregonus [HHP12]. cores [RFF01]. corixids [DS75]. cormorant [VBA22]. cormorants [KLG17]. Cornelis [VBK96]. corniculatum [QLZ20, QZL^{+}24]. Cornwall [RMW91]. coronatus [GSG24]. Corophiidae [BHG18a, BHG18b]. Corophium [Kle84]. Correction [BHG18a, BZG^{+}22, DHEG21a, FNS^{+}24a, FWR^{+}23a, FASdPN^{+}25, GPGSMM^{+}23b, KGM22, PSSKŽ24a, PKE22a, STTS18a, ŠDT^{+}24a, TK19b, WBS24]. Corrections [Ash92]. correlate [CCS20]. Correlation [MJ93, MKA79, CV95, LLL^{+}23]. Correlations [Rus92, GSM07]. correspond [BD09]. coscinodisci [WP91]. Coscinodiscus [Baa88a, WP91]. Cosmarium [Coe97, SH11]. cost [BRP15, RKLB22, RKLB23]. cost-effective [BRP15]. Costa [CDW92, WH92, Sol97]. costs [SB97]. Coulter [Ker70]. count [SJ00b]. counter [Ker70]. counters [BPT85, BM85, Ker85]. counting [FO98, TV85]. countries [MDP08]. Coupling [MFS03, AAPdG10, AC16, HKO14, KT15]. course [Bes76, SB08]. cover [AFC20, AAR^{+}24, CCS20, GBM10, IGS10]. coverage [ZWY^{+}22]. covered [KEH09, TGZ09, Zdo09]. COVID [SDN25]. Crab [LMM05, Dor79, FTH22, NNJ22, RPSSDS23, SRF21, SLS15, TK19b, TK19c, VH08, YSY^{+}22, dSSR^{+}23]. crabs [FBNF08, HdC24, RMH22, TC15]. Crambionella [BJS22]. Crangon [Spa71]. crassinervis [Lar92]. crassipes [LW11, LCL16, qLlZhY16, RRRRA07, TK88b]. Crassostrea [YRR09]. Craterocephalus [WHS13]. crayfish [AP00, AFR14, CBNF19, GBA20, HGM22, HKB19, HGB20, MLG20, OLM09, PdE^{+}24, RHS18, SLS15, VEK20, YYS^{+}23, vdHD17]. crayfishes [OLM09]. created [DBG13, GD91]. creek [QSA09]. Creeks [Gys72b, Gys72a, CFP08]. Crimean [MNK04]. crispus [WYZ13]. Critical [HFS08, IBV16, ONA^{+}23, ZD21, vLJA07]. critically [BSF11, KGC10]. croaker [MCV08]. Croatia [PHB09, ŽG09]. cross [COC18, MS07, EBA96, OSO09]. cross-designed [MS07]. cross-scale [COC18]. crucian [KG03]. crude [DPL03]. crush [GD09]. Crustacae [MJ93]. Crustacea [BHG18a, KAY20, KÇA22, ÖE04, PSSKŽ24a, PSSKŽ24b, PP95, BHG18b, DGD06, Dor74, Dor77, DHEG21a, DHEG21b, FvM03, FBNF08, Gee75, KG94, Kle84, KKM09b, LXH21, MLTH15, NZS16, Pin75, Slu81, VV85, VSB10]. Crustacean [Ano97e, Jüt05, VV05, CDL13, DKI05, FLP08, GAAT^{+}24, HEB00, HEPH09, IDP19, KTJ10, OBF10, SPA12, WCT00, ŽAL24]. crustaceans [YO02, dSCT23]. cryoconite [ZBB19, ZBA19]. cryptic [ARG20]. Cryptochironomus [GKK92]. Cryptomonas [GTI99, NHiN15, PA98]. Cryptophyceae [NHiN15]. crystallinus [Par69]. Ctenophora [FNW94]. Cu [RAP22]. cuccini [BW92]. cucullata [ADS25]. cucumber [DGGW20, ECGC25, PAAJM17]. cucumbers [ABF21]. cue [THS17a, THS17b]. cues [ÅDB00, BJ99, BH22, GBPR23, Mar17, MLZ10, SAM07, WV02, Zad03, ZL07]. Culex [MMBV24]. cultivars [OMC22]. cultivated [NC20, TK89, WRS05]. cultivation [SY02]. culturable [HYH03]. culture [Deg10, FDB97, PB04, PPH16, SD09, Som83, WGL21, WZL19, YYS^{+}23]. cultured [NJ00]. cultures [dSFFA03, GBM75, HJV08, JB83, PA98]. Cumberland [KGB04, GB82, Gru11]. Cumella [GCT^{+}23]. Curimatidae [PCB10]. Current [GGV86, HW13, OLJB20, WP12, CPP22, Cla84, GBT19, LZZ^{+}25, Soi05a, Soi05b, SAM07, Vie20, Wil07]. Current-mediated [WP12]. currents [DVK81, LLL22, WLN12]. Curtis [VC89]. curvature [WV02]. curves [Fli85, VHtB17, VB98]. cutthroat [Kae20, VK12]. cuttings [DZG17]. Cyanobacteria [BSA20a, FSdO20, FA18, LRG21, RCV91, AMSN07, BWS19, BAdOFC21, GH04, HEF13, Kap80, Kap82, KG03, LWdSD16, MVI16, MdMdLB21, NMS21, PPL16, SPH16, WCD17, WYW^{+}24, XWC15, YP17, Zev82, vGVV11]. Cyanobacteria-shrimp [LRG21]. Cyanobacterial [IFV16a, MS11, BH16, BMJ16, CGO19, CZZ^{+}25, DKI05, FAC16, GL19, HSP16, IBV16, IUK22, KK12a, Kim99, MJM16, MCB22, PLdFC22, RHA00, SK16, TK19a, TSV16, VIH16, ZGZ20, ZQH16, IFV16b]. cyanobacterium [CEH12, DCL19, MHG16, NR97, RTL17, VvGV12, YZY09]. cyanocidal [MJM16]. CYANOCOST [IFV16a, IBV16, IFV16b]. cyanoHABs [BAdOFC21]. cyanophlyctis [SG20]. Cyanophyceae [Nan00]. Cyanoprokaryota [KPV16]. cyanotoxins [BSA20a, BAdOFC21, RKLB22, RKLB23]. cycle [AvdV90, BVM15, Cup94, EK93, FDH94, Gil20, GBM75, HDF88, KBO74, LSY20, Nie74, Nie79a, Nie79b, NH92, Par69, PB04, PR92, SLG09a, Sch77, VC89, Ver76a]. cycles [Goo79, JVG95, JBS05]. cycling [AH82, BBM09, YZY09, dSSB24, Cad99]. cyclopoid [NBM98, RK02]. cyclopoids [LRSN22]. Cyclops [NBM98]. cylindracea [CCS09]. Cylindrospermopsis [KPV16, LZH^{+}25]. Cymatia [DS75]. Cymodocea [AKL15]. Cyperus [Cha87, Cha89]. cyprinid [BB12]. Cyprinidae [DBO11, TP09, HPT95, DT23a, SNO98]. cyprinids [MH13, RHH17, VHN00]. Cyprinus [Raa89, SD09]. Cyprus [MTC^{+}24]. cysts [Bak94, HVB97]. cytometry [VGV94]. Czech [ČN10]. Czechia [ŠDT^{+}24a, ŠDT^{+}24b]. Czerniavsky [HGR11].

D [Ano05b, Cad08, Gos98b, Gul98, Gul09b, Laa97, SNG21, vdV99]. D. [GAAT^{+}24]. Daily [FG91]. dam [CYK13, HLL^{+}25, PdCF^{+}25, WKK24, YKJ12]. damaged [TKR10]. dammed [BFB15, RHL02]. damming [ŽG09]. dams [Pee74]. damselflies [Hig68]. damselfly [HP06]. danae [HdC24]. Danish [Møh99]. dans [Bil73, Joi73, Pod74a, Pod74b, Pod74c]. Daphinia [Ker78a]. Daphnia [AL09, BMR07, BDH09, BDP21b, BDP21a, BDG06, BV84, BHP05, BJS12, CB00a, CvDG07, Coe97, DA99, De 97, DPT04, Fli70, GMVCLGGP24, GMD19, GH06, GSK08, GX11, GH04, INJ20, Ker70, Ker78b, Ker91, KvE21, Lin78, MS07, MCvE09, NZS16, NJ00, OFCP22, PBG03, Pri03, RR98, RR03, RW98, Rin70a, Rin70b, RvGB05a, RvGB05b, RKLB22, RKLB23, SB97, TZD21, VV07, WV02, WCD17, WF12, ZLT22, ZWF11, vGR98]. dark [CFRNT21, GBM75, Ken11]. darkness [RP22, WPvB17]. dasycneme [RHO99]. Data [ABV72, Cad80, CH91a, Cla82, COC18, DH09, DG84, DÖD21, Hig79, Hog69, Kla88, Lee84, MDP08, ODI08, OB13, SMR08, SKV02, Sto76, WSvA98]. Database [Sto76]. databases [MDP08]. dataseries [PV95]. dataset [PPC08]. dates [ABF01]. daughter [MLL21]. day [Vij91]. day-time [Vij91]. DDT [FV95b, GV99]. dead [MMC14, Roo82, dSCT23]. death [Swe71]. d’eau [Bil73]. debris [HLL^{+}25, VNKdOR15, VNSFG16]. decade [KF95]. Decapoda [VV85, dCKL23, Gee75, PNC22, RPSSDS23]. decapods [JLG06, MTC^{+}24]. Decay [VM79]. Deciphering [SXC^{+}25]. Decision [CPP22, DDD07, Cra83]. Decision-making [CPP22, Cra83]. Decline [GL19, JGH97, Kae20, NEP05a, NEP05b, PPK^{+}23]. declines [CPM09, MLL21]. decomposed [BH82]. decomposing [FVW03]. Decomposition [BMA00, BVD82, LAM06, Pel82, AH82, BLB^{+}23, BZD82, Bla82, BT23, FCG16, HC09, LDDS82, MR87, OMC22, PLC22, RGF21, SMC21, SGJ^{+}24, VVB82]. deconstructed [TCH18]. decrease [EK98]. decreasing [vdTS97]. deduction [KM83]. deel [Ano68a]. deep [BBD17, BBM21, CMB07, Cap75, DT23a, GSB13, GCN^{+}23, HKB19, JGM^{+}23, KPS09, Kor14, LPT21, LDW93, SJJA23, TFT10, VK80, VGV82]. deep-sea [TFT10]. defence [TBBS17]. defences [VvGV12]. defense [DF11, Gil12, KvE21, LSY20, MNS05a, MNS05b]. Defenses [Van97, BA14, SHP21, YYR20]. defensive [GBPR23]. define [PDE08, ZWY20]. defoliation [qLlZhY16]. deforestation [VNKdOR15]. deformities [LA99]. Degradation [GEG91, CZZ^{+}25, DSN09, DWB10, PdCF^{+}25, PPY^{+}22, VV90]. degraded [PCGW20]. deinsten [vdB70]. Delta [EAO20, Ver80a, ZZL16, AvdV90, GD91, Nwa84, PP95, dG71, DVK78, GYW17, HHV93, Kre79, NE85, PA78, SB78, SVK94, SVW95, Vaa75, vdVP76]. Delta-area [Vaa75]. deltagebieden [dG71]. deltaic [GCS93]. DeMeester [Gos98a]. demersal [YO02]. demersum [Bes79, DCG19, MB81]. Demmerik [HV89, VH89]. demographic [GFSN07, HLSN24, LRSN22, Nan00]. demography [NEGS07, NS21, NMS21, ZXZ09]. dendrobatidis [NGSOAD20, dNdSBJdCeS21]. Denitrification [CKK99, CSV19, GGT14]. denitrifier [VL18]. denitrifying [KIJ11, PRW08]. Denmark [KN86, Lar92]. densa [DDY^{+}23, LSdS^{+}24, PMSS18, TGD16]. densiflora [CF09, SBS12]. Densities [KBR10, BPP21, GFSN04, ZBA19]. Density [De 77a, Dor87, EMT00, RvGB05a, RvGB05b, Caz82, DBG02a, Df82, GZN20, JTGS24, LB20, LDW93, NHL18, PFLM15, Pen00, RB08, SAN99, ST20, SCC19, TNC08, VvOT09, ZZL16]. denticornis [DK80]. depend [SG20]. dependant [Rin73]. dependence [HHN03]. dependent [CCH11, CBN94, EMT00, GFSN07, KPV16, LZN14, OB13, Pri03]. depletion [FN94, KBS05, Møh99, PPH16]. deposit [DL89]. deposited [BHM03]. deposition [DH22, Pay10]. deposits [VD94]. depressa [WF05a, WF05b]. depression [HC09]. Depth [CS13, GBZ21, Nwa87, BDP21b, Bui95, JGG09, KJS15, LNHN08, NvdVdlM01, PFDM99, PFT12, RvGB05a, RvGB05b, SHK92, TCL22, ZG02, ZZL16]. Depth-related [GBZ21]. depths [LZD16]. deribae [Kre79]. derivatives [lCdB70]. derived [CDL13, Har82a]. Description [MG06, LZH^{+}25, Meu91, PMD10, SS92, SLH87, vdB70]. Descriptions [Jac92]. descriptors [BMS20]. desert [DF11, Lah98]. desiccation [Dor76b, KC22, KC82, MGT17]. design [GFZ15]. designed [MS07]. Desmarest [MCV08]. desmaresti [VV85]. desmid [Coe78, dMCFS^{+}23, MNT16]. Desmidiaceae [Coe97]. Desmidiales [ČN10, MNT16]. desmids [Coe75, Coe77, CB94, Coe97, M.75, NČŠ12]. Desmodesmus [AND25]. destruction [KÇA22, Sch78]. detect [CBF19, Cra91, JGM^{+}23, dBPG00]. detected [BPA01]. Detecting [TBF09, NKKG07]. Detection [VGV94, Alt98, CJS^{+}23, GCN^{+}23, MFP01, PV95, SJJA23]. deter [SRG18]. detergent [De 80a]. detergents [Kro80]. Determinant [KND^{+}22, ASLT15]. Determinants [MŚSJ16]. Determination [MRR10, MV83, PPCH22, Sch78, WCT00]. determine [DD17, HEPH09]. determined [ABF01, HHP12, HH13, MS07, MNK04]. determines [LMBM12, TS17, WXL16]. determining [BB00, HVK93, MCvE09, SAR21]. deterministic [Gra12, SCC19, TLH18]. Detrimental [nWdLjZ21]. detrital [YS98]. detritivore [DCC20]. Detritus [SFA05, BCN12, CYW24, DCC20, DL89, HMB88, Kle84, TZD21]. detritus-based [BCN12]. developed [ÁBZ08]. Developing [Nie98]. Development [DGD06, GH04, Hos80, LXH21, MLG95, SMC08, VS09, BV78, DWB10, DvZA90, DHL16, FM99, FV74, Gul09b, HGR11, Har74, HB75, KPS09, LB20, MR11, RHO09, Roo79, SB08, SN03, TG88, Wol97, YSY^{+}22]. developmental [BDP21a, Gil12]. Developments [Ano05b, Ste83, DDF93, HBB13]. device [SR85]. devices [RHA00, WXZ12]. Diadema [SHP21]. diagnosis [ŠV09]. diagrams [DDF93]. Diamesinae [Ker92, Mou92]. diapause [DM04, LOZ24, SOC12, ZL07]. diapausing [PM00]. Diaphanosoma [DHEG21a, DHEG21b, LXH21, NMS21]. Diaptomus [PM00]. Diatom [De 84b, Den94, GL01, BCP10, BBP08, BHG16, BR87, DSN09, ERA17, FDP01, GEG91, JB83, Jan75, KND^{+}22, LKSK15, MvAV20, MNT16, RJ11, SLM17, SK04, Soi05a, Soi05b, Van82a, WPB21]. diatom-based [RJ11]. diatomaceous [GEÓ04b]. Diatoms [AV92, Jüt05, Van79a, Baa79, Baa81, Baa85, Baa88a, Baa88b, DDC20, DVK81, Hou94, Mol76, ODR22, PDD07, REH92, TKD01, Van74, Van84a, VMS94, Van96, VD94, WD00, ZNP13, dPNF91, vDSB80]. diazinon [VSA97]. Dictyosphaerium [IM06]. did [BRJ97]. Didemnum [KHG20]. Didymosphenia [BCP10, BBP08, BHG16, GBT19, OFCP22]. dieback [Har94]. Diego [Bau00]. Diel [BdSCH09, BNPC07, GL92, Grz92, LFR75, MAM13, NYT09a, Ops80, RS85, RFR91, SZH21, WB10, XCL10, CCH11, DTG09, JM08, Jar08, Ker81, KK89, Rin81a, WFFM10]. Diep [Pee74, Ver80a]. Diet [IT09, BV84, CHV91, EDLC24, GBZ21, GS04, IDP19, IRT06, JBS05, LWO01, NMS21, NBT21, NAT08, ST09, SHK92, Tol02, UOJGRL21, Vel91]. dietary [ÅDB00, HF08, KT15, MMM11]. diethyl [VGP^{+}23]. Diets [KTJ10, AMSN07, NS07, TBA03, dLFTB22]. differ [LDK20, SGD22, TCH18, TDB13, ZL10]. difference [FSL^{+}23, Rin93]. Differences [KTH14, YF18, dCKL23, BD09, CMB07, Cru10, Den74, FBL13, FBM22, HWK11, HKT16, KGB04, LCL16, NMS21, RDG86, SKH19, WHW14]. Different [EDLC24, ÁBZ08, BJ99, BHPT20, BVD82, CGC19, CMG95, CDL14, Coa07, CHF09, DTH20, DB85, GFSN04, GJW20, Hal76, HWK11, HV79, JvDS06, KA16, KMS82a, KPV16, KMK09, KÇA22, LLX11, LZD16, LOZ24, MvAV20, MOCPG08, NvZvG16, NRB21, NMP12, PPV13, PPC08, PCD11, PES99, RM00, Roi81a, RHH17, STTS18a, STTS18b, SSK14, SRI20, TŠB17, TET23, TKR10, TGV00, VRH^{+}24, VHN00, WYZ13, WDP15, WRS05, ZLT22, ZKN09a, ZWY^{+}22, ZNP13, ZGZ20, ZL10]. Differential [AP00, Irv86, LA19, MSB15, RAD21, RR98, SVM08, Sir17, AND25, Bol07, DD17, OMNC24, Dar73]. differentiates [GKM17]. Differentiation [JB87, BHG18a, BHG18b, DHW81, KHK17, dSGD17]. differentielles [Dar73]. Differently [VV07]. differing [CAPGA08, DG84, Dor87, FCG16, HFS08, PVH17, PGT20b, RR98, RDG97]. differs [KJP14]. diffuse [Bui95]. diffusion [Lee82]. diffusive [GEÓ04b]. digging [Van82b]. digitifer [SS92]. Dikerogammarus [CB16, KKM09a, PDvdV06a, PDvdV06b]. dimensional [GBM10, PMD10, TNC08]. dimensions [dCAP23]. diminished [TIF10]. Dinagat [ECGC25]. Dinaric [VIM17]. dinoflagellate [HVB97]. dinoflagellates [RP22, XCL10]. dinotefuran [LWW^{+}25]. dioxide [DGB17, MD92, RHS18, THS17a, THS17b]. Dipleuchlanis [NS21]. Diplostomum [VvOT09]. Diptera [CDW92, MCC21, SKM05a, SKM05b, WH92, WMV09, BST89, BW92, Dvô70, FLA09, GEÓ04a, GL92, GPA98, IÓG04, Jac92, Ker92, KSW92, KML19, Kre79, Lar92, LM92, Mic92, MF92, Mou92, MO92, PGV95, Par69, PR92, Sæt92, Sch92, SB88, SKK92, SS92, WW84, WAW92, WB10, Ash92]. Direct [CLRdSR23, SSB21, AL09, TJB98]. directions [CPP22, Vie20]. Directive [SRH08, SG08]. Directives [EFD99]. Dirkswijde [Sch68a]. dirt [NB22]. Dirty [AFRS23]. disappearance [PHD13]. disaster [RCSF93]. disc [SEB11]. disc-based [SEB11]. discarded [TFH^{+}24]. discharge [DSG^{+}24, Ess03, MH13, QZL^{+}24, ST09, Wil79, WMTR07]. discharged [NHF96]. discharges [CAJ11, Koo73, RU87]. discharging [Bij75]. discrimination [MR17]. discussie [BM70]. discussion [MTdS^{+}24, BM70]. disease [DR96, Nie94]. disentangle [LPT21]. Disentangling [TWC13]. disjunct [VH08]. Dispersal [Cad04, AFR14, BdPP17, BSA20b, CMBG20, HD21, KTH14, NPH03, NRB21, PBM21, PGT20b, Sir17, SPA12]. dispersal- [KTH14]. Dispersion [TE04, RPN15, TGW06]. disruptors [GMVCLGGP24]. Dissolved [DM04, MHYL23, ZV93, Dan84, DDV91, DF90, GCL99, GM16, JM08, KK07, KTJ10, MGL22, MV83, PMP94, PFDM99, RKS14, SKW21, Sch78, TGZ09, VdJ94, YBZ20]. distans [MF07]. distichophylla [DNG22]. distinct [KD82, SGJ^{+}24]. distinctness [CAPGA08]. Distribution [CMBI25, CB94, DRN09, DDT24, DG06, EK93, HWK11, HK76, Hof99, HD79b, HWS22, HvB85, ISJ10, KPS09, KIJ11, NCDS^{+}25, PAAJM17, Sch77, SHK92, VS96b, VH08, YTD02, AAA^{+}24, Bak80, BMR07, BJS22, Bes81, BBP08, BZ97, BBM21, BS95, Bro84, CZC14, CV91, CFRNT21, DDV91, DDV84, DV92, EAO20, ECGC25, Fli70, FCDA09, GJ06, Gre05, GDT08, HD93, HWG77, HHEM19, HZC95, HH91, HH13, Irv89, JLG06, KND^{+}22, KKV92a, KBR10, KCR92, Kre82, KP97, Lee75, LWO01, LLX11, LMP02, MHYL23, MDH93, MR81, MMBP21, MHG16, MJ93, MOCPG08, MOCPG09, MO92, Nie82, NDW96, dCNGF25, Nwa84, PD02, PGM^{+}24, PMVdC^{+}23, PS06, RB08, RP92, RSK93, RMW91, Roo82, RCV91, RTW17, RFW17, SGRMP20, SB08, SMGAGG00, SAACR09, SMS^{+}24, SPS22, VdJ94]. distribution [VV85, VH88, VvdMP12, Ver80a, VEK20, Vij91, VTE19, WW84, WTC09, ZG02, ZTG10, Zet97a, ŽG09, ZGZ20, vdHD17]. distributions [Bau00, FDH94, MFL09, RZG09, SJ00b, TFR91, TZ05, WHP76, WTG95]. district [BN24, De 84b, KP24]. Disturbance [FWR^{+}23b, CDL14, DN23, NAM19, PVB20, SLM17, FWR^{+}23a]. disturbed [AANO25, SGRMP20]. ditch [Bel76a, Bel76b, De 76b, Hig76, HK76, MR81, Mol76, NRVV76, RM81, TNtK00, TNtK03, Van76]. ditches [Bel79, BR81, Bel87, Hig89, HV89, KK89, Kla80, RB82, Sim94, Van89, VH89, WvLA14, vLJA07]. diurnal [CDD20]. Diuron [Blo70, HAB98, KvdM70, Rin70a, Wil70, vdW70, Blo70]. diurone [KvdM70]. Divergent [KJS15, ZYL22]. Diverse [MK05b, KBN20]. diversicolor [FSC06, Rii94]. diversion [LRMdF19]. Diversity [Alt13, DHPN^{+}24, HEB00, PPV13, PEREBM^{+}23, PSSKŽ24b, SAB05, TAN14, VHB09, AT24, ASLT15, AKL15, BSA24, BDF20, BFB15, CCM21, CBdO^{+}24, CPA15, CMV91, CCS20, CRR22, DEP22, dSDMD20, DN23, Df82, FSdO20, FO98, GPGSMM^{+}23a, GPGSMM^{+}23b, Hal76, HJ23, HLK22, JAM15, KA16, KNH07, KP24, Kop21, LTH22, LMD11, LSdS^{+}24, MKG20, NAM19, NMP12, Pad13, PAP24, Rav01, RASA13, RAN24, SEP09, SLM17, Soi05a, Soi05b, TXW^{+}24, TRP22, The02, Tok92, TDB13, TBSZ22, Van82a, Wol74, YS98, ZCZ14, dSCT23, vGVV11, PSSKŽ24a]. diving [EG04, GE04]. division [Ste94]. d’Ivoire [AGB94, KKO08]. Djibouti [BOB20]. DNA [BVM15, GPGSMM^{+}23a, GPGSMM^{+}23b, JVCJ25, OHA08]. Do [AYM09, KLG17, LLF99, SH16, SCN^{+}24, SG20, TKH20, WZL18, Wey09, AFRS23, BD09, CB00a, CB00b, CDL13, DM04, GTPH08, RBMCP21, Zev82]. Dobra [ŽG09]. DOC [Fra00, RKS14]. docking [GMVCLGGP24]. Does [MKT23, SAB10, SA95, CLL10, CPA15, MMV09, SPA12, VFSH12]. Dollard [Bar77, CV77, Dan84, De 78, ECL94, GB82, Sch77, Van77a, Van77b]. dolphin [MFTM20, PBDJ20]. domain [dNdSBJdCeS21]. domestic [Cha87, ST09]. Dominance [GBA20, KMR20]. dominant [GIK21, GSM07, Irv86, Kat92, MGT17, TTD02, YL08, Zev82]. dominated [Bro81, BV96, KKI04, Nat05, SPH16, ZBA19, vdV81]. dominates [HKB19]. domingensis [SGM22]. Dominica [VTE19]. dominicanus [SCB13]. Doñana [ES09]. done [CH83]. Donk [Dum69]. Donk-meer [Dum69]. doodsoorzaak [Swe71]. door [Ano81a, Gel70]. Doppler [PJH06]. d’Orbigny [MSDDD22]. Dordrecht [Ano05b, Cad05e]. Dore [LCS01]. dormancy [RC98]. dormant [GMD19, VPS21]. dorsal [VP12]. dortmanna [PKR12]. doses [PA98]. d’Ostende [Pod74a, Pod74b, Pod74c, Joi73]. Dotilla [LMM05]. Douala [TAN14]. douce [Bil73]. Down [VV05, AVdNM19, BBM21, DdNM22, DGBL15, dNFdNM21, FBS19, VFSH12, WT14]. downingi [SS24]. downstream [FBM22, RAD21]. downward [GEK97, vGR98]. downwelling [FAM03]. Dr. [Eng74, NB80]. dragonfly [Cru10, RIT04, TNtK00]. drainable [Raa89]. Drainage [Gos99, WvLA14]. draining [WDA15]. drawdown [LCL16, MdMdLB21]. Dreijen [Ano81a]. Dreissena [BKD92, DG84, ISJ10, KK11, Lee75, MLB16, OTS05, OMM05a, OMM05b, RVK96, WFBJ17]. Dreissenidae [Ken11]. Drenthian [De 84b]. drift [Dor76a, GPA98, Grz92, HKJ12, Hof99, SWB13, WPB08]. drifting [GL92]. Drinking [Ver72, PAP24, Ano72b]. Drinkwatervoorziening [Ano72b, Ver72]. drive [BKD^{+}24, dFBdGGMLT21, LLL22, PKE22a, PKE22b, VSM19a]. driven [dSCdMM09, CZZ^{+}25, Gra12, SNO98, TDP06]. Drivers [RAN24, FBS19, GSB13, SZH21, TKH20]. drives [CCM21, NRB21]. driving [PLM^{+}22, VNSFG16]. drought [xHCjX^{+}23, TK19a, WHS13]. droughts [GL19]. Drs [Gee76]. dry [CM95a, CPM09, FWR^{+}23a, FWR^{+}23b, vOCRM19, SPH16]. dry-land [CPM09]. drying [Mag00, PVB20]. Drymonema [NNB^{+}24]. dual [BVM15, WFFM10]. dubia [AMSN07]. dubium [LXH21]. Dublin [WF05a, WF05b]. ducklings [HF08]. ducks [EG04, GE04]. duckweed [MD05]. due [ASA06, BWS21, BBD17, Koo73, MS02, PP95]. dufresnei [GOB24]. Dugesia [Van75a]. Duinigermeer [VKG95]. Dulk [dS68]. duration [PKE22a, PKE22b, SOU14]. durations [VL18]. during [Bak79, BdSCH09, BEZ91, Bel76a, CM95a, CZZ^{+}25, CV77, FDH94, FR09, Fli85, Flo01, FG91, GEG91, HLL14, JLS09, KLG17, KPS09, LSR08, MS02, MKA79, OSM13, RC98, RFR91, RvGB05a, RvGB05b, RRL09, SDN25, SY02, STH21, VGG89, VET20, VS09, Ver76a, ZQH16]. Dursban(R) [TJB98]. Dutch [Cad78, Sto76, Van77a, Van78a, Ade71, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68h, Ano68f, Ano68g, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano70b, Ano70c, Ano70e, Ano70f, Ano71b, Ano72a, Ano72b, Ano72c, Ano72d, Ano72f, Ano72i, Ano72g, Ano72h, Ano77c, Ano97b, Ano97c, Ano97d, Ano98e, ABV72, BM70, BH71, Bak72, Beu71b, Beu71a, BKO92, Blo70, Cla82, Cra85, Cra87, Dav71, DBV96, De 00, Dee69, Dor70b, Dor71, Dor72, Dre68, DM71, Dum69, Ess71, Fli70, Fon71, Gar71, GvdH68, GB69, Gee69, GD72, GR72, Gel70, Gie71, GdK69, Gys72b, Gys72a, HVS87, HHV93, Hig68, Hig69, Hig71, Hig76, Hig79, Hog69, HZC95, Kap80, Kat92, Ker70, KKV92a, KvdM70, Lam89, Lin72, Moe71, Mol72]. Dutch [Mur69, NZ75, Nie74, NvZvG16, Oom72, Osk72, Par68, PV68, Par69, Pee71, Per69, Pis69, Rin70a, Rin70c, Rin70b, Ros69a, Ros69b, Sch68b, Sch68a, Sch72a, Sch72b, Spa71, Sto76, Swe71, VS96a, Van82b, Vel72, VH89, VHC92, Ver72, Vis88, Vos87, Web71b, Wil70, Zim78, Zoe72, dB69, lCdB70, dG71, dK71, dP71, dR92, dW71a, dW71b, hvRB69, tH71, vDSB80, vG72, vdB70, vdE70, vdW70]. Dvojno [Bra99]. Dwarf [RM98]. dweller [SM12]. dwelling [Hou94, MSW10]. Dybowski [ŠDT^{+}24a, ŠDT^{+}24b]. dybowskii [CB00b]. Dynamic [xHCjX^{+}23, DKI05, JBB07, SKV02, TAN14, TD87a, vdHD17].

Dynamics [Cad99, CB00a, De 80c, MK05a, RDE93, BM03, BDH09, BBR01, BHP05, BC23, BLG99, BF12, BHPT20, BWM86, BN24, CB00b, CP10, DMM97, DYS86, EBA96, FNS^{+}24a, FNS^{+}24b, FNW94, GMD19, GH06, GAA24, GPA98, HYH03, HAS22, HC09, KKI04, KK20, KSF95, Kir22, KM83, KIG06, LLL^{+}23, LM92, LXH21, MGCC93, dCMMB15, MBM15, MAA22, MAD10, iNTT05a, iNTT05b, NSB24, OAS13, ONA^{+}23, PLM^{+}22, PVH17, PNC22, RMS11, RM81, RFO94, SNGV06, SNKN^{+}24, Sol97, SLC10, SPH16, SGK07, TYN02, TDBW95, Van91, WCZ11, WZM^{+}23, WF05a, WF05b, WF12, YL08, Zet97b, ZLC12, ZSD94]. Dytiscidae [DB02].

early [DR96, FG91, GLGU22, Nwa95, TCL22, VSM19a, VD94]. earth [MHYL23]. earthquakes [Bra21]. East [PVC96, STT01, FD94, GS04, RNJ97, DTH20, ECGC25, FLP08]. Eastern [BV78, CP10, GL92, GK07, Mun94, WB10, BVM15, BVD82, DWB10, DDF93, Ess71, EPL^{+}23, HHN03, JLG06, MTC^{+}24, PFSPdS24, SAL16, WMM94, AANO25, KK09, KP99, Lov74, Ole97, PPCH22, Par78, Str88, Van77a]. Ebro [vOCRM19]. ecads [PVC96]. Echinodermata [ABF21]. Echinogammarus [SDRM16, ŽG09]. Echinoidea [GBZ21]. echinulata [CEH12]. Ecklonia [WKT05]. Ecological [Ano99a, Ano00d, Ano03b, Ano04d, BSW^{+}24, Cad01, CBE20, Cor78a, FBNF08, Gie83, Goe07, HPPL00, HM84, Hil73, KHK17, MSDDD22, SRH08, TLH18, VDB81, Waa80, WWS20, WMM11, AKF^{+}23, AAA^{+}24, BPP21, BvdH90, CSM08, CPP22, CMBI25, Coa07, COC18, EAO20, EBM18, FDB97, Gil20, GGV86, GYW17, HSCR10, Koo76, KWO11, LCS01, Lah98, Lam97, LDL11, LBKV18, MVI16, MFTM20, MRR10, MCSV21, NRS19, ONA^{+}23, PDE08, Rin80, RA07, RAA98, SEP09, SG08, Spa71, TJZ^{+}24, TKH20, VMS94, VVK96, VS96b, Van75a, VPP16, VGV82, VHC92, VSM19b, WWG^{+}25, ŽG09, vD83, NZ75, Per69]. ecologically [De 97]. ecologist [VBK96]. ecologists [Cra83, Cra85]. Ecology [Ano97b, Ano97c, Ano97d, Ano98e, Ano99c, Ano99d, ET16, IM06, Kim99, Kül05, Pol73, Sar05d, Ver97, ŻD06, AK09, BBP08, BOB20, BLC12, Coe77, CH83, Cra87, Den06, DL89, DMD19, DW14, ESG04, FDG97, Gro22, HBB13, JB83, KGC10, KWO11, Mit74, MH13, NNJ22, NudWB^{+}25, OTS05, PdCF^{+}25, PBDJ20, PSSKŽ24a, PSSKŽ24b, QSA09, Rin93, RNJB20, SD09, SL24, SCB13, SSN24, Sim94, TAN14, TB22, Van82a, VH88, VGP^{+}23, Wol97, dCKL23, dHC76, IFV16a, IFV16b, VGM93, Cad08, Gul07b, Moo98]. ecometabolism [Bil73]. ecometabolisme [Bil73]. ecomorphological [SNO98]. economic [KKS11, MAA22]. Ecophysiological [Muu74, TIF10, Sch97b, SS99]. ecophysiology [Ano83b, Dor74]. ecoregion [EAO20]. ecoregions [BDF20, PU15]. Ecosystem [DBG02a, HG05, BDR97, Bas20, BTP24, BH92, BR07, BBV78, BBR01, BBH05, CMH14, CHP06, Clo99, CSV06, CKK99, CM99, DJD92, DBG02b, DN23, Ess03, Flo01, GFF17, GVV86, Hui88, JAM15, KBN20, Ker83, LPT21, LSO10, LTH22, MS11, MTJ10, NudWB^{+}25, Nie78, PPY^{+}22, PSD97, PES99, PMD10, RBPF11, RCSF93, Rin76, ST09, SDW^{+}24, TFR91, TXW^{+}24, VK90, WPB21, YAC21, ZLC12, dBPG00, vdTS97]. Ecosystems [Cad99, Adm90, BVI16, Ber24, Bro98, CGC19, Cad01, CGO19, Cap75, CKK99, DSN09, DP97, DBO11, De 79c, DHL16, FBV02, GB82, GV89, Gul09b, KSY19, Ker83, Kro86, Kur02, LCS01, MR11, Nor98, SKW21, SCEGL12, Sur01, VB98, WHW14, ZBA19, dCKL23, dWvG88, Gul99]. Ecotoxicity [BKD92, Blo83, MD05]. Ecotoxicological [HAB98, AFRS23, AdAPC20, BZL22, BZG^{+}22]. ecotoxicology [Gro22]. ecto [SMS^{+}24]. ecto- [SMS^{+}24]. Ectoparasite [WNZ24, PFLM15]. Ecuador [LBKV18, MT09]. edaphic [BTP24]. eddy [MML^{+}22, MML^{+}23]. Eden [DMD95]. edge [AKL15, ZD21]. edibility [Coe97]. edible [GOB24]. Edited [Art99, Cad99, Cad05e, Gos98a, Gos98b, Gos99, Gul99, Nie98, Shi98, War97, vdV99]. Editorial [Ano83a, Ano84a, Ano84b, Dor92, Dor95, Dor96, IS13, IFV16a, IFV16b, Ano04b]. Editors [Ano68f, Ano68g, Ano69e, Ano69f]. Edku [KHF01]. Eds [Ano05b, Cad01, Gul09b]. edule [PSP91]. edulis [MGCC93, PSP91, Rei99, SS99]. eel [DTC19, OSM13]. eelgrass [Bak80, Bol07, BH82, Har82a, JLG06, LDDS82, Nie82, Nie94, NDW96, Pel82]. eels [ACL13]. Effect [AMSN07, AOLLJ24, BHL09, BAG94, Gil20, GPS^{+}23, JGG09, KML02, KK17, KC82, KMS82a, LTL09, LRMdF19, MSB10, Møh99, MML^{+}22, MML^{+}23, NS07, PBF06, RHO09, RZG09, SMGAGG00, SSI21, ST20, SK84, VGP^{+}23, VV07, WLN12, YN08a, dSSB24, BMR07, BDH09, BP94, BT92, BDG10, CB00a, DTG09, Don79, FSC06, Gib86, Gil12, GD09, HVS87, HÅB08, INJ20, Ker75, Ker78a, Ker78b, NAM19, PLdFC22, PMZMJ16, Pra78, Rao10, RC98, TK88a, TGW06, TGD16, TG19, VNKdOR15, WPvB17, WXW11, Web71b, ZG02, ZLT22, ZWY^{+}22, dSSR^{+}23, vdTS97]. effective [BRP15]. Effectiveness [KJP14, KK11]. Effects [AdAPC20, BGM08, Bel87, BHM03, BG15, BdSM20, BA07, BHB01, CYW24, CF09, CDL14, CVM21, CAFA04, De 84a, De 79b, De 78, DTA22, EPL^{+}23, dSFFA03, FGF13, GGT14, GBM75, GVV86, GK07, GANA00, HJV08, JJG^{+}25, KC22, Kla80, KMG^{+}23, KCG05, LMV09, MM96, MS08, Mar17, MR17, NN05a, NN05b, Nor98, NC20, OCS20, OLTV19, PC14, PLC22, Pen00, PMSS18, PD10, PDP02, PIM95, PES99, RVK96, RHL02, RPSSDS23, RU87, SNGV06, SLM17, SZY19, SMS^{+}24, STV06, SLC10, TBA03, TFH^{+}24, TŠB17, TM84, TNC08, TNtK00, VV05, Wan79, YYS^{+}23, ZKN09a, ZKN09b, ZNP13, ZXZ09, ZWJ19, ZQH16, dNdSBJdCeS21, AL09, AL24, Bau00, BM98, BGL13, BT01, BBM21, Bri99, BA14, CKM19, CBF19, CF96, CDL13, CYK13, CLH06, CB13, Cra91, CDD20, De 79c]. effects [DH22, DdNM22, DGB17, DCG19, EMT00, FGA^{+}24, FGB20, dNFdNM21, FC93, FBS19, GBT19, GYW17, GX11, HZZ22, HBL13, HPS11, HLL^{+}25, IAHB18, JB83, JC80, JJP24, KA97, KGM13, KGM22, Ker83, KvE21, KM17, KM83, Koo73, Kop21, LB20, LLG10, LWW^{+}25, LZY21, LAZ21, LRSN22, MAS14, dCMMB15, MLG20, Met78, NW24, ODI08, PMA22, PGA^{+}22, PS06, PT14, RKS14, RASA13, RAD21, RPS20, RW98, REH92, RAP22, RTL17, SGRMP20, SWW02, SAGN07, SvFN15, SEE22, SGJ^{+}24, SSB21, TRP22, TDHH23, TFP06, TDP06, TCH18, TJB98, Van82b, VVW95, VGG89, Van77b, Van79b, WCT00, WRH98, WPB21, WF12, WMD07, XYH11, YAC21, YYR20, Zad03, dSPdS10, vGR98, IDP19]. efficacy [MNS05a, MNS05b]. efficiencies [PBG03]. Efficiency [YKJ15, ÅDB00, BWS19, Beu71b, CMB07, DWA11, HWG77, HPS11, Ker78a, Ker78b]. efficientie [Beu71b]. effluent [HVS87, Kou83, QLZ20, SN03]. effluents [Bol04, De 79b, TK95, Wan79]. effort [BW92]. Egeria [DDY^{+}23, LSdS^{+}24, PMSS18, TGD16]. Egg [MMBP21, BGC00, DHPN^{+}24, FG91, GMD19, HEB00, PM00, PVB20, RS85, dSSSES21, VSB10, YN08a]. eggs [FV74, KNH07, LAZ21, LOZ24, MM96]. Egirdir [YAC21]. Egypt [GE80a, KHF01]. Egyptian [HEZ03]. Ehrenberg [Ste94]. Eichhornia [ATB07, LW11, LCL16, qLlZhY16, RRRRA07, TK88b]. eight [FAC16, Hal76, Lah98]. eighteen [VS96a]. Einige [De 79a]. einiger [Ban75]. Ekologie [Pol73]. Elat [PGSL02]. Elbe [BKW94, Bro94, ECL94, KSF95, ŠDT^{+}24a, ŠDT^{+}24b, WS94, WGK95]. electric [Koo73]. electrical [LDK20]. Electricity [Kel73]. Electron [VKK73, SvFN15, SVB98, Vos82]. electronic [BPT85, BM85, Ker85]. elegans [KP97]. element [FKB92, Haa74, MWC94, WHP76]. Elemental [Cad99, KBS05, BJJ12, GSM07, VK12]. Elements [SGA06, De 76a, HCS02, MHYL23, RAP22, Ver76a]. Eleocharis [HD21]. Elevated [PKR12, THS17a, CBP22, LWX19, RHS18, THS17b]. elevation [DN23, NGSOAD20]. eleven [PR84]. elicited [RMR21]. Elmidae [MB14]. Elodea [Bes76, PSZ79]. elongation [HHN03]. elucidation [RP22]. embayment [BWH02]. Embryonale [Ban75]. Embryonic [GBPR23, HGR11, Ban75]. Embryos [GK07, MV14]. embryotrophe [KBO74]. Emergence [SR92, Del92, FWB05, GL92, Pen00, SKM05a, SKM05b, VIM17, WAW92]. emergent [dMCFS^{+}23, Dvo69, LLX11, MR87, SCN^{+}24, TIF10, WvLA14]. emergent-wetland [TIF10]. emerging [MJM16, Sch92]. Emerita [FCPCMB^{+}25]. emersum [SMPK09]. EMMY [SS99]. emphasis [Gon79, HD93, Kro86, Lam79, Ste94, WH92, vL86]. emphasizing [GM16]. empirical [PVB20, WSvA98]. Employing [VPP16]. Ems [Bar77, Bro84, CV77, Dan84, DE93, De 78, EK93, ER94, Sch77, Van77a, VdJ94, Van77b, YME98]. Emydura [SSV21]. Enantioselective [LWW^{+}25]. enclosed [dSLML21, SVK94, SVW95]. enclosure [TZ05]. enclosures [LJL05, SZH20]. encroachment [JCF17]. end [LWdSD16]. end-of-pipe [LWdSD16]. endangered [BSF11, DT23a, KGC10, SYW09, WHS13, XWC15]. endangerment [ŽRP15]. endemic [GCT^{+}23, JBS05, RSM24, RPSSDS23, RTW17, RFW17, SNO98, ŽG09]. endo [PMB^{+}24]. endo-helminth [PMB^{+}24]. Endochironomus [Mic92]. endocrine [ECR20, GMVCLGGP24]. endoparasite [SMS^{+}24]. endophagous [HVK93]. endpoint [SVB98]. energetic [RKLB22, RKLB23, XYH11]. Energy [GAAT^{+}24, YL11, AL24, BH92, BPP21, FSL^{+}23, GOB24, MNK04, SA95, TC15]. engage [VPP16]. Engagement [MTC^{+}24]. Engaño [GBJM22]. Engineering [RB78, Sha78, Cor78b, Mit78, PP95, Pra78, SB78]. engineers [BR07]. enginering [VW78]. England [FED95, MJ93, PJ95, SRG18, WFH93]. English [BM03, DDF93, ZPD06]. Engraulis [IT09]. enhance [PD10]. Enhanced [VRR17, ADS25, LvBR06]. enhancement [dJG07]. enhances [DF11, HKO14, MTC^{+}24]. enige [Hig71]. Enkele [ABV72, GD72, Gys72b, dW71b, DM71, Web71b]. enriched [TK88a, TK89, TK95]. enrichment [Clo99, DGB17, HMH05, MA95, PHB09, RNJB20, TLF19, TBF99, VBD96]. enrichments [Kop21]. ensis [TYN02]. Enteromorpha [KFS04, MSK93, Sch95]. entirely [Kae20]. Environment [Ver73, AXR15, BD74, BCN12, Beu74, Bil73, Blo70, BP92, CL11, Coe77, Del92, Dor70b, Har73b, Hei88, JBB07, KPJP23, LZZ^{+}25, Mur74b, OSR88, Rao10, Sha78, SdAdC^{+}25, VIM17, WKT05, ZXH15, Bak72, dP71]. Environmental [BB00, Coe78, DGN21, DG06, GWF92, HG05, HAS22, MAS14, Mar01, Met78, MTS11, MQC93, PA78, SB78, TCH18, AOLLJ24, Bal09, BSA24, BPA01, BAdOFC21, CSK10, CGL21, COC18, CKK99, Cra83, Cra85, CM95b, dSDLdA22, ECR20, FGL19, FDP01, GIK21, HV89, HNiN03, JVCJ25, LPT22, LLL^{+}23, MGCC93, MFL09, MJJ21, MKS22, MBFD24, MAD10, NRS19, NAM19, NMP12, OLJB20, ONA^{+}23, PBM21, PJF01, PVV07, PMZMJ16, PGT20b, RdMJN22, Rei92, RMH21, RTW17, RFW17, SMGAGG00, SVM08, SZH21, SVW95, ŠNZ21, Str93, TWC13, TJZ^{+}24, TMM21, TAN14, TBF09, TDL23, TLH18, VW78, VM25, WBR17, WZL18, WLB22, Web71b, WFBJ17, ZCZ14]. environments [ARP14, AV92, BZL22, BZG^{+}22, FASdPN^{+}25, GBZ21, GPV19, JDF23, Lov74, PCGW20, SASA^{+}24, SSK14, SCB13, TBS18, Vos82, dSGD17]. enzyme [TW03, VM25, nWdLjZ21]. enzyme-clay [TW03]. enzymes [xHCjX^{+}23]. ephemerae [Jac92]. Ephemeral [ABA^{+}24, AYM09, CDD20, KJF00]. ephemerality [VSB10]. ephemerella [JGH97]. Ephemeroptera [HH08, Hig69]. ephippia [Sir17]. Ephoron [BKO92]. Ephydatia [GPS^{+}23]. epibenthic [HHV93, PIM95]. epibionts [DKI05]. epidemiology [Haa73]. epifauna [SMGAGG00]. epigean [MSB10, TP09]. epilithic [PFT12, TKR10, WD00]. epipelic [WD00]. Epipelon [TLF19]. Epiphyte [RTJ11]. epiphytes [JTL09, NLF10]. Epiphytic [PDD07, PLdFC22, FGF13, Gon79, MNT16, Tok92, WD00]. epiphytism [HEF13]. epiphyton [SAF10]. Episodic [MFS03]. Epoicocladius [Jac92]. EPS [PPL16]. equal [OTJ19]. equally [RBN20]. equation [Ald79, DGBL15, PLM^{+}22]. equatorial [FRP08]. equivalence [SA95]. erecta [RPS20]. Erie [MFL09, Bou07, Rat14]. Erodibility [NLC06]. erosa [PA98]. erosion [NLC06, PVB20, TFP06, TDP06]. Errata [Ano78b, Ano92a]. Erratum [IFV16b, THS17b]. erythraeum [WYW^{+}24]. Esox [Gri89]. especially [Mur69, Nie79a]. Essai [Pod74a, Pod74b, Pod74c]. essential [GSK08]. establishment [BDH09, RG07, SPA12]. esters [Kow94, ZXZ09]. estimates [BSA14, GJ06, Gud04]. Estimating [Bir78, AMTSJ15, SGM99]. Estimation [But81, FK81, KP99, FWS^{+}23, KKS11, ŽRP15]. estimations [Lin78]. Estonia [BHL09, NLF10]. Estonia/Russia [BHL09]. estuarial [Cor78b, RB78]. Estuaries [Cad05d, BV78, Cor78a, DMD95, DMN88, ECL94, FGL19, FED95, GKK95, MA95, MM95, MCV08, Mit78, PN94, RSK93, SCG09, ŠNZ21, TGV00, WTG95, YME98, zEGB13]. Estuarine [ER94, HVK93, ABA^{+}24, AMTSJ15, AXR15, BD74, BTA94, BSA14, Bar21, BSA20a, BS95, CV95, DSN09, Dav93, De 77a, ED95, EFD99, EK98, Ess03, FSN21, FC93, FA99, GV99, GV89, Hui88, HDF88, KFS04, KSF95, LH93, LV93, MV14, McL93, MDH93, MQC93, Nie74, NC20, NMP12, Pee74, Pra78, RB83, RAN24, RAA98, ST09, SHS93, SC99, YMB93, dHC76]. Estuary [AP12, DMD95, EAB08, PDP02, RMW91, Bak77, BV78, BTV88, Bar77, BKW94, BLS88, BZV93, Bro94, BC95, Bro84, BGC00, CM95a, CCS99, CN94, CB00c, CV77, CDA03, Dan84, DW93, DAR99, DE93, De 78, DT23b, DDPT24, DDT24, Doo82, EK93, ER94, FvM03, FDH94, FV95b, FCDA09, GSH17, HF97, HJ23, Hei88, HPPL00, HBR12, HMB88, JLG06, KP99, LEB93, LDW93, LA99, Mar93, MHE93, MJ93, Møh99, MEB93, Mor99, MSK93, Nie78, PMP94, Par78, PP95, PM94, PJ95, QPP95, RCB95, RQ93, RFO94, RDE93, SCN^{+}24, SVB97, Sch77, SHS93, Sol97, SCC19, Str93, TYN02, Tay93, TLC93, Vaa79, Van77a, VdJ94, Van77b, VOD93, VML93, VV93b, VD87, WFH93, Wel74, WDP15, WBB93, WS94, Wol74, Wol79, YMB93, ZV93, BRG02, GIK21]. Estuary [GM93, LSR08, LME10, WGK95, WMM11]. etc [Ano76c]. eternity [Ker97]. Ethiopia [vGVV11, SNO98, TK19a, WMD07]. etiology [Haa73]. Etroplus [JDJ^{+}25]. ETS [SvFN15]. Étude [Bil73]. Euchlanidae [NS21]. Eucyclops [NS07]. Eudiaptomus [NBM98, RG07, RFR91]. Euglena [ZWY20]. eukaryotes [DGN21, GBJM22]. eukaryotic [BWS19, CCK10, KKI04]. Euphlyctis [SG20]. EUR [Cad05e]. Eurasian [MNJ21, SRG18, WKK24]. Euro [Ano05b]. Europe [Bra21, SKM05a, SKM05b, BRJ97, BZ97, KG94, MQC93, PLC08, Ver79, Ano05b]. European [FLP08, AR09, BGM08, BKG16, BFS21, CSM08, DBO11, DTC19, ED95, EFD99, FED95, GUT21, HGM22, MDP08, ODI08, OTJ19, OSM13, PME08, PDE08, PU15, PPC08, Pol76, QPP95, SRH08, SG08, TŠB17, WBS22, WBS24]. Eurycercus [BJ99]. Eurytemora [GD91, NCT91]. Eutrofhication [Lee80]. eutrofiering [GdK69, Pee71]. eutrofieringsproces [dK71]. eutrophic [BV84, BASJ94, CYW24, CAFA04, DD99, DZT13, DKI05, GSM07, GGV86, HK76, Hos89, KS94, Lam89, MGL22, NHiN15, Rat17, Ric86, Sch89, SB04, SB14, SPS22, TG04, VK90, Vij91, VGM93, VHB09, XCL10]. Eutrophication [Ber75, Cad05d, Ess06, Lij80, Rin80, AvdV90, ASA06, BBD17, DBV96, De 80b, Dew80, Eng86, GSB13, GdK69, Los80, Par80, Pee71, PME08, PLC08, Sim94, SG08, SLC10, Van86, Ver80b, Wil80, dK71, vLJA07, Cad99]. eutrophied [Roe96]. evaluate [WDP15]. evaluating [dSPdS10]. Evaluation [LWdSD16, TK89, WSvA98, dJG07, Ara01, Beu74, Coe75, Gri82, HHC21, KSY19, Ker75, M.75, Mar01, NAT08, VK80]. evenness [The02]. event [HLL14, KP97, WBS22, WBS24]. events [Bra21, dCMMB15, PVB20]. Everglades [AP00]. Evidence [HSCR10, Jüt05, VNSFG16, Bra21, FDP01, GUT21, HVK93, JDF23, JBS05, KK12a, KT15, Nau00, OAS13, RG07, RAN24, SXY20, SZH20, ZTNW21, ZD21, dSGD17, dCAP23, OR06, PFLM15]. evidences [CCS09]. evident [Coe78]. evolution [HLL14, Joi73, Joi73]. evolutionary [OTS05]. evolving [JAM15]. examination [BR07, Wil07]. examinations [ÁBZ08]. examine [KW12]. Examining [LKM01]. example [BvdH90, ER94, Mic92, SAL16]. exampled [Lee84]. Examples [DvZA90, AD89, Rin97]. Exceptional [NNN22]. Exchange [DMN88, RCB95, SHV96, SOU14]. Exclusion [KM17, MCB22, RG07]. execution [Meu91]. exercise [GTPH08]. existence [HMB88, KKM09b]. Existing [MJM16, VHN00]. exogenous [LTL09]. exopolymers [PPL16]. exopolysaccharide [WYW^{+}24]. Exotic [PMA22, WvRvdV03a, CYW24, CBNF19, IIMRC21, OLTV19, RBAB24, UET12, WZL18]. expansion [BF12, OMM05a, OMM05b, PR16]. expansive [GKM17]. experience [RSK09]. experiences [FAC16, MDP08]. experiment [BVM15, CNA07, ERA17, IS06, MS07, RMC95, TŠB17, WGL21, WPB08, ZJN^{+}23]. Experimental [BM85, Bor73, DD99, FM10, Geu84, HVK93, Nau00, BEQP18, CW04, CVM21, HD21, KOA00, KS94, KWO11, KMG^{+}23, KCG05, LCL16, qLlZhY16, MML^{+}22, MML^{+}23, PES99, TIF10, VB98, VGJ20, ZD21, ZWY^{+}24, dSSB24, Baa79]. experimentally [Bir78]. experimentele [Web71a]. experimenten [Blo70]. experiments [Blo70, Bre85, Deg10, HB85, OR06, Rin97, RP22, SMPK09, SKRB09, TV85]. experts [Cra83]. explain [RTW17]. explaining [VvdMP12]. explains [FCS15]. explanation [Irv86]. exploitation [KEL09, dBPG00]. exploiting [JAM15]. Exploration [RJ11]. exploratory [Dre68, MvAV20]. Exploring [EAO20, JAM15, WL24, SCEGL12]. Export [Bou07, SCG09, vG84]. exposed [Dor74, GL19, HLSN24, KK12a, OMNC24, Rei99, SN03, TGV00, VS96a]. Exposure [GFSN07, Bri99, BJS12, GBPR23, JDJ^{+}25, NMS21, RW98, dSSSES21, SGJ^{+}24, VL18, VRH^{+}24]. expression [PMZMJ16]. Extension [Bou07, KG94, MTdS^{+}24, Van75a]. extensive [ODI08]. External [Roi88a, HWS22, REH92, SOD92, SCO09, dR92, vIR92]. Extinctie [Lin72]. extinction [Bra99, RG07, Lin72]. extirpation [WWS20]. Extracellular [TW03, CGO19, ŠV09]. extracting [Roi81a]. Extraction [WBB00]. extracts [AdAPC20, DDC20, OFCP22]. extreme [CGD19, HLL^{+}25, dCMMB15, MT09]. extremes [MS11]. extremophile [FTP08]. extrinsic [dNdSBJdCeS21]. exuviae [GL92, KKV92b, WW84, WB10]. eye [Har94]. eye-witness [Har94]. eyes [BSMVP24].

F [Ano05b, BJ99, Hod98, NB80, Ver97, YZ08, Mic92, Nwa95]. F. [GAAT^{+}24, Gul09b]. Fabricius [NNJ22]. facilitates [BSA20b]. facilitation [FA18, HÅB08, HW13, MLB16]. facing [NMP12, RKLB22, RKLB23]. factor [DDF93, Kap80, ŻKŻ10]. Factors [BWH02, Gys72b, HG05, HD93, KP97, PM00, RP92, SB08, Van84b, WMV09, BB00, CF09, CGL21, CSV19, CM95b, DDV84, Deg10, ECR20, FGL19, GIK21, GDB19, HJV08, HNiN03, KND^{+}22, LPT22, LLL^{+}23, Mag00, MTA15, MB14, MR11, MGC19, ONA^{+}23, PBM21, PLM^{+}22, PPK^{+}23, PVV07, PMZMJ16, PGT20b, RdMJN22, SMGAGG00, SAACR09, TAN14, TDL23, VHC92, WZL18, WLB22, Web71b, WFBJ17, dNdSBJdCeS21]. faecal [RDE93]. fails [SRG18]. fairy [Sch83, VSB10]. fairy-tale [Sch83]. Faktoren [Gys72b]. Fal [RMW91]. falsemussel [Ken11]. far [CGD19, RBMCP21]. farm [MS08, TNtK03]. farmed [LPB07, TNtK00]. farms [RMS11]. fasciatus [TET23]. fast [GXY14, WL24]. fast-growing [WL24]. fasting [MSB10]. fat [KMS82a]. Fate [Roo82, MMC14, TJB98, Vos87]. father [FDG97]. fatty [GSK08, MMM11, MKG03, MCvE09, NAT08, ST09, ŠNZ21, SGK07, TBA03, TSZ10, VK12, WLB22, WBB00, ZWY^{+}24]. Fauna [DG06, BvdH90, BMS20, BP92, Cha71, CF96, Cla87, Doo82, Haa74, HHV93, Hig81, LR92, MLB16, Mol80, PIM95, PP95, Vaa79, Van77a, Van82b, VSM19b, WH92, dSPdS10]. faunal [BPA01, Bol07, HCS02, RFF01]. Faunistic [Gys72b]. Faunistische [Gys72b]. Fausch [vdV99]. favourably [OTJ19]. Faxonius [HGM22, HGB20]. Fe [CMG95, WA84]. feasibility [PPY^{+}22, Ric86]. feast [BFS21]. features [FED95, FAL07, Lac91, TGZ09]. fecal [MNJ21]. Fecundity [BW92, KC22]. Fed [RR03, CLL10, GTI99, GC08, HLSN24, PHM21, VDB81]. feedbacks [PSD97]. feeders [DL89, vdTS97]. Feeding [AR05, BSB97, GA10, KGC10, KML19, PBDJ20, PDvdV06a, PDvdV06b, RR03, SD09, SFB04, SCB13, TSZ10, CMH14, CCC19, CS13, DHW81, FAB18, FSL^{+}23, FNW94, GUT21, Geu84, GKM22, GSK08, Gul89, GD91, Hof99, Irv86, JJG^{+}25, KKO08, Mas88, MCV08, NvdVdlM01, OCS20, PFT12, QCF19, Rii94, SH16, SvFN15, SHP21, TTD02, VEK20, VH08, ZPD06, ZYL22, ZSD94, dCdCP19, dCKL23]. female [GBA20, KBO74, LRSN22]. females [LAZ21, LOZ24, YN08b, ZLT22]. femorata [Leh04]. fenestrata [LMM05]. fens [VBD96]. fern [SZY19]. ferns [GDB19]. Ferrissia [HVS87, Van91]. fertilisers [SMPK09]. fertilization [TNC08]. festivus [CM05, Fer92b, MF92]. Fiction [SV92]. fidelity [WD00]. Fieber [Kre82]. Field [GK07, RSK93, Sim94, WP91, CYW24, DPL03, GS11, OLM09, Rin97, SKRB09, SY02, SZH20, VN22, WPB08, vdB70]. fields [Fer92a, KGM13, KGM22, LPB07, WRS05]. fighting [Dew80]. figs [Har73a]. figured [Bak94]. filamentous [BHM03, Hil73, MMV09, PIM95, WCD17]. filholi [TK19b, TK19c]. filipendula [JTL09]. filled [BDH09, PT14]. filling [Van78b]. Filter [Rii94, Hof99, JJG^{+}25, WAW92]. Filter-feeding [Rii94, Hof99, JJG^{+}25]. filterfeeders [Gul76b]. filtering [TCH18, TLH18]. filters [PMVdC^{+}23]. filtration [MFP01, PHM21, zEGB13]. fin [VP12]. final [Roo82]. Finding [SKRB09]. Findley [SR92]. Finds [Hol70]. Fine [GC05a, GC05b, ZBB19, Dro84, FN94, KCG05, SFA05]. fine-grained [FN94]. Fine-scale [GC05a, GC05b, ZBB19]. fingerlings [Gri82]. Finland [HHP12, KPS09, VS09, JLS09, NSK08, VvOT09]. Finnish [SK04]. fire [AOLLJ24]. First [FA99, GTPH08, KKO08, MRD89, RBAB24, SS24, ŠDT^{+}24a, ŠDT^{+}24b, BHPT20, HLL14, KA86, MSDDD22, VKG95, VGG89, VHC92, ŽRP15]. first-order [BHPT20]. Fischer [NS07]. Fish [DEP22, FWS^{+}23, GvZ71, HG05, Had79, HvAV83, IRT06, KT15, NvdVdlM01, OSO09, SFB04, WHW14, ZZL16, dSPdS10, vD83, ARP14, ATB07, ABDPPP^{+}24, AMTSJ15, AL24, BMR07, BMA00, BGM08, BJ99, BMO15, BKW94, BSF11, BKC11, Bra21, CCM21, CBF19, CHP06, CCC19, CJS^{+}23, CC99, DTA22, DdNM22, DTG09, DT23a, DT23b, DDT24, Doo82, DMD19, DSL12, DBG13, ED95, FGL19, FGB20, dNFdNM21, FWR^{+}23a, FWR^{+}23b, FTP08, GCN^{+}23, GUT21, GKM22, GJ06, GRL20, HBO03, HKO14, Hig09, Irv89, JCS01, JM84, JBS05, KSY19, KBN20, KJP14, KvdM70, LMA10, LLF99, LJL05, Mag00, MS08, Mar17, MKV15, MMBV24, MLG95, MGC19, MTS11, MMC14, NMdAM21, Nwa84, NMP12, OLJB20, PFLM15, PPK^{+}23, PSW11, PLC22, PS06, PIM95, PCGW20, QCF19, RIT04, RAD21]. fish [RMS11, RNJB20, SGRMP20, SD09, SJJA23, SN03, SWB13, SDRM16, SGD22, SB97, SCC19, Sur01, TBA03, TM84, TC15, TBSZ22, TH16, Vaa79, VDV86, Van76, Vel91, VvdMP12, WXZ12, WT14, WHS13, WWS20, Wil70, Wil79, WPB08, WFFM10, WVM18, YAC21, YKJ12, ZPD06, ZCZ14, dSGD17, dSSB24, dCdCP19, dCAP23, dJG07, vGR98, vdV99]. fish-farm [MS08]. Fisheries [Sar05d, FBV02, Gri89, MŚSJ16]. Fishermen [VBA22]. fishers [MTC^{+}24]. Fishery [Wil80, IRT06, KEL09, Ste83, vD83, Dee69]. Fishery-aspects [Wil80]. fishes [BB12, CMH14, DW14, ECB09, FdLY^{+}24, JLG06, KCR92, LHV07, PBM21, PPCH22, PW73, QCF15, THS17a, THS17b, YO02, YKJ15]. Fishing [Lin72, NHL18, MOM24, NZ75, NP09]. Fishing-induced [NHL18]. fishless [MBMV20, PCD11]. fishway [YKJ12, YKJ15]. fissa [SAGN07]. fitness [KWO11, WPvB17]. five [GPGSMM^{+}23a, GPGSMM^{+}23b, NDW96, SWF17, TIF10, VK90, VHN00, VMS09, WBR17, YN08a]. five-year [VMS09]. Fixation [Bre85, APS81, Bes79, PGSL02, SK84]. fixing [Zev82]. fjord [LCPM15]. fjord-type [LCPM15]. flagellate [BHB01, SFA05]. flagellates [KBR10]. flagship [OTJ19]. Flanders [DH01, GGD07]. flat [LEB93, Zim78]. flatfish [CCC19]. flats [BC95, Cad78, DVK81, FN94, Van78a, VS96a, Van82b]. flatworm [CB00a, Van75a]. flavescens [KW12, RVB08]. flea [BJS12, GMD19]. fleas [BRA^{+}23]. Flemish [Ano99a, Ano00d, Ano03b, Ano04d, Gys72b]. flesus [FCDA09]. Flevoland [Ber75]. flexibility [MNJ21]. flies [Hig80]. floating [Har04, Kor98, MMRG03, SZY19, SXY20]. floating-leaved [Kor98]. floc [ECL94]. flocks [KLG17]. flocs [FDH94]. flood [BEQP18, LSR08, LZY21, SLM17, dSPdS10]. Flooded [ŻD06, BEQP18, NMdAM21, Żur06a, Żur06b]. flooding [DT23a, HLL^{+}25, SJD22, TCL22, WHS13]. Floodplain [ARP14, ATB07, BdPP17, BFB15, CZC14, IRT06, KRJ09, LMA10, LSR08, MLTV18, MS11, QCF15, SKW21, WWS20]. floodplains [VVK96]. flora [Coe78]. floral [DN23]. Florida [GD09]. floristic [CSK10, KPJP23]. flos [KIG06]. flos-aquae [KIG06]. flounder [FCDA09]. flow [BH92, Boe86, DTH20, Del92, EAK11, FBL13, FCW94, Gru11, JvDS06, KSY19, LRMdF19, PKE22a, PKE22b, RPS20, RL05, RCL94, VGV94, YL11]. flowcytometer [Van85]. flowing [SM12]. flows [COC18, Sta06, VGJ20]. Fluctuating [KAN20]. fluctuation [SOU14]. fluctuations [BH16, Bar77, CMV91, DZL21, DCC09, Dor87, GEÓ04a, GB07, JM08, MG91a, PHB09, RHL02, Rod10, SEE22, WCZ11, dRFC21]. fluid [MHG16]. flume [CLH06, MG06, NLC06, OR06]. flumes [JvDS06]. fluminea [HPM20, PHM21]. fluorescence [BWS21, KP99, Rat14, Rat17, Vri80]. fluorescent [FO98]. fluorescing [GEG91]. fluorochrome [TFT10]. fluorometry [BWS21]. flushing [FV95a, HM86, Zim78]. fluvial [FAM03, YKJ12]. fluviatilis [GPS^{+}23, HVS87, NHL18, VM25, WHS13]. flux [TGE04]. Fluxes [CDA03, BNV03, GEÓ04b, JA90, SB04, SGJ^{+}24, WFH93]. flying [NNN22]. Focke [Vij91]. focus [HP06, MJJ21, NSK08, VBA22]. focussed [GBM10]. Foerskal [KMS82a, KMS82b]. foliosa [GANA00]. foliose [WKT05]. Folkert [Ess06]. following [Bra99, CPA15, EK98, HXR09, PPH16, SMJ95, YHM05, YKJ12]. Food [Cad99, CMH14, GE04, IÓG04, MTSJA19, PCGW20, RB83, SCEGL12, Sar05f, SFB04, Shi98, UKS05, VGG89, VGM93, AL09, ASA06, BTV88, BMR07, BJ99, BGL13, Beu71a, BV84, BHG18a, BHG18b, BASJ94, BA07, CB00a, CLL10, Coe97, CSV06, DWA11, dSDLdA22, DJD92, DGBL15, ECB09, EG04, FAL07, FBM22, GFSN04, Gie71, Gul76b, GX11, HKO14, HXR09, HMB88, HH13, INJ20, IDP19, JVG95, Ker70, Ker78a, Ker91, KvE21, KKD02, KMG^{+}23, LCPM15, MM96, MKG03, MCvE09, MR17, MBMV20, MAM13, MDV03, iNTT05a, iNTT05b, NR97, NBD^{+}25, NJ00, OLTV19, PBG03, PC14, QCF15, SH16, SAN99, SAGN07, STTS18a, STTS18b, SFA05, SCS12, SMJ03, TSZ10, TJB98, VCT13, VHN00, WXL16, WCD17, WHW14, YL11, YZ08, YYR20, ZWF11, LB10]. Food-web [VGM93]. Foodweb [HG05]. foodwebs [Fra00, HBO03]. Foraging [HKT16, CLRdSR23, FSL^{+}23, KLG17, MS02, SAH10, ZCA20]. foraminifera [WGL21, ABT21]. forbesi [GIK21]. force [EMN04]. forces [NZ98, SCEGL12]. foregoing [CH91b]. Forest [AFC20, HHN03, IM06, RASA13, SHV96, dRFC21, VNSFG16, dNdSBJdCeS21]. Foreword [DeA10]. form [LWY^{+}24, ST20, SY02]. Formation [DBG02b, BHB01, JCL09, MMV05, RVK96, Rao10, WYZ13, vdE70]. formed [FM99]. former [Den94, Doo82]. forming [BCP10, DCL19, GXY14, VvGV12]. Formosa [FV95a, CMG95, DGGW20, FV95b, BR87]. forms [dSFFA03]. Forskål [KS79]. forth [MHE93, Tay93]. Fortnightly [SC99]. fortunei [CB00c, PCB10]. forwards [SDN25]. Fossaria [HKT16]. fossarum [MSB10, SvFN15, SZH20]. Fototrofe [vG72]. Fouling [GE80a, GE80b, RNJ97, Hod98]. found [VKD77]. Four [VV07, AdAPC20, BV14, IAHB18, Jac92, Ker92, KH92, MCV08]. Fourth [Ano97e]. FOX [SXC^{+}25]. Foz [DMD95]. fractionated [FBR87]. fractionation [LKM01, MHYL23]. fractioned [ZWY20]. fractions [DB81, FAS05a, FAS05b, GDT08]. Fragment [LZD16]. fragmentation [RASA13]. fragments [TGD16, UET12]. Framework [SRH08, SG08, FED95]. France [BDR97, BVM15, Dor76a, DK80, GL92, RDE93, ZSD94, BGC00, DD99, LCS01, Mar93, MBFD24, SLL09, TLC93]. Francisco [GIK21, LSR08, LME10, WMM11]. Free [Pol73, BV96, MMRG03, MKA79, TFH^{+}24, WZL19]. free-floating [MMRG03]. free-living [WZL19]. Freeman [Kre79]. freezing [SLG09a]. French [Bil73, Dar73, Joi73, Pod74a, Pod74b, Pod74c, Pol73]. frequency [DDF93, Lee84, SPS22, WZL18]. fresh [Coe75, Cra85, Cra87, Dor74, GV89, Hal76, M.75, Wol00]. fresh-water [Cra87, Dor74, GV89, Hal76]. Freshwater [Bou07, GRL20, Gul99, HD79b, Kir22, Kül05, Ros92a, Sar05d, SR85, Ver97, YP17, vD97, AOLLJ24, BMA00, BV14, BPP21, BP21, Bil73, BH22, BG15, BA14, BPD15, CSA23, DT23a, DHL16, ECR20, EPL^{+}23, EDLC24, GL19, GPS^{+}23, GD91, HVS87, HGR11, HAB98, Hil73, HV79, HPM20, HPS11, HWS22, JGM^{+}23, KK12b, KTH14, KT09, KNH07, KIJ11, KKM09a, LB14, LTL09, LME10, MKG03, MLTH15, MR11, MBFD24, MFP01, NS07, NBD^{+}25, NW24, OR11, OSM13, PPCH22, PMA22, PNC22, PSSKŽ24a, PSSKŽ24b, PLM18, RBMCP21, Rat17, RHO09, RPSSDS23, RNJB20, RAA98, SGRMP20, SNGV06, SLS15, Sch89, SEB11, SHS93, SC00, SGD22, Sim94, Slu81, SVB98, SYW09, TAN14, TGW06, The02, THS17a, THS17b, Tol84, TJB98, VFSH12, Van84a, VMS94, VL18, WT14]. freshwater [WWS20, WvRvdV03a, YYR20, YKJ15, ZKN09b, ZNP13, ZXZ09]. freshwater-estuarine [RAA98]. freshwaters [CW04, GDD07, SNG21, SR88]. FRG [GB82]. Frisian [Cla87, FD94]. frisii [VPS13]. Fritsch [GBM75]. Friuli [BPP21]. Frog [GK07, CB13, KGM13, KGM22, ZL10]. frogs [CBNF19, LB20, ZYS14]. frontalis [MFTM20]. frustulum [LKSK15]. fucoid [Dor76b]. Fucus [BHM03, EMT00, HMH05, NEP05a, NEP05b]. fuhrmanni [RPSSDS23]. functie [Web71b]. function [BGC00, CC99, HGR11, HXR09, LLF99, Mar93, PDP02, VL18, WWS21, Web71b]. Functional [CPA15, AKF^{+}23, AT24, AVdNM19, ASLT15, BdSCH09, dFBdGGMLT21, CH12, DEP22, GPV19, HHA92, Hig09, IAHB18, MCC21, MKG20, MOCPG08, NMP12, PPV13, PR16, QLZ20, SWF17, STTS18a, STTS18b, SEP09, SRI20, TBSZ22, TH16, WCZ11, ZWJ19, ZWY20]. functional-approach [GPV19]. functioning [BCN12, BBH05]. functions [RA07]. fundamental [Roi88b, Sch74]. Funde [Hol70]. Fundulus [TGV00]. fungal [BR87]. fungoid [DDC20]. fungus [WP91, dNdSBJdCeS21]. Funka [LKM01]. furnieri [MCV08]. Further [Dor76a, NvZvG16, Kae20]. Furuike [HNiN03]. fuscipes [Gie83]. Future [Sar05f, Ver72, CPP22, LZZ^{+}25, dCMMB15, Vie20]. fuzzy [DDT24]. Fysico [Dum69, Gys72b]. Fysico-chemie [Dum69]. Fysico-Chemische [Gys72b]. fysische [Hog69]. Fytoplanktonwaarnemingen [Sch72b].

G [Gul05, Nie98]. G. [Par69]. Gadopsis [BLC12]. Gaeta [CCC19]. Galápagos [MT09]. Galaxias [MLZ10, MAM13]. Galaxiidae [MAM13]. galeata [SB97, CB00a, Coe97, MS07, Pri03, WV02]. galeata/hyalina [Coe97]. Galicia [FBV02]. GAM [ZCZ14]. Gambusia [WHS13]. gamete [TNC08]. gametogenesis [Zad03]. gammarid [WvRvdV03a]. Gammaridae [FvM03, YTD02, ŽG09]. gammarideans [KKM09b]. gammariden [Dor70b]. gammarids [Dor70b, Dor74, KKM09a, SDRM16]. Gammarus [ÅDB00, DGD06, Den74, Dor77, IDP19, MKG03, MSB10, Pin75, SvFN15, SZH20, TET23, Wil07, YTD02, ZG02, ZTG10]. gamogenesis [NSK08]. Ganga [DSL12, DSG^{+}24, TK88a, TK88b]. Gangqu [WTC09]. gaps [VGM93]. Garâat [KHF01]. garbage [TFH^{+}24]. gargarizans [ZWJ19]. Garhwal [IUK22]. gariepinus [Nwa95]. Garnalen [Spa71]. Garonne [SLL09]. Garra [TP09]. Gasterosteus [BM98, HBL13, SDRM16, VRR17]. gastric [SPB^{+}24]. gastropod [CBN94, GJW20, dSLML21, SRF21]. Gastropoda [Ban75, AAMS24, Ban75, CM05, GSG24, GD09, KC22, Van91, VTE19, WF05a, WF05b]. gastropods [GL19, Kir22]. Gazi [SHV96]. GB [Cad05e]. gear [NP09]. gebied [vdB70]. Gebruik [Ker70]. gedrag [dG71]. Gegevens [ABV72, Hog69]. Gehouden [Ano72i, Gel70]. Gejin [War97]. Geller [Gos99]. geminata [BCP10, BBP08, BHG16, GBT19, OFCP22]. gender [WCT00]. gene [PMZMJ16]. genealogies [SSS98]. genecology [Gra74]. General [BV78, Meu91, Rin81b, Eng74, Lam97, PMD10]. generalists [YF18]. generality [BR07]. generalizations [RGD10]. generalized [MKV15]. generating [Met78]. generation [Kel73, Baa81]. generations [YN08b]. genes [SXC^{+}25]. Genetic [AKL15, LSdS^{+}24, SSS98, SYW09, vGVV11, De 97, GFW24, Pad13]. genetics [ISJ10]. genome [SPB^{+}24]. genotypes [RvGB05a, RvGB05b]. genotypic [NHiN15]. genotyping [OHA08]. genuine [vdVP76]. genus [DS75, Den74, HEF13, Hey92, KSW92, Zet97a, ŽRP15]. geo [SCC19]. geo-statistical [SCC19]. Geochemical [dG71]. Geochemisch [dG71]. Geochemistry [LEB93, GFF17, LTH22, PJ95, ZV93]. geographic [CGL21, SH11]. Geographical [HD79b, Lee75, LMP02, PD02, WTG95, Zet97a]. geographically [TDB13]. Geographische [HD79b]. geological [PD02]. geologically [HLK22]. Geometry [NWL11]. geomorphologies [CAPGA08]. Geophysical [ON88]. georganiseerd [Gel70]. geotaxis [YSY^{+}22]. geothermal [CAJ11]. German [Ban75, De 79a, GKK95, Hol70, HD79b, ISJ10, WGK95]. Germany [SKM05a, SKM05b, Bra01, HH08, KEH09, WS94]. germination [HD21]. germlings [BHM03]. Gerromorpha [NCDS^{+}25]. Gervais [Dor77]. geschikt [Gie71]. Gevolgen [Blo70, NZ75]. gezond [vdW70]. Gezuiverd [Zoe72]. Ghana [NAM19]. Ghats [BP21]. ghost [FGA^{+}24, NNJ22, PGA^{+}22]. Giant [ZL10, BFS21]. gibba [MD05]. gibbosus [MMM11]. gibelio [RRK15]. Gieskes [War97]. giftigheid [Rin70a, Wil70]. gigantea [KT09]. gigas [YRR09]. Gijster [HvB85]. Gill [FGL19, vdE70]. Ginsburg [FMdCFdCV24]. Girard [Van75a]. girgensohnii [HHN03]. Gironde [BGC00]. Girt [NB22]. Giulia [BPP21]. glacier [ZBA19, FM99]. glanis [WBS22, WBS24]. glauca [GGT14]. glehnii [HHN03]. glenii [RPN15, ŠDT^{+}24a, ŠDT^{+}24b]. global [Del92, WF12]. globosa [JB87]. globosus [KC22]. Gloeotrichia [CEH12]. Glossiphonia [AR05, Cup94]. Glossiphoniidae [Cup94]. Glumsø [KN86]. glutathione [GX11, LSY20]. glutathione-related [LSY20]. glycogen [HDF88]. Glyptotendipes [FLA09, Hey92]. Gmel [VVB82]. Gmel. [BHV82]. Gmelin [GSG24]. Gmelinoides [TET23]. Goa [KBN20]. gobies [BKG16, PR16]. gobiid [JBS05]. Gobiidae [ECR20, FV74, LA99]. goby [BF12, FV74]. Godey [Baa85]. Gölköy [Kül05]. Gom. [Ber75, Bij75, VZM75]. Gomont [VM79]. gonad [Rei99]. Gonionemus [Bak80]. Gonyaulax [VPF21]. Good [MCvE09, TMM21]. goosanders [KLG17]. Gorge [XCL10]. gorgo [NNB^{+}24]. Gosse [NS21]. Gotland [Ole97]. governance [TJZ^{+}24]. governing [HD93, NZ98, PVV07]. government [Dew80]. governmental [dR92]. grab [Beu71b, HWG77]. Gracilaria [GXY14, SY02]. gracilentus [IÓG04]. gracilis [NBM98, RG07, RFR91, SY02]. gradient [BGM08, Dav93, De 77a, Den94, FC93, IUK22, KFS04, KJS15, KRATA20, LV93, MTA15, MHE93, NvdVdlM01, PFT12, RBPF11, WiNDF22, YS98, YMB93]. gradients [AT24, FAM03, HVK93, LKSK15, LDW93, McL93, MGCC93, MDH93, MJJ21, MJ93, MKS22, MQC93, RBN20, RQ93, VSM19a, WBB93, Art99]. grained [FN94]. grandient [TLC93]. grandis [Gar05]. granii [WP91]. Grass [LB10, DSN09, Van75b, ZYL22]. grasses [VGJ20]. grassland [BEQP18, CPA15]. gravel [Sch92]. grazer [FCS15, WP12]. Grazers [Jüt05, HW13, TBS18]. Grazing [CLL10, VV05, Bre85, DB85, EMT00, EMN04, GEG91, GTI99, Gul76b, Gul85, HB85, HÅB08, KML02, KiNS05, KMG^{+}23, MNS05a, MNS05b, PSD97, TV85, Van97, Van85, VvGV12, ZLT22]. Great [MFL09, KLG17, VBA22, LH93, PK08, PJ95]. greater [SXY20, WB10]. grebe [TE04]. Greece [DDD07]. green [AMSN07, Bij75, CB13, DCL19, GXY14, KZV02, KFS04, Kap76, KA97, dCMMB15, MMV05, NW24, RVJ98, RMH22, RPSSDS23, TNC08, WBB00, AC86]. greenhouse [DPL03]. Greening [RVJ06]. Greenland [LM92]. gregarium [RP22]. Grevelingen [BV89, BV78, Bak78, Bak79, Bak80, Doo82, HWG77, Lam79, LDDS82, Nie78, Nie79a, Nie79b, Nie82, NDW96, OI89, Pee71, Sep79, Vaa79, Veg79]. Greville [JB83]. grey [KHF01, MAS14]. griseopunctatum [SS92]. groei [Mur69, dW71a]. groepsvorming [vdE70]. Grote [vdV76]. ground [QCF19]. Groundwater [GFF17, KÁ04, BLB^{+}23, BMS20, MGL22]. group [Beu71a, DB02, KSW92, Kir22, KAY20, vdE70]. groups [AVdNM19, BdSCH09, GPGSMM^{+}23a, GPGSMM^{+}23b, KA16, MOCPG08, PPV13, SRF21, SRI20, TCH18, TKH20, WCZ11, vdB70]. growing [JM84, Rin93, WL24, dCAP23]. grown [MJEC05]. Growth [BZD81, DDY^{+}23, DDL19, DZT13, GIK21, GSK08, Ker78a, KG03, LCPM15, qLlZhY16, PBF06, Raa89, SNKN^{+}24, SH11, TTD02, TGD16, TGV00, VZM75, WYZ13, AND25, AP00, ADS25, Ald79, AMSN07, ACB12, Ara01, BV14, BB12, BT23, BWM86, Caz82, CLL10, Clo99, De 84a, DA99, Don79, EAK11, FAS05a, FAS05b, Gil12, GD09, GBM75, GPS^{+}23, GCL99, GX11, Har82a, HJV08, HMH05, HdC24, HPS11, INJ20, IGS10, JB83, JJP24, JGG09, JBS05, KFS04, KML02, KK17, Kap80, KiNS05, Ker78b, KS79, KMS82a, KMS82b, KJS15, KGB04, KA86, KPV16, KHF01, LMV09, LB20, LLG10, LW11, LLX11, LZN14, LZD16, LWY^{+}24, LYM21, MM96, MCvE09, MHG16, MKA79, MMV05, Mur69, ME84, NR97, NS21, NC20, Nwa95, OTJ19, OSM13, PFLO14, PC14, Per69, RVJ98, RNJB20, STHN03]. growth [SSR13, SB04, SZY19, ST20, SK84, Sol97, TFT10, TBA03, TZD21, TBBS17, TBF99, UET12, VVW95, Van84b, VPM82, VC89, VHN00, VRR17, WPvB17, nWdLjZ21, XWJ22, YTD02, YN08a, ZWF11, ZKN09a, Zet97b, dW71a]. growth-limiting [Kap80]. Grunow [LKSK15]. GST [GX11]. gudgeon [SAM07]. Guide [Sar05d, Van84a]. guideline [Ver76b]. Guiding [DHL16]. guild [CMH14, ZPD06]. guilds [NvdVdlM01]. Gulati [SNG21]. Gulf [CFP08, DPL03, MS08, UOJGRL21, ABT21, AAPdG10, AAMS24, CCC19, GCS93, GPGSMM^{+}23a, GPGSMM^{+}23b, Leh04, MB14, PGSL02, PEREBM^{+}23, RSM24, SS24, SAACR09, Sör82, VH08, ZLC12]. gull [GB69]. gulls [Swe71, SCB13]. gunnari [WNZ24]. guppy [GR08]. Gurney [CB00b]. gut [HF08, JTGS24, KG03, NAT08]. Gyrinus [vdE70].

H [APS81, Cad99, Gos98b, Gos99, Nie98, War97, EAK11]. H. [Dor80a]. Habitat [AAA^{+}24, CHF09, DHW81, DGGW20, Fer92b, KH92, KÇA22, PMVdC^{+}23, SAB05, Sch92, Alb04, ACL13, AXR15, BH22, BTB09, BSF11, BSA20b, BLC12, Caz82, Cup94, DTC19, DTA22, DSL12, GIK21, GC05a, GC05b, GM16, HWK11, HHP12, KLG17, KT15, KAY20, LB14, LWO01, LSR08, LTH22, MSB15, Nat05, NGSOAD20, Nie73, OR11, PPK^{+}23, Pri03, QCF19, RPN15, RA07, SOW01, SZH20, TRP22, TVD91, VSB10, VPS13, VEK20, WXW11, WD00, ZBA19, dJG07, vdHD17]. habitats [AAMS24, BA97, BHPT20, BVD82, DWB10, DD84, Dor74, DBG13, GRDPP09, JJG^{+}25, JLG06, KÇA22, LMA10, LRMdF19, MMBV24, NMdAM21, PGV95, PF07, RAN24, Rus92, SSV21, SJ00a, SM12]. habits [FSL^{+}23, KKO08, SH16, ZYL22, dCdCP19]. Habrotrocha [Dev09]. Haematococcus [DCL19]. haemobaphes [CB16]. haemoglobine [Web71b]. Hagiwara [Shi98]. Haimovici [UOJGRL21]. Hald [Lar92]. Halimeda [HEF13]. Halocaridina [CCH11]. Halodule [PEREBM^{+}23]. halophytes [BVD82, HVK93]. Halopteris [SMGAGG00]. Haloragaceae [JGH97]. hammerhead [TRHHAR10]. Han [YBZ20]. Handbuch [Gos98b]. Hangzhou [STH21]. Haq [Nie98]. Har [JB87]. Harbor [YKJ15]. Harbor-type [YKJ15]. Harbour [Rei92, SMJ95, WKH82, WA84]. hard [Waa80]. Hardbound [Cad05e]. Haringvliet [Pee74, Ver80a]. harmful [HSP16, VPP16]. harness [TC15]. harpacticoid [CGD19, KRATA20]. Harpacticoida [NudWB^{+}25]. harpacticoids [NudWB^{+}25]. hartog [VBK96]. Harv [Nie74, dHC76]. harvesting [CMB07]. hatched [LOZ24]. hatchery [KAN20]. Hatching [PFLO14, VSB10, GMD19, dSSSES21]. Hatchling [SSV21]. hatchlings [KNH07, MV14]. havanaensis [LRSN22]. haven [Per69]. Hawaiian [CCH11, JBS05]. hazard [Lah98]. HCH [VV93b]. Headley [Gul98]. headstarted [NPH03]. headwater [BLB^{+}23, CPA15, Gru11, IJ06, KCG05, RB08]. headwaters [FSN21]. health [CPM09, KWO11, LDL11, STT01, WPB21]. healthy [CM99, vdW70]. heat [BDP21b, Don79, JLS09]. heated [HVS87]. Heavy [DV92, SGM22, WKH82, Fon71, JC80, MSK93, PPK^{+}23, Sur01, WWG^{+}25, WGK95, dG71]. Held [Ano72i, Gel70]. Heliozoa [Roi88a]. Heller [PNC22]. Helminth [RHH17, PMB^{+}24, WCT00]. helophytes [CV96]. help [Bir78]. hemimetabolous [VALRdF20]. Hemiptera [NCDS^{+}25, TMM21]. hemoglobin [PGV95, Web71b]. hepatic [HPT95]. Heptagenia [PBF06]. herbicide [KA97, SVB98]. herbicides [BAdOFC21, HAB98]. Herbivore [STTS18a, STTS18b, LWX19, RTJ11]. herbivorous [CHV91, iNTT05a, iNTT05b, STHN03]. Herbivory [PLM18, HMH05, JGH97, MLTV18, YF18]. Hering [Ano05b]. Heritage [WB10, MA95]. hermit [FBNF08, SRF21, TC15, TK19b, TK19c, dSSR^{+}23]. herring [Swe71]. Heterobranchus [Nwa95]. heteroclitus [TGV00]. heterocyst [PLdFC22]. Heterogeneity [NMdAM21, BNMH12, DBG02b, RAN24, Soi05a, Soi05b, ZBB19]. heterogeneous [Del92, HLK22, MS07]. Heteroptera [NCDS^{+}25]. heteropteran [SAH10]. Heterotrissocladius [Sæt92]. Heterotrophic [BEZ91, QJD02, DGN21, GCE11, LKM01, LMP02, NYT09b, dCNGF25, SFA05]. heterotrophy [BWH02, DWA11]. hiantina [WA84]. Hidalgo [CAPGA08]. hidden [FSdO20]. Hierarchy [CSV19]. High [SRG18, Bra21, BN24, DGN21, FAC16, FGB20, FASdPN^{+}25, JTGS24, MM96, NGSOAD20, PAAJM17, RMC95, SASA^{+}24, SJD22, SR85, TBS18, TLH18, VALRdF20, WCZ11, Wet93, WMD07, WYW^{+}24]. high-altitude [BN24]. high-latitude [TLH18]. high-mountain [Bra21]. high-nutrient [TBS18]. high-salinity [FASdPN^{+}25, SASA^{+}24]. high-speed [SR85]. high-throughput [DGN21]. high-value [PAAJM17]. highland [BdSM20, MIMRRZ08]. Highly [BLP16, BGC00, RNJB20, SGRMP20, ŠDT^{+}24a, ŠDT^{+}24b, TGE04]. Himalaya [IUK22]. Himalayan [BBP08]. Himalayas [PPCH22]. hinumensis [GCT^{+}23]. Hippocampus [AXR15, FMdCFdCV24]. Hira [NHiN15]. Hiroya [vdV99]. hirsuticornis [Hol70]. Hirudinea [Cup94, VH88]. Hirudo [UT16, ŽRP15, ŽRP16]. historic [Kae20, WSvA98, zEGB13]. Historical [CH91a, Cla84, Kla88, COC18, Das07, Eng74]. histories [DMM97]. History [Cad08, AL09, BDG06, CvDG07, CRR22, DD17, GR08, MS07, NS07, NJ00, Par80, RR98, RR03, RG07, RC98, RHH17, VSB10, VSM19a, YZ08]. Hogarth [vdV01]. holbrooki [WHS13]. hole [Dev09, ZBA19]. holes [PT14, SZH21, ZBB19]. Holistic [HG05]. Holland [VD94, BV76]. Hollands [Pee74, Ver80a]. Holocene [VD94]. hololimnetic [PNC22]. Holothuria [DGGW20]. Holothuroidea [ABF21]. Home [OLM09, CYW24]. Home-field [OLM09, CYW24]. homogenization [BSA20b]. Homology [FKB92]. Hong [CM05]. Hongfeng [PLM^{+}22, WLB22]. Honour [vdV99, SNG21]. Hoogkerk [Ess71]. Hoplias [SMS^{+}24]. Horizontal [RQ93, WBB93, DTG09, HH91]. hormogonia [PLdFC22]. horned [TE04]. horrens [PAAJM17]. Host [MNT16, DH22, FGL19, HVK93, KSY19, MSDDD22, PFLM15, RVB08, VvOT09, WNZ24, WCT00, ZTNW21]. host-parasite [ZTNW21]. hosts [MSW10, WCT00]. Hot [IUK22, SPH16]. hotspot [AKL15]. houden [vdW70]. Hudson [BRG02]. Hudson/Raritan [BRG02]. Huetzalin [NEGS07]. human [BGRM^{+}24, BPA01, SSB21, TB92, Wol00, dBPG00]. human-impact [SSB21]. human-induced [Wol00]. Humber [FED95, GM93, MWC94, PM94]. humic [DDV91]. Humus [dH71a]. Hungarian [SFV09]. Hungary [PPV13, VMS09]. hunting [Joi73, Pod74a, Pod74b, Pod74c]. Huron [PHD13]. Hurricane [BDG10, CFP08]. hurricanes [Rod10]. hutchinsoni [MJEC05]. huttoni [DDL19]. hyalina [CB00a, BV84, Coe97]. hybrid [BDP21a]. hybrids [Nwa95]. Hydrachna [DS75]. Hydrachnidia [MSW10, VS96b]. Hydraulic [Coa07, Mit78, VW78]. hydraulics [HM84]. Hydrilla [KC82]. hydro [BDF20, vdHD17]. hydro-dynamic [vdHD17]. hydro-ecoregions [BDF20]. Hydroacoustic [DTG09]. Hydrobia [ACH09]. Hydrobiologia [Ano05b]. Hydrobiological [ABV72, BVV80, Bel76a, Hig89, Roi88b, Koo73, vdB70, Ano72i, GL88, Gel70]. Hydrobiologie [Ano81a]. hydrobiologisch [vdB70]. Hydrobiologische [ABV72, Gel70, Ano72i]. Hydrobiologist [NL91]. Hydrobiology [Ano05b]. Hydrobius [Gie83]. hydrocarbon [AOS^{+}24, KZV02]. hydrocarbons [OMNC24]. Hydrocharitaceae [LSdS^{+}24]. hydrochemical [DVK78, PD02]. hydrochemistry [Pos74]. Hydrocotyle [HHEM19]. hydrodynamic [CV96, JvDS06]. Hydrodynamics [dSCdMM09, Van87, Ver80b]. Hydrodynamics-driven [dSCdMM09]. hydroelectric [AL24]. hydrogen [BVM15, SVB97]. hydrogeological [BMS20]. hydrogeomorphic [ES09]. Hydrografie [Par68]. hydrographic [LCPM15]. Hydrography [DJD92, Pos74, Par68]. hydrological [ARP14, AP00, CBC99, Eng86, HD21, JBS05, MS11, WDP15]. hydrology [OBF10, RDG97]. hydromedusa [PFSPdS24]. hydromorphological [LB14, SOW01]. hydropeaking [AL24]. hydroperiod [BdSM20, LB20]. Hydrophilidae [Gie83]. hydrophylla [MSD08]. hydrophysical [BGK02]. hydrosystem [VIM17]. hydrous [MD92]. Hydrozoa [Bak80, MAA22]. hydrozoans [WMM11]. hyper [Alt98, DZT13]. hyper-eutrophic [DZT13]. hyperbenthic [MDH93]. hypereutrophic [AVdNM19, FBS19, HNiN03]. hypersaline [dSDMD20, FMdCFdCV24, MNK04]. hypertidal [PMP94]. hypertrophic [CMM05, JA90, KN86, Kap80, LZN14, OMV14, SLC10, Zev82]. hyphomycetes [CKM19, FCG16]. hypogean [MSB10, TP09]. hypolimnetic [GBT19]. hypotheses [CHP06, Wil07]. Hypothesis [Bou07, CDA03]. hypoxia [KP97, SVB97, nWdLjZ21]. hypoxic [WBS22, WBS24].

Iberia [PMM10]. Iberian [FCDA09, GKM22, RRK15]. Iboe [EAB08]. Ice [KK20, YKJ15, GBM10, JLS09, JGG09, KPS09, KEH09, TGZ09, VS09, Zdo09]. ice-covered [KEH09, TGZ09, Zdo09]. ice-off [KK20]. Ice-on [KK20]. Iceland [DRN09, EÖ04, GEÓ04b, GS04, KÁ04, Mun94, ÖE04, TE04, TGE04, TG04]. Ichkeul [KHF01]. ichthyofauna [GFZ15]. ICOLLS [CSV19]. identical [BD74]. Identification [NHiN15, Sar05d, BKC11, LGCS09, OB13, Tol84, Van84a, VKK73, YBZ20]. identified [Bak94]. identify [JGM^{+}23]. identity [BZ97]. Idotea [EMT00]. IDP [GL01]. II [Ano68a, Ker78b, Los80, Par69, SVB98, Veg79]. III [Bak79, Hal76, MD92]. IJssel [Dee69, GdK69, Ros69a, Van78c, dK71]. IJsselmeer [DV92, Dee69, GdK69, Ros69a]. Ijsselmeergebied [dK71]. ilan [ECGC25]. ilicifolius [PCD11]. illegal [JDF23]. illumination [WPvB17]. image [CNN^{+}24, RRS^{+}22, RRS^{+}23, Sta06]. imbecillis [HPM20]. imbricata [Baa88b]. immature [SS92]. immigrant [Van75a, vdVP76]. Immigrants [Vaa75]. immunis [HGM22, HGB20]. immunity [VRR17]. immunochemical [VGV94]. immunological [JDJ^{+}25]. Impact [AvdV90, BH16, Bol07, CV96, DWB10, DYS86, GFZ15, MOM24, PPK^{+}23, QSA09, Sar05d, SLL09, ŠNZ21, VSA97, WYW^{+}24, AIS02, BVI16, Bas20, BR05, BPA01, BKC11, CNA07, CFP08, CB16, DT23a, EFD99, GH06, Kir22, LPT21, MAA22, MHL05, PPL16, PVB20, PGT20a, RTJ11, SDD^{+}23, SEB11, SSB21, THS17a, THS17b, VW78, nWdLjZ21, ŽAL24, dBPG00, vDSB80]. impacted [DEP22]. Impacts [LBY17, dSSSES21, SL24, TK19c, VL18, WXZ12, AOS^{+}24, Ber24, BZL22, BZG^{+}22, BB23, CDD20, DD99, DÖD21, EDLC24, FGB20, GCT^{+}23, HSCR10, MCC21, OTS05, SGD22, TK19b]. impairment [VvdMP12]. Impedance [HS85, HS86]. Imperative [Nie98]. impingement [HvAV83]. implementation [SRH08, VSM19b]. implication [KMG^{+}23]. Implications [GK07, MA95, RPS20, BBH05, Cor78a, DBG02a, DBG13, ERA17, Gil20, JJG^{+}25, KML19, Mit78, MQC93, Mor99, OB13, ONA^{+}23, RASA13, Rin80, RA07, Sch97a, SC99, ŠNZ21, TC15, TRP22, TBS18, VGJ20, WHS13, WKT05, ZPD06, zEGB13]. implying [GL19]. Import [Dro84, Cad80, vG84]. Importance [VRH^{+}24, Clo99, DTG09, DGBL15, DW14, FV95b, GRDPP09, KFS04, LV93, MLTV18, MS11, Rat17, RCB95, RFGLSB20, VvGV12]. important [BNMH12, Kir22, PR16, SAR21]. imposed [BTB09]. impoundment [BWH02, RHL02, SB04, SMJ95]. improved [DTC19, FWS^{+}23]. improvement [BSF11, TJZ^{+}24]. improvements [CPM09, NvZvG16]. improves [PFMD00]. in-lake [BVI16]. In-stream [ECB09, DWB10, DGD06]. inadequate [RBAB24]. iNaturalist [RAO^{+}24]. incapacitating [MLB16]. incidence [CHP06]. Incisocalliope [FvM03]. included [RFP01]. Including [HG05, GCL99]. incomplete [DZG17]. Incorporating [HKJ12]. Incorporation [SKW21]. increase [LMBM12, MGT17, PIM95, dCAP23]. Increased [Rei99, BJS12, MSB15]. increases [FM10, HKH^{+}23, TGD16, VALRdF20, ZYS14, ZZL16]. increasing [KRATA20, LvBR06, MLL21]. incubation [BSA14, RMC95]. incubator [FK81]. independence [MLL21]. Index [GL01, Ano98a, Ano03a, Clo99, DH01, DGH^{+}24, HHC21, VPS13, dSSR^{+}23]. indexed [PPCH22]. India [BN24, FR09, AR05, ABA^{+}24, BJS22, BSA24, BTP24, BP21, DSL12, IUK22, KBN20, KP24, NSB24, NudWB^{+}25, QSA09, RNJ97, RSM24, RAN24, SD09, SC00]. Indian [CP10, Hod98, BBP08, CMH14, FRP08, TK95, VSM19b, Wel74]. indica [SC00]. indicate [MMV09, dCdCP19]. indicated [FBR87]. indicates [TDL23]. indication [WLB22]. Indications [Van96]. indicator [CSK10, JTGS24, Koo76, KAY20, MJJ21, MRR10, RJ11, VMS94]. indicators [AA95, AV92, BB87, BMS20, ER94, KEL09, KWO11, LDL11, MCSV21, PME08, SMR08, TMM21]. indices [PDE08, Rav01, VPS13, ZCZ14]. Indies [VTE19]. indigenous [BDH09, BG15, GAA24, KHG20, PCB10, Pin75, WvRvdV03a]. Indirect [KGM13, CLRdSR23, TJB98, KGM22]. individual [Ara01, DMM97, DW14, FCS15, dLFTB22]. individuals [FGB20, KM83, MFS95]. indole [EAK11]. Indonesia [KCR92, STT01, SDW^{+}24]. induce [DM04, SCN^{+}24, WHW14]. induced [AGB94, BWS21, BDG06, BAG94, Bra21, DSG^{+}24, FA99, Gil12, GR08, KC22, KS94, KK11, Lee82, MLZ10, MOCPG09, NZS16, NHL18, PHB09, SXC^{+}25, Van97, VvdMP12, Wol00, dPNF91, TK19a]. Inducible [KvE21, BA14]. inducing [ECR20]. Induction [GLGU22, PM00, SB97, WFBJ17, ZL07]. industrial [CMBI25]. industry [De 80a]. infauna [Lu05]. Infaunal [QPP95, GCT^{+}23]. infected [WCT00, WP91]. infection [NGSOAD20]. infections [NKKG07]. inferred [CB00c, KTJ10, MFTM20, PFLO14]. infestation [Swe71, WNZ24]. infilled [DMD95]. inflows [FAC16]. Influence [ÅDB00, ARP14, ATB07, BSA14, DVK81, ERA17, EAK11, FAS05a, FAS05b, FAL07, GDB19, Gru11, GM16, Gys72b, IGS10, KK12b, KKM09b, LW11, LYM21, MF07, MFS03, MHG16, RdMJN22, RNJB20, SOU14, AAA^{+}24, AVdNM19, BR07, BR87, CvDG07, Cru10, DWA11, Dor74, Dor79, DSL12, Ess71, FV74, HBO03, HZZ22, Jar08, JCL09, LPT22, LWO01, LSR08, LDW93, LWX19, Mar93, MCC21, MGC19, MSF21, MOCPG09, MH13, Mur69, ME84, NYT09b, OBF10, PGT20b, RVJ98, RM81, Rin73, RBN20, Rod10, SL07, Sch68a, SSR13, Sha78, SY02, Str88, TK19a, TKR10, TBSZ22, VPM82, WLB22, Wil79, Wil07, WMTR07, WFFM10, ZPD06, ZXH15, dW71a]. influenced [ACA02, CMBI25, GFZ15, GGT14, KS94, ZZL16]. Influences [TB92, AFC20, AP12, HLK22, HW13, MB14, MFL09, MTS11, RMH21]. Influencing [PMA22, ONA^{+}23, PM00, RP92, Van84b]. infochemicals [NZS16, SWB13, Van97]. inform [BRP15]. informatics [Goe07]. Information [Ano04f, Ano05c, Ano05d, BV04, Ano98h, DW14]. Infralittoral [ABT21]. infrared [Das07]. ing [Hod98]. ingestion [BGK85, PSP91, SMM23, YL08]. ingredient [TJB98]. Ingvar [Par96]. inhabitant [Dev09]. inhabiting [CMH14, MKV15, PPCH22, QCF15, SKM05a, SKM05b, TKR10, YRR09]. Inherent [BRG02]. Inhibition [SVB98, MMV05]. inhibitory [JJP24]. Inhoud [Ano68a]. Initial [VVB82, CF96, SAN99]. injury [SXC^{+}25]. Inland [KKS11, ABDPPP^{+}24, CRR22, De 74, DDV84, De 00, GEK97, Kre82, RDG97, SKV02]. Inleiding [Rin70c, Sch68b, Sch72a, Vel72]. inlet [BVV80, LCPM15]. Innate [BWS19, BWS21, VRR17]. inoculum [CB13]. inorganic [AP12, CCS99, Dan84, FGB20, GC23, LME10, NN05a, NN05b, SS99, VdJ94]. input [CYW24, GGD07]. inputs [BLB^{+}23, FV95b, KCG05, PDM95, SB04]. insect [FWB05, GMW09, HVK93, KCG05, SLL09, WvLA14]. Insecta [WW84, WB10]. insecticide [Bri99]. insecticides [Lah98]. insects [Gar71, HH08, NNN22, PCD11, RASA13, TBS18]. insekten [Gar71]. Insights [dCNGF25, AND25, PCGW20, SNKN^{+}24, TNC08, VM25, WMM11, NNB^{+}24]. installed [DLG20]. instar [PAY03]. instars [IÓG04]. Institute [VGM93, Ano68e, Ano69b, Ano69c, Ano69d, Ano70e, Ano72f, Dor71, GR72]. instituten [vdB70]. Institutes [vdB70]. Instituut [Ano72f, GR72, Ano68e, Ano69b, Ano69c, Ano69d, Ano70e, Dor71]. instream [BSF11, dJG07]. Instructions [Ano04c]. insularis [dSDLdA22, LLL22, UOJGRL21]. intact [BH82]. intake [Had79, Ker70]. integer [Lin07, MFS95]. integrate [GRL20]. integrated [BBV78, Flo01, LTH22, Ano05b]. integrating [GS11, MGVC91, MTJ10]. integration [qLlZhY16]. integrative [Bro98, PFSPdS24]. integrity [GM16, HHC21]. intensification [SSB21]. intensities [KMS82a, TKR10]. Intensity [NSK08, SB02, DN23, GPS^{+}23, PE78, SG20]. intensively [TNtK00]. Inter [FCS15, FCW94, Kir22, KGB04, Rat14, SEE22]. inter- [Kir22]. inter-annual [SEE22]. inter-basin [Rat14]. Inter-individual [FCS15]. inter-site [KGB04]. inter-tidal [FCW94]. interact [CMBG20]. interacting [GL19]. Interaction [AZD02, Eng86, EMN04, LZZ^{+}25, MMV09, KS79, KÁ04, LWX19, MAS14, MSDDD22, NC20, SCS12, SRI20, XWC15]. Interactions [BV14, BT23, CCS99, Dew80, HG05, dWvG88, Alb04, BNV03, Bri99, CW04, dSDLdA22, DH22, DCC20, DTA22, EMF14, GR97, HKB19, Jar08, KW12, KML19, LSO10, LJL05, MSDDD22, MGCC93, MGVC91, Nau00, OBF10, PB04, Pri03, RMH22, RTJ11, ST20, STV06, Vie20, WP12, YAC21, ZTNW21]. interactive [DGB17, GYW17, KA97]. interacts [LWY^{+}24]. Interannual [MOdSCP07, SB04]. interbasin [FdLY^{+}24]. intercalibration [CSM08, GTPH08]. interface [BNV03, KIJ11]. intermittent [AYM09, BC23, CAPGA08, FdLY^{+}24, FNS^{+}24a, FNS^{+}24b, iNTT05a, iNTT05b, RdMJN22, RFGLSB20, SMC21]. intermittently [CSV19]. Internal [TG04, BMJ16, BNMH12, KC82, KEH09, Lij86, SCO09]. International [Ano97e]. interplay [BCN12, DN23, Rin97]. Interpretation [BPT85]. Interrelationships [BKW94]. interseasonal [NZS16]. Intersexuality [MFS95]. interspecific [ACE10, Den74, GUT21, KK17]. interstitial [GEÓ04b, Goo79]. Intertidal [Dor76b, MK05a, AAMS24, BS95, BC95, Cra91, DVK81, DRN09, Den94, ECGC25, FV95a, FSC06, GJW20, GCT^{+}23, LEB93, LV93, MHE93, Mor99, OSR88, PJ95, QSA09, RMC95, SC99, SA95, SRI20, STG18, TFH^{+}24, TFP06, TDP06, WGL21, YME98, dBPG00]. intestinalis [KFS04, MSK93]. Intra [BLB^{+}23, OMC22, Kir22]. Intra-annual [BLB^{+}23]. Intra-specific [OMC22, Kir22]. Intraguild [IS06]. Intraspecific [FLA09, SKH19, BNMH12, Gar05, MGCC93, SZY19]. intrinsic [dNdSBJdCeS21]. introduced [AR05, BFS21, FvM03, GCT^{+}23, HSY18, OLM09, PCB10, DAO24, YAC21]. Introduction [Bel76a, Cor78b, DJD92, Har82b, HB88, Rin80, Rin81c, San76, Bra99, De 77b, Ein04, Eng74, Flo01, GV89, Pin75, Rin81b, TRP22, Rin70c, Sch68b, Sch72a, Vel72]. intrusion [HKO14, iNTT05a, iNTT05b]. Inundated [DG06, SGA06, SGGO06]. Inundation [Bau00, KRJ09]. invaded [BRA^{+}23, SCS12]. invader [GKM22, NAT08, vOCRM19, RRK15, RG07]. invaders [CW04]. Invasion [BC23, OTS05, GGT14, ISJ10]. invasions [DÖD21, PLM18, Wol00]. Invasive [LDK20, SJD22, BLP16, BG15, BHG18a, BHG18b, BT23, BHG16, BF12, CF09, CB16, DD17, DNG22, GGT14, GBA20, GFW24, HKT16, HGM22, HGB20, HBR12, JGH97, KMR20, Ken11, KJP14, LZD16, LWY^{+}24, MV14, MSB15, OLM09, PK08, PR16, PDvdV06a, PDvdV06b, PHD13, RNJB20, RTL17, SMM23, SLS15, SDRM16, SM12, SAL16, ST20, ŠDT^{+}24a, ŠDT^{+}24b, TET23, TBBS17, YAC21, dSCT23, dSSB24, dCdCP19]. invasives [KKM09a]. inventories [LB14]. inversus [BNPC07]. Invertebrate [PSW11, RRRRA07, ACE10, AD89, BD09, BPP21, CNA07, FSN21, FM99, FBV02, HCS02, HKJ12, HKH^{+}23, LNHN08, MLB16, Mol80, NB22, OB13, PU15, PGT20b, SWB13, SC09, TRP22, The02, WRS05, WPB08]. invertebraten [Sto76]. Invertebrates [LPT22, Sto76, AdAPC20, AAR^{+}24, BT01, DRM18, DLG20, HWK11, HS85, HS86, JBB07, JM84, KWO11, Kor98, Kor14, Lah98, MBFD24, ODI08, PB04, SOW01, SEB11, SL24, TDB13, VK12, ZKN09b, ZBB19]. invertivorous [HKH^{+}23]. investigated [NLC06]. Investigating [HP06, CYW24, Rat17]. Investigation [Koo73, AD89, BGK02, NBM98, SKV02, SKM05a, SKM05b, TZ05]. Investigations [Lam79, Baa79, Baa81, Baa85, Baa88a, Baa88b, Bak77, Bor73, Hil73, RSK93, Ver76b, Waa80]. investment [GANA00, VRR17].

Invloed [Gys72b, dW71a, Ess71, Mur69, Sch68a]. involving [Alb04, GAAT^{+}24]. inwerking [Web71b]. ionic [WvRvdV03a, ZKN09a]. ipkea [MTdS^{+}24]. Iran [MKG20, MNJ21]. Ireland [BDR97, WMV09, MRR10, MO92, WBB93, WF05a, WF05b]. Irish [WMM94]. Iron [JCL09, BVI16, DWV81, De 84a, MD92, VdJ94, VvdMP12]. iron-induced [VvdMP12]. irradiance [BWS21, GEK97, JGG09, KT09, Rat14, RDG97, SY02, Str88, SGR86]. irradiance-induced [BWS21]. irradiated [SMC08]. irrigated [WRS05]. irrigation [KH92, MTS11]. isabelleana [MSDDD22]. ISBN [Ano05b, Cad05e]. Ischenko [Gul98]. Ischnura [HP06]. Iseo [LPT21]. Island [AAMS24, ECGC25, LME10, NYT09a, SS24, Sch95, WF05a, WF05b, FBM22, NB22, Bal09, JVCJ25]. Islands [ECGC25, MT09, RMS11, WNZ24, KBN20, CRR22, LRG21]. isle [DD84]. Isolate [tH71]. isolated [ČN10, Cla82, LKSK15, SH11]. Isolation [DL89]. Isoleren [tH71]. isopod [EMT00, EMN04]. Isopoda [HJV08]. isopods [KML02]. isotope [AMTSJ15, BPD15, CCS09, FVW03, HF08, KTJ10, KT15, MNJ21, PLC22, PFT12, RCSF93, VP12, WRS05, WVM18, XYH11]. isotope-activity [RCSF93]. Isotopes [SFB04, BVM15, UOJGRL21, dCdCP19]. isotopic [CSV06, CS13, GSH17, KSY19, MR17, SKH19, VEK20, WMM94]. Issue [IFV16a, IFV16b, ŻD06, DeA10, Goe07, GMD10, IBV16, SNG21, TB22]. Italian [Fer92a]. Italy [ABT21, BB87, BHP05, CCC19, DCC20, Mas85, Rav01, RG07, SB88, TBF09, BST89, BPP21, LPT21, NBT21, SB08]. itime [Nie98]. IV [Nie79a]. Ivan [CFP08]. Ivory [BASJ94].

J [Art99, Gee76, Gul98, Gul09b, Nie98, Ver97, War97, vD97, vdV01]. Jaar [Bak72]. Jade [LKE94]. Japan [LKM01, vdV99, HHN03, Har04, IT09, ISY10a, ISY10b, KKS11, NHiN15, WiNDF22, YO02]. Japanese [IT09, ISY10a, ISY10b, KGM13, KGM22, OHA08, SYW09]. japonicus [IT09, LLL^{+}23, YSY^{+}22]. jar [SPB^{+}24]. Java [STT01]. jellyfish [CNN^{+}24, NNB^{+}24]. jenkinsi [Dor87]. Jezero [Bra99]. Jiulong [MHYL23]. Johor [Lu05]. join [SJ00b]. join-count [SJ00b]. Joint [YYR20]. Jong [Ess06]. Joop [FDG97]. Journal [Ano25, Gul07b]. Julian [MB06]. jumbo [FSL^{+}23]. junius [Cru10]. juvenile [BH22, DD17, FD94, FCDA09, IT09, KKS11, NBM98, SSK14, TRHHAR10, WT14]. juvenile-stage [DD17]. juveniles [DT23b, DDT24, KJS15, PAAJM17].

K.N.N.V [Har73a]. Kaart [Ano68b]. Kabalebo [VGV82]. Kabeljauw [NZ75]. kairomone [BDG06]. kairomones [vGR98]. Kalimantan [KCR92]. Kalpakkam [RNJ97]. Kansas [GWF92]. Kariba [PDD07]. karst [BD09, KJF00, MBM15, MB06, PPY^{+}22, VIM17, BMS20]. karyotype [GKK92, MF92, SKK92]. Karyotypes [Ker92, KSW92]. Katelysia [WA84]. Katrina [BDG10]. Katsuwonus [WXZ12]. Kawanabe [vdV99]. keeping [vdW70]. Kees [NL91]. Kellicottia [OLTV19]. Kelp [SCB13]. Kenya [SHV96]. Ketelven [Pis69]. Key [MK05a, GSB13, MVI16, VHC92]. keystone [FCS15]. Khakasia [PMD10, QJD02, SKV02, TSZ10, Tol02, YTD02, DG02, KKR02, PD02]. Khakassia [ZTG10]. Khambhat [RSM24]. Kiefer [BNPC07]. Kieffer [Jac92]. Killarney [MO92]. kin [Spa71]. kin. [Spa71]. kinase [GMVCLGGP24]. kind [Rin76]. kinds [RMR21]. kindtii [PWG05a, PWG05b, Vij91]. Kinetic [AND25]. kinetics [VV90]. Kingdom [Cor78b]. Kingfisher [NBT21]. Kish [SS24]. Kivu [WDA15]. Kjerfve [Cad01]. Klapper [Gos98b, Gos99]. Kleine [VDB81]. Kluwer [Ano05b, Cad05e]. knife [ZD21]. knife-edge [ZD21]. know [AFRS23, RBMCP21]. knowledge [LZZ^{+}25, MTC^{+}24, Vie20, VBA22, VGM93]. Known [De 79a, NNB^{+}24]. Knysna [Bar21]. Koitajoki [HHP12]. Koka [TK19a]. Kokerjuffers [Hig68]. Kong [CM05]. Kopacki [PHB09]. koper [Ade71]. Korea [YKJ15]. Korean [Kim99, OHA08, YKJ15]. koreni [OR06]. korte [vdB70]. Krabbenkreek [vG84]. kranswieren [Har73a]. Krauss [TK19b, TK19c]. Kreken [Gys72b, Gys72a]. Kristensen [Par96]. Kristiansen [Art99, vD97]. Kromme [BB78]. kugenumaensis [YP17]. kukunoris [ZYS14]. Kullenberg [Nie98]. kunnen [Ros69b]. Kuntz [MSD08]. Kuntze [BHV82, VVB82]. Kuosa [Cad99]. Kursiumarios [Pus94]. Kurz [BHP05]. kust [dW71b]. kutum [VPS13, VPS13]. Kützing [BI09, LKSK15]. Kuwait [AAMS24]. kwantitatieve [Beu71b]. kynetics [BV14].

L [Gos98a]. L. [Ano05b, Bak77, BR05, BM98, BHM03, BJJ12, Caz82, Cha87, Cha89, DR96, DGD06, DCG19, EMT00, Gee75, Geu84, GS04, Gri89, HVS87, Har04, Har82a, HBL13, HHEM19, JLG06, KN86, KBO74, LMV09, LDDS82, MB81, MTdS^{+}24, Mit74, MD05, OMC22, PB84, PGV95, Pel82, PSP91, RVJ98, SD09, SB14, Spa71, TBA03, TE04, Van82b, VSM19a, ŻKŻ10, dW71a]. Laboratorium [Ano72f, GR72, Pol73, Ano68e, Ano69b, Ano69c, Ano69d, Ano70e, Dor71]. Laboratory [BTA94, Gul74, Pol73, SKM05a, SKM05b, Ano68e, Ano69b, Ano69c, Ano69d, Ano70e, Ano72f, Ano83b, AD89, BZD82, CBP22, DSN09, Dor71, EMF14, dSFFA03, FLA09, GR72, GS11, GLGU22, HJV08, IS06, KC22, Ker97, MS07, MSB10, NEGS07, NBM98, PC14, PWG05a, PWG05b, PA98, RSK93, RDG86, Rin97, RP22, SZH20, TBA03, YSY^{+}22, YTD02, ZG02]. laboratory-induced [KC22]. lactuca [WKH82]. lacustrine [Mas88]. lacustris [HVS87, MKG03, TET23, YTD02, ZG02, ZTG10]. lady [VH08]. laevis [GMVCLGGP24]. Lagarosiphon [PCD11]. Lagenisma [WP91]. Lagerheim [JB87]. lagocephalus [ECR20]. lagoon [AGB94, BV89, BAG94, DGG15, DGB17, DGGW20, DBdF02, FMdCFdCV24, LGCS09, MKG20, MGC19, MTSJA19, NP09, NDW96, STH21, AOS^{+}24, CVM21, Pus94, RM00]. Lagoonal [RRRRA07, AR97]. lagoons [CSV19, KML19]. lags [Gil12]. Laguna [AAV07, CAFA04]. Laguncularia [WZL18, WL24]. Lake [AC86, Bak72, BVI16, CCK10, CZZ^{+}25, DZT13, Kap80, KiNS05, KHM04, KHF01, MFS03, PDD07, PHD13, RP08, SCO09, UKS05, ÁBZ08, AT24, AVdNM19, AK09, AR97, AAV07, Bak77, BM03, BDH09, BR05, BM98, BBD17, BBH05, Bla82, BTB09, BV84, BI09, BJJ12, Bra99, BBM21, BV96, BA07, CYW24, dSCdMM09, CSM08, CNA07, CMM05, CC94, CvNB02, CS13, CAFA04, Das07, DG02, DH09, EKK21, FSdO20, FC93, Fra00, FBS19, FAM03, GTPH08, GH06, GVV86, GDT08, GD91, HBO03, HBB13, HWG77, Hof99, HXR09, HEPH09, HvB85, IM06, JVG95, JCF17, JJG^{+}25, KK12a, Kae20, KJF00, KKD02, KRZ03, KBS05, Kro86, Kur02, LZH^{+}25, Lee82, Lee84, LPT21, LZN14, LJL05, LM92, LFR75, LZY21, MGL22, MSRB23, dCMMB15, MMRG03, MRD89, MAM13, MTJ10, MD92]. lake [MNK04, Nat05, NRB21, OBF10, OAS13, Ola92, PPV13, PBG03, PPC08, Pri03, PLC08, QCF15, RCL94, STTS18a, STTS18b, SMR08, SEE22, SC00, SMJ03, SG08, SPH16, SK16, TRP22, TKH20, Van78b, VV91, Vij91, VHN00, WWS21, WHP76, WBS22, WBS24, Wey09, WFFM10, WMD07, YAC21, ZPD06, ZWY^{+}22, ŽAL24, ZXH15, ZPW13, ZPW15, AZD02, Bak72, BD74, BV78, Bak78, Bak79, Bak80, BV89, BEZ91, BVM15, BGK02, BBM10, BBV78, BZD81, Bes81, Bes87, BJ04, BHL09, BHP05, Bou07, But81, CDL13, CV91, CHV91, CMV91, CMM05, DHW81, DSD87, DTD94, De 79b, Dee69, DBG02b, DG02, DGH^{+}24, DK80, DD81, Dum69, ESG04, Ein04, EÖ04, EHM91, FK81, Fli85, FBR87, FAM03, GZG02, GBM10, GE80a, GEÓ04b, GdK69, Gon79, Gud04, GR91, Gul89]. Lake [Hig81, HH91, HRP91, HM86, JLS09, JGG09, Kae20, KGS02, Kap76, KBR10, KPS09, KEH09, KHM04, KD82, KKR02, KÁ04, Kül05, LCF91, Lac91, Lam79, Lar92, LPT21, LSO10, LDDS82, LHV87, LTH22, LMP02, LBKV18, MKG03, MG91a, Mas85, MOM24, MR87, Meu91, MFL09, MGVC91, MG91b, NEGS07, Nie78, Nie79a, Nie79b, Nie82, NHiN15, NO85, Nwa87, ON88, OBC82, OI89, Ops80, ÖE04, PPV13, Pee71, PR92, Pri03, PMD10, QJD02, Rat14, RG07, Rin81a, RvGB05a, RvGB05b, Rod10, RGD10, Roo79, Roo82, RCV91, Ros69a, SCS12, SB88, Sep79, SR92, SKV02, SNO98, SLC10, SLH87, TTD02, TD87a, TGZ09, TE04, TGE04, TG04, TDL23, Tol02, Vaa79, Van84b, VGG89, Veg79, VS09, Vel91, Ver76a, VKK02, VHB09, VMS09]. Lake [WBR17, WiNDF22, WMTR07, WMD07, WDA15, YAC21, YSYZ23, YTD02, YKU01, Zdo09, ZGZ20, ZQH16, dK71]. lake-outlet [Hof99]. Lakes [Gos99, MFL09, APS81, ABF01, AR09, BMR07, BVV80, Bas20, BVM15, Ber75, BRA^{+}23, BPA01, BBR01, BV76, BMJ16, Bra21, BSA20b, BN24, CGC19, CMB07, CZC14, CEH12, CCS20, CSV19, DHPN^{+}24, DeA10, DDV91, DZG10, DTG09, DG84, Dor87, DGBL15, DÖD21, FAK01, FLP08, Flo01, FDP01, GSB13, Gul76b, Gul76a, GMD10, HB03, Hal76, HKH^{+}23, HLK22, HD93, HFS08, Hos89, KTJ10, KIJ11, KHF01, LPT22, Lam89, Lij91, MVI16, MKT23, MLG95, MLTV18, MJJ21, MDP08, MHL05, MOdSCP07, MOPCP08, MT09, MDV03, NMB09, NLF10, NAT08, NKKG07, NRB21, ODI08, PD02, PBR01, PME08, PDE08, PFDM99, PFMD00, PJF01, PCGW20, RFP01, REP01, RFF01, RPN15, Roe96, RZG09, RJ96, SLG09a, SSI21, SEE22, SB02, Slu81, SYW09, SRH08, SPS22, SR85, TŠB17]. lakes [TSZ10, TZ05, VK80, VS96a, VK90, VGM93, VvOT09, Wor90, ZTG10, ZPD06, Zev82, Ano87, Boe86, CB94, Eng86, GGV86, GVV86, Irv86, JA90, Kal86, Ker81, KA86, Kro86, PK08, Rin81b, SGR86, TVD91, VDV86, Van87, Van86, VVA86, Moo98]. lakes/lagoons [CSV19]. LakeWeb [HBO03]. LakeWeb-model [HBO03]. lamellatus [BJ99]. lamellibranch [Mit74]. Laminariales [WKT05]. Lamouroux [KS79, KMS82a, KMS82b]. Land [APS81, CCS20, FN94, IAHB18, AC16, CPM09, DSN09, MSRB23, MvAV20, PBR01, SSI21, SEB11, SLL09, WiNDF22, WA84]. land-locked [WA84]. Land-use [IAHB18, PBR01]. Landbouwhogeschool [Ano81a]. Lande [CB00b]. landing [Hig68]. landlocked [MLZ10, WKH82]. landmark [SJJA23]. Landscape [KA16, Oom72, HLK22, vdHD17]. Landscape-level [KA16]. landscapes [dCNGF25]. Landschap [Oom72]. Langs [Gys72a, dW71b]. Lanka [DGH^{+}24, JGUF23]. Large [CMBG20, VW78, BR05, BBF17, BRA^{+}23, BHL09, BFS21, BNMH12, BWH02, CGL21, CCK10, DDPT24, DLG20, Grz92, JM08, KCG05, MMBP21, NNB^{+}24, PPV13, PPC08, RPN15, Rod10, SB08, SKRB09, SOC12, WWS20, WPB08, WMD07]. Large- [CMBG20]. large-bodied [SOC12]. large-scale [KCG05, SKRB09, WPB08]. largely [DMD95]. largemouth [SXC^{+}25]. larger [CMB07, PSW11]. Larus [SCB13]. larva [HÅB08, SKK92]. Larvae [GK07, BSB97, CB00c, Cru10, DH22, DT23b, DDT24, Dvô70, EMF14, ECR20, Fer92b, FLA09, HD93, HVK93, IS06, IÓG04, Jac92, Jar08, JGH97, KKV92a, KKV92b, LHV07, MM96, MBM^{+}22, MKL10, Ola92, ÓP04, PFLM15, PGV95, PCB10, PAY03, RIT04, SMC08, TD87a, TNtK00]. Larval [MSW10, SFB04, ZCA20, Ban75, CBNF19, GPA98, IT09, ISY10a, ISY10b, LCPM15, PFLO14, Pen00, Sch92, SB88, TYN02, YSY^{+}22, dRFC21]. larvale [Ban75]. laser [MG06]. Late [MFL09, PAY03, SNG21, TE04]. Late-summer [MFL09]. latifolia [LAM06, TCL22]. Latin [Cad01]. latitude [TLH18]. latitudinal [BGM08]. Latreille [CB13]. Laubach [RRRRA07]. laundry [Kla80]. Laurentian [MFL09]. Lauwersmeer [Bij75]. lavaretus [HHP12]. Lawrence [FAM03, VH08]. Laxá [ESG04, Ein04, GS04]. layer [Den94, KEH09]. layers [CMB07, KS94]. leachates [SK84]. lead [BKD92, TKH20]. leads [BSA20b, BNMH12]. leaf [AdAPC20, BT01, BT23, CKM19, CB16, DH22, DCC20, FCG16, FGF13, Har82a, IAHB18, LLG10, LA19, MSF21, OMC22, QZL^{+}24, RGF21, STTS18a, STTS18b, SK84, VVB82, WL24]. leaf-associated [IAHB18]. leaf-derived [Har82a]. leaf-marking [VVB82]. learned [SDN25]. Learning [AV92, DT23a, DDPT24, JGM^{+}23, JJP24, SJJA23]. leave [OR06]. leaved [Kor98]. leaves [BZD82, BH82, LB10]. Lebanese [Mou92]. Lecane [VKD77]. lectures [Ano70c]. Leden [Ano70b, Ano72c]. leech [AR05, Cup94, UT16, ŽRP16]. Leite [UOJGRL21]. Lelystad [RU87]. lemaneiformis [GXY14]. Lemna [MRBQ^{+}25, MD05]. length [RM00, SB97, WXL16]. leniusculus [OLM09]. Lentibulariaceae [CBE20]. lentic [PGV95, SM12]. Lepadomorpha [CY95]. Lepidocephalichthys [VGP^{+}23]. Lepidoptera [JGH97]. Lepomis [EBM18, KW12, MMM11]. Leptodora [PWG05a, PWG05b, Vij91]. leptopus [DM04, PM00]. less [DEP22]. lessons [DTC19, JCF17, SDN25]. Lestagella [DRM18]. lethal [VP12]. letourneuxi [KGC10]. leucas [RMH21]. leucophaeata [Ken11]. leucopterus [GKM17]. Levee [WXW11]. level [BH16, CKK99, DZL21, DCC09, GB07, GYW17, HXR09, JM84, KA16, Ker83, LMA10, LLX11, LWY^{+}24, MLG20, PHB09, RHL02, Rod10, SGRMP20, SB88, SLC10, SOU14, TŠB17, WCZ11, ZJN^{+}23]. levels [BSW^{+}24, Cha87, GB07, HAB98, Kra95, LLG10, NvZvG16, PMP94, Rin73, RVB08, SXY20]. levenscyclus [Par69]. lewini [TRHHAR10]. Leydig [HLSN24, KG94, Vis88]. leydigii [Mas88]. Liberty [LME10]. lichtverdeling [Fli70]. Life [Cup94, DD17, FDG97, MS07, MT09, NJ00, NH92, PR92, RR03, VC89, AL09, BDG06, CGD19, CF09, CvDG07, DMM97, DVZ94, Dvo69, EK93, FGM20, GFSN07, Gil20, GR08, KNH07, LSY20, NS07, Nie74, OSM13, Par69, RR98, RG07, RC98, RHH17, Sch97b, VSB10, VSM19a, YZ08, ZXZ09]. life-cycle [LSY20]. Life-history [RR03, RG07, RC98, VSM19a]. life-stages [KNH07]. life-table [ZXZ09]. Life-time [FDG97]. Light [RP08, AC86, Bir78, Bui95, CCS99, CMB07, Clo99, DZG17, DGG15, EHM91, Fli70, FAM03, GBM75, GPS^{+}23, GCE11, Gud04, HPM20, KMS82a, KGB04, KPV16, LTL09, LW11, LJL05, ME84, Nor98, PE78, PLdFC22, RDG86, SC99, SFV09, SH11, TP09, TKR10, VRR17, WPvB17, WZM^{+}23, XWJ22, vGR98]. light-dark [GBM75]. light-limited [AC86, GBM75]. light-temperature [SH11]. lightweight [SJJA23]. lignocellulosics [BMA00]. lii [LL77]. Like [SPH16, AL09, HKO14]. Limapontia [WF05a, WF05b]. Limburg [DM71]. Limestone [MB06]. liming [BHP05]. limit [TBS18]. limitation [BR87, BA07, Clo99, GE04, HKT16, IJ06, KGB04, KP99, PHB09, Rat14, Roe96, TBF99, Van97, WiNDF22, XWC15]. limitations [ŠV09]. limited [AC86, GBM75]. limiting [DGG15, Hos80, Kap80, TDL23]. limits [DRM18, GE04, HFS08]. limnetic [SCB13]. Limnocnida [SC00]. Limnological [Cla82, Hig79, Klo76, MG91b, BV76, BN24, Cla84, Gul76a, vdV76]. Limnologie [Gos98b]. Limnology [Gos99, Meu91, Lam97, MIMRRZ08, ON88, Rin93, SLG09a]. Limnomedusae [Bak80, SC00]. Limnomysis [HGR11]. Limnoperna [CB00c, PCB10]. limosus [Gee75]. limpet [FCS15]. limpets [HVS87]. Linde [Cla84, Dav71, Hig71, Moe71]. Lindevallei [Dav71, Dvô70, Hig71, Moe71]. line [MG06]. Linear [Kop21, TXW^{+}24]. lineatus [PCB10]. linkages [CHF09, PF07]. Linking [LMA10, Wet93]. links [PU15]. linnaei [BI09, SYW09]. Linnaeus [AIS02, Gie83, NNJ22, SDRM16, VBA22]. linuron [SVB98]. linza [GXY14]. Lions [GCS93]. Lipid [KZV02, ODR22, SSR13, DDC20, xHCjX^{+}23, SNKN^{+}24]. Lipids [Gul99]. Lipiniella [MF92, SKK92, SHK92]. lipped [KHF01]. List [Ano92b, Ano95b, VHC92, Ano68d]. literature [CPP22, Ken11, KD82, Van84a]. Lithobates [CB13, CBNF19]. Lithoglyphus [VM25]. Lithuania [Pus94]. litter [AANO25, AdAPC20, BT01, BT23, BHV82, CYW24, CB16, DB81, DH22, DCC20, FCG16, LA19, MSF21, RGF21, SMC21, SGJ^{+}24]. Little [De 79a, NNB^{+}24, BM03]. Little-Known [De 79a, NNB^{+}24]. Littoral [Mas85, BR05, BT01, Cla87, DHPN^{+}24, DSD87, Dvo69, Fra00, GR91, Har73b, Hig81, KOA00, LMA10, LHV07, LMD11, MKT23, MR87, MAM13, NEGS07, NHL18, Ola92, QCF15, STTS18a, STTS18b, SVK94, SGK07, TD87a, TCH18, TKH20, VS96a, WMTR07]. littoralis [MFP01]. littorals [BBH05]. Liu [XCL10]. Live [SFB04, JGM^{+}23, Shi98]. liver [HPT95]. Living [SGA06, DVK81, GAAT^{+}24, TFH^{+}24, WZL19, dSCT23]. Liza [KHF01]. load [FAC16, FCPCMB^{+}25, Lij91, Møh99, REH92, SOD92, dR92, vIR92]. loading [BB12, BMJ16, HM86, Lij86, PES99, SGR86, TG04, TGV00]. loadings [WGK95]. loads [HSP16, SS99, SCO09, vdTS97]. Lobelia [PKR12]. Lobophora [Vie20]. lobster [ZLC12]. Local [RASA13, ZXH15, CGC19, Del92, GRL20, MB14, SKW21, SON11, dL74]. localities [DF83]. localization [FKB92]. Locations [MK05a, RSM24]. locked [WA84]. Locust [Lah98]. loggerhead [NPH03]. logging [KCG05]. logic [DDT24]. Loire [Mar93]. Lombola [Pri03]. Long [DDF93, DÖD21, EÖ04, JCS01, NEP05a, NEP05b, SN03, VSW00, BBM09, De 79c, DH09, xHCjX^{+}23, JCF17, Mar01, MKV15, SHP21, SGK07]. long-chain [SGK07]. long-spined [SHP21]. Long-term [DDF93, DÖD21, EÖ04, JCS01, NEP05a, NEP05b, SN03, VSW00, BBM09, DH09, xHCjX^{+}23, JCF17, Mar01, MKV15]. longer [CKM19]. longicornis [BTV88, BV89, FG91]. longifilis [Nwa95]. longifolia [SK84]. longimanus [KG94, NAT08, Vis88]. longispina [BDP21a]. Longitudinal [FWB05, Nwa84]. longwave [CC94]. Lonnberg [WNZ24]. Looking [ECR20]. loop [ZJN^{+}23]. Loosdrecht [Boe86, Eng86, GGV86, GVV86, Irv86, JA90, Kal86, KA86, Kro86, SGR86, VDV86, Van86, VVA86, Ano87]. loss [KiNS05, LAM06, LMBM12]. lost [BA97]. Lota [BR05]. lotic [ARP14, BH92, CDW92, CHF09, Gre05, SM12, WH92, Wil07, dSGD17]. Lough [BDR97, FDB97, Gib86]. Louisiana [BDG10, GFW24]. Lour. [MSD08]. low [BMR07, CEH12, CF09, CBP22, CvNB02, CAFA04, DZG17, Gil20, JM08, NJ00, Van89, WYW^{+}24]. low-canopy [CvNB02]. low-moor [Van89]. low-nutrient [CEH12]. low-quality [NJ00]. Lower [LKE94, NE85, CB00c, DBG13, ET16, Nwa84, vOCRM19, ZZL16, GZN20, VVK96]. lowland [BB12, CZC14, ČN10, DMD19, GPA98, Grz92, Kop21, Kou83, PF07, PVV07, SB08, SKM05a, SKM05b, Tol82, Van79a, VDB81, ZPD06]. LR [LSY20]. Lubzens [Shi98]. lucioperca [LMV09]. lucius [Gri89]. Łuczański [KK20]. lugworms [Van82b]. lutea [ŻKŻ10]. Lutra [MNJ21]. Lymnaea [VvOT09]. Lyngb. [Baa85]. Lyngbye [BBP08, OFCP22]. lysing [CDL13].

[Ano81a, Ano05b, Art99, Gee76, Gul98, Hod98, Nie98, Ver97, vdV99, KPS09]. Maarssen [Slu81]. Maarsseveen [Bes81, Bes87, But81, CB94, DHW81, DSD87, DTD94, DD81, FK81, Fli85, FBR87, HD93, Hig81, Ker81, KD82, LHV87, MR87, Rin81a, Rin81b, RvGB05a, RvGB05b, Roo79, Roo82, RCV91, SLH87, TD87a, TVD91, Van84b, Van87]. machine [DDPT24, JJP24]. Macoma [dW71a]. macquarii [SSV21]. Macro [JBB07, HKH^{+}23, MLB16, SOW01, ŽAL24]. macro-crustacean [ŽAL24]. macro-invertebrate [HKH^{+}23, MLB16]. Macro-invertebrates [JBB07, SOW01]. macroalga [GXY14, KFS04]. macroalgae [HEF13, Sim94, WTG95]. macroalgal [EMN04, KKS11]. Macrobenthic [BPP21, AAR^{+}24, Bar21, Cra91, GGD07, Lu05, SA95, Van77a]. Macrobenthos [CM95b, PW73, HF97, Kop21, Van78c, YME98]. Macrobrachium [PNC22]. macrochirus [EBM18, KW12]. macrocopa [AMSN07, GFSN07, JTGS24, LAZ21, LOZ24, Zad03]. Macrofauna [DDV84, HZC95, Mol72, BR81, Bel87, DB81, Dvo69, Gar71, Hig74, Hig76, MHE93, MQC93, YMB93, Mol72]. Macrofauna-Onderzoek [Mol72]. Macrofaunal [Dav93]. macrofaunasampling [RB82]. macrofossil [BPA01]. Macroinvertebrate [AANO25, DDD07, MB06, SMR08, SPS22, AKF^{+}23, AVD07, AAA^{+}24, BR05, BDF20, BV96, Cla87, CAJ11, EAO20, Gru11, HGM22, KK07, KNH07, KF95, KM17, LDL11, LPB07, LJL05, LRMdF19, MKT23, MBMV20, MAD10, NRB21, NHL18, Rav01, RHL02, RAA98, STTS18a, STTS18b, SK04, SBS12, SZH20, TBF09, TCH18, TLH18, VN22, VSW00, WVM18]. Macroinvertebrates [HV89, PKE22b, VH89, CTO01, CMBI25, CYK13, CMBG20, DWB10, FCG16, dNFdNM21, GDD07, GBT19, HLL^{+}25, HvB85, IAHB18, JGM^{+}23, JM08, KA16, KK12b, LTH22, MS02, NWL11, Nwa84, OMNC24, PF07, Tol82, TKD01, VNKdOR15, VSA97, VHC92, VALRdF20, PKE22a]. macrolitter [SHV96]. macronutrients [PMSS18, TK95]. macrophye [Nat05]. macrophye-dominated [Nat05]. Macrophyte [LMD11, ATB07, BV96, Cha87, DHPN^{+}24, DZL21, DdNM22, DDY^{+}23, EAK11, FGF13, GM16, HLK22, JCF17, JGH97, KT09, LLG10, LLX11, LZD16, LBY17, Mar17, MMV05, MMV09, NMdAM21, NC20, PDE08, PDD07, Pip87, PKR12, QCF19, RIT04, RHL02, STTS18a, STTS18b, SB14, SAF10, TRP22, Wor90, XWJ22]. macrophyte-free [BV96]. macrophyte-mediated [DdNM22]. macrophyte-specific [Mar17]. Macrophytes [BdPP17, Bes76, Bes81, Bes87, BPD15, CZC14, CvNB02, dMCFS^{+}23, DK80, FVW03, dNFdNM21, GTPH08, Gre05, HWK11, KK17, KH92, Kor98, LMBM12, LZL22, MLTH15, MdMdLB21, Nie79a, NLF10, PME08, PCD11, PR84, PLM18, RPS20, RSK09, SCN^{+}24, TK89, TKD01, UET12, VV91, WVM18]. macrophytic [ZWY^{+}22, dL74]. macroplastic [AANO25]. Macrothrix [Hol70]. macrotidal [RDE93, TLC93]. macrovegetation [De 76b]. Macrozoobenthic [GZN20, SVK94, SVW95]. macrozoobenthos [CMM05, Lam79, Mar01, ŻKŻ10]. Mactra [MSDDD22]. maculata [GFW24, MV14]. maculatus [MLZ10, MAM13]. made [AvdV90, BV76, DTG09, dCMMB15, VK80, Van78b, ZPW13, ZPW15]. maenas [FTH22, RMH22]. maeotica [PFSPdS24]. magazine [Dor72]. magellanicus [GBZ21]. magic [Sch83]. magna [CvDG07, Fli70, GMD19, INJ20, Ker70, Ker78a, Ker78b, Ker91, Lin78, MCvE09, OFCP22, Rin70a, Rin70b, RKLB22, RKLB23, TZD21, WCD17, ZLT22, ZWF11]. magnitude [SOU14, ZQH16]. Main [VGM93, BWH02, CCS20, PMVdC^{+}23]. Maine [ZLC12]. maintained [BFB15]. Major [BGL13, FdLY^{+}24, MTA15]. Making [JvDS06, CPP22, Cra83]. Makrofauna [Hig71, Gar71, Hig71]. malabaricus [SMS^{+}24]. malacostraca [BSA24, CCH11]. Malaysia [NYT09a]. male [FTH22, GS11, GBA20, LRSN22, WZL18]. males [RM98, SS92]. Malesia [PVC96]. Malloch [SS92]. Malmgren [OR06]. mammal [MŚSJ16]. man [AvdV90, BV76, DTG09, dCMMB15, NNJ22, VK80, Van78b, ZPW13, ZPW15, TK19b, TK19c]. man-made [AvdV90, BV76, DTG09, dCMMB15, VK80, Van78b, ZPW13, ZPW15]. man-of-war [NNJ22]. managed [OTJ19]. Management [AAV07, FBV02, Nie98, AVD07, BH16, CC99, DBG02a, dSDLdA22, DGD06, DBdF02, EFD99, FED95, Goe07, Gri89, IBV16, Klo76, Lee80, LWM89, MVI16, Mor99, OI89, PCGW20, Ric86, Ste83, TNtK00, TNtK03, VVK96, VGJ20, WRS05, vD83, Ver97]. Managua [BEZ91, CV91, CHV91, CMV91, EHM91, GR91, HH91, HRP91, LCF91, Lac91, MG91a, MGVC91, MG91b, Vel91, Meu91]. manatees [CFRNT21]. Manchurian [TCL22]. mandibleplate [NCT91]. manganese [MD92]. Mangrove [HvdGvS10, BTP24, dSDMD20, EAB08, GC05a, GC05b, SHV96, WZL18]. Mangrove-sponge [HvdGvS10]. mangroves [DDPT24, vdV01]. manipulated [LJL05]. manipulation [AL09, VGG89]. Manora [GSG24]. manual [Van82b]. Manyara [MOM24]. map [Ano68b]. mapping [Sta06, VP77]. Mar [Nie98]. Mar-itime [Nie98]. Maranhense [PMVdC^{+}23]. marbled [HKB19, HGB20]. Marennes [ZSD94, BDR97]. Marenzelleria [BRJ97, BZ97, BRB97, Boc97, BSB97, EK93, KP97, Sch97b, SVB97, Zet97a, Zet97b]. marginal [PDD07]. Margins [Nie98]. Mariana [LRG21]. Marimo [SYW09]. marina [Bol07, BH82, DR96, Har82a, Har94, JLG06, LDDS82, Nie79a, Nie82, NDW96, Pel82, Van82b]. Marine [Cad04, FO98, Hou94, KBN20, McL93, AV92, AAR^{+}24, Baa79, Baa81, Baa85, Baa88a, Baa88b, BV89, BH22, Cad01, CBN94, CPP22, CLL10, CL11, CSV06, DSN09, DDC20, Dor74, Dro84, ECR20, Haa74, HSCR10, KAY20, LLF99, MGC19, MKA79, NNJ22, OMNC24, ODR22, PEREBM^{+}23, PES99, Rao10, RHO09, RAO^{+}24, RTJ11, SWW02, SCG09, SSR13, SL24, SCB13, SAR21, TNC08, VD87, Vos82, WP91, dPNF91, Ess06]. marinus [vdE70]. maritima [CF09]. mark [DB02, Cad08]. markers [De 97, LXH21, MFTM20, SSS98, TSZ10]. market [RBAB24]. marking [TFT10, VVB82]. marl [AP00]. marmorata [ACL13]. Marsdiep [CH91a, CH91b, FG91]. marsh [BVD82, CFP08, DSN09, Dan84, DPL03, Gra74, HYH03, HVK93, KK12b, Nie73, Ran74, RA07, Wol79]. marshes [ACA02, CF09, Har73b, Hui88, LLF99, MTA15]. Mart. [RRRRA07, TK88b]. Marten [Moo98]. mass [ACE10, BH82, HK82, Kal86, LAM06]. Massachusetts [TBF99, TGV00]. Masses [RRRRA07]. Massief [Zoe72]. Massif [Zoe72]. mat [BTP24, Har04, Kim99, MNK04]. match [TKH20]. mate [GBA20]. material [Bla82, KSY19, KM17, LME10, LKE94, PS06, Wol79]. Maternal [AL09, Zad03]. Mathematical [AC86, BGK02, BS95, DBG02b, PE78, PB04]. mats [RHA00]. Matter [Alt98, Bou07, BHM03, BKW94, Bro94, Cad78, CV95, CCS09, CLL10, CZZ^{+}25, DSN09, Dan84, ECL94, GCS93, GCL99, GM16, HK82, KSBR94, KSBŁ95, LLF99, MGL22, OBC82, PPL16, RFO94, ST09, SCG09, TLC93, VdJ94, Vos87, WLN12, WS94, WGK95]. matters [DLG20]. maturity [HdC24]. Mauremys [DDZ^{+}24, RBAB24]. Maurer [FvM03]. maxima [GKK95]. maximum [DW93, KSF95, KD82, Mar93]. May [Gel70, CGD19, KMR20, MMBP21, TBS18]. mayflies [YL08]. Mayfly [BKO92, PBF06, VIM17, DRM18, HÅB08, IGS10, OB13]. Mayo [WMV09]. Mazatlán [TRHHAR10]. McLeod [IJ06]. MCNN [FWS^{+}23]. mcyA [PMZMJ16]. meadow [SB14]. meaning [Eng74]. means [BMO15, Hig74, VP77]. measure [AAPdG10, KWO11, SEP09]. measured [Ber84, Bir78]. measurement [BGK85, CGZ83]. Measurements [Don79, vG84, BRG02, BTA94, Bir78, FBR87, GR97, HvdH84, RDG86, VPF21, Vri80, WDP15, WVM18]. measures [LWdSD16, SK16, Van82a]. Measuring [TRP22, TDP06, YB08]. mechanical [Van82b, VPF21]. mechanism [MNS05a, MNS05b, WL24]. mechanismen [Spa71]. Mechanisms [RRL09, SAL16, DBG02b, DG02, GA10, Spa71]. mechanistic [VN22]. Meded [Har73a]. Mededeling [Ano72d]. Mededelingen [Ano68c, Ano69a]. media [CNN^{+}24, MJEC05, NNJ22, SSN24]. mediated [AFR14, DdNM22, EMF14, KSY19, MSB15, PMA22, SWB13, SSB21, SPA12, WP12]. mediating [BA14]. mediators [ODR22]. medicinal [UT16, ŽRP16]. Mediterranean [Bal09, GCS93, LAC99, LZH^{+}25, AYM09, CCC19, CC94, CVM21, FGA^{+}24, GUT21, GRDPP09, HEZ03, HPPL00, KK09, LR92, LPB07, MKT23, MAA22, MTC^{+}24, MOdSCP07, MOPCP08, MOCPG08, NRB21, OR11, PGA^{+}22, SGRMP20, SCEGL12, SEE22, VSM19a, VALRdF20, ŽAL24]. Mediterranean-type [SGRMP20]. Medium [MKV15, BMO15, HH08]. Medium- [MKV15]. medium-sized [BMO15, HH08]. medroxyprogesterone [MRBQ^{+}25]. medusa [SC00]. Medusae [MAA22]. Meer [Bak72, Kap76, Pee71, Dum69]. Meeting [Ano79]. meetings [Ano76b]. meets [Gro22]. meeuwenven [GB69]. mei [Gel70]. meio [DB81]. meio- [DB81]. meiobenthic [WT14]. meiobenthos [Kur02]. Meiofauna [BHPT20, NHF96, DRN09, GCT^{+}23, HWG77, HHA92, MT09, PT14, PGT20a, RB08]. meiofaunal [PGT20a, RMS11]. Mej. [Gee76]. melanostomus [BF12]. Melita [FvM03]. members [Den74, Ano70b, Ano72c]. Memoriam [Eng69, Gul05, Dor80a, Gee76, NB80, Ros71, Sch79, Sch85, SvdG92]. Menidia [TGV00]. Mensurative [KW12]. Menyanthaceae [BHV82, VVB82]. Menyanthes [Har04]. Merapi [SDW^{+}24]. Mercenaria [Mit74]. mercure [Bil73]. Mercury [LCF91, Bil73, LEB93, PDM95, TK95]. Mere [BM03]. Merja [KHF01]. meromictic [KBR10, KRZ03, RZG09, RGD10, WHP76, WWŻ06, ZTG10]. meroplancton [Dar73]. meroplankton [Dar73]. Merrier [Lin07]. mesh [PGT20a]. Mesocosm [CNA07, DdNM22, ERA17, SMPK09, TŠB17, ZJN^{+}23]. mesocosms [BM98, SS99]. Mesocyclops [RK02]. mesograzer [SRG18]. Mesohabitat [GBT19, WHW14]. mesohabitat-specific [WHW14]. Mesopodopsis [MJ93]. mesopredator [FGB20]. mesopredators [PSW11]. mesotidal [RFO94]. mesotrophic [Fli85, PBG03, TLF19]. Mesozooplankton [FR09, SDD^{+}23]. message [DT92]. meta [WWS21]. meta-analysis [WWS21]. metabarcoding [FSdO20, GPGSMM^{+}23a, GPGSMM^{+}23b]. metabolic [MAM13, SVB97]. metabolism [CFP08, De 80c, DBdF02, EAK11, Gul76b, RBN20, SNKN^{+}24]. metabolites [Har82a]. metacercariae [RVB08]. metacommunities [Alt13, PGT20b]. Metacommunity [PVH17, RFGLSB20, PBM21, RdMJN22, Sir17, TCH18, TH16]. metagenomics [CBdO^{+}24]. metal [CBN94, De 84a, GPA98, Gul98, JCL09, dSLML21, SGM22, WWG^{+}25, WGK95, ZV93]. metalen [Fon71, dG71, dW71b]. metalimnetic [QJD02]. metallica [IS06]. metals [DV92, Fon71, GM93, JC80, MSK93, PPK^{+}23, Sur01, TG04, TVD91, WFH93, WKH82, dG71, dW71b]. metamorphosis [ZYS14]. Metapenaeus [TYN02]. metapopulation [OAS13]. metazooplankton [KK20]. Meteorological [MGC19]. meteorology [Møh99]. methabenzthiazuron [WRH98]. methane [BVM15]. methane-oxidizing [BVM15]. method [Cra91, DB81, HS85, HS86, JBB07, Kle84, LPT22, LB14, LKM01, SGM99, Sch78, TV85, TZ05, VVB82]. methodical [Mar01]. methodological [BHV82, KPV16]. methodology [STT01]. methods [AMTSJ15, BPT85, BMJ16, BKC11, Bra01, CKK99, DB85, GR97, Gul74, Gul85, HCS02, HM86, Roi81a, Ros92b, vIR92]. methyl [GFSN04]. Methylmercury [RVB08]. metolachlor [HAB98]. metrics [ODI08]. Meuse [SVK94, SVW95, VSA97, BKO92, DvZA90, HvAV83, HL95, HPT95, KF95, Kra95, MLB16, PW73, PP95, TDBW95, VVK96, VVW95, Wol74, dHC76]. Mexican [HSCR10, NGSOAD20]. mexicana [FTP08]. Mexico [ABDPPP^{+}24, CAPGA08, CFP08, CFRNT21, DPL03, GPGSMM^{+}23a, GPGSMM^{+}23b, LGCS09, MB14, NEGS07, RRRRA07, RJD07, SAACR09, TRHHAR10, UOJGRL21]. MFLD [SJJA23]. MFLD-net [SJJA23]. Mg [Lar92, PR92]. Michael [Gul99]. Michigan [AC86, MFS03, SCO09]. Micrasterias [NW24]. Micro [FLP08, Ker83, KS94, DGN21, DB81, De 79c, DM71, Møh99, PDvdV06a, PDvdV06b, RM81, Rin76]. micro- [DB81]. micro-algae [PDvdV06a, PDvdV06b, RM81]. Micro-benthic [FLP08]. micro-ecosystem [Rin76]. Micro-ecosystems [Ker83, De 79c]. micro-eukaryotes [DGN21]. Micro-organism [KS94]. micro-organismen [DM71]. micro-organisms [DM71]. micro-tidal [Møh99]. microalga [GAAT^{+}24, KZV02, NW24]. Microalgae [EAB08, HZZ22, LZZ^{+}25, RBMCP21, RHO09, RRL09, SSR13, TPT22]. microalgal [IUK22, PVV07]. microbe [EMF14]. microbe-mediated [EMF14]. microbenthos [De 77a]. microbes [GBJM22, JTGS24]. Microbial [KRZ03, LGCS09, LLL^{+}23, UKS05, YS98, AA95, AH82, ASA06, BGRM^{+}24, BLS88, BG15, CBdO^{+}24, CCK10, DSN09, Deg10, DCC20, FO98, Har82a, KRJ09, KKD02, dSLM94, MSF21, MDV03, iNTT05a, iNTT05b, NZ98, OLTV19, Wet93, YYS^{+}23, ZGZ20, ZJN^{+}23]. microbially [CZZ^{+}25]. Microbiological [Laa97]. microbiology [Ano83b]. microcalorimetry [MNK04]. Microchironomus [Kre79]. Microcommunities [Wet93]. microcontaminants [RHO99]. microcosm [EMF14, HBB13, TIF10]. microcosms [CH12, HAB98, TJB98, WRH98]. microcrustacea [BM03]. microcrustacean [GFF17]. Microcystin [PMZMJ16, LSY20, MdMdLB21, PPH16]. microcystin-LR [LSY20]. Microcystins [CDL13]. Microcystis [BHB01, CDL13, DCL19, DZT13, dSFFA03, HLL14, JJP24, KBS05, LZN14, MF07, MHG16, NR97, Nan00, PMZMJ16, PPH16, STH21, SBH17, TDL23, VvGV12, WZL19, YP17, YSYZ23, YZY09, ZWY20, vGVV11]. Microdistribution [Tol82, Ola92]. microecosystem [Ker97]. MicroExtraction [RHA00]. Microfauna [GD72, LMD11, GvdH68, GD72]. microflora [TTD02]. microgradient [Mer79]. microgradients [Wet93]. micrograzers [CLL10]. Microhabitat [RJD07, Gar05, IS06, PAAJM17]. microhabitats [BHPT20, CBP22]. microorganisms [Blo83, HWS22, RMC95]. microphytes [Dre68]. microphytobenthic [BTP24]. Microphytobenthos [BNV03, BC95, ABA^{+}24, SC99]. microphytoplankton [Bal09]. Microplankton [LAC99]. Microplastics [RBMCP21, HZZ22, HLSN24, LYM21, MRBQ^{+}25, ONA^{+}23]. Micropogonias [MCV08]. micropollutant [VVW95]. microsatellite [LXH21]. Microscale [BNMH12, NČŠ12]. microscopical [BPT85, VKK73]. microstructure [PFLO14]. microsystems [Ker75]. Microtox(R) [QPP95]. middens [dBPG00]. middle [VD94, ARP14]. midge [BW92, Pen00, SB88, WFFM10]. midges [MSP13, WAW92]. midget [DDL19]. Mignone [BB87]. migrating [BFS21]. Migration [ACL13, BNPC07, CCH11, DGD06, MS02, DAO24, RFR91, SZH21, VTE19, WFFM10, XCL10]. migrations [Dar73, OSM13, Cad04]. migratory [Cru10, RR98]. Mikrofauna [Dav71, GvdH68, Dav71]. Mikrofauna-Onderzoek [Dav71]. Mikrofyten [Dre68]. Milieu [Bak72, dP71, Bil73, Blo70, Dor70b]. milieufactoren [Web71b]. mill [SN03]. Millet [VV85]. Mine [DG06, Gos99, SGA06, ŻD06, BDH09, dSSSES21, SGGO06, Żur06a, Żur06b]. mined [VPS21]. mineral [DE93, FAL07, SMPK09, TW03]. Mineralization [Van77b, OBC82, RMC95, Sep79, SB02, VV93a]. Minho [FCDA09, VBA22]. minimal [KM17, Sch89]. minimally [AANO25]. Minimum [SPB97]. Mining [Gos99, DEP22]. minor [MRBQ^{+}25, VC89, VKD77]. minute [VKK73]. Mira [CM95b]. mire [HHN03]. mires [RKS14]. Mirolli [HVS87, Van91]. Miscellanea [Ano73a, Ano73b, Ano74a, Ano75a, vBHvdVH75]. Miscellaneous [Ano78c, Ano78d, Ano81c, Ano84c, Ano84d, Ano85, Ano88a, Ano88b, Ano89b, TD87b]. mission [Cra87]. mission-orientation [Cra87]. Missouri [DBG13]. mites [DS75, DHW81, DTD94, MSW10, VS96b]. Mitigating [HSP16]. mitigation [IFV16a, IBV16, IFV16b, MJM16]. Mitilidae [PCB10]. mitochondrial [OHA08]. mixed [AMSN07, MJ93, NBM98, PB04]. mixing [KS94, MOCPG09, RGD10, VS09, VIH16]. mixotrophic [BHB01, VCT13]. mixture [HAB98]. mixtures [Her83, LA19]. Mizorogaike [Har04]. Mjóifjördur [Mun94]. ML [Ber84]. Mn [CMG95, WA84]. mobile [FKB92, HHV93, dSCT23]. Mochokidae [KKO08]. mode [PGT20b]. Model [Los80, vDG94, BV14, BS95, Bui95, CFRNT21, DB85, De 80b, DGD06, DBG02b, GBM10, Gos98a, HBO03, HEZ03, JJP24, MD92, Nie82, PLM^{+}22, PE78, PB04, PMD10, Sch89, SS99, SMC08, TJB98, VK80, VPS13, Wor90, vLJA07, vdTS97, vdV76]. Modeling [SWW02, ZLC12, BBM10, DGBL15, NRS19]. Modelling [BLS88, BHP05, HB03, HG05, JA90, JVG95, KN86, NAM19, Pos80, TYN02, VOD93, ZSG14, AND25, AC86, CN94, Coa07, DYS86, FDB97, JBB07, KA86, MTJ10, PN94, SPB97]. Models [DDD07, BDR97, BGK02, GCN^{+}23, GR97, JAM15, MAS14, MKV15, Ver76b, ZCZ14]. Moderate [ACB12, SXY20]. modern [FDG97]. modernization [KGM13, KGM22]. modestus [DDPT24, ET16]. Modification [WKT05]. modified [AGB94, PDP02, Rao10, YKJ15]. modifies [WP12]. modify [WV02]. modulated [Rat14]. Modulation [PMP94, PPH16]. Mogelijk [Gie71]. Mogelijkheden [Pee71, vdW70]. Moina [AMSN07, GFSN07, JTGS24, LAZ21, LOZ24, MJEC05, Zad03, ZL07]. Moinidae [MJEC05]. moisture [CBP22]. Molecular [VSM19b, ZTNW21, GMVCLGGP24, GC23, HBR12, OMNC24, SNO98, ŽRP15]. Mollendo [FCPCMB^{+}25]. Mollusca [Ban75, Ban75, DG84, MSDDD22]. molluscan [CW04, dSDMD20, PEREBM^{+}23]. molluscivorous [KK11]. Molluscs [MK05a, TFT10]. mollusk [FNS^{+}24a, FNS^{+}24b]. molly [FTP08]. Mona [SCO09]. Mondsee [GDT08]. mongolianum [NMS21]. Monitoring [DF90, HCS02, BSMVP24, BMO15, BRP15, CNN^{+}24, Coe77, COC18, dSDLdA22, DTC19, ECB09, JCS01, Mar01, MDP08, SMR08, SD86, SK04, TKD01, VSM19b, VSW00]. monitorization [CHF09]. monogonont [RM98]. monooxygenase [HPT95]. Monoporeia [Leh04]. monsoon [AAPdG10, FR09]. monsoonal [SCC19]. monsterpunten [Ano68b, Ano68d]. montane [IJ06, PPCH22]. month [ACB12, SSV21, ZJN^{+}23]. Monthly [Kal86]. Moog [Ano05b]. moor [Van89]. moorland [Coe78, De 84b, Hig79, Van96, vDB78, vDSB80].

Moreau [Gul09b]. Moroccan [MAA22]. Morocco [KHF01]. morph [Gil12]. morpho [AVdNM19]. morpho-functional [AVdNM19]. morphodynamics [BGL13]. Morphofunctional [FBL13]. Morphological [SAF10, BA14, ČN10, DDY^{+}23, FAL07, Gie83, HWK11, HSY18, KvE21, LLX11, PU15, VvGV12, YYR20]. morphologically [PCD11]. morphologies [GJW20]. morphology [BdSM20, CBNF19, GJW20, KSW92, LZH^{+}25, LB20, MNS05a, MNS05b, MFS95, MMBP21, PR16, PMVdC^{+}23, QLW^{+}23, Roi88a, SK09, SY02]. morphology-main [PMVdC^{+}23]. Morphometric [KK09, RJD07]. morphometry [RDG97, SJJA23, VHN00]. morphospecies [YSYZ23]. morphs [KJS15]. mortality [ABDPPP^{+}24, BJS12, Had79, ISY10a, ISY10b, nWdLjZ21]. Mosaic [DN23]. mosquito [EMF14, MMBV24, PD10, RHA00, RHO09, SAH10]. mosselen [dW71b]. mosses [LB10]. most [IRT06]. mother [MLL21]. motility [OFCP22]. motion [MHG16]. Moulting [EG04]. Mount [SDW^{+}24]. mountain [Bra21, CGL21, Mou92, SB02, VALRdF20]. Mountains [SR92, WB10]. mouth [WMTR07]. movement [GC05a, GC05b, Pen00, TS17, YKJ12, dSSR^{+}23]. movements [BMO15, HHP12, MGC19, RMH21]. moving [PV95]. Mozambique [LMM05]. mucus [NYT09b, OR06]. mud [ACH09, HBR12, LDK20]. mudflat [CMH14, NLC06]. mudflats [TC15, WBB93]. mudskipper [EBA96]. mudsnail [HKT16]. mudsnails [BT23]. Müggelsee [KEH09]. Mugilidae [KHF01]. Müller [BJ99, HVS87, HLSN24, NNB^{+}24, SAN99, YZ08]. mullet [CCC19, KHF01]. Multi [RMH21, BWS21, FWS^{+}23, GCN^{+}23, PCGW20]. multi-beam [FWS^{+}23]. multi-species [GCN^{+}23]. multi-stressor [PCGW20]. multi-wavelength [BWS21]. Multi-year [RMH21]. multicellular [HK76]. multichannel [HvdH84]. multidecadal [VvdMP12]. multigenerational [OCS20]. multilocus [RvGB05a, RvGB05b]. Multiple [CHP06, LBKV18, BSW^{+}24, BHG16, CDD20, GCN^{+}23, HSY18, LZY21, LMP02, MAS14, PF07, TC15, TDHH23]. multiple-source [TC15]. multiplier [dPBCP22]. Multiscale [SGRMP20]. multispecies [NBD^{+}25]. multitrophic [SSB21]. Multivariate [VB98, HH08]. Mundeswari [PMB^{+}24]. muratensis [SHK92]. Murphy [Ver97]. musica [Ste94]. Mussel [MGCC93, BKD92, BDG10, FD94, ISJ10, KK11, MFP01, NBD^{+}25, RVK96, RVJ98, Rei99, WFBJ17]. mussels [BRA^{+}23, BC23, KJP14, Kra95, MLB16, NBD^{+}25, OTS05, OMM05a, OMM05b, SS99, VD85, dW71b]. Mutual [KKM09a]. mutualism [Wet93]. Muymuy [FCPCMB^{+}25]. Myotis [RHO99]. Myriophyllum [CvDG07, CM95b, JJP24, JGH97, LZD16, SRG18]. Mysid [Lin07, Alb04, HGR11, nWdLjZ21]. Mysidacea [Alb04, MFS95, MJ93]. mysids [BKW94, PC14]. Mysis [NAT08, SCS12]. mystery [OLM09]. Mytilacea [MGCC93]. Mytilopsis [Ken11]. Mytilus [MGCC93, PSP91, Rei99, SS99]. Myvatn [EÖ04, GEÓ04b, ÖE04, TE04, TGE04, TG04, BJ04, ESG04, Ein04, GEÓ04a, GE04, Gud04, KHM04, KÁ04].

N [BHP05, VDB81, BBM09, BZV93, BPD15, GCL99, MF07, PHB09, PMZMJ16, PES99, TLF19, TKR10, Zev82]. N- [PES99]. N.W [Gar71]. N.W. [Pol76]. N.W.-Overijssel [Gar71]. N/P [MF07]. Naamlijst [Ano68d]. naar [Dre68, DM71, Ker70]. NaCl [SNGV06]. Nagabhushanam [Hod98]. Naididae [Mas88]. Nakamura [vdV99]. namaycush [BTB09]. Namil [AAMS24]. nana [NSK08]. nannoplankton [Mom73]. nanoflagellates [LKM01]. nanophytoplankton [VET20]. nanoplankton [CLL10]. narrowest [SDD^{+}23]. nascent [KHG20]. Nassariidae [CM05]. Nassarius [CM05]. natalis [Pad13]. natans [MLL21, TZD21]. Natantia [VV85]. naticoides [VM25]. National [AP00, CAJ11, ES09, FM99, HP06, KKO08, BST89]. Nations [Nie98]. Native [SGD22, SZH20, Ber24, BLP16, BKG16, CYW24, CF09, DD17, DBG13, FO98, GBPR23, GKM22, HKT16, HLK22, HSY18, KKM09a, KKM09b, KJP14, KHK17, LDK20, LA19, OLM09, vOCRM19, PGM^{+}24, PK08, PDP02, SLS15, SDRM16, SAR21, TET23, UET12, WMM11]. natives [SJD22]. Natural [Cad08, CHV91, RR03, Vel91, AL09, BBM09, BWM86, CAJ11, Den06, Don79, dSFFA03, FAL07, Fli85, GTI99, KSBŁ95, LHV07, MA95, MKL10, PPK^{+}23, PLC22, Rao10, RPSSDS23, SLM17, SZH20, VPS21]. natural-like [AL09]. naturally [FCG16]. Nature [Mit78, DPT04, Dvô70, LH93, Lee80, NMS21, PPY^{+}22, TSV16, TNtK03, vDG94, PHB09]. nature-based [TSV16]. navigation [Ano25, CvNB02]. Neagh [Gib86]. Near [PJH06, Alb04, Das07, De 76b, KHG20, Kuk92, MGL22, Mol76, Slu81, Van91, WB10, vDB78, PPK^{+}23]. Near-bottom [PJH06, Alb04]. near-pristine [WB10]. near-shore [Kuk92]. nearctic [Jac92]. nearshore [Den06, GBZ21, PFLM15, SCG09]. necked [SSV21]. NECOV [Ano99a, Ano00d, Ano03b, Ano04d]. Nederland [Ano81a, Ano72b, Bak72, Ver72, Vis88, lCdB70, Ano81a, Har73a]. Nederlands [Ano04d]. Nederlandse [Par69, Sto76, dW71b]. need [Bro98, YB08]. Negative [FGA^{+}24, PGA^{+}22, CDL13, HSCR10, TDHH23, VGG89]. negatively [CLRdSR23]. neglected [RAA98]. neighbours [ST20]. nekton [RA07]. Nelson [GWF92]. nematode [BvdH90, Bro84, EK98, TFH^{+}24, WMTR07]. nematodes [ER94, LV93, PFT12, PGT20a, RB83]. Nemi [Mas85]. Neogobius [BF12]. Neolissochilus [DT23a]. Neomysis [Lin07, MFS95, nWdLjZ21]. neonate [RW98]. Neonates [VV07]. neophyte [LvBR06]. Neotropical [QCF15, AOLLJ24, CLRdSR23, MCC21, dSGD17, CDW92]. neozoon [Zet97b]. Nereina [VTE19]. Nereis [FSC06, Rii94]. Neritidae [VTE19]. nest [GB69]. nested [CCM21]. nesting [GS11]. net [SV92, SJJA23]. nete [VDB81]. Netherlands [Bel79, BZD81, Bij75, DD84, De 76b, GB82, HWG77, HvB85, MRD89, Pol75, PN94, SOD92, Str88, TFR91, Van91, Ver80a, VD94, WP91, dR92, vG84, Ano72b, Ano99a, Ano99c, Ano99d, Ano00d, Ano03b, Bak72, BD74, Bak77, BV78, Bak78, Bak80, BTV88, BV89, Bel76a, BZV93, BI09, BP92, CN94, Cla84, CH83, Cra83, Cup94, Dan84, DWV81, De 84b, DV92, ER94, FvM03, Fli85, FBR87, FDG97, Gee75, GV89, GD91, Haa73, HvAV83, HHV93, Hig80, HM84, HV89, Hos80, HM86, Hos89, Hou94, Kap82, Kla80, Kre79, Kre82, KKV92b, Lee84, LL77, MHL05, Nie79b, NDW96, NHF96, Par80, Pee74, PW73, Pin75, Ric86, REH92, Roi81b, RJ96, Roi88b, SB78, Sep79, Slu81, Ste83, SR88, TVD91, Tol84, TG88, Vaa75, VW78, Vaa79]. Netherlands [Van79a, Van84a, VMS94, VV85, VSA97, VS96b, Van75a, VH88, VV91, VV93b, VD87, Ver72, VGM93, Vis88, Wol97, Wor90, ZV93, lCdB70, vIR92, vdTS97]. netplankton [Mom73]. nets [FGA^{+}24, PGA^{+}22]. network [Alt13, BLB^{+}23, BCN12, DdNM22, GCN^{+}23, JBB07, JJP24, SJJA23, SRF21, RRS^{+}22, RRS^{+}23]. networks [DT23b, GGD07, GDD07, GRDPP09, KA16, RFGLSB20, SXC^{+}25, SOW01]. neumanni [MMBV24]. Neural [RRS^{+}22, RRS^{+}23, SOW01, DT23b, GGD07, GCN^{+}23, GDD07, JBB07, JJP24]. Neuse [PDP02]. Neusiedler [DH09]. Neutral [De 97]. neutralizing [HFS08]. newly [GD91, OLM09, ZPW13, ZPW15]. newly-created [GD91]. newts [ZCA20]. Niche [BHG18b, NBD^{+}25, VvdMP12, GSH17, KJS15, KHK17, NRS19, SAACR09, SKH19, SBLdAM23, dSGD17, BHG18a]. niches [BBH05, FAB18, ZYL22]. Nichupte [LGCS09]. Nichupte-Bojorquez [LGCS09]. niei [XCL10]. Niels [Gul07a]. niet [Gar71]. niet-insekten [Gar71]. Nieuw [Dor72]. Niger [NE85, EAO20, Nwa84, Gie83, GKM17]. Nigeria [AKF^{+}23, AAA^{+}24, EAO20, EAB08, EBA96, Nwa84, NO85, Nwa87, ON88, OSO09]. Nigerian [AAA^{+}24]. nigroculus [DRM18]. nigromaculata [LWW^{+}25]. Nijmegen [Van91]. Nikanorov [Gul98]. Nile [AIS02]. niloticus [AIS02, MOM24, TBA03]. nine [ABF01, BPA01, BBR01, FDP01, PBR01, PJF01, RFP01, RFF01]. Niño [TK19a]. Niño-induced [TK19a]. Niphargus [MSB10]. Nippoleucon [GCT^{+}23]. Nitellopsis [HLK22]. nitida [FvM03]. nitrate [CVG03, LW11, PMSS18, STV06, YBZ20]. Nitrification [STG18, SB14]. Nitrite [GK07, SXC^{+}25]. nitrite-induced [SXC^{+}25]. Nitrogen [AH82, Bro81, APS81, CCS99, CBP22, CVG03, CAFA04, FVW03, HSP16, HWS22, KTJ10, KT09, LWY^{+}24, LZZ^{+}25, MF07, NN05a, NN05b, NC20, PGSL02, PFT12, Pod74a, Pod74b, Pod74c, PPH16, QLW^{+}23, SCN^{+}24, SK84, SJD22, SBH17, TGV00, Van89, ZNP13]. nitrogen-phosphorus [CAFA04]. Nitzschia [LKSK15]. nm [DDV91]. No [BI09, CB13, SXY20, DF11, ZD21]. noise [BDP21a]. Nolan [War97]. nomenclature [Zet97a]. Non [KAY20, VP12, BDH09, Ber24, Bes79, DWB10, Gar71, GAA24, GEG91, GKM17, GBA20, HLK22, KLG17, KKM09b, KHG20, KHK17, LA19, MSP13, PMA22, PGM^{+}24, PRW08, SGD22, SBH17, WMM11]. non-biting [MSP13]. non-breeding [KLG17]. non-consumptive [PMA22, SGD22]. non-expansive [GKM17]. non-fluorescing [GEG91]. non-indigenous [BDH09, GAA24, KHG20]. non-insects [Gar71]. Non-lethal [VP12]. Non-marine [KAY20]. non-native [Ber24, HLK22, KKM09b, KHK17, LA19, PGM^{+}24, SGD22, WMM11]. non-sterilised [GBA20]. non-structural [Bes79]. non-toxic [SBH17]. non-urban [DWB10, PRW08]. noncoding [OHA08]. nonlinear [Kop21]. nonnative [ET16, Kae20]. Nonnetje [dW71a]. nonpoint [SON11]. Noordoost [Gys72b]. Noordoost-Vlaamse [Gys72b]. Noordzee [Gie71, NZ75]. Noordzee-Kabeljauw [NZ75]. Nordsee [Ban75]. Norman [Hol70, Baa88b, Hol70]. normanii [RSK93]. North [Ban75, BB23, FD94, FN94, FG91, PDP02, Sch95, AAMS24, GS04, HHN03, NBT21, PFSPdS24, SAL16, YO02, YKU01, ZXH15, Alt98, ABF01, AR09, BZ97, BPA01, BBR01, DT92, DBV96, DJD92, EK93, FvM03, FAK01, FLP08, Flo01, FDP01, Fra80, FNW94, Gie71, HF97, HHA92, Hey92, KHF01, Mom73, NZ75, NMB09, Pad13, PBR01, PJF01, RFP01, REP01, RFF01, RJ11, Rei92, TB92, VD94, WMM94]. North-American [FvM03]. north-east [GS04, FLP08]. north-eastern [HHN03, PFSPdS24, SAL16]. north-western [AAMS24]. North1 [AP12]. NorthCentral [CFP08]. Northeast [BPP21, VSM19a, Gys72b]. Northeastern [KND^{+}22, CEH12, FGL19, NCDS^{+}25]. Northern [DDD07, DN23, Gri89, MM95, Rav01, RG07, WS94, vGVV11, BD09, DT23a, DT23b, FBNF08, GIK21, HJV08, KPJP23, KÁ04, KHK17, MOM24, Pad13, PAAJM17, SS24, TWC13, TRP22, VHB09, BP21, MRR10, NAM19, PLC08, SB08, Van79a]. northernmost [Sör82]. Northwest [HBR12]. Northwestern [Zdo09, CPM09]. Norway [AK09, Pri03, SMR08]. Norwegian [HFS08]. Note [Kre82, BST89, Mas88, VKD77]. Notes [MJEC05, VH88, DK80, FNW94, ON88, Pol75, ŽG09]. Nothobranchius [MMBV24]. Notopterus [PPK^{+}23]. novel [DH22]. nuclear [JCS01]. nudiventris [SHK92]. numbers [GMD19, KD82]. Numerical [BBM10, PN94, Rei92, Ano68d, CN94, TNC08]. numerieke [Ano68d]. numidicus [CB00b]. Nuphar [BH82, ŻKŻ10]. nursery [CC99, KKS11, Mar93]. Nutrient [BBM09, Bes76, BZV93, BLG99, Cha89, ODI08, SFB04, TBF99, WiNDF22, WDP15, dPNF91, AP12, ACE10, BB12, BBF17, BHL09, BNMH12, CEH12, Cha87, CP10, Clo99, DGB17, GGT14, GCL99, HMH05, IJ06, JVG95, KFS04, KK17, KGB04, Kop21, KBS05, KP99, LAM06, LLG10, MA95, MJJ21, Møh99, MSD08, iNTT05a, iNTT05b, NvZvG16, PPL16, PHB09, PPC08, PDP02, PMZMJ16, Rat14, RPS20, REH92, Rod10, RNJB20, SMPK09, SS99, SAB10, SB04, SOD92, Som83, SOU14, ŠV09, TŠB17, TBS18, Van97, VBD96, Wet93, WMD07, YHM05, YYR20, ZJN^{+}23, dSSB24, dR92, vLJA07, vdTS97]. Nutrient-induced [dPNF91]. nutrient-rich [KK17]. Nutrients [VD87, BGM08, DGG15, GXY14, GCE11, KFS04, KRJ09, KHD99, MTA15, PLdFC22, RDE93, SZY19, TG04, VK12, WFH93, WBB93, Cad05d]. nutrition [Har04, SS98]. nutritional [DL89]. NVAE [Ano97b, Ano97c, Ano97d, Ano98e]. NVE [Ano99c, Ano99d]. NW [Bra99, CM95b, FBV02, MB06, FCDA09, SB02]. nycthemeral [Dar73]. nycthemerales [Dar73]. Nymphaea [PB84]. Nymphaeaceae [PB84]. nymphaeid [Bro81, vdV81]. nymphaeid-dominated [Bro81, vdV81]. Nymphoides [BHV82, BV96, MSD08, VVB82].

O [Ano05b, EAK11, Raa89]. O. [ARG20, BHV82, MSD08, VVB82, YZ08, BJ99]. objectives [zEGB13]. obliquus [DDY^{+}23, MMV05, WBB00]. observation [SS24]. Observations [Hig76, MO92, Sch72b, VV85, BV84, Dor76a, Dor87, Gie83, Gul76a, KHF01, NNJ22, RSK93, Rin97, WP91, CGD19]. observed [MMM11]. obtusa [NJ00, BHP05, HLK22]. obtusal [Pis69]. obtusale [Pis69]. occupancy [SLS15, SPS22]. Occurrence [CEH12, DTC19, HL95, LA99, PAP24, RPN15, RCV91, Ber75, Bij75, CB00c, Cru10, DR96, Dre68, HVS87, HKH^{+}23, Hou94, HHY21, Kat92, KÇA22, MGT17, MOdSCP07, PFSPdS24, SC00, VV85, ZSG14, dW71b]. occurrences [DT23b]. occurring [GSH17, GDB19, PMA22, SKH19]. Ocean [Den06, Hod98, RMS11, SL24, TK19b, TK19c, FRP08]. oceanic [LLL22]. Ochiishi [HHN03]. Ochromonas [BHB01]. octocoral [dPBCP22]. octopus [DTH20, DDL19, ARG20, DTH20, dSDLdA22, DDL19, LLL22, UOJGRL21]. octopuses [SRI20]. Ocypode [NNJ22]. Ocypodidae [LMM05]. Odessia [PFSPdS24]. Odonata [HP06, PMVdC^{+}23]. odonate [Jar08, dRFC21]. Odontobutidae [ŠDT^{+}24a, ŠDT^{+}24b]. Oecologische [NZ75, Per69, Spa71]. off [FBNF08, GBZ21, KK20, KBN20, PFLM15, SAACR09, TRHHAR10]. offered [NR97]. offs [INJ20, TBBS17]. offsetting [TRP22]. Oguta [ON88, NO85, Nwa87]. Ohio [KGB04, Sæt92]. oil [DPL03, SDD^{+}23]. Oisterwijk [Coe78, vDB78]. Okaichi [Cad05e]. Okanagan [SCS12]. Okeechobee [Rod10]. Oklahoma [Ber24]. Old [Bak72, ACB12]. older [KNH07]. Oléron [BDR97, ZSD94]. olfactory [Mar17, RMH22]. oligo [PBG03]. oligo-mesotrophic [PBG03]. Oligochaeta [KMK09, Mas88]. oligochaetes [KKV92a, RP92, Ver80a]. Oligotricha [MM96]. oligotrichous [CLL10]. oligotrophic [BN24, CS13, FRP08, Rat17, VIM17, dCAP23]. Olivier [BKO92]. Ologe [AOS^{+}24]. Olympia [zEGB13]. Om [Gys72a]. Oman [AAPdG10]. omgeving [Hog69]. omnivorous [dNFdNM21, dCdCP19]. omschrijving [vdB70]. Oncorhynchus [KAN20, VK12]. onder [Ess71]. ondergedompelde [Per69]. Onderzoek [Dav71, De 77b, Mol72, lCdB70, Dre68, DM71, Fon71, Ker70, vdB70]. One [KF95, GBM10, GB07, KKM09a, PMD10, SSV21]. one-dimensional [GBM10, PMD10]. onoverdekte [lCdB70]. Ontario [EPL^{+}23]. onto [MD92]. Ontogenetic [Gar05, ABF21, SRF21]. Ontogenic [DTH20]. ontogeny [OCS20, VEK20, VALRdF20]. Onychopoda [KG94, NAT08]. Oomycetes [WP91]. oostelijke [Ess71, Bel76a, Gul76a]. Oostende [Per69]. Oosterschelde [Ban75, BTV88, Ban75, TFR91, VD87, WP91, vG84, vdTS97]. Ooy [Oom72]. Ooypolder [Ano72i, ABV72, GD72, Mol72, Sch72b]. Op [Gys72b, vdB70, Fli70, Fon71, Gel70, Mur69, NZ75, Per69, Sch68a, Web71b, dW71a]. Open [REP01, BV78, CSV19, KD82, LB10, OBC82, TBS18]. open-canopy [LB10]. open-water [KD82, OBC82]. Opencast [DG06, SGA06, ŻD06, SGGO06, Żur06a, Żur06b]. Opening [Gin80]. operation [YKJ12]. opilio [YSY^{+}22]. Opisthobranchia [WF05a, WF05b]. opmerkingen [dW71b]. oppervlaktewater [Fon71]. Opportunistic [SAH10, dCdCP19]. opportunities [MTJ10, SJ00a]. opposing [KA16]. opposite [RTW17]. Oproep [Ano81a]. Opslag [Zoe72]. Optical [GEK97, AR09, BRG02, BP94]. optima [ODI08]. Optimising [WV02]. optimization [ADS25]. optimum [Rin70b]. options [BH16, IBV16, dJG07, Pee71]. oranges [YB08]. Orconectes [CBNF19, Gee75]. order [Ano68d, BHPT20]. ordered [PV95]. Ordination [Aag92, Ros92b]. ore [PdCF^{+}25]. Oregon [KHG20]. Oreochromis [AIS02, MOM24, TBA03]. Organic [Bou07, Bro94, TLC93, AP12, BT01, Cad78, Cad80, CH91b, CV95, CCS09, CZZ^{+}25, DSN09, Dan84, DDV91, De 78, DF90, ER94, EK98, Ess03, FGB20, GCS93, GCL99, GC23, GM16, HK82, HFS08, HH80, KTJ10, LLF99, LME10, LZN14, MGL22, Mas85, MV83, NN05a, NN05b, Nie79b, OBC82, PPL16, PFDM99, RKS14, SMPK09, SKW21, ST09, SCG09, Sch78, Sep79, VVW95, Vos87, WLN12, Wol79]. organisation [Ker83, TCH18]. organism [KS94, TKH20]. organismen [DM71]. organisms [BVI16, DV92, DM71, ECB09, Hod98, LZY21, NRS19, PGT20a, RMR21, Rin97, Tol84, VRH^{+}24, Gos98a, Cad04]. organization [Hig09, RFGLSB20, TH16]. organizations [vdB70]. organized [Gel70]. organochlorine [PJF01]. Organotin [DAR99]. organs [Cha89, LAM06]. orientatie [Fli70]. orientation [Cra87, Fli70, FCS15]. orienterend [Dre68]. Origin [PVV07, Cad78, DCC20, KSBŁ95, KÁ04, ON88, RW98, ST09, Van86]. originating [SSK14]. origins [KAN20, RGF21]. Orkney [WNZ24]. ornatus [HdC24]. Orta [BHP05]. orthoclad [Sæt92]. Osaka [TYN02, YO02]. Oscillations [Df82]. Oscillatoria [Ber75, Bij75, VZM75, VM79]. osmoconcentration [Dor79]. osmoconforming [Dor79]. osmoregulatie [Spa71]. osmoregulating [Dor79]. osmoregulation [Spa71]. Ostend [Joi73, Pod74a, Pod74b, Pod74c]. Ostracoda [ABT21, KAY20, KÇA22, MLTH15, OR11, PSSKŽ24a, PSSKŽ24b, RFF01, Slu81]. Ostracods [Kül05, Slu81, MLTH15]. Ostreopsis [TPT22]. Ostroumoff [PFSPdS24]. Osun [AAA^{+}24]. Other [DT92, DDV84, FDP01, GB07, MFS95, MSD08, RFF01]. otolith [PFLO14]. otoliths [KAN20]. Otranto [LAC99]. Ottelia [XWC15]. otter [MNJ21]. Oud [Bak72]. Ouse [PJ95]. outcompetes [GXY14]. outdoor [WRH98, lCdB70]. outlet [Hof99]. outline [Vos87]. output [BP94, IIMRC21]. outstanding [VBK96]. Outwelling [Bou07, CDA03]. ovata [RSM24, TPT22]. ovatoxins [TPT22]. overband [Gie71]. Overijssel [Gar71]. overlap [BNPC07, GSH17, KLG17, SAACR09, SKH19]. override [HLL^{+}25]. overrides [ZXH15]. Overview [BGRM^{+}24, BMO15, BP21, DMN88, McL93, MR87, RL05, Hog69]. overwintering [Sch95]. Overzicht [Hog69]. ovipositing [MMBV24]. oviposition [RHA00]. oxbow [BV96, RPN15]. oxidative [INJ20, SCN^{+}24, WFBJ17]. oxidizing [BVM15, HWS22, YYS^{+}23]. oxygen [BNPC07, HD93, JM08, KK07, Ker81, KK89, LMBM12, LFR75, MR81, dSLM94, Møh99, PMP94, RM81, SHK92, TGZ09, TGE04, Van77b, VL18]. oyster [FDB97, OHA08, YL11, YRR09]. oysters [ACB12, zEGB13].

P [Ano05b, Art99, Gos98a, Ver97, BZV93, FAC16, Fli85, GCL99, JC83, MF07, PHB09, PMZMJ16, PES99, TLF19, VALRdF20, ZD21]. P-loading [PES99]. P-supplementation [ZD21]. P. [dCKL23]. Pääjärvi [JLS09, VS09]. Pacifastacus [OLM09]. Pacific [HBR12, OHA08, YRR09, zEGB13]. pacifica [ACL13]. padanus [RG07]. Paguroidea [FBNF08]. Pagurus [TK19b, TK19c]. PAHs [VV93b]. Päijänne [KPS09]. pairing [Wil07]. Pakistan [GSG24]. Palaearctic [Ash92]. Palaemon [ET16]. Palaemonidae [PNC22]. Palaeoecological [KK12a, AV92, Den94]. Palaeoenvironmental [VD94]. Palaeolimnological [BPA01]. paleocommunity [LPT21]. Pallas [BKD92, DG84, ECR20, EMT00, Lee75, SAN99, YN08b, ZXZ09]. paludosa [Cup94, GD09]. Paluxy [MH13]. PAM [BWS21, Rat14]. Pampean [CTO01, FGF13, GL01, SSI21]. pan [LKSK15, TŠB17]. pan-European [TŠB17]. pans [SFV09]. Pantanal [dSPdS10]. papilio [EBA96]. papyrus [Cha87, Cha89]. paradigms [Lam97]. paradoxa [Pol76]. Paraguay [dSPdS10]. Paraíba [GDB19]. paralarvae [DTH20]. paramecium [SFA05]. Paramelita [DRM18]. Parameter [Rin73]. parameters [AL09, BN24, CvDG07, FGF13, GX11, KEL09, LWW^{+}25, LSY20, LHV87, Rat17, WWŻ06, ŻKŻ10]. parametrization [PMD10]. Paraná [ARP14, ATB07, CB00c]. parasitaire [Swe71]. parasite [BR87, BJS12, DBO11, DH22, FdLY^{+}24, KSY19, MSDDD22, NKKG07, RHH17, RVB08, VvOT09, WNZ24, ZTNW21]. parasite-mediated [KSY19]. parasites [FGL19, Sur01]. parasitic [FCPCMB^{+}25, Swe71]. parasitism [MSW10]. parathion [GFSN04]. parental [BDP21a]. paripes [FLA09]. Park [AP00, BST89, CAJ11, ES09, FM99, HP06, KKO08, PHB09]. parks [LZL22]. part [Ano68a, Ash92, HvB85, SAL16, Sör82]. parthenogenesis [Zad03]. parthenogenetic [YN08b]. Partial [CGO19, Van96, LLG10]. partially [FCS15, MJ93]. partially-mixed [MJ93]. participants [Ano92b, Ano95b]. Particle [HH80, MWC94, BPT85, BM85, BP94, BHG18a, BHG18b, Ker85, SL07, Sta06, VGV94, YS98]. particles [BTA94, GVV86, vDG94]. particular [MSW10, ŠV09]. Particulate [Alt98, PM94, BKW94, Dan84, DL89, GCL99, KSBR94, KSBŁ95, LKE94, PPL16, RFO94, ST09, SFA05, WMM94, WGK95, Wol79, ZV93]. partitioning [CMH14, KT09, KML19, NBD^{+}25, SRF21, SNO98]. parts [BKO92, VVA86, ZZL16]. parva [SAM07]. parvum [ABDPPP^{+}24]. passage [BKC11, KG03, YKJ12]. passes [BMO15]. passing [YKJ15]. passive [AFR14, NP09, VL18]. past [AV92, BVM15, KK12a]. Pasture [SBLdAM23]. Patagonia [GBZ21, DAO24]. Patagonian [GBJM22, BMR07, LCPM15, MAM13, NMB09, SRI20]. patch [CF96, Sch92]. patches [vdHD17]. patchiness [BNMH12]. patchy [BCN12, GFZ15]. path [TJZ^{+}24]. pathways [XYH11]. pattern [DSG^{+}24, FWB05, KMR20, KMS82b, MSP13]. patterning [NWL11]. Patterns [Bar21, DPT04, FCDA09, GSH17, HV79, NZ98, RK02, AXR15, BM03, BLB^{+}23, BMA00, BLP16, CGC19, CHP06, CS13, CMBG20, DTD94, DGGW20, DLG20, EAO20, FCS15, FGM20, GMW09, GJW20, GL92, HF97, HHY21, KA16, KJS15, KÇA22, LMV09, Mar17, MLZ10, MAM13, NPH03, NLC06, Ops80, Pad13, PFLO14, PPV13, RTW17, RA07, SSS98, SMC21, SGJ^{+}24, SHK92, Soi05a, Soi05b, SKM05a, SKM05b, SCC19, SPS22, TS17, TGW06, The02, VSM19a, VIM17, VTE19, Wey09, WS94, ZYL22, dSSR^{+}23]. patulus [HLSN24, SAN99]. paulensis [dCKL23]. Paulo [FBNF08]. Pauw [Gul07a]. Pavin [DK80]. Pb [WMM94]. PCB [CV95, FV95b, HPT95]. PCB-organic [CV95]. PCBs [VV93b]. PCDitch [vLJA07]. PCR [NHiN15, OHA08]. PCR-RFLP [OHA08]. Pearl [BDG10, ZYL22, ZZL16]. Peat [Hig89, Beu71b, ČN10, Har04, Hig71, NČŠ12, RD77, TNtK00, TNtK03, Van89]. peat-bog [RD77]. peatland [HKH^{+}23, Pay10, WvLA14]. Pectinaria [OR06]. Peipsi [BHL09, NLF10]. Pelagic [Jüt05, ASA06, Baa80, Bak78, Bre80, BSB97, DWA11, HKO14, JJG^{+}25, LNHN08, RR98, TFR91, TYN02, ZTG10]. pelamis [WXZ12]. Pelecypoda [Sol97]. pellets [NBT21]. peltata [BHV82, BV96, VVB82]. peltata-dominated [BV96]. Penaeidae [dCKL23]. Penaeus [LLL^{+}23, dCKL23]. penetration [EHM91]. penicillata [DRM18]. Peninsula [FCDA09, RRK15]. Pennsylvania [HC09]. peptidase [Nau00]. Perca [KW12, NHL18, RVB08]. Perccottus [RPN15, ŠDT^{+}24a, ŠDT^{+}24b]. perceived [ZCA20]. perception [PGM^{+}24]. Perch [SGA06, BRA^{+}23, KW12, NHL18, RVB08]. Perciformes [ŠDT^{+}24a, ŠDT^{+}24b]. Perdido [GPGSMM^{+}23a, GPGSMM^{+}23b]. perennial [CAPGA08, Gil20]. performance [ADS25, CKM19, CB13, EBM18, LZD16, LBY17, OTJ19, PJH06, RJ11, SJD22]. peri [GSB13]. peri-Alpine [GSB13]. Peridiniopsis [XCL10]. Period [LMM05, BdSCH09, Bes87, RFR91, RvGB05a, RvGB05b, RGD10, VS09]. periods [AAPdG10, FWR^{+}23a, FWR^{+}23b, WPvB17, ZGZ20]. Periophthalmus [DDPT24, EBA96]. periphytic [dMCFS^{+}23, dNFdNM21, WP12]. Periphyton [Gon79, LJL05, Mol76, PBF06, SON11, BGM08, GC08, GCE11, HB75, IJ06, MS08, PR84, Rod10, Roo79, SGM99, TZD21, VK12, ZWY^{+}24]. Periphyton-macroinvertebrate [LJL05]. permanence [Gru11]. permanent [DB02, SVM08, VP77]. permutation [SJ00b, VHtB17]. Perna [RVJ98, RVJ06]. peroxidation [xHCjX^{+}23]. Persian [SS24]. persistence [ZQH16]. persleiding [Ess71]. perspective [Alt13, JAM15, MG91b, PLM18, VDV86, VN22, VSW00, WL24, Ess06]. Perspectives [SLG09a, VV91, CC99, GGV86, Hos89, LZZ^{+}25, MJM16, VGM93, Gul09b]. perturbed [Adm90]. Peru [FCPCMB^{+}25]. pesticide [BJS12, GFSN04]. pesticides [JGUF23, lCdB70]. pet [BSMVP24, PGM^{+}24, RBAB24]. Peter [vdV01]. petgaten [Hig71]. pH [ČN10, GD09, HHN03, MJJ21, PFMD00, RKS14, RMH22, WGL21, WYW^{+}24, YN08a, ZKN09b]. pH-related [ČN10]. pH/high [WYW^{+}24]. Phaeocystis [JB87]. Phaeophyta [JTL09]. phaeopigment [HB85]. Phalacrocorax [VBA22]. Phalaris [CDL14]. phantom [WFFM10]. pharmaceutical [CMBI25]. Phase [RHA00, DHL16, OSM13, QPP95, SSR13, Van78b, Vie20]. Phases [DvZA90, MOdSCP07, MOPCP08]. phenolic [GM16, SRG18]. Phenology [Kül05, GWF92, Har04, Lar92, DAO24, TS17, VSB10]. Phenotype [ZGZ20]. Phenotypic [BDP21a, RR03, GRL20]. pheromone [FTH22]. Philippines [AAR^{+}24, CAFA04, DN23, ECGC25, PAAJM17]. philopatry [RMH21]. philoxeroides [LWY^{+}24, SJD22, YF18]. phlorotannins [HMH05]. Phocoena [Rei92]. phoenicia [Mou92]. Phormidium [MCB22]. phoronid [MBM^{+}22]. phosphatase [Nau00, NN05a, NN05b, ŠV09]. Phosphate [SGR86, BVV80, DE93, Hos80, PMSS18, VdJ94]. phosphates [Kro80]. phospholipid [ŠNZ21]. phosphomonoesterase [HPPL00]. phosphomonoesters [HPPL00]. Phosphorus [Kou83, Lij86, MGT17, YZY09, AC86, Boe86, BMJ16, Bro81, BR87, CAFA04, DPT04, DYS86, DCL19, DHL16, Gib86, HSP16, HKT16, HM86, JA90, JCL09, Lij91, MF07, PM94, PJ95, QLW^{+}23, RKS14, Rat17, RJ11, Roe96, RTL17, SCN^{+}24, SCO09, SBH17, SGR86, VPM82, Van89, WYZ13, ZWY^{+}24]. phosphorus-and [AC86]. photic [Baa80]. photodegradation [TW03]. photographic [SGM99]. photoperiod [AL09, MH13]. Photoresponses [PAY03]. Photosynthate [PA98]. Photosynthesis [Bes79, HHN03, KG03, KP99, PE78]. Photosynthetic [LKSK15, PSZ79, EHM91, GZG02, GBJM22, MRBQ^{+}25, SH11, ZNP13]. photosynthetically [GEK97]. photosystem [SVB98]. Phototactic [TP09]. phototaxis [YSY^{+}22]. phototrophic [DKI05, RZG09, dWvG88, vG72]. Phragmites [AH82, BZD81, BZD82, GYW17, HHY21]. phreatic [NHF96]. Phryganeidae [VC89]. phthalate [VGP^{+}23, ZXZ09]. phycosphere [ZWY20]. Phyllognathopus [CGD19]. phylogenetic [OTS05]. phylogenetically [TBBS17]. phylogenetics [VSM19b]. phylogeny [DHEG21a, DHEG21b]. Phylogeography [UT16, ŽRP16]. Physa [BA14, ZKN09a]. Physalia [NNJ22]. physalis [NNJ22]. Physical [BD09, Lac91, MIMRRZ08, MFS03, Bel76b, BMJ16, DZG10, HEZ03, Hog69, JB83, KPJP23, LHV87, ÓP04, PF07, PVB20, SZY19, SBS12, WKT05, Wey09]. Physical-Biological [MFS03, HEZ03]. Physico [Gys72b, BG15, CM95a, KSBŁ95, MJ93, REH92, Dum69]. Physico-Chemical [Gys72b, BG15, CM95a, KSBŁ95, MJ93, REH92]. Physico-chemistry [Dum69]. Physicochemical [BHG16, RBN20]. Physio [JDJ^{+}25]. Physio-biochemical [JDJ^{+}25]. physiochemical [KJF00]. Physiological [DZL21, KWO11, SAM07, Bes79, CBP22, HBR12, Leh04, LWW^{+}25, qLlZhY16, Rat17, VvdMP12, WPvB17, WvRvdV03a, WYW^{+}24, YF18, ZBA19]. Physiology [JC80, SL24]. phyto [BV78, BLS88]. phyto- [BV78, BLS88]. phytobenthos [GR91]. Phytoflagellates [VV05]. phytoindicator [MSRB23]. phytophilous [Dvô70]. Phytoplankton [Bak79, CBC99, CH91b, CP10, DD81, FAK01, GSB13, HH91, HRP91, KGB04, KKR02, MFS03, NO85, RR03, RMW91, Sch72b, SPH16, Tol02, TDBW95, Veg79, WWŻ06, YHM05, ZPW13, ZPW15, AAPdG10, AP12, AVdNM19, AR97, BGK85, BKD^{+}24, BV14, BdSCH09, BWS19, BKW94, Bre85, BLG99, BNMH12, BWM86, BN24, CM95a, CH91a, CGL21, CL11, CV91, Clo99, CVG03, CCS20, CVM21, CAFA04, DWV81, De 84a, DG02, DYS86, DH09, DGG15, DGB17, DBG13, EHM91, FAS05a, FAS05b, Fli85, FBR87, FBL13, FBS19, GZG02, GR97, GTI99, GSM07, GCL99, GDT08, Hal76, HHC21, KKI04, Kat92, KPS09, Kla88, KD82, KP99, LSR08, LTH22, MTA15, dCMMB15, MFL09, Mom73, MOCPG08, MOCPG09, ME84, MDV03, NE85, Ops80, PPV13, PLM^{+}22, PDP02, PVV07, PHM21, PSP91]. phytoplankton [PLC08, Rat14, Roi81a, SB08, SSI21, SWF17, SEE22, SAL16, SLC10, STH21, SR88, SGR86, TXW^{+}24, TLF19, TBF99, VVW95, Van84b, Van97, VD87, VS09, VGV94, Vri80, WPvB17, Wan79, WCZ11, WZM^{+}23, WDP15, ZWY^{+}22, ZQH16, BD74, Art99]. phytoplankton-based [SR88]. Phytoplankton/zooplankton [Bak79]. phytoplanktononderzoekers [Ano81a]. phytoremediation [LZL22, PMSS18, SGM22]. pias [Hig69]. Piaseczno [DG06, GJ06, SGGO06, WWŻ06]. Picea [HHN03]. pico [VET20]. pico- [VET20]. picoeukaryotes [SFV09]. picophytoplankton [CMB07, KiNS05]. Picoplankton [UKS05, HNiN03, VMS09]. Pierre [FAM03]. Pigment [Hal76, WS94, WGK95, GDT08]. pigmentation [HH13]. pigments [MRBQ^{+}25, WBB00]. pike [Gri82, Gri89]. pikeminnow [Gar05]. pikeperch [LMV09]. pileus [SAB10]. pilot [OAS13]. pink [dCKL23]. pinnata [SK84, TK95]. pipe [LWdSD16]. pipeline [Ess71]. Pisces [LA99, HPT95, FV74, KHF01, MAM13, Nwa95, PCB10]. Pisidium [DF83, Df82, Pis69]. pit [BDH09]. pits [KMK09]. PIV [Sta06]. plain [Den94, GKK95, Nie73, SC09, LLX11]. PLALM [MG91b]. plan [PP95]. plana [Sol97]. Planch [LSdS^{+}24]. Planch. [PMSS18]. planctoniques [Joi73]. Plankton [Bak72, BTV88, DCC09, AT24, BD74, Bak77, BHL09, BSA20b, dSCdMM09, CB00c, DSG^{+}24, DvZA90, DD99, GB69, GW73, GW74, Hof99, HAS22, JB83, JM84, JJG^{+}25, Kle84, KA86, LHV87, Pee74, QJD02, REH92, SCS12, TTD02, WLB22, WRH98, dP71, Osk72]. Planktonbeheersing [Osk72]. Planktonic [Cad99, dSLM94, BM98, CCS99, FNW94, HJ23, Joi73, KRJ09, KKD02, KMG^{+}23, NudWB^{+}25, Rin97, RRL09, Roi88a, SAGN07, VRH^{+}24, nWdLjZ21]. planktonicum [BR87]. planktononderzoek [GB69]. planning [CPP22]. Planorbidae [KC22]. Plant [Cha87, DCC20, LLX11, BM03, Bau00, BGM08, BJ99, BA97, Bes79, BPA01, Bla82, CBE20, DSN09, DSG^{+}24, DB81, FWR^{+}23a, FWR^{+}23b, HSY18, JCS01, JC83, KMR20, Kou83, LWY^{+}24, LWX19, QLW^{+}23, SFA05, SXY20, TBF09, VBD96]. plant-associated [BM03, BJ99]. plants [AL24, Bel87, Bol04, CvDG07, Gra74, HVK93, JCL09, Koo73, MMRG03, MB81, ST20, SVB98, TG19, TIF10, ZWY^{+}22, Ver97]. plaque [JCL09]. plas [Dre68]. Plasticity [BDP21b, RR03, BDP21a, BDG06, GBZ21, GRL20, GR08, NW24, VEK20]. Plata [CB00c]. Plateau [PKE22a, PKE22b, PLM^{+}22, SD09, ZYS14, Gru11]. platform [RAO^{+}24]. Plathelminthes [Van75a]. Platichthys [FCDA09]. Plationus [HLSN24]. Plecoptera [HH08]. Pleomorphism [Ste76]. Pleurobrachia [FNW94, SAB10]. pleustonic [KPJP23]. plicatilis [YZ08]. plume [MWC94]. plumes [SCG09]. Plumosus [PGV95, MKL10]. plutonium [Bol04]. pluvialis [DCL19]. Poaceae [AH82, IIMRC21]. Podiceps [TE04]. podon [Gie71]. podonspecies [Gie71]. Poecilia [FTP08, GR08]. Poeciliidae [FTP08]. poelen [DM71]. points [Ano68b, Ano68d]. Poland [KK20, KMK09, DG06, DF83, Df82, GJ06, Grz92, KPJP23, Kuk92, MŚSJ16, Nat05, SGGO06, WWŻ06]. polder [Hig89, KK89, Van89]. polders [Pol75]. policies [RBAB24, SWW02]. policy [Dew80]. Political [Par78]. pollen [BPA01, PBR01]. Pollicipes [CY95]. pollutants [BT92]. polluted [HF97, LZZ^{+}25, SVW95, YRR09]. pollution [AIS02, AOS^{+}24, CMBI25, Cra91, DGH^{+}24, EAO20, ER94, EK98, JTGS24, dSLML21, Mas85, MOM24, Ros69a, Ros69b, SON11, Str93, TDBW95, dB69, BH71, Ess71]. Polyalthia [SK84]. Polychaeta [BRJ97, BZ97, Boc97, BSB97, KP97, Sch97b, SVB97, Zet97b]. polychaete [EK93, Rii94, VSM19b]. polychaetes [AFRS23, Lam79, QSA09]. polyculture [LLL^{+}23]. polycyclic [OMNC24]. polyethylene [MRBQ^{+}25]. Polygonum [QCF19]. polymeric [CGO19]. polymorpha [BKD92, DG84, KK11, Lee75, RVK96, WFBJ17]. Polymorphism [YN08b]. polyp [WMM11]. Polypedilum [SS92]. Polyphenols [OFCP22]. polyps [RAP22, TKR10]. polystyrene [LYM21, VD85]. polytene [FKB92]. Polyunsaturated [Jüt05, MCvE09, SGK07]. Pomacea [AR05, DD17, GD09, GFW24, MV14, SMM23, SM12, YB08]. pond [BMA00, BASJ94, FM10, GB69, GRDPP09, Lin78, LTH22, MŚSJ16, MSD08, Par69, PD10, RHO99, SV92, Van91, Har04, HNiN03]. ponds [BdSM20, CGL21, CMBG20, DB02, Dvo69, FWR^{+}23a, FWR^{+}23b, FGM20, GRDPP09, HGM22, JM84, LPT22, Lah98, LLL^{+}23, MvAV20, MTS11, MKS22, NJ00, Nwa95, OTJ19, OR11, PCD11, PVB20, Raa89, SVM08, SR92, SSB21, SJ00b, TK95, VvOT09]. Ponto [BKG16, BHG18a, BHG18b, PDvdV06a, PDvdV06b]. Pontoporeia [Leh04]. pool [Coe78, Hig79, Nor98, Sir17, VSB10]. pools [BKD^{+}24, Bau00, BST89, CCH11, De 84b, DM71, GGT14, Mag00, Sæt92, Sim94, The02, Van96, VC89, lCdB70, vDB78, vDSB80]. poor [ZWF11]. Poppe [GD91]. populatie [Pis69]. Population [GEÓ04a, GMVCLGGP24, GFW24, LMM05, MFTM20, MK05a, NS21, PNC22, STHN03, TD87a, Van91, WF05a, WF05b, Zet97b, AMSN07, BM03, BHP05, Caz82, DMM97, DW14, EBA96, FMdCFdCV24, GSG24, GCT^{+}23, Gil12, HJV08, JTGS24, Kae20, Kap80, KHG20, KM83, LLL22, LM92, Lin78, LXH21, MOM24, NR97, PPK^{+}23, Pis69, DAO24, QJD02, SAN99, SNGV06, Sol97, TYN02, TP09, VH08, VvOT09, Wel74, WF12, vdHD17]. population-scale [DW14]. Populations [GK07, BWM86, But81, Deg10, Don79, Dor87, DF83, EÖ04, dSFFA03, GMD19, HP06, Irv89, Joi73, KM83, NEP05a, NEP05b, NHiN15, OHA08, Pad13, RPSSDS23, ŠV09, TB92, ZL10]. pore [CMG95]. porewater [BZV93]. Porifera [GPS^{+}23]. porosity [LMBM12]. porpoises [Rei92]. port [Per69, YO02]. Portage [SPA12]. Portugal [CM95b, FED95, PN94, RCB95, WNS08, BC95, CM95a, CY95, CV95, DGGW20, FV95a, GV99, MTdS^{+}24, MQC93, PDM95, RBAB24, RQ93, VBA22]. Portuguese [NNJ22]. Portunus [LLL^{+}23]. position [Hey92, Van75a]. positions [AMTSJ15]. Positive [TXW^{+}24, VGG89, KGM13, KGM22]. positively [dSDMD20]. possibilities [Rin76, vdW70]. possible [BBH05, HvdGvS10, Ker91, ON88, ŠV09, TK89, Vij91, Gie71]. Post [YKJ12, BEQP18, BHP05, GLGU22, LOZ24]. post-diapause [LOZ24]. post-flood [BEQP18]. post-liming [BHP05]. Post-passage [YKJ12]. post-settlement [GLGU22]. postlarvae [OR06]. potamodromous [BMO15]. Potamogeton [CM95b, PSZ79, WYZ13]. Potamopyrgus [Dor87, HKT16, HBR12, LDK20]. potassium [Lee80]. Potential [DT23b, HGB20, AA95, Bol07, BvdH90, CCS09, DA99, DCG19, FNW94, GA10, HLL14, HHEM19, JTGS24, KW12, KG03, MCSV21, PGM^{+}24, RHA00, SFA05, SNKN^{+}24, THS17a, THS17b, WPvB17, YP17]. Potomida [MFP01]. pouchetii [JB87]. power [AL24, DSG^{+}24, De 79b, GMD19, Had79, HvAV83, JCS01, Koo73, Met78, Sha78, Wan79]. Powys [CM05]. Poyang [LTH22]. Pozzuoli [ABT21]. pp [Cad05e]. practical [Hos80, LCS01]. practice [Dew80]. Pradesh [BJS22, BN24, KP24, SD09]. prairie [AP00, FM10]. Prasiola [RJD07]. Prasiolales [RJD07]. prawn [ET16, PNC22]. pre [BHP05]. pre- [BHP05]. precipitants [BVV80]. precipitation [SEE22]. Predaceous [TBS18]. Predation [ACH09, MM96, NBM98, SRI20, BMR07, BR05, Bol07, BFS21, CB00a, CNA07, DF11, FGB20, Gar05, GRL20, HH13, IS06, Kae20, KKM09a, NZS16, NHL18, OCS20, RIT04, SDRM16, SOC12, SG20, WV02, WHW14, YYR20]. Predator [Bri99, CCH11, Gil12, GR08, MLZ10, ÅDB00, dFBdGGMLT21, CBNF19, CDD20, DD17, DH22, GBPR23, GAAT^{+}24, HKO14, KK11, LB20, PMA22, PK08, RW98, RMH22, STHN03, Sch97a, SAM07, TS17, VGP^{+}23, ZCA20]. Predator-dependent [CCH11]. Predator-induced [Gil12, GR08, MLZ10]. Predator-prey [Bri99]. predators [FNW94, HKJ12, HEPH09, KGM13, KJP14, LNHN08, PS06, Rei99, SAH10, SGD22, SAR21, TET23, ZYS14, KGM22]. predatory [CDD20, HLSN24, KG94]. predict [JJP24, SPA12]. predicted [DDT24]. predicting [GDD07, PPK^{+}23]. Prediction [DDD07, PFDM99]. predictions [COC18, KA16, PFMD00]. predictive [CFRNT21, VK80]. predictors [BHG16, OLJB20]. Preface [Ano97f, Ano98f, Ano98g, Ano99e, Ano99f, Ano00e, Ano04e, AG07, DH92, Gul97, GD02, GMD10, MQR95, RIMG91, SLG09b, SNG21, Sma97, SG08, TB22, vDJ06, vdVB96, PV68]. Prefecture [WiNDF22]. preference [AXR15, Dor77, GM16, KKM09b, LB10, Mar17, SSK14, ZL10]. preferences [AAA^{+}24, EAO20, Fer92b, HF08, KJS15, KH92, NNN22, RPN15, Sch92, SHP21, ZBA19]. Preliminary [DDC20, DG84, Gib86, Har73b, SDW^{+}24, SR85, GB82, Kal86, NE85, PDD07, RS85, WH92]. presence [BKG16, Hig80, HPM20, Irv89, KK12b, KK11, MŚSJ16, ZYS14]. Presentation [Pol73, Pol73]. Preservation [MV83, XYH11, Gys72a]. preserved [BFB15]. preserving [GRDPP09]. pressure [Ess71, NHL18]. pressures [CBF19, JVCJ25]. Prevalence [FCPCMB^{+}25, SOC12, VvOT09, dNdSBJdCeS21]. prevent [Pee71]. prevention [IFV16a, IBV16, IFV16b]. preventive [SK16]. previous [NMS21, OLJB20]. previously [HLSN24, MBM^{+}22]. Prey [LHV07, PCB10, UOJGRL21, Bri99, DH22, GAAT^{+}24, HBL13, HLSN24, IT09, KKS11, KK12b, KJP14, NBM98, NBT21, OR11, RK02, RMH22, SAR21, SOC12, TS17, VGP^{+}23]. preying [RHO99]. Price [Cad05e]. prima [SKK92]. primaire [Dum69]. Primary [GAA24, NLF10, OMV14, VVA86, WRH98, AC86, Ber84, BC95, CH91b, CZC14, CFP08, CV77, CGZ83, De 78, Dum69, FK81, FBR87, Gol71, GDT08, MGT17, NMB09, PES99, RDG86, RCB95, SAB10, SC99, Str88, Veg79, VD87, WiNDF22, Wet93, WDP15, WMD07, dCAP23]. Primorye [SY02]. principal [DGH^{+}24, VHtB17, Zet97a, VB98]. principles [DHL16, Sta06]. prior [BPD15]. priority [PPCH22]. pristine [dSDLdA22, WB10]. probes [FO98]. probleem [dB69]. Problem [vIR92, Dew80, Had79, Sch74, dB69]. Problems [But81, CGZ83, Cra83, BGK85, BM85, Gul85, Har74, Kel73, Ker85, Muu74]. Procambarus [AP00, AFR14, HGB20, MLG20, RHS18]. procedure [GGD07]. procedures [But81, VDB81]. Proceedings [Ano87]. process [AFR14, CPP22, HÅB08, Rin80, RPSSDS23, VPM82, dK71]. processed [BMA00]. processes [ASLT15, AC16, BGK02, BLS88, DMN88, FSN21, MBMV20, Pel82, PSD97, RBPF11, RCB95, RMC95, SKV02, SAL16, SB02, TCL22, TLH18, WMM94]. processing [CB16]. Prochilodus [PCB10]. Procladius [Lar92]. produced [LSR08]. producer [MGT17, PB04]. producers [SAB10, dCAP23]. producing [DCL19]. productie [Dum69]. productieonderzoek [Beu71a]. Production [CMB07, KKS11, LDDS82, MR87, PPL16, TK88b, Ald78, Ald79, ACA02, Ara01, AC86, BSA14, BEZ91, Ber84, BZD81, Beu71a, Bir78, BC95, BHPT20, BR87, BGC00, Cad80, CH91b, CV77, CGZ83, De 78, Dum69, FRP08, FK81, FBR87, FTP08, FG91, GE04, GMD19, GZN20, Gol71, Gon79, GDT08, Gul74, HB03, KKR02, LAC99, LM92, Mom73, MdMdLB21, NMB09, Nie82, NLF10, Nwa95, OMV14, PBG03, PM00, PLdFC22, PMZMJ16, PA98, PR92, PES99, Raa89, RDG86, RCB95, SMPK09, Sol97, Str88, TK88a, TC15, VVA86, Veg79, VD87, WiNDF22, Wet93, WMD07, YL08, YHM05, Zet97b]. productive [TGE04]. Productivity [TDB13, Bes81, Beu74, CFP08, Das07, NO85, Ops80, SC99, TXW^{+}24, Ver79, WDP15, WF05a, WF05b]. Professor [Gul05]. profiles [NAT08]. profilometry [MG06]. profundal [BW92, RP92, SB88]. progeny [GS11]. program [dR92]. programme [Van86]. programmes [BGK02]. programs [MSP13]. progress [SS98]. project [De 77b, SB78, vdV81, ABF01, BPA01, BBR01, FAK01, Flo01, FDP01, PA78, PBR01, PJF01, RFP01, REP01, RFF01, SB78, WXW11, Meu91, SRH08]. projected [VGV82]. projects [VW78, VV91]. prokaryotic [AOS^{+}24, KKI04]. prolarvae [ECR20]. prolifera [KS79, KMS82a, KMS82b]. proliferation [FAC16]. proliferations [GL19]. prolonged [RP22]. promote [XWJ22]. Promoting [CvNB02]. propagation [CDL14]. propagule [QLZ20]. propagules [WGL21]. propatula [NS21]. properties [BTA94, BV14, BGL13, BG15, Das07, DN23, Har04, KPJP23, ÓP04, OB13, RBPF11, SH16, Żur06a]. proportion [MJJ21]. proportions [SEB11]. Proposal [LB14]. proposed [Cor78a, LWdSD16]. Prosobranchier [Ban75]. prosobranchs [Ban75]. protected [GSH17, HSCR10, PEREBM^{+}23, SWW02]. protection [TW03]. protects [RIT04]. protein [JM84, TK89]. proteins [BDP21b]. Proteobacteria [HYH03]. Protists [UKS05, DDC20, MLTV18]. protoplast [Ste94]. protozoa [dCNGF25]. protozoan [KKR02, NKKG07]. protozoans [SFA05]. protozooplankton [KMG^{+}23]. protuberans [GBM75]. provide [OAS13, PCGW20]. Province [ECGC25, Rav01, SSI21, Bel79, BV76, DT23a, Van79a]. proximity [ZCA20]. Prussian [RRK15]. Prymnesium [ABDPPP^{+}24]. Pseudechinus [GBZ21]. Pseudodiamesa [NH92]. Pseudodiaptomus [GIK21]. pseudokarstic [OBF10]. Pseudomonas [YZY09]. Pseudorasbora [SAM07]. Pseudothelphusidae [RPSSDS23]. PSII [BWS19, SVB98]. Pterygoplichthys [PGM^{+}24]. Ptychocheilus [Gar05]. public [PGM^{+}24]. Publication [Ano77b]. Publications [Ano76a, BED87, DDD84, DD79, Dav81, Dor73, Dor80b, DL82, DT83, Dor86, DDC88, HP75, Mur74a, Rep79, SD83, VK82, vD84, vDDS87, vDDD88, vGPD78, vGDD85, Tol84]. Publishers [Ano05b, Cad05e]. PUFA [GSM07]. pulchellum [IM06]. pulex [RW98, ÅDB00, DGD06, KvE21]. pulicaria [NZS16]. pulp [SN03]. pulse [LZY21, Rat14, YHM05, dSPdS10]. pulse-amplitude [Rat14]. pulsed [SBH17]. pulses [CVG03]. Pumping [MFP01, MKL10, RL05]. pumpkinseed [MMM11]. punctatum [QCF19]. punctipennis [PAY03]. punctulata [VTE19]. pupal [BW92, GL92, KKV92b, Rus92, WW84]. Purified [Zoe72]. purpurea [Gra12]. Pygospio [KP97]. Pyrénées [GL92]. pyrenoidosa [HZZ22]. pyriformis [SFA05]. Pyrolysis [BH82, HK82].

Qua [EAB08]. quadrangula [GTI99]. quadrat [VP77]. quadrata [NNJ22]. quadricauda [Ste76]. quagga [BC23, ISJ10, OTS05, OMM05a, OMM05b]. Qualitative [CMV91, RB82]. Quality [Laa97, PBF06, AA95, Ano87, BMR07, BB87, BB78, BB23, CKM19, CMBI25, CLL10, CZZ^{+}25, DH01, DGH^{+}24, DVK78, FBM22, GS11, Gri89, GX11, HVK93, JJG^{+}25, KCR92, Klo76, LSR08, LTH22, MM96, MR17, Mol80, MSF21, ME84, NJ00, NvZvG16, PBG03, PC14, Ric86, RDG86, RBN20, SD86, SDN25, SdAdC^{+}25, TBF09, TG88, Van79a, VK90, Van87, VOD93, VDB81, Ver76b, VHN00, WDA15, ZWF11, dL74]. Quantification [MKL10, BKC11, TGW06]. quantify [BSMVP24]. Quantifying [BWS21, zEGB13, ZCZ14]. Quantitative [Jan75, Kor98, PLC08, SG08, vGR98, Beu71b, CMV91, ES09, Her83, Kle84, Kor14]. quantity [GX11, KvE21, LSR08, RBN20]. quantum [GEK97]. quiet [OSM13]. quinquefasciatus [MMBV24]. Quotas [HG05].

R [Ano81a, Gos98b, Gul98, Hod98, Ver97]. R. [STV06]. R.-D [Gos98b]. R.Br. [CBE20, TK95]. racemosa [CCS09, WZL18, WL24]. raciborskii [KPV16, LZH^{+}25]. radial [LMBM12]. radiata [WKT05]. Radiation [JLS09, CC94, De 00, DGB17, GEK97, HPS11, HH13, KA97, PA98]. radioactive [Gul85]. radiocaesium [KSBR94]. radiocesium [KSBŁ95]. Radiometrically [ABF01]. radionuclides [Bol04]. Radix [LWX19]. Raf. [Gee75]. rain [TFP06, VDB81]. rain-fed [VDB81]. Rainfall [BSA20b, PVH17]. ramada [KHF01]. Ramesh [SNG21]. ramets [MLL21]. ramosissima [HL95]. Rana [GK07, LB20, LWW^{+}25, STV06, ZYS14]. Range [BF12, DN23, KG94, OMM05a, OMM05b, AKL15, MTdS^{+}24, PR16, SS99, Van75a]. Range-extension [Van75a]. ranges [BD74]. ranids [STV06]. rank [DDF93]. rank-frequency [DDF93]. rapid [MFP01, DAO24]. Rapporten [Ano72e]. Rare [BBF17, MHYL23, MTdS^{+}24, NNB^{+}24]. rarely [PFMD00]. Raritan [BRG02]. rate [BB12, BWM86, GSK08, GX11, MM96, dSLM94, MKL10, MFP01, TK88b, WDP15, YN08a, ZLT22, ZCA20]. rates [ABF01, Bes79, Bir78, CSV19, DB85, Dev09, FNW94, GIK21, GDT08, HPM20, KiNS05, LAC99, LMA10, PHM21, RGF21, RC98, RCL94, TGV00, UET12]. ratio [LZN14, MF07, PPL16, PMZMJ16, Rei99, RS85]. ratios [CAFA04, PDP02, WMM94, XYH11]. Ravella [Rav01]. rDNA [vGVV11]. re [xHCjX^{+}23, WPvB17]. re-illumination [WPvB17]. re-watering [xHCjX^{+}23]. reach [SHS93]. reaches [CM95b]. Reaching [JDF23]. reaction [MLZ10]. reactive [SK16]. reactivity [MWC94]. reality [SV92, SS98]. realm [KTH14]. rearing [BVM15, OCS20]. REBECCA [MDP08, SRH08]. rebuilding [WKK24]. recapture [DB02]. receiving [Cha87, Kou83]. recently [BDH09, FM99, LvBR06, vGVV11]. Reciprocal [SGJ^{+}24]. reclaimed [LZL22]. reclamation [FN94, Gos99]. recommendations [BGK85, CGZ83, RSK09]. reconsidered [SR85]. Reconstruction [Das07, RP08, XYH11]. record [Gud04, MSDDD22, PBR01, ŠDT^{+}24a, ŠDT^{+}24b]. recorded [Rus92]. records [RBAB24, WZ92]. Recovery [SOD92, XWC15, KN86, NEP05a, NEP05b, RC98, Van96]. recreational [CvNB02, NKKG07]. recruitment [AOLLJ24, BH22, PD10, RTJ11, RNJB20, TCL22]. recurrent [GL19]. Recycling [Ker91, ACE10, BBF17, SB04]. Red [PGSL02, GXY14, HKB19, KK09, MLG20, YSY^{+}22, Cad05e]. redactie [Ano68f, Ano68g, Ano69e, Ano69f]. redefine [PGT20a]. Redeke [Eng74]. redescribed [ARG20]. Redescription [SS92, Jac92]. redox [Har04, VM25]. reduce [BMJ16, KT15, SBLdAM23, TSV16, WPB08]. Reduced [SWB13, DZG17, Ess03, SHK92, VVW95, WZM^{+}23]. reducing [Gar05, HKH^{+}23]. Reduction [MS02, DBG02a, Gib86, KRZ03, Lij91, REH92, SOD92, dR92, vIR92]. reed [Roo79, Roo82]. reed-stems [Roo79]. reef [GFZ15, JCS01, NvdVdlM01, NYT09a, TXW^{+}24, TDHH23, YL11]. Reeuwijk [VK90]. reevesii [DDZ^{+}24]. reference [CSM08, De 84a, KBO74, Mar01, MSW10, iNTT05a, iNTT05b, SKRB09]. refining [GR97]. reflect [FBL13]. reflectance [Das07, RDG97]. reflected [LLX11]. reflections [Bak78, HBB13, NCT91]. reflects [PdCF^{+}25]. Refuge [HEPH09, WHS13]. refugia [DRM18]. regard [PGT20a]. regarding [Ald78, HKJ12]. regeling [Blo70]. regeneration [CVG03, GC23, TGD16, UET12, Wet93]. regime [BRB97, FBL13, Kae20, Ker81, KK89, LRMdF19, SH11, ZXH15]. regimes [DGN21, LTL09, LFR75, WRS05]. Region [NAM19, BKD^{+}24, FNS^{+}24a, FNS^{+}24b, GSB13, GPGSMM^{+}23a, GPGSMM^{+}23b, GE80b, GZN20, Kre79, LR92, OHA08, OBC82, VNSFG16, Ver80a, dK71]. Regional [AXR15, CGC19, CMBG20, MB14, RFGLSB20, SON11, TJZ^{+}24, VS96b, ZXH15]. Regional-scale [AXR15]. Regionally [ZYL22]. regions [Gil20, QLZ20, SH11, WBR17, dG71]. regression [MAS14]. regulate [NHL18]. regulated [VGJ20]. regulates [GFF17]. regulating [BWH02, Deg10]. regulation [Dor77, SAL16, SB14]. regulatory [SXC^{+}25]. rehabilitate [RSK09]. rehabilitation [JJG^{+}25]. Rehm. [SMPK09]. reidi [AXR15, FMdCFdCV24]. reintroductions [WWS20]. related [ČN10, CPM09, GSB13, GBZ21, KTH14, KML19, LSY20, MMM11, MBM15, NMS21, PFLM15, PPY^{+}22, QPP95, SH16, SHK92, Soi05a, Soi05b, SG08, TAN14, VvdMP12, VBD96, WWS20, YN08a, ZWY20]. relatie [Dor70b]. Relation [Bro84, DG06, AP00, BV78, Bes79, Caz82, CM05, DDV84, Dor70b, FGL19, HPT95, JVG95, LNHN08, dSLML21, Mas85, MGCC93, NMS21, OR11, SAN99, SGR86, TNtK03, Van87, VSB10, WvRvdV03a, YSY^{+}22, YZ08]. Relations [DS75]. Relationship [HH08, VvOT09, CZZ^{+}25, Gie71, GCE11, LKM01, MBFD24, OMV14, PE78, SVW95, TFR91, TXW^{+}24, VIM17, WNZ24]. Relationships [ZPD06, ABF21, Bak79, CZC14, DJD92, DCC09, FAB18, Her83, KPJP23, MLB16, NB22, PPC08, PHM21, TLF19, TDB13, WWŻ06, ZCZ14]. Relative [SDRM16, WCT00, BBM09, Clo99, ECGC25, HdC24, KFS04, RB08, SKH19, ZYL22]. relatively [YRR09]. Release [VPM82, Van89, BSMVP24, Boe86, Cha89, KGM13, KGM22, KSBR94, NYT09b, WLN12]. released [GM16, ZG02]. Relevance [FRP08, Coe75, EBM18, FBS19, GBJM22, KEL09, M.75, RG07]. relevant [De 97, Sta06]. Relic [CF96]. relicta [NAT08]. relies [JBS05]. relocation [Gar71]. remains [RFF01]. remarks [Bro98, dW71b, vL86]. remediation [SDN25, TBS18]. remedy [Had79]. Remoray [BVM15]. remote [Alt98, DDPT24, DDT24, ECB09, JCF17, SD86]. removal [CYK13, HBO03, HW13, HLL^{+}25, Kou83, LLG10, WP12]. renewable [AL24]. Repellents [Jüt05]. Replacement [Kro80]. Reply [WvRvdV03b, ŽRP16]. Report [JTGS24, PFSPdS24, Van76, RS85, Ano72i]. Reports [Gel70]. representative [SV92]. representatives [vDG94]. Reproduction [Sol97, Ald78, Ald79, Boc97, FAS05a, FAS05b, Gil20, GKM17, HVS87, INJ20, KRATA20, LYM21, MCvE09, MJEC05, RAP22, TZD21, nWdLjZ21, YN08a]. reproductions [YYR20]. Reproductive [GSG24, HdC24, AOLLJ24, BW92, CY95, CB00c, CM05, DW14, FTH22, GIK21, GBA20, IIMRC21, KBO74, MKA79, MH13]. Republic [ČN10, DG02]. requirements [Mar01, SPB97]. Research [Adm90, Bel79, Dav71, MB81, Mol72, MK05b, Ano87, Bel76a, BB78, Beu71a, Beu74, Cla84, CRR22, De 77b, Dew80, Fra80, GB69, Geu84, GGV86, HS85, HS86, Hig89, Klo76, Lam97, PA78, Rin76, Rin81b, Rin81c, Rin97, Roi88b, Van76, Van86, VD94, vdB70, vdV76, Fon71, lCdB70]. reservation [Dvô70]. reserves [PPY^{+}22]. Reservoir [DG06, Kuk92, WLB22, XCL10, BB00, BdSCH09, BNPC07, BC23, BLC12, DD99, DCC09, DKI05, dNFdNM21, GSM07, GSK08, GJ06, HHC21, HWS22, KH92, KIG06, MS07, MKG03, MIMRRZ08, MdMdLB21, MOCPG08, SGGO06, TLF19, TK19a, VGV82, VEK20, WCZ11, WWŻ06, CB00a, CB00b, PLM^{+}22]. reservoir-bay [WCZ11]. reservoirs [CGC19, De 76a, GB07, KKI04, KEL09, KKV92a, Klo76, KKV92b, LSdS^{+}24, MCSV21, PAP24, RP92, SWF17, Ver76b, vGVV11, vdV76]. resident [GC05a, GC05b, MGC19]. residues [PJF01]. Resilience [Flo01, FWR^{+}23a, FWR^{+}23b]. resins [VVW95]. resistance [BWS19, LMBM12, LMP02, MRR10, SAR21]. resistances [YF18]. resistant [MRR10]. resolution [ODI08, RRS^{+}22, RRS^{+}23]. resolved [DGN21]. Resource [GE04, SRF21, WXL16, BRA^{+}23, GA10, GSH17, KFS04, KML19, LA19, MLG20, SXY20, SNO98]. resources [CS13, NBD^{+}25, QCF15]. respect [KC82]. Respiratie [Dor70b]. respiration [BBF17, Dor70b]. Respiratory [Vos82, LAC99, VvdMP12]. Response [CVG03, Ess03, KOA00, Lij91, LCL16, MOPCP08, QZL^{+}24, RP22, TDL23, VB98, VHtB17, ZWY^{+}24, BGRM^{+}24, BJ99, Cra85, DDC20, DBG02a, DZL21, FSN21, GIK21, HKT16, KT09, KRJ09, KP97, LLG10, LRSN22, MS07, MSB15, Mur74b, NP09, Pay10, Rat14, SVB97, SS99, SOC12, SHK92, SON11, TP09, TPS09, UT16, VPF21, WYZ13, WHS13, WBS22, WBS24, WDP15, WYW^{+}24]. Responses [CAJ11, HMH05, Nan00, vOCRM19, QLW^{+}23, TCL22, VGJ20, WZM^{+}23, AND25, BPA01, dFBdGGMLT21, CLL10, CBP22, COC18, DPL03, DDY^{+}23, EAO20, GMVCLGGP24, GRL20, HS85, HS86, HLSN24, HXR09, HWS22, xHCjX^{+}23, JVCJ25, JDJ^{+}25, KNH07, LLX11, LYM21, qLlZhY16, OMNC24, PS06, PLC08, RR98, Rei92, SGRMP20, STTS18a, STTS18b, SVM08, SRH08, SG08, SAF10, SAM07, SG20, TLF19, TET23, THS17a, THS17b, TK19b, TK19c, TIF10, WGL21, ZWF11, vGR98]. Resting [Bak94, BdPP17, LAZ21, LOZ24, NBT21, OSM13, dSSSES21]. Restoration [BA97, Kro86, MRD89, REH92, Roe96, SOD92, dR92, vIR92, AVD07, AAV07, BVI16, CMM05, DGD06, DTC19, GGV86, HM86, Hos89, Mos97, PPY^{+}22, SK16, VV91, Van86, VPS21, ZPD06, zEGB13]. restore [YO02]. restricted [Kra95]. Results [KD82, Baa79, But81, FA99, FO98, Gib86, HH80, Kal86, KKO08, MRD89, SRH08, SR85, TV85, Van77a, VKG95, VGM93]. Resuspension [GVV86, MFS03, AGB94, BAG94, KMG^{+}23, PMP94, SLC10]. Retention [KHD99]. reticulata [GR08]. Retirement [Gul07a]. returning [CDA03]. Reunion [JVCJ25]. reveal [AKF^{+}23, DÖD21, KA16]. revealed [FSdO20, PFSPdS24]. Revealing [DAO24, SPB^{+}24]. reversals [FC93]. Review [Ano81a, Art99, Cad04, Cad99, Cad01, Cad05c, Cad05d, Cad08, DCH93, Ess06, Gos98a, Gos98b, Gos99, Gul98, Gul99, Gul09a, Gul09b, Har73a, Hod98, Laa97, Moo98, Nie98, Sar05d, Sar05e, Sar05f, Shi98, Ver97, War97, vD97, vdV99, vdV01, Ano00a, Ano00b, Ano00c, Ano04a, Ano05a, Ano05b, Bar21, Bas20, Boc97, BMJ16, Cad80, Cad05e, Cad05a, Cad05b, CBP22, CKK99, De 77b, DV92, GPV19, Hey92, IBV16, Ken11, KD82, Kra95, MA95, MHL05, OTS05, ONA^{+}23, PSD97, PLM18, RVJ06, Rii94, Sar05a, Sar05b, Sar05c, SL24, Sur01, VIH16, VDD77, WF12]. reviewed [LR92]. Reviews [AdJD95, Dor89, DAR97, MDB91, RDS92, VRP94, vGDR96, Ano97a, Ano98b, Ano98c, Ano98d, Ano99b]. revised [KvE21]. Reynolds [Art99]. RFLP [OHA08]. rheocrene [PKE22a, PKE22b]. Rhine [AvdV90, BKO92, BvdH90, DBV96, DvZA90, DF90, GD91, HvAV83, PW73, PP95, SVK94, SVW95, VV90, VVK96, VVW95, WW84, Wol74, Zoe72, dHC76]. Rhizoclonium [Nie74, dHC76]. rhizome [TCL22]. Rhizophydium [BR87]. rhizopod [HD79a]. Rhizopoden [De 79a, HD79b]. Rhizopods [De 79a, HD79b]. Rhizosolenia [Baa88a, Baa88b]. rhizosphere [HHY21]. rhodani [IGS10]. Rhodeus [DBO11]. Rhodophycean [HL95]. rhythm [Jar08]. rhythms [OCS20]. Ria [CM95b, AA95, ACA02, CMG95, DGGW20, DMD95, FV95a, FV95b, MQC93, PDM95]. Rica [CDW92, WH92]. rice [Fer92a, KGM13, KGM22, LAM06, LPB07, SD09, TCL22, VN22, WRS05, YYS^{+}23]. rich [KK17, Kle84]. Richard [GW73]. Richards [GW74]. richness [CDW92, DEP22, dSDMD20, HXR09, JM08, NB22, Pip87, SLM17, WB10, ZCZ14]. Rieth [LL77]. riffle [MB14]. riffles [PF07]. Riga [Leh04]. riihimakiensis [KSW92]. riihimakiensis-group [KSW92]. riisei [dPBCP22]. Rijkdom [Gys72b]. Rijn [BB78]. Rijnland [Kla88]. Rijnwater [Zoe72]. Ring [SEB11]. Ring-based [SEB11]. Ringelberg [FDG97]. rings [FKB92]. Río [CB00c]. Riparian [TBSZ22, DN23, EPL^{+}23, FCG16, HEB00, HC09, IGS10, QLW^{+}23, SMC21, SGJ^{+}24, VNKdOR15]. riparium [Nie74, dHC76]. riparius [GPA98]. rise [RKS14]. rises [De 79c]. Rising [BBD17]. risk [AdAPC20, FGB20, MLG20, RHO99, SOC12, SG20, WWG^{+}25, YYR20]. risks [Str93]. Risso [KHF01, vdVP76]. Rit [PHB09]. River [ATB07, AP12, BDG10, ET16, EBA96, GL92, GYW17, IJ06, MH13, NBD^{+}25, NE85, OSO09, PDP02, ŠDT^{+}24a, ŠDT^{+}24b, SC09, STT01, TJZ^{+}24, TKD01, VGV82, VBA22, WXW11, WTC09, YBZ20, ZZL16, AVD07, AANO25, AL24, AYM09, BB00, BB12, BB78, BSA24, BFB15, BF12, BWH02, CB00c, CPM09, CMBI25, COC18, CFRNT21, DSG^{+}24, DBV96, DGD06, DMD19, ECL94, FdLY^{+}24, FBM22, GA10, GPA98, Grz92, HPPL00, HAS22, HHP12, JBB07, JGUF23, KA16, KSF95, KKO08, KHD99, LMA10, MHYL23, MCC21, MS11, MB06, Nwa84, PU15, RdMJN22, Rav01, RAD21, RPN15, RHL02, RFGLSB20, Rus92, SB08, SC00, SB04, SDN25, Soi05a, Soi05b, SLL09, SdAdC^{+}25, SDW^{+}24, SGK07, TK88a, TK88b, TMM21, TBF09, TH16, VV93a, VVK96, Van78c, WWS20, WMTR07]. river [ŽG09, dSPdS10, dCAP23, dJG07, BGRM^{+}24, BB87, Bol04, BvdH90, CSK10, Cla84, DDD07, DSL12, DBG13, ESG04, Ein04, FAM03, GS04, GZN20, HL95, HBR12, HPT95, JCF17, KF95, Kra95, LCS01, MLB16, NBT21, Nwa84, vOCRM19, OMM05a, OMM05b, PMB^{+}24, SB08, SSI21, SDN25, TDBW95, VV90, VSA97, VVW95, WW84, ŽG09, ZYL22]. river* [VSW00]. river-floodplain [MS11]. river-floodplains [VVK96]. river-shed [BB00]. Riverine [SHS93, Alt13, BdPP17, BBF17, BTP24, DRM18, DBO11, DvZA90, HLL^{+}25, LSR08, ONA^{+}23, RFGLSB20, SAACR09]. Rivers [KGB04, AL24, AYM09, BMO15, BBP08, BKO92, BHG16, CB16, DvZA90, EAO20, FNS^{+}24a, FNS^{+}24b, FBL13, GL01, HvAV83, ISJ10, Mar01, MRR10, PPCH22, PW73, SK04, TWC13, VGJ20, WXL16, WiNDF22, WDA15, dCAP23, dHC76]. riverscape [RBN20]. RNA [LZN14]. RNA-to-total [LZN14]. Roach [SGA06, BJJ12, KK11, NBM98, Raa89, ZSG14]. road [KCG05, LAZ21, PD10]. Robarts [Gul98]. Robert [Gul05]. robust [HEF13]. robusta [Baa88b]. rock [BKD^{+}24, Nor98, Sir17, The02, VSB10]. rockfish [KKS11, LCPM15]. rocky [AAMS24, LWM89, NEP05a, NEP05b, SRI20]. Rodó [CMM05]. roeseli [Dor77]. Roijackers [Ano81a]. Role [BASJ94, RMC95, ACE10, BH92, BNPC07, Cap75, Coe97, DBO11, DE93, DDPT24, FM99, FNW94, GS11, GSK08, GCL99, HD93, Hig80, HV89, Hui88, HvdGvS10, Ker91, Kur02, Lam79, LCPM15, Lee80, Lij80, Mag00, MCvE09, MMRG03, Mol80, Nie79a, OR06, PBM21, RMH22, SNG21, SFV09, SHP21, SBS12, VdJ94, Ver80b, VEK20, WvLA14, WFBJ17]. roles [BAdOFC21, HHA92, MKS22, TWC13]. Rome [CSK10]. room [OB13]. Root [HSY18, RRRRA07, xHCjX^{+}23, LMBM12, QLW^{+}23]. rooted [SB14]. roots [DHPN^{+}24, PLdFC22, QLW^{+}23]. rostriformis [ISJ10, OTS05, OMM05a, OMM05b]. Roth [Nie74, dHC76]. Rotifer [OBF10, BB00, Dev09, GFSN04, HLSN24, LSY20, LYM21, NR97, OLTV19, VHB09, WWS21, YN08a, YYR20, ZXZ09]. Rotifera [EKK21, GFSN04, LRSN22, NS21, SAN99, SAGN07, VKD77, YZ08]. rotifers [Gil12, Gil20, MNS05a, MNS05b, Nan00, Nat05, NKKG07, RM98, RS85, STHN03, SNGV06, SAGN07]. round [BF12, PR16]. Rousselet [OLTV19]. routes [VRH^{+}24]. Rozema [War97]. rRNA [FO98]. rubens [SAGN07]. rubra [CCH11]. Rudd [SGA06]. Rug [vdV76]. ruimtelijke [Fli70]. rule [SA95]. run [HKB19]. running [Aag92, Ano05b, DTC19, HM84]. runoff [FV95b]. Ruppia [Ver79]. rupture [PdCF^{+}25]. Rushes [MSRB23]. ruspoliana [MTdS^{+}24]. Russia [KBR10, PMD10, TSZ10, Zdo09, BHL09, DZG10, FLP08, GZN20, Ker92, MKG03, RZG09, RGD10, SY02, TGZ09, ZTG10]. Russian [Gul98, Bol04]. Rutilus [BJJ12, KK11, Raa89, VPS13, ZSG14]. Rwandan [WDA15].

S [Art99, Gul09b, Nie98, Shi98, SC09, Van78b, War97, CMG95, GX11, dSLM94]. S-transferase [GX11]. S. [HHN03, Vaa79, ZV93]. S.E [PJ95]. S.W [BV78, TFR91, vG84]. S.W. [BD74, Bak80, BTV88, BV89, Pee74, VV93b]. S.W.-Netherlands [BD74, BTV88, BV89]. Sabine [NBD^{+}25]. Saccostrea [ADS25]. Sacramento [Gar05]. Sado [FV95b, RCB95, RQ93]. Sagartaal [KP24]. Saginaw [PHD13]. Sagitta [FNW94]. Sahel [Lah98]. Sailfin [PGM^{+}24]. Saint [FAM03]. Saint-Pierre [FAM03]. Sakai [YO02]. Sakai-Semboku [YO02]. Sakarya [SEE22]. Salado [SSI21]. Salamander [GK07]. salina [GAAT^{+}24]. Saline [MR81, Das07, DG02, DZG10, GZN20, KBR10, LDDS82, MKG03, MAD10, PD02, QJD02, Sep79, Vaa79]. saliniteit [Dor70b]. salinities [ZLT22]. Salinity [ADS25, HBR12, LOZ24, MTA15, AND25, AT24, BD74, Bro84, CM05, DDT24, Dor70b, Dor74, FV74, FC93, FASdPN^{+}25, KK12b, KNH07, KS79, KRATA20, Mar93, MJEC05, MHE93, NC20, PC14, RVK96, RP22, SASA^{+}24, SCN^{+}24, SNGV06, SSR13, TCL22, YZ08, ZLT22, ZKN09b]. salinity-temperature [Dor74]. salinization [ŽAL24]. salinus [TSZ10, Tol02, ZG02]. Salix [DZG17]. Salmo [AK09, GS04, HFS08]. Salmon [KAN20, GA10, DAO24]. Salomons [Gos99]. Salt [Dor77, Nie73, ACA02, BVD82, CF09, DSN09, Dan84, Gra74, GYW17, HYH03, Har73b, Hui88, LLF99, LAZ21, MTA15, Mos97, Ran74, RA07, TSZ10, TG19, Wol79]. Salt-marsh [Nie73]. salt-marshes [Har73b]. salterns [GAAT^{+}24]. saltmarsh [CF96, GC05a, GC05b, VP77]. saltmarshes [Pol76]. salts [PD10]. saltwater [Wol00]. salve [Sch83]. Salvelinus [AK09, BTB09]. Salvinia [HHEM19, PLdFC22]. Samborombón [SC09]. same [AYM09]. Samenvatting [BM70]. Samenvattingen [Ano70c]. Samoroda [GZN20]. sample [Ano68d, HCS02, KZV02, RJ11, SV92, Ano68b]. sampler [Kor98, Kor14, Mer79]. samples [BH82, Kle84, MV83, Roi81a]. Sampling [BR81, Goo79, HV79, LPT22, MBM15, SR85, VP12, WW84]. San [Bau00, GIK21, LSR08, LME10, WMM11]. Sanchez [Art99]. Sanchez-Castillo [Art99]. Sand [LMM05, DRN09, FA99, KMK09, Mas85, VS96a, WAW92]. Sand-Bubbler [LMM05]. sand-flats [VS96a]. Sander [LMV09]. Sandin [Ano05b]. sandy [AAMS24, BGL13, BZV93, HWG77, JLG06, Mas88, SKM05a, SKM05b]. sanguinea [ZWY20]. Sanjiang [LLX11]. Sanyati [PDD07]. São [FBNF08]. Sap [AR97]. saprobity [Sch88]. Saprolegniales [MKG20]. Sargassum [JTL09, dSLML21]. Sargassum- [dSLML21]. Sarracenia [Gra12]. Sars [Pri03, YTD02]. satellite [SKV02]. satiation [MLG20]. saturation [FM10]. savanna [MCC21]. saving [Kle84]. Saxonian [LKE94]. Say [MF92]. Scale [CCM21, AXR15, Bau00, COC18, CMBG20, DW14, GAA24, GC05a, GC05b, KEH09, KCG05, LWO01, Roi88a, SKRB09, VS96b, WPB08, ZBB19]. scale-bearing [Roi88a]. Scale-sensitive [CCM21]. scaled [Roi88a]. scales [BHG16, CMG95, Coa07, CHF09, HV79, PF07, SEB11]. scalloped [TRHHAR10]. scans [MG06]. scarcity [MGT17]. scavenger [EPL^{+}23]. scavenging [NNJ22]. scenarios [HHEM19]. Scenedesmus [DDY^{+}23, GBM75, KA97, MNS05a, MNS05b, MMV05, Mur69, Mur74b, Ste76, WBB00]. Schale [Ban75]. Scheffer [Moo98]. Schelde [BLS88, HB88, YMB93, YME98]. Scheldeboorden [Gys72a]. Scheldt [BV78, Str88, BH71, BZV93, BVD82, CN94, FvM03, Gys72a, KP99, LEB93, MEB93, Par78, PP95, PN94, Str93, VOD93, VML93, VV93b, ZV93, dP71, dHC76]. scheme [Blo70, SMJ95]. schemes [Cor78a, Cor78b, Met78, Mit78, Pra78, RB78]. Scherffel [JB87]. Schiemer [Gul09b]. Schmidt [BBP08, OFCP22]. science [EFD99, MTC^{+}24, MTdS^{+}24, PFSPdS24]. Scientific [Ano76b, Dew80]. scientist [BSMVP24]. Sciomyzidae [WMV09]. scleractinian [WLN12]. scoparia [SMGAGG00]. scope [Ara01, Cra87, VSM19b]. Scotland [DMD95, MA95]. Scottish [Pay10]. Scrobicularia [Sol97]. sculpin [CS13]. Scyphozoa [MAA22, NNB^{+}24]. SE [GBT19]. Sea [ABT21, Bal09, Ban75, FG91, GCS93, KKS11, LAC99, Leh04, MS08, MTC^{+}24, PGSL02, Sch95, Van78a, ABF21, CDA03, DGGW20, ECGC25, GBJM22, HEZ03, iNTT05a, iNTT05b, PAAJM17, Sch78, SHP21, SRI20, TFT10, WKH82, WA84, Alb04, Alt98, BHM03, BSB97, Cad78, Cad80, DGN21, DBV96, De 80c, DJD92, EMT00, EMN04, Ess71, FD94, FN94, Fra80, FNW94, GUT21, Gie71, HF97, HHA92, HZC95, KIJ11, KSBŁ95, Kow94, KP97, KHK17, LMV09, LKE94, MAS14, Mom73, MML^{+}22, MML^{+}23, NZ75, Nau00, NEP05a, NEP05b, Ole97, OMM05a, OMM05b, Pus94, Rei92, Sch97b, Sch95, SAL16, Sör82, TB92, Van77a, Van82b, VET20, VPS13, Vos87, WWG^{+}25, Web71b, WMM94, Zet97b, Zim78, DT92]. seabass [GUT21, ISY10a, ISY10b]. seabird [TB92]. Seagrass [AAR^{+}24, AKL15, Bar21, KKS11, PEREBM^{+}23, SHV96, WWG^{+}25]. seagrasses [PGSL02]. seahorse [AXR15, FMdCFdCV24]. seahorses [VP12]. seals [dLFTB22]. Seamounts [LLL22]. search [Aag92, CDD20, Ker83]. Searching [CFRNT21]. season [JLS09, KLG17, LSR08, RS85]. Seasonal [AAMS24, AVdNM19, AR97, Bal09, Bar77, BTB09, BN24, CCK10, DKI05, GOB24, GS04, GSM07, GPA98, Har04, HNiN03, HHP12, KKI04, KMS82b, KHG20, KIG06, Leh04, Lu05, MG91a, MDH93, MK05a, MAD10, NSB24, PDM95, QCF15, RNJ97, RM00, SWF17, SC00, SB04, SB14, SGK07, Wel74, WMD07, YKU01, dRFC21, AAPdG10, BDH09, CB00b, CM05, CVG03, Cru10, DRM18, DGGW20, FNW94, FWB05, Gie71, HH91, HEPH09, IRT06, JJP24, LXH21, LZY21, LMP02, Mar93, NW24, Nie82, PMB^{+}24, PEREBM^{+}23, PD10, SB08, VV85, VEK20, WNS08]. Seasonality [HLL^{+}25, RAA98, CB94, MBM^{+}22, PKE22a, PKE22b, TBSZ22]. seasonally [BEQP18, GFZ15]. seaweed [GJW20, RSM24]. Sebastes [KKS11]. sebifera [MSF21]. Secchi [Bui95]. secondary [BHPT20, De 78, Gul74, PES99, SAB10, WRH98]. secretion [OR06]. section [Ola92]. Sediment [KSF95, MFS03, PJ95, QPP95, SMM23, ACA02, BNV03, BZV93, BSA20a, CMG95, CN94, CDL14, CLH06, DH01, DV92, Dro84, Fer92b, FN94, FCW94, GEÓ04b, KFS04, Kap76, KIJ11, KHM04, KMG^{+}23, KCG05, LMBM12, LLG10, LZD16, LBY17, LTH22, PMP94, PN94, Pus94, RFF01, RHO99, RCB95, SMPK09, SL07, SHK92, SVW95, SOU14, SGGO06, TDP06, VV90, VOD93, VPM82, VML93, WBR17, WBB93, YS98, Zdo09, dWvG88]. sediment-phosphorus [VPM82]. sediment-polluted [SVW95]. sediment-water [BNV03, RCB95]. sedimentary [GCS93, Van89]. Sedimentation [AGB94, ABF01, RCL94]. sediments [AH82, ABF01, BBF17, Boe86, BvdH90, BS95, CV95, Cha89, DHPN^{+}24, Das07, De 77a, FV95a, FSC06, GV99, GM93, HYH03, HWS22, Kow94, KIG06, Lij80, Lij86, MHE93, NLC06, OAS13, ÓP04, PJF01, RMC95, Sch95, SB88, SC99, SVK94, TFP06, VV93a, ŻKŻ10, ZGZ20, ZV93]. See [DH09]. Seed [BEQP18, HD21]. seedlings [LLX11]. seeking [RHS18]. Seeliger [Cad01]. seepage [MR81]. Segregation [dLFTB22, VvdMP12]. Seine [BGRM^{+}24, TLC93]. seizoenssuccessie [Gie71]. Sektie [Ano81a]. select [SSV21]. selecta [Kre82]. selected [EAO20, KK07, MNS05a, MNS05b, NR97, ODI08, RSM24, RHO09, SK09, VV93b, ZKN09b, ŻKŻ10]. selection [BDP21b, BH22, BTB09, BV84, Den06, GGD07, HBL13, IÓG04, IT09, LHV07, Nat05, NBT21, PCB10, RRL09, RB83, Sch92, VHtB17]. Selective [GTI99, ISY10a, ISY10b, PSP91, SOC12]. selectivity [RK02]. selengensis [xHCjX^{+}23]. Semboku [YO02]. semi [BV89, BKD^{+}24, DB02, DN23, FdLY^{+}24, dSLML21, dCMMB15, MdMdLB21, WSvA98]. semi-arid [BKD^{+}24, FdLY^{+}24, dCMMB15, MdMdLB21]. semi-empirical [WSvA98]. semi-enclosed [dSLML21]. semi-permanent [DB02]. semi-stagnant [BV89]. semi-subsistence [DN23]. semiaquatic [NCDS^{+}25]. semiarid [BSA20b, dSDMD20, FNS^{+}24a, FNS^{+}24b, MAD10]. seminal [EBA96]. sensing [Alt98, DDPT24, DDT24, JCF17, SD86]. sensitive [CCM21]. Sensitivity [DPL03, Clo99, LDK20, LOZ24, TP09, ZNP13]. sensor [BP94, HvdH84]. sensory [BT92]. separating [DB81]. separation [De 77a, KD82, SEB11, SS92, WXW11]. sequence [HEPH09]. sequences [FO98]. sequencing [DGN21, SPB^{+}24]. Serbia [PSSKŽ24a, PSSKŽ24b, ŽRP15]. Sergipe [NCDS^{+}25]. Series [RP08, JBB07, LHV87, RD77, SCS12]. serrulata [AKL15]. serrulatus [NS07]. service [NNJ22, SSN24]. services [SDW^{+}24]. sessile [REH92]. Seston [DA99, PBG03, ZSD94, BAG94, BGC00, GSM07, GSK08, TSZ10]. Setae [WCD17]. setigera [Baa88a]. Seto [KKS11]. sets [RJ11, WSvA98]. Setting [HG05, vLJA07]. Settlement [RVJ98, DG84, GLGU22, RNJ97]. settling [FDH94]. seven [CSA23]. several [KKM09a, LWdSD16, RR98, SH11]. severe [SVB97]. sewage [JTGS24, Kou83, TK88a]. sewage-enriched [TK88a]. sewer [RU87]. Sex [IDP19, dLFTB22]. Sexton [Pin75]. sexual [IIMRC21, IDP19, YYR20]. Sfax [GAAT^{+}24]. shade [LLX11]. Shading [KMR20]. shadows [CJS^{+}23]. shads [BFS21]. shag [GUT21]. Shallow [Moo98, OAS13, RP08, ÁBZ08, AGB94, BVI16, BM03, BM98, BHL09, BAG94, dSCdMM09, CCK10, CP10, CvNB02, CSV06, CCS20, CAFA04, DH09, DGB17, DBG13, FM10, FBS19, GVV86, Gri89, HvdH84, Hos89, HvB85, IM06, JVG95, JJG^{+}25, KK12a, KS94, Kro86, LPT22, LZH^{+}25, Lee82, LZN14, LJL05, LCL16, LZY21, MMRG03, Mer79, MAM13, MOdSCP07, MOPCP08, MDV03, NMB09, NLF10, PPV13, PDD07, PIM95, Ric86, RCB95, Roe96, RJ96, RFO94, SSI21, Sch89, SEE22, SFV09, TŠB17, TGZ09, TRP22, TG04, TK19a, VV91, Vij91, VGM93, VHB09, WBS22, WBS24, Wor90, ZPD06, Zdo09, ZWY^{+}22, Zev82, ŽAL24, ZPW13, ZPW15]. Shanghai [ZPW13, ZPW15]. Shangrila [WTC09]. Shannon [WBB93]. shape [FGF13, Fli85, JB83, NW24]. shaped [Kae20, VEK20]. shaping [BDF20, TWC13]. shared [DD17]. shark [BOB20, TRHHAR10]. sharks [GSH17, RMH21, SKH19]. shed [BB00]. shelduck [ACH09]. shelf [SDD^{+}23]. shell [Ban75, GJW20, GD09, SRF21, dSSR^{+}23, dBPG00]. shells [MSDDD22]. Shelter [RHS18, BKG16, SLS15]. Shelter-seeking [RHS18]. shelters [KJP14]. Shield [BTB09, PFDM99, PFMD00]. shift [DAO24, ZXH15]. Shifts [Bra21, BBD17, Gar05, HLK22, LRMdF19, SWF17, Vie20]. Shiga [WiNDF22]. ship [HVB97]. shipwrecks [dPBCP22]. Shira [PMD10, QJD02, RGD10, SKV02, Tol02, YTD02, AZD02, BGK02, BBM10, DBG02b, DG02, DZG10, GZG02, GBM10, KGS02, KKR02, LMP02, MKG03, TTD02, VKK02, ZTG10]. shoals [FGB20]. shock [BDP21b, Van79b]. shocks [Don79]. shoot [CDL14]. shore [DK80, FTH22, JTL09, Kuk92, RMH22]. shorebirds [FAB18]. Shorelines [Cad08]. shores [GFZ15, Mas88, NEP05a, NEP05b, SRI20]. Short [BAG94, CYK13, DGB17, KRJ09, PK08, RKS14, ŠV09, CRR22, PLC22, RFR91, SSV21, VKD77, WPvB17]. short-necked [SSV21]. Short-term [BAG94, CYK13, DGB17, KRJ09, PK08, RKS14, ŠV09, PLC22]. shotgun [CBdO^{+}24, SPB^{+}24]. Should [PGT20a]. shredder [CLRdSR23]. shredders [BBH05, CKM19, LB10]. shredding [ÅDB00, BT01]. shrimp [CCH11, KK09, LRG21, QLZ20, QZL^{+}24, Spa71, TYN02, VSB10]. shrimps [Alb04, KHK17, dCKL23]. shrubsolei [Baa88b]. shrubsoli [KP97]. Shunet [KBR10, DZG10, ZTG10]. Si [BZV93]. Siberia [DZG10, GKK92, KBR10, QJD02, SKV02, TSZ10, YTD02, DG02, KKR02, MKG03, PD02, RZG09].

Siberian [DeA10, ET16, GSM07, GSK08, GMD10, KKI04, KSW92, KKD02, KIG06]. Sicilian [DNG22]. Sicily [MS08]. Sicydiinae [ECR20]. Sicyopterus [ECR20, JBS05]. sides [RTW17]. sieboldii [HLSN24]. Sigara [DS75, Kre82]. signal [OLM09]. signature [EAO20]. signatures [Coa07, CSV06, FVW03, PFT12]. signatus [Lar92]. Significance [PFDM99, HPPL00, Spa71, VVK96, Vij91]. silent [RRK15]. Silesia [KMK09]. silicate [AvdV90, Van78a]. Silt [PBF06]. Siluridae [IRT06]. Siluriformes [IRT06, KKO08]. Silurus [WBS22, WBS24]. Silver [SGA06, DD99]. similar [FSL^{+}23, ST20]. similarity [Rav01, VB98]. Simon [Gul09b]. simple [Clo99, Kle84, LB14, PE78]. simulated [PA98, SON11]. simulation [BEQP18, KOA00, Nie82, Ver76b]. simulations [BV14, SS99, TNC08, dJG07]. simuliid [GEÓ04a]. since [GGT14]. sinensis [DTH20, JB83, RBAB24, WXW11]. Singapore [Lu05]. Single [RAP22, FO98, NBM98, NHiN15]. single-cell [NHiN15]. singular [Bra21]. sink [LME10]. sinking [MOCPG09, YSY^{+}22]. Sinonovacula [LLL^{+}23]. Siphonophores [SAACR09]. site [BV96, GMVCLGGP24, KGB04, MOdSCP07]. sites [AANO25, FDP01, NBT21, OTJ19, PJF01, PJ95, RFP01, SKRB09]. situ [Ber84, CBP22, NYT09b, TFT10, VPF21, APS81, Bir78, BHV82, MD05, PB84, PGSL02, TGE04, TZ05]. situated [HvAV83]. Six [RMS11, Bes87, GSB13, RFW17]. Six-year [RMS11]. sixty [Hig79]. Size [CBN94, DDV91, FDH94, LZN14, MMM11, Pri03, TFR91, VV07, Wil07, ZWY20, BJS22, BHG18b, BBM21, BW92, Cru10, DG84, ECL94, FBR87, GS11, GDT08, HCS02, HJ23, IDP19, Jar08, KEL09, KM17, LKM01, MLTH15, Nat05, NSB24, OB13, Pen00, PGT20b, PGT20a, RB08, RJ11, RW98, SL07, SOC12, TŠB17, VCT13, VSB10, WXZ12, WCT00, YSYZ23, YS98, ZYS14, BHG18a]. Size-assortative [Wil07]. Size-dependent [LZN14, Pri03, OB13]. size-fractionated [FBR87]. size-fractionation [LKM01]. Size-fractioned [ZWY20]. size-fractions [GDT08]. Size-related [MMM11]. size-selective [SOC12]. size-structured [Pri03]. Sized [VV07, BMO15, GBJM22, HH08]. sizes [WZL19]. Skealoghan [WMV09]. sketchy [Vos87]. skin [Rus92]. skipjack [WXZ12]. slabberi [MJ93]. sleeper [RPN15, ŠDT^{+}24a, ŠDT^{+}24b]. slope [CCM21, HC09]. sloped [FCS15]. Slotconclusies [Ano70d]. Slovenia [Bra99, Bra21, MB06, SB02]. slow [WAW92]. sluice [OI89]. Sm [ŻKŻ10]. small [AL24, Bau00, BTB09, Bra99, Cla84, CVM21, CMBG20, DCC09, GAA24, HH08, HPPL00, KKI04, KIG06, LPT22, LWO01, LFR75, Lin78, LRMdF19, MB14, NH92, OSM13, PdCF^{+}25, PFDM99, PCD11, QCF15, Rav01, SOC12, SR85, WHW14, dSGD17]. small- [SOC12]. small-regional-scale [CMBG20]. small-scale [Bau00, GAA24, LWO01]. smallmouth [GS11]. Smith [Baa88a, FvM03]. snail [AR05, ACH09, BA14, CSA23, GOB24, GD09, GRL20, HKT16, HÅB08, HBR12, MV14, OLM09, vOCRM19, RFW17, SMM23, SM12, SGD22, VvOT09]. snails [DD17, GFW24, LDK20, PMA22, RTW17, RAN24, TAN14]. Snake [ET16].

Snell [Shi98]. snow [JGG09, YSY^{+}22]. Social [CNN^{+}24, NNJ22, SSN24, vdE70]. Sociale [vdE70]. Society [Ano97b, Ano97c, Ano97d, Ano98e, Ano99a, Ano99c, Ano99d, Ano00d, Ano03b, Ano04d, GL88, Ano72i, Ano98h, Ano04f, Ano05c, Ano05d, Bel76a, BV04, Gel70]. socio [VPP16]. socio-ecological [VPP16]. soda [ÁBZ08, LKSK15, SFV09]. sodium [Dor77]. soft [HHV93, MQC93, Wor90]. soft-water [Wor90]. softwater [Roe96]. soil [CB13, DN23, OSR88]. soils [KCR92]. solar [CC94]. sole [VSM19a]. Solea [VSM19a]. solid [DHL16, QPP95, RHA00]. solid-phase [DHL16, QPP95]. solids [DM04]. Solms [RRRRA07, TK88b]. Solms-Laubach [RRRRA07]. soluble [Van78a]. solutes [LDW93]. solution [SGM22, TSV16]. Somatic [NR97, INJ20, TZD21]. Somatochlora [IS06]. Sombreiro [Nwa84]. Some [Bak78, Bel76b, BV84, BHV82, Cap75, De 79a, GD72, Gul76a, Gys72b, KHF01, NCT91, Pol75, RGD10, TGZ09, Van77a, VGV82, Bak94, Ban75, Bij75, BV76, BH82, CvDG07, CDW92, CC99, De 84b, DM71, Dvô70, Hig71, HNiN03, HH80, JC80, Kat92, KCR92, KBO74, MKA79, ME84, MO92, Rin80, RD77, TK89, THS17a, THS17b, Ver76a, WFH93, Web71b, ABV72, Beu74, dW71b]. sometimes [SPH16]. Somme [RDE93]. sonar [FWS^{+}23]. Songkhla [AR97]. Soorten [Ros69b]. Sorption [KSBR94]. sorting [VNSFG16]. Sound [BB23]. soundscape [KBN20]. sour [Mos97]. source [CCS09, GCL99, HJV08, KKS11, LTL09, LME10, PKE22a, PKE22b, SFA05, SON11, TK89, TC15, TG19, Van78a, YBZ20]. sources [AP12, BJ99, DPT04, MKG03, MWC94]. South [AANO25, PVC96, WB10, DWB10, HHV93, KML19, Kre82, Mom73, SB78, BSA24, BV76, Bra01, DZG10, DGGW20, DM71, Dvo69, HdC24, LM92, NP09, SM12, SMS^{+}24, SY02, SK04, TDHH23, TSZ10, WNZ24, Wel74, YKJ15, dSSR^{+}23, dLFTB22]. South-Central [TDHH23]. South-East [PVC96]. South-Eastern [WB10]. South-Finnish [SK04]. south-west [HHV93, Kre82]. south-western [SB78]. Southeast [MHYL23, FC93, NEP05a, NEP05b, WZL18]. Southeastern [GFZ15, ON88, KK07, PNC22, PMM10, CSV06, NO85, Nwa87]. Southern [DG06, FCPCMB^{+}25, FG91, GJ06, KMK09, PD02, SMGAGG00, SGGO06, WWŻ06, APS81, AR97, BdSCH09, DF83, EMN04, FAB18, GE80b, Kow94, KP97, MAS14, MS08, MSDDD22, MM95, MCV08, MQC93, NSK08, Pad13, PEREBM^{+}23, QPP95, SAACR09, UT16, VPS13, VH08, Zet97b, ŽRP16, AAR^{+}24, BC23, Df82, FD94, FN94, LMM05, MFS03, MB06, SdAdC^{+}25]. Southwest [Bak72, WLB22, ABA^{+}24, BTP24, CY95, PLM^{+}22, AKF^{+}23, LLL22]. sow [Ara01]. Sowinsky [PDvdV06a, PDvdV06b]. sp [BHB01, HKT16, MR11, RAP22, SNKN^{+}24, VvOT09]. sp.n [Mou92, Sæt92]. sp.nov [SKK92]. Spaak [Gos98a]. Spaar [Osk72]. Spaarbekkens [Osk72]. space [ESG04, MAM13, MKS22, SZY19]. Spain [DMD95, ES09, FBV02, PR92, dJG07, BD09, RFO94, SMGAGG00, Sol97]. span [CF09]. Spanish [RP92]. spanning [AKL15]. Sparganium [SMPK09].

Spartina [CF09, IIMRC21, LvBR06, SBS12]. Spatial [BM03, BSA24, BLC12, CM95a, CGL21, CCC19, dMCFS^{+}23, FSL^{+}23, GMW09, GJW20, HHV93, IJ06, JLG06, LPB07, LNHN08, MGL22, Mag00, ÖE04, RA07, SCO09, TS17, WTC09, Zdo09, BKD^{+}24, BNV03, BR07, Bes81, BNPC07, BHG16, BB23, DTD94, DSL12, FR09, Fli70, GS04, GM93, HD93, Irv89, KLG17, KKV92a, MDH93, MOCPG08, PF07, PGT20b, RdMJN22, RMW91, RAN24, SMGAGG00, SWF17, SAACR09, SEB11, SCC19, TWC13, TCH18, VIM17, WNS08, WBR17, YSYZ23, YBZ20, ZBB19]. spatially [BLB^{+}23, BC23]. Spatio [GC08, MSP13, MBFD24, NČŠ12]. Spatio-temporal [GC08, MSP13, MBFD24, NČŠ12]. Spatiotemporal [Hig09, RFW17, MAA22, dCNGF25, PLM^{+}22, RSM24]. spawning [BFS21, DAO24, SS24, WXW11]. speciaal [Mur69]. Special [Ano98g, Goe07, IFV16a, IFV16b, ŻD06, DeA10, De 84a, GMD10, HD93, HP06, IBV16, KBO74, KP97, Lam79, Mar01, MJJ21, iNTT05a, iNTT05b, PGT20a, SNG21, Sim94, TB22]. specialist [CDD20]. specialists [YF18]. Speciation [SNO98, De 84a, Mic92]. Species [Bes81, CDW92, ECGC25, GR91, MB14, PPCH22, Pip87, PVC96, RR03, VV07, dSCT23, AdAPC20, ARG20, Baa81, Bak80, Bak94, BV14, Bel87, Bes87, BZ97, BLP16, BSF11, BHG18a, BHG18b, Bri99, Bro84, CH91b, CBN94, ČN10, CSA23, CC99, DRM18, DTA22, DF83, EAK11, EDLC24, FKB92, GKK92, GSH17, GCN^{+}23, GJW20, GAA24, Gie71, GSM07, GZN20, GCL99, HEF13, Hey92, HLL^{+}25, HSY18, IUK22, IRT06, Jac92, JC83, KMR20, Kat92, Ker92, KSW92, KRATA20, KH92, KML19, KAY20, KÇA22, LHV07, LKSK15, MŚSJ16, MNS05a, MNS05b, MF92, MTdS^{+}24, MCB22, MOCPG09, MSD08, MCSV21, Mur74b, ME84, NR97, NEGS07, NNB^{+}24, NHiN15, NC20, Pin75, PR84, PHM21, RM00, RJD07, Ros92b, RFW17, SGRMP20, STTS18a, STTS18b, SWB13, SK09, SLM17]. species [SMS^{+}24, SOC12, SJ00b, SPS22, SZH20, TTD02, TBBS17, UET12, VNSFG16, VvdMP12, VIM17, VKD77, VGV94, WPvB17, WZL18, WT14, WHS13, WvRvdV03a, Wor90, XCL10, YAC21, YP17, YN08a, ZL07, ŽG09, ZNP13, dSGD17, dSSR^{+}23, dSSB24, dPNF91]. species-level [SGRMP20]. species-specific [ZL07]. Specific [Ker85, DLG20, Kir22, LDK20, Mar17, OMC22, WSvA98, WHW14, ZL07]. specificity [KPJP23, MNT16, PMVdC^{+}23]. spectra [BPT85, BWS21, RDG97, TSZ10]. Spectral [FAM03, Alt98, BTA94, Str88]. spectrometric [HK82]. spectrometry [BH82]. speed [SEE22, SR85, Wil07]. Sperm [FTP08]. Spermatophyta [Har82a]. spermatozoa [YRR09]. sphacelata [HD21]. Sphagnum [CBP22, HHN03]. Sphyrna [TRHHAR10]. spicatum [JJP24, JGH97, LZD16, SRG18]. spill [SDD^{+}23]. spine [SB97]. spined [BLC12, SHP21, VRR17]. spinifera [VPF21]. spinosa [ZL10]. spinulosa [DZL21]. spiny [BRA^{+}23, ZL10]. spiny-frog [ZL10]. Spionid [EK93]. Spionidae [BRJ97, BZ97, Boc97, BSB97, KP97, Sch97b, SVB97, Zet97b]. spiralis [LLG10, LWX19, SB14]. spleen [SXC^{+}25]. split [PV95]. SPME [RHA00]. sponge [GPS^{+}23, HvdGvS10, STG18]. Spongillidae [GPS^{+}23]. spores [Sch95]. Sporulation [Sch95]. spotted [MFTM20]. spp [BRJ97, BJS22, DD17, KML02, MOM24, PC14, PGM^{+}24, Sch95, VSB10, WF12]. spread [LvBR06]. Spring [IUK22, Baa85, Bak79, Ber24, BHL09, CBC99, DCC20, FG91, GC08, GD91, KPS09, MSW10, MAD10, Mou92, MNK04, NB22, RTW17, VET20, WZM^{+}23]. spring-dwelling [MSW10]. spring-fed [GC08]. spring-time [GD91]. springs [BD09, GMW09, HP06, MB06, NHF96, PKE22a, PKE22b, RCL94, RFW17]. squarrosum [HHN03]. squid [FSL^{+}23]. Sri [DGH^{+}24, JGUF23]. St [FAM03, VH08]. St-Lawrence [FAM03]. stability [BGL13, CL11, FSC06]. stabilizing [SCEGL12]. Stable [AMTSJ15, FVW03, SFB04, WRS05, WVM18, dCdCP19, Aag92, BPD15, CCS09, GB07, HF08, KSY19, KTJ10, KT15, MNJ21, PFT12, Sch89, UOJGRL21, VP12, XYH11]. Stackhouse [SY02]. stage [Bes79, DD17, WMM11]. stages [BdPP17, FGM20, JBS05, KNH07, MKA79, SS92, VPS21]. Staghorn [GLGU22]. stagnalis [VvOT09]. stagnant [BD74, BV89, Vaa79, vdW70]. stakeholders [VPP16]. stand [dL74]. standardize [YB08]. standing [PMM10]. starfish [Rei99]. starry [HLK22]. stars [SRI20]. start [GTPH08]. started [Kae20]. starvation [IDP19]. stasipatric [Mic92]. State [AAA^{+}24, FBNF08, DB85, DGH^{+}24, DG84, GDB19, PPV13, PAP24, RJ96, VK12, WWS21]. states [NCDS^{+}25, NvZvG16, Sch89, CEH12]. station [De 79b, Wan79]. stations [CP10, Had79, HvAV83]. statistic [SJ00b]. Statistical [TGW06, GM93, SCC19]. status [BPP21, CTO01, CSM08, DGH^{+}24, EAB08, HHEM19, KMK09, MCSV21, PDE08, ŠV09, VH89, VM25]. statuses [CDL14]. Staurastrum [Coe97]. Staurozoa [MTdS^{+}24]. stay [OR06]. steady [DB85]. steady-state [DB85]. Steentoft [SY02]. steeply [FCS15]. Steinberg [Gos98b]. Stel [Nie98]. stem [YN08b]. stems [Roo79, Roo82]. Stenella [MFTM20]. step [KA86, VHC92]. steppe [KRZ03]. steps [GTPH08]. stepwise [GGD07]. sterilised [GBA20]. sterols [MCvE09]. Steryl [Kow94]. Stichopus [PAAJM17]. stickleback [VvdMP12]. sticklebacks [BM98, VRR17]. Still [HKB19, OR06]. stilstaand [vdW70]. stimpsoni [JBS05]. stimulated [KG03]. stimulation [DZT13, RMR21]. stimuli [VPF21, vGR98]. stochastic [SCC19]. stock [Gud04]. stocking [Bra21, DTC19, Gri82]. stoichiometric [EDLC24, ZD21]. Stoichiometry [ACE10, GC23, QLZ20, SAB10]. stolons [SXY20]. Stomach [TRHHAR10, MMM11, UOJGRL21, dCdCP19]. stomatal [QZL^{+}24]. stonewort [HLK22]. stony [Ros92b]. Storage [Zoe72, Cor78a, De 76a, HvB85, KKV92a, Klo76, KKV92b, MV83, Rod10, RB78]. stormwater [MvAV20, RU87]. story [RVJ06]. Strain [NMS21, KZV02, MCB22, PMZMJ16, SBH17]. Strain-related [NMS21]. strains [KA97, PLdFC22, SH11]. Strait [LAC99, Lu05]. stranding [AL24]. Strategies [DCL19, CCC19, DHW81, FBV02, GA10, HSP16, MBM15, MCV08, QZL^{+}24, RHH17, VCT13]. Strategy [Lin07, FSL^{+}23, LW11, PFT12, SK16]. Stratification [Van78b, BdSCH09, BBM10, DZG10, GBM10, HLL14, RGD10, TGW06, VK80]. stratified [BBM21, GH06, KIJ11, KKR02, KKD02]. stratigraphic [FDP01]. Stratiotes [MMV05, MMV09, RIT04]. stratiotetum [Dre68, Hig69]. Straus [Ker91]. Stream [Bra01, RBPF11, SAB05, WD00, AD89, BLB^{+}23, BDF20, BRP15, BHPT20, CCM21, CAPGA08, CKM19, CYK13, DWB10, DGD06, DEP22, ECB09, FGF13, FM99, GC08, Hig09, HM84, HW13, Hof99, IGS10, Kim99, KM17, Kou83, KCG05, LDL11, Mag00, MAD10, MMC14, NH92, OSM13, PBM21, PF07, Rav01, SL07, SLM17, SMC21, SGJ^{+}24, SKM05a, SKM05b, TDB13, TBSZ22, TH16, VALRdF20, WPB21, WPB08, WHW14, YL08, dNdSBJdCeS21, dRFC21]. stream-fish [Hig09, TH16]. streams [AKF^{+}23, AAA^{+}24, AFC20, CTO01, CPA15, DWB10, FCG16, FWB05, GKM22, GL01, HH08, IAHB18, IJ06, JM08, JBS05, KK07, LB10, MB14, MRR10, dCNGF25, OMC22, PSW11, PRW08, PVV07, RB08, RASA13, RSK09, RJD07, Ros92b, Sæt92, SCEGL12, SEB11, SK04, TAN14, Tol82, TLH18, Van79a, VS96b, VK12, WB10, ZSG14]. Streblospio [KP97]. Strengeriana [RPSSDS23]. strength [SRI20]. strengths [LMV09, SCS12]. Streptophyta [SH11]. Stress [Flo01, RC98, BWS19, Bro98, CGD19, GYW17, INJ20, LSY20, MGT17, QLW^{+}23, SCN^{+}24, SVB98, TG19, Van97, VB98, WCD17, WFBJ17, ZNP13]. stressed [KM83, NSB24, NudWB^{+}25, QSA09]. stressor [PCGW20]. stressors [GL19, LPT21, LBKV18]. striata [DS75, MM96]. Strong [dNFdNM21, WBS22, WBS24]. Structural [AT24, Bes79, DDF93, DGBL15, FGF13, HSY18, IAHB18, PLM^{+}22, SH16]. Structure [FNS^{+}24b, Gul89, KKD02, UKS05, AP12, AKL15, AOS^{+}24, BB00, Bak78, BCN12, Bol07, BV96, CGC19, CMH14, CMV91, dMCFS^{+}23, Cru10, CM95b, De 97, DF83, ED95, FBS19, FGM20, GZG02, GFF17, GPGSMM^{+}23a, GPGSMM^{+}23b, GSG24, GFW24, HBO03, HHV93, HJ23, HLK22, HHA92, Her83, HV89, Irv86, JVG95, KMG^{+}23, LSO10, LLL22, MFTM20, MBMV20, MTSJA19, NCT91, NČŠ12, NMdAM21, NAM19, NHL18, ÓP04, OLJB20, PDP02, PIM95, PVV07, PHD13, QPP95, RB08, RdMJN22, SCEGL12, SEE22, SMJ95, SYW09, SK04, SBS12, TŠB17, TGZ09, TK19a, Tok92, TBSZ22, TH16, Van82a, Van91, VL18, VHB09, VSW00, WWS21, WP12, YYS^{+}23, YHM05, ZPD06, Zdo09, dSSB24, dCAP23, dRFC21, FNS^{+}24a]. structure-activity [Her83]. structured [Pri03, VCT13]. structures [BSF11, WXZ12, WHW14, ZWY^{+}22]. structuring [EMN04, MTA15, MMRG03, SBS12, WvLA14, ŻKŻ10]. studie [Per69]. studied [BV14, FO98, ÖE04, VVB82]. Studies [Ano83b, BZD82, Cad01, Gul76b, SFB04, Vaa79, YZ08, AFRS23, Baa79, BHV82, DL89, FAC16, Flo01, GS11, GPV19, HBO03, Her83, HBR12, Koo76, NS21, ONA^{+}23, PB84, VW78, Van85, VGM93, dJG07, vD83]. Study [GWF92, AKF^{+}23, AAA^{+}24, AAV07, BVM15, BBV78, BV76, CZC14, CBE20, CSA23, CW04, DSN09, DDC20, DB02, DWV81, DNG22, Den94, DDPT24, DD99, Dre68, DM71, DMD95, EBM18, EBA96, FA18, FMdCFdCV24, GFSN07, GAA24, GX11, HH08, HEZ03, HLL^{+}25, HHC21, JGUF23, JCF17, KK09, Ker70, KKO08, LPT22, Lee82, MHYL23, MSW10, MSB10, MvAV20, MML^{+}22, MML^{+}23, Nan00, NEGS07, Nie78, Nie79b, OAS13, PC14, Per69, PJF01, PJH06, RMS11, Rin81a, RKLB22, RKLB23, SWW02, SEB11, SKV02, SK04, SdAdC^{+}25, TBA03, TGE04, TBF99, TIF10, VK90, VVW95, VP77, VMS09, YTD02, ZWY^{+}22, Bil73]. Studying [Boe86, DDF93, Kir22, YB08]. sturgeon [WXW11]. stygius [MSB10]. sub [BBD17, GA10, HAS22, PPCH22, PC14]. sub-adult [PC14]. sub-alpine [BBD17]. sub-Arctic [GA10]. sub-tropical [HAS22, PPCH22]. subadult [RMH21]. subalpine [LPT21, PBG03]. subarctic [AK09, GS04, Pri03, TGE04]. subboreal [Den94]. subfossil [BBD17, CZC14]. subhabitats [EKK21]. subject [RMC95, ZPD06]. subjected [GB07, ST09]. Sublethal [GK07]. Sublittoral [SVW95, Waa80]. Submerged [MdMdLB21, Bes81, Bes87, DZL21, DDY^{+}23, JC83, KMR20, KT09, KM17, Kor14, LLG10, LZD16, LBY17, MLL21, MS02, MMRG03, Per69, PDD07, PCD11, PR84, RPS20, SAF10, TM84, UET12, VNKdOR15, VV91, WVM18, XWJ22]. submergence [DZG17, LCL16, VGJ20]. submersed [JC80, JGH97, Kor98]. Subscribers [Ano72c, Ano70b]. subsistence [DN23]. subspecies [PBDJ20]. substances [CGO19, DCG19]. substrata [ÁBZ08, Cla87, DRN09, FCS15, NRB21, SGM99]. Substrate [BCP10, BHPT20, DLG20, ERA17, KKM09b, MKT23, NC20, RB08, TG19, dPBCP22]. substrate-specific [DLG20]. substraten [Per69]. substrates [BG15, GC23, HB75, HV79, HSY18, JBB07, MS08, Per69, TET23, Waa80]. substratum [OR06]. Subsurface [RDG97]. ’subtidal [DR96, DR96, MHE93, RTJ11]. subtle [De 79c]. subtropical [AMTSJ15, BdPP17, BdSCH09, dFBdGGMLT21, CM05, CCK10, CL11, CMM05, CCS20, DCC09, FBS19, FGM20, GSH17, JTL09, KK07, MHYL23, MMRG03, MTSJA19, Rod10, SZH21, SPH16, WCZ11, WXL16, YL08, YL11, ZYL22]. suburban [TAN14]. success [GS11, GKM22, Ker97, Lam89, Pen00, RVJ06, RMH22, RTL17, SAR21, SPA12, TNC08]. Successful [VCT13, CW04, MMBP21]. Succession [CZZ^{+}25, CMM05, Gra12, Baa85, Gie71, HEPH09, Lov74, RNJ97, RCV91, RD77, Van84b, YKU01, ZPW13]. successive [YN08b, vGR98]. such [DMN88, NNB^{+}24]. Suez [GE80b]. sugar [KMS82a]. Suitability [VPS13, Coe77, Van74, WXW11]. suitable [Gie71, YO02]. suivant [Dar73]. sulfate [KRZ03, LKSK15, PKR12]. sulfate- [LKSK15]. Sulfidevorming [Vos71]. sulphate [KCR92]. sulphide [BRB97, SVB97]. Sulphur [DG06, SGA06, ŻD06, Sch77, SGGO06, Żur06a, Żur06b]. sulphur-cycle [Sch77]. sulphurea [PBF06]. summaries [Ano83b, Ano70c]. Summary [BM70, Vie20]. summer [BBM10, DGG15, FR09, Gil20, MFL09, RFR91, RvGB05a, RvGB05b, SLC10, SGR86, TE04, VGJ20, WBS22, WBS24, XWJ22, YHM05, ZQH16, Hal76]. sun [dSCT23]. Sundarbans [CMH14]. sunfish [EBM18]. sunlight [BWS19]. supplant [HGB20]. supplementation [ZD21]. supplies [KFS04]. Supply [DV92, Ver72, Ano72b, EG04, iNTT05a, iNTT05b, SBH17, TIF10]. Support [MLL21, AVD07, MdMdLB21, SCG09]. supports [TZD21]. suratensis [JDJ^{+}25]. surf [MKV15]. Surface [SGGO06, BGRM^{+}24, BB23, Fon71, GBJM22, Hos80, KK07, RMC95, SEE22, VHC92, ZGZ20]. surficial [GV99]. Surinam [Van78b]. Suriname [VGV82]. surrounding [MSRB23, ZPW15]. Survey [De 76b, Bra21, Eng74, Har73b, KJF00, LB14, Van77a, WH92, WW84]. Survival [CBNF19, PBF06, ACB12, BHM03, DF11, Don79, FV74, GLGU22, JBS05, KC22, KK12b, KC82, MJEC05, Nwa95, PWG05a, PWG05b, RHO09, TS17, WPvB17, WV02, YN08a, ZLT22, ZKN09a]. survivorship [SMC08]. Suspended [Alt98, RFO94, BTA94, BKW94, CN94, CLH06, ECL94, GCS93, PS06, PN94, VdJ94, WS94, WGK95]. suspension [vdTS97]. Süsswasser [HD79b]. Sustainability [Flo01]. Sustainable [Sar05e, Sar05f, FBV02]. sustained [Wet93]. Sutcliffe [Laa97]. SW [ES09, WFH93, Bak77, Bak78, BZV93, Nie79b, SLL09, vdTS97]. SW-Netherlands [Bak78, BZV93]. SW. [Vaa75]. swamp [Cha89, EAB08, HKB19, MLG20]. Swarming [Lin07, Ker91]. Sweden [NEP05a, NEP05b, SMR08, Wey09]. Swedish [Hof99, PIM95]. sweet [Mos97]. swept [MIMRRZ08]. swim [OR06]. Swimming [NPH03, DTH20, EBM18, lCdB70, vGR98]. switching [Zad03]. Sylt [Sch95]. sylvatica [GK07, LB20]. sylvaticus [CBNF19]. symmetry [NW24]. sympatric [KJS15, dSSR^{+}23]. Symposia [Ano74b, Ano75b, Ano76c]. Symposium [Gel70, Gel70, HB88, Ano87]. Syndiniales [ZTNW21]. Synodontis [KKO08]. synopsis [BVI16, IBV16]. synthesis [SRH08]. Synuraceae [Roi81b]. System [RRRRA07, STT01, ABA^{+}24, AMTSJ15, AC86, Baa80, BSA14, Boe86, Bro81, BF12, DWA11, DEP22, GD91, Hig80, HAS22, Kal86, LSY20, MLTV18, MQC93, NHF96, QLW^{+}23, Rin73, RTJ11, SD09, SAACR09, SvFN15, SLH87, TVD91, Vos82]. systematic [GPV19, Hey92, PLM18]. Systematics [Pol73]. Systematiek [Pol73]. Systems [Nie98, BH22, BT92, CHF09, Eng86, GV99, JJG^{+}25, KSY19, Mos97, ONA^{+}23, Rat17, RDG86, Sch89, SAR21, VPP16, vdV81].

T [Cad99, Gul99, Shi98]. table [GFSN07, ZXZ09]. Tadorna [ACH09]. Tadpole [Jar08, CB13, GBPR23]. tadpoles [BdSM20, KGM13, KGM22, LWW^{+}25, STV06, SG20, ZWJ19, dNdSBJdCeS21]. tagged [MMC14]. tagging [HHP12]. Tagus [BC95, CBC99, CM95a, CCS99, CC99, DAR99, FED95, FA99, LA99, Mor99, PN94]. Taihu [CDL13, LSO10, SLC10, CCK10, CZZ^{+}25, DZT13, TDL23, WBR17, YSYZ23, ZGZ20, ZQH16]. Tail [SB97]. tailings [PdCF^{+}25, dSSSES21]. tailwater [GBT19]. Taiwan [CYK13, YRR09]. taking [CBdO^{+}24]. tale [PMB^{+}24, PBDJ20, Sch83]. tallow [MSF21]. Tamar [DW93, FDH94, MJ93, WFH93]. Tamaru [Shi98]. Tamminen [Cad99]. Tana [WMD07, SNO98]. tannins [HvdGvS10]. Tanytarsus [IÓG04]. Tanzania [MMBV24, MOM24]. Tapovan [IUK22].

Tardigrada [BP21, ZBA19]. target [PPCH22]. Tawang [BN24]. Taxa [DDD07, HYH03, IIMRC21, KTH14, MGT17, RB08, Ver79, WVM18]. taxifolia [DNG22]. Taxonomic [CAPGA08, RRS^{+}22, RRS^{+}23, TH16, Bor73, DEP22, Hig09, MTdS^{+}24, ODI08, OB13, Roi88a, TBSZ22, WB10]. Taxonomical [MKG20]. Taxonomy [PVC96, PFSPdS24]. Tay [SVB97]. te [Gel70]. Technical [Gul85]. technique [BHV82, FK81]. Techniques [BGK85, AVD07, SD86, YB08]. Technology [Hod98]. teledetection [GEK97]. telemetry [HHP12]. Teleostei [ECR20, FTP08, KBO74]. tell [DMM97]. telson [MFS95]. Tembladera [LBKV18]. temminckii [YKJ12]. Temnopleuridae [GBZ21]. Temora [BTV88, BV89, FG91]. temperate [Bar21, CH12, CSV19, DH09, DGG15, Gil20, GC08, MBM^{+}22, MS11, MSD08, RGF21, VHB09, ZXH15]. Temperature [BDG06, HVB97, Lin72, SvFN15, SSK14, AOLLJ24, BB12, BRP15, BA07, CCS99, CM05, CBP22, Cru10, DWA11, De 79c, Dor74, Dor79, FV74, HH08, HGR11, HvdH84, IUK22, JM08, Ker78b, Ker81, KS79, LMV09, LKSK15, MR17, MH13, NW24, Nor98, RVK96, Rao10, RBPF11, Rin70b, Rin73, RTL17, SSR13, SC99, SFV09, SH11, TBA03, TGD16, Van79b, VSM19a, WvRvdV03a, YZ08, dW71a]. temperature- [LKSK15]. temperatures [Gil20, IGS10, LvBR06, RMC95]. temperatuur [dW71a]. temperatuuroptimum [Rin70b]. Temporal [BKD^{+}24, BBH05, CSK10, CL11, DRM18, DTD94, HLL14, HJ23, HYH03, JTL09, JM84, KKV92a, The02, WBR17, YL08, YBZ20, BSA24, BR07, BB23, BHPT20, CM95a, GC08, HD93, KD82, LPB07, LV93, Mag00, MKV15, MSP13, MBFD24, NČŠ12, ÖE04, RA07, SCO09, TDP06, VIM17, YSYZ23, Zdo09]. Temporary [SJ00a, BST89, CMBG20, DB02, ES09, FWR^{+}23a, FWR^{+}23b, FGM20, GRDPP09, Lah98, MMBP21, NJ00, OR11, PVB20, QCF19, SVM08, SJ00b, VC89]. Ten [Bak72, Gul07b]. Ten-Year-Old [Bak72]. Tenagomysis [PC14]. tench [BM98]. tendencies [BV78]. tendens [Mic92]. Tenerife [RMS11]. Tennessee [KGB04]. ter [Pee71]. term [BBM09, BAG94, CYK13, DDF93, De 79c, DH09, DGB17, DÖD21, EÖ04, xHCjX^{+}23, JCS01, JCF17, KRJ09, Mar01, MKV15, NEP05a, NEP05b, PLC22, PK08, RKS14, SN03, ŠV09, VSW00]. Tern [GKM17]. Terpsinoe [Ste94]. Terrestrial [Del92, PBR01, KTH14, RMR21]. territory [ŽRP15]. test [Beu71b, COC18, PVB20, Pod74a, Pod74b, Pod74c, SJ00b, TZ05]. Testate [Pay10, ASLT15]. Testing [HKH^{+}23, TG19, CHP06, ODI08, QPP95, VHtB17]. tests [Blo83]. testudinum [HSCR10]. Tetrahymena [SFA05]. teweeg [Ros69b]. Texas [Hig09, MH13, NBD^{+}25, TH16]. Texel [DD84, Pol75]. textile [CMBI25]. Thailand [AR97, SAB05]. Thalassia [HSCR10]. Thale [AR97]. Thau [CVM21]. Thecamoeba [RD77]. Their [Sha78, BGRM^{+}24, BTV88, Bas20, ECB09, EAO20, EG04, FGL19, Hig80, Hou94, HWS22, Hui88, IBV16, JTGS24, KNH07, LPT22, LMBM12, LWX19, MAS14, MVI16, MAA22, MKA79, NLF10, Nwa95, OSM13, PDD07, SAACR09, STHN03, SGD22, SVW95, SRH08, TPS09, VW78, Van87, VSB10, WPvB17]. them [Gie71, RHO99]. Theodoxus [VM25]. Theoretical [Ald78]. theories [Nie94]. theory [Deg10, HKJ12]. Thermal [KPV16, VK80, DRM18, DSG^{+}24, De 79b, INJ20, Kim99, SSK14, TGZ09, TGW06, Wan79, Zdo09, ZL10]. Thermal-dependent [KPV16]. thermalis [VGP^{+}23]. thermally [GH06]. Thermische [dH71b]. thermocline [PFDM99]. thermocyclopoides [RK02]. Thermocyclops [BNPC07, CB00b]. these [Coe77, SCEGL12]. thickening [WCD17]. thienemanni [Dev09]. thin [KHF01]. thin-lipped [KHF01]. Thiobacillus [tH71]. thomasiana [NW24]. Thompson [Hod98]. Thorea [HL95]. those [SAR21]. thought [MBM^{+}22]. Thoughts [Sch88]. thraustochytrids [DDC20]. thread [OR06, RVK96]. threat [CBNF19, OCS20]. threatened [ECB09, PPK^{+}23]. three [ABF21, BB12, dFBdGGMLT21, BVD82, CZC14, CV95, DB85, DDF93, Den74, DTA22, EAK11, GKK92, GSH17, GV99, HYH03, JC83, KK17, KKM09a, KHF01, LHV07, MT09, NEGS07, NRB21, PU15, QLZ20, RJ11, RCSF93, SMPK09, SEE22, TNC08, VRR17, YAC21, ZNP13, ZXZ09, XCL10]. three-compartment [DB85]. three-dimensional [TNC08]. Three-Gorge [XCL10]. three-spined [VRR17]. Threshold [LAZ21, HAB98, SRH08, TFP06, TDP06]. throughout [Goo79, RS85, ŻKŻ10]. throughput [DGN21]. Tiber [CSK10]. Ticino [NBT21]. Tidal [BKW94, ECL94, FV95a, SHV96, Zim78, BD74, Bak77, Cad78, CFP08, CDA03, Dro84, FDH94, FN94, FCW94, Goo79, HVK93, HZC95, LME10, Met78, Møh99, PN94, Ran74, RHL02, SHS93, Sha78, Van78a, Van82b, VD87, vG84]. tidally [TDP06]. tidally-driven [TDP06]. tide [GXY14, ZZL16]. tide-forming [GXY14]. Tides [Cad05e]. Tien [Bak72]. Tienhoven [Bel76a, Bel76b, De 76b, Mol76, NRVV76, Van76]. Tiger [GK07]. Tigray [vGVV11]. tigrina [Van75a]. tigrinum [GK07, GK07]. tigrinus [Pin75]. Tijdschrift [Dor72]. Tilapia [FASdPN^{+}25, SASA^{+}24, AIS02, ST09, MOM24]. Time [LHV87, RP08, WKK24, CMG95, CKM19, ESG04, FDG97, GFSN07, GGT14, GD91, HGR11, JBB07, Kle84, MKS22, SCS12, Vij91]. time-dependent [GFSN07]. time-saving [Kle84]. time-series [SCS12]. times [Baa81]. timing [HKJ12]. Timsah [GE80a]. Tinca [BM98]. tiny [OB13]. Tioman [NYT09a]. tissue [KFS04, TKR10, WA84]. tissues [PLC22]. Tjeukemeer [vD83, DWV81, Gul76b]. toad [DH22, ZWJ19]. Toekomstige [Ver72]. Tolerance [Den74, MV14, AND25, Dor77, GH04, KNH07, Rin73, WvRvdV03a, ZKN09b, ZL10]. tolerances [HVB97, ODI08, RTW17]. Tolo [WKH82, WA84]. Tomotoshi [Cad05e]. tonsa [BTV88]. Tool [SFB04, CBF19, Cla87, DGD06, KSY19, MG06, MFP01, Pos80]. tools [Gri89, PDD07, SCC19]. tooth [PBDJ20]. Top [BBM21, DdNM22, VV05, WT14, AVdNM19, DGBL15, dNFdNM21, FBS19]. Top-Down [VV05, BBM21, DdNM22, WT14, AVdNM19, DGBL15, dNFdNM21, FBS19]. Topics [Ano98g, Zet97a]. topmouth [SAM07]. tot [Dor70b]. Total [CV95, Ald79, LZN14, MRR10, NC20, SCN^{+}24]. tourism [HSCR10, PPY^{+}22]. Townes [SS92]. Toxic [Kap82, dSFFA03, GCL99, GH04, HVB97, HLL14, HZZ22, MCB22, SBH17, VGV94, YP17]. toxicant [Kra95]. toxicants [Ker75, Ker83]. toxicite [Ade71]. Toxicity [MRBQ^{+}25, VV07, Ade71, CEH12, DF90, DDZ^{+}24, GFSN04, GFSN07, Her83, MFL09, QPP95, Rin70a, VRH^{+}24, Wil70]. toxins [KK12a]. Trace [GM93, SGA06, TVD91, WHP76, De 76a, MWC94, RAP22, WFH93, ZV93]. tracers [Gul85]. Trade [TBBS17, INJ20, JDF23]. Trade-offs [TBBS17, INJ20]. traded [JDF23]. traditional [IRT06]. training [RJ11]. trait [CCM21, DSG^{+}24, OMC22, PPV13, PKE22a, PKE22b, VN22]. trait-based [CCM21, PPV13, VN22]. Traits [RR03, AKF^{+}23, AAA^{+}24, DD17, De 97, EAO20, EBM18, EDLC24, FBL13, GR08, HdC24, KTH14, MVI16, MCC21, QLZ20, QZL^{+}24, RG07, RC98, STTS18a, STTS18b, TCH18, WL24, YF18, ZWJ19]. tramoserica [FCS15]. trans [NGSOAD20]. trans-Mexican [NGSOAD20]. transboundary [CFRNT21]. transcriptional [SXC^{+}25]. transect [WMV09]. Transfer [RHO99, BT01, FdLY^{+}24, JLS09, Pod74a, Pod74b, Pod74c]. transferase [GX11]. transfers [GAAT^{+}24]. transfert [Pod74a, Pod74b, Pod74c]. transformation [ECR20]. transformations [Gil12]. transition [Ran74, RAA98, dRFC21]. transitional [CBF19, VNSFG16]. transmission [DBO11]. transmitters [MMC14]. transparency [WMD07, ZZL16]. Transplanting [RSK09]. Transport [Pus94, Wol79, FCW94, FA99, KHM04, LDW93, ONA^{+}23, PN94, SvFN15, SVB98, TLC93, Vos82, WMM94]. transportation [YKJ12]. Transports [Dan84]. traps [ÖE04]. treat [LZL22]. treating [FAC16]. treatment [HVB97, Kou83, SMPK09]. Tree [DDD07, AdAPC20, Dev09, KGM13, KGM22, PT14]. tree-hole [Dev09]. trematode [VvOT09]. trend [Mar01]. trend-monitoring [Mar01]. Trends [NMP12, BB23, MKV15]. trends* [Bra01]. Triadica [MSF21]. triandroides [DZG17]. tributaries [BFB15, ŽG09]. tributary [NWL11, RPN15, SB04]. tribute [Gul07a]. Trichodesmium [WYW^{+}24]. Trichoptera [HH08, Hig68, Hig69, Hof99, MO92, VC89]. Trichostegia [VC89]. triclad [Van75a]. trifoliata [Har04]. trigger [LBKV18]. Triggers [PdE^{+}24]. triphosphate [Koo76]. triplefin [PFLO14]. trituberculatus [LLL^{+}23]. Trontelj [ŽRP16]. TROPH [AMTSJ15]. Trophic [Alb04, Art99, CH12, DMD19, FAB18, FDB97, MNJ21, PdCF^{+}25, SB88, dCAP23, AMTSJ15, dSDLdA22, DGH^{+}24, DTA22, DdNM22, DG84, Hal76, HXR09, KMK09, LSO10, MCSV21, OR11, PPV13, PAP24, RJ96, RNJB20, SWW02, SNO98, SBLdAM23, TFR91, TSZ10, VEK20, WWS21, WNZ24, ZPD06, ZYL22, dSGD17]. trophic-level [HXR09]. trophically [Dor87]. Tropical [LLL22, Sar05d, AVdNM19, ACL13, AGB94, AKL15, AXR15, Baa80, BSA24, BNPC07, BASJ94, BAG94, BFB15, BSA20b, BA07, CGC19, CKM19, CNA07, Cha87, CH12, CSV06, dMCFS^{+}23, dSDLdA22, DHPN^{+}24, DSL12, FVW03, dSFFA03, FAS05a, FAS05b, FAL07, dNFdNM21, FMdCFdCV24, GCE11, GANA00, HJ23, HEF13, HAS22, IRT06, JGUF23, KBN20, LBKV18, MLTV18, MIMRRZ08, MdMdLB21, MSD08, NYT09a, NSB24, NudWB^{+}25, PPCH22, PAP24, QSA09, RGF21, RASA13, SDD^{+}23, SWF17, SCN^{+}24, STTS18a, STTS18b, SDN25, SCC19, TK19a, VN22, WMD07, dCKL23, dCAP23]. trout [AK09, BTB09, GS04, HKH^{+}23, HFS08, Kae20, LWO01, SH16, VK12]. trutta [AK09, GS04, HFS08]. tshawytscha [KAN20]. Tubastraea [dSCT23]. tube [Hou94, ÓP04, Hig68]. tube-building [ÓP04]. tube-dwelling [Hou94]. tubes [CBdO^{+}24]. tuna [WXZ12]. tundra [FLP08]. tunes [GBPR23]. Tunisia [GAAT^{+}24, KHF01, DTG09]. Turbellaria [Van75a]. Turbid [RP08, BGC00, GEK97, GB82, HEPH09, LJL05, MOPCP08, MDV03, Soi05a, Soi05b, SFV09]. turbidity [BKW94, CH91b, DW93, FGB20, GKK95, HBL13, KSF95, LNHN08, Mar93]. turbulence [CLH06, EAK11, Gre05, ZQH16]. turion [WYZ13]. Turkey [Kül05, SEE22]. Turkish [YAC21]. Turlough [WMV09]. turnover [MR17]. Turp [Ste76]. turtle [DDZ^{+}24, JDF23]. turtlegrass [HSCR10]. turtles [AOLLJ24, NPH03, SSV21]. tussen [Gie71]. twee [Gie71]. Twenty [NDW96]. Twenty-five [NDW96]. Two [HEF13, MK05a, PF07, RR03, Rin97, ARP14, AKF^{+}23, BV78, BGL13, BZ97, BSF11, Bri99, BLC12, CBN94, ČN10, CLL10, CW04, DRM18, DD84, DZG10, Dor76a, Dor87, DMD95, DF83, Df82, EDLC24, FGL19, GMD19, GJW20, GB07, Gie71, GB82, GKK95, Gul89, HVS87, JGUF23, KKI04, KMR20, KEL09, KA97, KK89, KKV92a, KIJ11, KKM09b, KJS15, KRATA20, KML19, KKV92b, LPB07, LLX11, MOdSCP07, MOPCP08, MB06, MDV03, NLF10, NvZvG16, NC20, PPV13, PGV95, PMA22, PBDJ20, PCD11, Pol75, PHM21, PJ95, RDG86, RJD07, RZG09, Roi81a, RAP22, RHH17, STTS18a, STTS18b, SWB13, SR92, Slu81, STV06, SLC10, ŠNZ21, TBBS17, TSZ10, TZ05, Van87, VGG89, VvdMP12, WPvB17, WT14, WLN12, Wor90, YL11, ZCZ14, dSSR^{+}23, dPNF91]. two-compartment [TZ05]. two-spined [BLC12]. type [AYM09, GIK21, LCPM15, LBY17, PMZMJ16, SGRMP20, YKJ15]. types [CSM08, HM84, KAY20, LZD16, NvZvG16, PPC08, VHC92, Ros69b]. Typha [GGT14, SGM22]. typical [HMB88, JCF17, WWG^{+}25]. typological [VH89]. typology [Aag92, BvdH90, Coe75, M.75, SR88]. Tyrrhenian [ABT21, MS08].

U [Cad01, Gul09b, Nie98, Hol70]. U.K. [RB78]. U.S.A. [Sæt92, SR92]. U.V. [Sch78]. Uit [Ano68e, Ano69b, Ano69c, Ano69d, Ano70e, Ano72f, Dor71, GR72]. uiversnest [GB69]. UK [RMW91, CB16, SMR08]. ultraoligotrophic [CMB07, LM92]. Ultrastructural [YRR09, HPT95]. ultrastructure [WL24]. ultraviolet [De 00, DGB17, HPS11, KT09]. ultraviolet-B [KT09]. Ulva [GXY14, KML02, MR11, RSM24, WKH82]. ulvae [ACH09]. Ulvophyceae [MR11, PVC96, SYW09]. umbellata [HHEM19]. Umm [AAMS24]. unattached [SY02]. unavoidably [CRR22]. uncertainties [BWS21]. uncharted [GBJM22]. unconditioned [PA98]. under-ice [JGG09, KPS09, VS09]. Underestimation [FBR87]. underground [RCL94]. underpin [MMBP21]. understand [BGRM^{+}24]. Understanding [BAdOFC21, DG02, DW14, MVI16, VN22]. Underwater [EHM91, RP08, WZM^{+}23]. Unexpected [OR11]. unexplored [MT09]. unfolding [PMB^{+}24]. unicellular [dSFFA03]. uniformity [GM93]. unimpacted [DSL12]. Unionidae [LB14]. unique [MSDDD22]. uniqueness [BSW^{+}24, TLH18]. United [CEH12, Cor78b]. Universiteit [Pol73]. University [GWF92, Pol73, Ano83b]. Unlocking [SPB^{+}24]. unpredictable [BLP16]. unprocessed [BMA00]. Unraveling [EMF14, PBM21]. unsaturated [MBM15]. unseasonal [VGJ20]. Unveiling [NudWB^{+}25]. upland [BLC12, IGS10, NH92, WB10]. upon [GCT^{+}23, Roe96, SY02, TM84, Wil79]. upper [CMBI25, CM95b, DRM18, MHE93, SGK07, YBZ20, ZZL16, dSPdS10, ATB07, KMK09]. Upstream [FBM22, PR16, VTE19, BMO15, YKJ12]. Upstream/downstream [FBM22]. uptake [CVG03, GXY14, JC83, JCL09, KK17, KA97, KSBŁ95, SGR86, WDP15]. Upwelling [HKO14]. Upwelling-like [HKO14]. urban [dMCFS^{+}23, DWB10, EAO20, FSdO20, LZL22, MMC14, dCNGF25, PRW08, RU87, STHN03, TAN14, WDP15]. urban-agricultural [dCNGF25]. urbanization [RPSSDS23]. urbanized [CHP06]. urchin [SHP21]. urdaibai [RFO94]. urea [ZWJ19, lCdB70]. ureumderivaten [lCdB70]. Uruguay [CMM05]. US$ [Cad05e]. USA [AC86, AP00, CAJ11, ET16, GBT19, Gru11, HBR12, KK07, KGB04, PDP02, PKE22a, PKE22b, Rod10, TBF99, zEGB13]. Use [ASLT15, CTO01, CSV06, DDF93, MSK93, Van85, ACL13, AD89, BVM15, BRA^{+}23, BKG16, BLC12, CPM09, COC18, DTC19, GSH17, HHP12, IS06, IAHB18, Ker75, Koo76, LZN14, MSRB23, MAM13, MSP13, Mor99, NKKG07, PV95, PBR01, PMSS18, PD10, PDD07, QCF15, RHA00, RA07, SSI21, SLL09, Sur01, Van82a, VPS21, WiNDF22, WD00, Ker70]. used [CSM08, Lah98, MCSV21]. useful [KWO11]. uses [MvAV20]. Using [AKF^{+}23, BRP15, PLM^{+}22, PDE08, VVW95, BM85, Boe86, Das07, DGH^{+}24, Fra00, GGD07, GCN^{+}23, GPGSMM^{+}23a, GPGSMM^{+}23b, Gul85, HBO03, JBB07, JVCJ25, Ker85, KP99, LKM01, LZL22, MG06, NLC06, NHiN15, NAT08, ODI08, OHA08, PPC08, RDG86, RCL94, SJJA23, SB88, SK04, STT01, TV85, TFT10, TSZ10, UOJGRL21, VHtB17, WSvA98, WDP15, dJG07, vLJA07]. USSR [GKK92]. Utevsky [ŽRP16]. utilisation [CCS99, GA10, LTL09]. utilization [DSN09, DCL19, PSZ79, RKS14, SFA05]. Utrecht [De 76b, Bel79]. Utricularia [CBE20, GANA00, VKD77]. Uttar [KP24]. Uttarakhand [IUK22]. Utterbackia [HPM20]. UV [HH13, KA97, PA98, SMC08, Van97, War97]. UV-B [KA97, SMC08, Van97, War97]. UVB [VRR17]. UVR [PAY03].

V [Gul98, Lam79]. Vaas [NB80]. Val. [Nwa95]. Valencia [KGC10]. Valenciidae [KGC10]. Validity [AA95]. Valley [Dav71, Hig71, Moe71]. valleys [AAV07]. Vallisneria [DZL21, KT09, LLG10, LWX19, MLL21, PDD07, PCD11, SB14, TZD21].

Valuation [Sch72b]. value [KKS11, PAAJM17, PPCH22, TNtK03]. values [Hos80, PLC22, VMS94, vLJA07]. Vansickle [Ald79]. vanuit [Ess71]. var [CCS09, DNG22]. Varia [Ano68h, Ano69g, Ano69h, Ano69i, Ano70f, Ano71b, Ano72g, Ano72h, Ano77c]. variabilities [BTP24]. Variability [Nie74, Vij91, BNV03, BWS21, BJ04, FR09, KHG20, LPB07, LNHN08, MOdSCP07, OMC22, PDM95, SC99, SCO09, TS17, Wey09]. variable [BLB^{+}23, BLP16, GGD07, KP99, MFS95, Rat17, VHtB17]. variables [Bal09, BSA24, CMBG20, KRATA20, MAS14, PU15, TWC13, ZCZ14]. variance [BBD17, IJ06, RFW17]. Variation [DSG^{+}24, ESG04, IIMRC21, RW98, AAMS24, ACE10, AR97, BB12, BdSCH09, BSA24, BKW94, BJJ12, BN24, CM95a, CCK10, CCC19, CHF09, dMCFS^{+}23, CS13, DRM18, De 97, DPT04, DSL12, FCS15, GS04, GC08, HJ23, Hig09, JTL09, Kat92, LWO01, LV93, Lu05, MGL22, Mag00, NYT09a, NW24, ÖE04, PMB^{+}24, PEREBM^{+}23, QCF15, RS85, RJD07, SCN^{+}24, SVM08, SB04, SYW09, ŠV09, STH21, SAM07, TDP06, VEK20, WNS08, WTG95, WMD07, WB10, YSYZ23]. Variations [CMG95, MHE93, MK05a, AVdNM19, Bal09, BKD^{+}24, BAG94, ČN10, CSK10, CVG03, ECL94, GKK95, IRT06, KMS82b, Leh04, Mar93, Nwa87, RSM24, RM00, RvGB05a, RvGB05b, SY02, WBR17, WGL21, YBZ20, Zdo09]. variegatum [BH82]. various [CC94, RMR21, SH11, VVA86, VPF21, WZL19, Ano68h, Ano69g, Ano69h, Ano69i, Ano70f, Ano71b, Ano72g, Ano72h, Ano77c]. vary [PFMD00]. varying [DGN21, GX11, KTJ10, MJEC05, MvAV20, VL18]. Vaucheria [LL77, SK09]. Vechten [BZD81, BBV78, Gon79, Gul76b, OBC82, Ops80, Ver76a]. vechtplassen [VBD96]. Vechtplassengebied [Gul76a]. vectors [BdPP17, LLF99]. veen [Beu71b]. veen-bodemhapper [Beu71b]. Veenkoloniaal [De 77b]. Veeno [HWG77]. Veere [BD74, Bak77, BV78]. Veerse [Bak72]. vegetation [DPL03, Dvo69, FCG16, Gri89, Har73b, Hig81, JCF17, KOA00, Kor14, MS02, Mun94, NAM19, Pol75, SSV21, TM84, TBSZ22, TNtK03, VP77, VKD77, WvLA14, dL74]. vegetations [RD77]. vegetative [CDL14, MKA79]. vein [QZL^{+}24]. veins [GvdH68, Hig68, Hig69]. veligers [PCB10]. Velocimeter [PJH06]. velocimetry [Sta06]. Velocity [Gre05, FDH94, GBT19, HW13, RVJ98, RPS20, SAM07, YSY^{+}22]. Veluwe [HM86]. Vendée [MBFD24]. Vendyurskoe [TGZ09, Zdo09]. venematen [Dre68, GvdH68, Hig68, Hig69, Gar71, Gee69]. Venemates [Gar71]. Venezia [BPP21]. Venezia-Giulia [BPP21]. Venezuelan [LMA10]. Venice [RM00]. venous [Dre68]. venture [CGD19]. venturing [FASdPN^{+}25, SASA^{+}24]. Veracruz [RRRRA07]. verbana [UT16, ŽRP16]. Verbreitung [HD79b]. Verdonschot [Ano05b]. VERENIGING [Gel70, Ano72i]. verlandingszone [Gar71, Hig68]. vernal [Bau00, Sæt92]. verontreiniging [BH71, dH71b]. Verrill [EK93, Zet97b]. Versatile [TC15, AD89]. Verslag [Ano72i]. Verslagen [Gel70]. versus [AMTSJ15, BH22, CH12, GBA20, KKM09a, MCvE09, MMRG03, PCB10, SEB11, VvGV12]. vertebral [LA99]. vertens [Bak80]. Vertical [CV91, DZG10, GZG02, HWG77, HH13, Ola92, PS06, BBM10, BNPC07, BS95, DTH20, DBG02b, GBM10, GEK97, HLL14, MR81, MHG16, MOCPG09, PMD10, RFR91, RGD10, SZH21, TZ05, Vij91, WBB93, WFFM10, XCL10, ZTG10]. verticillata [KC82]. verticillatum [CvDG07]. Vervuiling [Ess71]. vesiculosus [BHM03, EMT00, HMH05]. Veulume [Zoe72]. Veulume-Massief [Zoe72]. vexillum [KHG20]. VI [Sep79]. via [DGN21, FO98, GXY14]. Viability [VET20, OFCP22, YSYZ23, YN08a]. viable [AFR14, KG03]. vicinity [Hog69]. vicinus [NBM98]. Victoria [APS81]. video [ECB09]. Vietnam [ACL13, DT23a, DT23b, DDPT24, TDHH23]. vigor [xHCjX^{+}23]. viguieri [CGD19]. vijver [Par69]. Villerest [DD99]. villosus [KKM09a, PDvdV06a, PDvdV06b]. vinifera [OMC22]. virescens [PR92, TK19b, TK19c]. Virgin [CRR22]. virginalis [HGB20]. virgo [BKO92, IS06]. Viridiplantae [ČN10]. viridis [BRB97, Boc97, EK93, KP97, RVJ98, RVJ06, RIT04]. virilis [CBNF19]. Virological [AA95]. viruses [CVM21]. vis [KvdM70, Wil70]. visible [Das07, De 00]. visible-near-infrared [Das07]. visserij [NZ75]. Visserijkundige [Dee69]. visual [MLZ10]. Viswater [Lin72]. Vitis [OMC22]. vivida [HP06]. viviparus [KBO74]. vivo [Vri80, WCT00, PB84]. Vlaamse [Gys72b]. VNIR [Das07]. voedsel [Gie71]. voedselketen [Beu71a]. voedselketen- [Beu71a]. voedselopname [Ker70]. Vogt [MTdS^{+}24]. volcanic [NGSOAD20, SDW^{+}24]. volcano [MT09]. Volga [GZN20, OMM05a, OMM05b]. volgorde [Ano68d]. Volkerak [GD91, Pee74]. Volkerak-Zoommeer [GD91]. Volume [Ano98i, Ano03c]. volutator [Kle84]. Volutidae [GOB24]. voor [Blo70, Gie71, Pol73, Rin70a, Wil70, vdW70]. voordrachten [Ano70c]. voorkomen [Dre68, dW71b]. voorkoming [Pee71]. Voorwoord [PV68]. Võrtsjärv [NLF10, VHB09]. Vossenbelt [Sch68a]. Vrije [Pol73]. vs [Fli85]. vulgaris [AND25, ARG20, DCG19, GFSN04, GCT^{+}23]. vulnerability [DD17].

W [Gos98b, Gos99, Laa97, RQ93, Shi98, Vaa79, War97, ZV93, dSLM94]. W. [Baa88a]. Waardering [Sch72b]. Wadden [Cad78, DBV96, FD94, HZC95, LKE94, Van77a, Van78a, Van82b, Vos87, Zim78, Cad80, De 80c, Ess71, Sch95, Web71b]. Waddensea [Har94]. Waddenzee [Ess71]. Wade [Ver97]. wadpieren [Web71b]. Wageningen [Ano81a]. Waikato [DÖD21]. Wainman [Gul99]. Wallago [IRT06]. Walleye [BRA^{+}23]. Waquoit [TGV00, TBF99]. war [NNJ22]. warm [Bar21]. warm-temperate [Bar21]. warmer [Wey09]. Warming [dFBdGGMLT21, GKM22, CVM21, FM10, KOA00, LCL16, vOCRM19, WF12, ZYS14]. was [LME10]. Washington [ET16, VK12]. waste [De 77b, De 78, Ess03, Ros69b]. waste-water [De 77b]. wastewater [Cha87, Kim99, Kla80, QZL^{+}24, ST09, TBS18]. wastewaters [STHN03]. wasting [DR96, Nie94]. Water [ABV72, Bak72, BB78, KCR92, Lin72, PBF06, RL05, SDN25, SG08, VSM19a, Ver72, ZBA19, Aag92, Ade71, AIS02, Ano72b, Ano87, AOS^{+}24, AC16, BGRM^{+}24, BD74, BH16, BVV80, BNV03, BB87, BRA^{+}23, BZV93, BT01, BSA20a, BB23, BJS12, Cap75, Cha89, CZZ^{+}25, Cor78a, Cra85, Cra87, DWA11, DTH20, Dan84, DS75, DHW81, DTD94, DB02, DD84, De 77b, DF90, DGH^{+}24, DZL21, DCC09, Dor74, DV92, DV74, DZT13, DVK78, DBG13, FdLY^{+}24, Fon71, GMD19, GB07, GEÓ04b, Goe07, Goo79, Gri89, GYW17, GCE11, GV89, HVS87, Hal76, HVB97, HGR11, Har74, HJV08, Hig74, HM84, HvB85, JTGS24, JJG^{+}25, JvDS06, KGS02, KFS04, KK17, Kel73, KPS09, KIJ11, KS94, Klo76, KD82, Koo73, KKR02, KIG06, KBO74, LMV09, LMA10, LLX11, LZD16]. water [LZL22, LLL^{+}23, LZZ^{+}25, LTH22, LRMdF19, MSW10, Mol80, MSF21, MOdSCP07, MOPCP08, MFP01, MD92, Mur69, Mur74b, Muu74, NYT09a, NvZvG16, Nor98, OBC82, OR06, PAP24, PHB09, PMM10, PT14, RVJ98, REP01, RHL02, Ric86, Rin81a, RCB95, Rod10, Ros69a, Ros69b, RB78, Sch78, Sch68a, Sch74, Sch88, SM12, SD86, SKV02, Soi05a, Soi05b, SSB21, SOU14, TK88a, TK88b, TK89, TŠB17, The02, TG88, TIF10, Van74, Van79a, VS96b, VK90, VVW95, Van87, VOD93, VDB81, VHC92, Ver76b, Ver73, VKK02, VHN00, WCZ11, WYZ13, WvRvdV03a, WLN12, Wil79, WMD07, Wor90, WDA15, ŻKŻ10, Zdo09, Zet97b, ZZL16, ZJN^{+}23, Zoe72, Żur06a, ZV93, dB69, dL74, vdW70, SRH08, Laa97]. water-chemistry [SSB21]. water-column [BT01]. water-filled [PT14]. water-level [BH16, RHL02, SOU14]. Waterbird [AFR14, HKH^{+}23, KT15]. Waterbird-mediated [AFR14]. waterbirds [MOPCP08]. waterborne [MRR10]. waterchemistry [Kla88]. watercourses [BPP21, DH01, Kop21]. watercress [Ber24]. Wateren [GD72, Sch72b]. waterfowl [Haa73, MS02]. watering [xHCjX^{+}23]. watermilfoil [SRG18]. Waters [GD72, Sch72b, ABDPPP^{+}24, Ano05b, Bij75, CBC99, CMG95, CBF19, Cla87, Coe75, CFRNT21, De 74, DDV84, De 00, DTC19, DVK78, FO98, GBJM22, GEK97, Gri89, GV89, HdC24, HvdH84, Hos80, HKB19, Kat92, Kre82, LLF99, M.75, MAS14, MBM^{+}22, Mer79, dSLM94, MMBP21, Mit74, PMM10, Pos74, RVJ98, RDG97, Sch97b, VHC92, Wol00, Cad05d]. watershed [AC16, BC23, BSW^{+}24, BKC11, BRP15, dMCFS^{+}23, Gru11, LRMdF19]. waterverontreiniging [Ros69a, Ros69b, dB69]. waterway [VDB81]. Watling [FvM03]. wautieri [HVS87, Van91]. wave [EMN04, FCW94]. wave-tidal [FCW94]. wavelength [BWS21]. waves [DVK81, Den06, KEH09]. way [Ker97, PMA22]. ways [SDN25]. weak [OLJB20]. Wealth [Gys72b]. weather [CC94, GH06]. Web [Cad99, UKS05, ASA06, BGL13, Coe97, CSV06, DWA11, dSDLdA22, DJD92, HKO14, HXR09, JVG95, KKD02, KMG^{+}23, MBMV20, MTSJA19, MDV03, OLTV19, PCGW20, SCEGL12, SCS12, TJB98, VGG89, VGM93, WHW14, YL11]. weberi [AR05]. webs [DGBL15, iNTT05a, iNTT05b, SH16, STTS18a, STTS18b, SMJ03, VCT13]. weed [Bel87, LW11, Van75b]. Weekly [WCZ11]. weight [GD09]. weir [YKJ15]. Wenig [De 79a]. were [DR96]. Werff [Bak94, DVZ94, Har94, Nie94]. werkgroep [Beu71a]. werkgroepen [vdB70]. Werkkamp [Ano72i]. werkzaam [vdB70]. wesenbergii [DCL19]. Weser [MSK93, RCSF93, SHS93]. West [VTE19, HHV93, Kre82, CDW92, GKK92, KEL09, KKO08, RAO^{+}24, RAN24]. Western [BP21, AAMS24, Cad80, DBV96, MŚSJ16, MFL09, SB78, VK12, BH71, Bar21, CN94, Den94, FvM03, HF08, KG94, PKE22a, PKE22b, SAB05, UOJGRL21, Ver79, dP71]. Westerschelde [BH71, BD74, Bak77, DMN88, Hei88, Hui88, HMB88, LV93, MDH93, OSR88, dP71]. Wetensch [Har73a]. Wetland [Bou07, Gul98, MNJ21, ABF01, BBM09, CDD20, FAK01, Flo01, FDP01, GGT14, JCL09, KRJ09, KHF01, KÇA22, LME10, MGT17, NAM19, PJF01, RFP01, REP01, RFF01, TIF10, dSPdS10, vdHD17]. wetlands [AP00, ES09, HEB00, HC09, MBMV20, MBFD24, NB22, OSO09, SOU14, VPS21]. wetted [NB22]. Wetzel [Gul05]. WFD [SRH08]. whale [BOB20, SBLdAM23]. Wheels [Gys72a]. Where [BRJ97, GTPH08]. whether [HKH^{+}23]. whirling [DVK81]. White [GKM17, Geu84, JBB07]. White-winged [GKM17]. whitefish [HHP12]. Whitemouth [MCV08]. who [OR11]. whole [SPB^{+}24]. whole-genome [SPB^{+}24]. whom [CH83]. widely [Hal76, JDF23]. width [SBLdAM23]. Wielen [Gys72a]. Wierzejski [DK80]. Wijk [Dor80a]. Wijnhoven [TKM03, WvRvdV03b]. wild [KZV02, KAN20, LAM06, TCL22]. wildfire [RAD21]. Wilken [Gos98b]. Wille [Pol76]. Winchester [KHG20]. Wind [FA99, GKM17, Lee82, AGB94, BAG94, GVV86, KS94, MIMRRZ08, MOCPG09, PVB20, SEE22, Sir17, WMTR07]. Wind-induced [FA99, BAG94, MOCPG09]. wind-swept [MIMRRZ08]. window [PV95]. winged [GKM17]. Winter [AK09, SFV09, BTB09, BHL09, DH09, FG91, GBM10, LWO01, LCL16, RZG09, SLG09a, VHB09, VMS09]. winters [Wey09]. wireni [BSB97, Sch97b, SVB97, Zet97b]. within [BC23, CMH14, Flo01, GSH17, GEÓ04b, HÅB08, KKM09a, KM17, MS07, MKS22, Ola92, TH16, WCT00, WKH82, ZBB19, ZWY20]. within-host [WCT00]. without [NAT08]. witness [Har94]. Wittrockiella [Pol76]. Woluwe [TKD01]. Wood [GK07, CBNF19, DLG20, KK07, LB20, SKM05a, SKM05b, IM06]. wood-inhabiting [SKM05a, SKM05b]. woodland [Ran74, VC89]. woody [TRP22, VNKdOR15, VNSFG16]. Work [DVZ94, Eng74]. workers [AFRS23]. working [Beu71a, vdB70]. works [PP95]. Workshop [Ano72i]. World [WB10]. worms [Web71b]. wrightii [PEREBM^{+}23]. Württemberg [Mar01].

XAD [VVW95]. Xanthophyceae [SK09]. xiii [Cad05e]. Xochimilco [NEGS07]. Xolotlán [BEZ91, CV91, CHV91, CMV91, EHM91, GR91, HH91, HRP91, LCF91, Lac91, MG91a, MGVC91, MG91b, Vel91].

Yamuna [SC00]. Yangtze [JCF17, WXW11]. Year [Bak72, Kat92, Bes76, Bes87, Bra21, CM95a, Fli85, GH06, LMV09, MMM11, OLJB20, RMH21, RMS11, Ver76a, VMS09, ŻKŻ10]. year-class [LMV09]. Year-to-year [Kat92, GH06]. Yearly [CB00c, VV85]. years [CH91b, Gul89, Gul07b, Hig79, KK12a, NDW96, PVH17, RCSF93, Rin97, VK90, VGG89]. Yellow [WWG^{+}25, BRA^{+}23, KW12, Pad13, RVB08, GYW17, TJZ^{+}24]. Yellowstone [CAJ11, Kae20]. Yenisei [Bol04, SGK07]. Yergelijkend [GB69]. young [MMM11]. young-of-the-year [MMM11]. YouTube [BSMVP24]. Yucatan [ABDPPP^{+}24]. Yuma [vdV99].

Zacco [YKJ12]. zaddachi [IDP19, Wil07]. Zealand [BT23, BHG16, DÖD21, HKT16, HBR12, LDK20]. Zebra [VD85, BRA^{+}23, BKD92, KK11, KJP14, Kra95, RVK96, WFBJ17]. Zeeland [Pee71]. Zeeuwse [Pee71]. Zegrzyńsky [Kuk92]. Zerga [KHF01]. Zett [Lar92]. Zhulidov [Gul98]. Zidona [GOB24]. zijnde [vdB70]. zilvermeeuwen [Swe71]. Zimbabwe [CPM09]. Zinc [VV07, MEB93, RMW91, SAGN07]. Zivić [UT16]. Zizania [LAM06, TCL22]. Zn [WA84]. Zoarces [KBO74]. Zonation [Lov74, YME98, CV96, Dor76b, Dvo69, Ole97]. Zone [Nie98, Baa80, Bak78, Bal09, BR05, BKW94, DHPN^{+}24, DSD87, ECGC25, FSN21, FCW94, Gar71, Hig68, KOA00, KSBR94, KD82, Kuk92, MKV15, MBM15, MR87, MAM13, Ola92, QCF15, RTW17, RAA98, STTS18a, STTS18b, TD87a, ZTG10, dRFC21]. zones [EPL^{+}23, NvdVdlM01]. zoning [SWW02]. zoobenthos [HB03, Mas85, RL05]. Zoom [MRD89]. Zoommeer [GD91]. zooplanktivorous [HBL13]. Zooplankton [BDH09, BFB15, BV71, Fra80, GR97, GB07, GPV19, Gul85, KP24, Lam97, MDV03, PFMD00, RFF01, Sar05d, SS98, VV05, WHS13, Żur06b, Gee69, BV78, Bak79, BGK85, Bar77, BdPP17, BKW94, BASJ94, Bra21, BBM21, Bre85, But81, CGC19, CNA07, CDL13, CHV91, CMV91, DB85, De 79b, DH09, Don79, DKI05, DÖD21, DBG13, FGM20, FNW94, GPGSMM^{+}23a, GPGSMM^{+}23b, GJ06, GANA00, Gul76b, Gul89, GM16, HEB00, HEPH09, HH13, Irv86, Irv89, KTJ10, KJF00, LHV07, MG91a, MCSV21, NYT09a, PCB10, PVB20, PHD13, Pri03, REP01, RdMJN22, dSSSES21, SNGV06, SNG21, SCG09, SVM08, SPA12, SR85, TLF19, TŠB17, Tay93, TTD02, TGW06, TZ05, VFSH12, Van79b, VPS21, Wel74, WWŻ06, YKU01, ZTNW21, ZWY^{+}22, Gee69]. Zooplankton-phytoplankton [GR97]. zooplanktonic [AFC20]. zoospore [BR87]. Zostera [Bol07, BH82, DR96, Har82a, Har94, JLG06, LDDS82, Nie79a, Nie82, NDW96, Pel82]. Zschach [Gie83]. Zuid [DM71]. Zuid-Limburg [DM71]. Zuidwest [Bak72]. Zuidwest-Nederland [Bak72]. zware [Fon71, dG71]. zwembaden [lCdB70]. Zwemlust [Gul89, VGG89]. Zygnematophyceae [SH11].

References

Alcantara:1995:VQR

[AA95] Fernanda Alcântara and Maria Adelaide Almeida. Virological quality of the Ria de Aveiro: Validity of potential microbial indicators. Aquatic Ecology, 29(3-4):419–425, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02084240.

Akinpelu:2024:HCA

[AAA^{+}24] O. T. Akinpelu, F. O. Arimoro, A. V. Ayanwale, V. I. Chukwuemeka, T. A. Oladeji, and E. O. Akindele. Habitat characteristics and anthropogenic activities influence the distribution of macroinvertebrate traits and ecological preferences in Nigerian streams: a case study of Osun State, Nigeria. Aquatic Ecology, 58(3):833–852, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10108-3.

Aagaard:1992:OTS

[Aag92] K. Aagaard. Ordination or typology — the search for a stable classification of running water communities. Aquatic Ecology, 26(2-4):441–445, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255273.

Ali:2024:SVI

[AAMS24] Mohammad Ali, Matrah Al-Mutairi, and M. N. V. Subrahmanyam. Seasonal variation of intertidal Gastropoda at rocky and sandy habitats of Umm Al-Namil Island, Kuwait, north-western Arabian Gulf. Aquatic Ecology, 58(3):931–948, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10116-3.

Ali:2025:MCM

[AANO25] Andrew A. Ali, Frank C. Akamagwuna, Chika F. Nnadozie, and Oghenekaro N. Odume. Macroinvertebrate colonisation of macroplastic litter in minimally disturbed river sites, Eastern Cape, South Africa. Aquatic Ecology, 59(1):233–245, March 2025. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10159-6.

Al-Azri:2010:CMS

[AAPdG10] Adnan R. Al-Azri, Sergey A. Piontkovski, and Helga R. do Gomes. Chlorophyll-a as a measure of seasonal coupling between phytoplankton and the monsoon periods in the Gulf of Oman. Aquatic Ecology, 44(2):449–461, June 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9303-2.

Arriesgado:2024:SCA

[AAR^{+}24] Dan Arriesgado, Elgen Arriesgado, Elnor Roa, Aida Perpetua, Ruby Gonzales, Rodrigo Acuña, Rustan Eballe, Delyn Bucay, Mayenne Balaba, Lyndon Roa, Bopp Benjack Janolino, Richard Muallil, Julius Mingoc, and Marnelle Sornito. Seagrass cover and associated macrobenthic marine invertebrates in Southern Philippines. Aquatic Ecology, 58(3): 643–657, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10095-5.

Ayala:2007:MLA

[AAV07] Rosmery Ayala, Francisca Acosta, and Paul A. Van Damme. Management of Laguna Alalay: a case study of lake restoration in Andean valleys in Bolivia. Aquatic Ecology, 41(4): 621–630, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9123-1.

Amal:2024:EAM

[ABA^{+}24] K. S. Amal, Niya Benny, P. Ajayakumar, Lathika Cicily Thomas, and K. B. Padmakumar. Ephemeral analysis on microphytobenthos in an estuarine system along the southwest coast of India. Aquatic Ecology, 58(4):1049–1065, December 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10123-4.

Almazan-Becerril:2024:CFM

[ABDPPP^{+}24] Antonio Almazán-Becerril, Benjamín Delgado-Pech, Jorge Carlos Peniche-Pérez, Juan Manuel Arana-Ravell, and José Adán Caballero-Vázquez. A case of fish mortality caused by Prymnesium parvum in inland waters in Yucatan, Mexico. Aquatic Ecology, 58(4):1281–1285, December 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10141-2.

Appleby:2001:RDD

[ABF01] P. G. Appleby, H. H. Birks, and A. A. Fathi. Radiometrically determined dates and sedimentation rates for recent sediments in nine North African wetland lakes (the CASSA-RINA Project). Aquatic Ecology, 35(3-4):347–367, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011938522939.

AzevedoeSilva:2021:ARA

[ABF21] F. Azevedo e Silva, A. C. Brito, and P. M. Félix. Allometric relationships to assess ontogenetic adaptative changes in three NE Atlantic commercial sea cucumbers (Echinodermata, Holothuroidea). Aquatic Ecology, 55(2):711–720, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09856-3.

Aiello:2021:IOB

[ABT21] Giuseppe Aiello, Diana Barra, and Marco Trifuoggi. Infralittoral Ostracoda and benthic Foraminifera of the Gulf of Pozzuoli (Tyrrhenian Sea, Italy). Aquatic Ecology, 55(3): 955–998, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09874-1.

Asman:1972:EHC

[ABV72] W. A. H. Asman, B. G. H. J. Beltman, and J. T. A. Verhoeven. Enkele hydrobiologische en chemische gegevens van het water in de Ooypolder. (Dutch) [Some hydrobiological and chemical data of the water in the Ooypolder]. Aquatic Ecology, 6(2):53–61, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02304095.

Acs:2008:CAB

[ÁBZ08] Éva Ács, A. K. Borsodi, and Gy. Záray. Comparative algological and bacteriological examinations on biofilms developed on different substrata in a shallow soda lake. Aquatic Ecology, 42(4):521–531, December 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9132-0.

Auer:1986:MMP

[AC86] Martin T. Auer and Raymond P. Canale. Mathematical modelling of primary production in Green Bay (Lake Michigan, USA), a phosphorus-and light-limited system. Aquatic Ecology, 20(1-2):195–211, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02291163.

Atkinson:2016:BAC

[AC16] Carla L. Atkinson and Joshua T. Cooper. Benthic algal community composition across a watershed: coupling processes between land and water. Aquatic Ecology, 50(2):315–326, June 2016. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9580-5.

Almeida:2002:BPR

[ACA02] M. A. Almeida, M. A. Cunha, and F. Alcântara. Is bacterioplankton production in the Ria de Aveiro influenced by salt marshes and bed sediment? Aquatic Ecology, 36(4):469–482, December 2002. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1021104402018.

Amaral:2012:MAA

[ACB12] Valter Amaral, Henrique N. Cabral, and Melanie J. Bishop. Moderate acidification affects growth but not survival of 6month-old oysters. Aquatic Ecology, 46(1):119–127, March 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-011-9385-5.

Alves:2010:SBI

[ACE10] João M. Alves, Adriano Caliman, and Francisco A. Esteves. Stoichiometry of benthic invertebrate nutrient recycling: interspecific variation and the role of body mass. Aquatic Ecology, 44(2):421–430, June 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9302-3.

Anders:2009:PST

[ACH09] N. R. Anders, T. Churchyard, and J. G. Hiddink. Predation of the shelduck Tadorna tadorna on the mud snail Hydrobia ulvae. Aquatic Ecology, 43(4):??, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9216-5.

Arai:2013:MHU

[ACL13] Takaomi Arai, Naoko Chino, and Dung Quang Le. Migration and habitat use of the tropical eels Anguilla marmorata and A. bicolor pacifica in Vietnam. Aquatic Ecology, 47(1):57– 65, March 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-012-9424-x.

Armitage:1989:VLS

[AD89] Patrick Armitage and Anthony Davies. A versatile laboratory stream with examples of its use in the investigation of invertebrate behaviour. Aquatic Ecology, 23(2):151–160, December 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256732.

Alonso:2020:ELL

[AdAPC20] A. Alonso, B. R. Vázquez de Aldana, and M. E. Pérez-Corona. Effects of leaf litter extracts from four tree species on aquatic invertebrates: an ecotoxicological risk assessment approach. Aquatic Ecology, 54(4):1155–1168, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09800-x.

Aabjornsson:2000:IPD

[ÅDB00] Kajsa Åbjörnsson, Jonas Dahl, and Christer Brönmark. Influence of predator and dietary chemical cues on the behaviour and shredding efficiency of Gammarus pulex. Aquatic Ecology, 34(4):379–387, December 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011442331229.

Adema:1971:TIV

[Ade71] D. M. M. Adema. De toxicite it van koper in water. (Dutch) [The toxicity of copper in water]. Aquatic Ecology, 5(4):177, December 1971. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185959.

Admiraal:1995:R

[AdJD95] Wim Admiraal, Victor N. de Jonge, and Jaap Dorgelo. Reviews. Aquatic Ecology, 29(2):183–197, June 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02336049.

Admiraal:1990:RPA

[Adm90] W. Admiraal. Research on perturbed aquatic ecosystems. Aquatic Ecology, 24(1):3, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256743.

Akbary:2025:SOE

[ADS25] Paria Akbary, Sachinandan Dutta, and Salim Sharifian. Salinity optimization for enhanced growth and biochemical performance of Saccostrea cucullata. Aquatic Ecology, 59(1):67–81, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10148-9.

AlcantaraMissiasGomes:2020:FCI

[AFC20] Ana Caroline Alcântara Missias Gomes, Leonardo Fernandes Gomes, and Ludgero Cardoso Galli Vieira. Forest cover influences zooplanktonic communities in Amazonian streams. Aquatic Ecology, 54(4):1067–1078, December 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09794-6.

Anastacio:2014:WMP

[AFR14] Pedro M. Anastácio, Miriam P. Ferreira, and João E. Rabaça. Waterbird-mediated passive dispersal is a viable process for crayfish (Procambarus clarkii). Aquatic Ecology, 48(1):1–10, March 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-013-9461-0.

Aguiar:2023:DWW

[AFRS23] Thayanne Medeiros Aguiar, Vanessa Fernández-Rodríguez, and Cinthya Simone Gomes Santos. Dirty workers: What do we know about the ecotoxicological studies in polychaetes? Aquatic Ecology, 57(3):783–796, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10049-3.

Arcifa:2007:P

[AG07] Marlene Sofia Arcifa and Ramesh D. Gulati. Preface. Aquatic Ecology, 41(4):509–510, December 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9124-0.

Arfi:1994:SMW

[AGB94] Robert Arfi, Daniel Guiral, and Marc Bouvy. Sedimentation modified by wind induced resuspension in a shallow tropical lagoon (Côte d’Ivoire). Aquatic Ecology, 28(3-4):427–431, September 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334213.

Andersen:1982:NCM

[AH82] Frede Østergaard Andersen and Jette Ingrid Hansen. Nitrogen cycling and microbial decomposition in sediments with Phragmites australis (Poaceae). Aquatic Ecology, 16(1):11– 19, June 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255409.

Adham:2002:BCN

[AIS02] Khadiga G. Adham, Hania M. Ibrahim, and Ramadan A. Saleh. Blood chemistry of the Nile tilapia, Oreochromis niloticus (Linnaeus, 1757 ) under the impact of water pollution. Aquatic Ecology, 36(4):549–557, December 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1021137122046.

Amundsen:2009:WEA

[AK09] Per-Arne Amundsen and Rune Knudsen. Winter ecology of Arctic charr (Salvelinus alpinus) and brown trout (Salmo trutta) in a subarctic lake, Norway. Aquatic Ecology, 43(3): 765–775, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9261-8.

Adedapo:2023:UMF

[AKF^{+}23] Abiodun M. Adedapo, Esther D. Kowobari, Ibrahim R. Fagbohun, Tolulope A. Oladeji, Timothy O. Amoo, and Emmanuel O. Akindele. Using macroinvertebrate functional traits to reveal ecological conditions of two streams in Southwest Nigeria — a case study. Aquatic Ecology, 57(2):281– 297, June 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10010-4.

Arriesgado:2015:GDS

[AKL15] Dan M. Arriesgado, Hiroyuki Kurokochi, and Chunlan Lian. Genetic diversity and structure of the tropical seagrass Cymodocea serrulata spanning its central diversity hotspot and range edge. Aquatic Ecology, 49(3):357–372, September 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9529-0.

Alekseev:2009:MDE

[AL09] Victor R. Alekseev and Dmitry L. Lajus. Maternal and direct effects of natural-like changes of photoperiod and food condition manipulation on life history parameters in Daphnia. Aquatic Ecology, 43(2):415–421, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9168-9.

AmorimReis-Filho:2024:BRE

[AL24] José Amorim Reis-Filho and Antoine O. H. C. Leduc. Balancing renewable energy and river conservation: effects of hydropeaking from small hydroelectric power plants on fish stranding in small Brazilian rivers. Aquatic Ecology, 58(3): 551–569, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10090-w.

Albertsson:2004:TII

[Alb04] Jan Albertsson. Trophic interactions involving mysid shrimps (Mysidacea) in the near-bottom habitat in the Baltic Sea. Aquatic Ecology, 38(3):457–469, September 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000035163.30037.38.

Aldenberg:1978:TCR

[Ald78] Tom Aldenberg. Theoretical considerations regarding the calculation of production in the case of continuous reproduction. Aquatic Ecology, 12(2):119–126, July 1978. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260712.

Aldenberg:1979:CPR

[Ald79] Tom Aldenberg. The calculation of production, reproduction and total growth in the autonomous Vansickle equation. Aquatic Ecology, 13(1):3–12, April 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260302.

Althuis:1998:SPM

[Alt98] Ij. A. Althuis. Suspended particulate matter detection in the North Sea by hyper spectral airborne remote sensing. Aquatic Ecology, 32(1):93–98, March 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009946711742.

Altermatt:2013:DRM

[Alt13] Florian Altermatt. Diversity in riverine metacommunities: a network perspective. Aquatic Ecology, 47(3):365–377, September 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9450-3.

Alva-Martinez:2007:EMD

[AMSN07] Alejandro Federico Alva-Martínez, S. S. S. Sarma, and S. Nandini. Effect of mixed diets (cyanobacteria and green algae) on the population growth of the cladocerans Ceriodaphnia dubia and Moina macrocopa. Aquatic Ecology, 41 (4):579–585, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9115-1.

Amezcua:2015:SIA

[AMTSJ15] Felipe Amezcua, Víctor Muro-Torres, and Martín F. Soto-Jiménez. Stable isotope analysis versus TROPH: a comparison of methods for estimating fish trophic positions in a subtropical estuarine system. Aquatic Ecology, 49(2):235– 250, June 2015. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-015-9517-4.

Abbas:2025:KMS

[AND25] Mohamed Abbas, Lixiao Ni, and Cunhao Du. Kinetic modelling and salinity tolerance in Chlorella vulgaris and Desmodesmus communis (Chlorophyta): insights into differential growth responses. Aquatic Ecology, 59(1):145–158, March 2025. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10153-y.

Anonymous:1968:IDI

[Ano68a] Anonymous. Inhoud deel I en II. (Dutch) [Contents part I and II]. Aquatic Ecology, 2(3):85, September 1968. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185816.

Anonymous:1968:KMM

[Ano68b] Anonymous. Kaart met monsterpunten. (Dutch) [Sample points map]. Aquatic Ecology, 2(3):94, September 1968. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185819.

Anonymous:1968:MVB

[Ano68c] Anonymous. Mededelingen van het bestuur. (Dutch) [Board announcements]. Aquatic Ecology, 2(3):82, September 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185814.

Anonymous:1968:NVM

[Ano68d] Anonymous. Naamlijst van de monsterpunten in numerieke volgorde. (Dutch) [List of sample points in numerical order]. Aquatic Ecology, 2(3):92–93, September 1968. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185818.

Anonymous:1968:UIL

[Ano68e] Anonymous. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 2(2):41, June 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185806.

Anonymous:1968:VRDa

[Ano68f] Anonymous. Van de redactie. (Dutch) [From the Editors]. Aquatic Ecology, 2(2):41, June 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185805.

Anonymous:1968:VRDb

[Ano68g] Anonymous. Van de redactie. (Dutch) [From the Editors]. Aquatic Ecology, 2(3):81–82, September 1968. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185813.

Anonymous:1968:V

[Ano68h] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 2(2): 41, June 1968. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185807.

Anonymous:1969:MVB

[Ano69a] Anonymous. Mededelingen van het bestuur. (Dutch) [Board announcements]. Aquatic Ecology, 3(1):1, March 1969. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185834.

Anonymous:1969:UILa

[Ano69b] Anonymous. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 3(1):4, March 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185836.

Anonymous:1969:UILb

[Ano69c] Anonymous. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 3(2):33, July 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185846.

Anonymous:1969:UILc

[Ano69d] Anonymous. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 3(3):61, November 1969. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185852.

Anonymous:1969:VRDa

[Ano69e] Anonymous. Van de redactie. (Dutch) [From the Editors]. Aquatic Ecology, 3(1):1, March 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185833.

Anonymous:1969:VRDb

[Ano69f] Anonymous. Van de redactie. (Dutch) [From the Editors]. Aquatic Ecology, 3(2):33, July 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185845.

Anonymous:1969:Va

[Ano69g] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 3(1): 4, March 1969. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185837.

Anonymous:1969:Vb

[Ano69h] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 3(2):33–34, July 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185847.

Anonymous:1969:Vc

[Ano69i] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 3 (3):61–62, November 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185853.

Anonymous:1970:C

[Ano70a] Anonymous. Congressen. Aquatic Ecology, 4(3):130, November 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185870.

Anonymous:1970:LAP

[Ano70b] Anonymous. Leden en abonnees per 1-1-1971. (Dutch) [Members and subscribers as of 1-1-1971]. Aquatic Ecology, 4 (3):137–143, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185873.

Anonymous:1970:SVV

[Ano70c] Anonymous. Samenvattingen van voordrachten. (Dutch) [Summaries of lectures]. Aquatic Ecology, 4(3):144, November 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185874.

Anonymous:1970:S

[Ano70d] Anonymous. Slotconclusies. Aquatic Ecology, 4(1):46–47, March 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185867.

Anonymous:1970:UIL

[Ano70e] Anonymous. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 4(3):127–128, November 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185868.

Anonymous:1970:V

[Ano70f] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 4 (3):131, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185871.

Anonymous:1971:C

[Ano71a] Anonymous. Congressen. Aquatic Ecology, 5(2):68, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185940.

Anonymous:1971:V

[Ano71b] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 5(2):68–70, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185941.

Anonymous:1972:C

[Ano72a] Anonymous. Congressen. (Dutch) [congresses]. Aquatic Ecology, 6(4):160–162, December 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334723.

Anonymous:1972:DND

[Ano72b] Anonymous. Drinkwatervoorziening in Nederland. (Dutch) [Drinking water supply in The Netherlands]. Aquatic Ecology, 6(3):115, September 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02314879.

Anonymous:1972:LAP

[Ano72c] Anonymous. Leden en abonnees per 1-1-1973. (Dutch) [Members and subscribers as of 1-1-1973]. Aquatic Ecology, 6 (4):164–171, December 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334725.

Anonymous:1972:MVB

[Ano72d] Anonymous. Mededeling van het bestuur. (Dutch) [Board announcements]. Aquatic Ecology, 6(4):159, December 1972.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334720.

Anonymous:1972:R

[Ano72e] Anonymous. Rapporten. Aquatic Ecology, 6(4):159–160, December 1972. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334722.

Anonymous:1972:UIL

[Ano72f] Anonymous. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 6(4):159, December 1972. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334721.

Anonymous:1972:Va

[Ano72g] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 6(1):6–8, March 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02336204.

Anonymous:1972:Vb

[Ano72h] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 6(4): 162–163, December 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334724.

Anonymous:1972:VVW

[Ano72i] Anonymous. Verslag van het werkkamp 1971, van de Hydrobiologische Vereniging, gehouden in de Ooypolder. (Dutch) [Report of the workshop 1971, of the Hydrobiological Society, held in the Ooypolder]. Aquatic Ecology, 6(2):47, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02304092.

Anonymous:1973:Ma

[Ano73a] Anonymous. Miscellanea. Aquatic Ecology, 7(1):44–46, March 1973. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02279921.

Anonymous:1973:Mb

[Ano73b] Anonymous. Miscellanea. Aquatic Ecology, 7(3):136–139, September 1973. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02275615.

Anonymous:1974:M

[Ano74a] Anonymous. Miscellanea. Aquatic Ecology, 8(3):292–293, December 1974. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257509.

Anonymous:1974:S

[Ano74b] Anonymous. Symposia. Aquatic Ecology, 8(3):291, December 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257508.

Anonymous:1975:M

[Ano75a] Anonymous. Miscellanea. Aquatic Ecology, 9(3):141, December 1975. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02263335.

Anonymous:1975:S

[Ano75b] Anonymous. Symposia. Aquatic Ecology, 9(3):140, December 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02263334.

Anonymous:1976:P

[Ano76a] Anonymous. Publications. Aquatic Ecology, 10(2):123, September 1976. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02282220.

Anonymous:1976:SM

[Ano76b] Anonymous. Scientific meetings. Aquatic Ecology, 10(3):178, December 1976. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02263367.

Anonymous:1976:SE

[Ano76c] Anonymous. Symposia, etc. Aquatic Ecology, 10(2):123–124, September 1976. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02282221.

Anonymous:1977:C

[Ano77a] Anonymous. Congress. Aquatic Ecology, 11(2):62, September 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02265711.

Anonymous:1977:P

[Ano77b] Anonymous. Publication. Aquatic Ecology, 11(2):62–63, September 1977. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02265712.

Anonymous:1977:V

[Ano77c] Anonymous. Varia. (Dutch) [Various]. Aquatic Ecology, 11(1):30, April 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02282214.

Anonymous:1978:C

[Ano78a] Anonymous. Congresses. Aquatic Ecology, 12(1):61–62, April 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260796.

Anonymous:1978:E

[Ano78b] Anonymous. Errata. Aquatic Ecology, 12(2):160, July 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260720.

Anonymous:1978:Ma

[Ano78c] Anonymous. Miscellaneous. Aquatic Ecology, 12(2):153– 158, July 1978. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260719.

Anonymous:1978:Mb

[Ano78d] Anonymous. Miscellaneous. Aquatic Ecology, 12(3-4):351– 352, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259195.

Anonymous:1979:M

[Ano79] Anonymous. Meeting. Aquatic Ecology, 13(1):70, April 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260313.

Anonymous:1980:C

[Ano80] Anonymous. Congresses. Aquatic Ecology, 14(3):227, December 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260135.

Anonymous:1981:BRO

[Ano81a] Anonymous. Book review: Oproep aan de phytoplanktononderzoekers in Nederland door R. M. M. Roijackers (Sektie Hydrobiologie, Landbouwhogeschool, De Dreijen 12, 6703 BC Wageningen, Nederland). Aquatic Ecology, 15(3):200– 202, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255184.

Anonymous:1981:C

[Ano81b] Anonymous. Congresses. Aquatic Ecology, 15(3):203, December 1981. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255185.

Anonymous:1981:M

[Ano81c] Anonymous. Miscellaneous. Aquatic Ecology, 15(1-2):94, October 1981. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260264.

Anonymous:1982:C

[Ano82] Anonymous. Congresses. Aquatic Ecology, 16(1):116, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255422.

Anonymous:1983:E

[Ano83a] Anonymous. Editorial. Aquatic Ecology, 17(1):3, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255187.

Anonymous:1983:SEA

[Ano83b] Anonymous. Studies of the ecophysiology of algae in the laboratory of microbiology of the University of Amsterdam (summaries). Aquatic Ecology, 17(2):169–178, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02280825.

Anonymous:1984:Ea

[Ano84a] Anonymous. Editorial. Aquatic Ecology, 18(1):2, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256666.

Anonymous:1984:Eb

[Ano84b] Anonymous. Editorial. Aquatic Ecology, 18(2):84, December 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257046.

Anonymous:1984:Ma

[Ano84c] Anonymous. Miscellaneous. Aquatic Ecology, 18(1):81, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02256688.

Anonymous:1984:Mb

[Ano84d] Anonymous. Miscellaneous. Aquatic Ecology, 18(2):174, December 1984. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257057.

Anonymous:1985:M

[Ano85] Anonymous. Miscellaneous. Aquatic Ecology, 19(2):225–226, December 1985. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02270770.

Anonymous:1987:PWQ

[Ano87] Anonymous. Proceedings of the water quality research Loosdrecht Lakes Symposium. Aquatic Ecology, 21(1):111, July 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255460.

Anonymous:1988:Ma

[Ano88a] Anonymous. Miscellaneous. Aquatic Ecology, 22(1):106–107, August 1988. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256793.

Anonymous:1988:Mb

[Ano88b] Anonymous. Miscellaneous. Aquatic Ecology, 22(2):210, December 1988. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256826.

Anonymous:1989:C

[Ano89a] Anonymous. Congresses. Aquatic Ecology, 23(2):223, December 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256741.

Anonymous:1989:M

[Ano89b] Anonymous. Miscellaneous. Aquatic Ecology, 23(2):223, December 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256742.

Anonymous:1991:A

[Ano91] Anonymous. Acknowledgements. Aquatic Ecology, 25(2): 180, June 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291253.

Anonymous:1992:E

[Ano92a] Anonymous. Errata. Aquatic Ecology, 26(1):89, November 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02298030.

Anonymous:1992:LP

[Ano92b] Anonymous. List of participants. Aquatic Ecology, 26 (2-4):97–98, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255229.

Anonymous:1995:A

[Ano95a] Anonymous. Announcement. Aquatic Ecology, 29(2):197, June 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02336050.

Anonymous:1995:LP

[Ano95b] Anonymous. List of participants. Aquatic Ecology, 29(3-4):202, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02084216.

Anonymous:1996:A

[Ano96] Anonymous. Announcements. Aquatic Ecology, 30(1):69, July 1996. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02092149.

Anonymous:1997:BR

[Ano97a] Anonymous. Book reviews. Aquatic Ecology, 31(4):429–431, December 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017140324458.

Anonymous:1997:DSAa

[Ano97b] Anonymous. Dutch Society for Aquatic Ecology (NVAE). Aquatic Ecology, 31(1):111–112, March 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/A:
1017198602103.

Anonymous:1997:DSAb

[Ano97c] Anonymous. Dutch Society for Aquatic Ecology (NVAE). Aquatic Ecology, 31(3):345–346, September 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1017160503238.

Anonymous:1997:DSAc

[Ano97d] Anonymous. Dutch Society for Aquatic Ecology (NVAE). Aquatic Ecology, 31(4):433–434, December 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1017198522331.

Anonymous:1997:FIC

[Ano97e] Anonymous. Fourth International Crustacean Congress. Aquatic Ecology, 31(2):259, June 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1023/A:1017167809149.

Anonymous:1997:P

[Ano97f] Anonymous. Preface. Aquatic Ecology, 31(4):347, December 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1017187906769.

Anonymous:1998:AI

[Ano98a] Anonymous. Authors index. Aquatic Ecology, 32(4):378, December 1998. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017156010685.

Anonymous:1998:BRa

[Ano98b] Anonymous. Book reviews. Aquatic Ecology, 32(1):99–100, March 1998. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017141117402.

Anonymous:1998:BRb

[Ano98c] Anonymous. Book reviews. Aquatic Ecology, 32(1):100–101, March 1998. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017136900564.

Anonymous:1998:BRc

[Ano98d] Anonymous. Book reviews. Aquatic Ecology, 32(2):191, September 1998. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017109414550.

Anonymous:1998:DSA

[Ano98e] Anonymous. Dutch Society for Aquatic Ecology (NVAE). Aquatic Ecology, 32(4):379–380, December 1998. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1017241505878.

Anonymous:1998:P

[Ano98f] Anonymous. Preface. Aquatic Ecology, 32(2):105, September 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1017190217276.

Anonymous:1998:PST

[Ano98g] Anonymous. Preface — special topics. Aquatic Ecology, 32(1):1, March 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1017113232424.

Anonymous:1998:SI

[Ano98h] Anonymous. Society information. Aquatic Ecology, 32(3): 259–260, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1017146127369.

Anonymous:1998:VC

[Ano98i] Anonymous. Volume contents. Aquatic Ecology, 32(4): 375–377, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1017173626615.

Anonymous:1999:ANF

[Ano99a] Anonymous. Announcement — Netherlands–Flemish Ecological Society (NECOV). Aquatic Ecology, 33(4):401–402, December 1999. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017266429658.

Anonymous:1999:BR

[Ano99b] Anonymous. Book reviews. Aquatic Ecology, 33(1):117–118, March 1999. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017232827394.

Anonymous:1999:NSEa

[Ano99c] Anonymous. Netherlands Society for Ecology (NVE). Aquatic Ecology, 33(2):221–222, September 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1017290915805.

Anonymous:1999:NSEb

[Ano99d] Anonymous. Netherlands Society for Ecology (NVE). Aquatic Ecology, 33(3):323–324, October 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/A:
1017246825706.

Anonymous:1999:Pa

[Ano99e] Anonymous. Preface. Aquatic Ecology, 33(1):1, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1017287608720.

Anonymous:1999:Pb

[Ano99f] Anonymous. Preface. Aquatic Ecology, 33(3):223–233, October 1999. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017283930490.

Anonymous:2000:BRa

[Ano00a] Anonymous. Book review. Aquatic Ecology, 34(2):209–210, June 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017351209310.

Anonymous:2000:BRb

[Ano00b] Anonymous. Book review. Aquatic Ecology, 34(2):211–212, June 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017314313756.

Anonymous:2000:BRc

[Ano00c] Anonymous. Book review. Aquatic Ecology, 34(3):323–328, September 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1017302019174.

Anonymous:2000:NFE

[Ano00d]  Anonymous.  Netherlands–Flemish Ecological Society

(NECOV). Aquatic Ecology, 34(1):101–102, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1017269814080.

Anonymous:2000:P

[Ano00e] Anonymous. Preface. Aquatic Ecology, 34(1):1, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1017221402222.

Anonymous:2002:A

[Ano02] Anonymous. Announcement. Aquatic Ecology, 36(2):341, April 2002. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1016306802723.

Anonymous:2003:AI

[Ano03a] Anonymous. Author index. Aquatic Ecology, 37(4):457, October 2003. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000007017.25181.a9.

Anonymous:2003:NFE

[Ano03b]  Anonymous.  Netherlands–Flemish Ecological Society

(NECOV). Aquatic Ecology, 37(4):451–452, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000007080.50945.33.

Anonymous:2003:VC

[Ano03c] Anonymous. Volume contents. Aquatic Ecology, 37(4): 453–455, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/B:AECO.0000007016.53478.
1b.

Anonymous:2004:BR

[Ano04a] Anonymous. Book review. Aquatic Ecology, 38(1):101, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000021174.21138.69.

Anonymous:2004:E

[Ano04b] Anonymous. editorial. Aquatic Ecology, 38(1):1–2, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000021202.35158.85.

Anonymous:2004:IA

[Ano04c] Anonymous. Instructions for authors. Aquatic Ecology, 38(1):103–106, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000021176.46688.69.

Anonymous:2004:NFE

[Ano04d] Anonymous. Nederlands–Flemish Ecological Society (NECOV). Aquatic Ecology, 38(1):107–108, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/B:
AECO.0000021205.34883.38.

Anonymous:2004:P

[Ano04e] Anonymous. Preface. Aquatic Ecology, 38(2):109, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032134.94332.a5.

Anonymous:2004:SI

[Ano04f] Anonymous. Society information. Aquatic Ecology, 38 (2):349–350, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/B:AECO.0000032139.47806.
be.

Anonymous:2005:BR

[Ano05a] Anonymous. Book review. Aquatic Ecology, 39(1):119–120, March 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-4932-y.

Anonymous:2005:BRI

[Ano05b] Anonymous. Book review: Integrated assessment of running waters in Europe, Hering, D., P. F. M. Verdonschot, O. Moog & L. Sandin (Eds.) Developments in Hydrobiology 175/Hydrobiologia 516 (2004) Kluwer Academic Publishers, Dordrecht. ISBN 1-4020-1804-5, Euro 110. Aquatic Ecology, 39(3):377–378, September 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-1951-2.

Anonymous:2005:SIa

[Ano05c] Anonymous. Society information. Aquatic Ecology, 39 (2):261–262, June 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-5248-2.

Anonymous:2005:SIb

[Ano05d] Anonymous. Society information. Aquatic Ecology, 39(3): 379–380, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-1952-1.

Anonymous:2025:JN

[Ano25] Anonymous. Journal navigation. Aquatic Ecology, 59(1):??, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

Arcanjo-Oliveira:2024:EFE

[AOLLJ24] Betânia Arcanjo-Oliveira, Luciano B. Lima, and Dilermando P. Lima-Junior. Effect of fire and environmental temperature on the reproductive recruitment of neotropical freshwater turtles. Aquatic Ecology, 58(4):1161–1174, December 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10131-4.

Ashade:2024:AIC

[AOS^{+}24] Ahmeed Olalekan Ashade, Oluwafemi Sunday Obayori, Lateef Babatunde Salam, Muibat Omotola Fashola, and Francisca Obiageri Nwaokorie. Assessments of the impacts of chemical and hydrocarbon pollution on Ologe Lagoon water and its prokaryotic community structure. Aquatic Ecology, 58(3):741–757, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10102-9.

Acosta:2000:DGC

[AP00] Charles A. Acosta and Sue A. Perry. Differential growth of crayfish Procambarus alleni in relation to hydrological conditions in marl prairie wetlands of Everglades National Park, USA. Aquatic Ecology, 34(4):389–395, December 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011472828639.

Altman:2012:CIO

[AP12] Julia C. Altman and Hans W. Paerl. Composition of inorganic and organic nutrient sources influences phytoplankton community structure in the New River Estuary, North1 Carolina. Aquatic Ecology, 46(3):269–282, September 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9398-8.

Allnutt:1981:LSN

[APS81] F. C. Thomas Allnutt, Bruce C. Parker, and George M. Simmons, Jr. In situ nitrogen (C_{2}H_{2}) fixation in lakes of southern Victoria Land, Antarctica. Aquatic Ecology, 15(3): 99–109, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255169.

Angsupanich:1997:SVP

[AR97] Saowapa Angsupanich and Suphaphorn Rakkheaw. Seasonal variation of phytoplankton community in Thale Sap Songkhla, a lagoonal lake in southern Thailand. Aquatic Ecology, 30(4):297–307, May 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02085873.

Aditya:2005:FLG

[AR05] Gautam Aditya and Srimanta K. Raut. Feeding of the leech Glossiphonia weberi on the introduced snail Pomacea bridgesii in India. Aquatic Ecology, 39(4):465–471, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-9006-2.

Arst:2009:AOC

[AR09] Helgi Arst and Anu Reinart. Application of optical classifications to North European lakes. Aquatic Ecology, 43(4): 789–801, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9225-4.

Arakelova:2001:EIP

[Ara01] Katherine S. Arakelova. The evaluation of individual production and scope for growth in aquatic sow bugs (Asellus aquaticus). Aquatic Ecology, 35(1):31–42, March 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011446224456.

Avendano:2020:OAC

[ARG20] Otilio Avendaño, Álvaro Roura, and Ángel Guerra. Octopus americanus: a cryptic species of the O. vulgaris species complex redescribed from the Caribbean. Aquatic Ecology, 54(4): 909–925, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09778-6.

Abrial:2014:IHC

[ARP14] Elie Abrial, Ana P. Rabuffetti, and Aldo R. Paira. Influence of hydrological changes on the fish community in two lotic environments of the Middle Paraná Floodplain, Argentina. Aquatic Ecology, 48(3):337–349, September 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9488-x.

Arts:1999:BRP

[Art99] Michael T. Arts. Book review: Phytoplankton and Trophic Gradients. Edited by M. Alvarez-Cobelas, C. S. Reynolds, P. Sanchez-Castillo and J. Kristiansen. Aquatic Ecology, 33(2): 218–219, September 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009906930942.

Andersson:2006:CPM

[ASA06] Agneta Andersson, Kristina Samuelsson, and Jan Albertsson. Changes in the pelagic microbial food web due to artificial eutrophication. Aquatic Ecology, 40(3):299–313, September 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-006-9041-7.

Ashe:1992:CCP

[Ash92] P. Ashe. Corrections to the Chironomidae part of the Catalogue of Palaearctic Diptera. Aquatic Ecology, 26(2-4):215– 221, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255244.

Arrieira:2015:UFD

[ASLT15] Rodrigo Leite Arrieira, Leilane Talita Fatoreto Schwind, and Fábio Amodêo Lansac-Tôha. Use of functional diversity to assess determinant assembly processes of testate amoebae community. Aquatic Ecology, 49(4):561–571, December 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9546-z.

Afonina:2024:SFD

[AT24] E. Yu. Afonina and N. A. Tashlykova. Structural and functional diversity of plankton communities along lake salinity gradients. Aquatic Ecology, 58(3):717–740, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10101-w.

Agostinho:2007:IME

[ATB07] Angelo A. Agostinho, Sidinei M. Thomaz, and Solma L. S. M. A. Baltar. Influence of the macrophyte Eichhornia azurea on fish assemblage of the Upper Paraná River floodplain (Brazil). Aquatic Ecology, 41(4):611–619, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9122-2.

Anderson:1992:LPD

[AV92] N. John Anderson and Peter Vos. Learning from the past: Diatoms as palaeoecological indicators of changes in marine environments. Aquatic Ecology, 26(1):19–30, November 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02298025.

Adriaenssens:2007:ACT

[AVD07] V. Adriaenssens, P. F. M. Verdonschot, and N. De Pauw. Application of clustering techniques for the characterization of macroinvertebrate communities to support river restoration management. Aquatic Ecology, 41(3):387–398, September 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-2836-0.

Amorim:2019:SVM

[AVdNM19] Cihelio Alves Amorim, Celina Rebeca Valença, and Ariadne do Nascimento Moura. Seasonal variations of morphofunctional phytoplankton groups influence the top-down control of a cladoceran in a tropical hypereutrophic lake. Aquatic Ecology, 53(3):453–464, September 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09701-8.

Admiraal:1990:IES

[AvdV90] W. Admiraal and J. C. van der Vlugt. Impact of eutrophication on the silicate cycle of man-made basins in the Rhine delta. Aquatic Ecology, 24(1):23–36, July 1990. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256746.

Aylesworth:2015:RSP

[AXR15] L. A. Aylesworth, J. H. Xavier, and I. L. Rosa. Regional-scale patterns of habitat preference for the seahorse Hippocampus reidi in the tropical estuarine environment. Aquatic Ecology, 49(4):499–512, December 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
015-9542-3.

Anna:2009:DIE

[AYM09] Argyroudi Anna, Chatzinikolaou Yorgos, and Lazaridou Maria. Do intermittent and ephemeral Mediterranean rivers belong to the same river type? Aquatic Ecology, 43(2):465– 476, June 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9176-9.

Adamovich:2002:ICL

[AZD02] Vera V. Adamovich, Tatiana A. Zotina, and Andrey G. Degermendzhy. Interaction coefficients in the Lake Shira algal-bacterial community. Aquatic Ecology, 36(2):261–270, April 2002. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1015638420766.

Beltman:1997:RLA

[BA97] B. Beltman and C. Allegrini. Restoration of lost aquatic plant communities: New habitats for Chara. Aquatic Ecology, 30(4):331–337, May 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02085876.

Bunioto:2007:EFL

[BA07] Taís Cristina Bunioto and Marlene Sofia Arcifa. Effects of food limitation and temperature on cladocerans from a tropical Brazilian lake. Aquatic Ecology, 41(4):569–578, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9114-2.

Bukowski:2014:ECM

[BA14] Steven J. Bukowski and Josh R. Auld. The effects of calcium in mediating the inducible morphological defenses of a freshwater snail, Physa acuta. Aquatic Ecology, 48(1):85– 90, March 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9468-6.

Baars:1979:AIM

[Baa79] J. W. M. Baars. Autecological investigations on marine diatoms. I. Experimental results in biogeographical studies. Aquatic Ecology, 13(2-3):123–137, December 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284748.

Baars:1980:PSP

[Baa80] Martien A. Baars. The pelagic system in the photic zone of the tropical Atlantic. Aquatic Ecology, 14(3):223–224, December 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260130.

Baars:1981:AIM

[Baa81] J. W. M. Baars. Autecological investigations on marine diatoms. 2. Generation times of 50 species. Aquatic Ecology, 15(3):137–151, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255172.

Baars:1985:AIM

[Baa85] J. W. M. Baars. Autecological investigations on marine diatoms. 4: Biddulphia aurita (Lyngb.) Brebisson et Godey — a succession of spring diatoms. Aquatic Ecology, 19(2): 109–116, December 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02270756.

Baars:1988:AIMa

[Baa88a] J. W. M. Baars. Autecological investigations on marine diatoms. 5: Coscinodiscus concinnus W. Smith and Rhizosolenia setigera Brightwell. Aquatic Ecology, 22(2):147–155, December 1988. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256818.

Baars:1988:AIMb

[Baa88b] J. W. M. Baars. Autecological investigations on marine diatoms 6: Rhizosolenia robusta Norman, Rhizosolenia imbricata Brightwell and Rhizosolenia shrubsolei Cleve. Aquatic Ecology, 22(2):157–162, December 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256819.

Breda-Alves:2021:UER

[BAdOFC21] Fernanda Brêda-Alves, Valéria de Oliveira Fernandes, and Mathias Ahii Chia. Understanding the environmental roles of herbicides on cyanobacteria, cyanotoxins, and cyanoHABs. Aquatic Ecology, 55(2):347–361, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09849-2.

Bouvy:1994:STV

[BAG94] Marc Bouvy, Robert Arfi, and Daniel Guiral. Short-term variations of seston characteristics in a shallow tropical lagoon: Effect of wind-induced resuspension. Aquatic Ecology, 28(3-4):433–440, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334214.

Bakker:1972:MPV

[Bak72] C. Bakker. Milieu en plankton van het Veerse Meer, een tien jaar oud brakwatermeer in Zuidwest-Nederland. (Dutch) [Environment and plankton of the Lake Veerse. A ten-year-old brackish water lake in the southwest of The Netherlands]. Aquatic Ecology, 6(1):15–38, March 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02336207.

Bakker:1977:CPI

[Bak77] C. Bakker. Comparative plankton investigations in a brackish lake (L. Veere) and a brackish tidal estuary (Westerschelde) in the SW Netherlands. Aquatic Ecology, 11(1):18– 19, April 1977. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02282210.

Bakker:1978:SRA

[Bak78] C. Bakker. Some reflections about the structure of the pelagic zone of the brackish Lake Grevelingen (SW-Netherlands). Aquatic Ecology, 12(2):67–84, July 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260708.

Bakker:1979:PZR

[Bak79] C. Bakker. Phytoplankton/zooplankton relationships during spring in the brackish Lake Grevelingen 1976–1978 (III). Aquatic Ecology, 13(2-3):78–79, December 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284730.

Bakker:1980:DGV

[Bak80] C. Bakker. On the distribution of ‘Gonionemus vertens’ a. Agassiz (Hydrozoa, Limnomedusae), a new species in the eelgrass beds of Lake Grevelingen (S.W. Netherlands). Aquatic Ecology, 14(3):186–195, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260120.

Bakker:1994:RCS

[Bak94] C. Bakker. Resting cysts of some Chaetoceros species, identified and figured by A. Van der Werff. Aquatic Ecology, 28(1):71–75, March 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334247.

Balkis:2009:SVM

[Bal09] Neslihan Balkis. Seasonal variations of microphytoplankton assemblages and environmental variables in the coastal zone of Bozcaada Island in the Aegean Sea (NE Mediterranean Sea). Aquatic Ecology, 43(2):249–270, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9175-x.

Bandel:1975:ELS

[Ban75] Klaus Bandel. Embryonale und larvale Schale einiger Prosobranchier (Gastropoda, Mollusca) der Oosterschelde (Nordsee). (German) [Embryonic and larval shell of some prosobranchs (Gastropoda, Mollusca) of the Oosterschelde (North Sea)]. Aquatic Ecology, 9(1):3–22, April 1975. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02257515.

Baretta:1977:SFZ

[Bar77] J. W. Baretta. Seasonal fluctuations in the zooplankton of the Ems–Dollard estuary. Aquatic Ecology, 11(1):12–13, April 1977. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02282205.

Barnes:2021:PSM

[Bar21] R. S. K. Barnes. Patterns of seagrass macrobenthic biodiversity in the warm-temperate Knysna estuarine bay, Western Cape: a review. Aquatic Ecology, 55(2):327–345, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09848-3.

Bashinskiy:2020:BLR

[Bas20] I. V. Bashinskiy. Beavers in lakes: a review of their ecosystem impact. Aquatic Ecology, 54(4):1097–1120, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09796-4.

Bouvy:1994:RBF

[BASJ94] Marc Bouvy, Robert Arfi, and Lucien Saint-Jean. Role of bacteria as food for zooplankton in a eutrophic tropical pond (Ivory Coast). Aquatic Ecology, 28(2):167–174, June 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02333987.

Bauder:2000:IES

[Bau00] Ellen T. Bauder. Inundation effects on small-scale plant distributions in San Diego, California vernal pools. Aquatic Ecology, 34(1):43–61, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009916202321.

Beltman:1978:WQR

[BB78] B. Beltman and W. Bleuten. Water quality research in the area of the river Kromme Rijn. Aquatic Ecology, 12(1):39– 51, April 1978. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260792.

Bazzanti:1987:CWQ

[BB87] Marcello Bazzanti and Franco Bambacigno. Chironomids as water quality indicators in the River Mignone (Central Italy). Aquatic Ecology, 21(2):213–222, December 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255447.

Baiao:2000:EFD

[BB00] Célia Baião and Maria-José Boavida. Environmental factors determining the structure of rotifer communities in a rivershed reservoir. Aquatic Ecology, 34(4):369–377, December 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011406018683.

Beardsley:2012:CTN

[BB12] Helen Beardsley and J. Robert Britton. Contribution of temperature and nutrient loading to growth rate variation of three cyprinid fishes in a lowland river. Aquatic Ecology, 46 (1):143–152, March 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9387-3.

Brentjens:2023:BSS

[BB23] Emma T. Brentjens and Anika R. Bratt. Beneath the surface: spatial and temporal trends in water quality and its impacts on algal community composition in the Albemarle Sound, North Carolina. Aquatic Ecology, 57(2):243–262, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10008-y.

Belle:2017:RVA

[BBD17] Simon Belle, Virgile Baudrot, and Vasilis Dakos. Rising variance and abrupt shifts of subfossil chironomids due to eutrophication in a deep sub-alpine lake. Aquatic Ecology, 51(2):307–319, June 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9618-3.

Benelli:2017:RLB

[BBF17] Sara Benelli, Marco Bartoli, and Elisa Anna Fano. Rare but large bivalves alter benthic respiration and nutrient recycling in riverine sediments. Aquatic Ecology, 51(1):1–16, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-016-9590-3.

Bjelke:2005:TNS

[BBH05] Ulf Bjelke, Irene M. Bohman, and Jan Herrmann. Temporal niches of shredders in lake littorals with possible implications on ecosystem functioning. Aquatic Ecology, 39(1):41– 53, March 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-3524-1.

Bernot:2009:NCR

[BBM09] Melody J. Bernot, Randall J. Bernot, and James T. Morris. Nutrient cycling relative to δ^{15}N and δ^{13}C natural abundance in a coastal wetland with long-term nutrient additions. Aquatic Ecology, 43(4):803–813, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9221-8.

Belolipetsky:2010:NMV

[BBM10] Pavel V. Belolipetsky, Victor M. Belolipetskii, and Wolf M. Mooij. Numerical modeling of vertical stratification of Lake Shira in summer. Aquatic Ecology, 44(3):561–570, September 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-010-9330-z.

Braun:2021:TBE

[BBM21] Lisa-Marie Braun, Sandra Brucet, and Thomas Mehner. Top-down and bottom-up effects on zooplankton size distribution in a deep stratified lake. Aquatic Ecology, 55(2):527– 543, June 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09843-8.

Bhatt:2008:BDE

[BBP08] Jay Prakash Bhatt, Arun Bhaskar, and Maharaj Krishan Pandit. Biology, distribution and ecology of Didymosphenia geminata (Lyngbye) Schmidt an abundant diatom from the Indian Himalayan rivers. Aquatic Ecology, 42(3):347–353, September 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9106-2.

Birks:2001:RED

[BBR01] Hilary H. Birks, H. J. B. Birks, and M. Ramdani. Recent ecosystem dynamics in nine North African lakes in the CAS-SARINA Project. Aquatic Ecology, 35(3-4):461–478, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011997820776.

Best:1978:TIS

[BBV78] Elly P. H. Best, Marianne C. I. Blaauboer, and H. Verdouw. Towards an integrated study of the ecosystem of Lake Vechten. Aquatic Ecology, 12(2):107–118, July 1978. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260711.

Brotas:1995:MPP

[BC95] Vanda Brotas and Fernando Catarino. Microphytobenthos primary production of Tagus estuary intertidal flats (Portugal). Aquatic Ecology, 29(3-4):333–339, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084232.

Booth:2023:IDQ

[BC23] Michael T. Booth and Carolynn S. Culver. Invasion dynamics of quagga mussels within a Southern California reservoir and its spatially intermittent watershed. Aquatic Ecology, 57(2):499–522, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10025-x.

Bellisario:2012:IBN

[BCN12] Bruno Bellisario, Fulvio Cerfolli, and Giuseppe Nascetti. The interplay between network structure and functioning of detritus-based communities in patchy aquatic environment. Aquatic Ecology, 46(4):431–441, December 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9412-1.

Bergey:2010:SCA

[BCP10] Elizabeth A. Bergey, Joshua T. Cooper, and Barret C. Phillips. Substrate characteristics affect colonization by the bloom-forming diatom Didymosphenia geminata. Aquatic Ecology, 44(1):33–40, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9247-6.

Bakker:1974:CBW

[BD74] C. Bakker and N. De Pauw. Comparison of brackish water plankton assemblages of identical salinity ranges in an estuarine tidal (Westerschelde) and stagnant (Lake Veere) environment (S.W.-Netherlands). I. Phytoplankton. Aquatic Ecology, 8(1-2):179–189, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02254918.

Barquin:2009:PCD

[BD09] José Barquín and Russell G. Death. Physical and chemical differences in karst springs of Cantabria, northern Spain: do invertebrate communities correspond? Aquatic Ecology, 43(2):445–455, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9170-2.

Bo:2020:CBD

[BDF20] Tiziano Bo, Alberto Doretto, and Stefano Fenoglio. Contribution of beta diversity in shaping stream macroinvertebrate communities among hydro-ecoregions. Aquatic Ecology, 54 (4):957–971, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09786-6.

Bernot:2006:TKI

[BDG06] Randall J. Bernot, Walter K. Dodds, and Christopher S. Guy. Temperature and kairomone induced life history plasticity in coexisting Daphnia. Aquatic Ecology, 40(3):361–372, September 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-006-9035-5.

Brown:2010:EHK

[BDG10] Kenneth M. Brown, Wesley Daniel, and Gerry George. The effect of Hurricane Katrina on the mussel assemblage of the Pearl River, Louisiana. Aquatic Ecology, 44(1):223–231, March 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9255-6.

Balvert:2009:ZSD

[BDH09] Sheree F. Balvert, Ian C. Duggan, and Ian D. Hogg. Zooplankton seasonal dynamics in a recently filled mine pit lake: the effect of non-indigenous Daphnia establishment. Aquatic Ecology, 43(2):403–413, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9165-z.

Bernatowicz:2021:PPD

[BDP21a] Piotr Bernatowicz, Piotr Dawidowicz, and Joanna Pijanowska. Phenotypic plasticity and developmental noise in hybrid and parental clones of Daphnia longispina complex. Aquatic Ecology, 55(4):1179–1188, December 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09898-7.

Bernatowicz:2021:PDS

[BDP21b] Piotr Bernatowicz, Piotr Dawidowicz, and Joanna Pijanowska. Plasticity in depth selection behavior and heat shock proteins in Daphnia. Aquatic Ecology, 55(4):1171– 1178, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09895-w.

Battauz:2017:MDV

[BdPP17] Yamila S. Battauz, Susana B. José de Paggi, and Juan C. Paggi. Macrophytes as dispersal vectors of zooplankton resting stages in a subtropical riverine floodplain. Aquatic Ecology, 51(2):191–201, June 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
016-9610-3.

Bacher:1997:ACM

[BDR97] C. Bacher, P. Duarte, and O. Raillard. Assessment and comparison of the Marennes–Oléron Bay (France) and Carlingford Lough (Ireland) carrying capacity with ecosystem models. Aquatic Ecology, 31(4):379–394, December 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009925228308.

Becker:2009:DVP

[BdSCH09] Vanessa Becker, Luciana de Souza Cardoso, and Vera Lúcia M. Huszar. Diel variation of phytoplankton functional groups in a subtropical reservoir in southern Brazil during an autumnal stratification period. Aquatic Ecology, 43(2):285– 293, June 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9164-0.

Boelter:2020:EHM

[BdSM20] Thaíse Boelter, Fernanda Marcelia dos Santos, and Leonardo Maltchik. Effects of hydroperiod on morphology of tadpoles from highland ponds. Aquatic Ecology, 54(4):1145–1153, December 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09799-1.

Best:1987:P

[BED87] E. P. H. Best, D. Eisma, and C. Davids. Publications. Aquatic Ecology, 21(2):235–246, December 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255449.

Beltman:1976:IRC

[Bel76a] B. Beltman. Introduction to the research carried out in a ditch of the “Oostelijke Binnenpolder” at Tienhoven during a camp of the Netherlands Hydrobiological Society. Aquatic Ecology, 10(1):31–34, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02308497.

Beltman:1976:SPC

[Bel76b] B. Beltman. Some physical and chemical characteristics of the ditch at Tienhoven. Aquatic Ecology, 10(1):35–36, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02308498.

Beltman:1979:RDC

[Bel79] B. Beltman. Research in ditches and canals in the province of Utrecht (The Netherlands). Aquatic Ecology, 13(1):29, April 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260305.

Beltman:1987:EWC

[Bel87] B. Beltman. Effects of weed control on species composition of aquatic plants and bank plants and macrofauna in ditches. Aquatic Ecology, 21(2):171–179, December 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255443.

Bao:2018:SBS

[BEQP18] Francielli Bao, Tracy Elsey-Quirk, and Arnildo Pott. Seed bank of seasonally flooded grassland: experimental simulation of flood and post-flood. Aquatic Ecology, 52(1):93–105, March 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-017-9647-y.

Berger:1975:EOO

[Ber75] C. Berger. Eutrophication and occurrence of Oscillatoria agardhii Gom. in the lakes bordering on Flevoland. Aquatic Ecology, 9(2):60–61, September 1975. CODEN AQECF9.

ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02257637.

Berger:1984:SPP

[Ber84] C. Berger. In situ primary production measured with 250 ML and 130 ML bottles. Aquatic Ecology, 18(1):17–22, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02256669.

Bergey:2024:INN

[Ber24] Elizabeth A. Bergey. The impacts of non-native watercress in Oklahoma spring ecosystems. Aquatic Ecology, 58(2):411– 427, June 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10081-3.

Best:1976:NCA

[Bes76] P. H. Best. Nutrient content of the aquatic macrophytes Elodea canadensis and Ceratophyllum in the course of the year. Aquatic Ecology, 10(1):15–16, April 1976. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02308493.

Best:1979:PCD

[Bes79] Elly P. H. Best. Photosynthesis in Ceratophyllum demersum. Carbon fixation rates in relation to the plants’ physiological stage, and the contents of chlorophylland non-structural carbohydrates. Aquatic Ecology, 13(2-3):112, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284742.

Best:1981:SAM

[Bes81] Elly P. H. Best. The submerged aquatic macrophytes in Lake Maarsseveen i: Species composition, spatial distribution and productivity. Aquatic Ecology, 15(1-2):72–81, October 1981. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260260.

Best:1987:SML

[Bes87] Elly P. H. Best. The submerged macrophytes in Lake Maarsseveen I: Changes in species composition and biomass over a six year period. Aquatic Ecology, 21(1):55–60, July 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255455.

Beukema:1971:WVP

[Beu71a] J. J. Beukema. De werkgroep voedselketen- en productieonderzoek. (Dutch) [The food chain and production research working group]. Aquatic Ecology, 5(2):98, June 1971. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185947.

Beukema:1971:KTV

[Beu71b] J. J. Beukema. Een kwantitatieve test van de efficientie van de van veen-bodemhapper. (Dutch) [A quantitative test of the efficiency of the peat grab]. Aquatic Ecology, 5(2):98– 99, June 1971. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185948.

Beukema:1974:EPR

[Beu74] J. J. Beukema. An evaluation of productivity research in the aquatic environment. Some comments (abstract). Aquatic Ecology, 8(3):269, December 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02257501.

Bell:1991:HBP

[BEZ91] Russell T. Bell, Rolf Erikson, and Argentina Zelaya. Heterotrophic bacterial production in Lake Xolotlán (Managua) during 1988–1989. Aquatic Ecology, 25(2):145–149, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02291246.

Brownscombe:2012:RED

[BF12] Jacob W. Brownscombe and Michael G. Fox. Range expansion dynamics of the invasive round goby (Neogobius melanostomus) in a river system. Aquatic Ecology, 46 (2):175–189, June 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9390-3.

Braghin:2015:ZDD

[BFB15] Louizi S. M. Braghin, Bruno R. S. Figueiredo, and Claudia C. Bonecker. Zooplankton diversity in a dammed river basin is maintained by preserved tributaries in a tropical floodplain. Aquatic Ecology, 49(2):175–187, June 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9514-7.

Bouletreau:2021:GFP

[BFS21] Stéphanie Boulêtreau, Thomas Fauvel, and Frédéric Santoul. “The giants’ feast”: predation of the large introduced European catfish on spawning migrating allis shads. Aquatic Ecology, 55(1):75–83, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09811-8.

Boeker:2015:EII

[BG15] Carolin Boeker and Juergen Geist. Effects of invasive and indigenous amphipods on physico-chemical and microbial properties in freshwater substrates. Aquatic Ecology, 49(4): 467–480, December 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9539-y.

Burdloff:2000:CEP

[BGC00] Didier Burdloff, Stephane Gasparini, and Jacques Castel. Is the copepod egg production in a highly turbid estuary (the Gironde, France) a function of the biochemical composition of seston? Aquatic Ecology, 34(2):165–175, June 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009903702667.

Bakker:1985:MIP

[BGK85] C. Bakker, R. D. Gulati, and K. Kersting. The measurement of ingestion of phytoplankton by zooplankton: Techniques, problems and recommendations. Aquatic Ecology, 19(1):3–4, November 1985. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255086.

Belolipetskii:2002:MMC

[BGK02] V. M. Belolipetskii, S. N. Genova, and L. A. Kompaniets. Mathematical models and computer programmes for the investigation of hydrophysical processes in Lake Shira. Aquatic Ecology, 36(2):143–152, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1015608110845.

Bergamino:2013:MFW

[BGL13] Leandro Bergamino, Julio Gómez, and Diego Lercari. Major food web properties of two sandy beaches with contrasting morphodynamics, and effects on the stability. Aquatic Ecology, 47(3):253–261, September 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9440-5.

Becares:2008:ENF

[BGM08] Eloy Bécares, Joan Gomá, and Brian Moss. Effects of nutrients and fish on periphyton and plant biomass across a European latitudinal gradient. Aquatic Ecology, 42(4):561–574, December 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9126-y.

Bagagnan:2024:OMC

[BGRM^{+}24] Sadia Bagagnan, Sabrina Guerin-Rechdaoui, Anthony Marconi, Vincent Rocher, Stéphanie Giusti-Miller, Régis Moilleron, and My Dung Jusselme. Overview of microbial communities in the surface water of the Seine River to understand their response to climate change and human activities. Aquatic Ecology, 58(4):1067–1089, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10124-3.

Bakker:1971:VVW

[BH71] C. Bakker and G. R. Heerebout. De verontreiniging van de Westerschelde. (Dutch) [Pollution of the Western Scheldt].

Aquatic Ecology, 5(1):53–64, March 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185888.

Boon:1982:PMS

[BH82] Jaap J. Boon and Johan Haverkamp. Pyrolysis mass spectrometry of intact and decomposed leaves of Nuphar variegatum and Zostera marina, and some archeological eelgrass samples. Aquatic Ecology, 16(1):71–82, June 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255415.

Berg:1992:RCE

[BH92] Martin B. Berg and Ronald A. Hellenthal. The role of Chironomidae in energy flow of a lotic ecosystem. Aquatic Ecology, 26(2-4):471–476, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255277.

Bakker:2016:IWL

[BH16] Elisabeth S. Bakker and Sabine Hilt. Impact of water-level fluctuations on cyanobacterial blooms: options for management. Aquatic Ecology, 50(3):485–498, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9556-x.

Bilodeau:2022:CCA

[BH22] Stephanie M. Bilodeau and Mark E. Hay. Chemical cues affecting recruitment and juvenile habitat selection in marine versus freshwater systems. Aquatic Ecology, 56(2):339– 360, June 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09905-x.

Burkert:2001:EMF

[BHB01] U. Burkert, P. Hyenstrand, and P. Blomqvist. Effects of the mixotrophic flagellate Ochromonas sp. on colony formation in Microcystis aeruginosa. Aquatic Ecology, 35(1):11– 17, March 2001. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1011454313607.

Bray:2016:PPI

[BHG16] Jonathan Bray, Jon S. Harding, and Philippe Gerbeaux. Physicochemical predictors of the invasive diatom Didymosphenia geminata at multiple spatial scales in New Zealand rivers. Aquatic Ecology, 50(1):1–14, March 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9543-2.

Borza:2018:CND

[BHG18a] Péter Borza, Thomas Huber, and Wolfram Graf. Correction to: Niche differentiation among invasive Ponto–Caspian Chelicorophium species (Crustacea, Amphipoda, Corophiidae) by food particle size. Aquatic Ecology, 52(2-3):191–192, September 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-018-9659-2. See [BHG18b].

Borza:2018:NDA

[BHG18b] Péter Borza, Thomas Huber, and Wolfram Graf. Niche differentiation among invasive Ponto–Caspian Chelicorophium species (Crustacea, Amphipoda, Corophiidae) by food particle size. Aquatic Ecology, 52(2-3):179–190, September 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s10452-018-9653-8. See correction [BHG18a].

Blank:2009:EWC

[BHL09] Kätlin Blank, Juta Haberman, and Reet Laugaste. Effect of winter conditions on spring nutrient concentrations and plankton in a large shallow Lake Peipsi (Estonia/ Russia). Aquatic Ecology, 43(3):745–753, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9283-2.

Berger:2003:EFA

[BHM03] Rita Berger, Emma Henriksson, and Torleif Malm. Effects of filamentous algae and deposited matter on the survival of Fucus vesiculosus L. germlings in the Baltic Sea. Aquatic Ecology, 37(1):1–11, January 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1022136900630.

Bonacina:2005:MPD

[BHP05] Carla Bonacina, Waleed Hamza, and Andrea Pasteris. Modelling the population dynamics of Daphnia obtusa (Kurz) in Lake Orta (N. Italy) under pre- and post-liming conditions. Aquatic Ecology, 39(1):93–106, March 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-3855-y.

Bruchner-Huttemann:2020:MSH

[BHPT20] Henrike Brüchner-Hüttemann, Christoph Ptatscheck, and Walter Traunspurger. Meiofauna in stream habitats: temporal dynamics of abundance, biomass and secondary production in different substrate microhabitats in a first-order stream. Aquatic Ecology, 54(4):1079–1095, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09795-5.

Brock:1982:LSS

[BHV82] Th. C. M. Brock, C. A. M. Huijbregts, and M. A. Vlassak. In situ studies on the breakdown of Nymphoides peltata (Gmel.) O. Kuntze (Menyanthaceae); some methodological aspects of the litter bag technique. Aquatic Ecology, 16(1):35–49, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255411.

Boedeker:2009:NML

[BI09] Christian Boedeker and Anne Immers. No more lake balls (Aegagropila linnaei Kützing, Cladophorophyceae, Chlorophyta) in The Netherlands? Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9231-1.

BijDeVaate:1975:OBG

[Bij75] A. Bij De Vaate. The occurrence of the blue-green alga Oscillatoria agardhii Gom. in some waters discharging into the Lauwersmeer (The Netherlands). Aquatic Ecology, 9(2): 55–59, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257636.

Billen:1973:EED

[Bil73] Gilles Billen. Étude de l’ecometabolisme du mercure dans un milieu d’eau douce. (French) [Study of the ecometabolism of mercury in a freshwater environment]. Aquatic Ecology, 7(2):60–68, June 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02282199.

Birnbaum:1978:ELS

[Bir78] E. L. Birnbaum. Estimating in situ algal production rates with the help of light measurements and experimentally measured production rates. Aquatic Ecology, 12(2):127–133, July 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260713.

Beklioglu:1999:BRP

[BJ99] Meryem Beklioglu and Erik Jeppesen. Behavioural response of plant-associated Eurycercus lamellatus (Ö. F. Müller) to different food sources and fish cues. Aquatic Ecology, 33(2): 167–173, September 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009934828802.

Bjornsson:2004:CCV

[BJ04] Halldór Björnsson and Trausti Jónsson. Climate and climatic variability at Lake Myvatn. Aquatic Ecology, 38(2):129–144, June 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000032061.51508.e6.

Boros:2012:BLV

[BJJ12] Gergely Boros, Jussi Jyväsjärvi, and Roger I. Jones. Between-lake variation in the elemental composition of roach (Rutilus rutilus L.). Aquatic Ecology, 46(4):385–394, December 2012. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-012-9402-3.

Buser:2012:CEP

[BJS12] Claudia C. Buser, Mieke Jansen, and Piet Spaak. Combined exposure to parasite and pesticide causes increased mortality in the water flea Daphnia. Aquatic Ecology, 46(2):261– 268, June 2012. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-012-9397-9.

Behera:2022:BSD

[BJS22] Pralaya Ranjan Behera, M. A. Jishnudev, and Raju Saravanan. Biometry and size distribution of Crambionella annandalei and Chrysaora spp. along the coast of Andhra Pradesh, India. Aquatic Ecology, 56(1):227–237, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09911-z.

Bourne:2011:BFP

[BKC11] Christina M. Bourne, Dan G. Kehler, and David Cote. Barriers to fish passage and barriers to fish passage assessments: the impact of assessment methods and assumptions on barrier identification and quantification of watershed connectivity. Aquatic Ecology, 45(3):389–403, September 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9362-z.

Bleeker:1992:ELZ

[BKD92] E. A. J. Bleeker, M. H. S. Kraak, and C. Davids. Ecotoxicity of lead to the zebra mussel Dreissena polymorpha, Pallas. Aquatic Ecology, 25(3):233–236, August 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270808.

Barbosa:2024:TSV

[BKD^{+}24] Luciana Gomes Barbosa, Mihir R. Kulkarni, Ênio Wocyli Dantas, Danielle Lima de Oliveira, Fátima dos Santos Silva, and Jascieli Carla Bortolini. Temporal and spatial variations drive the phytoplankton communities in rock pools in a semi-arid region. Aquatic Ecology, 58(3):659–674, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10097-3.

Blonska:2016:CPA

[BKG16] D. Błońska, J. Kobak, and J. Grabowska. Can the presence of alien Ponto–Caspian gobies affect shelter use by the native European bullhead? Aquatic Ecology, 50(4):653–665, December 2016. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9584-1.

BijDeVaate:1992:MEV

[BKO92] Abraham Bij De Vaate, Alexander Klink, and Frank Oosterbroek. The mayfly, Ephoron virgo (Olivier), back in the Dutch parts of the rivers Rhine and Meuse. Aquatic Ecology, 25(3):237–240, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02270809.

Bernat:1994:TVB

[BKW94] N. Bernát, B. Köpcke, and K. Wolfstein. Tidal variation in bacteria, phytoplankton, zooplankton, mysids, fish and suspended particulate matter in the turbidity zone of the Elbe estuary; interrelationships and causes. Aquatic Ecology, 28(3-4):467–476, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334218.

Blake:1982:CDP

[Bla82] G. Blake. Characterization of decomposition of plant material in an alpine lake. Aquatic Ecology, 16(1):5–9, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255408.

Banks:2023:IAP

[BLB^{+}23] L. K. Banks, I. Lavoie, M. P. Boreux, S. L. Kroeze, N. Gotkowski, C. E. Robinson, J. W. Roy, and A. G. Yates. Intra-annual patterns in biofilm communities and cellulose decomposition in a headwater stream network with spatially variable groundwater inputs. Aquatic Ecology, 57(3): 681–700, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10038-6.

Broadhurst:2012:SEH

[BLC12] Ben T. Broadhurst, Mark Lintermans, and Rhian C. Clear. Spatial ecology and habitat use of two-spined blackfish Gadopsis bispinosus in an upland reservoir. Aquatic Ecology, 46(3):297–309, September 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
012-9401-4.

Brett:1999:NCB

[BLG99] Michael T. Brett, Fred S. Lubnow, and Charles R. Goldman. Nutrient control of bacterioplankton and phytoplankton dynamics. Aquatic Ecology, 33(2):135–145, September 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009998915078.

Blok:1970:DGV

[Blo70] E. Blok. Diuron: Gevolgen voor het aquatisghe milieu en een regeling voor experimenten. (Dutch) [Diuron: Consequences for the aquatic environment and a scheme for experiments]. Aquatic Ecology, 4(1):10–13, March 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185861.

Blok:1983:ETM

[Blo83] J. Blok. Ecotoxicity tests with microorganisms. Aquatic Ecology, 17(1):87, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255196.

Bierbach:2016:HVU

[BLP16] David Bierbach, Kate L. Laskowski, and Martin Plath. Highly variable, unpredictable activity patterns in invasive, but not native amphipod species. Aquatic Ecology, 50 (2):261–271, June 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9573-4.

Billen:1988:MMP

[BLS88] G. Billen, C. Lancelot, and P. Servais. Modelling microbial processes (phyto- and bacterioplankton) in the Schelde estuary. Aquatic Ecology, 22(1):43–55, August 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256781.

Bakker:1970:SVD

[BM70] C. Bakker and L. Mur. Samenvatting van de discussie. (Dutch) [Discussion summary]. Aquatic Ecology, 4(1):43– 46, March 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185866.

Baretta:1985:EPU

[BM85] J. W. Baretta and J. F. P. Malschaert. Experimental problems using electronic particle counters. Aquatic Ecology, 19 (1):21–27, November 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255089.

Beklioglu:1998:ETT

[BM98] Meryem Beklioglu and Brian Moss. The effects of tench (Tinca tinca (L.)) and sticklebacks (Gasterosteus aculeatus L.) on planktonic and benthic communities in mesocosms in a shallow lake. Aquatic Ecology, 32(3):229–240, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009946505397.

Balayla:2003:SPP

[BM03] D. J. Balayla and B. Moss. Spatial patterns and population dynamics of plant-associated microcrustacea (Cladocera) in an English shallow lake (Little Mere, Cheshire). Aquatic Ecology, 37(4):417–435, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/B:
AECO.0000007045.85315.dc.

Barik:2000:DPU

[BMA00] Shyam Kumar Barik, Snehasish Mishra, and Subbanna Ayyappan. Decomposition patterns of unprocessed and processed lignocellulosics in a freshwater fish pond. Aquatic Ecology, 34(2):185–204, June 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009981319515.

Bormans:2016:CIP

[BMJ16] Myriam Bormans, Blahoslav Maršálek, and Daniel Jančula. Controlling internal phosphorus loading in lakes by physical methods to reduce cyanobacterial blooms: a review. Aquatic Ecology, 50(3):407–422, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9564-x.

Benitez:2015:OPF

[BMO15] Jean-Philippe Benitez, Billy Nzau Matondo, and Michaël Ovidio. An overview of potamodromous fish upstream movements in medium-sized rivers, by means of fish passes monitoring. Aquatic Ecology, 49(4):481–497, December 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9541-4.

Balseiro:2007:DDA

[BMR07] Esteban Balseiro, Beatriz Modenutti, and Mariana Reissig. Daphnia distribution in Andean Patagonian lakes: effect of low food quality and fish predation. Aquatic Ecology, 41(4): 599–609, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9113-3.

Brancelj:2020:GFC

[BMS20] Anton Brancelj, Nataša Mori, and Fabio Stoch. The groundwater fauna of the Classical Karst: hydrogeological indicators and descriptors. Aquatic Ecology, 54(1):205–224, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09737-w.

Bushi:2024:SVL

[BN24] Dhoni Bushi and Gibji Nimasow. Seasonal variation of limnological parameters and phytoplankton dynamics in the high-altitude oligotrophic lakes of Tawang district, Arunachal Pradesh (India). Aquatic Ecology, 58(3):801–818, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10106-5.

Bucci:2012:MPL

[BNMH12] Vanni Bucci, Daliangelis Nunez-Milland, and Ferdi L. Hellweger. Microscale patchiness leads to large and important intraspecific internal nutrient heterogeneity in phytoplankton. Aquatic Ecology, 46(1):101–118, March 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9384-6.

Bezerra-Neto:2007:DVM

[BNPC07] José Fernandes Bezerra-Neto and Ricardo Motta Pinto-Coelho. Diel vertical migration of the copepod Thermocyclops inversus (Kiefer, 1936) in a tropical reservoir: the role of oxygen and the spatial overlap with Chaoborus. Aquatic Ecology, 41(4):535–545, December 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9119-x.

Bartoli:2003:MAF

[BNV03] Marco Bartoli, Daniele Nizzoli, and Pierluigi Viaroli. Microphytobenthos activity and fluxes at the sediment-water interface: interactions and spatial variability. Aquatic Ecology, 37(4):341–349, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000007040.43077.5f.

Boldrocchi:2020:EWS

[BOB20] G. Boldrocchi, Moussa Omar, and R. Bettinetti. The ecology of the whale shark in Djibouti. Aquatic Ecology, 54(2):535– 551, June 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09758-w.

Bochert:1997:MVP

[Boc97] Ralf Bochert. Marenzelleria viridis (Polychaeta: Spionidae): a review of its reproduction. Aquatic Ecology, 31(2):163– 175, June 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009951404343.

Boers:1986:SPR

[Boe86] P. C. M. Boers. Studying the phosphorus release from the Loosdrecht Lakes sediments, using a continuous flow system. Aquatic Ecology, 20(1-2):51–60, November 1986. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291150.

Bolsunovsky:2004:ARA

[Bol04] Alexander Bolsunovsky. Artificial radionuclides in aquatic plants of the Yenisei River in the area affected by effluents of a Russian plutonium complex. Aquatic Ecology, 38(1):57– 62, March 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000020950.43944.ec.

Bologna:2007:IDP

[Bol07] Paul A. X. Bologna. Impact of differential predation potential on eelgrass (Zostera marina) faunal community structure. Aquatic Ecology, 41(2):221–229, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9059-x.

Borsje:1973:ETI

[Bor73] W. J. Borsje. Experimental taxonomic investigations on algae. Aquatic Ecology, 7(1):28–35, March 1973. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02279918.

Bouchard:2007:EOM

[Bou07] Virginie Bouchard. Export of organic matter from a coastal freshwater wetland to Lake Erie: an extension of the outwelling hypothesis. Aquatic Ecology, 41(1):1–7, March 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9044-4.

Buskens:1992:NEC

[BP92] R. F. M. Buskens and H. K. M. Moller Pillot. The Netherlands as an environment for chironomid fauna. Aquatic Ecology, 26(2-4):223–228, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255245.

Benns:1994:EPC

[BP94] E. J. Benns and D. A. Pilgrim. The effect of particle characteristics on the beam attenuation coefficient and output from an optical backscatter sensor. Aquatic Ecology, 28(3-4): 245–248, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334191.

Bhakare:2021:OFT

[BP21] Kalyani C. R. Bhakare and Kalpana Pai. An overview of freshwater Tardigrada in Northern Western Ghats of India. Aquatic Ecology, 55(4):1327–1338, December 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09886-x.

Birks:2001:PRN

[BPA01] Hilary H. Birks, Sylvia M. Peglar, and H. M. A. Abdelzaher. Palaeolimnological responses of nine North African lakes in the CASSARINA Project to recent environmental changes and human impact detected by plant macrofossil, pollen, and faunal analyses. Aquatic Ecology, 35(3-4):405–430, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1013265310922.

Burke:2015:CAA

[BPD15] Samantha M. Burke, Anurani D. Persaud, and Peter J. Dillon. A case against acidifying freshwater macrophytes prior to C and N stable isotope analysis. Aquatic Ecology, 49 (2):251–261, June 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9519-2.

Bertoli:2021:MIE

[BPP21] Marco Bertoli, Gabriele Piazza, and Elisabetta Pizzul. Macrobenthic invertebrate energy densities and ecological status in freshwater watercourses (Friuli Venezia-Giulia, Northeast Italy). Aquatic Ecology, 55(2):501–518, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09840-x.

Bakker:1985:IPS

[BPT85] C. Bakker, T. C. Prins, and M. L. M. Tackx. Interpretation of particle spectra of electronic counters by microscopical methods. Aquatic Ecology, 19(1):49–59, November 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255093.

Beltman:1981:SMD

[BR81] B. Beltman and W. Rietveld. Sampling macrofauna in ditches. Aquatic Ecology, 15(3):153–159, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255173.

Bruning:1987:IPL

[BR87] K. Bruning and J. Ringelberg. The influence of phosphorus limitation of the diatom Asterionella formosa on the zoospore production of its fungal parasite Rhizophydium planktonicum. Aquatic Ecology, 21(1):49–54, July 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255454.

Baumgartner:2005:IPB

[BR05] Daniel Baumgärtner and Karl-Otto Rothhaupt. The impact of predation by burbot (Lota lota L.) on the macroinvertebrate community in the littoral zone of a large lake. Aquatic Ecology, 39(1):79–92, March 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
004-1907-y.

Berkenbusch:2007:EST

[BR07] Katrin Berkenbusch and Ashley A. Rowden. An examination of the spatial and temporal generality of the influence of ecosystem engineers on the composition of associated assemblages. Aquatic Ecology, 41(1):129–147, March 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9053-3.

Brancelj:1999:EAA

[Bra99] Anton Brancelj. The extinction of Arctodiaptomus alpinus (Copepoda) following the introduction of charr into a small alpine lake Dvojno Jezero (NW Slovenia). Aquatic Ecology, 33(4):355–361, December 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009972108485.

Braukmann:2001:SAS

[Bra01] Ulrich Braukmann. Stream acidification in South Germany — chemical and biological assessment methods and trends*. Aquatic Ecology, 35(2):207–232, June 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011452014258.

Brancelj:2021:SZC

[Bra21] Anton Brancelj. Shifts in zooplankton communities in highmountain lakes induced by singular events (fish stocking, earthquakes): evidence from a 20-year survey in Slovenia (Central Europe). Aquatic Ecology, 55(4):1253–1271, December 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09858-1.

Bethke:2023:WYP

[BRA^{+}23] Bethany J. Bethke, Heidi M. Rantala, Tyler D. Ahrenstorff, Holly A. Wellard Kelly, Katya E. Kovalenko, Ryan P. Maki, Jodene K. Hirsch, Joshua D. Dumke, Valerie J. Brady, Jaime F. LeDuc, and Gretchen J. A. Hansen. Walleye and yellow perch resource use in large lakes invaded by spiny water fleas and zebra mussels. Aquatic Ecology, 57(3):571–584, September 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10030-0.

Bochert:1997:MCV

[BRB97] Anke Bochert, Doreen Richard, and Ralf Bochert. Marenzelleria cf. viridis and the sulphide regime. Aquatic Ecology, 31(2):223–231, June 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009919901870.

Breteler:1980:BPC

[Bre80] W. C. M. Klein Breteler. Breeding pelagic copepods. Aquatic Ecology, 14(3):225, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260132.

Breteler:1985:FAP

[Bre85] W. C. M. Klein Breteler. Fixation artifacts of phytoplankton in zooplankton grazing experiments. Aquatic Ecology, 19 (1):13–19, November 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255088.

Bagheri:2002:IAO

[BRG02] Sima Bagheri, Machteld Rijkeboer, and Herman J. Gons. Inherent and apparent optical measurements in the Hudson/ Raritan Estuary. Aquatic Ecology, 36(4):559–562, December 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1021114112617.

Bridges:1999:PPI

[Bri99] Christine M. Bridges. Predator-prey interactions between two amphibian species: effects of insecticide exposure. Aquatic Ecology, 33(2):205–211, September 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009922825444.

Bastrop:1997:WDM

[BRJ97] Ralf Bastrop, Matthias Röhner, and Karl Jürss. Where did Marenzelleria spp. (Polychaeta: Spionidae) in Europe come from? Aquatic Ecology, 31(2):119–136, June 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009994102526.

Brock:1981:NPB

[Bro81] Th. C. M. Brock. Nitrogen and phosphorus budgets of a nymphaeid-dominated system. Aquatic Ecology, 15(3): 191–192, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255179.

Brouwer:1984:RBS

[Bro84] Chris Brouwer. Relation between salinity and distribution of nematode species from the Ems estuary. Aquatic Ecology, 18(1):67–68, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256676.

Brockmann:1994:OME

[Bro94] U. H. Brockmann. Organic matter in the Elbe estuary. Aquatic Ecology, 28(3-4):371–381, September 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334207.

Brock:1998:ACS

[Bro98] Theo C. M. Brock. Assessing chemical stress in aquatic ecosystems: remarks on the need of an integrative approach. Aquatic Ecology, 32(2):107–111, September 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009920612418.

Braun:2015:UWC

[BRP15] Douglas C. Braun, John D. Reynolds, and David A. Patterson. Using watershed characteristics to inform cost-effective stream temperature monitoring. Aquatic Ecology, 49(3): 373–388, September 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9531-6.

Brotas:1995:MMV

[BS95] Vanda Brotas and João Serôdio. A mathematical model for the vertical distribution of chlorophyll-A in estuarine intertidal sediments. Aquatic Ecology, 29(3-4):315–321, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02084230.

Baptista:2014:IIC

[BSA14] I. Baptista, A. L. Santos, and A. Almeida. Influence of incubation conditions on bacterial production estimates in an estuarine system. Aquatic Ecology, 48(3):327–336, September 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-014-9487-y.

Bormans:2020:CCE

[BSA20a] Myriam Bormans, Véronique Savar, and Zouher Amzil. Cyanobacteria and cyanotoxins in estuarine water and sediment. Aquatic Ecology, 54(2):625–640, June 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09764-y.

Brasil:2020:RLH

[BSA20b] Jandeson Brasil, Juliana B. O. Santos, and José Luiz Attayde. Rainfall leads to habitat homogenization and facilitates plankton dispersal in tropical semiarid lakes. Aquatic Ecology, 54(1):225–241, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
019-09738-9.

Benchamin:2024:STV

[BSA24] Dani Benchamin, R. Sreejai, and M. S. Arya. Spatial and temporal variation in diversity and abundance of malacostraca associated with environmental variables in a tropical river in South India. Aquatic Ecology, 58(4):1013–1031, December 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10121-6.

Burckhardt:1997:FBP

[BSB97] Roger Burckhardt, Rhena Schumann, and Ralf Bochert. Feeding biology of the pelagic larvae of Marenzelleria cf. wireni (Polychaeta: Spionidae) from the Baltic Sea. Aquatic Ecology, 31(2):149–162, June 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009947320273.

Boavida:2011:AIS

[BSF11] I. Boavida, J. M. Santos, and M. T. Ferreira. Assessment of instream structures for habitat improvement for two critically endangered fish species. Aquatic Ecology, 45 (1):113–124, March 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-010-9340-x.

BarrosoMagalhaes:2024:SEM

[BSMVP24] André Lincoln Barroso Magalhães, Cristiano Schetini de Azevedo, Alberto Maceda-Veiga, and Jiří Patoka. The scientist eyes: monitoring YouTube to quantify aquatic pet release in Brazil. Aquatic Ecology, 58(1):73–88, March 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10059-1.

Bazzanti:1989:NCD

[BST89] Marcello Bazzanti, Marco Seminara, and Claudio Tamorri. A note on chironomids (Diptera) of temporary pools in the National Park of Circeo, Central Italy. Aquatic Ecology, 23 (2):189–193, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256736.

Borquez:2024:EUA

[BSW^{+}24] Jéssica Bórquez, Sandra Sampertegui, Britt N. Wallberg, Diana Coral-Santacruz, Víctor H. Ruiz, Paul B. Samollow, Nicolas Gouin, and Angéline Bertin. Ecological uniqueness across multiple levels of biodiversity in a Chilean watershed. Aquatic Ecology, 58(2):139–158, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10051-9.

Blaxter:1992:EPS

[BT92] J. H. S. Blaxter and C. C. Ten Hallers-Tjabbes. The effect of pollutants on sensory systems and behaviour of aquatic animals. Aquatic Ecology, 26(1):43–58, November 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02298027.

Bohman:2001:ESI

[BT01] Irene M. Bohman and Lars J. Tranvik. The effects of shredding invertebrates on the transfer of organic carbon from littoral leaf litter to water-column bacteria. Aquatic Ecology, 35(1):43–50, March 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1011425905036.

Bovee:2023:IBI

[BT23] Emily N. Bovee and Scott D. Tiegs. Interactions between invasive New Zealand mudsnails and leaf litter: growth and decomposition. Aquatic Ecology, 57(2):523–528, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10026-w.

Bale:1994:LMS

[BTA94] A. J. Bale, M. D. Tocher, and J. Aiken. Laboratory measurements of the spectral properties of estuarine suspended particles. Aquatic Ecology, 28(3-4):237–244, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334190.

Blanchfield:2009:SHS

[BTB09] Paul J. Blanchfield, Lori S. Tate, and Ken G. Beaty. Seasonal habitat selection by lake trout (Salvelinus namaycush) in a small Canadian Shield lake: constraints imposed by winter conditions. Aquatic Ecology, 43(3):777–787, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9266-3.

Benny:2024:ACE

[BTP24] Niya Benny, Lathika Cicily Thomas, and K. B. Padmakumar. Analysis of climatic and edaphic variabilities on the microphytobenthic mat characteristics of a riverine mangrove ecosystem along the southwest coast of India. Aquatic Ecology, 58(2):467–485, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10084-0.

Bakker:1988:PCT

[BTV88] C. Bakker, M. L. M. Tackx, and P. Van Rijswijk. Plankton copepods Temora longicornis and Acartia tonsa, and their food in the Oosterschelde estuary (S.W.-Netherlands).

Aquatic Ecology, 22(1):75–78, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256786.

Buiteveld:1995:MCD

[Bui95] H. Buiteveld. A model for calculation of diffuse light attenuation (PAR) and Secchi depth. Aquatic Ecology, 29(1):55– 65, April 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02061789.

Butter:1981:EZP

[But81] Maureen E. Butter. Estimation of zooplankton populations in Lake Maarsseveen i: Problems, procedures and results. Aquatic Ecology, 15(1-2):51–59, October 1981. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260258.

Bromley:1971:Z

[BV71] H. J. Bromley and J. Vijverberg. Zooplankton. Aquatic Ecology, 5(2):80–84, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185944.

Bolier:1976:LSS

[BV76] G. Bolier and A. N. Van Breemen. A limnological study of some man-made lakes in the province of South Holland. Aquatic Ecology, 10(1):19–29, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02308495.

Bakker:1978:GTP

[BV78] C. Bakker and F. Vegter. General tendencies of phyto- and zooplankton development in two closed estuaries (Lake Veere and Lake Grevelingen) in relation to an open estuary (Eastern Scheldt), S.W. Netherlands. Aquatic Ecology, 12(3-4): 226–245, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259185.

Bloem:1984:SOD

[BV84] J. Bloem and J. Vijverberg. Some observations on the diet and food selection of Daphnia hyalina (Cladocera) in an eutrophic lake. Aquatic Ecology, 18(1):39–45, June 1984. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256672.

Bakker:1989:MCC

[BV89] C. Bakker and P. Van Rijswijk. A marine calanoid copepod (Temora longicornis) in a semi-stagnant coastal lagoon (Lake Grevelingen, S.W.-Netherlands). Aquatic Ecology, 23 (2):179–188, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256735.

Brock:1996:AMC

[BV96] Theo C. M. Brock and Gerard Van Der Velde. Aquatic macroinvertebrate community structure of a Nymphoides peltata-dominated and macrophyte-free site in an oxbow lake. Aquatic Ecology, 30(2-3):151–163, October 1996. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02272235.

Bisseling-Visser:2004:SI

[BV04] Annette Bisseling-Visser. Society information. Aquatic Ecology, 38(3):473–474, September 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000035216.45556.7e.

Barreiro:2014:IBA

[BV14] Aldo Barreiro and Vitor Manuel Vasconcelos. Interactions between allelopathic properties and growth kynetics in four freshwater phytoplankton species studied by model simulations. Aquatic Ecology, 48(2):191–205, June 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9475-2.

Buth:1982:DTH

[BVD82] G. J. C. Buth, P. F. M. Verdonschot, and L. De Wolf. Decomposition of three halophytes in different habitats of an eastern Scheldt salt marsh. Aquatic Ecology, 16(1):103–112, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255418.

Bongers:1990:PBE

[BvdH90] Tom Bongers and Jan van de Haar. On the potential of basing an ecological typology of aquatic sediments on the nematode fauna: an example from the River Rhine. Aquatic Ecology, 24(1):37–45, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256747.

Bakker:2016:LRL

[BVI16] Elisabeth S. Bakker, Ellen Van Donk, and Anne K. Immers. Lake restoration by in-lake iron addition: a synopsis of iron impact on aquatic organisms and shallow lake ecosystems. Aquatic Ecology, 50(1):121–135, March 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9552-1.

Belle:2015:CSP

[BVM15] Simon Belle, Valérie Verneaux, and Michel Magny. A case study of the past CH_{4} cycle in lakes by the combined use of dual isotopes (carbon and hydrogen) and ancient DNA of methane-oxidizing bacteria: rearing experiment and application to Lake Remoray (eastern France). Aquatic Ecology, 49(3):279–291, September 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9523-6.

Bannink:1980:HCA

[BVV80] B. A. Bannink, J. H. M. Van Der Meulen, and J. C. Van Der Vlugt. Hydrobiological consequences of the addition of phosphate precipitants to inlet water of lakes. Aquatic Ecology, 14(1-2):73–89, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260275.

Butler:1992:FRE

[BW92] Malcolm G. Butler and Lorn Walker. Fecundity, reproductive effort, and pupal size in the profundal midge Chironomus cuccini (Diptera: Chironomidae). Aquatic Ecology, 26 (2-4):263–267, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255250.

Bukaveckas:2002:FRA

[BWH02] Paul A. Bukaveckas, John Jason Williams, and Susan P. Hendricks. Factors regulating autotrophy and heterotrophy in the main channel and an embayment of a large river impoundment. Aquatic Ecology, 36(3):355–369, August 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1016515518032.

Burger-Wiersma:1986:CDG

[BWM86] Tineke Burger-Wiersma and Luuc R. Mur. Carbohydrate dynamics and growth rate of natural phytoplankton populations. Aquatic Ecology, 20(1-2):87–92, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291153.

Beecraft:2019:IRP

[BWS19] Laura Beecraft, Susan B. Watson, and Ralph E. H. Smith. Innate resistance of PSII efficiency to sunlight stress is not an advantage for cyanobacteria compared to eukaryotic phytoplankton. Aquatic Ecology, 53(3):347–364, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09694-4.

Beecraft:2021:QUM

[BWS21] Laura Beecraft, Susan B. Watson, and Ralph E. H. Smith. Quantifying the uncertainties in multi-wavelength PAM fluorometry due to innate and irradiance-induced variability of fluorescence spectra. Aquatic Ecology, 55(1):169–186, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09821-6.

Bick:1997:IDT

[BZ97] Andreas Bick and Michael L. Zettler. On the identity and distribution of two species of Marenzelleria (Polychaeta, Spionidae) in Europe and North America. Aquatic Ecology, 31 (2):137–148, June 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009995119364.

Best:1981:GPP

[BZD81] Elly P. H. Best, M. Zippin, and J. H. A. Dassen. Growth and production of Phragmites australis in Lake Vechten (the Netherlands). Aquatic Ecology, 15(3):165–173, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255175.

Best:1982:SDP

[BZD82] Elly P. H. Best, M. Zippin, and J. H. A. Dassen. Studies on decomposition of Phragmites australis leaves under laboratory conditions. Aquatic Ecology, 16(1):21–33, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255410.

Botte:2022:CAA

[BZG^{+}22] Audrey Botté, Mariem Zaidi, Joëlle Guery, Denis Fichet, and Vincent Leignel. Correction to: Aluminium in aquatic environments: abundance and ecotoxicological impacts. Aquatic Ecology, 56(4):1359, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09965-7. See [BZL22].

Botte:2022:AAE

[BZL22] Audrey Botté, Mariem Zaidi, and Vincent Leignel. Aluminium in aquatic environments: abundance and ecotoxicological impacts. Aquatic Ecology, 56(3):751–773, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09936-4. See correction [BZG^{+}22].

Boderie:1993:NBW

[BZV93] Pascal M. A. Boderie, John J. G. Zwolsman, and Cornelis H. Van der Weijden. Nutrient biogeochemistry in the water column (N, P, Si) and porewater (N) of sandy sediment of the Scheldt estuary (SW-Netherlands). Aquatic Ecology, 27 (2-4):309–318, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334794.

Cadee:1978:OOM

[Cad78] G. C. Cadée. On the origin of organic matter accumulating on tidal flats of Balgzand, Dutch Wadden Sea. Aquatic Ecology, 12(2):145–150, July 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260717.

Cadee:1980:RDP

[Cad80] C. C. Cadee. A review of data on production and import of organic carbon in the western Wadden Sea. Aquatic Ecology, 14(3):225–226, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260133.

Cadee:1999:BRE

[Cad99] Gerhard C. Cadée. Book review: Eutrophication in Planktonic Ecosystems: Food Web Dynamics and Elemental Cycling. Edited by T. Tamminen and H. Kuosa. Aquatic Ecology, 33(2):214–215, September 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009902813195.

Cadee:2001:BRC

[Cad01] Gerhard C. Cadée. Book review: Coastal marine ecosystems of Latin America. Ecological Studies 144, U. Seeliger & B. Kjerfve (Eds). Aquatic Ecology, 35(1):87, March 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011472103764.

CadEe:2004:BRM

[Cad04] Gerhard C. CadÉe. Book review: Migrations and Dispersal of Marine Organisms. Aquatic Ecology, 38(3):471–472, September 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000035214.09264.21.

Cadee:2005:BRa

[Cad05a] Gerhard C. Cadée. Book review. Aquatic Ecology, 38(4): 611–612, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-1174-6.

Cadee:2005:BRc

[Cad05b] Gerhard C. Cadée. Book review. Aquatic Ecology, 38(4): 616–617, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-3428-8.

Cadee:2005:BRB

[Cad05c] Gerhard C. Cadee. Book review: Biology of the Baltic. Aquatic Ecology, 38(4):611–612, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-1174-y.

Cadee:2005:BRN

[Cad05d] Gerhard C. Cadee. Book review: Nutrients and Eutrophication in Estuaries and Coastal Waters. Aquatic Ecology, 38 (4):616–617, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-004-3428-0.

Cadee:2005:BRR

[Cad05e] Gerhard C. Cadée. Book review: Red Tides. Edited by Tomotoshi Okaichi, 2004, xiii + 439 pp. (Hardbound). Kluwer Academic Publishers, Dordrecht. ISBN 1-4020-1512-7. Price: EUR 90.00, US$ 99.00, GB £62.00. Aquatic Ecology, 39(2): 259, June 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-5202-3.

Cadee:2008:BRM

[Cad08] Gerhard C. Cadée. Book review: Mark D. Bertness, Atlantic Shorelines, Natural History and Ecology. Aquatic Ecology, 42(1):177–178, March 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9094-2.

Cuvin-Aralar:2004:ELN

[CAFA04] Maria Lourdes Cuvin-Aralar, Ulfert Focken, and Emiliano V. Aralar. Effects of low nitrogen-phosphorus ratios in the phytoplankton community in Laguna de Bay, a shallow eutrophic lake in the Philippines. Aquatic Ecology, 38(3):387–401, September 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000035174.35091.0b.

Clements:2011:RBM

[CAJ11] William H. Clements, Jeffrey L. Arnold, and Cathie Jean. Responses of benthic macroinvertebrate communities to natural geothermal discharges in Yellowstone National Park, USA. Aquatic Ecology, 45(1):137–149, March 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9342-8.

Cappenberg:1975:SAR

[Cap75] Th. E. Cappenberg. Some aspects on the role of bacteria in deep water ecosystems. Aquatic Ecology, 9(1):45–47, April 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257518.

Campbell:2008:TDA

[CAPGA08] W. Bruce Campbell, Roberto Arce-Pérez, and José Antonio Gómez-Anaya. Taxonomic distinctness and aquatic Coleoptera: comparing a perennial and intermittent stream with differing geomorphologies in Hidalgo, Mexico. Aquatic Ecology, 42(1):103–113, March 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
007-9086-2.

Cazemier:1982:GBA

[Caz82] W. G. Cazemier. The growth of bream (Abramis brama L.) in relation to habitat and population density. Aquatic Ecology, 16(2-3):269–277, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255380.

Coesel:1994:DSD

[CB94] Peter Coesel and Hanny Kooijman-Van Blokland. Distribution and seasonality of desmids in the Maarsseveen Lakes area. Aquatic Ecology, 28(1):19–24, March 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334241.

Caramujo:2000:DDH

[CB00a] Maria-José Caramujo and Maria-José Boavida. Dynamics of Daphnia hyalina × galeata in Castelo-do-Bode Reservoir: The effect of food availability and flatworm predation. Aquatic Ecology, 34(2):155–163, June 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009956621502.

Caramujo:2000:SDC

[CB00b] Maria-José Caramujo and Maria-José Boavida. The seasonal dynamics of Copidodiaptomus numidicus (Gurney, 1909) and Thermocyclops dybowskii (Lande, 1890) in Castelo-do-Bode Reservoir. Aquatic Ecology, 34(2):143–153, June 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009982422411.

Cataldo:2000:YRA

[CB00c] Daniel H. Cataldo and Demetrio Boltovskoy. Yearly reproductive activity of Limnoperna fortunei (Bivalvia) as inferred from the occurrence of its larvae in the plankton of the lower Paraná river and the Río de la Plata estuary (Argentina). Aquatic Ecology, 34(3):307–317, September 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009983920942.

Cohen:2013:NAE

[CB13] Jillian S. Cohen and Bernd Blossey. No apparent effects of soil inoculum on green frog (Lithobates clamitans Latreille) tadpole performance. Aquatic Ecology, 47(4):425–431, December 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9455-y.

Constable:2016:IIA

[CB16] Drew Constable and Nina J. Birkby. The impact of the invasive amphipod Dikerogammarus haemobaphes on leaf litter processing in UK rivers. Aquatic Ecology, 50(2):273–281, June 2016. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9574-3.

Cabecadas:1999:PSB

[CBC99] L. Cabeçadas, M. J. Brogueira, and G. Cabeçadas. Phytoplankton spring bloom in the Tagus coastal waters: hydrological and chemical conditions. Aquatic Ecology, 33(3):243–250, October 1999. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009941313169.

Ceriello:2024:BSM

[CBdO^{+}24] Hellen Ceriello, Gustavo R. Brito, Bruno Francesco Rodrigues de Oliveira, Marielton dos Passos Cunha, and Sérgio N. Stampar. Bullseye: shotgun metagenomics taking aim at the microbial diversity associated with tubes of Ceriantharia. Aquatic Ecology, 58(4):1287–1300, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10144-z.

Ceschin:2020:ESA

[CBE20] Simona Ceschin, Amii Bellini, and Neil Thomas William Ellwood. Ecological study of the aquatic carnivorous plant Utricularia australis R.Br. (Lentibulariaceae). Aquatic Ecology, 54(1):295–307, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
019-09743-y.

Cavraro:2019:BCF

[CBF19] Francesco Cavraro, Nicola Bettoso, and Piero Franzoi. Body condition in fish as a tool to detect the effects of anthropogenic pressures in transitional waters. Aquatic Ecology, 53(1):21–35, March 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-018-09670-4.

Catsiki:1994:SDM

[CBN94] V. A. Catsiki, F. Bel, and A. Nicolaidou. Size dependent metal concentrations in two marine gastropod species. Aquatic Ecology, 28(2):157–165, June 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02333986.

Crane:2019:SBM

[CBNF19] Adam L. Crane, Kevin R. Bairos-Novak, and Maud C. O. Ferrari. Survival, behaviour, and morphology of larval wood frogs, Lithobates sylvaticus, under threat from an exotic crayfish predator, Orconectes virilis. Aquatic Ecology, 53(3): 383–392, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09696-2.

Chiapusio:2022:SPR

[CBP22] Geneviève Chiapusio, Philippe Binet, and Pierrick Priault. Sphagnum physiological responses to elevated temperature, nitrogen, CO_{2} and low moisture in laboratory and in situ microhabitats: a review. Aquatic Ecology, 56(2):429–445, June 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09924-8.

Colomer:1994:CSL

[CC94] J. Colomer and X. Casamitjana. Calculating solar and longwave radiation for a Mediterranean lake under various weather conditions. Aquatic Ecology, 28(2):143–148, June 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02333984.

Costa:1999:CTN

[CC99] M. J. Costa and H. N. Cabral. Changes in the Tagus nursery function for commercial fish species: some perspectives for management. Aquatic Ecology, 33(3):287–292, October 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009904621771.

Cicala:2019:SVF

[CCC19] Davide Cicala, Edoardo Calizza, and Maria Letizia Costantini. Spatial variation in the feeding strategies of Mediterranean fish: flatfish and mullet in the Gulf of Gaeta (Italy). Aquatic Ecology, 53(4):529–541, December 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09706-3.

Carey:2011:PDD

[CCH11] Cayelan C. Carey, Moana P. Ching, and Nelson G. Hairston, Jr. Predator-dependent diel migration by Halocaridina rubra shrimp (Malacostraca: Atyidae) in Hawaiian anchialine pools. Aquatic Ecology, 45(1):35–41, March 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9321-0.

Chen:2010:SVM

[CCK10] Meijun Chen, Feizhou Chen, and Fanxiang Kong. Seasonal variation of microbial eukaryotic community composition in the large, shallow, subtropical Taihu Lake, China. Aquatic Ecology, 44(1):1–12, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9254-7.

Caetano:2021:SSS

[CCM21] Vinne Caetano, Mateus Camana, and Adriano Sanches Melo. Scale-sensitive stream slope drives nested fish trait-based diversity. Aquatic Ecology, 55(3):1051–1063, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09881-2.

Cabrita:1999:ILT

[CCS99] Maria Teresa Cabrita, Fernando Catarino, and Gerd Slawyk. Interactions of light, temperature and inorganic nitrogen in controlling planktonic nitrogen utilisation in the Tagus estuary. Aquatic Ecology, 33(3):251–261, October 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009993300873.

Casu:2009:CRV

[CCS09] Daniela Casu, Giulia Ceccherelli, and Gianluca Sarà. Caulerpa racemosa var. cylindracea as a potential source of organic matter for benthic consumers: evidences from a stable isotope analysis. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9223-6.

Costa:2020:LCM

[CCS20] Ana Paula Tavares Costa, Luciane Oliveira Crossetti, and Fabiana Schneck. Land cover is the main correlate of phytoplankton beta diversity in subtropical coastal shallow lakes. Aquatic Ecology, 54(4):1015–1028, December 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09790-w.

Cunha:2003:FBB

[CDA03] M. A. Cunha, J. M. Dias, and F. Alcântara. Fluxes of bacterioplankton between a tidal estuary and the sea: returning to the “Outwelling Hypothesis”. Aquatic Ecology, 37(1):45–54, January 2003. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1022173730178.

Cuthbert:2020:AMP

[CDD20] Ross N. Cuthbert, Tatenda Dalu, and Jaimie T. A. Dick. Assessing multiple predator, diurnal and search area effects on predatory impacts by ephemeral wetland specialist copepods. Aquatic Ecology, 54(1):181–191, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09735-y.

Chen:2013:MDL

[CDL13] Feizhou Chen, Xi Dai, and Zhengwen Liu. Microcystins derived from lysing Microcystis cells do not cause negative effects on crustacean zooplankton in Lake Taihu, China. Aquatic Ecology, 47(4):379–387, December 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9451-2.

Chen:2014:ESB

[CDL14] Xinsheng Chen, Zhengmiao Deng, and Y. F. Li. Effects of sediment burial disturbance on the vegetative propagation of Phalaris arundinacea with different shoot statuses. Aquatic Ecology, 48(4):409–416, December 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9493-0.

Coffman:1992:SRS

[CDW92] W. P. Coffman, C. De La Rosa, and M. A. Wilzbach. Species richness in some Neotropical (Costa Rica) and afrotropical (West Africa) lotic communities of Chironomidae (Diptera). Aquatic Ecology, 26(2-4):229–237, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255246.

Carey:2012:OTC

[CEH12] Cayelan C. Carey, Holly A. Ewing, and James F. Haney. Occurrence and toxicity of the cyanobacterium Gloeotrichia echinulata in low-nutrient lakes in the northeastern United States. Aquatic Ecology, 46(4):395–409, December 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9409-9.

Chandrasekara:1996:ERF

[CF96] W. U. Chandrasekara and C. L. J. Frid. The effects of Relic fauna on initial patch colonisation in a British saltmarsh. Aquatic Ecology, 30(1):49–60, July 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02092147.

Castillo:2009:EAF

[CF09] Jesús M. Castillo and Enrique Figueroa. Effects of abiotic factors on the life span of the invasive cordgrass Spartina densiflora and the native Spartina maritima at low salt marshes. Aquatic Ecology, 43(1):51–60, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9159-2.

Cebrian:2008:IHI

[CFP08] Just Cebrian, C. Drew Foster, and Sean P. Powers. The impact of Hurricane Ivan on the primary productivity and metabolism of marsh tidal creeks in the NorthCentral Gulf of Mexico. Aquatic Ecology, 42(3):391–404, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9096-0.

Corona-Figueroa:2021:SMD

[CFRNT21] M. F. Corona-Figueroa, N. Ríos, and C. A. Niño-Torres. Searching for manatees in the dark waters of a transboundary river between Mexico and Belize: a predictive distribution model. Aquatic Ecology, 55(1):59–74, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09810-9.

Cabral:2019:PZC

[CGC19] Camila R. Cabral, Rafael D. Guariento, and Adriano Caliman. Are the patterns of zooplankton community structure different between lakes and reservoirs? An local and regional assessment across tropical ecosystems. Aquatic Ecology, 53(3):335–346, September 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09693-5.

Capezzuto:2019:HFM

[CGD19] Francesca Capezzuto, Diana Maria Paola Galassi, and Gianfranco D’Onghia. How far may life venture? Observations on the harpacticoid copepod Phyllognathopus viguieri under extreme stress conditions. Aquatic Ecology, 53(4):629–637, December 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09713-4.

Chang:2021:SEF

[CGL21] Chao Chang, Li Gao, and Ming Li. Spatial and environmental factors contributing to phytoplankton biogeography and biodiversity in mountain ponds across a large geographic area. Aquatic Ecology, 55(2):721–735, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09857-2.

Can:2019:PCC

[CGO19] Hatice Kaplan Can, Fatma Gurbuz, and Mehmet Odabaşı. Partial characterization of cyanobacterial extracellular polymeric substances for aquatic ecosystems. Aquatic Ecology, 53(3):431–440, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
019-09699-z.

Colijn:1983:MPP

[CGZ83] F. Colijn, W. W. C. Gieskes, and W. Zevenboom. The measurement of primary production: Problems and recommendations. Aquatic Ecology, 17(1):29–51, June 1983. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255190.

Cramer:1983:AEN

[CH83] Jacqueline Cramer and Rob Hagendijk. Aquatic ecology in The Netherlands: What is being done by whom? Aquatic Ecology, 17(1):77–85, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255195.

Cadee:1991:HPD

[CH91a] Gerhard C. Cadée and Jan Hegeman. Historical phytoplankton data of the Marsdiep. Aquatic Ecology, 24(2):111–118, April 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260427.

Cadee:1991:PPP

[CH91b] Gerhard C. Cadée and Jan Hegeman. Phytoplankton primary production, chlorophyll and species composition, organic carbon and turbidity in the Marsdiep in 1990, compared with foregoing years. Aquatic Ecology, 25(1):29–35, October 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02259586.

Coll:2012:TFC

[CH12] Marta Coll and Kevin Hargadon. Trophic and functional cascades in tropical versus temperate aquatic microcosms. Aquatic Ecology, 46(1):55–71, March 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9381-9.

Chambers:1971:BF

[Cha71] M. Chambers. Bottom fauna. Aquatic Ecology, 5(2):85– 91, June 1971. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185945.

Chale:1987:PBN

[Cha87] Francis M. M. Chale. Plant biomass and nutrient levels of a tropical macrophyte (Cyperus papyrus L.) receiving domestic wastewater. Aquatic Ecology, 21(2):167–170, December 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255442.

Chale:1989:NRS

[Cha89] Francis M. M. Chale. Nutrient release from swamp water, Cyperus papyrus L. organs and swamp sediments. Aquatic Ecology, 23(2):195–199, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256737.

Cortes:2009:HVD

[CHF09] Rui Manuel Vitor Cortes, Samantha Jane Hughes, and Maria Teresa Ferreira. Habitat variation at different scales and biotic linkages in lotic systems: consequences for monitorization. Aquatic Ecology, 43(4):1107–1120, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9228-9.

Chizinski:2006:MHT

[CHP06] C. J. Chizinski, C. L. Higgins, and K. L. Pope. Multiple hypotheses testing of fish incidence patterns in an urbanized ecosystem. Aquatic Ecology, 40(1):97–109, March 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9025-z.

Cisneros:1991:NDH

[CHV91] R. Cisneros, E. Hooker, and L. E. Velasquez. Natural diet of herbivorous zooplankton in Lake Xolotlán (Managua). Aquatic Ecology, 25(2):163–167, June 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291249.

Connolly:2023:SAF

[CJS^{+}23] R. M. Connolly, K. I. Jinks, A. Shand, M. D. Taylor, T. F. Gaston, A. Becker, and E. L. Jinks. Out of the shadows: automatic fish detection from acoustic cameras. Aquatic Ecology, 57(4):833–844, December 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09967-5.

Cornwell:1999:DCE

[CKK99] Jeffrey C. Cornwell, W. Michael Kemp, and Todd M. Kana. Denitrification in coastal ecosystems: methods, environmental controls, and ecosystem level controls, a review. Aquatic Ecology, 33(1):41–54, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009921414151.

Casotti:2019:LCB

[CKM19] Cinthia G. Casotti, Walace P. Kiffer, Jr., and Marcelo S. Moretti. The longer the conditioning, the better the quality? The effects of leaf conditioning time on aquatic hyphomycetes and performance of shredders in a tropical stream. Aquatic Ecology, 53(2):163–178, June 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09680-w.

Chen:2011:TSM

[CL11] Bingzhang Chen and Hongbin Liu. Temporal stability of marine phytoplankton in a subtropical coastal environment. Aquatic Ecology, 45(3):427–438, September 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9365-9.

Claassen:1982:LDI

[Cla82] T. H. L. Claassen. Limnological data of an isolated Dutch broad. Aquatic Ecology, 16(2-3):165–179, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255370.

Claassen:1984:HCL

[Cla84] T. H. L. Claassen. Historical and current limnological research in the small River Linde, The Netherlands. Aquatic Ecology, 18(1):23–34, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256670.

Claassen:1987:MFL

[Cla87] T. H. L. Claassen. The macroinvertebrate fauna of littoral and bottom substrata as a tool for biotypology of Frisian waters. Aquatic Ecology, 21(2):181–191, December 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255444.

Cloutier:2006:ESS

[CLH06] Danielle Cloutier, Magali N. LeCouturier, and Phil R. Hill. The effects of suspended sediment concentration on turbulence in an annular flume. Aquatic Ecology, 40(4):555–565, December 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-4810-2.

Chen:2010:GGR

[CLL10] Bingzhang Chen, Hongbin Liu, and Macy T. S. Lau. Grazing and growth responses of a marine oligotrichous ciliate fed with two nanoplankton: does food quality matter for micrograzers? Aquatic Ecology, 44(1):113–119, March 2010.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9264-5.

Cloern:1999:RIL

[Clo99] James E. Cloern. The relative importance of light and nutrient limitation of phytoplankton growth: a simple index of coastal ecosystem sensitivity to nutrient enrichment. Aquatic Ecology, 33(1):3–15, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009952125558.

Cararo:2023:DIC

[CLRdSR23] Emanuel Rampanelli Cararo, Cássia Alves Lima-Rezende, and Renan de Souza Rezende. Direct and indirect competition negatively affect the foraging activity of a neotropical shredder. Aquatic Ecology, 57(2):397–404, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10017-x.

Cabrita:1995:STV

[CM95a] M. T. Cabrita and M. T. Moita. Spatial and temporal variation of physico-chemical conditions and phytoplankton during a dry year in the Tagus estuary (Portugal). Aquatic Ecology, 29(3-4):323–332, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02084231.

Cunha:1995:MPM

[CM95b] M. R. Cunha and M. H. Moreira. Macrobenthos of Potamogeton and Myriophyllum beds in the upper reaches of Canal de Mira (Ria de Aveiro, NW Portugal): Community structure and environmental factors. Aquatic Ecology, 29(3-4):377–390, December 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02084237.

Costanza:1999:WHE

[CM99] Robert Costanza and Michael Mageau. What is a healthy ecosystem? Aquatic Ecology, 33(1):105–115, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009930313242.

Chan:2005:RBN

[CM05] K. Chan and Brian Morton. The reproductive biology of Nassarius festivus (Powys, 1835) (Gastropoda: Nassariidae) in relation to seasonal changes in temperature and salinity in subtropical Hong Kong. Aquatic Ecology, 39(2):213–228, June 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-6144-x.

Callieri:2007:PBP

[CMB07] Cristiana Callieri, Beatriz Modenutti, and Esteban Balseiro. Production and biomass of picophytoplankton and larger autotrophs in Andean ultraoligotrophic lakes: differences in light harvesting efficiency in deep layers. Aquatic Ecology, 41 (4):511–523, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9125-z.

Cunillera-Montcusi:2020:LSR

[CMBG20] David Cunillera-Montcusí, Dani Boix, and Stéphanie Gascón. Large- and small-regional-scale variables interact in the dispersal patterns of aquatic macroinvertebrates from temporary ponds. Aquatic Ecology, 54(4):1041–1058, December 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09792-8.

Chazanah:2025:DBM

[CMBI25] Nurul Chazanah, Barti Setiani Muntalif, Satria Wira Bagaskara, and Fikri Ilmi. Distribution of benthic macroinvertebrates as bioindicators of river ecological quality influenced by textile and pharmaceutical industrial pollution in the upper citarum river. Aquatic Ecology, 59(1):215–231, March 2025. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10158-7.

Caetano:1995:VMF

[CMG95] M. Caetano, M. J. Madureira, and M. L. Gonçalves. Variations of Mn, Fe and S concentrations in sediment pore waters of Ria Formosa at different time scales. Aquatic Ecology, 29 (3-4):275–281, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02084225.

Chaudhuri:2014:FPA

[CMH14] Atreyee Chaudhuri, Sudeshna Mukherjee, and Sumit Homechaudhuri. Food partitioning among carnivores within feeding guild structure of fishes inhabiting a mudflat ecosystem of Indian Sundarbans. Aquatic Ecology, 48(1):35–51, March 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-013-9464-x.

Clemente:2005:SCM

[CMM05] Juan María Clemente, Néstor Mazzeo, and Mariana Meerhoff. Succession and collapse of macrozoobenthos in a subtropical hypertrophic lake under restoration (Lake Rodó, Uruguay). Aquatic Ecology, 39(4):455–464, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9004-4.

Cisneros:1991:QQS

[CMV91] R. Cisneros, E. I. Mangas, and M. Van Maren. Qualitative and quantitative structure, diversity and fluctuations in abundance of zooplankton in Lake Xolotlán (Managua). Aquatic Ecology, 25(2):151–156, June 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291247.

Cancino:1994:NMC

[CN94] Leonor Cancino and Ramiro Neves. 3D-numerical modelling of cohesive suspended sediment in the Western Scheldt estuary (The Netherlands). Aquatic Ecology, 28(3-4):337–345, September 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334202.

Cerna:2010:PRM

[ČN10] Kateřina Černá and Jiří Neustupa. The pH-related morphological variations of two acidophilic species of Desmidiales (Viridiplantae) isolated from a lowland peat bog, Czech Republic. Aquatic Ecology, 44(2):409–419, June 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9296-x.

Castilho-Noll:2007:MEI

[CNA07] Maria Stela Maioli Castilho-Noll and Marlene Sofia Arcifa. Mesocosm experiment on the impact of invertebrate predation on zooplankton of a tropical lake. Aquatic Ecology, 41 (4):587–598, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9112-4.

Carneiro:2024:SMI

[CNN^{+}24] A. Carneiro, L. S. Nascimento, M. A. Noernberg, C. S. Hara, and A. T. R. Pozo. Social media image classification for jellyfish monitoring. Aquatic Ecology, 58(1):3–15, March 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10078-y.

Coarer:2007:HSE

[Coa07] Yann Le Coarer. Hydraulic signatures for ecological modelling at different scales. Aquatic Ecology, 41(3):451–459, September 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-9005-3.

Colloff:2018:UHE

[COC18] Matthew J. Colloff, Ian C. Overton, and Susan M. Cuddy. The use of historical environmental monitoring data to test predictions on cross-scale ecological responses to alterations in river flows. Aquatic Ecology, 52(1):133–153, March 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9650-y.

Coesel:1975:RDB

[Coe75] P. F. M. Coesel. The relevance of desmids in the biological typology and evaluation of fresh waters. Aquatic Ecology, 9 (3):93–101, December 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02263326.

Coesel:1977:EDS

[Coe77] P. F. M. Coesel. On the ecology of desmids and the suitability of these algae in monitoring the aquatic environment. Aquatic Ecology, 11(1):20–21, April 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02282211.

Coesel:1978:ECO

[Coe78] P. F. M. Coesel. Environmental changes in the Oisterwijk moorland pool area, evident from the composition of the desmid flora. Aquatic Ecology, 12(1):52–53, April 1978. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260793.

Coesel:1997:ESC

[Coe97] Peter F. M. Coesel. The edibility of Staurastrum chaetoceras and Cosmarium abbreviatum (Desmidiaceae) for Daphnia galeata/hyalina and the role of desmids in the aquatic food web. Aquatic Ecology, 31(1):73–78, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009987217642.

Corlett:1978:EIP

[Cor78a] John Corlett. Ecological implications of proposed water storage schemes in British estuaries. Aquatic Ecology, 12(3-4): 291–298, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259190.

Corlett:1978:IUK

[Cor78b] John Corlett. Introduction to United Kingdom estuarial engineering schemes. Aquatic Ecology, 12(3-4):273–276, December 1978. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02259188.

Choudhury:2010:PND

[CP10] Avik Kumar Choudhury and Ruma Pal. Phytoplankton and nutrient dynamics of shallow coastal stations at Bay of Bengal, Eastern Indian coast. Aquatic Ecology, 44(1):55–71, March 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9252-9.

Cibils:2015:FDA

[CPA15] Luciana Cibils, Romina Principe, and Ricardo Albariño. Functional diversity of algal communities from headwater grassland streams: How does it change following afforestation? Aquatic Ecology, 49(4):453–466, December 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9538-z.

Chakona:2009:CBA

[CPM09] Albert Chakona, Crispen Phiri, and Nikite Muller. Changes in biota along a dry-land river in northwestern Zimbabwe: declines and improvements in river health related to land use. Aquatic Ecology, 43(4):1095–1106, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9222-7.

Chamberlain:2022:DME

[CPP22] Debbie A. Chamberlain, Hugh P. Possingham, and Stuart R. Phinn. Decision-making with ecological process for coastal and marine planning: current literature and future directions. Aquatic Ecology, 56(1):1–19, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09896-9.

Cramer:1983:PEE

[Cra83] Jacqueline Cramer. Problems of ecologists as experts in environmental decision-making in The Netherlands. Aquatic Ecology, 17(2):157–167, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02280824.

Cramer:1985:BDF

[Cra85] Jacqueline Cramer. The behaviour of Dutch fresh water ecologists in response to environmental concern. Aquatic Ecology, 19(2):207–216, December 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02270768.

Cramer:1987:SMO

[Cra87] Jacqueline Cramer. The scope of mission-orientation in Dutch fresh-water ecology. Aquatic Ecology, 21(2):223–234, December 1987. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255448.

Craeymeersch:1991:AAB

[Cra91] J. A. Craeymeersch. Applicability of the abundance/biomass comparison method to detect pollution effects on intertidal macrobenthic communities. Aquatic Ecology, 24(2):133–140, April 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260430.

Cruz-Rivera:2022:USH

[CRR22] Edwin Cruz-Rivera and D. Christopher Rogers. An unavoidably short history of inland aquatic animal diversity research in the US Virgin Islands. Aquatic Ecology, 56(3):719–740, September 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09933-7.

Crumrine:2010:BST

[Cru10] Patrick W. Crumrine. Body size, temperature, and seasonal differences in size structure influence the occurrence of cannibalism in larvae of the migratory dragonfly, Anax junius. Aquatic Ecology, 44(4):761–770, December 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9314-z.

Cummings:2013:DVI

[CS13] Brittany M. Cummings and Daniel E. Schindler. Depth variation in isotopic composition of benthic resources and assessment of sculpin feeding patterns in an oligotrophic Alaskan lake. Aquatic Ecology, 47(4):403–414, December 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9453-0.

Chakraborty:2023:CSB

[CSA23] Anupam Chakraborty, Goutam K. Saha, and Gautam Aditya. A comparative study on the bioturbation ability of seven freshwater snail species. Aquatic Ecology, 57(1):35– 52, March 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-022-09991-5.

Ceschin:2010:TFV

[CSK10] Simona Ceschin, Giovanni Salerno, and Alma Kumbaric. Temporal floristic variations as indicator of environmental changes in the Tiber River in Rome. Aquatic Ecology, 44 (1):93–100, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9292-1.

Carvalho:2008:CRC

[CSM08] L. Carvalho, A. Solimini, and U. Mischke. Chlorophyll reference conditions for European lake types used for intercalibration of ecological status. Aquatic Ecology, 42(2):203– 211, June 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9189-4.

Corbisier:2006:UIS

[CSV06] T. N. Corbisier, L. S. H. Soares, and I. Valiela. Use of isotopic signatures to assess the food web in a tropical shallow marine ecosystem of Southeastern Brazil. Aquatic Ecology, 40(3): 381–390, September 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-006-9033-7.

Crawshaw:2019:HFC

[CSV19] Josie A. Crawshaw, Marc Schallenberg, and Robert Van Hale. Hierarchy of factors controls denitrification rates in temperate intermittently closed and open coastal lakes/ lagoons (ICOLLS). Aquatic Ecology, 53(4):719–744, December 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09721-4.

Capitulo:2001:UBM

[CTO01] Alberto Rodrigues Capítulo, Mariana Tangorra, and Carolina Ocón. Use of benthic macroinvertebrates to assess the biological status of Pampean streams in Argentina. Aquatic Ecology, 35(2):109–119, June 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1011456916792.

Cuppen:1994:LCH

[Cup94] J. G. M. Cuppen. Life cycle and habitat of Glossiphonia paludosa (Hirudinea: Glossiphoniidae), a new leech for the Netherlands. Aquatic Ecology, 28(2):193–197, June 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02333990.

Colijn:1977:BPP

[CV77] Franciscus Colijn and Loes Venekamp. Benthic primary production in the Ems–Dollard estuary during 1975. Aquatic Ecology, 11(1):16–17, April 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02282209.

Chow:1991:VDP

[CV91] Ninoska Chow and Luisa Amanda Vargas. Vertical distribution of phytoplankton in Lake Xolotlán (Managua). Aquatic Ecology, 25(2):133–136, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02291244.

Castro:1995:TPO

[CV95] Odete G. Castro and Carlos Vale. Total PCB-organic matter correlation in sediments from three estuarine areas of Portugal. Aquatic Ecology, 29(3-4):297–302, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084228.

Coops:1996:IHC

[CV96] Hugo Coops and Gerard Van Der Velde. Impact of hydrodynamic changes on the zonation of helophytes. Aquatic Ecology, 30(2-3):165–173, October 1996. CODEN AQECF9.

ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02272236.

Cerbin:2007:IMV

[CvDG07] Slawomir Cerbin, Ellen van Donk, and Ramesh D. Gulati. The influence of Myriophyllum verticillatum and artificial plants on some life history parameters of Daphnia magna. Aquatic Ecology, 41(2):263–271, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9091-5.

Collos:2003:RCP

[CVG03] Y. Collos, A. Vaquer, and N. Garcia. Response of coastal phytoplankton to ammonium and nitrate pulses: seasonal variations of nitrogen uptake and regeneration. Aquatic Ecology, 37(3):227–236, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1025881323812.

Courboules:2021:EEW

[CVM21] Justine Courboulès, Francesca Vidussi, and Behzad Mostajir. Effects of experimental warming on small phytoplankton, bacteria and viruses in autumn in the Mediterranean coastal Thau Lagoon. Aquatic Ecology, 55(2):647–666, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09852-7.

Coops:2002:PLC

[CvNB02] H. Coops, E. H. van Nes, and G. D. Butijn. Promoting lowcanopy macrophytes to compromise conservation and recreational navigation in a shallow lake. Aquatic Ecology, 36(4): 483–492, December 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1021177714293.

Cope:2004:CIB

[CW04] Neisha J. Cope and Michael J. Winterbourn. Competitive interactions between two successful molluscan invaders of freshwaters: an experimental study. Aquatic Ecology, 38(1):83–91, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000021018.20945.9d.

Cardoso:1995:ARB

[CY95] A. C. Cardoso and A. B. Yule. Aspects of the reproductive biology of Pollicipes pollicipes (Cirripedia; Lepadomorpha) from the southwest coast of Portugal. Aquatic Ecology, 29 (3-4):391–396, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02084238.

Chiu:2013:STE

[CYK13] Ming-Chih Chiu, Chao-Hsien Yeh, and Mei-Hwa Kuo. Shortterm effects of dam removal on macroinvertebrates in a Taiwan stream. Aquatic Ecology, 47(2):245–252, June 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9439-y.

Cao:2024:EED

[CYW24] Dandan Cao, Yongqing Yang, and Dong Wang. Effects of exotic detritus input on native litter breakdown in a eutrophic lake: investigating the home-field advantage. Aquatic Ecology, 58(2):501–513, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10086-6.

Cao:2014:PSR

[CZC14] Yanmin Cao, Enlou Zhang, and Guojun Cheng. A primary study on relationships between subfossil chironomids and the distribution of aquatic macrophytes in three lowland floodplain lakes, China. Aquatic Ecology, 48(4):481–492, December 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9499-7.

Chen:2025:SBC

[CZZ^{+}25] Jing Chen, Yongqiang Zhou, Yunlin Zhang, Quanzhong Guo, Shulan Zhang, and Guanghuan Ge. Succession of bacterial community during microbially driven cyanobacterial organic matter degradation and its relationship to water quality in Taihu Lake, China. Aquatic Ecology, 59(1):125–144, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10152-z.

DeLange:1999:SCP

[DA99] H. J. De Lange and M. T. Arts. Seston composition and the potential for Daphnia growth. Aquatic Ecology, 33(4): 387–398, December 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009920317186.

Dankers:1984:TWP

[Dan84] Norbert Dankers. Transports of water, particulate and dissolved organic and inorganic matter between a salt marsh and the Ems–Dollard estuary, the Netherlands. Aquatic Ecology, 18(1):77, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256686.

Prinzio:2024:RRS

[DAO24] Cecilia Y. Di Prinzio, Ivan Arismendi, and J. Andrés Olivos. Revealing a rapid shift in the phenology of the adult spawning migration of an introduced Chinook salmon population in Patagonia. Aquatic Ecology, 58(2):299–312, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10066-2.

Daro:1973:MND

[Dar73] M. H. Daro. Les migrations nycthemerales differentielles suivant l’age du meroplancton. (French) [Differential nycthemeral migrations according to the age of meroplankton]. Aquatic Ecology, 7(2):73–80, June 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02282201.

Dorgelo:1997:R

[DAR97] Jaap Dorgelo, Wim Admiraal, and Tineke Reede. Reviews. Aquatic Ecology, 30(4):339–348, May 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02085877.

DeBettencourt:1999:OTE

[DAR99] A. M. M. De Bettencourt, M. O. Andreae, and S. Rapsomanikis. Organotin in the Tagus estuary. Aquatic Ecology, 33 (3):271–280, October 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009966006325.

Das:2007:RHP

[Das07] Biplob Das. Reconstruction of historical productivity using visible-near-infrared (VNIR) reflectance properties from boreal and saline lake sediments. Aquatic Ecology, 41(2):209– 220, June 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-006-9071-1.

Davids:1971:MOL

[Dav71] C. Davids. Mikrofauna-onderzoek in de Lindevallei. (Dutch) [Mikrofauna research in the Linde Valley]. Aquatic Ecology, 5(3):115–125, September 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185889.

Davids:1981:P

[Dav81] C. Davids. Publications. Aquatic Ecology, 15(3):203, December 1981. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255186.

Davey:1993:MCB

[Dav93] John T. Davey. Macrofaunal community bioturbation along an estuarine gradient. Aquatic Ecology, 27(2-4):147–153, June 1993. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334777.

deBrabander:1969:PVW

[dB69] K. de Brabander. Het probleem van de waterverontreiniging in Belgie. (Dutch) [The problem of water pollution in Belgium]. Aquatic Ecology, 3(2):34–36, July 1969. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185848.

DeWolf:1981:AMS

[DB81] L. De Wolf and G. J. C. Buth. Asimple method for separating micro-, meio- and macrofauna, and plant litter fractions. Aquatic Ecology, 15(3):175–177, December 1981. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255176.

Daro:1985:CZG

[DB85] M. H. Daro and M. A. Baars. Calculations of zooplankton grazing rates according to a closed, steady-state, three-compartment model applied to different ^{14}C methods. Aquatic Ecology, 19(2):159–170, December 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270762.

Davy-Bowker:2002:MRS

[DB02] John Davy-Bowker. A mark and recapture study of water beetles (Coleoptera: Dytiscidae) in a group of semipermanent and temporary ponds. Aquatic Ecology, 36(3): 435–446, August 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1016550127986.

Duarte:2002:ACL

[DBdF02] P. Duarte, J. M. Bernardo, and L. Cancela da Fonseca. Analysis of coastal lagoon metabolism as a basis for management. Aquatic Ecology, 36(1):3–19, January 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1013394521627.

Dame:2002:ERB

[DBG02a] Richard Dame, David Bushek, and Leah Gregory. Ecosystem response to bivalve density reduction: management implications. Aquatic Ecology, 36(1):51–65, January 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1013354807515.

Degermendzhy:2002:FVH

[DBG02b] Andrei G. Degermendzhy, Victor M. Belolipetsky, and Ramesh D. Gulati. Formation of the vertical heterogeneity in the Lake Shira ecosystem: the biological mechanisms and mathematical model. Aquatic Ecology, 36(2):271–297, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1015621508971.

Dzialowski:2013:CZP

[DBG13] Andrew R. Dzialowski, Joseph L. Bonneau, and Todd R. Gemeinhardt. Comparisons of zooplankton and phytoplankton in created shallow water habitats of the lower Missouri River: implications for native fish. Aquatic Ecology, 47 (1):13–24, March 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9421-0.

Davidova:2011:REB

[DBO11] Martina Dávidová, Radim Blažek, and Markéta Ondračková. The role of the European bitterling (Rhodeus amarus, cyprinidae) in parasite accumulation and transmission in riverine ecosystems. Aquatic Ecology, 45(3):377–387, September 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-011-9361-0.

deBoer:2000:CRA

[dBPG00] W. F. de Boer, T. Pereira, and A. Guissamulo. Comparing recent and abandoned shell middens to detect the impact of human exploitation on the intertidal ecosystem. Aquatic Ecology, 34(3):287–297, September 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009957409421.

DeJonge:1996:RCC

[DBV96] V. N. De Jonge, J. F. Bakker, and M. Van Stralen. Recent changes in the contributions of river Rhine and North Sea to the eutrophication of the western Dutch Wadden Sea. Aquatic Ecology, 30(1):27–39, July 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02092145.

deCarvalho:2023:TSF

[dCAP23] Débora Reis de Carvalho, Carlos Bernardo Mascarenhas Alves, and Paulo Santos Pompeu. Trophic structure of fish assemblages from oligotrophic tropical rivers: evidence of growing assimilation of autochthonous primary producers with the increase in river dimensions. Aquatic Ecology, 57 (2):405–419, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10018-w.

Donagh:2009:PRU

[DCC09] María Elicia Mac Donagh, María Adela Casco, and María Cristina Claps. Plankton relationships under small water level fluctuations in a subtropical reservoir. Aquatic Ecology, 43(2):371–381, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
008-9197-4.

DiSabatino:2020:PDO

[DCC20] Antonio Di Sabatino, Bruno Cicolani, and Giovanni Cristiano. Plant detritus origin and microbial–detritivore interactions affect leaf litter breakdown in a Central Apennine (Italy) cold spring. Aquatic Ecology, 54(2):495–504, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09755-z.

deCarvalho:2019:SIS

[dCdCP19] Débora Reis de Carvalho, Diego Marcel Parreira de Castro, and Paulo Santos Pompeu. Stable isotopes and stomach content analyses indicate omnivorous habits and opportunistic feeding behavior of an invasive fish. Aquatic Ecology, 53(3): 365–381, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09695-3.

Dong:2019:AEP

[DCG19] Jing Dong, Mengyang Chang, and Yunni Gao. Allelopathic effects and potential active substances of Ceratophyllum demersum L. on Chlorella vulgaris Beij. Aquatic Ecology, 53 (4):651–663, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09715-2.

Dorgelo:1993:R

[DCH93] Jaap Dorgelo, Gerhard C. Cadée, and L. W. G. Higler. Review. Aquatic Ecology, 27(1):59–69, July 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02336929.

deCarvalho:2023:DFE

[dCKL23] Cássia de Carvalho, Karina A. Keunecke, and Helena P. Lavrado. Differences in feeding ecology of the pink shrimps Penaeus brasiliensis and P. paulensis (decapoda: Penaeidae) in Brazilian tropical ecosystems. Aquatic Ecology, 57(3): 701–714, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10040-y.

Ding:2019:SPU

[DCL19] Dongbo Ding, Shasha Chen, and Jie Li. Strategies of phosphorus utilization in an astaxanthin-producing green alga Haematococcus pluvialis, a comparison with a bloom-forming cyanobacterium Microcystis wesenbergii. Aquatic Ecology, 53 (4):679–688, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09718-z.

Medeiros:2015:FBG

[dCMMB15] Luciana de Castro Medeiros, Arthur Mattos, and Vanessa Becker. Is the future blue-green or brown? The effects of extreme events on phytoplankton dynamics in a semi-arid man-made lake. Aquatic Ecology, 49(3):293–307, September 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9524-5.

Nunez:2025:IHP

[dCNGF25] María del Carmen Núñez, María Florencia Gutierrez, and Diego Frau. Insights into heterotrophic protozoa assemblages spatiotemporal distribution in urban-agricultural landscapes streams. Aquatic Ecology, 59(1):169–183, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10155-w.

Davids:1979:P

[DD79] C. Davids and Jaap Dorgelo. Publications. Aquatic Ecology, 13(2-3):189–192, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02284754.

Dorgelo:1981:PLM

[DD81] Jaap Dorgelo and Ilse De Graaf Bierbrauwer. Phytoplankton in Lake Maarsseveen I (1977–1979). Aquatic Ecology, 15(1-2):29–40, October 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260256.

DeKroon:1984:CTN

[DD84] Hans De Kroon and Hans De Jong. The colonization of two new brackish water habitats on the isle of Texel (The Netherlands). Aquatic Ecology, 18(1):72, June 1984. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256680.

Domaizon:1999:ESI

[DD99] Isabelle Domaizon and Jean Devaux. Experimental study of the impacts of silver carp on plankton communities of eutrophic Villerest reservoir (France). Aquatic Ecology, 33(2): 193–204, September 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009984713183.

Davidson:2017:LHT

[DD17] Andrew T. Davidson and Nathan J. Dorn. Life history traits determine the differential vulnerability of native and invasive apple snails (Pomacea spp.) to a shared juvenile-stage predator. Aquatic Ecology, 51(3):331–341, September 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9620-9.

Dorgelo:1988:P

[DDC88] Jaap Dorgelo, C. Davids, and P. Coesel. Publications. Aquatic Ecology, 22(2):203–210, December 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256825.

Damare:2020:PSR

[DDC20] Varada S. Damare, Priya M. D’Costa, and Svetlana Cardozo. Preliminary study on the response of marine fungoid protists, the thraustochytrids, to lipid extracts of diatoms. Aquatic Ecology, 54(1):355–367, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09747-z.

Daan:1984:P

[DDD84] Niels Daan, Jaap Dorgelo, and C. Davids. Publications. Aquatic Ecology, 18(1):77–81, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256687.

Dakou:2007:DTM

[DDD07] Eleni Dakou, Tom D’heygere, and Niels De Pauw. Decision tree models for prediction of macroinvertebrate taxa in the River Axios (Northern Greece). Aquatic Ecology, 41(3): 399–411, September 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-006-9058-y.

Davoult:1993:LTC

[DDF93] D. Davoult, J. M. Dewarumez, and S. Frontier. Long-term changes (1979–1990) in three benthic communities (eastern English Channel): Use of factor analysis and rank-frequency diagrams for studying structural developments. Aquatic Ecology, 27(2-4):415–426, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02334803.

Donlon:2019:GAM

[DDL19] Erica M. Y. Donlon, Erin L. Damsteegt, and Miles D. Lamare. Growth and age of the midget octopus, Octopus huttoni. Aquatic Ecology, 53(4):689–706, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09719-y.

Diniz:2022:TBE

[DdNM22] Anamaria Silva Diniz and Ariadne do Nascimento Moura. Top-down and bottom-up effects of fish on a macrophytemediated trophic network: a mesocosm approach. Aquatic Ecology, 56(4):1157–1175, December 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09976-4.

Do:2024:ARM

[DDPT24] Anh Ngoc Thi Do, Tuyet Anh Thi Do, Long Van Pham, and Hau Duc Tran. Assessment of the role of mangroves for Periophthalmus modestus applying machine learning and remote sensing: a case study in a large estuary from Vietnam. Aquatic Ecology, 58(3):879–893, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10111-8.

Do:2024:DFL

[DDT24] Anh Ngoc Thi Do, Tuyet Anh Thi Do, and Hau Duc Tran. Distribution of fish larvae and juveniles on salinity in an estuary predicted from remote sensing and fuzzy logic approach. Aquatic Ecology, 58(3):983–998, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10119-0.

DeKroon:1984:MDB

[DDV84] Hans De Kroon, Hans De Jong, and Jos T. A. Verhoeven. Macrofauna distribution in brackish inland waters in relation to chlorinity and other factors. Aquatic Ecology, 18(1): 71, June 1984. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256679.

DeHaan:1991:SDD

[DDV91] H. De Haan, T. De Boer, and J. Voerman. Size distribution of dissolved (< 200 nm) organic carbon and aluminium in alkaline and humic lakes. Aquatic Ecology, 24(2):145–151, April 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260432.

Dong:2023:GMR

[DDY^{+}23] Jing Dong, Dujuan Dai, Yue Yang, Feihu Wang, Yang Zhang, Man Zhang, Yunni Gao, Xiaofei Gao, and Xuejun Li. Growth and morphological responses of Scenedesmus obliquus to submerged macrophyte Egeria densa. Aquatic Ecology, 57(1): 127–138, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-10001-x.

Dong:2024:ATC

[DDZ^{+}24] Aiguo Dong, Huidong Dong, Tianmiao Zhang, Xuejie Jing, Hui He, and Junfeng Huo. The acute toxicity of cadmium on turtle Mauremys reevesii. Aquatic Ecology, 58(4):1217–1223, December 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10136-z.

DeJonge:1974:CBC

[De 74] V. N. De Jonge. Classification of brackish coastal inland waters. Aquatic Ecology, 8(1-2):29–39, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254903.

DeGreef:1976:BTE

[De 76a] E. De Greef. The behaviour of trace elements in storage reservoirs. Aquatic Ecology, 10(3):155–163, December 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02263364.

DeLange:1976:SMD

[De 76b] L. De Lange. Survey of the macrovegetation in a ditch near Tienhoven (Utrecht, the Netherlands). Aquatic Ecology, 10(1):37–39, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02308499.

DeJonge:1977:DGS

[De 77a] V. N. De Jonge. Density gradient separation of microbenthos from estuarine sediments. Aquatic Ecology, 11(1):13– 14, April 1977. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02282206.

DeWolf:1977:BOV

[De 77b] P. De Wolf. “Biologisch Onderzoek Veenkoloniaal Afvalwater”: an introduction and a review of a waste-water research project. Aquatic Ecology, 11(1):7–11, April 1977. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02282203.

DeWolf:1978:EOW

[De 78] P. De Wolf. Effects of organic waste on primary and secondary production in the Ems–Dollard estuary. Aquatic Ecology, 12(3-4):260–272, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02259187.

DeGroot:1979:EWB

[De 79a] A. A. De Groot. Über Einige Wenig Bekannten Rhizopoden. (German) [On some little-known rhizopods]. Aquatic Ecology, 13(1):34–49, April 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260307.

DeNie:1979:ETE

[De 79b] H. W. De Nie. Effects of thermal effluents from the Bergum power station on the zooplankton in Lake Bergum. Aquatic Ecology, 13(2-3):100, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02284739.

DeZwart:1979:LTE

[De 79c] Dick De Zwart. The long term effects of subtle temperature rises on aquatic micro-ecosystems. Aquatic Ecology, 13 (2-3):96, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02284736.

DeOude:1980:CDI

[De 80a] N. T. De Oude. Comments of the detergent industry. Aquatic Ecology, 14(1-2):94–97, August 1980. CODEN AQECF9.

ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260277.

DeRooij:1980:ACM

[De 80b] N. M. De Rooij. Application of a chemical model to combat eutrophication. Aquatic Ecology, 14(1-2):106–115, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260279.

DeWilde:1980:DMB

[De 80c] P. A. W. J. De Wilde. Dynamics and metabolism of the benthos of the Wadden Sea. Aquatic Ecology, 14(3):216–218, December 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260124.

DeHaan:1984:EMS

[De 84a] H. De Haan. Effects of metal speciation on growth of phytoplankton with special reference to iron. Aquatic Ecology, 18 (2):85–94, December 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257047.

DeVries:1984:DAS

[De 84b] B. J. De Vries. Diatom assemblies in some moorland pools in the Drenthian district (the Netherlands). Aquatic Ecology, 18(1):3–10, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256667.

DeMeester:1997:NME

[De 97] Luc De Meester. Neutral markers, ecologically relevant traits, and the structure of genetic variation in Daphnia. Aquatic Ecology, 31(1):79–87, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009939301712.

DeLange:2000:AUV

[De 00] H. J. De Lange. The attenuation of ultraviolet and visible radiation in Dutch inland waters. Aquatic Ecology, 34(3):

215–226, September 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009943211779.

DeJonge:1993:RMC

[DE93] V. N. De Jonge and M. M. Engelkes. The role of mineral compounds and chemical conditions in the binding of phosphate in the Ems estuary. Aquatic Ecology, 27(2-4):227–236, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334786.

DeAngelis:2010:FSL

[DeA10] Donald L. DeAngelis. Foreword to the Siberian lakes special issue. Aquatic Ecology, 44(3):479, September 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9337-5.

Deelder:1969:VAV

[Dee69] C. L. Deelder. Visserijkundige aspecten van het Ijsselmeer. (Dutch) [Fishery aspects of Lake IJssel]. Aquatic Ecology, 3(1):6, March 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185839.

Degermendzhi:2010:CMP

[Deg10] A. Degermendzhi. Coexistence of microbial populations and autostabilization of regulating factors in continuous culture: theory and experiments. Aquatic Ecology, 44(3):541–560, September 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-010-9325-9.

Delettre:1992:TCC

[Del92] Y. R. Delettre. Terrestrial Chironomidae: Contribution of local emergence to global aerial flow in a heterogeneous environment. Aquatic Ecology, 26(2-4):269–271, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255251.

Dennert:1974:TDI

[Den74] Henk G. Dennert. Tolerance differences and interspecific competition in three members of the amphipod genus Gammarus. Aquatic Ecology, 8(1-2):109, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254910.

Denys:1994:DAA

[Den94] Luc Denys. Diatom assemblages along a former intertidal gradient: a palaeoecological study of a subboreal clay layer (Western coastal plain, Belgium). Aquatic Ecology, 28(1):85– 96, March 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334249.

Denny:2006:OWN

[Den06] Mark W. Denny. Ocean waves, nearshore ecology, and natural selection. Aquatic Ecology, 40(4):439–461, December 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-004-5409-8.

DosSantosSilva:2022:FFD

[DEP22] Nathália Carina Dos Santos Silva, Bruno Eleres Soares, and Miriam Pilz Albrecht. Fish functional diversity is less impacted by mining than fish taxonomic richness in an Amazonian stream system. Aquatic Ecology, 56(3):815–827, September 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-022-09946-w.

Devetter:2009:CRB

[Dev09] Miloslav Devetter. Clearance rates of the bdelloid rotifer, Habrotrocha thienemanni, a tree-hole inhabitant. Aquatic Ecology, 43(1):85–89, March 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
007-9160-9.

Dewaide:1980:IAS

[Dew80] J. H. Dewaide. Interactions among scientific research, government policy and actual practice in fighting the eutrophication problem. Aquatic Ecology, 14(1-2):125–132, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260281.

Dyduch-falniowska:1982:ODD

[Df82] Anna Dyduch-falniowska. Oscillations in density and diversity of Pisidium communities in two biotopes in Southern Poland. Aquatic Ecology, 16(2-3):123–132, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255366.

Dyduch-Falniowska:1983:ASP

[DF83] Anna Dyduch-Falniowska. Age structure of the populations of Pisidium species from two localities in southern Poland. Aquatic Ecology, 17(2):111–117, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02280820.

DeZwart:1990:MTO

[DF90] Dick De Zwart and Ariejan Folkerts. Monitoring the toxicity of organic compounds dissolved in Rhine water. Aquatic Ecology, 24(1):5–12, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256744.

Dayton:2011:AND

[DF11] Gage H. Dayton and Lee A. Fitzgerald. The advantage of no defense: predation enhances cohort survival in a desert amphibian. Aquatic Ecology, 45(3):325–333, September 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9356-x.

Bomfim:2021:WPD

[dFBdGGMLT21] Francieli de Fátima Bomfim, Maria da Graça Gama Melão, and Fábio Amodêo Lansac-Tôha. Warming and predator drive functional responses of three subtropical cladocerans. Aquatic Ecology, 55(3):903–914, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09870-5.

deGroot:1971:GGV

[dG71] A. J. de Groot. Geochemisch gedrag van zware metalen in deltagebieden. (Dutch) [Geochemical behavior of heavy metals in delta regions]. Aquatic Ecology, 5(4):172–176, December 1971. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185958.

Dorgelo:1984:PDS

[DG84] Jaap Dorgelo and Maarten Gorter. Preliminary data on size composition and settlement of Dreissena polymorpha (Pallas) (Mollusca: Bivalvia) in lakes differing in trophic state. Aquatic Ecology, 18(2):159–163, December 1984. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02257055.

Degermendzhy:2002:UMB

[DG02] Andrei G. Degermendzhy and Ramesh D. Gulati. Understanding the mechanisms of blooming of phytoplankton in Lake Shira, a saline lake in Siberia (the Republic of Khakasia). Aquatic Ecology, 36(2):333–340, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1016386720912.

Dumnicka:2006:DBF

[DG06] Elżbieta Dumnicka and Joanna Galas. Distribution of benthic fauna in relation to environmental conditions in an inundated opencast sulphur mine (Piaseczno Reservoir, Southern Poland). Aquatic Ecology, 40(2):203–210, June 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-6040-z.

Domingues:2017:STI

[DGB17] Rita B. Domingues, Cátia C. Guerra, and Ana B. Barbosa. Short-term interactive effects of ultraviolet radiation, carbon dioxide and nutrient enrichment on phytoplankton in a shallow coastal lagoon. Aquatic Ecology, 51(1):91–105, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-016-9601-4.

Du:2015:AIT

[DGBL15] Xue Du, Emili García-Berthou, and Zhongjie Li. Analyzing the importance of top-down and bottom-up controls in food webs of Chinese lakes through structural equation modeling. Aquatic Ecology, 49(2):199–210, June 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9518-3.

Dedecker:2006:DSM

[DGD06] Andy P. Dedecker, Peter L. M. Goethals, and Niels De Pauw. Development of an in-stream migration model for Gammarus pulex L. (Crustacea, Amphipoda) as a tool in river restoration management. Aquatic Ecology, 40(2):249–261, June 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-9022-2.

Domingues:2015:NLL

[DGG15] Rita B. Domingues, Cátia C. Guerra, and Helena M. Galvão. Are nutrients and light limiting summer phytoplankton in a temperate coastal lagoon? Aquatic Ecology, 49(2):127– 146, June 2015. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-015-9512-9.

Dominguez-Godino:2020:HAS

[DGGW20] Jorge A. Domínguez-Godino and Mercedes González-Wangüemert. Habitat associations and seasonal abundance patterns of the sea cucumber Holothuria arguinensis at Ria Formosa coastal lagoon (South Portugal). Aquatic Ecology, 54(1):337–354, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09746-0.

Dharmarathna:2024:APS

[DGH^{+}24] Dilshi Dharmarathna, Ridmi Galagedara, Sivaperumaan Himanujahn, Shiromi Karunaratne, and Bandunee Athapattu. Assessment of pollution state of Beira Lake in Sri Lanka using water quality index, trophic status, and principal component analysis. Aquatic Ecology, 58(2):159–174, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10052-8.

Das:2021:EAH

[DGN21] Priya Brata Das, Mangesh Gauns, and Syed Wajih Ahmad Naqvi. Environmental association of heterotrophic micro-eukaryotes in the varying biogeochemical regimes of the Arabian Sea, resolved via high-throughput sequencing. Aquatic Ecology, 55(3):807–824, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09862-5.

deHaan:1971:HC

[dH71a] H. de Haan. Humus compounds. Aquatic Ecology, 5(2):78– 79, June 1971. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185943.

denHartog:1971:TV

[dH71b] C. den Hartog. Thermische verontreiniging. Aquatic Ecology, 5(1):49–52, March 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185887.

denHartog:1972:B

[dH72] C. den Hartog. Boeken. Aquatic Ecology, 6(2):43–45, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02304091.

Davids:1992:P

[DH92] C. Davids and F. Heinis. Preface. Aquatic Ecology, 26(2-4): 96, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255228.

DePauw:2001:BIS

[DH01] Niels De Pauw and Steven Heylen. Biotic index for sediment quality assessment of watercourses in Flanders, Belgium. Aquatic Ecology, 35(2):121–133, June 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011478427152.

Dokulil:2009:ALT

[DH09] Martin T. Dokulil and Alois Herzig. An analysis of long-term winter data on phytoplankton and zooplankton in Neusiedler See, a shallow temperate lake, Austria. Aquatic Ecology, 43(3):715–725, September 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9282-3.

DiGiacopo:2022:ENL

[DH22] Devin G. DiGiacopo and Jessica Hua. The effects of novel leaf litter deposition on competitive, predator–prey and host–parasite interactions of American toad larvae. Aquatic Ecology, 56(1):59–73, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
021-09893-y.

denHartog:1976:BER

[dHC76] C. den Hartog and H. E. Coomans. Biosystematics and ecology of Rhizoclonium riparium (Roth) Harv. (Chlorophyceae: Cladophorales) in the estuarine area of the rivers Rhine, Meuse and Scheldt. Aquatic Ecology, 10(1):78–80, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02308506.

Dumont:2021:CTP

[DHEG21a] Henri J. Dumont, Bo-Ping Han, and Manuel Elias-Gutierrez. Correction to: Toward a phylogeny and biogeography of Diaphanosoma (Crustacea: Cladocera). Aquatic Ecology, 55 (4):1223, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09879-w.

Dumont:2021:TPB

[DHEG21b] Henri J. Dumont, Bo-Ping Han, and Manuel Elias-Gutierrez. Toward a phylogeny and biogeography of Diaphanosoma (Crustacea: Cladocera). Aquatic Ecology, 55(4):1207–1222, December 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09819-0.

Douglas:2016:GPD

[DHL16] G. B. Douglas, D. P. Hamilton, and M. Lurling. Guiding principles for the development and application of solidphase phosphorus adsorbents for freshwater ecosystems. Aquatic Ecology, 50(3):385–405, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9575-2.

Dare:2024:DCC

[DHPN^{+}24] Luana Daré, Raoul Henry, Gilmar Perbiche-Neves, Maria Carolina A. Castilho, and Jorge L. Portinho. Diversity and composition of cladoceran egg bank in sediments and macrophyte roots from the littoral zone of tropical lakes. Aquatic Ecology, 58(3):571–581, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10091-9.

Davids:1981:HDF

[DHW81] C. Davids, C. F. Heijnis, and J. E. Weekenstroo. Habitat differentiation and feeding strategies in water mites in Lake Maarsseveen I. Aquatic Ecology, 15(1-2):87–91, October 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260262.

DeWilde:1992:IEN

[DJD92] P. A. W. J. De Wilde, M. I. Jenness, and G. C. A. Duineveld. Introduction into the ecosystem of the North Sea: Hydrography, biota, and food web relationships. Aquatic Ecology, 26(1):7–18, November 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02298024.

deJalon:2007:EIH

[dJG07] Diego García de Jalón and Javier Gortázar. Evaluation of instream habitat enhancement options using fish habitat simulations: case-studies in the river Pas (Spain). Aquatic Ecology, 41(3):461–474, September 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9030-x.

deKloet:1971:EVI

[dK71] W. A. de Kloet. Het eutrofieringsproces van het Ijsselmeergebied. (Dutch) [The eutrophication process of the Lake IJssel region]. Aquatic Ecology, 5(1):23–38, March 1971. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185885.

Dorgelo:1980:AMA

[DK80] Jaap Dorgelo and Kees Koning. Avoidance of macrophytes and additional notes on avoidance of the shore by ‘Acanthodiaptomus denticornis’ (Wierzejski, 1887) from Lake Pavin (Auvergne, France). Aquatic Ecology, 14(3):196–208, December 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260121.

Dubovskaya:2005:SDP

[DKI05] Olga P. Dubovskaya, Elena P. Klimova, and Elena A. Ivanova. Seasonal dynamic of phototrophic epibionts on crustacean zooplankton in a eutrophic reservoir with cyanobacterial bloom. Aquatic Ecology, 39(2):167–180, June 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-5001-2.

deLange:1974:AWQ

[dL74] L. de Lange. The assessment of water quality by the local stand of macrophytic vegetation (abstract). Aquatic Ecology, 8(3):288–289, December 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02257505.

Dorgelo:1982:P

[DL82] Jaap Dorgelo and P. Leentvaar. Publications. Aquatic Ecology, 16(1):116–120, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255423.

Dorgelo:1989:IAP

[DL89] Jaap Dorgelo and Pim E. G. Leonards. Isolation and analysis of particulate detritus as an aid to studies on the nutritional ecology of deposit feeders. Aquatic Ecology, 23(2):201–206, December 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256738.

deLima:2022:SDS

[dLFTB22] Renan C. de Lima, Valentina Franco-Trecu, and Silvina Botta. Segregation of diets by sex and individual in South American fur seals. Aquatic Ecology, 56(1):251–267, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09915-9.

Dossi:2020:AMS

[DLG20] Florian Dossi, Patrick Leitner, and Wolfram Graf. Age matters: substrate-specific colonization patterns of benthic invertebrates on installed large wood. Aquatic Ecology, 54(3): 741–760, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09772-y.

Dresscher:1971:ONM

[DM71] Th. G. N. Dresscher and L. R. Mur. Een onderzoek naar de micro-organismen van enkele poelen in Zuid-Limburg. (Dutch) [A study into the micro-organisms of some pools in South Limburg]. Aquatic Ecology, 5(4):181–202, December 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185962.

DiLonardo:2004:DSD

[DM04] Steve S. Di Lonardo and Edward J. Maly. Dissolved solids do not induce diapause in the calanoid copepod Aglaodiaptomus leptopus. Aquatic Ecology, 38(3):425–432, September 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000035176.77771.06.

Costa:2023:SVP

[dMCFS^{+}23] Fabiana de Matos Costa, Carla Ferragut, Nadson Ressyé Simões, Geraldo José Peixoto Ramos, Daniela Mariano Lopes da Silva, and Carlos Wallace do Nascimento Moura. Spatial variation of periphytic desmid community structure on emergent macrophytes in a tropical urban watershed. Aquatic Ecology, 57(3):747–763, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10045-7.

Duck:1995:CST

[DMD95] R. W. Duck, J. McManus, and J. J. Diez. Comparative study of two largely infilled estuaries: The Eden Estuary (Scotland) and the Ria de Foz (Spain). Aquatic Ecology, 29 (3-4):203–210, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02084217.

Duarte:2019:TEB

[DMD19] Cleber Duarte, Anne E. Magurran, and Cláudia P. Deus. Trophic ecology of benthic fish assemblages in a lowland river in the Brazilian Amazon. Aquatic Ecology, 53(4):707–718, December 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09720-5.

DeRoos:1997:WIL

[DMM97] A. M. De Roos, E. McCauley, and W. W. Murdoch. What individual life histories can (and cannot) tell about population dynamics. Aquatic Ecology, 31(1):37–45, March 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009995420368.

Duursma:1988:EPE

[DMN88] E. K. Duursma, A. G. A. Merks, and J. Nieuwenhuize. Exchange processes in estuaries such as the Westerschelde, an overview. Aquatic Ecology, 22(1):7–20, August 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256777.

Dulnuan:2023:MIF

[DN23] Minerva M. Dulnuan and Jones T. Napaldet. Mosaic interplay of floral diversity, soil properties, disturbance intensity and elevation in the riparian ecosystem under semisubsistence agriculture of Cordillera Central Range, Northern Philippines. Aquatic Ecology, 57(3):613–631, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10033-x.

dasNeves-da-Silva:2021:EIE

[dNdSBJdCeS21] Dener das Neves-da Silva, Vitor Nelson Texeira Borges-Júnior, and Ana Maria Paulino Telles de Carvalho-e Silva. Effects of intrinsic and extrinsic factors on the prevalence of the fungus Batrachochytrium dendrobatidis (Chytridiomycota) in stream tadpoles in the Atlantic Forest domain. Aquatic Ecology, 55(3):891–902, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09869-y.

Filho:2021:STE

[dNFdNM21] Silvano Lima do Nascimento Filho and Ariadne do Nascimento Moura. Strong top-down effects of omnivorous fish and macroinvertebrates on periphytic algae and macrophytes in a tropical reservoir. Aquatic Ecology, 55(2):667–680, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09853-6.

Defranoux:2022:CSI

[DNG22] Fanny Defranoux, Simona Noè, and Paola Gianguzza. Chemoecological study of the invasive alga Caulerpa taxifolia var. distichophylla from the Sicilian coast. Aquatic Ecology, 56(2):447–457, June 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09925-7.

Duggan:2021:LTZ

[DÖD21] Ian C. Duggan, Deniz Özkundakci, and Bruno O. David. Long-term zooplankton composition data reveal impacts of invasions on community composition in the Waikato lakes, New Zealand. Aquatic Ecology, 55(4):1127–1142, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09803-8.

Donze:1979:MEH

[Don79] M. Donze. Measurements of the effect of heat shocks on survival and growth of natural zooplankton populations. Aquatic Ecology, 13(2-3):97, December 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284737.

Doornbos:1982:CFF

[Doo82] G. Doornbos. Changes in the fish fauna of the former Grevelingen estuary, before and after the clousure in 1971. Aquatic Ecology, 16(2-3):279–283, December 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255381.

Dorgelo:1970:B

[Dor70a] J. Dorgelo. Boeken. Aquatic Ecology, 4(3):129–130, November 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185869.

Dorgelo:1970:RVG

[Dor70b] J. Dorgelo. Respiratie van gammariden in relatie tot de saliniteit van het Milieu. (Dutch) [Respiration of gammarids in relation to the salinity of the environment]. Aquatic Ecology, 4(3):145–146, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185876.

Dorgelo:1971:UIL

[Dor71] J. Dorgelo. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 5(2):67–68, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185939.

Dorgelo:1972:NT

[Dor72] J. Dorgelo. Nieuw tijdschrift. (Dutch) [New magazine]. Aquatic Ecology, 6(1):5–6, March 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02336203.

Dorgelo:1973:P

[Dor73] Jaap Dorgelo. Publications. Aquatic Ecology, 7(3):127–136, September 1973. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02275614.

Dorgelo:1974:CEG

[Dor74] Jaap Dorgelo. Comparative ecophysiology of gammarids (Crustacea: Amphipoda) from marine, brackish and freshwater habitats, exposed to the influence of salinitytemperature combinations. Aquatic Ecology, 8(1-2):90–108, July 1974. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02254909.

Dorgelo:1976:FOD

[Dor76a] Jaap Dorgelo. Further observations on drift in two brooks in Auvergne (France). Aquatic Ecology, 10(1):30, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02308496.

Dorgelo:1976:IFZ

[Dor76b] Jaap Dorgelo. Intertidal fucoid zonation and desiccation. Aquatic Ecology, 10(2):115–122, September 1976. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02282219.

Dorgelo:1977:STS

[Dor77] Jaap Dorgelo. Salt tolerance, salt preference and blood sodium regulation in Gammarus roeseli Gervais, 1835 (Crustacea: Amphipoda). Aquatic Ecology, 11(2):56–61, September 1977. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02265710.

Dorgelo:1979:ITB

[Dor79] Jaap Dorgelo. The influence of temperature on the blood osmoconcentration of an osmoconforming and an osmoregulating crab. Aquatic Ecology, 13(1):22–28, April 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260304.

Dorgelo:1980:MHV

[Dor80a] Jaap Dorgelo. In memoriam H. Van Wijk. Aquatic Ecology, 14(3):228, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260136.

Dorgelo:1980:P

[Dor80b] Jaap Dorgelo. Publications. Aquatic Ecology, 14(3):226–227, December 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260134.

Dorgelo:1986:P

[Dor86] Jaap Dorgelo. Publications. Aquatic Ecology, 20(1-2):263– 269, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02291169.

Dorgelo:1987:DFP

[Dor87] Jaap Dorgelo. Density fluctuations in populations (1982– 1986) and biological observations of Potamopyrgus jenkinsi in two trophically differing lakes. Aquatic Ecology, 21(1):95– 110, July 1987. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255459.

Dorgelo:1989:R

[Dor89] Jaap Dorgelo. Reviews. Aquatic Ecology, 23(2):213–222, December 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256740.

Dorgelo:1992:E

[Dor92] Jaap Dorgelo. Editorial. Aquatic Ecology, 26(1):2, November 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02298022.

Dorgelo:1995:E

[Dor95] Jaap Dorgelo. Editorial. Aquatic Ecology, 29(1):II, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02061784.

Dorgelo:1996:E

[Dor96] Jaap Dorgelo. Editorial. Aquatic Ecology, 30(1):I, July 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02092141.

dePauw:1971:MPW

[dP71] N. de Pauw. Milieu en plankton in de Westerschelde. (Dutch) [Environment and plankton in the Western Scheldt]. Aquatic Ecology, 5(1):3–16, March 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185883.

Dame:1997:BCC

[DP97] Richard F. Dame and Theo C. Prins. Bivalve carrying capacity in coastal ecosystems. Aquatic Ecology, 31(4):409–421, December 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009997011583.

dePadua:2022:AOC

[dPBCP22] Stella Maris Feitosa de Pádua, Mônica Lúcia Botter-Carvalho, and Carlos Daniel Pérez. The alien octocoral Carijoa riisei is a biogenic substrate multiplier in artificial Brazilian shipwrecks. Aquatic Ecology, 56(1):183–200, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09908-8.

DeLaune:2003:SUG

[DPL03] R. D. DeLaune, S. R. Pezeshki, and C. W. Lindau. Sensitivity of US Gulf of Mexico coastal marsh vegetation to crude oil: Comparison of greenhouse and field responses. Aquatic Ecology, 37(4):351–360, October 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000007047.42636.96.

dePauw:1991:NIC

[dPNF91] Niels de Pauw and Eddy Naessens-Foucquaert. Nutrientinduced competition between two species of marine diatoms. Aquatic Ecology, 25(1):23–27, October 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02259585.

DeMott:2004:PSV

[DPT04] William R. DeMott, Bryn J. Pape, and Alan J. Tessier. Patterns and sources of variation in Daphnia phosphorus content in nature. Aquatic Ecology, 38(3):433–440, September 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000035183.53389.66.

deRuiter:1992:RBN

[dR92] Marien de Ruiter. Restoration of Botshol (The Netherlands) by reduction of external nutrient load: Botshol and the Dutch governmental program of nutrient reduction. Aquatic Ecology, 25(3):261–264, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02270811.

DeJonge:1996:HSW

[DR96] Victor N. De Jonge and Johan F. Ruiter. How subtidal were the ’subtidal beds’ of Zostera marina L. before the occurrence of the wasting disease in the early 1930’s? Aquatic Ecology, 30(2-3):99–106, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02272231.

Dresscher:1968:OON

[Dre68] Th. G. N. Dresscher. Een orienterend onderzoek naar het voorkomen van Mikrofyten in het stratiotetum van de plas venematen. (Dutch) [An exploratory study on the occurrence of microphytes in the stratiotetum of the venous venous]. Aquatic Ecology, 2(2):45–52, June 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185809.

deResende:2021:SFS

[dRFC21] Bethânia Oliveira de Resende, Victor Rennan Santos Ferreira, and Helena Soares Ramos Cabette. Seasonal fluctuations in the structure of the larval odonate community of a stream in the Cerrado–Amazon forest transition zone. Aquatic Ecology, 55(3):861–873, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09865-2.

Dallas:2018:TTR

[DRM18] Helen F. Dallas and Nicholas A. Rivers-Moore. Temporal thermal refugia and seasonal variation in upper thermal limits of two species of riverine invertebrates: the amphipod, Paramelita nigroculus, and the mayfly, Lestagella penicillata. Aquatic Ecology, 52(4):333–349, December 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9667-2.

Delgado:2009:DAM

[DRN09] Juan D. Delgado, Rodrigo Riera, and Jorge Núñez. Distribution and abundance of meiofauna in intertidal sand substrata around Iceland. Aquatic Ecology, 43(2):221–233, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9200-0.

Dronkers:1984:IFM

[Dro84] J. Dronkers. Import of fine marine sediment in tidal areas. Aquatic Ecology, 18(1):74–75, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256683.

duSaar:1968:D

[dS68] A. du Saar. A. Den Dulk. Aquatic Ecology, 2(3):82–83, September 1968. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185815.

Davids:1975:RBW

[DS75] C. Davids and C. J. F. Schoots. Relations between water mites of the genus Hydrachna and the corixids Sigara striata and Cymatia coleoptrata. Aquatic Ecology, 9(2):89, September 1975. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257643.

Cardoso:2009:HDP

[dSCdMM09] Luciana de Souza Cardoso and David da Motta Marques. Hydrodynamics-driven plankton community in a shallow lake. Aquatic Ecology, 43(1):73–84, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9151-x.

daSilva:2023:SDA

[dSCT23] A. R. da Silva, J. C. Creed, and M. Tavares. Species diversity and abundance of mobile crustaceans associated with living and dead colonies of the invasive sun coral Tubastraea. Aquatic Ecology, 57(2):529–541, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10027-9.

Davids:1987:CLZ

[DSD87] C. Davids, M. Stolp, and C. J. De Groot. The cladocerans of the littoral zone of Lake Maarsseveen I. Aquatic Ecology, 21(1):71–79, July 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255457.

Dantas:2022:TIO

[dSDLdA22] Renato Junqueira de Souza Dantas, Tatiana Silva Leite, and Cristiano Queiroz de Albuquerque. The trophic interactions of Octopus insularis in the food web of a pristine tropical atoll: a baseline for management and monitoring under environmental changes. Aquatic Ecology, 56(1):269–284, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09917-7.

Duarte:2020:ACP

[dSDMD20] Rafaela Cristina de Souza Duarte, Ellori Laíse Silva Mota, and Thelma Lúcia Pereira Dias. Algal complexity positively affects the abundance, richness and diversity of molluscan assemblages of a semiarid hypersaline mangrove. Aquatic Ecology, 54(4):1001–1013, December 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09789-3.

Ferrao-Filho:2003:EUC

[dSFFA03] Aloysio da Silva Ferrão-Filho and Sandra Maria F. O. Azevedo. Effects of unicellular and colonial forms of toxic Microcystis aeruginosa from laboratory cultures and natural populations on tropical cladocerans. Aquatic Ecology, 37 (1):23–35, January 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1022133329940.

Das:2024:VTT

[DSG^{+}24] Basanta Kumar Das, Dhruba Jyoti Sarkar, Pranab Gogoi, Saurav Kumar Nandy, Arghya Kunui, Manisha Bhor, and Amiya Kumar Sahoo. Variation in thermal trait and plankton assemblage pattern induced by coal power plant discharge in river Ganga. Aquatic Ecology, 58(3):759–773, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10103-8.

daSilva:2017:CSF

[dSGD17] Jislaine Cristina da Silva, Éder André Gubiani, and Rosilene Luciana Delariva. Coexisting small fish species in lotic neotropical environments: evidence of trophic niche differentiation. Aquatic Ecology, 51(2):275–288, June 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9616-5.

Dubey:2012:IHS

[DSL12] Vineet Kumar Dubey, Uttam Kumar Sarkar, and Wazir Singh Lakra. The influence of habitat on the spatial variation in fish assemblage composition in an unimpacted tropical River of Ganga basin, India. Aquatic Ecology, 46(2):165–174, June 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-012-9389-9.

Mesquita:1994:PMC

[dSLM94] Hilda de Souza Lima Mesquita. Planktonic microbial community oxygen consumption rate in Cananéia waters (25^{o}S 48^{o}W), Brazil. Aquatic Ecology, 28(3-4):441–451, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334215.

Longo:2021:SAG

[dSLML21] Pedro Augusto dos Santos Longo, Karine Ferreira Ribeiro Mansur, and Fosca Pedini Pereira Leite. Sargassum- associated gastropod and amphipod assemblages in relation to metal pollution in a semi-enclosed bay. Aquatic Ecology, 55 (2):623–646, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09851-8.

Dai:2009:LSB

[DSN09] Jihong Dai, Ming-Yi Sun, and John E. Noakes. A laboratory study on biochemical degradation and microbial utilization of organic matter comprising a marine diatom, land grass, and salt marsh plant in estuarine ecosystems. Aquatic Ecology, 43(4):825–841, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9211-x.

daSilva:2010:FCP

[dSPdS10] H. Pains da Silva, A. C. Petry, and C. J. da Silva. Fish communities of the Pantanal wetland in Brazil: evaluating the effects of the upper Paraguay river flood pulse on baía Caiçara fish fauna. Aquatic Ecology, 44(1):275–288, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9289-9.

daSilva:2024:EIF

[dSSB24] Jislaine Cristina da Silva, Claudemir Martins Soares, and Andréa Bialetzki. Effect of an invasive fish species on nutrient cycling and on the community structure: an experimental approach. Aquatic Ecology, 58(4):1143–1156, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10129-y.

daSilva:2023:CTS

[dSSR^{+}23] Alexandre R. da Silva, Renan A. P. Santos, Gabriel F. B. Rodrigues, Neida R. Vieira, Bruna K. Baroni, and Rafael A.

Gregati. Coexistence of two sympatric hermit crab species from South Brazil: the effect of the shell adequacy index and circadian movement patterns. Aquatic Ecology, 57(2):459– 469, June 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10022-0.

Santos:2021:IEM

[dSSSES21] Gleice de Souza Santos, Edissa Emi Cortez Silva, and Eneida M. Eskinazi-Sant’Anna. Impacts of exposure to mine tailings on zooplankton hatching from a resting egg bank. Aquatic Ecology, 55(2):545–557, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09844-7.

Dorgelo:1983:P

[DT83] Jaap Dorgelo and Harry H. Tolkamp. Publications. Aquatic Ecology, 17(1):91–93, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255200.

DeJonge:1992:MON

[DT92] V. N. De Jonge and C. C. Ten Hallers-Tjabbes. The message of ‘The Other North Sea’. Aquatic Ecology, 26(1):3–6, November 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02298023.

Do:2023:ADL

[DT23a] Anh Ngoc Thi Do and Hau Duc Tran. Application of deep learning in assessing the impact of flooding on the endangered freshwater fish Neolissochilus benasi (Cyprinidae) in a northern province of Vietnam. Aquatic Ecology, 57(4): 951–967, December 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10056-4.

Do:2023:PAA

[DT23b] Anh Ngoc Thi Do and Hau Duc Tran. Potential application of artificial neural networks for analyzing the occurrences of fish larvae and juveniles in an estuary in northern Vietnam. Aquatic Ecology, 57(4):813–831, December 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09959-5.

Dias:2022:EHC

[DTA22] Rosa Maria Dias, Raffael Marcos Tófoli, and Angelo Antonio Agostinho. Effects of habitat complexity on trophic interactions of three congeneric fish species. Aquatic Ecology, 56(3): 877–889, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09954-w.

Degerman:2019:OHU

[DTC19] Erik Degerman, Carl Tamario, and Olle Calles. Occurrence and habitat use of European eel (Anguilla anguilla) in running waters: lessons for improved monitoring, habitat restoration and stocking. Aquatic Ecology, 53(4):639–650, December 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09714-3.

Davids:1994:TSP

[DTD94] C. Davids, E. H. Ten Winkel, and C. J. De Groot. Temporal and spatial patterns of water mites in Lake Maarsseveen I. Aquatic Ecology, 28(1):11–17, March 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334240.

Djemali:2009:HFB

[DTG09] Imed Djemali, Rachid Toujani, and Jean Guillard. Hydroacoustic fish biomass assessment in man-made lakes in Tunisia: horizontal beaming importance and diel effect. Aquatic Ecology, 43(4):1121–1131, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9215-6.

Dan:2020:OCV

[DTH20] Shigeki Dan, Arata Takasugi, and Katsuyuki Hamasaki. Ontogenic change in the vertical swimming of East Asian common octopus Octopus sinensis paralarvae under different water flow conditions. Aquatic Ecology, 54(3):795–812, September 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09777-7.

Dumont:1969:FCP

[Dum69] H. J. Dumont. Fysico-chemie en primaire productie van het Donk-meer in Belgie. (Dutch) [Physico-chemistry and primary production of Lake Donk in Belgium]. Aquatic Ecology, 3(3):62–63, November 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185854.

Dresscher:1974:BWA

[DV74] Th. G. N. Dresscher and H. Van Der Mark. A biological water analysis (abstract). Aquatic Ecology, 8(3):289, December 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257506.

Dorgelo:1992:HMI

[DV92] Jaap Dorgelo and Liesbeth Van der Kamp. Heavy metals in the IJsselmeer area (The Netherlands): Supply, distribution and concentrations in water, sediment and organisms. A review. Aquatic Ecology, 25(3):191–216, August 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270805.

Duursma:1978:CHW

[DVK78] E. K. Duursma, F. Vegter, and P. Kelderman. Concepts of hydrochemical water quality of the Delta waters. Aquatic Ecology, 12(3-4):215–225, December 1978. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02259184.

DeJonge:1981:ICW

[DVK81] V. N. De Jonge, J. Van Den Bergs, and J. E. C. A. Koekenbier. Influence of currents and waves on the whirling up of benthic diatoms living on intertidal flats. Aquatic Ecology, 15(3):197, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255182.

Dvorak:1969:ZLC

[Dvo69] Jan Dvořák. The zonation of life conditions and the macrofauna of littoral emergent vegetation in ponds (South Bohemia). Aquatic Ecology, 3(1):12–22, March 1969. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185843.

Dvorak:1970:SPC

[Dvô70] Jan Dvôrák. On some phytophilous chironomid larvae of the nature reservation “De Lindevallei” (Chironomidae, Diptera). Aquatic Ecology, 4(3):172–174, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185882.

DeWolf:1994:WLA

[DVZ94] H. De Wolf, M. A. M. Van Hoorn, and E. Zevenhuizen. Work and life of Albert van der Werff (1903–1991). Aquatic Ecology, 28(1):1–10, March 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334239.

DeRuyterVanSteveninck:1990:PDR

[DvZA90] E. D. De Ruyter Van Steveninck, B. van Zanten, and W. Admiraal. Phases in the development of riverine plankton: Examples from the rivers Rhine and Meuse. Aquatic Ecology, 24(1):47–55, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256748.

deWilde:1971:IVT

[dW71a] P. A. W. J. de Wilde. Invloed van de temperatuur op de groei van het Nonnetje, Macoma balthica (L.). (Dutch) [Influence of temperature on the growth of the bivalve, Macoma balthica (L.)]. Aquatic Ecology, 5(2):101–103, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185950.

deWolf:1971:EOV

[dW71b] P. de Wolf. Enkele opmerkingen over het voorkomen van metalen in mosselen langs de Nederlandse kust. (Dutch) [Some remarks about the occurrence of metals in mussels along the Dutch coast]. Aquatic Ecology, 5(4):179–180, December 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185961.

Darbyshire:1993:TMT

[DW93] E. J. Darbyshire and J. R. West. The turbidity maximum in the Tamar estuary. Aquatic Ecology, 27(2-4):121–133, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334775.

Durham:2014:UCR

[DW14] Bart W. Durham and Gene R. Wilde. Understanding complex reproductive ecology in fishes: the importance of individual and population-scale information. Aquatic Ecology, 48(1):91–106, March 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-014-9469-0.

Dahlgren:2011:IAH

[DWA11] Kristin Dahlgren, Ann-Kristin Eriksson Wiklund, and Agneta Andersson. The influence of autotrophy, heterotrophy and temperature on pelagic food web efficiency in a brackish water system. Aquatic Ecology, 45(3):307–323, September 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-011-9355-y.

Davies:2010:IUD

[DWB10] Peter J. Davies, Ian A. Wright, and Shelley Burgin. Impact of urban development on aquatic macroinvertebrates in south eastern Australia: degradation of in-stream habitats and comparison with non-urban streams. Aquatic Ecology, 44 (4):685–700, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9307-y.

DeHaan:1981:ASA

[DWV81] H. De Haan, D. J. Wijbenga, and M. J. W. Veldhuis. Aspects of the study on the availability of iron for phytoplankton in Tjeukemeer, the Netherlands. Aquatic Ecology, 15(3):

193–194, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255180.

deWit:1988:IBP

[dWvG88] Rutger de Wit and Hans van Gemerden. Interactions between phototrophic bacteria in sediment ecosystems. Aquatic Ecology, 22(2):135–145, December 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256817.

Depinto:1986:IPA

[DYS86] J. V. Depinto, T. C. Young, and D. K. Salisbury. Impact of phosphorus availability on modelling phytoplankton dynamics. Aquatic Ecology, 20(1-2):225–243, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291165.

Degermendzhy:2010:VSP

[DZG10] Andrey G. Degermendzhy, Egor S. Zadereev, and Ramesh D. Gulati. Vertical stratification of physical, chemical and biological components in two saline lakes Shira and Shunet (South Siberia, Russia). Aquatic Ecology, 44(3):619–632, September 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-010-9336-6.

Ding:2017:AST

[DZG17] Xiaohui Ding, Jianfeng Zou, and Zixuan Ge. Acclimation of Salix triandroides cuttings to incomplete submergence is reduced by low light. Aquatic Ecology, 51(2):321–330, June 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-017-9619-2.

Ding:2021:PAS

[DZL21] Mingming Ding, Ruixue Zhou, and Lei Li. Physiological adaptations of the submerged macrophyte Vallisneria spinulosa in response to water level fluctuations. Aquatic Ecology, 55(1):33–45, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09808-3.

Du:2013:GSM

[DZT13] Jingjing Du, Guiying Zhao, and Xingjun Tian. Growth stimulation of Microcystis aeruginosa by a bacterium from hypereutrophic water (Taihu Lake, China). Aquatic Ecology, 47(3): 303–313, September 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-013-9445-0.

Essien:2008:MBB

[EAB08] Joseph Peter Essien, Sylvester P. Antai, and Nsikak U. Benson. Microalgae biodiversity and biomass status in Qua Iboe Estuary mangrove swamp, Nigeria. Aquatic Ecology, 42(1):71–81, March 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9083-5.

Ellawala:2011:IFT

[EAK11] Champika Ellawala, Takashi Asaeda, and Kiyoshi Kawamura. Influence of flow turbulence on growth and indole acetic acid and H_{2}O_{2} metabolism of three aquatic macrophyte species. Aquatic Ecology, 45(3):417–426, September 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-011-9364-x.

Edegbene:2020:EDP

[EAO20] A. O. Edegbene, F. O. Arimoro, and O. N. Odume. Exploring the distribution patterns of macroinvertebrate signature traits and ecological preferences and their responses to urban and agricultural pollution in selected rivers in the Niger Delta ecoregion, Nigeria. Aquatic Ecology, 54(2):553– 573, June 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09759-9.

Etim:1996:SSD

[EBA96] Lawrence Etim, Thomas Brey, and Wolf Arntz. A seminal study of the dynamics of a mudskipper (Periophthalmus papilio) population in the Cross River, Nigeria. Aquatic Ecology, 30(1):41–48, July 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02092146.

Ellerby:2018:AER

[EBM18] David J. Ellerby, Caroline G. Berlin, and Clinton J. Moran. Assessing the ecological relevance of swimming performance traits: a case study of bluegill sunfish (Lepomis macrochirus). Aquatic Ecology, 52(4):311–322, December 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-018-9665-4.

Ebner:2009:SBT

[ECB09] Brendan Ebner, Rhian Clear, and Matthew Beitzel. Instream behaviour of threatened fishes and their food organisms based on remote video monitoring. Aquatic Ecology, 43(2):569–576, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9192-9.

Escobedo:2025:SCR

[ECGC25] Salem Kit T. Escobedo, Warren G. Caneos, Reynald V. Gimena, and Chelmarie Joy S. Clavano. Species composition, relative abundance and distribution of sea cucumber in the intertidal zone of East Cab-ilan Island, Province of Dinagat Islands, Philippines. Aquatic Ecology, 59(1):295–306, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10163-w.

Eisma:1994:TVS

[ECL94] D. Eisma, S. Chen, and A. Li. Tidal variations in suspended matter floc size in the Elbe river and Dollard estuaries. Aquatic Ecology, 28(3-4):267–274, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334194.

Ellien:2020:LEE

[ECR20] Céline Ellien, Romain Causse, and Karine Rousseau. Looking for environmental and endocrine factors inducing the transformation of Sicyopterus lagocephalus (Pallas 1770) (Teleostei: Gobiidae: Sicydiinae) freshwater prolarvae into marine larvae. Aquatic Ecology, 54(1):163–180, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09734-z.

Elliott:1995:SCE

[ED95] M. Elliott and F. Dewailly. The structure and components of European estuarine fish assemblages. Aquatic Ecology, 29 (3-4):397–417, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02084239.

Evangelista:2024:DID

[EDLC24] C. Evangelista, M. Danger, R. Lassus, and J. Cucherousset. Different impacts of diet composition on the stoichiometric traits of two freshwater species. Aquatic Ecology, 58(2):249– 261, June 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10063-5.

Elliott:1999:IED

[EFD99] M. Elliott, T. F. Fernandes, and V. N. De Jonge. The impact of European directives on estuarine and coastal science and management. Aquatic Ecology, 33(3):311–321, October 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009960706750.

Einarsson:2004:MDD

[EG04] Árni Einarsson and Arnthor Gardarsson. Moulting diving ducks and their food supply. Aquatic Ecology, 38(2):297– 307, June 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000032055.49860.6f.

Erikson:1991:ULP

[EHM91] R. Erikson, E. Hooker, and M. Meija. Underwater light penetration, phytoplankton biomass and photosynthetic activity in Lake Xolotlán (Managua). Aquatic Ecology, 25(2):137– 144, June 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291245.

Einarsson:2004:LMR

[Ein04] Árni Einarsson. Lake Myvatn and the River Laxá: An introduction. Aquatic Ecology, 38(2):111–114, June 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032091.58691.45.

Essink:1993:DLC

[EK93] Karel Essink and Hans L. Kleef. Distribution and life cycle of the North American Spionid polychaete Marenzelleria viridis (Verrill, 1873) in the Ems estuary. Aquatic Ecology, 27(2-4):237–246, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334787.

Essink:1998:CEN

[EK98] Karel Essink and Harm Keidel. Changes in estuarine nematode communities following a decrease of organic pollution. Aquatic Ecology, 32(3):195–202, October 1998. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009901207217.

Ejsmont-Karabin:2021:RLS

[EKK21] Jolanta Ejsmont-Karabin and Maciej Karpowicz. Rotifera in lake subhabitats. Aquatic Ecology, 55(4):1285–1296, December 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09818-1.

Egizi:2014:UMM

[EMF14] Andrea Egizi, Peter J. Morin, and Dina M. Fonseca. Unraveling microbe-mediated interactions between mosquito larvae in a laboratory microcosm. Aquatic Ecology, 48(2):179– 189, June 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9474-3.

Engkvist:2004:IBI

[EMN04] Roland Engkvist, Torleif Malm, and Jonas Nilsson. Interaction between isopod grazing and wave action: a structuring force in macroalgal communities in the southern Baltic Sea. Aquatic Ecology, 38(3):403–413, September 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000035162.07481.1f.

Engkvist:2000:DDG

[EMT00] Roland Engkvist, Torleif Malm, and Stefan Tobiasson. Density dependent grazing effects of the isopod Idotea baltica Pallas on Fucus vesiculosus L. in the Baltic Sea. Aquatic Ecology, 34(3):253–260, September 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009919526259.

Engel:1969:M

[Eng69] H. Engel. In memoriam. Aquatic Ecology, 3(1):2–3, March 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185835.

Engel:1974:HSD

[Eng74] H. Engel. A historical survey of Dr. Redeke’s work and its meaning (general introduction). Aquatic Ecology, 8(1-2):5– 14, July 1974. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02254901.

Engelen:1986:IHS

[Eng86] G. B. Engelen. Interaction of hydrological systems and eutrophication of the Loosdrecht Lakes. Aquatic Ecology, 20 (1-2):17–25, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02291147.

Einarsson:2004:LTC

[EÖ04] Arni Einarsson and Erla Björk Örnólfsdóttir. Long-term changes in benthic Cladocera populations in Lake Myvatn, Iceland. Aquatic Ecology, 38(2):253–262, June 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032060.29256.95.

Etherington:2023:EAA

[EPL^{+}23] B. S. Etherington, M. L. Piczak, L. LaRochelle, A. J. Gallagher, and S. J. Cooke. Effects of anthropogenic activities on scavenger communities in freshwater riparian zones of eastern Ontario, Canada. Aquatic Ecology, 57(1):115–125, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-022-09993-3.

Essink:1994:ENI

[ER94] Karel Essink and Karin Romeyn. Estuarine nematodes as indicators of organic pollution; an example from the Ems estuary (The Netherlands). Aquatic Ecology, 28(2):213–219, June 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02333992.

Elias:2017:ICS

[ERA17] C. L. Elias, R. J. M. Rocha, and S. F. P. Almeida. Influence of the colonizing substrate on diatom assemblages and implications for bioassessment: a mesocosm experiment. Aquatic Ecology, 51(1):145–158, March 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9605-0.

Espinar:2009:QHA

[ES09] José L. Espinar and Laura Serrano. A quantitative hydrogeomorphic approach to the classification of temporary wetlands in the Doñana National Park (SW Spain). Aquatic Ecology, 43(2):323–334, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9162-7.

Einarsson:2004:ELM

[ESG04] Árni Einarsson, Gerdur Stefánsdóttir, and Arnthor Gardarsson. The ecology of Lake Myvatn and the River Laxá: Variation in space and time. Aquatic Ecology, 38(2):317–348, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000032090.72702.a9.

Essink:1971:VOW

[Ess71] K. Essink. Vervuiling in de oostelijke Waddenzee onder invloed van de persleiding vanuit Hoogkerk. (Dutch) [Pollution in the eastern Wadden Sea under the influence of the pressure pipeline from Hoogkerk]. Aquatic Ecology, 5(1):17–22, March 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185884.

Essink:2003:REE

[Ess03] Karel Essink. Response of an estuarine ecosystem to reduced organic waste discharge. Aquatic Ecology, 37(1):65–76, January 2003. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1022139600926.

Essink:2006:BRF

[Ess06] Karel Essink. Book review: Folkert de Jong, Marine Eutrophication in Perspective. Aquatic Ecology, 40(4):581–582, December 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-006-9063-1.

Erhardt:2016:ENS

[ET16] John M. Erhardt and Kenneth F. Tiffan. Ecology of nonnative Siberian prawn (Palaemon modestus) in the lower Snake River, Washington, USA. Aquatic Ecology, 50(4):607–621, December 2016. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9581-4.

Freire:1999:WIS

[FA99] Paula Freire and César Andrade. Wind-induced sand transport in Tagus estuarine beaches — first results. Aquatic Ecology, 33(3):225–233, October 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009911012260.

Fernandez:2018:CCC

[FA18] Rocío Fernández and Javier Alcocer. Cyanobacteria consumption by cladocerans: a case study on facilitation.

Aquatic Ecology, 52(4):245–254, December 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9660-9.

Faria:2018:TNF

[FAB18] Fernando Azevedo Faria, Edélti Faria Albertoni, and Leandro Bugoni. Trophic niches and feeding relationships of shorebirds in southern Brazil. Aquatic Ecology, 52(4): 281–296, December 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-018-9663-6.

Fastner:2016:CCP

[FAC16] Jutta Fastner, Sally Abella, and Ingrid Chorus. Combating cyanobacterial proliferation by avoiding or treating inflows with high P load — experiences from eight case studies. Aquatic Ecology, 50(3):367–383, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9558-8.

Fathi:2001:PCN

[FAK01] A. A. Fathi, H. M. A. Abdelzaher, and M. M. Kraïem. Phytoplankton communities of North African wetland lakes: the CASSARINA Project. Aquatic Ecology, 35(3-4):303–318, October 2001. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1011988722774.

Fileto:2007:IBM

[FAL07] Claudia Fileto, Marlene S. Arcifa, and Norberto P. Lopes. Influence of biochemical, mineral and morphological features of natural food on tropical cladocerans. Aquatic Ecology, 41 (4):557–568, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9117-z.

Frenette:2003:SGD

[FAM03] Jean-Jacques Frenette, Michael T. Arts, and Jean Morin. Spectral gradients of downwelling light in a fluvial lake (Lake Saint-Pierre, St-Lawrence River). Aquatic Ecology, 37 (1):77–85, January 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1022133530244.

Fileto:2005:IPFa

[FAS05a] Cláudia Fileto, Marlene S. Arcifa, and Lúcia Helena S. Silva. Influence of phytoplankton fractions on growth and reproduction of tropical cladocerans. Aquatic Ecology, 38(4): 503–514, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-004-4087-x.

Fileto:2005:IPFb

[FAS05b] Cláudia Fileto, Marlene S. Arcifa, and Lúcia Helena S. Silva. Influence of phytoplankton fractions on growth and reproduction of tropical cladocerans. Aquatic Ecology, 38(4): 503–514, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-4087-5.

Franco:2025:CTV

[FASdPN^{+}25] Ana Clara Sampaio Franco, Valter M. Azevedo-Santos, Marluce Aparecida Mattos de Paula Nogueira, Tommaso Giarrizzo, Rachel Ann Hauser-Davis, Erick Cristofore Guimarães, Roger Henrique Dalcin, Marcelo Soeth, Matheus Oliveira Freitas, Áthila A. Bertoncini, Vinícius Abilhoa, Almir Manoel Cunico, Johnatas Adelir-Alves, Bianca Bentes, José Luís Costa Novaes, Mauricio Hostim-Silva, Jonas Rodrigues Leite, Vagner Leonardo Macêdo dos Santos, and Jean Ricardo Simões Vitule. Correction: Tilapia venturing into high-salinity environments: a cause for concern? Aquatic Ecology, 59(1):393–394, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10143-0. See [SASA^{+}24].

Fraisse:2013:MTR

[FBL13] Stéphane Fraisse, Myriam Bormans, and Yvan Lagadeuc. Morphofunctional traits reflect differences in phytoplankton community between rivers of contrasting flow regime. Aquatic Ecology, 47(3):315–327, September 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9446-z.

Frotte:2022:UDF

[FBM22] Lou Frotté, Alexandre Bec, and Dominique Monti. Upstream/downstream food quality differences in a Caribbean island river. Aquatic Ecology, 56(1):311–317, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09887-w.

Fransozo:2008:EAH

[FBNF08] Adilson Fransozo, Giovana Bertini, and Maria Lucia Negreiros-Fransozo. Ecological aspects of hermit crabs (Crustacea, Anomura, Paguroidea) off the northern coast of São Paulo State, Brazil. Aquatic Ecology, 42(3):437–448, September 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9103-5.

Flik:1987:UPP

[FBR87] B. J. G. Flik, M. Bos, and J. Ringelberg. Underestimation of primary production as indicated by measurements with size-fractionated phytoplankton in Lake Maarsseveen I (The Netherlands). Aquatic Ecology, 21(1):39–47, July 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255453.

Frau:2019:ART

[FBS19] Diego Frau, Yamila Battauz, and Rodrigo Sinistro. Assessing the relevance of top-down and bottom-up effects as phytoplankton structure drivers in a subtropical hypereutrophic shallow lake. Aquatic Ecology, 53(2):265–280, June 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09687-3.

Freire:2002:MSS

[FBV02] J. Freire, C. Bernárdez, and P. Verísimo. Management strategies for sustainable invertebrate fisheries in coastal ecosystems of Galicia (NW Spain). Aquatic Ecology, 36(1):41–50, January 2002. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1013350723445.

Forbes:1993:BES

[FC93] A. T. Forbes and D. P. Cyrus. Biological effects of salinity gradient reversals in a southeast African estuarine lake. Aquatic Ecology, 27(2-4):483–488, June 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334808.

Freitas:2009:PAD

[FCDA09] Vânia Freitas, Sérgia Costa-Dias, and Carlos Antunes. Patterns in abundance and distribution of juvenile flounder, Platichthys flesus, in Minho estuary (NW Iberian Peninsula). Aquatic Ecology, 43(4):??, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9237-8.

Ferreira:2016:AHB

[FCG16] Verónica Ferreira, José Castela, and Manuel A. S. Graça. Aquatic hyphomycetes, benthic macroinvertebrates and leaf litter decomposition in streams naturally differing in riparian vegetation. Aquatic Ecology, 50(4):711–725, December 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-016-9588-x.

Flores-Carpio:2025:PPL

[FCPCMB^{+}25] Alessandra Flores-Carpio, Elian F. Pariapaza-Coaquira, Elí Tomás Martinez-Barrios, César R. Luque-Fernández, and Pastor Coayla-Peñaloza. Prevalence and parasitic load in Emerita analoga “Muymuy” on the beaches of Mollendo (Arequipa), Southern Peru. Aquatic Ecology, 59(1):339–348, March 2025. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10166-7.

Fraser:2015:IIV

[FCS15] Clarissa M. L. Fraser, Ross A. Coleman, and Frank Seebacher. Inter-individual variation partially explains patterns of orientation on steeply sloped substrata in a keystone grazer, the limpet Cellana tramoserica. Aquatic Ecology, 49 (2):189–197, June 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9515-6.

Freeman:1994:CST

[FCW94] D. P. Freeman, L. E. Coates, and J. R. West. Cohesive sediment transport on an inter-tidal zone under combined wavetidal flow. Aquatic Ecology, 28(3-4):283–288, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334196.

Flemming:1994:BJM

[FD94] B. W. Flemming and M. T. Delafontaine. Biodeposition in a juvenile mussel bed of the east Frisian Wadden Sea (Southern North Sea). Aquatic Ecology, 28(3-4):289–297, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334197.

Ferreira:1997:TCC

[FDB97] J. G. Ferreira, P. Duarte, and B. Ball. Trophic capacity of Carlingford Lough for oyster culture — analysis by ecological modelling. Aquatic Ecology, 31(4):361–378, December 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009952729216.

Flik:1997:LTC

[FDG97] Ben Flik, Niels Daan, and Ramesh Gulati. Life-time contributions of Joop Ringelberg to new approaches in aquatic ecology, father of modern aquatic ecology in the Netherlands. Aquatic Ecology, 31(1):1–8, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009959808073.

Fennessy:1994:SSV

[FDH94] M. J. Fennessy, K. R. Dyer, and D. A. Huntley. Size and settling velocity distributions of flocs in the Tamar estuary during a tidal cycle. Aquatic Ecology, 28(3-4):275–282, September 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334195.

Falkenberg:2024:CPC

[FdLY^{+}24] Julia Martini Falkenberg, Vitória Maria Moreira de Lima, Fábio Hideki Yamada, Telton Pedro Anselmo Ramos, and Ana Carolina Figueiredo Lacerda. Changes in parasite communities of fishes from an intermittent river in the Brazilian semi-arid, after a major interbasin water transfer. Aquatic Ecology, 58(3):895–916, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10112-7.

Flower:2001:REC

[FDP01] R. J. Flower, S. Dobinson, and S. T. Patrick. Recent environmental change in North African wetland lakes: diatom and other stratigraphic evidence from nine sites in the CASSA-RINA Project. Aquatic Ecology, 35(3-4):369–388, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011984627760.

Fernandes:1995:MEE

[FED95] T. F. Fernandes, M. Elliott, and M. C. Da Silva. The management of European estuaries: a comparison of the features, controls and management framework of the Tagus (Portugal) and Humber (England). Aquatic Ecology, 29(3-4):459–468, December 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02084244.

Ferrarese:1992:CIR

[Fer92a] Uberto Ferrarese. Chironomids of Italian rice fields. Aquatic Ecology, 26(2-4):341–346, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255260.

Ferrington:1992:HSP

[Fer92b] Leonard C. Ferrington, Jr. Habitat and sediment preferences of Axarus festivus larvae. Aquatic Ecology, 26(2-4):347–354, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255261.

Fransz:1991:DEP

[FG91] H. G. Fransz and S. R. Gonzalez. Daily egg production of Temora longicornis (Copepoda, Calanoida) during winter and early spring in the Marsdiep (Southern North Sea). Aquatic Ecology, 25(1):61–64, October 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02259590.

FernandesPerroca:2024:NEG

[FGA^{+}24] Júlia Fernandes Perroca, Tommaso Giarrizzo, Ernesto Azzurro, Jorge Luiz Rodrigues-Filho, Carolina V. Silva, Marlene S. Arcifa, and Valter M. Azevedo-Santos. Negative effects of ghost nets on Mediterranean biodiversity. Aquatic Ecology, 58(1):131–137, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09985-3.

Figueiredo:2020:CEP

[FGB20] Bruno R. S. Figueiredo, Rafaela V. Granzotti, and Evanilde Benedito. Cascading effects of predation risk under high organic and inorganic turbidity: impacts on individuals and shoals of a mesopredator fish. Aquatic Ecology, 54(3):855– 868, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09782-w.

Ferreiro:2013:EMA

[FGF13] Nicolás Ferreiro, Adonis Giorgi, and Claudia Feijoó. Effects of macrophyte architecture and leaf shape complexity on structural parameters of the epiphytic algal community in a Pampean stream. Aquatic Ecology, 47(4):389–401, December 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9452-1.

Falkenberg:2019:GPF

[FGL19] Julia M. Falkenberg, Jéssica Emília S. A. Golzio, and Ana C. F. Lacerda. Gill parasites of fish and their relation to host and environmental factors in two estuaries in northeastern Brazil. Aquatic Ecology, 53(1):109–118, March 2019.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09676-6.

Freiry:2020:CSC

[FGM20] Raquel F. Freiry, Andressa Gouvea, and Leonardo Maltchik. Community structure and concordance patterns among zooplankton life stages in subtropical temporary ponds. Aquatic Ecology, 54(1):257–270, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
019-09740-1.

Flik:1981:EPP

[FK81] B. J. G. Flik and A. Keyser. Estimation of the primary production in Lake Maarsseveen I with an incubator technique. Aquatic Ecology, 15(1-2):41–50, October 1981. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260257.

Filippova:1992:HBR

[FKB92] M. A. Filippova, I. I. Kiknadze, and A. G. Blinov. Homology of Balbiani rings among chironomid species and localization of a new mobile element on the polytene chromosomes. Aquatic Ecology, 26(2-4):123–128, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255232.

Frouz:2009:ICG

[FLA09] Jan Frouz, Richard J. Lobinske, and Arshad Ali. Intraspecific competition of Glyptotendipes paripes (Diptera: Chironomidae) larvae under laboratory conditions. Aquatic Ecology, 43(2):487–500, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9178-7.

Flik:1970:OVD

[Fli70] B. J. G. Flik. De orientatie van Daphnia magna op een ruimtelijke lichtverdeling. (Dutch) [The orientation of Daphnia magna on a spatial light distribution]. Aquatic Ecology, 4 (3):146–150, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185877.

Flik:1985:ASP

[Fli85] Ben J. G. Flik. Analysis of the shape of P vs I curves of a natural phytoplankton assemblage during the year in the mesotrophic Lake Maarsseveen (i), The Netherlands. Aquatic Ecology, 19(2):117–122, December 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02270757.

Flower:2001:CSS

[Flo01] R. J. Flower. Change, stress, sustainability and aquatic ecosystem resilience in North African wetland lakes during the 20th century: an introduction to integrated biodiversity studies within the CASSARINA Project. Aquatic Ecology, 35(3-4):261–280, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1011978420737.

Fefilova:2008:MBC

[FLP08] Elena B. Fefilova, Olga A. Loskutova, and Sergey V. Pestov. Micro-benthic crustacean communities in tundra lakes of North-East European Russia. Aquatic Ecology, 42(3):449– 461, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9109-z.

Flory:1999:RCI

[FM99] Elizabeth A. Flory and Alexander M. Milner. The role of competition in invertebrate community development in a recently formed stream in Glacier Bay National Park, Alaska. Aquatic Ecology, 33(2):175–184, September 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009961000854.

Flanagan:2010:EWI

[FM10] Kyla M. Flanagan and Edward McCauley. Experimental warming increases CO_{2} saturation in a shallow prairie pond. Aquatic Ecology, 44(4):749–759, December 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9313-0.

Freret-Meurer:2024:BSS

[FMdCFdCV24] Natalie Villar Freret-Meurer, Tatiane do Carmo Fernández, and Amanda do Carmo Vaccani. Baseline study of the seahorse Hippocampus reidi Ginsburg, 1933 population in a tropical hypersaline lagoon. Aquatic Ecology, 58(1):117–123, March 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10039-5.

Flemming:1994:LRC

[FN94] B. W. Flemming and N. Nyandwi. Land reclamation as a cause of fine-grained sediment depletion in backbarrier tidal flats (Southern North Sea). Aquatic Ecology, 28(3-4): 299–307, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334198.

FeitosaLeal:2024:CSD

[FNS^{+}24a] Manuella Feitosa Leal, Miodeli Nogueira Júnior, Karina Ketelen Silva Dantas, Edson Lourenço Silva, Tamaris Gimenez Pinheiro, and Ana Carolina Figueiredo Lacerda. Correction: Structure and dynamics of mollusk communities from intermittent rivers in Brazilian semiarid region. Aquatic Ecology, 58(4):1191, December 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10142-1. See [?].

FeitosaLeal:2024:SDM

[FNS^{+}24b] Manuella Feitosa Leal, Miodeli Nogueira Júnior, Karina Ketelen Silva Dantas, Edson Lourenço Silva, Tamaris Gimenez Pinheiro, and Ana Carolina Figueiredo Lacerda. Structure and dynamics of mollusk communities from intermittent rivers in Brazilian semiarid region. Aquatic Ecology, 58(4):1175–1189, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-

024-10133-2. See correction [FNS^{+}24a].

Frid:1994:FRP

[FNW94] C. L. J. Frid, L. C. Newton, and J. A. Williams. The feeding rates of Pleurobrachia (Ctenophora) and Sagitta (Chaetognatha), with notes on the potential seasonal role of planktonic predators in the dynamics of North Sea zooplankton communities. Aquatic Ecology, 28(2):181–191, June 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02333989.

Fuhrman:1998:MMD

[FO98] Jed A. Fuhrman and Cleber C. Ouverney. Marine microbial diversity studied via 16S rRNA sequences: cloning results from coastal waters and counting of native archaea with fluorescent single cell probes. Aquatic Ecology, 32(1):3–15, March 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009974817127.

Fonds:1971:OOZ

[Fon71] A. W. Fonds. Het onderzoek op zware metalen in het oppervlaktewater. (Dutch) [Research on heavy metals in surface water]. Aquatic Ecology, 5(4):178, December 1971. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185960.

Fernandes:2009:MCB

[FR09] Veronica Fernandes and N. Ramaiah. Mesozooplankton community in the Bay of Bengal (India): spatial variability during the summer monsoon. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9209-4.

Fransz:1980:ZRN

[Fra80] H. G. Fransz. Zooplankton research in the North Sea. Aquatic Ecology, 14(3):224–225, December 1980. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260131.

France:2000:CAL

[Fra00] R. L. France. Comparing δ^{13}C among littoral foodwebs using lake DOC. Aquatic Ecology, 34(4):445–448, December 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011418926126.

Fernandes:2008:RBA

[FRP08] Veronica Fernandes, Veera Rodrigues, and Jane T. Paul. Relevance of bacterioplankton abundance and production in the oligotrophic equatorial Indian Ocean. Aquatic Ecology, 42(4): 511–519, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9142-y.

Fernandes:2006:NDE

[FSC06] Sandra Fernandes, Paula Sobral, and Maria Helena Costa. Nereis diversicolor effect on the stability of cohesive intertidal sediments. Aquatic Ecology, 40(4):567–579, December 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-8000-z.

Fabrin:2020:CUL

[FSdO20] Thomaz Mansini Carrenho Fabrin, Bruno Henrique Mioto Stabile, and Alessandra Valéria de Oliveira. Cyanobacteria in an urban lake: hidden diversity revealed by metabarcoding. Aquatic Ecology, 54(2):671–675, June 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09763-z.

Feng:2023:SDF

[FSL^{+}23] Yixuan Feng, Wei Song, Dongming Lin, André E. Punt, and Xinjun Chen. Spatial difference in feeding habits but similar foraging strategy for energy acquisition in jumbo squid. Aquatic Ecology, 57(3):653–666, September 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10035-9.

Feio:2021:HEZ

[FSN21] Maria João Feio, Sónia R. Q. Serra, and João M. Neto. From headwaters into the estuarine zone: changes in processes and invertebrate communities in response to abiotic conditions. Aquatic Ecology, 55(1):149–168, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09820-7.

Fletcher:2022:PBC

[FTH22] Nichola Fletcher, John A. Terschak, and Jorg D. Hardege. A pheromone bouquet controls the reproductive behaviour of the male shore crab, Carcinus maenas. Aquatic Ecology, 56(2):419–427, June 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09930-w.

Franssen:2008:SPE

[FTP08] Courtney M. Franssen, Michael Tobler, and Martin Plath. Sperm production in an extremophile fish, the cave molly (Poecilia mexicana, Poeciliidae, Teleostei). Aquatic Ecology, 42(4):685–692, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
007-9128-9.

Fonds:1974:ITS

[FV74] M. Fonds and G. Van Buurt. The influence of temperature and salinity on development and survival of goby eggs (Pisces, Gobiidae). Aquatic Ecology, 8(1-2):110–116, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02254911.

Falcao:1995:TFA

[FV95a] Manuela Falcão and Carlos Vale. Tidal flushing of ammonium from intertidal sediments of Ria Formosa, Portugal. Aquatic Ecology, 29(3-4):239–244, December 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084221.

Ferreira:1995:IRD

[FV95b] A. M. Ferreira and C. Vale. The importance of runoff to DDT and PCB inputs to the Sado estuary and Ria Formosa. Aquatic Ecology, 29(3-4):211–216, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084218.

Faasse:2003:NAA

[FvM03] Marco Faasse and Godfried van Moorsel. The North-American amphipods, Melita nitida Smith, 1873 and Incisocalliope aestuarius (Watling and Maurer, 1973) (Crustacea: Amphipoda: Gammaridae), introduced to the Western Scheldt estuary (The Netherlands). Aquatic Ecology, 37 (1):13–22, January 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1022120729031.

Fellerhoff:2003:SCN

[FVW03] Claudia Fellerhoff, Maren Voss, and Karl Matthias Wantzen. Stable carbon and nitrogen isotope signatures of decomposing tropical macrophytes. Aquatic Ecology, 37(4):361–375, October 2003. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000007049.25535.12.

Fureder:2005:LSP

[FWB05] Leopold Füreder, Manfred Wallinger, and Rainer Burger. Longitudinal and seasonal pattern of insect emergence in alpine streams. Aquatic Ecology, 39(1):67–78, March 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-2969-6.

Fontanilles:2023:CDR

[FWR^{+}23a] Adeline Fontanilles, Alexander Wezel, Soraya Rouifed, Mathieu Guerin, Marie Vanacker, Florent Arthaud, and Joël Robin. Correction: Disturbance and resilience of aquatic plant communities in fish ponds after temporary dry periods. Aquatic Ecology, 57(3):611, September 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10047-5. See [FWR^{+}23b].

Fontanilles:2023:DRA

[FWR^{+}23b] Adeline Fontanilles, Alexander Wezel, Soraya Rouifed, Mathieu Guerin, Marie Vanacker, Florent Arthaud, and Joël Robin. Disturbance and resilience of aquatic plant communities in fish ponds after temporary dry periods. Aquatic Ecology, 57(3):597–609, September 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s10452-023-10032-y. See correction [FWR^{+}23a].

Feng:2023:FAE

[FWS^{+}23] Yihan Feng, Yaoguang Wei, Shuo Sun, Jincun Liu, Dong An, and Jia Wang. Fish abundance estimation from multi-beam sonar by improved MCNN. Aquatic Ecology, 57(4):895–911, December 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10007-z.

Gabler:2010:FSR

[GA10] Heidi-Marie Gabler and Per-Arne Amundsen. Feeding strategies, resource utilisation and potential mechanisms for competitive coexistence of Atlantic salmon and alpine bullhead in a sub-Arctic river. Aquatic Ecology, 44(2):325–336, June 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9243-x.

Giacobbe:2024:PCS

[GAA24] Salvatore Giacobbe, Cosentino Andrea, and Rosso Antonietta. Primary colonization and small-scale dynamics of nonindigenous benthic species: a case study. Aquatic Ecology, 58(4):1237–1254, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
024-10138-x.

Guermazi:2024:ETP

[GAAT^{+}24] Wassim Guermazi, Khaled Athmouni, Neila Annabi-Trabelsi, Jannet Elloumi, Habib Ayadi, and Vincent Leignel. Energy transfers in a predator–prey context involving D. salina (microalga), F. salina (ciliate) and A. salina (crustacean), living in salterns of Sfax (Tunisia). Aquatic Ecology, 58(2):429–449, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10082-2.

Guisande:2000:EZC

[GANA00] C. Guisande, C. Andrade, and M. Núñez-Avellaneda. Effects of zooplankton and conductivity on tropical Utricularia foliosa investment in carnivory. Aquatic Ecology, 34(2):137– 142, June 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009966231287.

Gardeniers:1971:MNI

[Gar71] J. J. P. Gardeniers. De makrofauna (niet-insekten) in een verlandingszone van de Venematen (N.W.-Overijssel). (Dutch) [The macrofauna (non-insects) in a relocation zone of the Venemates (N.W.-Overijssel)]. Aquatic Ecology, 5(3):146– 160, September 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185892.

Gard:2005:OMS

[Gar05] Mark F. Gard. Ontogenetic microhabitat shifts in Sacramento pikeminnow, Ptychocheilus grandis: reducing intraspecific predation. Aquatic Ecology, 39(2):229–235, June 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-004-1912-1.

Geelen:1969:YPU

[GB69] J. F. M. Geelen and A. M. J. Bouwhuis. Yergelijkend planktononderzoek in “uiversnest” en “meeuwenven”. (Dutch) [Comparative plankton research in “gull’s nest” and “gull pond”]. Aquatic Ecology, 3(1):23–31, March 1969. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185844.

Gordon:1982:PCT

[GB82] Donald C. Gordon, Jr. and Job W. Baretta. A preliminary comparison of two turbid coastal ecosystems: The Dollard (Netherlands–FRG) and the Cumberland Basin (Canada).

Aquatic Ecology, 16(2-3):255–267, December 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255379.

Geraldes:2007:ZAT

[GB07] Ana Maria Geraldes and Maria-José Boavida. Zooplankton assemblages in two reservoirs: one subjected to accentuated water level fluctuations, the other with more stable water levels. Aquatic Ecology, 41(2):273–284, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9057-z.

Green:2020:DRB

[GBA20] Nicky Green, J. Robert Britton, and Demetra Andreou. Dominance, reproductive behaviours and female mate choice in sterilised versus non-sterilised invasive male crayfish. Aquatic Ecology, 54(3):813–822, September 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09779-5.

Giaccardi:2022:AMS

[GBJM22] Laura I. Giaccardi, Misael A. Badenas, Leandro R. Jones, and Julieta M. Manrique. Abundant microbes of surface sea waters of the uncharted Engaño Bay at the Atlantic Patagonian Coast: relevance of bacteria-sized photosynthetic eukaryotes. Aquatic Ecology, 56(4):1217–1230, December 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09962-w.

Gons:1975:ELD

[GBM75] Herman J. Gons, Daan Barug, and Luuc R. Mur. Effects of a light-dark cycle on the growth of Scenedesmus protuberans fritsch in light-limited continuous cultures. Aquatic Ecology, 9(2):71–80, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257639.

Genova:2010:ODM

[GBM10] S. N. Genova, V. M. Belolipetskii, and W. M. Mooij. A onedimensional model of vertical stratification of Lake Shira focussed on winter conditions and ice cover. Aquatic Ecology, 44(3):571–584, September 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
010-9327-7.

Gazzola:2023:EEN

[GBPR23] Andrea Gazzola, Alessandro Balestrieri, and Daniele Pellitteri-Rosa. Embryonic exposure to native and alien predator cues tunes tadpole defensive behaviour. Aquatic Ecology, 57(2):421–431, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
023-10019-9.

Green:2019:MCV

[GBT19] Matthew W. Green, Peter W. Blum IV, and Shea R. Tuberty. Mesohabitat current velocity effects on Didymosphenia geminata and macroinvertebrates in a SE USA hypolimnetic tailwater. Aquatic Ecology, 53(4):607–628, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09712-5.

Gil:2021:DRP

[GBZ21] Damian G. Gil, Alicia L. Boraso, and Héctor E. Zaixso. Depth-related plasticity in the diet composition of Pseudechinus magellanicus (Echinoidea, Temnopleuridae) in nearshore environments off central Patagonia, Argentina. Aquatic Ecology, 55(2):589–606, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09847-4.

Guest:2005:FSMa

[GC05a] Michaela A. Guest and Rod M. Connolly. Fine-scale movement and assimilation of carbon in saltmarsh and mangrove habitat by resident animals. Aquatic Ecology, 38(4):599–609, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-0442-9.

Guest:2005:FSMb

[GC05b] Michaela A. Guest and Rod M. Connolly. Fine-scale movement and assimilation of carbon in saltmarsh and mangrove habitat by resident animals. Aquatic Ecology, 38(4):599–609, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-0442-1.

Godwin:2008:STV

[GC08] Casey M. Godwin and Hunter J. Carrick. Spatio-temporal variation of periphyton biomass and accumulation in a temperate spring-fed stream. Aquatic Ecology, 42(4):583–595, December 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9133-z.

Guo:2023:MTS

[GC23] J. Guo and M. Cherif. More than stoichiometry: the molecular composition of inorganic and organic substrates controls ammonium regeneration by bacteria. Aquatic Ecology, 57(2):543–555, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10028-8.

Guariento:2011:HLN

[GCE11] Rafael D. Guariento, Luciana S. Carneiro, and Francisco A. Esteves. How light and nutrients affect the relationship between autotrophic and heterotrophic biomass in a tropical black water periphyton community. Aquatic Ecology, 45(4): 561–569, November 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9377-5.

Graneli:1999:RCP

[GCL99] Edna Granéli, Per Carlsson, and Catherine Legrand. The role of C, N and P in dissolved and particulate organic matter as a nutrient source for phytoplankton growth, including toxic species. Aquatic Ecology, 33(1):17–27, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009925515059.

Garcia:2023:ADN

[GCN^{+}23] Guglielmo Fernandez Garcia, Thomas Corpetti, Marie Nevoux, Laurent Beaulaton, and François Martignac. AcousticIA, a deep neural network for multi-species fish detection using multiple models of acoustic cameras. Aquatic Ecology, 57(4):881–893, December 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
023-10004-2.

Gadel:1993:BSS

[GCS93] F. Gadel, B. Charriere, and L. Serve. Behaviour of suspended and sedimentary organic matter in the deltaic areas of the Gulf of Lions (Mediterranean Sea). Aquatic Ecology, 27(2-4):437–447, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334805.

Gerwing:2023:CPI

[GCT^{+}23] Travis G. Gerwing, Lily Campbell, Hanna M. Thomson, Sarah E. Dudas, and Francis Juanes. Community and population impacts of an introduced meiofauna (Nippoleucon hinumensis) upon an intertidal infaunal community and its closest endemic analog (Cumella vulgaris). Aquatic Ecology, 57(1):85–97, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09996-0.

Geelen:1972:MVE

[GD72] J. F. M. Geelen and C. Davids. Microfauna van enkele wateren in de Ooypolder. (Dutch) [Microfauna of some waters in the Ooypolder]. Aquatic Ecology, 6(2):89–99, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02304098.

Gulati:1991:STA

[GD91] Ramesh D. Gulati and Anja Doornekamp. The spring-time abundance and feeding of Eurytemora affinis (Poppe) in Volkerak-Zoommeer, a newly-created freshwater lake system in the Rhine delta (The Netherlands). Aquatic Ecology, 25 (1):51–60, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259589.

Gulati:2002:P

[GD02] Ramesh D. Gulati and Andrei G. Degermendzhy. Preface. Aquatic Ecology, 36(2):105–106, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1016394520003.

Glass:2009:ECP

[GD09] Nancy H. Glass and Philip C. Darby. The effect of calcium and pH on Florida apple snail, Pomacea paludosa (Gastropoda: Ampullariidae), shell growth and crush weight. Aquatic Ecology, 43(4):??, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9226-3.

Goetz:2019:IAF

[GDB19] Milena Nunes Bernardes Goetz, Ênio Wocyli Dantas, and Iva Carneiro Leão Barros. Influence of abiotic factors on the composition and abundance of aquatic ferns occurring in the state of Paraíba, Brazil. Aquatic Ecology, 53(4):557–567, December 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09708-1.

Goethals:2007:AAN

[GDD07] Peter L. M. Goethals, Andy P. Dedecker, and Niels De Pauw. Applications of artificial neural networks predicting macroinvertebrates in freshwaters. Aquatic Ecology, 41(3):491–508, September 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9093-3.

Golterman:1969:EVI

[GdK69] H. L. Golterman and W. de Kloet. De eutrofiering van het Ijsselmeer. (Dutch) [The eutrophication of Lake IJssel]. Aquatic Ecology, 3(1):7–9, March 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185840.

Greisberger:2008:CPS

[GDT08] Sonja Greisberger, Martin T. Dokulil, and Katrin Teubner. A comparison of phytoplankton size-fractions in Mondsee, an alpine lake in Austria: distribution, pigment composition and primary production rates. Aquatic Ecology, 42(3): 379–389, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9095-1.

Ghobashy:1980:FLT

[GE80a] A. F. A. Ghobashy and M. M. El Komy. Fouling in Lake Timsah (Egypt). Aquatic Ecology, 14(3):169–178, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260118.

Ghobashy:1980:FSR

[GE80b] A. F. A. Ghobashy and M. M. El Komy. Fouling in the southern region of the Suez Canal. Aquatic Ecology, 14(3): 179–185, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260119.

Gardarsson:2004:RLD

[GE04] Arnthor Gardarsson and Arni Einarsson. Resource limitation of diving ducks at Myvatn: Food limits production. Aquatic Ecology, 38(2):285–295, June 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032058.83651.2c.

Geelen:1969:ZVV

[Gee69] J. M. F. Geelen. Zoöplankton van de Venematen. (Dutch) [Zooplankton of the Venematen]. Aquatic Ecology, 3(3): 67–74, November 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185857.

Geelen:1975:OLR

[Gee75] J. F. M. Geelen. Orconectes limosus (Raf.) and Astacus astacus L. (Crustacea, Decapoda) in The Netherlands. Aquatic Ecology, 9(3):109–113, December 1975. CODEN AQECF9.

ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02263328.

Geelen:1976:MMD

[Gee76] Hannie Geelen. In memoriam Mej. Drs. A. M. J. Bouwhuis. Aquatic Ecology, 10(1):78, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02308505.

Gieskes:1991:DDC

[GEG91] W. W. C. Gieskes, M. M. Engelkes, and G. W. Grakaay. Degradation of diatom chlorophyll to colourless, nonfluorescing compounds during copepoo grazing. Aquatic Ecology, 25(1):65–72, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02259591.

Gons:1997:OTV

[GEK97] Herman J. Gons, Jeannine Ebert, and Jacco Kromkamp. Optical teledetection of the vertical attenuation coefficient for downward quantum irradiance of photosynthetically available radiation in turbid inland waters. Aquatic Ecology, 31(3):299–311, September 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009902627476.

Geleide:1970:VVS

[Gel70] Ten Geleide. Verslagen van een SYMPOSIUM, georganiseerd door de HYDROBIOLOGISCHE VERENIGING en gehouden op 2 mei 1969 te Amsterdam. (Dutch) [Reports of a symposium, organized by the Hydrobiological Society and held on 2 May 1969 in Amsterdam]. Aquatic Ecology, 4(1):1– 2, March 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185859.

Gardarsson:2004:PFC

[GEÓ04a] Arnthor Gardarsson, Árni Einarsson, and Jón S. Ólafsson. Population fluctuations of chironomid and simuliid Diptera at Myvatn in 1977–1996. Aquatic Ecology, 38(2):209–217, June 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000032051.14118.e1.

Gislason:2004:CCI

[GEÓ04b] Sigurdur Reynir Gíslason, Eydís Salome Eiríksdóttir, and Jón Sigurdur Ólafsson. Chemical composition of interstitial water and diffusive fluxes within the diatomaceous sediment in Lake Myvatn, Iceland. Aquatic Ecology, 38(2):163– 175, June 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000032063.15765.e4.

Geursen:1984:ERF

[Geu84] J. Geursen. Experimental research on the feeding of bream (Abramis brama L.) and white bream (Blicca bjoerkna L.). Aquatic Ecology, 18(1):69, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256677.

Galassi:2017:GBC

[GFF17] Diana M. P. Galassi, Barbara Fiasca, and Simone Fattorini. Groundwater biodiversity in a chemoautotrophic cave ecosystem: how geochemistry regulates microcrustacean community structure. Aquatic Ecology, 51(1):75–90, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-016-9599-7.

Gama-Flores:2004:ACT

[GFSN04] José Luis Gama-Flores, S. S. S. Sarma, and S. Nandini. Acute and chronic toxicity of the pesticide methyl parathion to the rotifer Brachionus angularis (Rotifera) at different algal (Chlorella vulgaris) food densities. Aquatic Ecology, 38(1):27–36, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000020986.92471.32.

Gama-Flores:2007:ETD

[GFSN07] José Luis Gama-Flores, S. S. S. Sarma, and S. Nandini. Exposure time-dependent cadmium toxicity to Moina macrocopa (Cladocera): a life table demographic study. Aquatic Ecology, 41(4):639–648, December 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9121-3.

Greufe:2024:PGS

[GFW24] Casey Greufe, Allyse Ferrara, and Justine Whitaker. Population genetic structure of invasive apple snails Pomacea maculata in Louisiana. Aquatic Ecology, 58(2):487–500, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10085-7.

Gatts:2015:IAP

[GFZ15] Pedro Gatts, Marcos Franco, and Ilana Zalmon. Impact of artificial patchy reef design on the ichthyofauna community of seasonally influenced shores at southeastern Brazil. Aquatic Ecology, 49(3):343–355, September 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9530-7.

Gabriels:2007:AMC

[GGD07] Wim Gabriels, Peter L. M. Goethals, and Niels De Pauw. Analysis of macrobenthic communities in Flanders, Belgium, using a stepwise input variable selection procedure with artificial neural networks. Aquatic Ecology, 41(3):427–441, September 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9081-7.

Geddes:2014:EIT

[GGT14] Pamela Geddes, Tanya Grancharova, and Nancy C. Tuchman. Effects of invasive Typha × glauca on wetland nutrient pools, denitrification, and bacterial communities are influenced by time since invasion. Aquatic Ecology, 48(3):247–258, September 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9480-5.

Gons:1986:ELL

[GGV86] Herman J. Gons, Ramesh D. Gulati, and Louis Van Liere. The eutrophic Loosdrecht Lakes: Current ecological research and restoration perspectives. Aquatic Ecology, 20(1-2):61–75, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291151.

Gustafsson:2004:DTA

[GH04] Susanne Gustafsson and Lars-Anders Hansson. Development of tolerance against toxic cyanobacteria in Daphnia. Aquatic Ecology, 38(1):37–44, March 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000020985.47348.5e.

George:2006:IYY

[GH06] D. G. George and D. P. Hewitt. The impact of year-to-year changes in the weather on the dynamics of Daphnia in a thermally stratified lake. Aquatic Ecology, 40(1):33– 47, March 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-9013-3.

Gibson:1986:PRP

[Gib86] C. E. Gibson. Preliminary results on phosphorus reduction in Lough Neagh — assessing the effect against a background of change. Aquatic Ecology, 20(1-2):173–182, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291161.

Gieskes:1971:MCV

[Gie71] W. W. Gieskes. Mogelijk causaal overband tussen de seizoenssuccessie van twee podonspecies (Cladocera) en die van voor hen geschikt voedsel in de Noordzee. (Dutch) [Possible causal relationship between the seasonal succession of two podon species (Cladocera) and that of food suitable for them in the North Sea]. Aquatic Ecology, 5(2):100–101, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185949.

Giesen:1983:EMO

[Gie83] Th. G. Giesen. Ecological and morphological observations on Hydrobius niger (Zschach, 1788) and Hydrobius fuscipes (Linnaeus, 1758) (Coleoptera, Hydrophilidae). Aquatic Ecology, 17(2):119–127, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02280821.

Gearty:2021:GRR

[GIK21] Austin J. Gearty, Toni R. Ignoffo, and Wim J. Kimmerer. Growth and reproductive rates of the dominant copepod Pseudodiaptomus forbesi in response to environmental factors and habitat type in the northern San Francisco Estuary. Aquatic Ecology, 55(3):825–848, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09863-4.

Gilbert:2012:PID

[Gil12] John J. Gilbert. Predator-induced defense in rotifers: developmental lags for morph transformations, and effect on population growth. Aquatic Ecology, 46(4):475–486, December 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-012-9416-x.

Gilbert:2020:ELT

[Gil20] John J. Gilbert. Effect of low temperatures on the reproduction of summer and perennial rotifers from temperate regions: ecological and life cycle implications. Aquatic Ecology, 54(3):711–719, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
020-09769-7.

Ginjaar:1980:OA

[Gin80] L. Ginjaar. Opening address. Aquatic Ecology, 14(1-2):3–4, August 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260266.

Godlewska:2006:AEF

[GJ06] Małgorzata Godlewska and Marek Jelonek. Acoustical estimates of fish and zooplankton distribution in the Piaseczno reservoir, Southern Poland. Aquatic Ecology, 40(2):211–219, June 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-3525-0.

Gemelli:2020:SPA

[GJW20] Federica Gemelli, Craig R. Johnson, and Jeffrey T. Wright. Spatial patterns of abundance and shell morphology of two gastropod species associated with different morphologies of an intertidal seaweed. Aquatic Ecology, 54(2):653–670, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09766-w.

Griffis-Kyle:2007:SEN

[GK07] Kerry L. Griffis-Kyle. Sublethal effects of nitrite on eastern tiger salamander (Ambystoma tigrinum tigrinum) and wood frog (Rana sylvatica) embryos and larvae: Implications for field populations. Aquatic Ecology, 41(1):119–127, March 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-006-9047-1.

G:1992:KTC

[GKK92] Istomina A. G., I. I. Kiknadze, and I. E. Kerkis. The karyotype of three Cryptochironomus species of West Siberia, USSR. Aquatic Ecology, 26(2-4):139–144, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255234.

Grabemann:1995:AVT

[GKK95] I. Grabemann, J. Kappenberg, and G. Krause. Aperiodic variations of the turbidity maxima of two German coastal plain estuaries. Aquatic Ecology, 29(3-4):217–227, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02084219.

Golawski:2017:WDR

[GKM17] Artur Golawski, Zbigniew Kasprzykowski, and Emilia Mroz. Wind differentiates reproduction in the non-expansive black tern Chlidonias niger and the expansive white-winged tern Chlidonias leucopterus. Aquatic Ecology, 51(2):235–245, June 2017. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9613-0.

Gkenas:2022:WAF

[GKM22] Christos Gkenas, Alexa Kodde, and Maria Filomena Magalhães. Warming affects the feeding success of invader and native fish in Iberian streams. Aquatic Ecology, 56 (1):319–324, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09888-9.

Geelen:1988:HCH

[GL88] J. F. M. Geelen and P. Leentvaar. From Hydrobiological club to Hydrobiological society. Aquatic Ecology, 22(1):67– 68, August 1988. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256784.

Gendron:1992:DEP

[GL92] J. M. Gendron and H. Laville. Diel emergence patterns of drifting chironomid (Diptera) pupal exuviae in the Aude River (Eastern Pyrénées, France). Aquatic Ecology, 26(2-4):273–279, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255252.

Gomez:2001:PDI

[GL01] Nora Gómez and Magdalena Licursi. The Pampean Diatom Index (IDP) for assessment of rivers and streams in Argentina. Aquatic Ecology, 35(2):173–181, June 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011415209445.

Gerard:2019:DFG

[GL19] Claudia Gérard and Emilie Lance. Decline of freshwater gastropods exposed to recurrent interacting stressors implying cyanobacterial proliferations and droughts. Aquatic Ecology, 53(1):79–96, March 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09674-8.

Gomez-Lemos:2022:ISC

[GLGU22] Luis Alonso Gómez-Lemos and Rocío García-Urueña. Induction of Staghorn coral settlement and early post-settlement survival in laboratory conditions. Aquatic Ecology, 56(3): 685–696, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09948-8.

Grant:1993:TMS

[GM93] Alastair Grant and Richard Middleton. Trace metals in sediments from the Humber Estuary: a statistical analysis of spatial uniformity. Aquatic Ecology, 27(2-4):111–120, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334774.

Gutierrez:2016:IMI

[GM16] María Florencia Gutierrez and Gisela Mayora. Influence of macrophyte integrity on zooplankton habitat preference, emphasizing the released phenolic compounds and chromophoric dissolved organic matter. Aquatic Ecology, 50 (1):137–151, March 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9561-0.

Gulati:2010:PSL

[GMD10] Ramesh D. Gulati, Wolf M. Mooij, and Andrey G. Degermendzhy. Preface to the Siberian lakes special issue. Aquatic Ecology, 44(3):481–483, September 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9338-4.

Geerts:2019:PND

[GMD19] Aurora N. Geerts, Kelle Moreau, and Luc De Meester. The power of numbers: dynamics of hatching and dormant egg production in two populations of the water flea Daphnia magna. Aquatic Ecology, 53(3):393–406, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09697-1.

Gayosso-Morales:2024:PRE

[GMVCLGGP24] Manuel Aaron Gayosso-Morales, Alejandro Valdez-Calderón, Isaac Lucas-Gómez, and Brenda Karen González-Pérez. Population responses of Daphnia laevis to endocrine disruptors: a molecular docking by binding active site to arginine kinase. Aquatic Ecology, 58(2):227–238, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10060-8.

Gathmann:2009:SPI

[GMW09] F. Oliver Gathmann, Lisa L. Manne, and D. Dudley Williams. Spatial patterns in insect community composition in coldwater springs. Aquatic Ecology, 43(2):501–512, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9166-y.

Gimenez:2024:SCE

[GOB24] Juliana Giménez, Mariel Ojeda, and Claudia Clementina Boy. Seasonal changes and energy allocation in the edible snail Zidona dufresnei (Caenogastropoda, Volutidae). Aquatic Ecology, 58(4):1097–1106, December 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10126-1.

Goethals:2007:SIE

[Goe07] Peter L. M. Goethals. Special issue ‘ecological informatics applications in water management’. Aquatic Ecology, 41(3): 371–372, September 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9100-8.

Golterman:1971:CPP

[Gol71] H. L. Golterman. Chemistry and primary production. Aquatic Ecology, 5(2):71–77, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185942.

Gons:1979:PLV

[Gon79] Herman J. Gons. Periphyton in Lake Vechten, with emphasis on biomass and production of epiphytic algae. Aquatic Ecology, 13(2-3):116, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02284744.

Goodman:1979:ASI

[Goo79] Keith S. Goodman. An apparatus for sampling interstitial water throughout tidal cycles. Aquatic Ecology, 13(1):30– 33, April 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260306.

Goser:1998:BRC

[Gos98a] Brigitte Goser. Book review: Cladocera: The Biology of model Organisms. Edited by A. Brancelj, L. DeMeester and P. Spaak. Aquatic Ecology, 32(4):369–370, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009928004970.

Goser:1998:BRH

[Gos98b] Brigitte Goser. Book review: Handbuch Angewandte Limnologie. Edited by C. Steinberg, W. Calmano, H. Klapper and R.-D. Wilken. Aquatic Ecology, 32(3):256–257, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009988117779.

Goser:1999:BRA

[Gos99] Brigitte Goser. Book review: Acidic Mining Lakes — Acid Mine Drainage, Limnology and Reclamation. Edited by W. Geller, H. Klapper, W. Salomons. Aquatic Ecology, 33(2): 213–214, September 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009914529124.

Groenendijk:1998:SDL

[GPA98] Dick Groenendijk, Jaap F. Postma, and Wim Admiraal. Seasonal dynamics and larval drift of Chironomus riparius (Diptera) in a metal contaminated lowland river. Aquatic Ecology, 32(4):341–351, December 1998. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009951709797.

Gasca-Pineda:2023:CSD

[GPGSMM^{+}23a] Jaime Gasca-Pineda, Clara E. Galindo-Sánchez, Miguel A. Martinez-Mercado, Sylvia P. A. Jiménez-Rosenberg, Clara María Hereu, Yasuhide Nakamura, Sharon Z. Herzka, Jesus C. Compaire, Ricardo Gomez-Reyes, Javier Robles-Flores, Anaid Saavedra-Flores, and Maria Clara Arteaga. Community structure and diversity of five groups of zooplankton in the Perdido region of the Gulf of Mexico using DNA metabarcoding. Aquatic Ecology, 57(1):149–164, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-

5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-022-10002-w. See correction

[GPGSMM^{+}23b].

GascauPineda:2023:CCS

[GPGSMM^{+}23b] Jaime Gasca-Pineda, Clara E. Galindo-Sánchez, Miguel A. Martinez-Mercado, Sylvia P. A. Jiménez-Rosenberg, Clara María Hereu, Yasuhide Nakamura, Sharon Z. Herzka, Jesus C. Compaire, Ricardo Gomez-Reyes, Javier Robles-Flores, Anaid Saavedra-Flores, and Maria Clara Arteaga. Correction: Community structure and diversity of five groups of zooplankton in the Perdido region of the Gulf of Mexico using DNA metabarcoding. Aquatic Ecology, 57(1):241, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10005-1. See [GPGSMM^{+}23a].

Gost:2023:EAL

[GPS^{+}23] Marc Gost, Samuel Pinya, Antoni Sureda, Silvia Tejada, and Pere Ferriol. Effect of alkalinity and light intensity on the growth of the freshwater sponge Ephydatia fluviatilis (Porifera: Spongillidae). Aquatic Ecology, 57(2):353–367, June 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10014-0.

Gomes:2019:ZFA

[GPV19] Leonardo Fernandes Gomes, Hasley Rodrigo Pereira, and Ludgero Cardoso Galli Vieira. Zooplankton functionalapproach studies in continental aquatic environments: a systematic review. Aquatic Ecology, 53(2):191–203, June 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09682-8.

Geelen:1972:UIL

[GR72] J. F. M. Geelen and J. Ringelberg. Uit instituut en laboratorium. (Dutch) [From institute and laboratory]. Aquatic Ecology, 6(1):3–5, March 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02336202.

Guerrero:1991:SCP

[GR91] Martha Guerrero and Lesbia Rodriguez. Species composition of phytobenthos in the littoral of Lake Xolotlán (Managua). Aquatic Ecology, 25(2):117–120, June 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291241.

George:1997:ZPI

[GR97] D. G. George and C. S. Reynolds. Zooplankton-phytoplankton interactions: the case for refining methods, measurements and models. Aquatic Ecology, 31(1):59–71, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009920214357.

Gosline:2008:PIP

[GR08] Anna K. Gosline and F. Helen Rodd. Predator-induced plasticity in guppy (Poecilia reticulata) life history traits. Aquatic Ecology, 42(4):693–699, December 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9138-7.

Gray:1974:GSM

[Gra74] A. J. Gray. The genecology of salt marsh plants. Aquatic Ecology, 8(1-2):152–165, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02254915.

Gray:2012:SAS

[Gra12] Sarah M. Gray. Succession in the aquatic Sarracenia purpurea community: deterministic or driven by contingency? Aquatic Ecology, 46(4):487–499, December 2012.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9417-9.

Gomez-Rodriguez:2009:MTP

[GRDPP09] Carola Gómez-Rodríguez, Carmen Díaz-Paniagua, and Alexandre Portheault. Mediterranean temporary ponds as amphibian breeding habitats: the importance of preserving pond networks. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9235-x.

Green:2005:VTD

[Gre05] Julian C. Green. Velocity and turbulence distribution around lotic macrophytes. Aquatic Ecology, 39(1):01–10, March 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-004-1913-0.

Grimm:1982:ESP

[Gri82] M. P. Grimm. The evaluation of the stocking of pike fingerlings. Aquatic Ecology, 16(2-3):285–286, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255382.

Grimm:1989:NPE

[Gri89] M. P. Grimm. Northern pike (Esox lucius L.) and aquatic vegetation, tools in the management of fisheries and water quality in shallow waters. Aquatic Ecology, 23(1):59–65, March 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02286427.

Goeppner:2020:FSR

[GRL20] Scott R. Goeppner, Maggie E. Roberts, and Barney Luttbeg. Freshwater snail responses to fish predation integrate phenotypic plasticity and local adaptation. Aquatic Ecology, 54 (1):309–322, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09744-x.

Gross:2022:ACE

[Gro22] Elisabeth M. Gross. Aquatic chemical ecology meets ecotoxicology. Aquatic Ecology, 56(2):493–511, June 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09938-2.

Grubbs:2011:IFP

[Gru11] Scott A. Grubbs. Influence of flow permanence on headwater macroinvertebrate communities in a Cumberland Plateau watershed, USA. Aquatic Ecology, 45(2):185–195, May 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9345-5.

Grzybkowska:1992:DDC

[Grz92] Maria Grzybkowska. Diel drift of Chironomidae in a large lowland river (Central Poland). Aquatic Ecology, 26(2-4):355–360, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255262.

Gislason:2004:SSV

[GS04] Gísli Már Gíslason and Stefán Ó. Steingrímsson. Seasonal and spatial variation in the diet of brown trout (Salmo trutta L.) in the subarctic River Laxá, north-east Iceland. Aquatic Ecology, 38(2):263–270, June 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032052.04874.fb.

Gingerich:2011:RPQ

[GS11] Andrew J. Gingerich and Cory D. Suski. The role of progeny quality and male size in the nesting success of smallmouth bass: integrating field and laboratory studies. Aquatic Ecology, 45(4):505–515, November 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9371-y.

Gallina:2013:PCS

[GSB13] Nicole Gallina, Nico Salmaso, and Martin Beniston. Phytoplankton configuration in six deep lakes in the peri-Alpine region: are the key drivers related to eutrophication and climate? Aquatic Ecology, 47(2):177–193, June 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9433-4.

George:2024:RBP

[GSG24] Nazish George, Ghazala Siddiqui, and Zarish George. Reproductive bionomics and population structure of Conus coronatus Gmelin, 1791 (Gastropoda, Conidae) of Buleji and Manora, Pakistan. Aquatic Ecology, 58(3):583–599, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10092-8.

Gallagher:2017:PRU

[GSH17] Austin J. Gallagher, David S. Shiffman, and N. Hammerschlag. Patterns of resource use and isotopic niche overlap among three species of sharks occurring within a protected subtropical estuary. Aquatic Ecology, 51(3):435–448, September 2017. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-017-9627-2.

Gladyshev:2008:GRD

[GSK08] Michail I. Gladyshev, Nadezhda N. Sushchik, and Galina S. Kalachova. Growth rate of Daphnia feeding on seston in a Siberian reservoir: the role of essential fatty acid. Aquatic Ecology, 42(4):617–627, December 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9146-7.

Gladyshev:2007:SCE

[GSM07] Michail I. Gladyshev, Nadezhda N. Sushchik, and Olesia N. Makhutova. Seasonal correlations of elemental and ω 3 PUFA composition of seston and dominant phytoplankton species in a eutrophic Siberian reservoir. Aquatic Ecology, 41(1):9– 23, March 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-006-9040-8.

Gladyshev:1999:SGC

[GTI99] Michail I. Gladyshev, Tatiana A. Temerova, and Elena A. Ivanova. Selective grazing on Cryptomonas by Ceriodaphnia quadrangula fed a natural phytoplankton assemblage. Aquatic Ecology, 33(4):347–353, December 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009916209394.

G-Toth:2008:FSC

[GTPH08] László G.-Tóth, Sandra Poikane, and Jenica Hanganu. First steps in the Central-Baltic intercalibration exercise on lake macrophytes: where do we start? Aquatic Ecology, 42 (2):265–275, June 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9184-9.

Gudbergsson:2004:ACL

[Gud04] Gudni Gudbergsson. Arctic charr in Lake Myvatn: The centennial catch record in the light of recent stock estimates. Aquatic Ecology, 38(2):271–285, June 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032100.28896.81.

Gulati:1974:LMS

[Gul74] R. D. Gulati. Laboratory methods in secondary production. Aquatic Ecology, 8(3):255–268, December 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02257500.

Gulati:1976:SLO

[Gul76a] R. D. Gulati. Some limnological observations in the lakes of the “Oostelijke Vechtplassengebied”. Aquatic Ecology, 10(1): 3–9, April 1976. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02308489.

Gulati:1976:SFG

[Gul76b] R. D. Gulati. Studies on the food, grazing and metabolism of a filterfeeders community of zooplankton in the lakes Vechten and Tjeukemeer. Aquatic Ecology, 10(1):10–12, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02308490.

Gulati:1985:ZGM

[Gul85] R. D. Gulati. Zooplankton grazing methods using radioactive tracers: Technical problems. Aquatic Ecology, 19(1):61–66, November 1985. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255094.

Gulati:1989:SFA

[Gul89] R. D. Gulati. Structure and feeding activities of zooplankton community in Lake Zwemlust, in the two years after biomanipulation. Aquatic Ecology, 23(1):35–48, March 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02286425.

Gulati:1997:P

[Gul97] Ramesh D. Gulati. Preface. Aquatic Ecology, 31(1):V, March 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009985918033.

Gulati:1998:BRA

[Gul98] Ramesh D. Gulati. Book review: Atlas of the Russian Wetland: Biogeography and metal concentrations. By Zhulidov, J. V. Headley, R. D. Robarts, A. M. Nikanorov and A. A. Ischenko. Aquatic Ecology, 32(4):372–373, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009932009996.

Gulati:1999:BRL

[Gul99] Ramesh D. Gulati. Book review: Lipids in Freshwater Ecosystems. Edited by Michael T. Arts and Bruce C. Wainman. Aquatic Ecology, 33(2):215–218, September 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009954830033.

Gulati:2005:MPR

[Gul05] Ramesh D. Gulati. In memoriam: Professor Robert G. Wetzel. Aquatic Ecology, 39(2):121–122, June 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-7264-z.

Gulati:2007:RNP

[Gul07a] R. D. Gulati. Retirement of Niels De Pauw: a tribute. Aquatic Ecology, 41(3):367–370, September 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9137-8.

Gulati:2007:TYJ

[Gul07b] Ramesh D. Gulati. Ten years of the journal, Aquatic Ecology: an appraisal. Aquatic Ecology, 41(2):149–151, June 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9099-x.

Gulati:2009:BR

[Gul09a] Ramesh D. Gulati. Book review. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9300-5.

Gulati:2009:BRF

[Gul09b] Ramesh D. Gulati. Book review: F. Schiemer, D. Simon, U. S. Amarasinghe & J. Moreau (eds): Aquatic ecosystems and development: comparative Asian perspectives. Aquatic Ecology, 43(1):193–195, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9234-y.

Gatti:2021:NEF

[GUT21] Roberto Cazzolla Gatti, Pero Ugarkovic, and Francesco Tiralongo. New evidence of a fish–bird interspecific feeding association between the European seabass and the European shag in the Mediterranean Sea. Aquatic Ecology, 55(3):1113–1119, September 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09868-z.

Gulati:1989:BNA

[GV89] R. D. Gulati and E. Van Donk. Biomanipulation in The Netherlands: Applications in fresh-water ecosystems and estuarine waters — an introduction. Aquatic Ecology, 23(1):1– 4, March 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02286421.

Gil:1999:DCS

[GV99] Odete Gil and Carlos Vale. DDT concentrations in surficial sediments of three estuarine systems in Portugal. Aquatic Ecology, 33(3):263–269, October 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009961901782.

Geelen:1968:MVV

[GvdH68] J. F. M. Geelen and J. v. d. Heide. Mikrofauna van de venematen. (Dutch) [Microfauna of the veins]. Aquatic Ecology, 2(2):53–56, June 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185810.

Gons:1986:EWS

[GVV86] Herman J. Gons, Roelof Veeningen, and Ronny Van Keulen. Effects of wind on a shallow lake ecosystem: Resuspension of particles in the Loosdrecht Lakes. Aquatic Ecology, 20 (1-2):109–120, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02291156.

Goldspink:1971:F

[GvZ71] Ch. Goldspink and N. P. van Zalinge. Fish. Aquatic Ecology, 5(2):92–97, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185946.

Grindley:1973:PRB

[GW73] J. R. Grindley and T. Woolridge. The plankton of Richard Bay. Aquatic Ecology, 7(3):141, September 1973. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02275616.

Grindley:1974:PRB

[GW74] J. R. Grindley and T. Wooldridge. The plankton of Richards Bay. Aquatic Ecology, 8(1-2):201–212, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254920.

Goldhammer:1992:CPC

[GWF92] D. S. Goldhammer, C. A. Wright, and L. C. Ferrington, Jr. Composition and phenology of Chironomidae from the Nelson Environmental Study Area, University of Kansas. Aquatic Ecology, 26(2-4):281–291, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255253.

Guo:2011:SEF

[GX11] Nichun Guo and Ping Xie. A study on the effects of food quantity and quality on glutathione S-transferase (GST) activity and growth rate parameters of Daphnia carinata varying in age. Aquatic Ecology, 45(1):63–73, March 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9324-x.

Gao:2014:GTF

[GXY14] Zhengquan Gao, Dong Xu, and Naihao Ye. The green tideforming macroalga Ulva linza outcompetes the red macroalga Gracilaria lemaneiformis via allelopathy and fast nutrients uptake. Aquatic Ecology, 48(1):53–62, March 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9465-9.

Gysels:1972:OBV

[Gys72a] H. Gysels. Om het behoud van de kreken en wielen langs de Scheldeboorden. (Dutch) [Preservation of the creeks and wheels along the Scheldt banks]. Aquatic Ecology, 6(4):

196–203, December 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334727.

Gysels:1972:IVE

[Gys72b] H. Gysels. Over de invloed van enkele fysico-chemische faktoren op de faunistische rijkdom van de Noordoost-Vlaamse kreken. (Dutch) [On the influence of some physico-chemical factors on the faunistic wealth of the Northeast Flemish creeks]. Aquatic Ecology, 6(4):172–195, December 1972. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334726.

Guan:2017:EAP

[GYW17] Bo Guan, Junbao Yu, and Xuehong Wang. The ecological adaptability of Phragmites australis to interactive effects of water level and salt stress in the Yellow River Delta. Aquatic Ecology, 51(1):107–116, March 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9602-3.

Gaevsky:2002:VSP

[GZG02] Nikolai A. Gaevsky, Tatiana A. Zotina, and Tamara B. Gorbaneva. Vertical structure and photosynthetic activity of Lake Shira phytoplankton. Aquatic Ecology, 36(2):165–178, April 2002. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1015609011227.

Golovatyuk:2020:MCS

[GZN20] Larisa V. Golovatyuk, Tatiana D. Zinchenko, and Larisa B. Nazarova. Macrozoobenthic communities of the saline Bolshaya Samoroda River (Lower Volga region, Russia): species composition, density, biomass and production. Aquatic Ecology, 54(1):57–74, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09726-z.

Haagsma:1973:EEB

[Haa73] J. Haagsma. The etiology and epidemiology of botulism in waterfowl in The Netherlands. Aquatic Ecology, 7(3):96–105, September 1973. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02275610.

Haahtela:1974:MEF

[Haa74] Ilpo Haahtela. The marine element in the fauna of the Bothnian Bay. Aquatic Ecology, 8(1-2):232–241, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254923.

Hartgers:1998:ETL

[HAB98] Elizabeth M. Hartgers, G. H. (René) Aalderink, and Theo C. M. Brock. Ecotoxicological threshold levels of a mixture of herbicides (atrazine, diuron and metolachlor) in freshwater microcosms. Aquatic Ecology, 32(2):135–152, September 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009968112009.

Hertonsson:2008:CFW

[HÅB08] Pia Hertonsson, Kajsa Åbjörnsson, and Christer Brönmark. Competition and facilitation within and between a snail and a mayfly larva and the effect on the grazing process. Aquatic Ecology, 42(4):669–677, December 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9129-8.

Hadderingh:1979:FIM

[Had79] R. H. Hadderingh. Fish intake mortality at power stations the problem and its remedy. Aquatic Ecology, 13(2-3):83–93, December 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02284734.

Hallegraeff:1976:PDF

[Hal76] Gustaaf M. Hallegraeff. Pigment diversity in fresh-water phytoplankton. III. Summer phytoplankton of eight lakes with widely different trophic characteristics. Aquatic Ecology, 10 (2):87–95, September 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02282215.

Hartog:1973:BRK

[Har73a] C. Den Hartog. Book review: De kranswieren (Charophyta) van Nederland. Wetensch. Meded. K.N.N.V., 93, 44p., 35 figs. Aquatic Ecology, 7(1):44, March 1973. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02279920.

Hartog:1973:PSA

[Har73b] C. Den Hartog. Preliminary survey of the algal vegetation of salt-marshes, a littoral border environment. Aquatic Ecology, 7(1):3–14, March 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02279915.

Hartog:1974:BWC

[Har74] C. Den Hartog. Brackish-water classification, its development and problems. Aquatic Ecology, 8(1-2):15–28, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02254902.

Harrison:1982:CMG

[Har82a] Paul Garth Harrison. Control of microbial growth on eelgrass (Zostera marina L.: Spermatophyta) by leaf-derived metabolites. Aquatic Ecology, 16(1):115, June 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255421.

Hartog:1982:I

[Har82b] C. Den Hartog. Introduction. Aquatic Ecology, 16(1):3–4, June 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255407.

Hartog:1994:DZM

[Har94] C. Den Hartog. The dieback of Zostera marina in the 1930’s in the Waddensea; an eye-witness account by A. van der Werff. Aquatic Ecology, 28(1):51–54, March 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334244.

Haraguchi:2004:SCR

[Har04] Akira Haraguchi. Seasonal changes in redox properties of peat, nutrition and phenology of Menyanthes trifoliata L. in a floating peat mat in Mizorogaike Pond, central Japan. Aquatic Ecology, 38(3):351–357, September 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000035171.00202.5c.

Hossain:2022:ECP

[HAS22] Md. Sanwar Hossain, Syeda Ayshia Akter, and Subrata Sarker. Environmental controls of plankton community dynamics in a sub-tropical river system of Bangladesh. Aquatic Ecology, 56(4):1271–1286, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09974-6.

Herder-Brouwer:1975:DPA

[HB75] S. J. Herder-Brouwer. The development of periphyton on artificial substrates. Aquatic Ecology, 9(2):81–86, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257640.

Helling:1985:CCC

[HB85] G. R. Helling and M. A. Baars. Changes of the concentrations of chlorophyll and phaeopigment in grazing experiments. Aquatic Ecology, 19(1):41–48, November 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255092.

Hummel:1988:ISS

[HB88] H. Hummel and C. Bakker. Introduction to the Schelde Symposium. Aquatic Ecology, 22(1):5, August 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256776.

Haakanson:2003:MPB

[HB03] Lars Håkanson and Viktor V. Boulion. Modelling production and biomasses of zoobenthos in lakes. Aquatic Ecology, 37(3):277–306, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1025854406257.

Hansson:2013:LMR

[HBB13] Lars-Anders Hansson, Jakob Brodersen, and Christer Brönmark. A lake as a microcosm: reflections on developments in aquatic ecology. Aquatic Ecology, 47(2):125–135, June 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9435-2.

Helenius:2013:ETP

[HBL13] Laura K. Helenius, Janica P. G. Borg, and Hannu Lehtonen. The effects of turbidity on prey consumption and selection of zooplanktivorous Gasterosteus aculeatus L. Aquatic Ecology, 47(3):349–356, September 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9448-x.

Haakanson:2003:IBF

[HBO03] Lars Håkanson, Viktor V. Boulion, and Alexander P. Ostapenia. The influence of biomanipulations (fish removal) on the structure of lake foodwebs, case studies using the LakeWebmodel. Aquatic Ecology, 37(1):87–99, January 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1022181302864.

Hoy:2012:SAI

[HBR12] Marshal Hoy, Bruce L. Boese, and Rusty Rodriguez. Salinity adaptation of the invasive New Zealand mud snail (Potamopyrgus antipodarum) in the Columbia River estuary (Pacific Northwest, USA): physiological and molecular studies. Aquatic Ecology, 46(2):249–260, June 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9396-x.

Hough:2009:ADD

[HC09] Zaneta Hough and Charles Andrew Cole. Aboveground decomposition dynamics in riparian depression and slope wetlands of central Pennsylvania. Aquatic Ecology, 43(2):335– 349, June 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9198-3.

Halse:2002:MCA

[HCS02] S. A. Halse, D. J. Cale, and R. J. Shiel. Monitoring change in aquatic invertebrate biodiversity: sample size, faunal elements and analytical methods. Aquatic Ecology, 36(3): 395–410, August 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1016563001530.

Hoogenaad:1979:CRA

[HD79a] H. R. Hoogenaad and A. A. De Groot. Comparison of rhizopod associations. Aquatic Ecology, 13(1):50–55, April 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260308.

Hoogenraad:1979:GVS

[HD79b] H. R. Hoogenraad and A. A. De Groot. Die Geographische Verbreitung der Süsswasser–Rhizopoden. (German) [The geographical distribution of freshwater rhizopods]. Aquatic Ecology, 13(2-3):152–171, December 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284751.

Heinis:1993:FGS

[HD93] F. Heinis and C. Davids. Factors governing the spatial and temporal distribution of Chironomid larvae in the Maarsseveen lakes with special emphasis on the role of oxygen conditions. Aquatic Ecology, 27(1):21–34, July 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02336926.

Higgisson:2021:SGD

[HD21] William Higgisson and Fiona Dyer. Seed germination and dispersal of Eleocharis acuta and Eleocharis sphacelata under experimental hydrological conditions. Aquatic Ecology, 55(1):21–32, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09806-5.

Herrera:2024:RTR

[HdC24] Daphine Ramiro Herrera and Rogerio Caetano da Costa. Reproductive traits, relative growth and maturity of blue crabs Callinectes danae and Callinectes ornatus in South Atlantic waters. Aquatic Ecology, 58(3):963–982, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10118-1.

Hummel:1988:ACG

[HDF88] H. Hummel, L. De Wolf, and A. W. Fortuin. The annual cycle of glycogen in estuarine benthic animals. Aquatic Ecology, 22(2):199–202, December 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256824.

Havel:2000:DCZ

[HEB00] John E. Havel, E. Matt Eisenbacher, and Alice A. Black. Diversity of crustacean zooplankton in riparian wetlands: colonization and egg banks. Aquatic Ecology, 34(1):63–76, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009918703131.

Hensley:2013:TSH

[HEF13] Nicholai M. Hensley, Omar L. Elmasri, and Peggy Fong. Two species of Halimeda, a calcifying genus of tropical macroalgae, are robust to epiphytism by cyanobacteria. Aquatic Ecology, 47(4):433–440, December 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9456-x.

Heip:1988:BAE

[Hei88] Carlo Heip. Biota and abiotic environment in the Westerschelde estuary. Aquatic Ecology, 22(1):31–34, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256779.

Horppila:2009:RAS

[HEPH09] Jukka Horppila, Pertti Eloranta, and Zeynep Pekcan-Hekim. Refuge availability and sequence of predators determine the seasonal succession of crustacean zooplankton in a clayturbid lake. Aquatic Ecology, 43(1):91–103, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9158-3.

Hermens:1983:AQS

[Her83] Joop Hermens. Applications of quantitative structureactivity relationships in aquatic toxicity studies of chemicals and mixtures of chemicals. Aquatic Ecology, 17(1):87– 88, June 1983. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255197.

Heyn:1992:RSP

[Hey92] Michael W. Heyn. A review of the systematic position of the North American species of the genus Glyptotendipes. Aquatic Ecology, 26(2-4):129–137, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255233.

Hamza:2003:PBM

[HEZ03] Waleed Hamza, Peeter Ennet, and Vladimir Zalesny. The 3D physical-biological model study in the Egyptian Mediterranean coastal sea. Aquatic Ecology, 37(3):307–324, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1025836929341.

Hall:1997:CPA

[HF97] J. A. Hall and C. L. J. Frid. Colonisation patterns of adult macrobenthos in a polluted North Sea estuary. Aquatic Ecology, 31(3):333–340, September 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009901502443.

Hornung:2008:CDP

[HF08] J. P. Hornung and A. L. Foote. Comparing dietary preferences of bufflehead ducklings in Western Canada through gut content and stable isotope analysis. Aquatic Ecology, 42(1):61–70, March 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9079-1.

Hesthagen:2008:CLA

[HFS08] Trygve Hesthagen, Peder Fiske, and Brit Lisa Skjelkvåle. Critical limits for acid neutralizing capacity of brown trout (Salmo trutta) in Norwegian lakes differing in organic carbon concentrations. Aquatic Ecology, 42(2):307–316, June 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9191-x.

Haakanson:2005:SFQ

[HG05] Lars Håkanson and Andreas Gyllenhammar. Setting fish quotas based on holistic ecosystem modelling including environmental factors and foodweb interactions — a new approach. Aquatic Ecology, 39(3):325–351, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-3418-x.

Hossain:2020:PMC

[HGB20] Md Shakhawate Hossain, Wei Guo, and Miloš Buřič. Potential of marbled crayfish Procambarus virginalis to supplant invasive Faxonius immunis. Aquatic Ecology, 54(1):45–56, March 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09725-0.

Herrmann:2022:ICF

[HGM22] Alexander Herrmann, Karsten Grabow, and Andreas Martens. The invasive crayfish Faxonius immunis causes the collapse of macroinvertebrate communities in central European ponds. Aquatic Ecology, 56(3):741–750, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09935-5.

Hanselmann:2011:EDT

[HGR11] Almut J. Hanselmann, René Gergs, and Karl-Otto Rothhaupt. Embryonic development time of the freshwater mysid Limnomysis benedeni Czerniavsky as a function of water temperature. Aquatic Ecology, 45(4):539–546, November 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-011-9374-8.

Hulsmann:1980:POC

[HH80] Adrie Hulsmann and Pieter Hengst. Particle organic carbon analysis: apparatus and some results. Aquatic Ecology, 14(3): 135–141, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260114.

Hooker:1991:PBL

[HH91] E. Hooker and S. Hernandez. Phytoplankton biomass in Lake Xolotlán (Managua): Its seasonal and horizontal distribution. Aquatic Ecology, 25(2):125–131, June 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291243.

Haidekker:2008:RBB

[HH08]  Alexandra Haidekker and Daniel Hering.  Relation-

ship between benthic insects (Ephemeroptera, Plecoptera, Coleoptera, Trichoptera) and temperature in small and medium-sized streams in Germany: a multivariate study. Aquatic Ecology, 42(3):463–481, September 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9097-z.

Hylander:2013:VDP

[HH13] Samuel Hylander and Lars-Anders Hansson. Vertical distribution and pigmentation of Antarctic zooplankton determined by a blend of UV radiation, predation and food availability. Aquatic Ecology, 47(4):467–480, December 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9459-7.

Heip:1992:CSF

[HHA92] Carlo Heip, Rony Huys, and Rob Alkemade. Community structure and functional roles of meiofauna in the North Sea. Aquatic Ecology, 26(1):31–41, November 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02298026.

Hu:2021:SEI

[HHC21] Jun Hu, Juxiang Hu, and Shiyun Chi. A study on evaluation index of the biotic integrity based on phytoplankton for a Chinese reservoir. Aquatic Ecology, 55(3):1065–1080, September 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09882-1.

Heneidy:2019:SPD

[HHEM19] Selim Z. Heneidy, Marwa Waseem A. Halmy, and Asmaa M. El-Makawy. The status and potential distribution of Hydrocotyle umbellata L. and Salvinia auriculata Aubl. under climate change scenarios. Aquatic Ecology, 53(4):509–528, December 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09705-4.

Haraguchi:2003:PDP

[HHN03] Akira Haraguchi, Tomoko Hasegawa, and Hiroki Nishijima. The pH dependence of photosynthesis and elongation of Sphagnum squarrosum and S. girgensohnii in the Picea glehnii mire forest in Cape Ochiishi, north-eastern Japan. Aquatic Ecology, 37(1):101–104, January 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1022130622499.

Huuskonen:2012:SMH

[HHP12] Hannu Huuskonen, Helena Haakana, and Jorma Piironen. Seasonal movements and habitat use of river whitefish (Coregonus lavaretus) in the Koitajoki River (Finland), as determined by Carlin tagging and acoustic telemetry. Aquatic Ecology, 46(3):325–334, September 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9403-2.

Hamerlynck:1993:MEF

[HHV93] Olivier Hamerlynck, Kris Hostens, and Paul A. Van Damme. The mobile epibenthic fauna of soft bottoms in the Dutch Delta (south-west Netherlands): Spatial structure. Aquatic Ecology, 27(2-4):343–358, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02334797.

Hu:2021:CCO

[HHY21] Siwen Hu, Rujia He, and Zhongbo Yu. Composition and co-occurrence patterns of Phragmites australis rhizosphere bacterial community. Aquatic Ecology, 55(2):695–710, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09855-4.

Higler:1968:KTV

[Hig68] L. W. G. Higler. Kokerjuffers (Trichoptera) in een verlandingszone van de venematen. (Dutch) [Tube damselflies (Trichoptera) in a landing zone of the veins]. Aquatic Ecology, 2(2):57–66, June 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185811.

Higler:1969:TES

[Hig69] L. W. G. Higler. Trichoptera en Ephemeroptera in het stratiotetum van de pias venematen. (Dutch) [Trichoptera and Ephemeroptera in the stratiotetum of the pias veins]. Aquatic Ecology, 3(3):75–82, November 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185858.

Higler:1971:MLE

[Hig71] L. W. G. Higler. Makrofauna in de Linde en enige petgaten van de Lindevallei. (Dutch) [Makrofauna in the Linde and some peat bogs in the Linde Valley]. Aquatic Ecology, 5(3): 126–141, September 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185890.

Higler:1974:BAB

[Hig74] L. W. G. Higler. The biological assessment of bodies of water by means of the macrofauna. Aquatic Ecology, 8(3):285–287, December 1974. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257504.

Higler:1976:OMD

[Hig76] L. W. G. Higler. Observations on the macrofauna of a Dutch ditch. Aquatic Ecology, 10(1):66–73, April 1976. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02308503.

Higler:1979:LDD

[Hig79] LW. G. Higler. Limnological data on a Dutch moorland pool through sixty years. Aquatic Ecology, 13(2-3):138–143, December 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02284749.

Higler:1980:PCF

[Hig80] L. W. G. Higler. The presence of caddis flies in the Netherlands and their role in the aquatic system. Aquatic Ecology, 14(3):209–212, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260122.

Higler:1981:BFL

[Hig81] L. W. G. Higler. Bottom fauna and littoral vegetation fauna in Lake Maarsseveen. Aquatic Ecology, 15(1-2):82–86, October 1981. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260261.

Highler:1989:HRP

[Hig89] L. W. G. Highler. Hydrobiological research in peat polder ditches. Aquatic Ecology, 23(2):105–109, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256727.

Higgins:2009:SVF

[Hig09] Christopher L. Higgins. Spatiotemporal variation in functional and taxonomic organization of stream-fish assemblages in central Texas. Aquatic Ecology, 43(4):1133–1141, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9233-z.

Hillebrand:1973:EIF

[Hil73] H. Hillebrand. Ecological investigations on filamentous freshwater algae. Aquatic Ecology, 7(1):25–27, March 1973. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02279917.

Hani:2023:TVD

[HJ23] P. M. Hani and K. J. Jayalakshmi. Temporal variation in diversity, abundance and size class structure of planktonic copepods from a tropical estuary. Aquatic Ecology, 57(1):199–216, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10003-3.

Hasu:2008:EGF

[HJV08] Tiina Hasu, Jukka Jokela, and E. Tellervo Valtonen. Effects of growth factors and water source on laboratory cultures of a northern Asellus aquaticus (Isopoda) population. Aquatic Ecology, 42(1):141–150, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9089-z.

Hillebrand:1976:DMB

[HK76] H. Hillebrand and S. P. Klapwijk. Distribution of multicellular benthic algae in an eutrophic ditch. Aquatic Ecology, 10(1):48–58, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02308501.

Haverkamp:1982:PMS

[HK82] J. Haverkamp and P. G. Kistemaker. Pyrolysis mass spectrometric analysis of complex bio-organic matter. Aquatic Ecology, 16(1):115, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255420.

Hossain:2019:SWR

[HKB19] Md Shakhawate Hossain, Jan Kubec, and Miloš Buřič. Still waters run deep: marbled crayfish dominates over red swamp crayfish in agonistic interactions. Aquatic Ecology, 53(1): 97–107, March 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09675-7.

Hancock:2023:TWR

[HKH^{+}23] Mark H. Hancock, Daniela Klein, Robert Hughes, Paul Stagg, Paul Byrne, Trevor D. Smith, Alison MacLennan, Paul P. J. Gaffney, and Colin W. Bean. Testing whether reducing brown trout biomass in peatland lakes increases macro-invertebrate biomass and invertivorous waterbird occurrence. Aquatic Ecology, 57(1):217–240, March 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-10000-y.

Hammock:2012:IIP

[HKJ12] Bruce G. Hammock, Nickilou Y. Krigbaum, and Michael L. Johnson. Incorporating invertebrate predators into theory regarding the timing of invertebrate drift. Aquatic Ecology, 46(2):153–163, June 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9388-x.

Hamaoka:2014:ULB

[HKO14] Hideki Hamaoka, Atsushi Kaneda, and Koji Omori. Upwelling-like bottom intrusion enhances the pelagic– benthic coupling by a fish predator in a coastal food web. Aquatic Ecology, 48(1):63–71, March 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9466-8.

Hansen:2016:FDB

[HKT16] Brenda K. Hansen, Amy C. Krist, and Teresa M. Tibbets. Foraging differences between the native snail, Fossaria sp. and the invasive New Zealand mudsnail (Potamopyrgus antipodarum) in response to phosphorus limitation. Aquatic Ecology, 50(2):297–306, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9578-z.

Hartog:1995:ORA

[HL95] C. Den Hartog and G. M. Lokhorst. Occurrence of the Rhodophycean alga Thorea ramosissima bory in the River Meuse. Aquatic Ecology, 29(1):117–120, April 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02061794.

Harrow-Lyle:2022:NNC

[HLK22] Tyler J. Harrow-Lyle and Andrea E. Kirkwood. The nonnative charophyte Nitellopsis obtusa (starry stonewort) influences shifts in macrophyte diversity and community structure in lakes across a geologically heterogeneous landscape. Aquatic Ecology, 56(3):829–840, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09950-0.

Hammou:2014:TEV

[HLL14] H. Ait Hammou, D. Latour, and M. Loudiki. Temporal evolution and vertical stratification of Microcystis toxic potential during a first bloom event. Aquatic Ecology, 48(2):219– 228, June 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9477-0.

Hsieh:2025:SEF

[HLL^{+}25] Tsung-Tse Hsieh, Jia-Yi Li, Che-Chang Lin, Mei-Hwa Kuo, and Ming-Chih Chiu. Seasonality and extreme flooding can override the effects of debris dam removal on riverine species: a case study of macroinvertebrates. Aquatic Ecology, 59 (1):203–214, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10157-8.

Hernandez-Lucero:2024:BDR

[HLSN24] José Antonio Hernández-Lucero, S. S. S. Sarma, and S. Nandini. Behavioral and demographic responses of the predatory rotifer Asplanchna sieboldii (Leydig, 1854) fed prey (Plationus patulus (Müller, 1786)) previously exposed to cadmium and microplastics. Aquatic Ecology, 58(2):239–248, June 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10061-7.

Higler:1984:ETR

[HM84] L. W. G. Higler and A. W. M. Mol. Ecological types of running water based on stream hydraulics in The Netherlands. Aquatic Ecology, 18(1):51–57, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256674.

Hosper:1986:CPL

[HM86] Harry Hosper and Marie-Louise Meyer. Control of phosphorus loading and flushing as restoration methods for Lake Veluwe, The Netherlands. Aquatic Ecology, 20(1-2):183–194, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291162.

Hummel:1988:CET

[HMB88] Herman Hummel, Gerard Moerland, and Cees Bakker. The concomitant existence of a typical coastal and a detritus food chain in the Westerschelde estuary. Aquatic Ecology, 22(1):35–41, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256780.

Hemmi:2005:RGP

[HMH05] Anne Hemmi, Anita Mäkinen, and Tuija Honkanen. Responses of growth and phlorotannins in Fucus vesiculosus to nutrient enrichment and herbivory. Aquatic Ecology, 39 (2):201–211, June 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-004-3526-z.

Hirose:2003:SCA

[HNiN03] Miho Hirose, Yuichiro Nishibe, and Shin ichi Nakano. Seasonal changes in the abundance of autotrophic picoplankton and some environmental factors in hypereutrophic Furuike Pond. Aquatic Ecology, 37(1):37–43, January 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1022128306570.

Hodson:1998:BRF

[Hod98] S. Hodson. Book review: Fouling organisms of the Indian Ocean: Biofoul-ing and Control Technology. By R. Nagabhushanam and M. F. Thompson. Aquatic Ecology, 32(4): 367–368, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009982612722.

Hoffsten:1999:DFF

[Hof99] Per-Ola Hoffsten. Distribution of filter-feeding caddisflies (Trichoptera) and plankton drift in a Swedish lake-outlet stream. Aquatic Ecology, 33(4):377–386, December 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009926509349.

Hogendijk:1969:OVC

[Hog69] C. J. Hogendijk. Overzicht van chemische en fysische gegevens in de omgeving van Beltschutsloot. (Dutch) [Overview of chemical and physical data in the vicinity of Beltschutsloot]. Aquatic Ecology, 3(2):50–59, July 1969. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185851.

Hollwedel:1970:FMH

[Hol70] Werner Hollwedel. Funde Von Macrothrix hirsuticornis Norman u. Brady Auf Ameland. (German) [Finds of Macrothrix hirsuticornis Norman and Brady from Ameland]. Aquatic Ecology, 4(3):170–171, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185881.

Hosper:1980:DPA

[Hos80] S. H. Hosper. Development and practical application of limiting values for the phosphate concentration in surface waters in the Netherlands. Aquatic Ecology, 14(1-2):64–72, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260274.

Hosper:1989:BNP

[Hos89] S. H. Hosper. Biomanipulation, new perspectives for restoration of shallow, eutrophic lakes in The Netherlands. Aquatic Ecology, 23(1):5–10, March 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02286422.

Houpt:1994:MTD

[Hou94] P. M. Houpt. Marine tube-dwelling diatoms and their occurrence in the Netherlands. Aquatic Ecology, 28(1):77–84, March 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334248.

Huntjens:1975:P

[HP75] J. L. M. Huntjens and S. Parma. Publications. Aquatic Ecology, 9(3):139–140, December 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02263333.

Hornung:2006:IDP

[HP06] Christine L. R. Hornung and Charlie Pacas. Investigating damselfly populations at springs in Banff National Park, Canada, with special focus on Argia vivida, Amphiagrion abbreviatum, and Ischnura cervula (Odonata: Coenagrionidae). Aquatic Ecology, 40(1):49–58, March 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9017-z.

Hills:2020:CRF

[HPM20] Amber Hills, Simon Pouil, and Teresa J. Mathews. Clearance rates of freshwater bivalves Corbicula fluminea and Utterbackia imbecillis in the presence and absence of light. Aquatic Ecology, 54(4):1059–1066, December 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09793-7.

Hernandez:2000:ESP

[HPPL00] Ignacio Hernández, Abraham Pérez-Pastor, and J. Lucas Pérez Lloréns. Ecological significance of phosphomonoesters and phosphomonoesterase activity in a small Mediterranean river and its estuary. Aquatic Ecology, 34 (2):107–117, June 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009930405572.

Hortnagl:2011:CEU

[HPS11] Paul Hörtnagl, María Teresa Pérez, and Ruben Sommaruga. Contrasting effects of ultraviolet radiation on the growth efficiency of freshwater bacteria. Aquatic Ecology, 45(1): 125–136, March 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-010-9341-9.

Hugula:1995:PCC

[HPT95] J. L. Hugula, J. C. Philippart, and J. P. Thome. PCB contamination of the common barbel, Barbus barbus (pisces, cyprinidae), in the River Meuse in relation to hepatic monooxygenase activity and ultrastructural liver change. Aquatic Ecology, 29(1):135–145, April 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02061796.

Hooker:1991:PCC

[HRP91] E. Hooker, M. Ruiz, and M. Pum. Phytoplankton community composition in Lake Xolotlán (Managua). Aquatic Ecology, 25(2):121–124, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02291242.

Heinis:1985:ICM

[HS85] F. Heinis and W. R. Swain. Impedance conversion as a method of research for assessing behavioural responses of aquatic invertebrates. Aquatic Ecology, 19(2):183–192, December 1985. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02270765.

Heinis:1986:ICM

[HS86] F. Heinis and W. R. Swain. Impedance conversion as a method of research for assessing behavioural responses of aquatic invertebrates. Aquatic Ecology, 20(1-2):261, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291168.

Herrera-Silveira:2010:ENI

[HSCR10] Jorge A. Herrera-Silveira, Just Cebrian, and Peter Ralph. Evidence of negative impacts of ecological tourism on turtlegrass (Thalassia testudinum) beds in a marine protected area of the Mexican Caribbean. Aquatic Ecology, 44(1):23–31, March 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9260-9.

Hamilton:2016:MHC

[HSP16] David P. Hamilton, Nico Salmaso, and Hans W. Paerl. Mitigating harmful cyanobacterial blooms: strategies for control of nitrogen and phosphorus loads. Aquatic Ecology, 50(3): 351–366, September 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9594-z.

Huang:2018:RMS

[HSY18] Xiaolong Huang, Nan Shen, and Dan Yu. Root morphological and structural comparisons of introduced and native aquatic plant species in multiple substrates. Aquatic Ecology, 52(1):65–76, March 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9645-0.

Huiskes:1988:SMW

[Hui88] H. L. Huiskes. The salt marshes of the Westerschelde and their role in the estuarine ecosystem. Aquatic Ecology, 22 (1):57–63, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256782.

Hillebrand:1979:PFA

[HV79] H. Hillebrand and C. P. J. Van Dijk. Patterns of freshwater algae on artificial substrates at different sampling scales. Aquatic Ecology, 13(2-3):118–119, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284746.

Higler:1989:MDD

[HV89] L. W. G. Higler and P. F. M. Verdonschot. Macroinvertebrates in the demmerik ditches (The Netherlands): The role of environmental structure. Aquatic Ecology, 23(2):143–150, December 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256731.

Hadderingh:1983:FIP

[HvAV83] R. H. Hadderingh, G. H. F. M. van Aerssen, and J. W. Van Der Stoep. Fish impingement at power stations situated along the rivers Rhine and Meuse in The Netherlands. Aquatic Ecology, 17(2):129–141, December 1983. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02280822.

Huisman:1985:DMS

[HvB85] Jolanda Huisman and L. W. C. A. van Breemen. Distribution of macroinvertebrates in the shallow part of ‘De Gijster’, a water storage lake in the Biesbosch (The Netherlands). Aquatic Ecology, 19(2):199–206, December 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270767.

Hallegraeff:1997:TTT

[HVB97] Gustaaf M. Hallegraeff, Joseph P. Valentine, and Christopher J. Bolch. Temperature tolerances of toxic dinoflagellate cysts: application to the treatment of ships’ ballast water. Aquatic Ecology, 31(1):47–52, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009972931195.

Hunting:2010:MSA

[HvdGvS10] Ellard R. Hunting, Harm G. van der Geest, and Rob W. M. van Soest. Mangrove-sponge associations: a possible role for tannins. Aquatic Ecology, 44(4):679–684, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9306-z.

Hillebrand:1984:MST

[HvdH84] Herman Hillebrand and Joop van der Heide. A multichannel sensor for temperature measurements in shallow waters. Aquatic Ecology, 18(1):35–38, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256671.

Hemminga:1993:EGD

[HVK93] M. A. Hemminga, J. Van Soelen, and W. Kiswara. Estuarine gradients determining the quality of tidal marsh halophytes as host plants for endophagous insect larvae: Experimental evidence. Aquatic Ecology, 27(2-4):377–382, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334799.

hommevanReine:1969:BDB

[hvRB69] W. F. Prud homme van Reine and C. Bakker. Boeken. (Dutch) [Books]. Aquatic Ecology, 3(1):4–6, March 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185838.

Hadderingh:1987:EHE

[HVS87] R. H. Hadderingh, G. Van Der Velde, and P. G. Schnabel. The effect of heated effluent on the occurrence and the reproduction of the freshwater limpets Ancylus fluviatilis Müller, 1774, Ferrissia wautieri (Mirolli, 1960) and Acroloxus lacustris (L., 1758) in two Dutch water bodies. Aquatic Ecology, 21(2):193–205, December 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255445.

Hintz:2013:CVI

[HW13] William D. Hintz and Todd Wellnitz. Current velocity influences the facilitation and removal of algae by stream grazers. Aquatic Ecology, 47(2):235–244, June 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9438-z.

Heip:1977:VDM

[HWG77] C. Heip, K. A. Willems, and A. Goossens. Vertical distribution of meiofauna and the efficiency of the Van Veeno grab on sandy bottoms in lake Grevelingen (the Netherlands). Aquatic Ecology, 11(2):35–45, September 1977. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02265707.

Hansen:2011:DDA

[HWK11] Joakim P. Hansen, Sofia A. Wikström, and Lena Kautsky. Distribution differences and active habitat choices of invertebrates between macrophytes of different morphological complexity. Aquatic Ecology, 45(1):11–22, March 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9319-7.

Huang:2022:DCA

[HWS22] Jingyu Huang, Xia Wang, and Shuang Song. Distribution characteristics of ammonia-oxidizing microorganisms and their responses to external nitrogen and carbon in sediments of a freshwater reservoir, China. Aquatic Ecology, 56(3): 841–857, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09943-z.

Hogsden:2009:AFW

[HXR09] Kristy L. Hogsden, Marguerite A. Xenopoulos, and James A. Rusak. Asymmetrical food web responses in trophic-level richness, biomass, and function following lake acidification. Aquatic Ecology, 43(2):591–606, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9169-8.

Hardwick:2003:TDT

[HYH03] Elaine O. Hardwick, Wenying Ye, and Robert E. Hodson. Temporal dynamics of three culturable γ-Proteobacteria taxa in salt marsh sediments. Aquatic Ecology, 37(1):55–64, January 2003. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1022137514010.

Hongguang:1995:MDB

[HZC95] Ma Hongguang, Yu Zhiying, and Gerhard C. Cadée. Macrofauna distribution and bioturbation on tidal confluences of the Dutch Wadden Sea. Aquatic Ecology, 29(2):167–176, June 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02336047.

He:2022:IMT

[HZZ22] Dandan He, Youmei Zeng, and Guangming Zhou. The influence of microplastics on the toxic effects and biodegradation of bisphenol A in the microalgae Chlorella pyrenoidosa. Aquatic Ecology, 56(4):1287–1296, December 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09966-6.

Iniguez-Armijos:2018:LUE

[IAHB18] Carlos Iñiguez-Armijos, Henrietta Hampel, and Lutz Breuer. Land-use effects on structural and functional composition of benthic and leaf-associated macroinvertebrates in four Andean streams. Aquatic Ecology, 52(1):77–92, March 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9646-z.

Ibelings:2016:CSI

[IBV16] Bastiaan W. Ibelings, Myriam Bormans, and Petra M. Visser. CYANOCOST special issue on cyanobacterial blooms: synopsis — a critical review of the management options for their prevention, control and mitigation. Aquatic Ecology, 50(3):595–605, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9596-x.

Ironside:2019:SFE

[IDP19] Joseph Edward Ironside, Samuel Thomas Dalgleish, and William Payne. Sex or food? Effects of starvation, size and diet on sexual cannibalism in the amphipod crustacean Gammarus zaddachi. Aquatic Ecology, 53(1):1–7, March 2019.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9668-1.

Ibelings:2016:CBE

[IFV16a] Bastiaan W. Ibelings, Jutta Fastner, and Petra M. Visser. Cyanobacterial blooms. Ecology, prevention, mitigation and control: Editorial to a CYANOCOST special issue. Aquatic Ecology, 50(3):327–331, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9595-y. See erratum [IFV16b].

Ibelings:2016:ECB

[IFV16b] Bastiaan W. Ibelings, Jutta Fastner, and Petra M. Visser. Erratum to: Cyanobacterial blooms. Ecology, prevention, mitigation and control: Editorial to a CYANOCOST special issue. Aquatic Ecology, 50(4):735, December 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9600-5. See [IFV16a].

Imholt:2010:IRC

[IGS10] C. Imholt, C. N. Gibbins, and C. Soulsby. Influence of riparian cover on stream temperatures and the growth of the mayfly Baetis rhodani in an upland stream. Aquatic Ecology, 44(4):669–678, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9305-0.

Infante-Izquierdo:2021:VSR

[IIMRC21] M. Dolores Infante-Izquierdo, Adolfo. F. Muñoz-Rodríguez, and Jesús M. Castillo. Variation in sexual reproductive output among exotic taxa of Spartina (Poaceae). Aquatic Ecology, 55(1):107–123, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
020-09815-4.

Irvine:2006:SVN

[IJ06] Robyn L. Irvine and Leland J. Jackson. Spatial variance of nutrient limitation of periphyton in montane, headwater streams (McLeod River, Alberta, Canada). Aquatic Ecology, 40(3):337–348, September 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
006-9036-4.

Irfanullah:2006:EDP

[IM06] Haseeb Md. Irfanullah and Brian Moss. Ecology of Dictyosphaerium pulchellum Wood (Chlorophyta, Chlorococcales) in a shallow, acid, forest lake. Aquatic Ecology, 40 (1):1–12, March 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-9011-5.

Im:2020:EFA

[INJ20] Hyungjoon Im, Joorim Na, and Jinho Jung. The effect of food availability on thermal stress in Daphnia magna: tradeoffs among oxidative stress, somatic growth, and reproduction. Aquatic Ecology, 54(4):1201–1210, December 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09804-7.

Nakano:2005:DMHa

[iNTT05a] Shin ichi Nakano, Yuji Tomaru, and Hidetaka Takeoka. The dynamics of microbial and herbivorous food webs in a coastal sea with special reference to intermittent nutrient supply from bottom intrusion. Aquatic Ecology, 38(4):485–493, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-0441-x.

Nakano:2005:DMHb

[iNTT05b] Shin ichi Nakano, Yuji Tomaru, and Hidetaka Takeoka. The dynamics of microbial and herbivorous food webs in a coastal sea with special reference to intermittent nutrient supply from bottom intrusion. Aquatic Ecology, 38(4):485–493, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-0441-2.

Ingvason:2004:FST

[IÓG04] Haraldur R. Ingvason, Jón S. Ólafsson, and Arnthor Gardarsson. Food selection of Tanytarsus gracilentus larvae (Diptera: Chironomidae): an analysis of instars and cohorts. Aquatic Ecology, 38(2):231–237, June 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032053.67992.03.

Islam:2006:FAT

[IRT06] Md. Shahidul Islam, Mohammed Mokhlesur Rahman, and Masaru Tanaka. Fish assemblage of a traditional fishery and the seasonal variations in diet of its most abundant species Wallago attu (Siluriformes: Siluridae) from a tropical floodplain. Aquatic Ecology, 40(2):263–272, June 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9023-1.

Irvine:1986:DFB

[Irv86] Kenneth Irvine. Differential feeding behaviour of the dominant Cladocera as an explanation of zooplankton community structure in the Loosdrecht Lakes. Aquatic Ecology, 20(1-2): 121–134, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02291157.

Irvine:1989:SDZ

[Irv89] Kenneth Irvine. The spatial distribution of zooplankton populations in the presence and absence of fish. Aquatic Ecology, 23(2):169–178, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256734.

Ilmonen:2006:IPC

[IS06] Jari Ilmonen and Jukka Suhonen. Intraguild predation, cannibalism, and microhabitat use in Calopteryx virgo and Somatochlora metallica larvae: a laboratory experiment. Aquatic Ecology, 40(1):59–68, March 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9019-x.

Ibelings:2013:E

[IS13] Bas Ibelings and Piet Spaak. Editorial. Aquatic Ecology, 47(2):123, June 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-013-9441-4.

Imo:2010:DIG

[ISJ10] Miriam Imo, Alfred Seitz, and Jes Johannesen. Distribution and invasion genetics of the quagga mussel (Dreissena rostriformis bugensis) in German rivers. Aquatic Ecology, 44(4): 731–740, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9311-2.

Islam:2010:SMLa

[ISY10a] Md. Shahidul Islam, Keita W. Suzuki, and Yoh Yamashita. Selective mortality of larval Japanese seabass in Ariake Bay, Japan. Aquatic Ecology, 44(1):309–316, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9271-6.

Islam:2010:SMLb

[ISY10b] Md. Shahidul Islam, Keita W. Suzuki, and Yoh Yamashita. Selective mortality of larval Japanese seabass in Ariake Bay, Japan. Aquatic Ecology, 44(1):317–324, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9281-4.

Islam:2009:DPS

[IT09] Md. Shahidul Islam and Masaru Tanaka. Diet and prey selection in larval and juvenile Japanese anchovy Engraulis japonicus in Ariake Bay, Japan. Aquatic Ecology, 43(2):549– 558, June 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9207-6.

Ikram:2022:CSC

[IUK22] Sana F. Ikram, Vaishnavi Uniyal, and Dhananjay Kumar. Changes in species composition of cyanobacterial and microalgal communities along a temperature gradient in Tapovan Hot Spring, Garhwal Himalaya, Uttarakhand, India. Aquatic Ecology, 56(3):573–584, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09921-x.

Janse:1990:MPF

[JA90] J. H. Janse and T. Aldenberg. Modelling phosphorus fluxes in the hypertrophic Loosdrecht Lakes. Aquatic Ecology, 24(1):69–89, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256750.

Jacobsen:1992:DLF

[Jac92] Richard E. Jacobsen. Descriptions of the larvae of four nearctic species of Epoicocladius (Diptera: Chironomidae) with a redescription of Epoicocladius ephemerae (Kieffer). Aquatic Ecology, 26(2-4):145–155, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255235.

Janssen:2015:EEE

[JAM15] Annette B. G. Janssen, George B. Arhonditsis, and Wolf M. Mooij. Exploring, exploiting and evolving diversity of aquatic ecosystem models: a community perspective. Aquatic Ecology, 49(4):513–548, December 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
015-9544-1.

Jansma:1975:QDA

[Jan75] M. J. Jansma. Quantitative diatom analysis as an archaeological aid. Aquatic Ecology, 9(2):87–88, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02257641.

Jara:2008:TOL

[Jar08] Fabián Gastón Jara. Tadpole–odonate larvae interactions: influence of body size and diel rhythm. Aquatic Ecology, 42(3):503–509, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
007-9110-6.

Jahnke:1983:CPE

[JB83] J. Jahnke and M. Baumann. Chemical and physical effects on the shape and growth of the diatom Biddulphia sinensis Greville in batch cultures: a contribution to bioindication in plankton ecology. Aquatic Ecology, 17(1):5–20, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255188.

Jahnke:1987:DBP

[JB87] J. Jahnke and M. E. M. Baumann. Differentiation between Phaeocystis pouchetii (Har.) Lagerheim and Phaeocystis globosa Scherffel. Aquatic Ecology, 21(2):141–147, December 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255439.

Jaarsma:2007:MID

[JBB07] N. G. Jaarsma, M. Bergman, and A. Bij de Vaate. Macroinvertebrates in a dynamic river environment: analysis of time series from artificial substrates, using a ‘white box’ neural network modelling method. Aquatic Ecology, 41(3): 413–425, September 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-9016-0.

Julius:2005:SSS

[JBS05] Matthew L. Julius, Richard W. Blob, and Heiko L. Schoenfuss. The survival of Sicyopterus stimpsoni, an endemic amphidromous Hawaiian gobiid fish, relies on the hydrological cycles of streams: evidence from changes in algal composition of diet through growth stages fish. Aquatic Ecology, 39(4): 473–484, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-9007-1.

Jana:1980:PSS

[JC80] Sasadhar Jana and Monojit A. Choudhuri. Physiology of some submersed aquatic angiosperms: effects of heavy metals. Aquatic Ecology, 14(3):152–157, December 1980. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260116.

Jana:1983:UTS

[JC83] Sasadhar Jana and Monojit A. Choudhuri. ^{32}P uptake in three submerged aquatic plant species. Aquatic Ecology, 17 (2):97–101, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02280818.

Jia:2017:CAM

[JCF17] Qiang Jia, Lei Cao, and Anthony David Fox. Combating aggressive macrophyte encroachment on a typical Yangtze River lake: lessons from a long-term remote sensing study of vegetation. Aquatic Ecology, 51(1):177–189, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9609-9.

Jiang:2009:IPF

[JCL09] Fu Ying Jiang, Xin Chen, and An Cheng Luo. Iron plaque formation on wetland plants and its influence on phosphorus, calcium and metal uptake. Aquatic Ecology, 43(4):879–890, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9241-z.

Jan:2001:LTM

[JCS01] Rong-Quen Jan, Jeng-Ping Chen, and Kwang-Tsao Shao. Long-term monitoring of the coral reef fish communities around a nuclear power plant. Aquatic Ecology, 35(2):233– 243, June 2001. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1011496117632.

Julian:2023:RNE

[JDF23] Arango-Lozano Julián, Patiño-Siro Dahian, and Toro-Cardona Felipe. Reaching new environments through illegal trade: evidence of a widely traded turtle in Colombia. Aquatic Ecology, 57(2):471–480, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10023-z.

Jayasree:2025:PBI

[JDJ^{+}25] Parvathy Ashokkumar Jayasree, Bini Cyril Das, Jifiriya Marathippallam Jamal, Chandhini Sathyajith, Tincy Varghese, Devika Pillai, and Rejish Kumar Vattiringal Jayadradhan. Physio-biochemical and immunological responses of Etroplus suratensis (Cichliformes) to chronic ammonia exposure. Aquatic Ecology, 59(1):263–275, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10161-y.

Jewson:2009:ESD

[JGG09] David H. Jewson, Nick G. Granin, and Ruslan Yu Gnatovsky. Effect of snow depth on under-ice irradiance and growth of Aulacoseira baicalensis in Lake Baikal. Aquatic Ecology, 43(3):673–679, September 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9267-2.

Johnson:1997:DIS

[JGH97] Robert L. Johnson, Elisabeth M. Gross, and Nelson G. Hairston, Jr. Decline of the invasive submersed macrophyte Myriophyllum spicatum (Haloragaceae) associated with herbivory by larvae of Acentria ephemerella (Lepidoptera). Aquatic Ecology, 31(3):273–282, September 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009960131857.

Jaballah:2023:DLA

[JGM^{+}23] Sami Jaballah, Guglielmo Fernandez Garcia, François Martignac, Nicolas Parisey, Stéphane Jumel, Jean-Marc Roussel, and Olivier Dézerald. A deep learning approach to detect and identify live freshwater macroinvertebrates. Aquatic Ecology, 57(4):933–949, December 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10053-7.

Jayawardana:2023:BPA

[JGUF23] J. M. C. K. Jayawardana, W. D. T. M. Gunawardana, E. P. N. Udayakumara, and S. Fernando. Biomonitoring of pesticides in agricultural river catchments: a case study from two river catchments in tropical Sri Lanka. Aquatic Ecology, 57(2):337–352, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10013-1.

Jin:2025:EFF

[JJG^{+}25] Zong’an Jin, Hui Jin, Baoyan Gao, Chunfu Tong, Erik Jeppesen, Lars G. Rudstam, Henri J. Dumont, María de los Ángeles González Sagrario, Vladimir Razlutskij, Zhengwen Liu, Yali Tang, and Xiufeng Zhang. Effects of filter-feeding bivalves in benthic and pelagic habitats on plankton community and water quality in shallow systems: implications for lake rehabilitation. Aquatic Ecology, 59(1):53–66, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10147-w.

Jeong:2024:ANN

[JJP24] Seonah Jeong, Sungbae Joo, and Sangkyu Park. Applying a neural network machine learning model to predict seasonal allelopathic inhibitory effects of Myriophyllum spicatum on the growth of Microcystis aeruginosa. Aquatic Ecology, 58 (2):349–361, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10073-3.

Joseph:2006:SDF

[JLG06] Venitia Joseph, Andrea Locke, and Jean-Guy J. Godin. Spatial distribution of fishes and decapods in eelgrass (Zostera marina L.) and sandy habitats of a New Brunswick estuary, eastern Canada. Aquatic Ecology, 40(1):111–123, March 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-9027-x.

Jakkila:2009:RTH

[JLS09] Juho Jakkila, Matti Leppäranta, and Kalevi Salonen. Radiation transfer and heat budget during the ice season in Lake Pääjärvi, Finland. Aquatic Ecology, 43(3):681–692, September 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9275-2.

Jana:1984:TCP

[JM84] B. B. Jana and A. K. Manna. Temporal changes in protein level of plankton, benthic invertebrates and carp growing in fish ponds. Aquatic Ecology, 18(1):59–66, June 1984. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256675.

Jacobsen:2008:BAS

[JM08] Dean Jacobsen and Rubén Marín. Bolivian Altiplano streams with low richness of macroinvertebrates and large diel fluctuations in temperature and dissolved oxygen. Aquatic Ecology, 42(4):643–656, December 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9127-x.

Joiris:1973:EPB

[Joi73] C. Joiris. Évolution des populations bacteriennes planctoniques dans le bassin de chasse d’Ostende (Belgique) en 1971 et 1972. (French) [Evolution of planktonic bacterial populations in the hunting basin of Ostend (Belgium) in 1971 and 1972]. Aquatic Ecology, 7(2):54–59, June 1973. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02282198.

Janakiraman:2024:RHD

[JTGS24] A. Janakiraman, S. Thangaraj, V. Gokula, and A. Stalin. Report on high density population of a cladoceran, Moina macrocopa and their gut microbes in anoxic sewage water: a potential bio-indicator of aquatic pollution. Aquatic Ecology, 58(4):1255–1265, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
024-10139-w.

Jacobucci:2009:TVA

[JTL09] Giuliano Buzá Jacobucci, Marcel Okamoto Tanaka, and Fosca Pedini Pereira Leite. Temporal variation of amphipod assemblages associated with Sargassum filipendula (Phaeophyta) and its epiphytes in a subtropical shore. Aquatic Ecology, 43(4):1031–1040, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9230-2.

Juttner:2005:EPA

[Jüt05] Friedrich Jüttner. Evidence that polyunsaturated aldehydes of diatoms are repellents for pelagic crustacean grazers. Aquatic Ecology, 39(3):271–282, September 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-3419-9.

Jannel:2025:ABR

[JVCJ25] Lou-Anne Jannel, Pierre Valade, Pascale Chabanet, and Philippe Jourand. Aquatic biodiversity on Reunion Island: responses of biological communities to environmental and anthropogenic pressures using environmental DNA. Aquatic Ecology, 59(1):363–391, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10168-5.

Jonsson:2006:MWF

[JvDS06] Per R. Jonsson, Luca A. van Duren, and Estelle Schaaff. Making water flow: a comparison of the hydrodynamic characteristics of 12 different benthic biological flumes. Aquatic Ecology, 40(4):409–438, December 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9049-z.

Janse:1995:MNC

[JVG95] J. H. Janse, E. Van Donk, and R. D. Gulati. Modelling nutrient cycles in relation to food web structure in a biomanipulated shallow lake. Aquatic Ecology, 29(1):67–79, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02061790.

Kouwenhoven:1986:FSM

[KA86] Peter Kouwenhoven and Tom Aldenberg. A first step in modelling plankton growth in the Loosdrecht Lakes. Aquatic Ecology, 20(1-2):135–145, November 1986. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291158.

Kasai:1997:IEU

[KA97] Fumie Kasai and Michael T. Arts. The interactive effects of UV-b radiation and a herbicide on uptake and allocation of carbon in two strains of the green alga Scenedesmus. Aquatic Ecology, 31(3):261–272, September 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02493513.

Kristmannsdottir:2004:GLM

[KÁ04] Hrefna Kristmannsdóttir and Halldór Ármannsson. Groundwater in the Lake Myvatn area, northern Iceland: Chemistry, origin and interaction. Aquatic Ecology, 38(2):115–128, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000032067.47495.71.

Kaelin:2016:LLP

[KA16] Katharina Kaelin and Florian Altermatt. Landscape-level predictions of diversity in river networks reveal opposing patterns for different groups of macroinvertebrates. Aquatic Ecology, 50(2):283–295, June 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
016-9576-1.

Kaeding:2020:NCR

[Kae20] Lynn R. Kaeding. New climate regime started and further shaped the historic Yellowstone Lake cutthroat trout population decline commonly attributed entirely to nonnative lake trout predation. Aquatic Ecology, 54(2):641–652, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09765-x.

Kal:1986:MMB

[Kal86] B. F. M. Kal. Monthly mass balances for compartments of the Loosdrecht Lakes system: Approach and preliminary results. Aquatic Ecology, 20(1-2):27–39, November 1986. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291148.

Koeberle:2020:FAA

[KAN20] Alex L. Koeberle, Ivan Arismendi, and David L. G. Noakes. Fluctuating asymmetry of adult Chinook salmon (Oncorhynchus tshawytscha) otoliths from wild and hatchery origins. Aquatic Ecology, 54(1):431–446, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09733-0.

Kappers:1976:BGA

[Kap76] F. I. Kappers. Blue-green algae in the sediment of the Lake Brielse Meer. Aquatic Ecology, 10(3):164–171, December 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02263365.

Kappers:1980:GLF

[Kap80] F. I. Kappers. The growth-limiting factor of a population of cyanobacteria in the hypertrophic Dutch Lake Brielle. Aquatic Ecology, 14(1-2):36–46, August 1980. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260271.

Kappers:1982:TCN

[Kap82] F. I. Kappers. Toxic cyanobacteria in the Netherlands. Aquatic Ecology, 16(2-3):291, December 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255387.

Kat:1992:YYV

[Kat92] Marie Kat. Year-to-year variation in the occurrence of some dominant phytoplankton species in Dutch coastal waters, 1973–1984. Aquatic Ecology, 25(3):225–231, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270807.

Kulkoyluoglu:2020:NMO

[KAY20] Okan Külköylüoğlu, Derya Akdemir, and Mehmet Yavuzatmaca. Non-marine Ostracoda (Crustacea) as indicator species group of habitat types. Aquatic Ecology, 54(2):519– 533, June 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09757-x.

Kandlakunta:2020:MSF

[KBN20] Laxminarsimha Chary Kandlakunta, Sreekanth G. B., and Sharma Nitin. Marine soundscape and fish chorus in an archipelago ecosystem comprising bio-diverse tropical islands off Goa Coast, India. Aquatic Ecology, 54(2):475–493, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09754-0.

Kristoffersson:1974:ACS

[KBO74] Rolf Kristoffersson, Saara Broberg, and Aimo Oikari. Annual changes in some blood constituents of female brackish water Zoarces viviparus (L.), Teleostei, with special reference to the reproductive cycle and the embryotrophe. Aquatic Ecology, 8(1-2):117–123, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02254912.

Khromechek:2010:DDF

[KBR10] E. B. Khromechek, Y. V. Barkhatov, and D. Y. Rogozin. Densities and distribution of flagellates and ciliates in the chemocline of saline, meromictic Lake Shunet (Siberia, Russia). Aquatic Ecology, 44(3):497–511, September 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9332-x.

Krivtsov:2005:ECM

[KBS05] V. Krivtsov, E. G. Bellinger, and D. C. Sigee. Elemental composition of Microcystis aeruginosa under conditions of lake nutrient depletion. Aquatic Ecology, 39(2):123–134, June 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-004-6833-5.

Kar:1982:EDI

[KC82] Rup Kumar Kar and Monojit A. Choudhuri. Effect of desiccation on internal changes with respect to survival of Hydrilla verticillata. Aquatic Ecology, 16(2-3):213–221, December 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255375.

Kalinda:2022:ELI

[KC22] Chester Kalinda and Moses J. Chimbari. Effects of laboratory-induced desiccation on fecundity and survival of Bulinus globosus (Gastropoda: Planorbidae). Aquatic Ecology, 56(1):143–152, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
021-09903-z.

Kulkoyluoglu:2022:HDW

[KÇA22] Okan Külköylüoğlu, Beyza Çelikbaş, and Alper Ataman. Habitat destruction in wetland affects Ostracoda (Crustacea) species occurrence patterns amid different aquatic habitats. Aquatic Ecology, 56(3):605–618, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09934-6.

Kreutzweiser:2005:EFS

[KCG05] David P. Kreutzweiser, Scott S. Capell, and Kevin P. Good. Effects of fine sediment inputs from a logging road on stream insect communities:a large-scale experimental approach in a Canadian headwater stream. Aquatic Ecology, 39(1):55–66, March 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-5066-y.

Klepper:1992:WQD

[KCR92] O. Klepper, G. T. Chairuddin, and H. D. Rijksen. Water quality and the distribution of some fishes in an area of acid sulphate soils, Kalimantan, Indonesia. Aquatic Ecology, 25 (3):217–224, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02270806.

Koning:1982:DTS

[KD82] Kees Koning and Jaap Dorgelo. A distinct temporal separation of maximum bacterial numbers and phytoplankton biomass in the open-water zone of Lake Maarsseveen I: Results and literature review. Aquatic Ecology, 16(2-3):181–190, December 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255371.

Kirillin:2009:BSI

[KEH09] Georgiy Kirillin, Christof Engelhardt, and Thomas Hintze. Basin-scale internal waves in the bottom boundary layer of ice-covered Lake Müggelsee, Germany. Aquatic Ecology, 43(3):641–651, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9274-3.

Keller:1973:EGC

[Kel73] K. J. Keller. Electricity generation and cooling water problems. Aquatic Ecology, 7(3):83–85, September 1973. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02275608.

Kantoussan:2009:RSP

[KEL09] Justin Kantoussan, Jean Marc Ecoutin, and Raymond Laë. The relevance of size parameters as indicators of fishery exploitation in two West African reservoirs. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9236-9.

Kennedy:2011:IDF

[Ken11] Victor S. Kennedy. The invasive dark falsemussel Mytilopsis leucophaeata (Bivalvia: Dreissenidae): a literature review. Aquatic Ecology, 45(2):163–183, May 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9344-6.

Kersting:1970:GVC

[Ker70] K Kersting. Gebruik van de Coulter counter bij het onderzoek naar de voedselopname van Daphnia magna. (Dutch) [Use of the Coulter counter in the study of the food intake of Daphnia magna]. Aquatic Ecology, 4(3):151–152, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185878.

Kersting:1975:UME

[Ker75] Kees Kersting. The use of microsystems for the evaluation of the effect of toxicants. Aquatic Ecology, 9(3):102–108, December 1975. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02263327.

Kersting:1978:GEDa

[Ker78a] Kees Kersting. Growth efficiency of Daphinia magna. I. The effect of food concentration. Aquatic Ecology, 12(1):3–20, April 1978. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260789.

Kersting:1978:GEDb

[Ker78b] Kees Kersting. The growth efficiency of Daphnia magna II. The effect of temperature. Aquatic Ecology, 12(2):99– 106, July 1978. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260710.

Kersting:1981:ADO

[Ker81] Kees Kersting. Annual and diel oxygen and temperature regime of the Lakes Maarsseveen. Aquatic Ecology, 15(1-2):10–28, October 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260255.

Kersting:1983:MET

[Ker83] Kees Kersting. Micro-ecosystems and toxicants: a search for effects on the ecosystem level of organisation. Aquatic Ecology, 17(1):89, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255199.

Kersting:1985:SPU

[Ker85] Kees Kersting. Specific problems using electronic particle counters. Aquatic Ecology, 19(1):5–12, November 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255087.

Kersting:1991:RFD

[Ker91] Kees Kersting. Recycling of food by Daphnia magna Straus (Cladocera): a possible role in swarming. Aquatic Ecology, 25(1):73–75, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259592.

Kerkis:1992:KFD

[Ker92] I. E. Kerkis. Karyotypes of four Diamesinae species (Diptera) from Russia. Aquatic Ecology, 26(2-4):157–162, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255236.

Kersting:1997:WES

[Ker97] Kees Kersting. On the way to eternity; the success of an aquatic laboratory microecosystem. Aquatic Ecology, 31 (1):29–35, March 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009991318551.

Ketelaars:1995:ODB

[KF95] Henk A. M. Ketelaars and Nicole M. L. H. F. Frantzen. One decade of benthic macroinvertebrate biomonitoring in the River Meuse. Aquatic Ecology, 29(1):121–133, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02061795.

Kamer:2004:RIS

[KFS04] Krista Kamer, Peggy Fong, and Kenneth Schiff. The relative importance of sediment and water column supplies of nutrients to the growth and tissue nutrient content of the green macroalga Enteromorpha intestinalis along an estuarine resource gradient. Aquatic Ecology, 38(1):45–56, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000021041.31385.19.

Ketelaars:1994:REP

[KG94] Henk A. M. Ketelaars and Luc Gille. Range extension of the predatory cladoceran Bythotrephes longimanus Leydig 1860 (Crustacea, Onychopoda) in Western Europe. Aquatic Ecology, 28(2):175–180, June 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02333988.

Kolmakov:2003:GPP

[KG03] Vladimir I. Kolmakov and Michail I. Gladyshev. Growth and potential photosynthesis of cyanobacteria are stimulated by viable gut passage in crucian carp. Aquatic Ecology, 37 (3):237–242, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1025801326088.

Koch:2004:PGO

[KGB04] Richard W. Koch, Debbie L. Guelda, and Paul A. Bukaveckas. Phytoplankton growth in the Ohio, Cumberland and Tennessee Rivers, USA: inter-site differences in light and nutrient limitation. Aquatic Ecology, 38(1):17–26, March 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000021082.42784.03.

Kalogianni:2010:FEC

[KGC10] Eleni Kalogianni, Sofia Giakoumi, and Yorgos Chatzinikolaou. Feeding ecology of the critically endangered Valencia letourneuxi (Valenciidae). Aquatic Ecology, 44(1):289–299, March 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9253-8.

Katayama:2013:IPE

[KGM13] Naoki Katayama, Tetsuo Goto, and Tadashi Miyashita. Indirect positive effects of agricultural modernization on the abundance of Japanese tree frog tadpoles in rice fields through the release from predators. Aquatic Ecology, 47 (2):225–234, June 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-013-9437-0.

Katayama:2022:CIP

[KGM22] Naoki Katayama, Tetsuo Goto, and Tadashi Miyashita. Correction to: Indirect positive effects of agricultural modernization on the abundance of Japanese tree frog tadpoles in rice fields through the release from predators. Aquatic Ecology, 56(1):331, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09916-8.

Kalacheva:2002:CAL

[KGS02] Galina S. Kalacheva, Vladimir G. Gubanov, and Sergei V. Savitsky. Chemical analysis of Lake Shira water (1997– 2000). Aquatic Ecology, 36(2):123–141, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1015695813280.

Kondo:1992:HPF

[KH92] Shigeo Kondo and Shigetaka Hamashima. Habitat preferences of four chironomid species associated with aquatic macrophytes in an irrigation reservoir. Aquatic Ecology, 26 (2-4):371–377, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255264.

Kronvang:1999:RNR

[KHD99] Brian Kronvang, Carl Christian Hoffmann, and Jesper Dørge. Retention of nutrients in river basins. Aquatic Ecology, 33(1):29–40, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009947907811.

Kraiem:2001:SOA

[KHF01] M. M. Kraïem, C. Ben Hamza, and R. J. Flower. Some observations on the age and growth of thin-lipped grey mullet, Liza ramada Risso, 1826 (Pisces, Mugilidae) in three North African wetland lakes: Merja Zerga (Morocco), Garâat Ichkeul (Tunisia) and Edku Lake (Egypt). Aquatic Ecology, 35(3-4):335–345, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1011900631096.

Knorek:2020:SVN

[KHG20] Zofia R. Knorek, Bruce P. Hansen, and Aaron W. E. Galloway. Seasonal variability in a nascent population of a non-indigenous colonial ascidian (Didemnum vexillum) near Winchester Bay, Oregon. Aquatic Ecology, 54(4):895–907, December 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09774-w.

Kuprijanov:2017:END

[KHK17] Ivan Kuprijanov, Kristjan Herkül, and Jonne Kotta. Ecological niche differentiation between native and non-native shrimps in the northern Baltic Sea. Aquatic Ecology, 51(3): 389–404, September 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9624-5.

Kjaran:2004:LCS

[KHM04] Snorri Páll Kjaran, Sigurdur Lárus Hólm, and Eric Matthew Myer. Lake circulation and sediment transport in Lake Myvatn. Aquatic Ecology, 38(2):145–162, June 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032049.94886.5a.

Kravchuk:2006:SDA

[KIG06] Elena S. Kravchuk, Elena A. Ivanova, and Michail I. Gladyshev. Seasonal dynamics of akinetes of Anabaena flos-aquae in bottom sediments and water column of small Siberian reservoir. Aquatic Ecology, 40(3):325–336, September 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9031-9.

Kim:2011:DDB

[KIJ11] Ok-Sun Kim, Johannes F. Imhoff, and Pilar Junier. Distribution of denitrifying bacterial communities in the stratified water column and sediment–water interface in two freshwater lakes and the Baltic Sea. Aquatic Ecology, 45(1):99–112, March 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-010-9335-7.

Kim:1999:ECM

[Kim99] Baik-Ho Kim. Ecology of a cyanobacterial mat community in a Korean thermal wastewater stream. Aquatic Ecology, 33(4):331–338, December 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009986606414.

Katano:2005:AGG

[KiNS05] Toshiya Katano, Shin ichi Nakano, and Masahito Sugiyama. Abundance, growth and grazing loss rates of picophytoplankton in Barguzin Bay, Lake Baikal. Aquatic Ecology, 39(4): 431–438, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-9000-8.

Kirsch:2022:FGI

[Kir22] Danielle R. Kirsch. Freshwater gastropods as an important group for studying the impact of inter- and intraspecific chemical communication on aquatic community dynamics. Aquatic Ecology, 56(2):361–375, June 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09961-x.

Kelley:2000:SPC

[KJF00] Randall Kelley, Jeffrey Jack, and Michael Fant. A survey of the physiochemical characteristics and the zooplankton communities in an ephemeral karst lake. Aquatic Ecology, 34(1):77–89, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009922828999.

Kobak:2014:EZM

[KJP14] Jarosław Kobak, Łukasz Jermacz, and Dariusz Pła̧chocki. Effectiveness of zebra mussels to act as shelters from fish predators differs between native and invasive amphipod prey. Aquatic Ecology, 48(4):397–408, December 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9492-1.

Knudsen:2015:DGP

[KJS15] Rune Knudsen, Helge Johnsen, and Sten Ivar Siikavuopio. Divergent growth patterns between juveniles of two sympatric Arctic charr morphs with contrasting depth gradient niche preferences. Aquatic Ecology, 49(1):33–42, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9502-y.

Kersting:1989:ADO

[KK89] Kees Kersting and Peter Kouwenhoven. Annual and diel oxygen regime in two polder ditches. Aquatic Ecology, 23(2): 111–123, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256728.

Kaller:2007:AMA

[KK07] Michael D. Kaller and William E. Kelso. Association of macroinvertebrate assemblages with dissolved oxygen concentration and wood surface area in selected subtropical streams of the southeastern USA. Aquatic Ecology, 41 (1):95–110, March 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-006-9046-2.

Kapiris:2009:MSR

[KK09] Kostas Kapiris and Stefanos Kavvadas. Morphometric study of the red shrimp Aristeus antennatus in the Eastern Mediterranean. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9224-5.

Kobak:2011:EIA

[KK11] Jarosław Kobak and Tomasz Kakareko. The effectiveness of the induced anti-predator behaviour of zebra mussel Dreissena polymorpha in the presence of molluscivorous roach Rutilus rutilus. Aquatic Ecology, 45(3):357–366, September 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-011-9359-7.

Kaczorowska:2012:PEC

[KK12a] Anna Kaczorowska and Ryszard Kornijów. Palaeoecological evidence for changes over the past 200 years in chironomid communities of a shallow lake exposed to cyanobacterial toxins. Aquatic Ecology, 46(4):465–473, December 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9415-y.

Kang:2012:ISP

[KK12b] Sung-Ryong Kang and Sammy L. King. Influence of salinity and prey presence on the survival of aquatic macroinvertebrates of a freshwater marsh. Aquatic Ecology, 46(4): 411–420, December 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9410-3.

Kankanamge:2017:EIC

[KK17] Champika Ellawala Kankanamge and Hasini Kodithuwakku. Effect of interspecific competition on the growth and nutrient uptake of three macrophytes in nutrient-rich water. Aquatic Ecology, 51(4):625–634, December 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9640-5.

Kalinowska:2020:IID

[KK20] Krystyna Kalinowska and Maciej Karpowicz. Ice-on and ice-off dynamics of ciliates and metazooplankton in the Łuczański Canal (Poland). Aquatic Ecology, 54(4):1121– 1134, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09797-3.

Kopylov:2002:SPM

[KKD02] Alexander I. Kopylov, Dmitriy B. Kosolapov, and Andrey G. Degermendzhy. Structure of planktonic microbial food web in a brackish stratified Siberian lake. Aquatic Ecology, 36 (2):179–204, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1015678918611.

Kalachova:2004:SDA

[KKI04] Galina S. Kalachova, Andzhela A. Kolmakova, and Elena A. Ivanova. Seasonal dynamics of amino acids in two small Siberian reservoirs dominated by prokaryotic and eukaryotic phytoplankton. Aquatic Ecology, 38(1):3–15, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000021044.55658.71.

Kinzler:2009:MPB

[KKM09a] Werner Kinzler, Axel Kley, and Gerhard Maier. Mutual predation between and cannibalism within several freshwater gammarids: Dikerogammarus villosus versus one native and three invasives. Aquatic Ecology, 43(2):457–464, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9206-7.

Kley:2009:ISP

[KKM09b] Axel Kley, Werner Kinzler, and Gerhard Maier. Influence of substrate preference and complexity on co-existence of two non-native gammarideans (Crustacea: Amphipoda). Aquatic Ecology, 43(4):1047–1059, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9242-y.

Kone:2008:FRS

[KKO08] Tidiani Koné, Essétchi P. Kouamélan, and Frans Ollevier. First results of a study of the feeding habits of Synodontis comoensis (Siluriformes: Mochokidae) in a West African river (Comoé River, Comoé National Park, Côte d’Ivoire). Aquatic Ecology, 42(1):35–42, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9074-y.

Kopylov:2002:PBP

[KKR02] Alexander I. Kopylov, Dmitriy B. Kosolapov, and Anna V. Romanenko. Phytoplankton, bacterial production and protozoan bacterivory in stratified, brackish-water Lake Shira (Khakasia, Siberia). Aquatic Ecology, 36(2):205–218, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1015611023296.

Kamimura:2011:PPS

[KKS11] Yasuhiro Kamimura, Akihide Kasai, and Jun Shoji. Production and prey source of juvenile black rockfish Sebastes cheni in a seagrass and macroalgal bed in the Seto Inland Sea, Japan: estimation of the economic value of a nursery. Aquatic Ecology, 45(3):367–376, September 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9360-1.

Ketelaars:1992:TSD

[KKV92a] H. A. M. Ketelaars, A. M. J. P. Kuijpers, and L. W. C. A. Van Breemen. Temporal and spatial distribution of chironomid larvae and oligochaetes in two Dutch storage reservoirs. Aquatic Ecology, 26(2-4):361–369, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255263.

Kuijpers:1992:CPE

[KKV92b] A. M. J. P. Kuijpers, H. A. M. Ketelaars, and L. W. C. A. Van Breemen. Chironomid pupal exuviae and larvae of two storage reservoirs in The Netherlands. Aquatic Ecology, 26 (2-4):379–383, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255265.

Klapwijk:1980:ELW

[Kla80] S. P. Klapwijk. Effects of laundry wastewater on benthic algae in ditches in the Netherlands. Aquatic Ecology, 14(3): 142–151, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260115.

Klapwijk:1988:HRD

[Kla88] S. P. Klapwijk. Historical and recent data of waterchemistry and phytoplankton in the Rijnland area. Aquatic Ecology, 22(1):93–94, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256790.

Kleef:1984:STS

[Kle84] H. L. Kleef. A simple and time-saving method for quantitative collection of Corophium volutator (Crustacea: Amphipoda) from plankton samples rich in detritus. Aquatic Ecology, 18(1):47–50, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256673.

Kajtoch:2017:DFG

[KLG17] Łukasz Kajtoch, Peter Lešo, and Robert Gwiazda. Do flocks of great cormorants and goosanders avoid spatial overlap in foraging habitat during the non-breeding season? Aquatic Ecology, 51(3):473–483, September 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9630-7.

Klomp:1976:LRM

[Klo76] R. Klomp. Limnological research for the management of water quality in storage reservoirs. Aquatic Ecology, 10(3): 129–135, December 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02263361.

Kooijman:1983:DCS

[KM83] S. A. L. M. Kooijman and J. A. J. Metz. On the dynamics of chemically stressed populations: The deduction of population consequences from effects on individuals. Aquatic Ecology, 17(1):88–89, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255198.

Knorp:2017:ESM

[KM17] Natalie E. Knorp and Justin N. Murdock. Exclusion size and material have minimal effects on stream benthic algae and macroinvertebrate colonization within submerged cages. Aquatic Ecology, 51(4):545–556, December 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9635-2.

Kousri:2023:EES

[KMG^{+}23] Kaouther Mejri Kousri, Marouan Meddeb, Boutheina Grami, Sondes Melliti Ben Garali, Oumayma Chkili, Inès Sahraoui, Catherine Gonzalez, Chrystelle Montigny, Olivier Pringault, and Asma Sakka Hlaili. Effects of experimental sediment resuspension on protozooplankton grazing activity: implication for the planktonic food web structure. Aquatic Ecology, 57(1):165–186, March 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09998-y.

Krodkiewska:2009:BOC

[KMK09] M. Krodkiewska and A. Michalik-Kucharz. The bottom Oligochaeta communities in sand pits of different trophic status in Upper Silesia (Southern Poland). Aquatic Ecology, 43(2):437–444, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9199-2.

Kamermans:2002:EGI

[KML02] Pauline Kamermans, Erik-Jan Malta, and Angela Tjin A. Lien. Effect of grazing by isopods and amphipods on growth of Ulva spp. (Chlorophyta). Aquatic Ecology, 36(3):425–433, August 2002. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1016551911754.

Kornijow:2019:FTC

[KML19] Ryszard Kornijów, Marina Markiyanova, and Evgenia Lange. Feeding by two closely related species of Chironomus (Diptera: Chironomidae) in south Baltic lagoons, with implications for competitive interactions and resource partitioning. Aquatic Ecology, 53(3):315–324, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09691-7.

Kankanamge:2020:SMA

[KMR20] Champika Ellawala Kankanamge, Fleur E. Matheson, and Tenna Riis. Shading may alter the colonization pattern and dominance between two invasive submerged aquatic plant species. Aquatic Ecology, 54(3):721–728, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09770-0.

Khaleafa:1982:EDL

[KMS82a] A. F. Khaleafa, A. F. Mohsen, and S. H. Shaalan. Effect of different light intensities on growth, amino-acid, fat and sugar concentrations in Caulerpa prolifera (Foerskal) Lamouroux. Aquatic Ecology, 16(2-3):207–212, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255374.

Khaleafa:1982:SVG

[KMS82b] A. F. Khaleafa, A. F. Mohsen, and S. H. Shaalan. Seasonal variations in the growth and amino acid pattern of Caulerpa prolifera (Foerskal) Lamouroux. Aquatic Ecology, 16(2-3): 201–206, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255373.

Kamp-Nielsen:1986:MRH

[KN86] Lars Kamp-Nielsen. Modelling the recovery of hypertrophic L. Glumsø (Denmark). Aquatic Ecology, 20(1-2):245–255, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291166.

Kaddeche:2022:DFD

[KND^{+}22] Hadjer Kaddeche, Faïza Noune, Sabrina Dzizi, Nadjla Chaib, Zine Eddine Boudjellab, and Saúl Blanco. Determinant factors of diatom assemblage’s distribution along the Coastal Central Constantine (Northeastern Algeria). Aquatic Ecology, 56(4):1245–1269, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09980-8.

Kefford:2007:STF

[KNH07] Ben J. Kefford, Dayanthi Nugegoda, and Kathryn L. Hassell. The salinity tolerance of freshwater macroinvertebrate eggs and hatchlings in comparison to their older life-stages: a diversity of responses. Aquatic Ecology, 41(2):335–348, June 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-006-9066-y.

Kankaala:2000:RLV

[KOA00] Paula Kankaala, Anne Ojala, and Lauri Arvola. Response of littoral vegetation on climate warming in the boreal zone; an experimental simulation. Aquatic Ecology, 34(4):433–444, December 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1011457815299.

Koops:1973:IHE

[Koo73] F. B. J. Koops. Investigation of hydrobiological effects due to cooling water discharges of electric power plants. Aquatic Ecology, 7(3):86–95, September 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02275609.

Kool:1976:UAT

[Koo76] H. J. Kool. The use of adenosine triphosphate (ATP) as an indicator of biological activity and biomass in ecological studies. Aquatic Ecology, 10(3):145–154, December 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02263363.

Koperski:2021:LNE

[Kop21] Paweł Koperski. Linear and nonlinear effects of nutrient enrichments on the diversity of macrobenthos in lowland watercourses. Aquatic Ecology, 55(3):1011–1031, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09878-x.

Kornijow:1998:QSC

[Kor98] Ryszard Kornijów. Quantitative sampler for collecting invertebrates associated with submersed and floating-leaved macrophytes. Aquatic Ecology, 32(3):241–244, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009972525967.

Kornijow:2014:QSC

[Kor14] Ryszard Kornijów. A quantitative sampler for collecting invertebrates associated with deep submerged vegetation. Aquatic Ecology, 48(4):417–422, December 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9494-z.

Kouwe:1983:PRS

[Kou83] F. A. Kouwe. Phosphorus removal in a sewage treatment plant and plant biomass in an effluent receiving lowland stream. Aquatic Ecology, 17(2):143–156, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02280823.

Kowalewska:1994:SCE

[Kow94] Grażyna Kowalewska. Steryl chlorin esters in sediments of the southern Baltic Sea. Aquatic Ecology, 28(2):149–156, June 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02333985.

Kube:1997:FCD

[KP97] Jan Kube and Martin Powilleit. Factors controlling the distribution of Marenzelleria cf. viridis, Pygospio elegans and Streblospio shrubsoli (Polychaeta: Spionidae) in the southern Baltic Sea, with special attention for the response to an event of hypoxia. Aquatic Ecology, 31(2):187–198, June 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009955505252.

Kromkamp:1999:EPP

[KP99] Jacco Kromkamp and Jan Peene. Estimation of phytoplankton photosynthesis and nutrient limitation in the Eastern Scheldt estuary using variable fluorescence. Aquatic Ecology, 33(1):101–104, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009900124650.

Khan:2024:ZCA

[KP24] Qasim Khan and Saltanat Parveen. Zooplankton composition, abundance and diversity in Sagartaal, Budaun district, Uttar Pradesh, India. Aquatic Ecology, 58(3):949–962, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10117-2.

Klosowski:2023:RBP

[KPJP23] Stanisław Kłosowski, Paweł Pawlikowski, Ewa Jabłońska, and Monika Podgórska. The relationships between the physical and chemical properties of an aquatic environment and the floristic specificity of pleustonic communities in northern Poland. Aquatic Ecology, 57(2):383–395, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10016-y.

Kiili:2009:DDU

[KPS09] Mari Kiili, Merja Pulkkanen, and Kalevi Salonen. Distribution and development of under-ice phytoplankton in 90-m deep water column of Lake Päijänne (Finland) during spring convection. Aquatic Ecology, 43(3):707–713, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9262-7.

Kovacs:2016:TDG

[KPV16] Attila W. Kovács, Mátyás Présing, and Lajos Vörös. Thermal-dependent growth characteristics for Cylindrospermopsis raciborskii (Cyanoprokaryota) at different light availabilities: methodological considerations. Aquatic Ecology, 50 (4):623–638, December 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9582-3.

Kraak:1995:ZMR

[Kra95] Michiel H. S. Kraak. Are zebra mussels restricted by toxicant levels in the River Meuse? A review. Aquatic Ecology, 29 (1):147–150, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02061797.

Kobbi-Rebai:2020:ARV

[KRATA20] Rayda Kobbi-Rebai, Neila Annabi-Trabelsi, and Habib Ayadi. Abundance and reproduction variables of two species of harpacticoid copepods along an increasing salinity gradient. Aquatic Ecology, 54(1):387–400, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09749-x.

Krebs:1979:MDF

[Kre79] B. P. M. Krebs. Microchironomus deribae (Freeman, 1957) (Diptera, Chironomidae) in the Delta region of The Netherlands. Aquatic Ecology, 13(2-3):144–151, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284750.

Krebs:1982:NDS

[Kre82] B. P. M. Krebs. Note on the distribution of Sigara selecta (Fieber) in the brackish inland waters of the south-west Netherlands. Aquatic Ecology, 16(2-3):159–164, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255369.

Kobayashi:2009:STR

[KRJ09] Tsuyoshi Kobayashi, Darren S. Ryder, and Stephen J. Jacobs. Short-term response of nutrients, carbon and planktonic microbial communities to floodplain wetland inundation. Aquatic Ecology, 43(4):843–858, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9219-2.

Kroes:1980:RPD

[Kro80] H. W. Kroes. Replacement of phosphates in detergents. Aquatic Ecology, 14(1-2):90–93, August 1980. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260276.

Kroes:1986:RSL

[Kro86] H. W. Kroes. Restoration of shallow lake ecosystems, with emphasis on Loosdrecht Lakes. Aquatic Ecology, 20(1-2):5–7, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291145.

Kosolapov:2003:MSR

[KRZ03] Dmitriy B. Kosolapov, Denis Yu. Rogozin, and Elena E. Zakharova. Microbial sulfate reduction in a brackish meromictic steppe lake. Aquatic Ecology, 37(3):215–226, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1025871300917.

Khaleafa:1979:IST

[KS79] A. F. Khaleafa and S. Shaalan. The interaction of salinity and temperature on the growth of Caulerpa prolifera (Forskål) Lamouroux. Aquatic Ecology, 13(2-3):172–176, December 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02284752.

Klinkenberg:1994:MOA

[KS94] G. Klinkenberg and R. Schumann. Micro-organism activity in aggregate layers in shallow eutrophic brackish water as influenced by wind induced mixing; an experimental approach. Aquatic Ecology, 28(3-4):421–426, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334212.

Knapinska-Skiba:1995:BPC

[KSBŁ95] D. Knapińska-Skiba, R. Bojanowski, and M. Łotocka. The biological and physico-chemical uptake of radiocesium by particulate matter of natural origin (Baltic Sea). Aquatic Ecology, 29(3-4):283–290, December 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084226.

Knapinska-Skiba:1994:SRR

[KSBR94] D. Knapińska-Skiba, R. Bojanowski, and Z. Radecki. Sorption and release of radiocaesium from particulate matter of the Baltic coastal zone. Aquatic Ecology, 28(3-4):413–419, September 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334211.

Kappenberg:1995:SDE

[KSF95] J. Kappenberg, G. Schymura, and H.-U. Fanger. Sediment dynamics and estuarine circulation in the turbidity maximum of the Elbe river. Aquatic Ecology, 29(3-4):229–237, December 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02084220.

Kiknadze:1992:SSR

[KSW92] I. I. Kiknadze, M. T. Siirin, and W. F. Wülker. Siberian species of the riihimakiensis-group in the genus Chironomus (Diptera, Chironomidae). 1. Karyotypes and morphology. Aquatic Ecology, 26(2-4):163–171, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255237.

Kanaya:2019:ASI

[KSY19] Gen Kanaya, Mikhail M. Solovyev, and Natalia I. Yurlova. Application of stable isotopic analyses for fish host–parasite systems: an evaluation tool for parasite-mediated material flow in aquatic ecosystems. Aquatic Ecology, 53(2):217–232, June 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09684-6.

Kasige:2009:CNP

[KT09] Anusha Kasige and Asaeda Takashi. Carbon and nitrogen partitioning in the freshwater submerged macrophyte Vallisneria gigantea in response to ultraviolet-b irradiance. Aquatic Ecology, 43(2):313–322, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9201-z.

Kloskowski:2015:FRH

[KT15] J. Kloskowski and A. Trembaczowski. Fish reduce habitat coupling by a waterbird: evidence from combined stable isotope and conventional dietary approaches. Aquatic Ecology, 49(1):21–31, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9501-z.

Kappes:2014:DDC

[KTH14] Heike Kappes, Oliver Tackenberg, and Peter Haase. Differences in dispersal- and colonization-related traits between taxa from the freshwater and the terrestrial realm. Aquatic Ecology, 48(1):73–83, March 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
013-9467-7.

Kankaala:2010:DCZ

[KTJ10] Paula Kankaala, Sami Taipale, and Roger I. Jones. Diets of crustacean zooplankton, inferred from stable carbon and nitrogen isotope analyses, in lakes with varying allochthonous dissolved organic carbon content. Aquatic Ecology, 44(4): 781–795, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-010-9316-x.

Kuklinska:1992:CCN

[Kuk92] Bozena Kuklińska. Chironomidae communities of the nearshore zone in Zegrzyńsky Reservoir, Poland. Aquatic Ecology, 26(2-4):385–392, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255266.

Kulkoyluoglu:2005:EPF

[Kül05] Okan Külköylüoğlu. Ecology and phenology of freshwater ostracods in Lake Gölköy (Bolu, Turkey). Aquatic Ecology, 39(3):295–304, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
005-0782-5.

Kurashov:2002:RML

[Kur02] Evgeny A. Kurashov. The role of meiobenthos in lake ecosystems. Aquatic Ecology, 36(3):447–463, August 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1016535921472.

Koeman:1970:AVD

[KvdM70] J. H. Koeman and H. L. van der Maas. Accumulatie van diuron in vis. (Dutch) [Accumulation of diurone in fish]. Aquatic Ecology, 4(1):38–42, March 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185865.

Klintworth:2021:IMD

[KvE21] Sandra Klintworth and Eric von Elert. Inducible morphological defense in Daphnia pulex: food quantity effects revised. Aquatic Ecology, 55(1):47–57, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09809-2.

Kaemingk:2012:MAE

[KW12] Mark A. Kaemingk and David W. Willis. Mensurative approach to examine potential interactions between age-0 yellow perch (Perca flavescens) and bluegill (Lepomis macrochirus). Aquatic Ecology, 46(3):353–362, September 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-012-9406-z.

Koop:2011:PIF

[KWO11] Jochen H. E. Koop, Carola Winkelmann, and Christian Ortmann. Physiological indicators of fitness in benthic invertebrates: a useful measure for ecological health assessment and experimental ecology. Aquatic Ecology, 45(4):547–559, November 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-011-9375-7.

Kalacheva:2002:LHC

[KZV02] Galina S. Kalacheva, Natalia O. Zhila, and Tatiana G. Volova. Lipid and hydrocarbon compositions of a collection strain and a wild sample of the green microalga Botryococcus. Aquatic Ecology, 36(2):317–331, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1015615618420.

LopesDaCunha:1999:OVD

[LA99] Pedro Lopes Da Cunha and Maria Margarida Antunes. Occurrence of vertebral deformities in Gobiidae (Pisces) from the Tagus estuary. Aquatic Ecology, 33(3):281–285, October 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009986504933.

Little:2019:DRC

[LA19] Chelsea J. Little and Florian Altermatt. Differential resource consumption in leaf litter mixtures by native and non-native amphipods. Aquatic Ecology, 53(2):151–162, June 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09679-3.

Laanbroek:1997:BRM

[Laa97] Riks Laanbroek. Book review: The Microbiological Quality of Water, D. W. Sutcliffe. Aquatic Ecology, 31(3):342–344, September 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009908127171.

Lacayo:1991:PCF

[Lac91] Mauricio Lacayo. Physical and chemical features of Lake Xolotlán (Managua). Aquatic Ecology, 25(2):111–116, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02291240.

LaFerla:1999:MRA

[LAC99] Rosabruna La Ferla, Maurizio Azzaro, and Giuseppina Chiodo. Microplankton respiratory activity and CO_{2} production rates in the Otranto Strait (Mediterranean Sea). Aquatic Ecology, 33(2):157–165, September 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009922403442.

Lahr:1998:EAH

[Lah98] Joost Lahr. An ecological assessment of the hazard of eight insecticides used in Desert Locust control, to invertebrates in temporary ponds in the Sahel. Aquatic Ecology, 32(2): 153–162, September 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009932427938.

Lambeck:1979:IML

[Lam79] R. H. D. Lambeck. Investigations on macrozoobenthos in Lake Grevelingen, with special emphasis on the role of polychaetes (V). Aquatic Ecology, 13(2-3):81, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284732.

Lammens:1989:CCS

[Lam89] E. H. R. R. Lammens. Causes and consequences of the success of bream in Dutch eutrophic lakes. Aquatic Ecology, 23(1):11–18, March 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02286423.

Lampert:1997:ZRC

[Lam97] Winfried Lampert. Zooplankton research: the contribution of limnology to general ecological paradigms. Aquatic Ecology, 31(1):19–27, March 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009943402621.

Lan:2006:DAB

[LAM06] Nguyen Kim Lan, Takashi Asaeda, and Jagath Manatunge. Decomposition of aboveground and belowground organs of wild rice (Zizania latifolia): mass loss and nutrient changes. Aquatic Ecology, 40(1):13–21, March 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9020-4.

Larsen:1992:PPC

[Lar92] Jørgen Erik Larsen. The phenology of Procladius crassinervis (Zett), Procladius signatus (Zett) and Procladius choreus (Mg) (Diptera, Chironomidae) at Lake Hald, Denmark. Aquatic Ecology, 26(2-4):293–295, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255254.

Lopatina:2021:TCR

[LAZ21] Tatiana Lopatina, Olesya Anishchenko, and Egor Zadereev. Threshold concentrations of the road salt for adverse effects on females and resting eggs of cladoceran Moina macrocopa. Aquatic Ecology, 55(1):283–297, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09830-z.

Leberfinger:2010:GMA

[LB10] Karolina Leberfinger and Irene Bohman. Grass, mosses, algae, or leaves? Food preference among shredders from opencanopy streams. Aquatic Ecology, 44(1):195–203, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9268-1.

Lamand:2014:PSH

[LB14] Florent Lamand and Jean-Nicolas Beisel. Proposal for a simple hydromorphological habitat survey method for freshwater bivalve (Unionidae) inventories. Aquatic Ecology, 48(2):237– 245, June 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9479-y.

Lent:2020:EHP

[LB20] Emily May Lent and Kimberly J. Babbitt. The effects of hydroperiod and predator density on growth, development, and morphology of wood frogs (Rana sylvatica). Aquatic Ecology, 54(1):369–386, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09748-y.

Lopez-Blanco:2018:MST

[LBKV18] Charo López-Blanco, William F. Kenney, and Andrés Varas. Multiple stressors trigger ecological changes in tropical Lake La Tembladera (Ecuador). Aquatic Ecology, 52(2-3):211–224, September 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-018-9656-5.

Liu:2017:IST

[LBY17] Lin Liu, Xiang-Qi Bu, and Fei-Hai Yu. Impacts of sediment type on the performance and composition of submerged macrophyte communities. Aquatic Ecology, 51(1):167–176, March 2017. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9607-y.

deBussy:1970:OUB

[lCdB70] I. J. le Cosquino de Bussy. Onderzoek over ureumderivaten als bestrijdingsmiddelen van algen in onoverdekte zwembaden in Nederland. (Dutch) [Research on urea derivatives as pesticides for algae in outdoor swimming pools in The Netherlands]. Aquatic Ecology, 4(1):14–28, March 1970. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185862.

Lacayo:1991:MCL

[LCF91] M. Lacayo, A. Cruz, and I. Fomsgaard. Mercury contamination in Lake Xolotlán (Managua). Aquatic Ecology, 25 (2):173–176, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02291251.

Liu:2016:RDE

[LCL16] Junchu Liu, Xiuwen Chen, and Chunhua Liu. Response differences of Eichhornia crassipes to shallow submergence and drawdown with an experimental warming in winter. Aquatic Ecology, 50(2):307–314, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9579-y.

Landaeta:2015:GCL

[LCPM15] Mauricio F. Landaeta, Jorge E. Contreras, and Alejandro Pérez-Matus. Growth and condition of larval rockfish in a Patagonian fjord-type inlet: role of hydrographic conditions and food availability. Aquatic Ecology, 49(4):573–584, December 2015. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-015-9547-y.

Lafont:2001:PCE

[LCS01] Michel Lafont, Jean-Claude Camus, and Eric Sourp. A practical concept for the ecological assessment of aquatic ecosystems: application on the River Dore in France. Aquatic Ecology, 35(2):195–205, June 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1011413806318.

Lindeboom:1982:PDE

[LDDS82] H. J. Lindeboom, H. A. J. De Klerk, V. D. Driessche, and A. J. J. Sandee. Production and decomposition of eelgrass (Zostera marina L.) in saline Lake Grevelingen. Aquatic Ecology, 16(1):93–102, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255417.

Larson:2020:IPA

[LDK20] Michele D. Larson, Janet C. Dewey, and Amy C. Krist. Invasive Potamopyrgus antipodarum (New Zealand mud snails) and native snails differ in sensitivity to specific electrical conductivity and cations. Aquatic Ecology, 54(1):103–117, March 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09729-w.

Lear:2011:CBC

[LDL11] Gavin Lear, Andrew Dopheide, and Gillian D. Lewis. A comparison of bacterial, ciliate and macroinvertebrate indicators of stream ecological health. Aquatic Ecology, 45(4):517–527, November 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-011-9372-x.

Lindsay:1993:IDG

[LDW93] P. Lindsay, E. J. Darbyshire, and J. R. West. The influence of density gradients on the transport of solutes in a deep estuary. Aquatic Ecology, 27(2-4):77–86, June 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334771.

Leermakers:1993:GMI

[LEB93] M. Leermakers, M. Elskens, and W. Baeyens. Geochemistry of mercury in an intertidal flat of the Scheldt estuary. Aquatic Ecology, 27(2-4):267–277, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02334790.

Leentvaar:1975:GDB

[Lee75] P. Leentvaar. Geographical distribution and biology of Dreissena polymorpha Pallas. Aquatic Ecology, 9(3):120–122, December 1975. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02263330.

Leentvaar:1980:ENM

[Lee80] P. Leentvaar. Eutrofhication, nature management and the role of potassium. Aquatic Ecology, 14(1-2):22–29, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260269.

Leenen:1982:WID

[Lee82] J. D. Leenen. Wind induced diffusion in a shallow lake, a case study. Aquatic Ecology, 16(2-3):231–240, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255377.

Leentvaar:1984:CAF

[Lee84] P. Leentvaar. Comparative abundance-frequency assessment exampled by data from the Bergumermeer, a lake in the Netherlands. Aquatic Ecology, 18(1):76–77, June 1984. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256685.

Lehtonen:2004:SVP

[Leh04] Kari K. Lehtonen. Seasonal variations in the physiological condition of the benthic amphipods Monoporeia affinis and Pontoporeia femorata in the Gulf of Riga (Baltic Sea). Aquatic Ecology, 38(3):441–456, September 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000035165.97619.78.

Lingeman:1975:DAO

[LFR75] Robert Lingeman, Ben J. G. Flik, and J. Ringelberg. Diel and annual oxygen regimes in a small lake. Aquatic Ecology, 9(1):35–44, April 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257517.

Leon-Galvan:2009:MIN

[LGCS09] Ma. Fabiola León-Galván, Noel Carbajal, and Leticia Santos. Microbial identification of the Nichupte-Bojorquez coastal lagoon in Cancun, Mexico. Aquatic Ecology, 43(2):197–205, June 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9171-1.

Laffoley:1993:CBE

[LH93] Daniel Laffoley and Keith Hiscock. The classification of benthic estuarine communities for nature conservation assessments in Great Britain. Aquatic Ecology, 27(2-4):181–187, June 1993. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334781.

Lingeman:1987:TSP

[LHV87] R. Lingeman, F. Heinis, and A. Veen. Time series of physical, chemical and plankton parameters in Lake Maarsseveen I: 1980–1986. Aquatic Ecology, 21(1):25–38, July 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255452.

Lehtiniemi:2007:PSL

[LHV07] Maiju Lehtiniemi, Tomi Hakala, and Markku Viitasalo. Prey selection by the larvae of three species of littoral fishes on natural zooplankton assemblages. Aquatic Ecology, 41(1):85– 94, March 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-006-9042-6.

Lijklema:1980:ERS

[Lij80] L. Lijklema. Eutrophication: The role of sediments. Aquatic Ecology, 14(1-2):98–105, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260278.

Lijklema:1986:PAS

[Lij86] Lambertus Lijklema. Phosphorus accumulation in sediments and internal loading. Aquatic Ecology, 20(1-2):213–224, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291164.

Lijklema:1991:RLR

[Lij91] Lambertus Lijklema. Response of lakes to the reduction of phosphorus load. Aquatic Ecology, 24(2):165–170, April 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260434.

Lingenman:1972:ETV

[Lin72] Rob Lingenman. Extinctie en temperature in viswater en Bizonbaai. (Dutch) [Extinction and temperature in fishing water and Bison Bay]. Aquatic Ecology, 6(2):62–68, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02304096.

Lingeman:1978:AED

[Lin78] Robert Lingeman. Abundancy estimations of a Daphnia magna population in a small artificial pond. Aquatic Ecology, 12(1):30–38, April 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260791.

Linden:2007:MMS

[Lin07] Eveliina Lindén. The more the merrier: Swarming as an antipredator strategy in the mysid Neomysis integer. Aquatic Ecology, 41(2):299–307, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9055-1.

Liboriussen:2005:PMI

[LJL05] Lone Liboriussen, Erik Jeppesen, and Majbritt F. Lassen. Periphyton-macroinvertebrate interactions in light and fish manipulated enclosures in a clear and a turbid shallow lake. Aquatic Ecology, 39(1):23–39, March 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-3039-9.

Liebezeit:1994:BCC

[LKE94] Gerd Liebezeit, Thomas Kraul, and Beate Everts. Bulk chemical characterization of particulate material from the Jade Bay, Lower Saxonian Wadden Sea. Aquatic Ecology, 28 (3-4):365–370, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334206.

Lee:2001:ERB

[LKM01] Choon Weng Lee, Isao Kudo, and Yoshiaki Maita. Examining the relationship between bacteria and heterotrophic nanoflagellates in Funka Bay (Japan) using the sizefractionation method. Aquatic Ecology, 35(1):1–8, March 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011417918888.

Lengyel:2015:PCB

[LKSK15] Edina Lengyel, Attila W. Kovács, and Csilla Stenger-Kovács. Photosynthetic characteristics of the benthic diatom species Nitzschia frustulum (Kützing) Grunow isolated from a soda pan along temperature-, sulfate- and chloride gradients. Aquatic Ecology, 49(4):401–416, December 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9533-4.

Luykx:1977:VLR

[LL77] R. W. Luykx and I. Laval. Vaucheria lii Rieth, a new alga for the Netherlands. Aquatic Ecology, 11(1):3–6, April 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02282202.

Lefeuvre:1999:DFC

[LLF99] Jean-Claude Lefeuvre, Pascal Laffaille, and Eric Feunteun. Do fish communities function as biotic vectors of organic matter between salt marshes and marine coastal waters? Aquatic Ecology, 33(3):293–299, October 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009956605842.

Li:2010:CGS

[LLG10] Kuanyi Li, Zhengwen Liu, and Binhe Gu. Compensatory growth of a submerged macrophyte (Vallisneria spiralis) in response to partial leaf removal: effects of sediment nutrient levels. Aquatic Ecology, 44(4):701–707, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9308-x.

Lima:2022:SOC

[LLL22] Françoise D. Lima, Tatiana S. Leite, and Sergio M. Q. Lima. Seamounts and oceanic currents drive the population structure of Octopus insularis in the Southwest Tropical Atlantic. Aquatic Ecology, 56(4):1143–1155, December 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-022-09955-9.

Li:2023:MCD

[LLL^{+}23] Junnan Li, Guangbin Liu, Qian Liu, Fang Wang, Hongwei Shan, Yicheng Xie, and Changjian Li. Microbial community dynamics and its correlation with environmental factors in the water of polyculture ponds containing Penaeus japonicus, Portunus trituberculatus and Sinonovacula constricta. Aquatic Ecology, 57(2):263–279, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10009-x.

Li:2011:PDC

[LLX11] Feng Li, Youzhi Li, and Yonghong Xie. Plant distribution can be reflected by the different growth and morphological responses to water level and shade in two emergent macrophyte seedlings in the Sanjiang Plain. Aquatic Ecology, 45(1):89– 97, March 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-010-9334-8.

Lindegaard:1992:APD

[LM92] Claus Lindegaard and Peter Mæhl. Abundance, population dynamics and production of Chironomidae (Diptera) in an ultraoligotrophic lake in South Greenland. Aquatic Ecology, 26(2-4):297–308, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255255.

Layman:2010:LFC

[LMA10] Craig A. Layman, Carmen G. Montaña, and Jacob E. Allgeier. Linking fish colonization rates and water level change in littoral habitats of a Venezuelan floodplain river. Aquatic Ecology, 44(1):269–273, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9256-5.

Lemoine:2012:AAM

[LMBM12] Damien G. Lemoine, Florian Mermillod-Blondin, and Emmanuel Malet. The ability of aquatic macrophytes to increase root porosity and radial oxygen loss determines their resistance to sediment anoxia. Aquatic Ecology, 46(2):191– 200, June 2012. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-012-9391-2.

Lucena-Moya:2011:MAA

[LMD11] Paloma Lucena-Moya and Ian C. Duggan. Macrophyte architecture affects the abundance and diversity of littoral microfauna. Aquatic Ecology, 45(2):279–287, May 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9353-0.

Lehman:2010:FTW

[LME10] P. W. Lehman, S. Mayr, and C. Enright. The freshwater tidal wetland Liberty Island, CA was both a source and sink of inorganic and organic material to the San Francisco Estuary. Aquatic Ecology, 44(2):359–372, June 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9295-y.

Litulo:2005:PBB

[LMM05] Carlos Litulo, Yolanda Mahanjane, and Fernando L. M. Mantelatto. Population biology and breeding period of the sand-bubbler crab Dotilla fenestrata (Brachyura: Ocypodidae) from Southern Mozambique. Aquatic Ecology, 39(3): 305–313, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-3443-9.

Lobova:2002:GSD

[LMP02] T. I. Lobova, E. Ye. Maksimova, and N. S. Pechurkin. Geographical and seasonal distribution of multiple antibiotic resistance of heterotrophic bacteria of Lake Shira. Aquatic Ecology, 36(2):299–307, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1015618820450.

Lappalainen:2009:EWT

[LMV09] Jyrki Lappalainen, Marco Milardi, and Ari Venäläinen. Effects of water temperature on year-class strengths and growth patterns of pikeperch (Sander lucioperca (L.)) in the brackish Baltic Sea. Aquatic Ecology, 43(1):181–191, March 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9150-y.

Liljendahl-Nurminen:2008:SVA

[LNHN08] Anne Liljendahl-Nurminen, Jukka Horppila, and Juha Niemistö. Spatial variability in the abundance of pelagic invertebrate predators in relation to depth and turbidity. Aquatic Ecology, 42(1):25–33, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9070-2.

Los:1980:AAB

[Los80] F. J. Los. Application of an algal bloom model (BLOOM II) to combat eutrophication. Aquatic Ecology, 14(1-2):116–124, August 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260280.

Lovric:1974:ZSB

[Lov74] Andrija Zelimir Lovrić. Zonation and succession in brackish environments of the Eastern Adriatic. Aquatic Ecology, 8 (1-2):166–171, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02254916.

Lopatina:2024:SSE

[LOZ24] Tatiana Lopatina, Natalia Oskina, and Egor Zadereev. Salinity sensitivity of Moina macrocopa post-diapause females hatched from resting eggs of different ages. Aquatic Ecology, 58(3):865–877, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
024-10110-9.

Leitao:2007:STV

[LPB07] S. Leitão, P. Pinto, and M. F. Brito. Spatial and temporal variability of macroinvertebrate communities in two farmed Mediterranean rice fields. Aquatic Ecology, 41(3):373–386, September 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9082-6.

Leoni:2021:CPD

[LPT21] Barbara Leoni, Martina Patelli, and Monica Tolotti. Cladocera paleocommunity to disentangle the impact of anthropogenic and climatic stressors on a deep subalpine lake ecosystem (Lake Iseo, Italy). Aquatic Ecology, 55(2):607– 621, June 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09850-9.

Labat:2022:ISS

[LPT22] Frédéric Labat, Christophe Piscart, and Gabrielle Thiébaut. Invertebrates in small shallow lakes and ponds: a new sampling method to study the influence of environmental factors on their communities. Aquatic Ecology, 56(3):585–603, September 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09939-1.

Laville:1992:CFM

[LR92] H. Laville and F. Reiss. The chironomid fauna of the Mediterranean region reviewed. Aquatic Ecology, 26(2-4):239–245, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255247.

Leber:2021:CSC

[LRG21] Christopher A. Leber, Andres Joshua Reyes, and William H. Gerwick. Cyanobacteria-shrimp colonies in the Mariana Islands. Aquatic Ecology, 55(2):453–465, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09837-6.

Lopez-Rodriguez:2019:ESH

[LRMdF19] Manuel Jesús López-Rodríguez, Carolina Márquez Muñoz, and José Manuel Tierno de Figueroa. Effect of shifts in habitats and flow regime associated to water diversion for agriculture on the macroinvertebrate community of a small watershed. Aquatic Ecology, 53(3):483–495, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09703-6.

Lopez-Rocha:2022:AEM

[LRSN22] Ana Nayeli López-Rocha, S. S. S. Sarma, and S. Nandini. Allelopathic effects of male and female calanoids and cyclopoids (Copepoda) on the demographic response of Brachionus havanaensis (Rotifera). Aquatic Ecology, 56(2):399–407, June 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09937-3.

Lucio:2024:GDE

[LSdS^{+}24] Léia Carolina Lucio, Bárbara Scorsim, Adrian Cesar da Silva, Alessandra Valéria de Oliveira, Sidinei Magela Thomaz, Sônia Maria Alves Pinto Prioli, and Alberto José Prioli. Genetic diversity of Egeria densa Planch. (Hydrocharitaceae) in Brazilian reservoirs. Aquatic Ecology, 58(3):707–716, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10100-x.

Li:2010:CTI

[LSO10] Yunkai Li, Bing Song, and Derek Olson. Changes in the trophic interactions and the community structure of Lake Taihu (China) ecosystem from the 1960s to 1990s. Aquatic Ecology, 44(2):337–348, June 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9293-0.

Lehman:2008:IFH

[LSR08] Peggy W. Lehman, Ted Sommer, and Linda Rivard. The influence of floodplain habitat on the quantity and quality of riverine phytoplankton carbon produced during the flood season in San Francisco Estuary. Aquatic Ecology, 42(3): 363–378, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9102-6.

Liang:2020:CLC

[LSY20] Ye Liang, Li Shao, and Jiaxin Yang. Changes in the lifecycle parameters and glutathione-related antioxidant defense system of rotifer Brachionus calyciflorus under the combined stress of microcystin-LR and ammonia. Aquatic Ecology, 54 (1):243–256, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09739-8.

Liu:2022:IAP

[LTH22] Wei Liu, Qiyang Tan, and Feng He. An integrated analysis of pond ecosystem around Poyang Lake: assessment of water quality, sediment geochemistry, phytoplankton and benthic macroinvertebrates diversity and habitat condition. Aquatic Ecology, 56(3):775–791, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09931-9.

Lear:2009:ELR

[LTL09] Gavin Lear, Susan J. Turner, and Gillian D. Lewis. Effect of light regimes on the utilisation of an exogenous carbon source by freshwater biofilm bacterial communities. Aquatic Ecology, 43(2):207–220, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9193-8.

Lu:2005:SVM

[Lu05] Lin Lu. Seasonal variation of macrobenthic infauna in the Johor Strait, Singapore. Aquatic Ecology, 39(1):107–111, March 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-7111-2.

Li:1993:TVI

[LV93] Jian Li and Magda Vincx. The temporal variation of intertidal nematodes in the Westerschelde I. The importance of an estuarine gradient. Aquatic Ecology, 27(2-4):319–326, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334795.

Loebl:2006:SNS

[LvBR06] Martina Loebl, Justus E. E. van Beusekom, and Karsten Reise. Is spread of the neophyte Spartina anglica recently enhanced by increasing temperatures? Aquatic Ecology, 40(3): 315–324, September 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-006-9029-3.

Li:2011:ILN

[LW11] Weiguo Li and Jianbo Wang. Influence of light and nitrate assimilation on the growth strategy in clonal weed Eichhornia crassipes. Aquatic Ecology, 45(1):1–9, March 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9318-8.

Lurling:2016:ESE

[LWdSD16] Miquel Lürling, Guido Waajen, and Lisette N. de Senerpont Domis. Evaluation of several end-of-pipe measures proposed to control cyanobacteria. Aquatic Ecology, 50(3): 499–519, September 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9563-y.

Leewis:1989:AMA

[LWM89] R. J. Leewis, H. W. Waardenburg, and A. J. M. Meijer. Active management of an artificial rocky coast. Aquatic Ecology, 23(1):91–99, March 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02286431.

Lehane:2001:ISS

[LWO01] B. M. Lehane, B. Walsh, and J. O’Halloran. The influence of small-scale variation in habitat on winter trout distribution and diet in an afforested catchment. Aquatic Ecology, 35(1):61–71, March 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1011467711628.

Li:2025:EED

[LWW^{+}25] Fangze Li, Yu Wang, Xiaohuan Wang, Hui Jiao, Lulu Ran, and Ping Lu. Enantioselective effects of dinotefuran on physiological parameters in Rana nigromaculata tadpoles. Aquatic Ecology, 59(1):95–106, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10150-1.

Lv:2019:IEC

[LWX19] Chaochao Lv, Hui Wang, and Chengxing Xia. The influence of elevated CO_{2} on Vallisneria spiralis, Radix auricularia, and their plant–herbivore interaction. Aquatic Ecology, 53 (1):137–150, March 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09678-4.

Li:2024:CIN

[LWY^{+}24] Dan Li, Xiao-Gai Wang, Jia-Xin Yang, Yu-Fei Hu, Wen-Yue Lyu, Jun-Cai Xin, Rui Zhang, Zhi-Huan Chen, and Chao Si. Cadmium interacts with nitrogen level and nitrogen form to affect the growth of an invasive plant Alternanthera philoxeroides. Aquatic Ecology, 58(2):289–297, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10065-3.

Liu:2021:DMM

[LXH21] Ping Liu, Shao-Lin Xu, and Bo-Ping Han. Development of microsatellite markers for Diaphanosoma dubium (Crustacea, Cladocera) and application to seasonal population dynamics. Aquatic Ecology, 55(4):1189–1206, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09812-7.

Liang:2021:IPM

[LYM21] Ye Liang, Xiaotong Yang, and Longjiang Mao. Influence of polystyrene microplastics on rotifer (Brachionus calyciflorus) growth, reproduction, and antioxidant responses. Aquatic Ecology, 55(3):1097–1111, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09885-y.

Li:2016:FGP

[LZD16] Feng Li, Lianlian Zhu, and Zhengmiao Deng. Fragment growth performance of the invasive submerged macrophyte Myriophyllum spicatum under conditions of different water depths and sediment types. Aquatic Ecology, 50(4):727–734, December 2016. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9589-9.

Lamia:2025:ADM

[LZH^{+}25] Seridi Lamia, Guellati Fatma Zohra, Touati Hassen, Kadri Skander, and Bensouilah Mourad. Advanced description of the morphology of Cylindrospermopsis raciborskii in a shallow mediterranean lake. Aquatic Ecology, 59(1):247–262, March 2025. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10160-z.

Li:2022:APT

[LZL22] Lanxin Li, Wenwen Zhang, and Wenyan Liang. Applications of phytoremediation to treat reclaimed water in urban parks using aquatic macrophytes. Aquatic Ecology, 56(1):75–88, March 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09894-x.

Li:2014:SDG

[LZN14] Ming Li, Wei Zhu, and Philip Nti Nkrumah. Sizedependent growth of Microcystis colonies in a shallow, hypertrophic lake: use of the RNA-to-total organic carbon ratio. Aquatic Ecology, 48(2):207–217, June 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9476-1.

Liu:2021:EFP

[LZY21] Yang Liu, Min Zhang, and Chenyu Yang. The effects of flood pulse on multiple aquatic organisms in a seasonal shallow lake. Aquatic Ecology, 55(2):379–399, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09829-y.

Li:2025:IMN

[LZZ^{+}25] Youwen Li, Xinxin Zhao, Zhaoqi Zhong, Fanfan Ju, Hui Liu, Ning Lin, and Xianghao Zha. Interaction of microalgae, nitrogen and arsenic in a water environment co-polluted by nitrogen and arsenic: current knowledge and future perspectives. Aquatic Ecology, 59(1):323–337, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10165-8.

M:1975:RDB

[M.75] Coesel P. F. M. The relevance of desmids in the biological typology and evaluation of fresh waters. Aquatic Ecology, 9(2):91, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257645.

Mathieson:1995:RNE

[MA95] Scot Mathieson and Stephen M. Atkins. A review of nutrient enrichment in the estuaries of Scotland: Implications for the natural heritage. Aquatic Ecology, 29(3-4):437–448, December 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02084242.

Mghili:2022:MSH

[MAA22] Bilal Mghili, Mohamed Analla, and Mustapha Aksissou. Medusae (Scyphozoa and Hydrozoa) from the Moroccan Mediterranean coast: abundance and spatiotemporal dynamics and their economic impact. Aquatic Ecology, 56 (1):213–226, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09910-0.

Moreno:2010:SDM

[MAD10] Jose L. Moreno, David G. Angeler, and Jorge De las Heras. Seasonal dynamics of macroinvertebrate communities in a semiarid saline spring stream with contrasting environmental conditions. Aquatic Ecology, 44(1):177–193, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9251-x.

Magoulick:2000:STV

[Mag00] Daniel D. Magoulick. Spatial and temporal variation in fish assemblages of drying stream pools: The role of abiotic and biotic factors. Aquatic Ecology, 34(1):29–41, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009914619061.

Milano:2013:DPS

[MAM13] Daniela Milano, Juana C. Aigo, and Patricio J. Macchi. Diel patterns in space use, food and metabolic activity of Galaxias maculatus (Pisces: Galaxiidae) in the littoral zone of a shallow Patagonian lake. Aquatic Ecology, 47(3):277–290, September 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9443-2.

Marchand:1993:ISS

[Mar93] Jocelyne Marchand. The influence of seasonal salinity and turbidity maximum variations on the nursery function of the Loire estuary (France). Aquatic Ecology, 27(2-4):427–436, June 1993. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334804.

Marten:2001:EMB

[Mar01] Michael Marten. Environmental monitoring in Baden– Württemberg with special reference to biocoenotic trendmonitoring of macrozoobenthos in rivers and methodical requirements for evaluation of long-term biocoenotic changes. Aquatic Ecology, 35(2):159–171, June 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011447307618.

Martin:2017:EMS

[Mar17] Charles W. Martin. Effects of macrophyte-specific olfactory cues on fish preference patterns. Aquatic Ecology, 51 (1):159–165, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9606-z.

Mastrantuono:1985:LSZ

[Mas85] Luciana Mastrantuono. Littoral sand zoobenthos and its relation to organic pollution in Lake Nemi (Central Italy). Aquatic Ecology, 19(2):171–178, December 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270763.

Mastrantuono:1988:NFA

[Mas88] Luciana Mastrantuono. A note on the feeding of Amphichaeta leydigii (Oligochaeta, Naididae) in lacustrine sandy shores. Aquatic Ecology, 22(2):195–198, December 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02256823.

Mahmoudi:2014:EVT

[MAS14] Nemat Mahmoudi, Mohammad Reza Ahmadi, and Jafar Seyfabadi. Environmental variables and their interaction effects on chlorophyll-a in coastal waters of the southern Caspian Sea: assessment by multiple regression grey models. Aquatic Ecology, 48(3):351–365, September 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9489-9.

Meulemans:1981:RCC

[MB81] John T. Meulemans and Elly P. H. Best. Research into the characteristics of C_{3} and C_{4} plants in Ceratophyllum demersum L. Aquatic Ecology, 15(3):199, December 1981. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255183.

Mori:2006:MCK

[MB06] Nataša Mori and Anton Brancelj. Macroinvertebrate communities of karst springs of two river catchments in the Southern Limestone Alps (the Julian Alps, NW Slovenia). Aquatic Ecology, 40(1):69–83, March 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9018-y.

McCreadie:2014:SCL

[MB14] John W. McCreadie and Chris Bedwell. Species composition of local riffle beetle (Coleoptera: Elmidae) assemblages in small coastal streams of the Gulf of Mexico: the influences of local and regional factors. Aquatic Ecology, 48(2):127– 141, June 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9471-6.

Moreau:2024:STA

[MBFD24] Axelle Moreau, Pierrick Bocher, Sébastien Farau, and Christine Dupuy. Spatio-temporal assessment of the relationship between freshwater invertebrates and environmental characteristics in Vendée wetlands, France. Aquatic Ecology, 58(3): 675–687, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10098-2.

Meleg:2015:CCC

[MBM15] Ioana N. Meleg, Karina P. Battes, and Oana T. Moldovan. Contrasting copepod community dynamics related to sampling strategies in the unsaturated zone of a karst aquifer. Aquatic Ecology, 49(4):549–560, December 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9545-0.

McGuinness:2022:ASP

[MBM^{+}22] Maria McGuinness, Hannah Brownlow, Rob McAllen, Luke Harman, Damien Haberlin, and Thomas K. Doyle. Abundance and seasonality of phoronid larvae in coastal temperate waters: More abundant than previously thought? Aquatic Ecology, 56(4):1315–1321, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09982-6.

Mermillod-Blondin:2020:BPC

[MBMV20] Florian Mermillod-Blondin, Pierre Marmonier, and Laurence Volatier. Bottom-up processes control benthic macroinvertebrate communities and food web structure of fishless artificial wetlands. Aquatic Ecology, 54(2):575–589, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09760-2.

Monteiro:2022:CET

[MCB22] N. Monteiro, J. Cuaresma, and A. Barreiro. Competitive exclusion of toxic cyanobacterial species by an allelopathic strain of Phormidium. Aquatic Ecology, 56(2):459–474, June 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09901-1.

Martins:2021:AII

[MCC21] Isabela Martins, Diego M. P. Castro, and Marcos Callisto. Anthropogenic impacts influence the functional traits of Chironomidae (Diptera) assemblages in a neotropical savanna river basin. Aquatic Ecology, 55(3):1081–1095, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09884-z.

McLusky:1993:MEG

[McL93] Donald S. McLusky. Marine and estuarine gradients — an overview. Aquatic Ecology, 27(2-4):489–493, June 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334809.

Munoz-Colmenares:2021:CZS

[MCSV21] Manuel E. Muñoz-Colmenares, Juan M. Soria, and Eduardo Vicente. Can zooplankton species be used as indicators of trophic status and ecological potential of reservoirs? Aquatic Ecology, 55(4):1143–1156, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09897-8.

Mendoza-Carranza:2008:WCM

[MCV08] Manuel Mendoza-Carranza and João Vieira. Whitemouth croaker Micropogonias furnieri (Desmarest, 1823) feeding strategies across four southern Brazilian estuaries. Aquatic Ecology, 42(1):83–93, March 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
007-9084-4.

Martin-Creuzburg:2009:GFV

[MCvE09] Dominik Martin-Creuzburg and Eric von Elert. Good food versus bad food: the role of sterols and polyunsaturated fatty acids in determining growth and reproduction of Daphnia magna. Aquatic Ecology, 43(4):943–950, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9239-6.

Mukai:1992:ACI

[MD92] T. Mukai and H. De Haan. Adsorptive characteristics of iron(III) onto hydrous manganese dioxide in model lake water under acidic conditions. Aquatic Ecology, 25(3):183–189, August 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02270804.

Mkandawire:2005:ALG

[MD05] Martin Mkandawire and E. Gert Dudel. Assignment of Lemna gibba L. (duckweed) bioassay for in situ ecotoxicity assessment. Aquatic Ecology, 39(2):151–165, June 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-5411-1.

Mur:1991:R

[MDB91] Luuc Mur, Jaap Dorgelo, and C. Bakker. Reviews. Aquatic Ecology, 24(2):171–185, April 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260435.

Mees:1993:SCS

[MDH93] Jan Mees, Ann Dewicke, and Olivier Hamerlynck. Seasonal composition and spatial distribution of hyperbenthic communities along estuarine gradients in the Westerschelde. Aquatic Ecology, 27(2-4):359–376, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02334798.

Monteiro:2021:SMS

[MdMdLB21] Flávia Morgana Monteiro, Gustavo Correia de Moura, and José Etham de Lucena Barbosa. Submerged macrophytes support cyanobacteria and microcystin production in a drawdown tropical semi-arid reservoir. Aquatic Ecology, 55(3): 875–890, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09866-1.

Moe:2008:RDE

[MDP08] S. Jannicke Moe, Bernard Dudley, and Robert Ptacnik. RE-BECCA databases: experiences from compilation and analyses of monitoring data from 5,000 lakes in 20 European countries. Aquatic Ecology, 42(2):183–201, June 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9190-y.

Muylaert:2003:ZPM

[MDV03] Koenraad Muylaert, Steven Declerck, and Wim Vyverman. Zooplankton, phytoplankton and the microbial food web in two turbid and two clearwater shallow lakes in Belgium. Aquatic Ecology, 37(2):137–150, April 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1023988702926.

Mur:1984:ILQ

[ME84] L. R. Mur and R. P. Elema. The influence of light quality on the growth of some phytoplankton species. Aquatic Ecology, 18(1):73–74, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256682.

Monteny:1993:BCZ

[MEB93] F. Monteny, M. Elskens, and W. Baeyens. The behaviour of copper and zinc in the Scheldt estuary. Aquatic Ecology, 27 (2-4):279–286, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334791.

Merks:1979:MSS

[Mer79] A. G. A. Merks. A microgradient sampler for shallow waters. Aquatic Ecology, 13(1):61–67, April 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260310.

Mettam:1978:EET

[Met78] C. Mettam. Environmental effects of tidal power generating schemes. Aquatic Ecology, 12(3-4):307–321, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02259192.

Meulemans:1991:GDE

[Meu91] John T. Meulemans. General description of the execution of the ‘Project Limnology Applied to Lake Managua’. Aquatic Ecology, 25(2):103–104, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02291238.

Michailova:1992:KAF

[MF92] Paraskeva Vl. Michailova and L. C. Ferrington, Jr. The karyotype of Axarus festivus (Say) and a comparison with species of Lipiniella (Chironomidae: Diptera). Aquatic Ecology, 26 (2-4):181–185, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255239.

ManziMarinho:2007:IPR

[MF07] Marcelo Manzi Marinho and Sandra Maria Feliciano de Oliveira e Azevedo. Influence of N/P ratio on competitive abilities for nitrogen and phosphorus by Microcystis aeruginosa and Aulacoseira distans. Aquatic Ecology, 41(4): 525–533, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9118-y.

Millie:2009:LSP

[MFL09] David F. Millie, Gary L. Fahnenstiel, and R. Wayne Litaker. Late-summer phytoplankton in western Lake Erie (Laurentian Great Lakes): bloom distributions, toxicity, and environmental influences. Aquatic Ecology, 43(4):915–934, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9238-7.

Mouabad:2001:PBF

[MFP01] A. Mouabad, M. Ait Fdil, and J. C. Pihan. Pumping behaviour and filtration rate of the freshwater mussel Potomida littoralis as a tool for rapid detection of water contamination. Aquatic Ecology, 35(1):51–60, March 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011499325045.

Mees:1995:AIN

[MFS95] Jan Mees, Nancy Fockedey, and Jean-Claude Sorbe. Aberrant individuals of Neomysis integer and other Mysidacea: Intersexuality and variable telson morphology. Aquatic Ecology, 29(2):161–166, June 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02336046.

Millie:2003:PBC

[MFS03] David F. Millie, Gary L. Fahnenstiel, and Oscar M. E. Schofield. Physical-biological coupling in Southern Lake Michigan: Influence of episodic sediment resuspension on phytoplankton. Aquatic Ecology, 37(4):393–408, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000007046.48955.70.

Mendez-Fernandez:2020:PSA

[MFTM20] Paula Méndez-Fernandez, Satie Taniguchi, and Rosalinda C. Montone. Population structure of the Atlantic spotted dolphin (Stenella frontalis) inferred through ecological markers. Aquatic Ecology, 54(1):21–34, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09722-3.

Mangas:1991:SFZ

[MG91a] Emma Mangas and Helen Garcia. Seasonal fluctuations of zooplankton biomass in Lake Xolotlán (Managua). Aquatic Ecology, 25(2):157–162, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02291248.

Montenegro-Guillen:1991:LPL

[MG91b] Salvador Montenegro-Guillen. Limnological perspective of Lake Xolotlán (Managua): The PLALM. Aquatic Ecology, 25(2):105–109, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02291239.

Michael:2006:DFC

[MG06] Friedrichs Michael and Graf Gerhard. Description of a flume channel profilometry tool using laser line scans. Aquatic Ecology, 40(4):493–501, December 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-8057-0.

Milardi:2019:MFI

[MGC19] Marco Milardi, Anna Gavioli, and Giuseppe Castaldelli. Meteorological factors influence marine and resident fish movements in a brackish lagoon. Aquatic Ecology, 53(2):251–263, June 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09686-4.

Mcgrorty:1993:MME

[MGCC93] S. Mcgrorty, J. D. Goss-Custard, and R. T. Clarke. Mussel Mytilus edulis (Mytilacea) dynamics in relation to environmental gradients and intraspecific interactions. Aquatic Ecology, 27(2-4):163–171, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334779.

Ma:2022:SVB

[MGL22] Ning Ma, Li Gao, and Ming Li. Spatial variation in bacterial community and dissolved organic matter composition in groundwater near a eutrophic lake. Aquatic Ecology, 56(3): 555–571, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09926-6.

Marazzi:2017:PSD

[MGT17] L. Marazzi, E. E. Gaiser, and F. A. C. Tobias. Phosphorus scarcity and desiccation stress increase the occurrence of dominant taxa in wetland benthic primary producer communities. Aquatic Ecology, 51(4):571–589, December 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9637-0.

Montenegro-Guillen:1991:BIL

[MGVC91] S. Montenegro-Guillen, K. Vammen, and R. Cisneros. Biotic interactions in Lake Xolotlán (Managua): an integrating approach. Aquatic Ecology, 25(2):177–179, June 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291252.

Munz:2013:IDP

[MH13] Jeremy T. Munz and Christopher L. Higgins. The influence of discharge, photoperiod, and temperature on the reproductive ecology of cyprinids in the Paluxy River, Texas. Aquatic Ecology, 47(1):67–74, March 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9425-9.

McLusky:1993:VIS

[MHE93] D. S. McLusky, S. C. Hull, and M. Elliott. Variations in the intertidal and subtidal macrofauna and sediments along a salinity gradient in the upper forth estuary. Aquatic Ecology, 27(2-4):101–109, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334773.

Missaghi:2016:IFM

[MHG16] Shahram Missaghi, Miki Hondzo, and Michele Guala. Influence of fluid motion on growth and vertical distribution of cyanobacterium Microcystis aeruginosa. Aquatic Ecology, 50 (4):639–652, December 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9583-2.

Mooij:2005:ICC

[MHL05] Wolf M. Mooij, Stephan Hülsmann, and Eddy H. R. R. Lammens. The impact of climate change on lakes in the Netherlands: a review. Aquatic Ecology, 39(4):381–400, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-9008-0.

Ma:2023:DRE

[MHYL23] Shunrong Ma, Guilin Han, Yiyun Yang, and Xiaoqiang Li. Dissolved rare earth elements distribution and fractionation in a subtropical coastal river: a case study from Jiulong River, Southeast China. Aquatic Ecology, 57(3):765–781, September 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10048-4.

Michailova:1992:ETF

[Mic92] Paraskeva Michailova. Endochironomus tendens (F.) (Chironomidae, Diptera) an example of stasipatric speciation. Aquatic Ecology, 26(2-4):173–180, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255238.

Merino-Ibarra:2008:PCL

[MIMRRZ08] Martin Merino-Ibarra, Emiliano Monroy-Ríos, and Jorge Ramírez-Zierold. Physical and chemical limnology of a windswept tropical highland reservoir. Aquatic Ecology, 42(3): 335–345, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9111-5.

Mitchell:1974:AEL

[Mit74] Roger Mitchell. Aspects of the ecology of the lamellibranch Mercenaria mercenaria (L.) in British waters. Aquatic Ecology, 8(1-2):124–138, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02254913.

Mitchell:1978:NCI

[Mit78] Roger Mitchell. Nature conservation implications of hydraulic engineering schemes affecting British estuaries. Aquatic Ecology, 12(3-4):333–350, December 1978. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02259194.

Moffat:1993:CDM

[MJ93] A. M. Moffat and M. B. Jones. Correlation of the distribution of Mesopodopsis slabberi (Crustacae, Mysidacea) with physico-chemical gradients in a partially-mixed estuary (Tamar, England). Aquatic Ecology, 27(2-4):155–162, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334778.

Martinez-Jeronimo:2005:NRS

[MJEC05] Fernando Martínez-Jerónimo and Félix Espinosa-Ch’z. Notes on the reproduction and survival of Moina hutchinsoni Brehm, 1937 (Moinidae: Anomopoda) grown in media of varying salinity. Aquatic Ecology, 39(1):113–118, March 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-004-3118-y.

Min:2021:CEI

[MJJ21] Cai Min, Liselotte Sander Johansson, and Erik Jeppesen. Copepods as environmental indicator in lakes: special focus on changes in the proportion of calanoids along nutrient and pH gradients. Aquatic Ecology, 55(4):1241–1252, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09877-y.

Matthijs:2016:EEC

[MJM16] Hans C. P. Matthijs, Daniel Jančula, and Blahoslav Maršálek. Existing and emerging cyanocidal compounds: new perspectives for cyanobacterial bloom mitigation. Aquatic Ecology, 50(3):443–460, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9577-0.

Misra:2005:SVP

[MK05a] Sudipta Misra and Rahul Kundu. Seasonal variations in population dynamics of key intertidal molluscs at two contrasting locations. Aquatic Ecology, 39(3):315–324, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-1779-9.

Moustakas:2005:HDA

[MK05b] Aristides Moustakas and Ioannis Karakassis. How diverse is aquatic biodiversity research? Aquatic Ecology, 39(3): 367–375, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-6041-y.

Mohsen:1979:CBF

[MKA79] A. F. Mohsen, A. F. Khaleafa, and Y. Azab. Correlation between free and conjugated amino acids in some marine algae during their vegetative growth and reproductive stages. Aquatic Ecology, 13(2-3):120–122, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284747.

Makhutova:2003:CFA

[MKG03] Olesia N. Makhutova, Galina S. Kalachova, and Michail I. Gladyshev. A comparison of the fatty acid composition of Gammarus lacustris and its food sources from a freshwater reservoir, Bugach, and the saline Lake Shira in Siberia, Russia. Aquatic Ecology, 37(2):159–167, April 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1023975101075.

Masigol:2020:TFD

[MKG20] Hossein Masigol, Seyed Akbar Khodaparast, and Hans-Peter Grossart. Taxonomical and functional diversity of Saprolegniales in Anzali lagoon, Iran. Aquatic Ecology, 54(1): 323–336, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09745-w.

Morad:2010:QPR

[MKL10] M. R. Morad, A. Khalili, and J. Lewandowski. Quantification of pumping rate of Chironomus plumosus larvae in natural burrows. Aquatic Ecology, 44(1):143–153, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9259-2.

Montana:2022:CAW

[MKS22] Carmen G. Montaña, Friedrich W. Keppeler, and Christopher M. Schalk. Community assembly within ponds: the roles of space, time, and environmental gradients. Aquatic Ecology, 56(1):123–142, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
021-09902-0.

Mavromati:2023:DLS

[MKT23] Efpraxia Mavromati, Dimitra Kemitzoglou, and Vasiliki Tsiaoussi. Does littoral substrate affect macroinvertebrate assemblages in Mediterranean lakes? Aquatic Ecology, 57(3): 667–679, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10037-7.

Martins:2015:MLT

[MKV15] Ana Carolina Braga Martins, Paul Gerhard Kinas, and João Paes Vieira. Medium- and long-term temporal trends in the fish assemblage inhabiting a surf zone, analyzed by Bayesian generalized additive models. Aquatic Ecology, 49 (1):57–69, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9504-9.

Marescaux:2016:BMI

[MLB16] Jonathan Marescaux, Adrien Latli, and Jean-Nicolas Beisel. Benthic macro-invertebrate fauna associated with Dreissena mussels in the Meuse River: from incapacitating relationships to facilitation. Aquatic Ecology, 50(1):15–28, March 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9540-5.

Meijer:1995:DFC

[MLG95] M.-L. Meijer, E. H. R. R. Lammens, and M. P. Grimm. Development of fish communities in lakes after biomanipulation. Aquatic Ecology, 29(1):91–101, April 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02061792.

Mergler:2020:ESL

[MLG20] Ciara J. Mergler, Alexandra N. Ludwig, and Brian G. Gall. The effects of satiation level and competition risk on resource acquisition in red swamp crayfish (Procambarus clarkii). Aquatic Ecology, 54(3):889–894, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09775-9.

Ma:2021:SMR

[MLL21] Xiaowen Ma, Yang Li, and Chunhua Liu. Support from mother ramets declines with increasing independence of daughter ramets in submerged clonal Vallisneria natans. Aquatic Ecology, 55(1):299–308, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09831-y.

Matsuda:2015:ABS

[MLTH15] Juliana Tiemi Matsuda, Fábio Amodêo Lansac-Tôha, and Janet Higuti. Association of body size and behavior of freshwater ostracods (Crustacea, Ostracoda) with aquatic macrophytes. Aquatic Ecology, 49(3):321–331, September 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9527-2.

Meira:2018:IHP

[MLTV18] Bianca Ramos Meira, Fernando Miranda Lansac-Toha, and Luiz Felipe Machado Velho. The importance of herbivory by protists in lakes of a tropical floodplain system. Aquatic Ecology, 52(2-3):193–210, September 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9654-7.

Milano:2010:PIR

[MLZ10] Daniela Milano, Mariana Lozada, and Horacio E. Zagarese. Predator-induced reaction patterns of landlocked Galaxias maculatus to visual and chemical cues. Aquatic Ecology, 44 (4):741–748, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-010-9312-1.

Mclusky:1995:NSE

[MM95] D. S. Mclusky and A. D. Mcintyre. Northern and southern estuaries and coastal areas. Aquatic Ecology, 29(3-4):469–471, December 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02084245.

Majecki:1996:POS

[MM96] Janusz Majecki and Katarzyna Majecka. Predation by Oligotricha striata caddis larvae on amphibian eggs: Effects of a high quality food on growth rate. Aquatic Ecology, 30(1):21– 25, July 1996. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02092144.

Meyer-Milne:2021:EMM

[MMBP21] Elizabeth Meyer-Milne, Luc Brendonck, and Tom Pinceel. Egg morphology may underpin the successful distribution of large branchiopods in temporary waters. Aquatic Ecology, 55(1):237–251, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09826-1.

Mataba:2024:EUQ

[MMBV24] Gordian Rocky Mataba, Linus Munishi, Luc Brendonck, and Bram Vanschoenwinkel. Culex quinquefasciatus mosquito avoids ovipositing in habitats with the annual fish (Nothobranchius neumanni) in Tanzania. Aquatic Ecology, 58(3): 775–787, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10104-7.

Muhametsafina:2014:FDF

[MMC14] A. Muhametsafina, J. D. Midwood, and S. J. Cooke. The fate of dead fish tagged with biotelemetry transmitters in an urban stream. Aquatic Ecology, 48(1):23–33, March 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9463-y.

Mosharova:2022:EAE

[MML^{+}22] Irina V. Mosharova, Alexander S. Mikaelyan, Anna V. Lifanchuk, Sergey A. Mosharov, Vladimir A. Silkin, and Alexey V. Fedorov. Effect of anticyclonic eddy on bacterioplankton in the Black Sea: an experimental study. Aquatic Ecology, 56(4):??, November 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
022-09987-1.

Mosharova:2023:EAE

[MML^{+}23] Irina V. Mosharova, Alexander S. Mikaelyan, Anna V. Lifanchuk, Sergey A. Mosharov, Vladimir A. Silkin, and Alexey V. Fedorov. Effect of anticyclonic eddy on bacterioplankton in the Black Sea: an experimental study. Aquatic Ecology, 57(1):1–13, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
022-09987-1.

Maazouzi:2011:SRD

[MMM11] Chafik Maazouzi, Vincent Médoc, and Gérard Masson. Sizerelated dietary changes observed in young-of-the-year pumpkinseed (Lepomis gibbosus): stomach contents and fatty acid analyses. Aquatic Ecology, 45(1):23–33, March 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9320-1.

Meerhoff:2003:SRF

[MMRG03] Mariana Meerhoff, Néstor Mazzeo, and Lorena Rodríguez-Gallego. The structuring role of free-floating versus submerged plants in a subtropical shallow lake. Aquatic Ecology, 37(4):377–391, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000007041.57843.0b.

Mulderij:2005:AGI

[MMV05] G. Mulderij, W. M. Mooij, and E. Van Donk. Allelopathic growth inhibition and colony formation of the green alga Scenedesmus obliquus by the aquatic macrophyte Stratiotes aloides. Aquatic Ecology, 39(1):11–21, March 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-1021-1.

Mulderij:2009:IBM

[MMV09] G. Mulderij, B. Mau, and E. Van Donk. Interaction between the macrophyte Stratiotes aloides and filamentous algae: does it indicate allelopathy? Aquatic Ecology, 43 (2):305–312, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9194-7.

Mirzajani:2021:TFE

[MNJ21] Alireza Mirzajani, Saeid Naderi, and Jamileh Javidpour. Trophic flexibility of Eurasian otter (Lutra lutra) in Anzali Wetland, Iran, assessed by fecal and stable isotope analysis. Aquatic Ecology, 55(2):401–415, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09832-x.

Mukhanov:2004:SEB

[MNK04] Vladimir S. Mukhanov, Olga G. Naidanova, and Richard B. Kemp. The spring energy budget of the algal mat community in a Crimean hypersaline lake determined by microcalorimetry. Aquatic Ecology, 38(3):375–385, September 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000035169.08581.10.

Mayeli:2005:ESMa

[MNS05a] S. M. Mayeli, S. Nandini, and S. S. S. Sarma. The efficacy of Scenedesmus morphology as a defense mechanism against grazing by selected species of rotifers and cladocerans. Aquatic Ecology, 38(4):515–524, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-0329-1.

Mayeli:2005:ESMb

[MNS05b] S. M. Mayeli, S. Nandini, and S. S. S. Sarma. The efficacy of Scenedesmus morphology as a defense mechanism against grazing by selected species of rotifers and cladocerans. Aquatic Ecology, 38(4):515–524, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-0329-9.

Mutinova:2016:HSE

[MNT16] Petra Thea Mutinová, Jiří Neustupa, and Antonio Terlizzi. Host specificity of epiphytic diatom (Bacillariophyceae) and desmid (Desmidiales) communities. Aquatic Ecology, 50(4): 697–709, December 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9587-y.

Murray:1992:ODS

[MO92] D. A. Murray and J. P. O’Connor. Observations on the distribution of some chironomids (Diptera) and caddisflies (Trichoptera) in the Killarney area, Ireland. Aquatic Ecology, 26 (2-4):393–396, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255267.

Moreno-Ostos:2008:SDD

[MOCPG08] Enrique Moreno-Ostos, Luis Cruz-Pizarro, and D. Glen George. The spatial distribution of different phytoplankton functional groups in a Mediterranean reservoir. Aquatic Ecology, 42(1):115–128, March 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
007-9087-1.

Moreno-Ostos:2009:IWI

[MOCPG09] Enrique Moreno-Ostos, Luis Cruz-Pizarro, and D. Glen George. The influence of wind-induced mixing on the vertical distribution of buoyant and sinking phytoplankton species. Aquatic Ecology, 43(2):271–284, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9167-x.

Moreno-Ostos:2007:IBS

[MOdSCP07] Enrique Moreno-Ostos, Sergio L. Rodrigues da Silva, and Luis Cruz-Pizarro. Interannual and between-site variability in the occurrence of clear water phases in two shallow Mediterranean lakes. Aquatic Ecology, 41(2):285–297, June 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-006-9072-0.

Moed:1971:CAL

[Moe71] J. R. Moed. Chemische analyses in de Lindevallei. (Dutch) [Chemical analyzes in the Linde Valley]. Aquatic Ecology, 5(3):142–145, September 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185891.

Mohlenberg:1999:EMN

[Møh99] Flemming Møhlenberg. Effect of meteorology and nutrient load on oxygen depletion in a Danish micro-tidal estuary. Aquatic Ecology, 33(1):55–64, March 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009956210537.

Mols:1972:MOO

[Mol72] E. J. P. J. Mols. Macrofauna-onderzoek in de Ooypolder. (Dutch) [Macrofauna research in the Ooypolder]. Aquatic Ecology, 6(2):100–112, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02304099.

Mols:1976:PDD

[Mol76] E. J. P. J. Mols. Periphyton diatoms in a ditch near Tienhoven. Aquatic Ecology, 10(1):40–47, April 1976. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02308500.

Mol:1980:RIF

[Mol80] A. Mol. The role of the invertebrate fauna in the biological assessment of water quality. Aquatic Ecology, 14(3):222–223, December 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260129.

Mommaerts:1973:PPN

[Mom73] J. P. Mommaerts. On the production of phytoplankton (nannoplankton and netplankton) in the south bight of the North Sea. Aquatic Ecology, 7(2):69–72, June 1973. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02282200.

Mataba:2024:IAP

[MOM24] Gordian Rocky Mataba, Fredrick Ojija, and Linus Munishi. Impact of anthropogenic pollution and artisanal fishing on the population of Tilapia spp. Oreochromis niloticus and Oreochromis amphimelas in Lake Manyara, northern Tanzania. Aquatic Ecology, 58(2):451–465, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10083-1.

Mooij:1998:BRE

[Moo98] W. M. Mooij. Book review: Ecology of Shallow Lakes, by Marten Scheffer. Aquatic Ecology, 32(3):255–256, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009997300032.

Moreno-Ostos:2008:RWA

[MOPCP08] Enrique Moreno-Ostos, Mariano Paracuellos, and Luis Cruz-Pizarro. Response of waterbirds to alternating clear and turbid water phases in two shallow Mediterranean lakes. Aquatic Ecology, 42(4):701–706, December 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9141-z.

Moreira:1999:UBI

[Mor99] Francisco Moreira. On the use by birds of intertidal areas of the Tagus estuary: implications for management. Aquatic Ecology, 33(3):301–309, October 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009908722680.

Moss:1997:SSS

[Mos97] Brian Moss. The sweet, the salt and the sour: Contrasts in the restoration of aquatic systems. Aquatic Ecology, 30 (4):255–264, May 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02085870.

Moubayed:1992:BPS

[Mou92] J. Moubayed. Boreoheptagyia phoenicia sp.n. (Diptera, Chironomidae, Diamesinae) from a Lebanese mountain spring. Aquatic Ecology, 26(2-4):187–190, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255240.

Moreira:1993:EGS

[MQC93] M. H. Moreira, H. Queiroga, and M. R. Cunha. Environmental gradients in a southern Europe estuarine system: Ria de Aveiro, Portugal implications for soft bottom macrofauna colonization. Aquatic Ecology, 27(2-4):465–482, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334807.

Moreira:1995:P

[MQR95] Maria Helena Moreira, Victor Quintino, and Ana Maria Rodrigues. Preface. Aquatic Ecology, 29(3-4):201, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02084215.

Merks:1981:SSV

[MR81] A. G. A. Merks and J. W. Rijstenbil. Saline seepage and vertical distribution of oxygen in a brackish ditch. Aquatic Ecology, 15(3):111–121, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255170.

Meulemans:1987:PDE

[MR87] John T. Meulemans and Pieter J. Roos. Production and decomposition in the emergent littoral zone of Lake Maarsseveen 1, an overview. Aquatic Ecology, 21(1):61–69, July 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255456.

Messyasz:2011:AFA

[MR11] Beata Messyasz and Andrzej Rybak. Abiotic factors affecting the development of Ulva sp. (Ulvophyceae; Chlorophyta) in freshwater ecosystems. Aquatic Ecology, 45(1):75–87, March 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-010-9333-9.

Masclaux:2017:ETF

[MR17] Hélène Masclaux and Nicole B. Richoux. Effects of temperature and food quality on isotopic turnover and discrimination in a cladoceran. Aquatic Ecology, 51(1):33–44, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-016-9592-1.

Martins:2025:TPM

[MRBQ^{+}25] Thaís Fabiane Gomes Martins, Bárbara Rani-Borges, Lucas Gonçalves Queiroz, Karen Ferreira de Souza, and Marcelo Pompêo. Toxicity of polyethylene microplastics combined with medroxyprogesterone on photosynthetic pigments of Lemna minor. Aquatic Ecology, 59(1):307–322, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10164-9.

Meijer:1989:RBL

[MRD89] M. L. Meijer, A. J. P. Raat, and R. W. Doef. Restoration by biomanipulation of lake Bleiswijkse Zoom (The Netherlands): First results. Aquatic Ecology, 23(1):49–57, March 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02286426.

Moore:2010:DTA

[MRR10] John E. Moore, Juluri R. Rao, and Paul J. Rooney. Determination of total antibiotic resistance in waterborne bacteria in rivers and streams in Northern Ireland: Can antibiotic-resistant bacteria be an indicator of ecological change? Aquatic Ecology, 44(2):349–358, June 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9294-z.

Marklund:2002:RBM

[MS02] Ola Marklund and Håkan Sandsten. Reduction of benthic macroinvertebrates due to waterfowl foraging on submerged vegetation during autumn migration. Aquatic Ecology, 36(4): 541–547, December 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1021133713777.

Machacek:2007:LHR

[MS07] Jiří Macháček and Jaromír Sed’a. Life history response of Daphnia galeata to heterogeneous conditions within a reservoir as determined in a cross-designed laboratory experiment. Aquatic Ecology, 41(1):55–66, March 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9045-3.

Mannino:2008:EFF

[MS08] Anna Maria Mannino and Gianluca Sara. Effects of fishfarm biodeposition on periphyton assemblages on artificial substrates in the southern Tyrrhenian Sea (Gulf of Castellammare, Sicily). Aquatic Ecology, 42(4):575–581, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9131-1.

Mihaljevic:2011:CBT

[MS11] Melita Mihaljević and Filip Stević. Cyanobacterial blooms in a temperate river-floodplain ecosystem: the importance of hydrological extremes. Aquatic Ecology, 45(3):335–349, September 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-011-9357-9.

Mezek:2010:EFH

[MSB10] Tadej Mezek, Tatjana Simčič, and Anton Brancelj. Effect of fasting on hypogean (Niphargus stygius) and epigean (Gammarus fossarum) amphipods: a laboratory study. Aquatic Ecology, 44(2):397–408, June 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9299-7.

Mendez:2015:DBC

[MSB15] María M. Mendez, Evangelina Schwindt, and Alejandro Bortolus. Differential benthic community response to increased habitat complexity mediated by an invasive barnacle. Aquatic Ecology, 49(4):441–452, December 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9536-1.

Mukhopadhyay:2008:CBN

[MSD08] Gautam Mukhopadhyay, Sourav Sengupta, and Anjana Dewanji. Changes in biomass and nutrient content of Nymphoides hydrophylla (Lour.) O. Kuntz. in a tropical pond: a comparison with other tropical and temperate species. Aquatic Ecology, 42(4):597–605, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9145-8.

Marini:2022:EIS

[MSDDD22] Keila Marini, Paula Spotorno, Débora Diniz, and Paula Dentzien-Dias. Ecological interactions on shells of Mactra isabelleana d’Orbigny, 1846 (Mollusca: Bivalvia) from southern Brazil: first record of a unique host–parasite interaction. Aquatic Ecology, 56(4):1205–1216, December 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09986-2.

Montez:2021:ICT

[MSF21] Raymond D. Montez, Daniel Saenz, and Kenneth Farrish. The influence of Chinese tallow (Triadica sebifera) leaf litter on water quality and microbial community composition. Aquatic Ecology, 55(1):265–282, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09828-z.

Muller:1993:UEI

[MSK93] M. Müller, M. Schirmer, and J. Kettler. Use of Enteromorpha intestinalis (Chlorophyceae) for active biomonitoring of heavy metals in the Weser estuary. Aquatic Ecology, 27 (2-4):189–195, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334782.

Milosevic:2013:STP

[MSP13] Djuradj Milošević, Vladica Simić, and Momir Paunović. Spatio-temporal pattern of the Chironomidae community: toward the use of non-biting midges in bioassessment programs. Aquatic Ecology, 47(1):37–55, March 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9423-y.

Maslanko:2023:RPL

[MSRB23] Weronika Maślanko, Joanna Sender, and Monika Różańska-Boczula. Rushes as a phytoindicator of a lake’s surrounding land use. Aquatic Ecology, 57(3):633–652, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10034-w.

Manikowska-Slepowronska:2016:DPC

[MŚSJ16] Brygida Manikowska-Ślepowrońska, Barbara Szydzik, and Dariusz Jakubas. Determinants of the presence of conflict bird and mammal species at pond fisheries in western Poland. Aquatic Ecology, 50(1):87–95, March 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9554-z.

Martin:2010:LPS

[MSW10] Peter Martin, Elisabeth Stur, and Sofia Wiedenbrug. Larval parasitism of spring-dwelling alpine water mites (Hydrachnidia, Acari): a study with particular reference to chironomid hosts. Aquatic Ecology, 44(2):431–448, June 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9301-4.

Muschiol:2009:LEM

[MT09] Daniel Muschiol and Walter Traunspurger. Life at the extreme: meiofauna from three unexplored lakes in the caldera of the Cerro Azul volcano, Galápagos Islands, Ecuador. Aquatic Ecology, 43(2):235–248, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9202-y.

Masmoudi:2015:SGN

[MTA15] Salma Masmoudi, Emmanuelle Tastard, and Habib Ayadi. Salinity gradient and nutrients as major structuring factors of the phytoplankton communities in salt marshes. Aquatic Ecology, 49(1):1–19, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
014-9500-5.

Michail:2024:EFC

[MTC^{+}24] Christina Michail, Valentina Tanduo, Fabio Crocetta, Ioannis Giovos, Sofia Litsiou, and Periklis Kleitou. Engagement of fishers in citizen science enhances the knowledge on alien decapods in Cyprus (eastern Mediterranean sea). Aquatic Ecology, 58(1):107–116, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10046-6.

Miranda:2024:EIE

[MTdS^{+}24] Lucília S. Miranda, Sílvia Tavares, Antonina dos Santos, Emanuel J. Gonçalves, Ester A. Serrão, and Márcio A. G. Coelho. L. ipkea ruspoliana Vogt, 1886 (Cnidaria: Staurozoa) in Portugal: the contribution of citizen science to range extension and taxonomic discussion of rare species. Aquatic Ecology, 58(1):31–45, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10070-6.

Mooij:2010:COI

[MTJ10] Wolf M. Mooij, Dennis Trolle, and Jan H. Janse. Challenges and opportunities for integrating lake ecosystem modelling approaches. Aquatic Ecology, 44(3):633–667, September 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-010-9339-3.

Mitsuo:2011:EIF

[MTS11] Yoshito Mitsuo, Hiroshi Tsunoda, and Yutaro Senga. Environmental influences on fish assemblages in irrigation ponds. Aquatic Ecology, 45(4):473–482, November 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9368-6.

Muro-Torres:2019:FWS

[MTSJA19] V. M. Muro-Torres, M. F. Soto-Jiménez, and F. Amezcua. Food web structure of a subtropical coastal lagoon. Aquatic Ecology, 53(3):407–430, September 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
019-09698-0.

Munda:1994:BAV

[Mun94] Ivka M. Munda. The benthic algal vegetation of the Mjóifjördur, Eastern Iceland. Aquatic Ecology, 28(1):35–49, March 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334243.

Mur:1969:IVB

[Mur69] L. R. Mur. De invloed van brak water op de groei van Chlorococcales, speciaal Scenedesmus. (Dutch) [The influence of brackish water on the growth of Chlorococcales, especially Scenedesmus]. Aquatic Ecology, 3(3):64, November 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185855.

Mur:1974:P

[Mur74a] L. R. Mur. Publications. Aquatic Ecology, 8(3):290–291, December 1974. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257507.

Mur:1974:RSS

[Mur74b] L. R. Mur. The response of Scenedesmus species to the brackish water environment. Aquatic Ecology, 8(1-2):224–231, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02254922.

Muus:1974:EPB

[Muu74] Bent Muus. Ecophysiological problems of the brackish water. Aquatic Ecology, 8(1-2):76–89, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02254908.

Merks:1983:PSS

[MV83] A. G. A. Merks and A. G. Vlasblom. Preservation and storage of samples for the determination of dissolved organic carbon. Aquatic Ecology, 17(1):53–58, June 1983. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255191.

Martin:2014:TEH

[MV14] Charles W. Martin and John F. Valentine. Tolerance of embryos and hatchlings of the invasive apple snail Pomacea maculata to estuarine conditions. Aquatic Ecology, 48(3): 321–326, September 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-014-9486-z.

Minelgaite:2020:ESB

[MvAV20] Greta Minelgaite, Nikki van Alst, and Jes Vollertsen. An exploratory study of benthic diatom communities in stormwater ponds of different land uses and varying biocide contamination. Aquatic Ecology, 54(3):761–774, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09773-x.

Mantzouki:2016:UKE

[MVI16] Evanthia Mantzouki, Petra M. Visser, and Bas W. Ibelings. Understanding the key ecological traits of cyanobacteria as a basis for their management and control in changing lakes. Aquatic Ecology, 50(3):333–350, September 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9526-3.

Millward:1994:PST

[MWC94] G. E. Millward, M. R. Williams, and R. J. Clifton. Particle sources and trace element reactivity in the Humber plume. Aquatic Ecology, 28(3-4):359–364, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334205.

Nsor:2019:MEE

[NAM19] Collins Ayine Nsor, Osei Owusu Antobre, and Foster Mensah. Modelling the effect of environmental disturbance on community structure and diversity of wetland vegetation in Northern Region of Ghana. Aquatic Ecology, 53(1):119–136, March 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09677-5.

Nandini:2000:RRC

[Nan00] S. Nandini. Responses of rotifers and cladocerans to Microcystis aeruginosa (Cyanophyceae): a demographic study. Aquatic Ecology, 34(3):227–242, September 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009986928706.

Natalia:2005:BSH

[Nat05] Kuczyńska-Kippen Natalia. On body size and habitat selection in rotifers in a macrophye-dominated lake Budzyńskie, Poland. Aquatic Ecology, 39(4):447–454, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9003-5.

Nordin:2008:EDM

[NAT08] L. J. Nordin, M. T. Arts, and W. D. Taylor. An evaluation of the diet of Mysis relicta using gut contents and fatty acid profiles in lakes with and without the invader Bythotrephes longimanus (Onychopoda, Cercopagidae). Aquatic Ecology, 42(3): 421–436, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9098-y.

Nausch:2000:EEI

[Nau00] Monika Nausch. Experimental evidence for interactions between bacterial peptidase and alkaline phosphatase activity in the Baltic Sea. Aquatic Ecology, 34(4):331–343, December 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011482815250.

Nienhuis:1980:MDK

[NB80] P. H. Nienhuis and C. Bakker. In memoriam Dr. K. F. Vaas. Aquatic Ecology, 14(3):228–231, December 1980. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260137.

Negus:2022:GDI

[NB22] Peter Negus and Joanna Blessing. Girt by dirt: island biogeography and relationships between invertebrate richness and wetted area of spring wetlands. Aquatic Ecology, 56 (1):299–310, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09920-y.

Netti:2025:NPF

[NBD^{+}25] Marissa Netti, Suzannah J. Bozarth, Jared W. Dickson, Marsha G. Williams, Lance R. Williams, and Matthew J. Greenwold. Niche partitioning of food resources by freshwater mussels in a multispecies mussel bed in the Sabine River, Texas. Aquatic Ecology, 59(1):39–51, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10146-x.

Nassal:1998:PJR

[NBM98] Brigitte Nassal, Wiebke Burghard, and Gerhard Maier. Predation by juvenile roach on the calanoid copepod Eudiaptomus gracilis and the cyclopoid copepod Cyclops vicinus: a laboratory investigation with mixed and single prey. Aquatic Ecology, 32(4):335–340, December 1998. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009995432725.

Nessi:2021:KAA

[NBT21] Alessandro Nessi, Alessandro Balestrieri, and Paolo Tremolada. Kingfisher (Alcedo atthis) diet and prey selection as assessed by the analysis of pellets collected under resting sites (River Ticino, north Italy). Aquatic Ecology, 55(1):135–147, March 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09817-2.

Nunes:2020:ESG

[NC20] Laís Samira Correia Nunes and Antonio Fernando Monteiro Camargo. Effects of salinity on growth, competitive interaction and total nitrogen content of two estuarine macrophyte species cultivated on artificial substrate. Aquatic Ecology, 54 (4):973–983, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09787-5.

Nery:2025:DCS

[NCDS^{+}25] Leticia Nery, Bruno Carvalho, Karina Dias-Silva, Thiago Bernardi Vieira, and Felipe Ferraz Figueiredo Moreira. Distribution and conservation of semiaquatic bugs (Hemiptera: Heteroptera: Gerromorpha) in the states of Alagoas and Sergipe, northeastern Brazil. Aquatic Ecology, 59(1):21–37, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10145-y.

Neustupa:2012:STC

[NČŠ12] Jiří Neustupa, Kateřina Černá, and Jan Št’astný. Spatiotemporal community structure of peat bog benthic desmids on a microscale. Aquatic Ecology, 46(2):229–239, June 2012.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9394-z.

N:1991:SRS

[NCT91] Revis N., J. Castel, and M. L. M. Tackx. Some reflections on the structure of the mandibleplate of Eurytemora affinis (Copepoda, Calanoida). Aquatic Ecology, 25(1):45–50, October 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02259588.

Nienhuis:1996:TFY

[NDW96] P. H. Nienhuis, B. H. H. De Bree, and E. G. J. Wessel. Twenty-five years of changes in the distribution and biomass of eelgrass, Zostera marina, in Grevelingen lagoon, The Netherlands. Aquatic Ecology, 30(2-3):107–117, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02272232.

Nwadiaro:1985:PCP

[NE85] C. S. Nwadiaro and E. O. Ezefili. A preliminary checklist of the phytoplankton of New Calabar River, Lower Niger Delta. Aquatic Ecology, 19(2):133–138, December 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270759.

Nandini:2007:LSD

[NEGS07] S. Nandini, C. Enríquez-García, and S. S. S. Sarma. A laboratory study on the demography and competition of three species of littoral cladocerans from Lake Huetzalin, Xochimilco, Mexico. Aquatic Ecology, 41(4):547–556, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9116-0.

Nilsson:2005:LTDa

[NEP05a] Jonas Nilsson, Roland Engkvist, and Lars-Eric Persson. Long-term decline and recent recovery of Fucus populations along the rocky shores of southeast Sweden, Baltic Sea. Aquatic Ecology, 38(4):587–598, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-5665-2.

Nilsson:2005:LTDb

[NEP05b] Jonas Nilsson, Roland Engkvist, and Lars-Eric Persson. Long-term decline and recent recovery of Fucus populations along the rocky shores of southeast Sweden, Baltic Sea. Aquatic Ecology, 38(4):587–598, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-5665-7.

Nava-Gonzalez:2020:BDI

[NGSOAD20] Bisbrian Alhelí Nava-González, Ireri Suazo-Ortuño, and Javier Alvarado-Díaz. Batrachochytrium dendrobatidis infection in amphibians from a high elevation habitat in the trans-Mexican volcanic belt. Aquatic Ecology, 54(1):75–87, March 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09727-y.

Nolte:1992:LCP

[NH92] Ulrike Nolte and Thomas Hoffmann. Life cycle of Pseudodiamesa branickii (Chironomidae) in a small upland stream. Aquatic Ecology, 26(2-4):309–314, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255256.

Notenboom:1996:MAD

[NHF96] Jos Notenboom, Wim Hendrix, and Arie-Jan Folkerts. Meiofauna assemblages discharged by springs from a phreatic aquifer system in the Netherlands. Aquatic Ecology, 30(1): 1–13, July 1996. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02092142.

Nishino:2015:ISG

[NHiN15] Hiroshi Nishino, Yoshikuni Hodoki, and Shin ichi Nakano. Identification of species and genotypic compositions of Cryptomonas (Cryptophyceae) populations in the eutrophic Lake Hira, Japan, using single-cell PCR. Aquatic Ecology, 49(3): 263–272, September 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9520-9.

Nurminen:2018:FIC

[NHL18] Leena Nurminen, Noora Hellén, and Hannu Lehtonen. Fishing-induced changes in predation pressure by perch (Perca fluviatilis) regulate littoral benthic macroinvertebrate biomass, density, and community structure. Aquatic Ecology, 52(1):1–16, March 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9641-4.

Nienhuis:1973:SMB

[Nie73] P. H. Nienhuis. Salt-marsh and beach plain as a habitat for benthic algae. Aquatic Ecology, 7(1):15–24, March 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02279916.

Nienhuis:1974:VLC

[Nie74] P. H. Nienhuis. Variability in the life cycle of Rhizoclonium riparium (Roth) Harv. (Chlorophyceae: Cladophorales) under Dutch estuarine conditions. Aquatic Ecology, 8(1-2):172– 178, July 1974. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02254917.

Nienhuis:1978:LGC

[Nie78] P. H. Nienhuis. Lake Grevelingen: a case study of ecosystem changes in a closed estuary. Aquatic Ecology, 12(3-4): 246–259, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259186.

Nienhuis:1979:RME

[Nie79a] P. H. Nienhuis. The role of macrophytes, especially Zostera marina, in the carbon cycle of Lake Grevelingen (IV). Aquatic Ecology, 13(2-3):79–80, December 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284731.

Nienhuis:1979:SOC

[Nie79b] P. H. Nienhuis. The study of the organic carbon cycle in brackish Lake Grevelingen (SW Netherlands) (I). Aquatic Ecology, 13(2-3):75–76, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02284728.

Nienhuis:1982:SMP

[Nie82] P. H. Nienhuis. A simulation model of production, seasonal changes in biomass and distribution of eelgrass (Zostera marina) in Lake Grevelingen. Aquatic Ecology, 16(2-3):286, December 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255383.

Nienhuis:1994:CEW

[Nie94] P. H. Nienhuis. Causes of the eelgrass wasting disease: Van der Werff’s changing theories. Aquatic Ecology, 28(1):55– 61, March 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334245.

Nienhuis:1998:BRC

[Nie98] P. H. Nienhuis. Book review: Coastal Zone Management Imperative for Mar-itime Developing Nations. Coastal Systems and Continental Margins. Edited by B. U. Haq, S. M. Haq, G. Kullenberg and J. H. Stel. Aquatic Ecology, 32(4):370–371, December 1998. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009995016356.

Nix:2000:LHC

[NJ00] M. H. Nix and D. G. Jenkins. Life history comparisons of Daphnia obtusa from temporary ponds, cultured with a lowquality food. Aquatic Ecology, 34(1):19–27, March 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009985529660.

Nowosad:2007:URD

[NKKG07] Piotr Nowosad, Natalia Kuczyńska-Kippen, and Thaddeus K. Graczyk. The use of rotifers in detecting protozoan parasite infections in recreational lakes. Aquatic Ecology, 41(1):47–54, March 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-006-9043-5.

Nienhuis:1991:KBH

[NL91] Piet H. Nienhuis and Rob H. D. Lambeck. Kees Bakker: Hydrobiologist. Aquatic Ecology, 25(1):4–8, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02259582.

Neumeier:2006:EEP

[NLC06] Urs Neumeier, Cathy H. Lucas, and Michael Collins. Erodibility and erosion patterns of mudflat sediments investigated using an annular flume. Aquatic Ecology, 40(4):543–554, December 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-0189-8.

Noges:2010:PPA

[NLF10] Tiina Nõges, Helen Luup, and Tõnu Feldmann. Primary production of aquatic macrophytes and their epiphytes in two shallow lakes (Peipsi and Võrtsjärv) in Estonia. Aquatic Ecology, 44(1):83–92, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9249-4.

Navarro:2009:BBP

[NMB09] Marcela Bastidas Navarro, Beatriz Modenutti, and Esteban Balseiro. Balance between primary and bacterial production in North Patagonian shallow lakes. Aquatic Ecology, 43(4): 867–878, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9220-9.

Nonato:2021:HMB

[NMdAM21] Flávia Alessandra Silva Nonato, Thaisa Sala Michelan, and Luciano Fogaça de Assis Montag. Heterogeneity of macrophyte banks affects the structure of fish communities in flooded habitats of the Amazon Basin. Aquatic Ecology, 55 (1):215–226, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09823-4.

Nyitrai:2012:TEF

[NMP12] Daniel Nyitrai, Filipe Martinho, and Miguel A. Pardal. Trends in estuarine fish assemblages facing different environmental conditions: combining diversity with functional attributes. Aquatic Ecology, 46(2):201–214, June 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9392-1.

Nandini:2021:SRD

[NMS21] S. Nandini, M. R. Miracle, and S. S. S. Sarma. Strain-related differences in bacterivory and demography of Diaphanosoma mongolianum (Cladocera) in relation to diet and previous exposure to cyanobacteria in nature. Aquatic Ecology, 55(4): 1225–1239, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09892-z.

Neddermann:2005:EOIa

[NN05a] Kirsten Neddermann and Monika Nausch. Effects of organic and inorganic nitrogen compounds on the activity of bacterial alkaline phosphatase. Aquatic Ecology, 38(4):475–484, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-5666-1.

Neddermann:2005:EOIb

[NN05b] Kirsten Neddermann and Monika Nausch. Effects of organic and inorganic nitrogen compounds on the activity of bacterial alkaline phosphatase. Aquatic Ecology, 38(4):475–484, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-004-5666-6.

Nascimento:2024:SRS

[NNB^{+}24] L. S. Nascimento, M. A. Noernberg, T. B. Bleninger, A. Lindner, and M. Nogueira Júnior. Not such a rare species, after all? Insights into Drymonema gorgo Müller 1883 (Cnidaria, Scyphozoa), a large and little-known jellyfish from Brazil. Aquatic Ecology, 58(1):17–30, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10074-2.

Nascimento:2022:SMS

[NNJ22] L. Silva Nascimento, M. Almeida Noernberg, and M. Nogueira Júnior. Social media in service of marine ecology: new observations of the ghost crab Ocypode quadrata (Fabricius, 1787) scavenging on Portuguese man-of-war Physalia physalis (Linnaeus, 1758). Aquatic Ecology, 56(3):859–864, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-022-09947-9.

Negishi:2022:ECP

[NNN22] Junjiro N. Negishi, Tomohiro Nakagawa, and Futoshi Nakamura. Exceptional color preferences for flying adult aquatic insects. Aquatic Ecology, 56(1):325–330, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09914-w.

Nwadiaro:1985:PPC

[NO85] C. S. Nwadiaro and E. O. Oji. Phytoplankton productivity and chlorophyll-A concentration of Oguta Lake in Southeastern Nigeria. Aquatic Ecology, 19(2):123–131, December 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02270758.

Norberg:1998:ETL

[Nor98] Jon Norberg. Effects of temperature and light on the composition of brackish-water rock pool ecosystems. Aquatic Ecology, 32(4):323–334, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/A:
1009947928529.

Netto:2009:BCR

[NP09] Sérgio A. Netto and Tiago J. Pereira. Benthic community response to a passive fishing gear in a coastal lagoon (South Brazil). Aquatic Ecology, 43(2):521–538, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9177-8.

Nagelkerken:2003:SBD

[NPH03] I. Nagelkerken, L. P. J. J. Pors, and P. Hoetjes. Swimming behaviour and dispersal patterns of headstarted loggerhead turtles Caretta caretta. Aquatic Ecology, 37(2):183– 190, April 2003. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1023924631480.

Nandini:1997:SPG

[NR97] S. Nandini and T. R. Rao. Somatic and population growth in selected cladoceran and rotifer species offered the cyanobacterium it Microcystis aeruginosa as food. Aquatic Ecology, 31(3):283–298, September 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009940127959.

Ntitslidou:2021:WDB

[NRB21] Chrysoula Ntitslidou, Bruno Rossaro, and Dimitra C. Bobori. What drives benthic macroinvertebrate dispersal in different lake substrata? The case of three Mediterranean lakes. Aquatic Ecology, 55(3):1033–1050, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09880-3.

Nori:2019:EBA

[NRS19] Javier Nori and Octavio Rojas-Soto. On the environmental background of aquatic organisms for ecological niche modeling: a call for caution. Aquatic Ecology, 53(4):595–605, December 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09711-6.

Notenboom-Ram:1976:CDT

[NRVV76] E. Notenboom-Ram, H. Van Der Hammen, and P. Van Der Vet-Kappetein. The Cladocera of a ditch at Tienhoven. Aquatic Ecology, 10(1):59–65, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02308502.

Nandini:2007:EAA

[NS07] S. Nandini and S. S. S. Sarma. Effect of algal and animal diets on life history of the freshwater copepod Eucyclops serrulatus (Fischer, 1851). Aquatic Ecology, 41(1):75–84, March 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-006-9051-5.

Nandini:2021:PGD

[NS21] S. Nandini and S. S. S. Sarma. Population growth, demography and competition studies on Dipleuchlanis propatula (Gosse, 1886) (Rotifera: Euchlanidae). Aquatic Ecology, 55(4):1305–1316, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
021-09883-0.

Nawaz:2024:SDB

[NSB24] Md. Anwar Nawaz, Kandhasamy Sivakumar, and Gurunathan Baskar. Seasonal dynamics of body size in calanoid copepods (Calanoida: Copepoda) from the stressed tropical coast of India, Chennai, Bay of Bengal. Aquatic Ecology, 58(2):363–373, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10075-1.

Nevalainen:2008:IAG

[NSK08] Liisa Nevalainen and Kaarina Sarmaja-Korjonen. Intensity of autumnal gamogenesis in chydorid (Cladocera, Chydoridae) communities in southern Finland, with a focus on Alonella nana (Baird). Aquatic Ecology, 42(1):151–163, March 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9090-6.

Nawaz:2025:UEP

[NudWB^{+}25] Md Anwar Nawaz, Mehraj ud din War, Gurunathan Baskar, Sivalingam Valliappan Meenakshi, and Kandhasamy Sivakumar. Unveiling the ecology of planktonic harpacticoids (Harpacticoida: Copepoda) in a stressed tropical coastal ecosystem, Bay of Bengal, India. Aquatic Ecology, 59(1): 159–167, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10154-x.

Nagelkerken:2001:FFG

[NvdVdlM01] I. Nagelkerken, G. van der Velde, and E. Cocheret de la Morinière. Fish feeding guilds along a gradient of bay biotopes and coral reef depth zones. Aquatic Ecology, 35 (1):73–86, March 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1011416902370.

Noordhuis:2016:FIW

[NvZvG16] Ruurd Noordhuis, Bastiaan G. van Zuidam, and Gerben J. van Geest. Further improvements in water quality of the Dutch Borderlakes: two types of clear states at different nutrient levels. Aquatic Ecology, 50(3):521–539, September 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9521-8.

Neustupa:2024:ETP

[NW24] Jiri Neustupa and Katerina Woodard. The effects of temperature on plasticity, shape symmetry and seasonal variation in the freshwater benthic green microalga Micrasterias thomasiana. Aquatic Ecology, 58(3):601–615, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10093-7.

Nwadiaro:1984:LDM

[Nwa84] C. S. Nwadiaro. The longitudinal distribution of macroinvertebrates and fish in a lower Niger delta river (River Sombreiro) in Nigeria. Aquatic Ecology, 18(2):133–140, December 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257052.

Nwadiaro:1987:DVC

[Nwa87] C. S. Nwadiaro. Depth variations in the chemistry of Oguta Lake in Southeastern Nigeria. Aquatic Ecology, 21(2): 133–139, December 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255438.

Nwadukwe:1995:APE

[Nwa95] Francis O. Nwadukwe. Analysis of production, early growth and survival of Clarias gariepinus (Burchell), Heterobranchus longifilis (Val.) (Pisces: Clariidae) and their F_{1} hybrids in ponds. Aquatic Ecology, 29(2):177–182, June 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02336048.

Wang:2021:DIH

[nWdLjZ21] Qiao ning Wang, Xiao dong Li, and Ming jiang Zhou. Detrimental impact of hypoxia on the mortality, growth, reproduction, and enzyme activities of planktonic mysid Neomysis awatschensis. Aquatic Ecology, 55(3):849–859, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09864-3.

Nally:2011:GBP

[NWL11] Ralph Mac Nally, Elizabeth Wallis, and P. Sam Lake. Geometry of biodiversity patterning: assemblages of benthic macroinvertebrates at tributary confluences. Aquatic Ecology, 45(1):43–54, March 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
010-9322-z.

Nakajima:2009:DVZ

[NYT09a] Ryota Nakajima, Teruaki Yoshida, and Tatsuki Toda. Diel variation of zooplankton in the tropical coral-reef water of Tioman Island, Malaysia. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9208-5.

Nakajima:2009:SRC

[NYT09b] Ryota Nakajima, Teruaki Yoshida, and Tatsuki Toda. In situ release of coral mucus by Acropora and its influence on the heterotrophic bacteria. Aquatic Ecology, 43(4):815–823, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9210-y.

N:1975:OGV

[NZ75] Daan N. and J. J. Zijlstra. Oecologische gevolgen van de visserij op Noordzee-Kabeljauw. (Dutch) [Ecological consequences of fishing for North Sea cod]. Aquatic Ecology, 9 (2):89–90, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257644.

Nold:1998:PGF

[NZ98] Stephen C. Nold and Gabriel Zwart. Patterns and governing forces in aquatic microbial communities. Aquatic Ecology, 32(1):17–35, March 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009991918036.

Nihongi:2016:BIA

[NZS16] Ai Nihongi, Joshua J. Ziarek, and J. Rudi Strickler. Behavioural interseasonal adaptations in Daphnia pulicaria (Crustacea: Cladocera) as induced by predation infochemicals. Aquatic Ecology, 50(4):667–684, December 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9585-0.

Okamura:2013:SLS

[OAS13] Beth Okamura, Karen Ayres, and Carl D. Sayer. Shallow lake sediments provide evidence for metapopulation dynamics: a pilot study. Aquatic Ecology, 47(2):163–176, June 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-013-9432-5.

Orlofske:2013:TMR

[OB13] Jessica M. Orlofske and Donald J. Baird. The tiny mayfly in the room: implications of size-dependent invertebrate taxonomic identification for biomonitoring data properties. Aquatic Ecology, 47(4):481–494, December 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9460-1.

Olie:1982:AMO

[OBC82] J. J. Olie, Marianne C. I. Blaauboer, and Th. E. Cappenberg. Aerobic mineralization of organic matter in the open-water region of Lake Vechten. Aquatic Ecology, 16(1):69–70, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255414.

Obertegger:2010:RCI

[OBF10] Ulrike Obertegger, Andrea Borsato, and Giovanna Flaim. Rotifer–crustacean interactions in a pseudokarstic lake: influence of hydrology. Aquatic Ecology, 44(1):121–130, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9285-0.

Olivares:2020:EMR

[OCS20] Manuel Olivares, Albert Calbet, and Enric Saiz. Effects of multigenerational rearing, ontogeny and predation threat on copepod feeding rhythms. Aquatic Ecology, 54(3):697–709, September 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09768-8.

OToole:2008:NOT

[ODI08] C. O’Toole, I. Donohue, and K. Irvine. Nutrient optima and tolerances of benthic invertebrates, the effects of taxonomic resolution and testing of selected metrics in lakes using an extensive European data base. Aquatic Ecology, 42(2):277– 291, June 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9185-8.

Orefice:2022:LMM

[ODR22] Ida Orefice, Valeria Di Dato, and Giovanna Romano. Lipid mediators in marine diatoms. Aquatic Ecology, 56(2):377– 397, June 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09932-8.

Ornolfsdottir:2004:STV

[ÖE04] Erla Björk Örnólfsdóttir and Árni Einarsson. Spatial and temporal variation of benthic Cladocera (Crustacea) studied with activity traps in Lake Myvatn, Iceland. Aquatic Ecology, 38(2):239–257, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000032059.99310.d3.

Olivares-Ferretti:2022:PED

[OFCP22] Pamela Olivares-Ferretti, Viviana Chavez, and Jorge Parodi. Polyphenols extracts from Didymosphenia geminata (Lyngbye) Schmidt altered the motility and viability of Daphnia magna. Aquatic Ecology, 56(1):35–45, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09890-1.

Okimoto:2008:PRG

[OHA08] Takane Okimoto, Kenji Hara, and Futoshi Aranishi. PCR-RFLP genotyping for Japanese and Korean populations of Pacific oyster using mitochondrial DNA noncoding region. Aquatic Ecology, 42(1):1–4, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9064-0.

Oorthuysen:1989:ABS

[OI89] W. Oorthuysen and C. W. Iedema. Active biological sluice management in Lake Grevelingen. Aquatic Ecology, 23(1):85– 89, March 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02286430.

Olafsson:1992:VMB

[Ola92] J. S. Olafsson. Vertical microdistribution of benthic chironomid larvae within a section of the littoral zone of a lake. Aquatic Ecology, 26(2-4):397–403, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255268.

Olenin:1997:BZE

[Ole97] Sergej Olenin. Benthic zonation of the Eastern Gotland Basin, Baltic Sea. Aquatic Ecology, 30(4):265–282, May 1997.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02085871.

Oliveira:2020:CEC

[OLJB20] Fagner Junior M. Oliveira, Dilermando P. Lima-Junior, and Luis Mauricio Bini. Current environmental conditions are weak predictors of fish community structure compared to community structure of the previous year. Aquatic Ecology, 54(3):729–740, September 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09771-z.

Olden:2009:HFA

[OLM09] Julian D. Olden, Eric R. Larson, and Meryl C. Mims. Home-field advantage: native signal crayfish (Pacifastacus leniusculus) out consume newly introduced crayfishes for invasive Chinese mystery snail (Bellamya chinensis). Aquatic Ecology, 43(4):??, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9244-9.

Oliveira:2019:EER

[OLTV19] Felipe Rafael Oliveira, Fernando Miranda Lansac-Tôha, and Luiz Felipe Machado Velho. Effects of the exotic rotifer Kellicottia bostoniensis (Rousselet, 1908) on the microbial food web components. Aquatic Ecology, 53(4):581–594, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09710-7.

Oliveira:2022:ISL

[OMC22] Ricardo Oliveira, Aingeru Martínez, and Cristina Canhoto. Intra-specific leaf trait variability controls leaf decomposition of Vitis vinifera L. cultivars in streams. Aquatic Ecology, 56(1):47–57, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09891-0.

Orlova:2005:REQa

[OMM05a] Marina I. Orlova, Jim R. Muirhead, and Hugh J. MacIsaac. Range expansion of quagga mussels Dreissena rostriformis bugensis in the Volga River and Caspian Sea basin. Aquatic Ecology, 38(4):561–573, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
005-0311-6.

Orlova:2005:REQb

[OMM05b] Marina I. Orlova, Jim R. Muirhead, and Hugh J. MacIsaac. Range expansion of quagga mussels Dreissena rostriformis bugensis in the Volga River and Caspian Sea basin. Aquatic Ecology, 38(4):561–573, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
004-0311-y.

Onyena:2024:CMD

[OMNC24] Amarachi P. Onyena, Cathrine S. Manohar, Joseph A. Nkwoji, and Lucian O. Chukwu. Characterization of the molecular differential responses in marine benthic macroinvertebrates exposed to polycyclic aromatic hydrocarbons. Aquatic Ecology, 58(2):263–288, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10064-4.

Onandia:2014:PPU

[OMV14] Gabriela Onandia, Maria Rosa Miracle, and Eduardo Vicente. Primary production under hypertrophic conditions and its relationship with bacterial production. Aquatic Ecology, 48(4):447–463, December 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
014-9497-9.

Odigi:1988:GLL

[ON88] M. I. Odigi and C. S. Nwadiaro. Geophysical limnology of Lake Oguta in Southeastern Nigeria with notes on its possible origin. Aquatic Ecology, 22(2):113–126, December 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256815.

Owowenu:2023:CRE

[ONA^{+}23] Enahoro Kennedy Owowenu, Chika Felicitas Nnadozie, Frank Akamagwuna, Xavier Siwe Noundou, Jude Edafe Uku, and Oghenekaro Nelson Odume. A critical review of environmental factors influencing the transport dynamics of microplastics in riverine systems: implications for ecological studies. Aquatic Ecology, 57(2):557–570, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10029-7.

Oomen:1972:LVO

[Oom72] H. C. J. Oomen. Het landschap van de Ooy. (Dutch) [the landscape of the Ooy]. Aquatic Ecology, 6(2):51–52, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02304094.

Olafsson:2004:ABS

[ÓP04] Jón S. Ólafsson and David M. Paterson. Alteration of biogenic structure and physical properties by tubebuilding chironomid larvae in cohesive sediments. Aquatic Ecology, 38(2):219–229, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000032050.10546.bb.

Opstelten:1980:DPP

[Ops80] W. Opstelten. Diel patterns of phytoplankton productivity in Lake Vechten. Aquatic Ecology, 14(3):219–220, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260126.

Olivier:2006:HLS

[OR06] Frédéric Olivier and Christian Retière. How to leave or stay on the substratum when you can’t swim? Evidence of the role of mucus thread secretion by postlarvae of Pectinaria koreni (Malmgren) in still water and flume experiments. Aquatic Ecology, 40(4):503–519, December 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-8139-z.

Ottonello:2011:OAT

[OR11] Dario Ottonello and Antonio Romano. Ostracoda and Amphibia in temporary ponds: who is the prey? Unexpected trophic relation in a Mediterranean freshwater habitat. Aquatic Ecology, 45(1):55–62, March 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9323-y.

Oskam:1972:PBS

[Osk72] G. Oskam. Planktonbeheersing in de Biesbosch-Spaarbekkens. (Dutch) [Plankton control in the Biesbosch Spaar Basins]. Aquatic Ecology, 6(3):141–156, September 1972. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02314882.

Ovidio:2013:BQB

[OSM13] Michaël Ovidio, Aurore L. Seredynski, and Billy Nzau Matondo. A bit of quiet between the migrations: the resting life of the European eel during their freshwater growth phase in a small stream. Aquatic Ecology, 47(3):291–301, September 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-013-9444-1.

Offem:2009:FCA

[OSO09] Benedict Obeten Offem, Yemi Akegbejo Samsons, and Isaac Tunde Omoniyi. Fish composition and abundance in the wetlands of Cross River, Nigeria. Aquatic Ecology, 43 (4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9240-0.

Oenema:1988:SEI

[OSR88] Oene Oenema, Rob Steneker, and Jaap Reynders. The soil environment of the intertidal area in the Westerschelde. Aquatic Ecology, 22(1):21–30, August 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256778.

Orchard:2019:AGE

[OTJ19] D. Orchard, G. Tessa, and R. Jehle. Age and growth in a European flagship amphibian: equal performance at agricultural ponds and favourably managed aquatic sites. Aquatic Ecology, 53(1):37–48, March 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-09671-3.

Orlova:2005:IEQ

[OTS05] Marina I. Orlova, Thomas W. Therriault, and Gregory Kh. Shcherbina. Invasion ecology of quagga mussels (Dreissena rostriformis bugensis): a review of evolutionary and phylogenetic impacts. Aquatic Ecology, 39(4):401–418, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-9010-6.

Peeters:1978:ERD

[PA78] J. C. H. Peeters and J. P. Al. Environmental research and the Delta Project. Aquatic Ecology, 12(3-4):203–214, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02259183.

Plante:1998:PPL

[PA98] A. J. Plante and M. T. Arts. Photosynthate production in laboratory cultures (UV conditioned and unconditioned) of Cryptomonas erosa under simulated doses of UV radiation. Aquatic Ecology, 32(4):297–312, December 1998. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009911811188.

Palomar-Abesamis:2017:DMA

[PAAJM17] N. Palomar-Abesamis, R. A. Abesamis, and M. A. Juinio-Meñez. Distribution and microhabitat associations of the juveniles of a high-value sea cucumber, Stichopus cf. horrens, in northern Philippines. Aquatic Ecology, 51(1):17–31, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-016-9591-2.

Padhi:2013:CPG

[Pad13] Abinash Padhi. Contrasting patterns of genetic diversity between the northern and southern populations of yellow bullhead catfish, Ameiurus natalis in North America. Aquatic Ecology, 47(3):357–363, September 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9449-9.

Pantoja-Agreda:2024:ODB

[PAP24] Fernando Pantoja-Agreda and Silvia Pajares. Occurrence and diversity of bacterioplankton in drinking water tropical reservoirs of contrasting trophic state. Aquatic Ecology, 58(2):515–530, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10087-5.

Parma:1968:HVB

[Par68] S. Parma. Hydrografie van de Biesbosch. (Dutch) [Hydrography of the Biesbosch]. Aquatic Ecology, 2(3):95–145, September 1968. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185820.

Parma:1969:ILC

[Par69] S. Parma. II. De levenscyclus can Chaoborus crystallinus (De G.) (Diptera, Chaoboridae) in een Nederlandse vijver. (Dutch) [II. The life cycle can Chaoborus crystallinus (De G.) (Diptera, Chaoboridae) in a Dutch pond]. Aquatic Ecology, 3(3):64–66, November 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185856.

Parma:1978:PAC

[Par78] S. Parma. Political aspects of the closure of the Eastern Scheldt estuary. Aquatic Ecology, 12(3-4):163–175, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02259180.

Parma:1980:HEC

[Par80] S. Parma. The history of the eutrophication concept and the eutrophication in the Netherlands. Aquatic Ecology, 14 (1-2):5–11, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260267.

Parma:1996:IK

[Par96] S. Parma. Ingvar Kristensen (1918–1996). Aquatic Ecology, 30(2-3):239, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02272243.

Persaud:2003:PLI

[PAY03] A. D. Persaud, M. T. Arts, and N. D. Yan. Photoresponses of late instar Chaoborus punctipennis larvae to UVR. Aquatic Ecology, 37(3):257–265, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1025863809405.

Payne:2010:TAR

[Pay10] Richard J. Payne. Testate amoeba response to acid deposition in a Scottish peatland. Aquatic Ecology, 44(2):373– 385, June 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9297-9.

Paffen:1984:LSL

[PB84] B. G. P. Paffen and Th. C. M. Brock. In situ and in vivo studies on the breakdown of Nymphaea alba L. (Nymphaeaceae). Aquatic Ecology, 18(1):72–73, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256681.

Pisman:2004:MMI

[PB04] Tamara I. Pisman and Oksana N. Bogdanova. A mathematical model of the interactions in the mixed culture of invertebrates and algae in the “producer–consumer” aquatic biotic cycle. Aquatic Ecology, 38(3):415–423, September 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000035182.15752.e3.

Pereira:2020:FET

[PBDJ20] Luiza B. Pereira, Silvina Botta, and Fábio G. Daura-Jorge. Feeding ecology of two subspecies of bottlenose dolphin: a tooth tale. Aquatic Ecology, 54(4):941–955, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09785-7.

Peeters:2006:ESW

[PBF06] Edwin T. H. M. Peeters, Bart T. M. J. Brugmans, and Rob J. M. Franken. Effect of silt, water and periphyton quality on survival and growth of the mayfly Heptagenia sulphurea. Aquatic Ecology, 40(3):373–380, September 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9026-y.

Park:2003:SFQ

[PBG03] Sangkyu Park, Michael T. Brett, and Charles R. Goldman. Seston food quality and Daphnia production efficiencies in an oligo-mesotrophic subalpine lake. Aquatic Ecology, 37 (2):123–136, April 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1023972425915.

Palheta:2021:URE

[PBM21] Giovanni S. Palheta, Naraiana L. Benone, and Luciano F. A. Montag. Unraveling the role of environmental factors and dispersal capacity in a metacommunity of Amazonian stream fishes. Aquatic Ecology, 55(1):227–236, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09824-3.

Peglar:2001:TPR

[PBR01] Sylvia M. Peglar, Hilary H. Birks, and M. Ramdani. Terrestrial pollen record of recent land-use changes around nine North African lakes in the CASSARINA Project. Aquatic Ecology, 35(3-4):431–448, October 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011988127627.

Paul:2014:ESF

[PC14] Sourav Paul and Gerard P. Closs. Effects of salinity and food quality on the growth of sub-adult mysids of Tenagomysis spp.: a laboratory study. Aquatic Ecology, 48(2):229– 235, June 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9478-z.

Paolucci:2010:PSL

[PCB10] Esteban M. Paolucci, Daniel H. Cataldo, and Demetrio Boltovskoy. Prey selection by larvae of Prochilodus lineatus (Pisces: Curimatidae): indigenous zooplankton versus veligers of the introduced bivalve Limnoperna fortunei (Bivalvia: Mitilidae). Aquatic Ecology, 44(1):255–267, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9263-6.

Phiri:2011:AIA

[PCD11] Crispen Phiri, Albert Chakona, and Jenny A. Day. Aquatic insects associated with two morphologically different submerged macrophytes, Lagarosiphon ilicifolius and Vallisneria aethiopica, in small fishless ponds. Aquatic Ecology, 45(3):405–416, September 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
011-9363-y.

Pingram:2020:FWC

[PCGW20] M. A. Pingram, K. J. Collier, and J. Garrett-Walker. Food web characteristics of fish communities across degraded lakes provide insights for management in multi-stressor environments. Aquatic Ecology, 54(1):401–419, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09750-4.

Parnachev:2002:GGH

[PD02] Valeri P. Parnachev and Andrei G. Degermendzhy. Geographical, geological and hydrochemical distribution of saline lakes in Khakasia, Southern Siberia. Aquatic Ecology, 36 (2):107–122, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1015670130522.

Petranka:2010:ERS

[PD10] James W. Petranka and Edward J. Doyle. Effects of road salts on the composition of seasonal pond communities: can the use of road salts enhance mosquito recruitment? Aquatic Ecology, 44(1):155–166, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9286-z.

Paiva:2025:TES

[PdCF^{+}25] Larissa Calais Paiva, Débora Reis de Carvalho, Frederico Fernandes Ferreira, Jorge A. Dergam, Marcelo Zacharias Moreira, Carlos Frankl Sperber, and Paulo Santos Pompeu. Trophic ecology of a small characid reflects the degradation of a basin after the rupture of an ore tailings dam. Aquatic Ecology, 59(1):349–362, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10167-6.

Phiri:2007:EDA

[PDD07] C. Phiri, J. Day, and E. Dhlomo. Epiphytic diatoms associated with a submerged macrophyte, Vallisneria aethiopica, in the shallow marginal areas of Sanyati Basin (Lake Kariba): a preliminary assessment of their use as biomonitoring tools. Aquatic Ecology, 41(2):169–181, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9073-z.

Penning:2008:UAM

[PDE08] W. Ellis Penning, Bernard Dudley, and Frauke Ecke. Using aquatic macrophyte community indices to define the ecological status of European lakes. Aquatic Ecology, 42(2):253– 264, June 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9183-x.

Peeters:2024:TAC

[PdE^{+}24] Edwin T. H. M. Peeters, Robin de Vries, Jesper Elzinga, Mercédesz Ludányi, Robbert van Himbeeck, and Ivo Roessink. Triggers affecting crayfish burrowing behaviour. Aquatic Ecology, 58(2):191–206, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10057-3.

Pereira:1995:SVM

[PDM95] M. E. Pereira, A. C. Duarte, and G. E. Millward. Seasonal variability in mercury inputs into the Ria de Aveiro, Portugal. Aquatic Ecology, 29(3-4):291–296, December 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084227.

Piehler:2002:EMN

[PDP02] Michael F. Piehler, Julianne Dyble, and Hans W. Paerl. Effects of modified nutrient concentrations and ratios on the structure and function of the native phytoplankton community in the Neuse River Estuary, North Carolina, USA. Aquatic Ecology, 36(3):371–385, August 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1016529015349.

Platvoet:2006:FMAa

[PDvdV06a] Dirk Platvoet, Jaimie T. A. Dick, and Gerard van der Velde. Feeding on micro-algae in the invasive Ponto– Caspian amphipod Dikerogammarus villosus (Sowinsky, 1894). Aquatic Ecology, 40(2):237–245, June 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9028-9.

Platvoet:2006:FMAb

[PDvdV06b] Dirk Platvoet, Jaimie T. A. Dick, and Gerard van der Velde. Feeding on micro-algae in the invasive Ponto– Caspian amphipod Dikerogammarus villosus (Sowinsky 1984). Aquatic Ecology, 40(2):247, June 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9034-6.

Peeters:1978:RBL

[PE78] J. C. H. Peeters and P. Eilers. The relationship between light intensity and photosynthesis — a simple mathematical model. Aquatic Ecology, 12(2):134–136, July 1978. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260714.

Peelen:1971:MTV

[Pee71] R. Peelen. Mogelijkheden ter voorkoming van eutrofiering van het Zeeuwse Meer en de Grevelingen. (Dutch) [Options to prevent eutrophication of the Zeeland Lake and the Grevelingen]. Aquatic Ecology, 5(1):39–48, March 1971. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185886.

Peelen:1974:CPE

[Pee74] R. Peelen. Changes in the plankton of the estuarine area of the Haringvliet–Hollands Diep–Biesbosch in the S.W. Netherlands caused by the dams through Volkerak and Haringvliet. Aquatic Ecology, 8(1-2):190–200, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254919.

Pellikaan:1982:DPE

[Pel82] G. C. Pellikaan. Decomposition processes of eelgrass, Zostera marina L. Aquatic Ecology, 16(1):83–92, June 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255416.

Pennuto:2000:ELM

[Pen00] Christopher M. Pennuto. Effects of larval movement behavior and density on emergence success and adult body size in a commensal midge. Aquatic Ecology, 34(2):177–184, June 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009985405789.

Persoone:1969:OSV

[Per69] G. Persoone. Oecologische studie van de aangroei op ondergedompelde substraten in de haven van Oostende. (Dutch) [Ecological study of the growth on submerged substrates in the port of Oostende]. Aquatic Ecology, 3(2):36– 37, July 1969. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185849.

Perez-Estrada:2023:DSV

[PEREBM^{+}23] C. J. Pérez-Estrada, R. Rodríguez-Estrella, F. G. Brun-Murillo, P. Gurgo-Salice, R. Valles-Jiménez, E. Morales-Bojórquez, and M. A. Medina-López. Diversity and seasonal variation of the molluscan community associated with the seagrass Halodule wrightii in a marine protected area in the southern Gulf of California. Aquatic Ecology, 57(2):299– 319, June 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10011-3.

Prins:1999:EDP

[PES99] Theo C. Prins, Vincent Escaravage, and Aad C. Smaal. Effects of different N- and P-loading on primary and secondary production in an experimental marine ecosystem. Aquatic Ecology, 33(1):65–81, March 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009929706903.

Pedersen:2007:TLS

[PF07] Morten Lauge Pedersen and Nikolai Friberg. Two lowland stream riffles — linkages between physical habitats and macroinvertebrates across multiple spatial scales. Aquatic Ecology, 41(3):475–490, September 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-1584-x.

Perez-Fuentetaja:1999:SDO

[PFDM99] Alicia Pérez-Fuentetaja, Peter J. Dillon, and Donald J. McQueen. Significance of dissolved organic carbon in the prediction of thermocline depth in small Canadian Shield lakes. Aquatic Ecology, 33(2):127–133, September 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009998118504.

Palacios-Fuentes:2015:EBR

[PFLM15] Pamela Palacios-Fuentes, Mauricio F. Landaeta, and Gabriela Muñoz. Is ectoparasite burden related to host density? Evidence from nearshore fish larvae off the coast of central Chile. Aquatic Ecology, 49(1):91–98, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9507-6.

Palacios-Fuentes:2014:HPL

[PFLO14] Pamela Palacios-Fuentes, Mauricio F. Landaeta, and F. Patricio Ojeda. Hatching patterns and larval growth of a triplefin from central Chile inferred by otolith microstructure analysis. Aquatic Ecology, 48(3):259–266, September 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-014-9481-4.

Perez-Fuentetaja:2000:ZBR

[PFMD00] Alicia Pérez-Fuentetaja, Donald J. McQueen, and Peter J. Dillon. Zooplankton biomass rarely improves predictions of chlorophyll concentration in Canadian Shield lakes that vary in pH. Aquatic Ecology, 34(2):127–136, June 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009934506481.

Pires:2024:ROH

[PFSPdS24] Rita F. T. Pires, Elsa Froufe, Giulia Secci-Petretto, and Antonina dos Santos. Report on the occurrence of the hydromedusa Odessia maeotica (Ostroumoff, 1896) in the northeastern Atlantic revealed by citizen science and integrative taxonomy. Aquatic Ecology, 58(2):323–334, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10071-5.

Peters:2012:CCC

[PFT12] Lars Peters, Carsten Faust, and Walter Traunspurger. Changes in community composition, carbon and nitrogen stable isotope signatures and feeding strategy in epilithic aquatic nematodes along a depth gradient. Aquatic Ecology, 46(3):371–384, September 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
012-9408-x.

Perroca:2022:NEG

[PGA^{+}22] Júlia Fernandes Perroca, Tommaso Giarrizzo, Ernesto Azzurro, Jorge Luiz Rodrigues-Filho, Carolina V. Silva, Marlene S. Arcifa, and Valter M. Azevedo-Santos. Negative effects of ghost nets on Mediterranean biodiversity. Aquatic Ecology, 56(4):??, November 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09985-3.

Paul:2024:SCE

[PGM^{+}24] Pranesh Paul, Neha Kumari Gupta, Debjit Mondal, Abhijit Sikary, and Gautam Aditya. The sailfin catfish (Pterygoplichthys spp.) as a non-native aquarium pet: assessment of the potential distribution and public perception. Aquatic Ecology, 58(1):89–105, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10055-5.

Pereg-Gerk:2002:SNF

[PGSL02] Lily Pereg-Gerk, Nehemia Sar, and Yaakov Lipkin. In situ nitrogen fixation associated with seagrasses in the Gulf of Elat (Red Sea). Aquatic Ecology, 36(3):387–394, August 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1016566019208.

Ptatscheck:2020:SWR

[PGT20a] C. Ptatscheck, S. Gehner, and W. Traunspurger. Should we redefine meiofaunal organisms? The impact of mesh size on collection of meiofauna with special regard to nematodes. Aquatic Ecology, 54(4):1135–1143, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09798-2.

Ptatscheck:2020:IES

[PGT20b] Christoph Ptatscheck, Birgit Gansfort, and Walter Traunspurger. The influence of environmental and spatial factors on benthic invertebrate metacommunities differing in size and dispersal mode. Aquatic Ecology, 54(2):447–461, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09752-2.

Panis:1995:HCC

[PGV95] Luc Int Panis, Boudewijn Goddeeris, and Rudolf Verheyen. The hemoglobin concentration of Chironomus cf. Plumosus L. (Diptera: Chironomidae) larvae from two lentic habitats. Aquatic Ecology, 29(1):1–4, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02061785.

Persic:2009:CPL

[PHB09] Vesna Peršić, Janja Horvatić, and Ivan Bogut. Changes in N and P limitation induced by water level fluctuations in Nature Park Kopački Rit (Croatia): nutrient enrichment bioassay. Aquatic Ecology, 43(1):27–36, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9156-5.

Pothoven:2013:CZC

[PHD13] Steven A. Pothoven, Tomas O. Höök, and Julianne Dyble. Changes in zooplankton community structure associated with the disappearance of invasive alewife in Saginaw Bay, Lake Huron. Aquatic Ecology, 47(1):1–12, March 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9420-1.

Pouil:2021:ARF

[PHM21] Simon Pouil, Amber Hills, and Teresa J. Mathews. Allometric relationships in the filtration rates of the Asian clam Corbicula fluminea fed two phytoplankton species. Aquatic Ecology, 55(3):915–923, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09871-4.

Pihl:1995:RIF

[PIM95] L. Pihl, I. Isaksson, and P.-O. Moksnes. Recent increase of filamentous algae in shallow Swedish bays: Effects on the community structure of epibenthic fauna and fish. Aquatic Ecology, 29(3-4):349–358, December 1995. CODEN AQECF9.

ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02084234.

Pinkster:1975:IAA

[Pin75] Sjouk Pinkster. The introduction of the alien amphipod Gammarus tigrinus Sexton, 1939 (Crustacea, Amphipoda) in The Netherlands and its competition with indigenous species. Aquatic Ecology, 9(3):131–138, December 1975. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02263332.

Pip:1987:SRA

[Pip87] Eva Pip. Species richness of aquatic macrophyte communites in Central Canada. Aquatic Ecology, 21(2):159–165, December 1987. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255441.

Pistoor:1969:PVP

[Pis69] S. J. Pistoor. Een populatie van Pisidium obtusale in het Ketelven. (Dutch) [A population of Pisidium obtusal in the Ketelven]. Aquatic Ecology, 3(2):38–49, July 1969. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185850.

Prastka:1995:SGP

[PJ95] K. E. Prastka and T. D. Jickells. Sediment geochemistry of phosphorus at two intertidal sites on the Great Ouse estuary, S.E. England. Aquatic Ecology, 29(3-4):245–255, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02084222.

Peters:2001:REC

[PJF01] A. J. Peters, K. C. Jones, and A. A. Fathi. Recent environmental change in North African wetland lakes: a baseline study of organochlorine contaminant residues in sediments from nine sites in the CASSARINA Project. Aquatic Ecology, 35(3-4):449–459, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1011980226851.

Precht:2006:NBP

[PJH06] Elimar Precht, Felix Janssen, and Markus Huettel. Nearbottom performance of the Acoustic Doppler Velocimeter (ADV) — a comparative study. Aquatic Ecology, 40(4): 481–492, December 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-004-8059-y.

Pennuto:2008:STP

[PK08] Christopher Pennuto and Dawn Keppler. Short-term predator avoidance behavior by invasive and native amphipods in the Great Lakes. Aquatic Ecology, 42(4):629–641, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9139-6.

Pond:2022:CMS

[PKE22a] Gregory J. Pond, Kelly J. G. Krock, and Leah F. Ettema. Correction to: Macroinvertebrates at the source: flow duration and seasonality drive biodiversity and trait composition in rheocrene springs of the Western Allegheny Plateau, USA. Aquatic Ecology, 56(1):333–335, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09919-5. See [PKE22b].

Pond:2022:MSF

[PKE22b] Gregory J. Pond, Kelly J. G. Krock, and Leah F. Ettema. Macroinvertebrates at the source: flow duration and seasonality drive biodiversity and trait composition in rheocrene springs of the Western Allegheny Plateau, USA. Aquatic Ecology, 56(1):99–121, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09900-2. See correction [PKE22a].

Pulido:2012:EAS

[PKR12] Cristina Pulido, Danny J. H. Keijsers, and Jan G. M. Roelofs. Elevated alkalinity and sulfate adversely affect the aquatic macrophyte Lobelia dortmanna. Aquatic Ecology, 46(3): 283–295, September 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9399-7.

Ptacnik:2008:QRL

[PLC08] R. Ptacnik, L. Lepistö, and Laurence Carvalho. Quantitative responses of lake phytoplankton to eutrophication in Northern Europe. Aquatic Ecology, 42(2):227–236, June 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9181-z.

Peiman:2022:EST

[PLC22] Kathryn S. Peiman, Hsien-Yung Lin, and Steven J. Cooke. Effects of short-term decomposition on isotope values of fish tissues under natural conditions. Aquatic Ecology, 56 (1):173–181, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09907-9.

Pimenta:2022:ECS

[PLdFC22] Larissa Langsdorff Pimenta, Gustavo Pereira Lima, and Flávia de Freitas Coelho. Epiphytic cyanobacterial strains in the roots of Salvinia auriculata and the effect of light and nutrients on the production of heterocyst, akinete and hormogonia. Aquatic Ecology, 56(3):543–553, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09922-w.

Pulzatto:2018:HFM

[PLM18] Mikaela Marques Pulzatto, Lucas Assumpção Lolis, and Roger Paulo Mormul. Herbivory on freshwater macrophytes from the perspective of biological invasions: a systematic review. Aquatic Ecology, 52(4):297–309, December 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9664-5.

Pan:2022:USE

[PLM^{+}22] Shaopu Pan, Qiuhua Li, Chunlan Meng, Mengshu Han, Yiming Ma, and Anton Brancelj. Using a structural equation model to assess the spatiotemporal dynamics and driving factors of phytoplankton in the plateau Hongfeng Reservoir in southwest China. Aquatic Ecology, 56(4):1297–1313, December 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-022-09963-9.

Prastka:1994:PPH

[PM94] Katherine E. Prastka and Stephen J. Malcolm. Particulate phosphorus in the Humber estuary. Aquatic Ecology, 28(3-4): 397–403, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334209.

Piercey:2000:FII

[PM00] David W. Piercey and Edward J. Maly. Factors influencing the induction of diapausing egg production in the calanoid copepod Diaptomus leptopus. Aquatic Ecology, 34(1):9–17, March 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009970719969.

Paul:2022:IWT

[PMA22] Pranesh Paul, Debasish Mondal, and Gautam Aditya. Influencing the way they compete: Exotic predator mediated non-consumptive effects on two co-occurring freshwater snails. Aquatic Ecology, 56(1):285–298, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09918-6.

Patra:2024:TMR

[PMB^{+}24] Bhairab Kumar Patra, Jit Marick, Roshmi Biswas, Soumyajit Banerjee, and Anirban Ash. A tale of Mundeswari River: unfolding endo-helminth biodiversity and seasonal variation. Aquatic Ecology, 58(4):1129–1141, December 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10128-z.

Prokopkin:2010:GOD

[PMD10] I. G. Prokopkin, W. M. Mooij, and A. G. Degermendzhy. A general one-dimensional vertical ecosystem model of Lake Shira (Russia, Khakasia): description, parametrization and analysis. Aquatic Ecology, 44(3):585–618, September 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9326-8.

Penning:2008:CAM

[PME08] W. Ellis Penning, Marit Mjelde, and Frauke Ecke. Classifying aquatic macrophytes as indicators of eutrophication in European lakes. Aquatic Ecology, 42(2):237–251, June 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9182-y.

Picazo:2010:CSW

[PMM10] F. Picazo, J. L. Moreno, and A. Millán. The contribution of standing waters to aquatic biodiversity: the case of water beetles in southeastern Iberia. Aquatic Ecology, 44 (1):205–216, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9279-y.

Parker:1994:MDO

[PMP94] W. R. Parker, L. D. Marshall, and A. J. Parfitt. Modulation of dissolved oxygen levels in a hypertidal estuary by sediment resuspension. Aquatic Ecology, 28(3-4):347–352, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334203.

Pestana:2018:ECU

[PMSS18] Inácio A. Pestana, Annaliza C. Meneguelli-Souza, and Cristina M. M. Souza. Effects of a combined use of macronutrients nitrate, ammonium, and phosphate on cadmium absorption by Egeria densa Planch. and its phytoremediation applicability. Aquatic Ecology, 52(1):51–64, March 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9644-1.

Pereira-Moura:2023:HSM

[PMVdC^{+}23] Lucas Pereira-Moura, Daniel Silas Veras, Fernando Geraldo de Carvalho, Leandro Juen, and Sheyla Regina Marques Couceiro. Habitat specificity and morphology-main filters for the distribution of Odonata in the Cerrado Maranhense, Brazil. Aquatic Ecology, 57(2):443–458, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10021-1.

Pineda-Mendoza:2016:MPM

[PMZMJ16] Rosa María Pineda-Mendoza, Gerardo Zúñiga, and Fernando Martínez-Jerónimo. Microcystin production in Microcystis aeruginosa: effect of type of strain, environmental factors, nutrient concentrations, and N : P ratio on mcya gene expression. Aquatic Ecology, 50(1):103–119, March 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9559-7.

Portela:1994:NMS

[PN94] Luis Ivens Portela and Ramiro Neves. Numerical modelling of suspended sediment transport in tidal estuaries: a comparison between the Tagus (Portugal) and the Scheldt (Belgium–the Netherlands). Aquatic Ecology, 28(3-4):329– 335, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334201.

Perroca:2022:PDH

[PNC22] Júlia Fernandes Perroca, Caio Santos Nogueira, and Rogerio Caetano Costa. Population dynamics of a hololimnetic population of the freshwater prawn Macrobrachium amazonicum (Heller, 1862) (Decapoda, Palaemonidae) in southeastern Brazil. Aquatic Ecology, 56(1):21–34, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09889-8.

Podamo:1974:EBAa

[Pod74a] Jo Podamo. Essai de bilan annuel du transfert de l’azote dans le bassin de chasse d’Ostende. (French) [Annual balance test of nitrogen transfer in the hunting basin of Ostend]. Aquatic Ecology, 8(1-2):46–52, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02254905.

Podamo:1974:EBAb

[Pod74b] Jo Podamo. Essai de bilan annuel du transfert de l’azote dans le bassin de chasse d’Ostende. (French) [Annual balance test of nitrogen transfer in the hunting basin of Ostend]. Aquatic Ecology, 8(1-2):53–66, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02254906.

Podamo:1974:EBAc

[Pod74c] Jo Podamo. Essai de bilan annuel du transfert de l’azote dans le bassin de chasse d’Ostende. (French) [Annual balance test of nitrogen transfer in the hunting basin of Ostend]. Aquatic Ecology, 8(1-2):67–75, July 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02254907.

Polk:1973:PDL

[Pol73] F. Polk. Presentation du “Laboratorium voor Ekologie en Systematiek” (Vrije Universiteit, Brussel). (French) [Presentation of the “Laboratory for Ecology and Systematics” (Free University, Brussels)]. Aquatic Ecology, 7(2):51–53, June 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02282197.

Polderman:1975:SNA

[Pol75] P. J. G. Polderman. Some notes on the algal vegetation of two brackish polders on Texel (The Netherlands). Aquatic Ecology, 9(1):23–34, April 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02257516.

Polderman:1976:WPW

[Pol76] P. J. G. Polderman. Wittrockiella paradoxa Wille (Cladophoraceae) in N.W. European saltmarshes. Aquatic Ecology, 10(2): 98–103, September 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02282217.

Postma:1974:HHB

[Pos74] H. Postma. Hydrography and hydrochemistry of brackish waters. Aquatic Ecology, 8(1-2):40–45, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254904.

Postma:1980:MT

[Pos80] Leo Postma. Modelling as a tool. Aquatic Ecology, 14(1-2):55–63, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260273.

Platvoet:1995:CAF

[PP95] Dirk Platvoet and Sjouk Pinkster. Changes in the amphipod fauna (crustacea) of the Rhine, Meuse and Scheldt estuary due to the ‘delta plan’ coastal engineering works. Aquatic Ecology, 29(1):5–30, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02061786.

Phillips:2008:CNR

[PPC08] G. Phillips, O.-P. Pietiläinen, and A. C. Cardoso. Chlorophyll– nutrient relationships of different lake types using a large European dataset. Aquatic Ecology, 42(2):213–226, June 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9180-0.

Panja:2022:SCT

[PPCH22] Soumyadip Panja, Anupam Podder, Munmun Chakrabarty, and Sumit Homechaudhuri. Species conservation target for freshwater fishes inhabiting Bengal sub-tropical montane rivers of Eastern Himalayas: an indexed value approach for priority determination. Aquatic Ecology, 56(4):1323–1346, December 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-022-09973-7.

Puddick:2016:MMC

[PPH16] Jonathan Puddick, Michèle R. Prinsep, and David P. Hamilton. Modulation of microcystin congener abundance following nitrogen depletion of a Microcystis batch culture. Aquatic Ecology, 50(2):235–246, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9571-6.

Parab:2023:IAF

[PPK^{+}23] Vaishnavi Parab, Jyoti Jagdish Prajapati, Samya Karan, Amiya Ranjan Bhowmick, and Joyita Mukherjee. Impact of abiotic factors and heavy metals in predicting the population decline of Near Threatened fish Notopterus chitala in natural habitat. Aquatic Ecology, 57(4):863–879, December 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-022-09995-1.

Pannard:2016:PEE

[PPL16] Alexandrine Pannard, Julie Pédrono, and Yvan Lagadeuc. Production of exopolymers (EPS) by cyanobacteria: impact on the carbon-to-nutrient ratio of the particulate organic matter. Aquatic Ecology, 50(1):29–44, March 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9550-3.

Palffy:2013:DPT

[PPV13] Károly Pálffy, Mátyás Présing, and Lajos Vörös. Diversity patterns of trait-based phytoplankton functional groups in two basins of a large, shallow lake (Lake Balaton, Hungary) with different trophic state. Aquatic Ecology, 47(2):195–210, June 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9434-3.

Pang:2022:CAE

[PPY^{+}22] Wanting Pang, Yangdong Pan, Qingmin You, Yong Cao, Lizhu Wang, Guiping Deng, and Quanxi Wang. Causes of aquatic ecosystem degradation related to tourism and the feasibility of restoration for karst nature reserves. Aquatic Ecology, 56(4):1231–1243, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09983-5.

Pip:1984:CAP

[PR84] Eva Pip and G. G. C. Robinson. A comparison of algal periphyton composition on eleven species of submerged macrophytes. Aquatic Ecology, 18(2):109–118, December 1984. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02257050.

Prat:1992:LCP

[PR92] Narcis Prat and Maria Rieradevall. Life cycle and production of Cladopelma virescens (Mg.) (Diptera: Chironomidae) in Lake Banyoles (NE Spain). Aquatic Ecology, 26(2-4):315– 320, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255257.

Pennuto:2016:URE

[PR16] C. M. Pennuto and S. M. Rupprecht. Upstream range expansion by invasive round gobies: Is functional morphology important? Aquatic Ecology, 50(1):45–57, March 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9551-2.

Prater:1978:EEE

[Pra78] A. J. Prater. The effect of estuarine engineering schemes on birds. Aquatic Ecology, 12(3-4):322–332, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02259193.

Primicerio:2003:SDH

[Pri03] Raul Primicerio. Size-dependent habitat choice in Daphnia galeata Sars and size-structured interactions among zooplankton in a subarctic lake (Lake Lombola, Norway). Aquatic Ecology, 37(2):107–122, April 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1023942931825.

Perryman:2008:ADC

[PRW08] Shane E. Perryman, Gavin N. Rees, and Christopher J. Walsh. Analysis of denitrifying communities in streams from an urban and non-urban catchment. Aquatic Ecology, 42 (1):95–101, March 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9085-3.

Persson:2006:VDB

[PS06] Anders Persson and Jonas M. Svensson. Vertical distribution of benthic community responses to fish predators, and effects on algae and suspended material. Aquatic Ecology, 40(1):85– 95, March 2006. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-9014-2.

Prins:1997:RFB

[PSD97] Theo C. Prins, Aad C. Smaal, and Richard F. Dame. A review of the feedbacks between bivalve grazing and ecosystem processes. Aquatic Ecology, 31(4):349–359, December 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009924624259.

Prins:1991:SIP

[PSP91] T. C. Prins, A. C. Smaal, and A. J. Pouwer. Selective ingestion of phytoplankton by the bivalves Mytilus edulis L. and Cerastoderma edule (L.). Aquatic Ecology, 25(1):93–100, October 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02259595.

Pokrajac:2024:CDE

[PSSKŽ24a] Jovo Pokrajac, Katarina Stojanović, Sran Stamenković, and Tamara Karan-Žnidaršič. Correction: Diversity and ecology of freshwater Ostracoda (Crustacea) in Central Serbia. Aquatic Ecology, 58(3):641, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10096-4. See [PSSKŽ24b].

Pokrajac:2024:DEF

[PSSKŽ24b] Jovo Pokrajac, Katarina Stojanović, Sran Stamenković, and Tamara Karan-Žnidaršič. Diversity and ecology of freshwater Ostracoda (Crustacea) in Central Serbia. Aquatic Ecology, 58(3):617–640, September 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-

024-10094-6. See correction [PSSKŽ24a].

Parks:2011:IML

[PSW11] Jennifer M. B. Parks, Julie Sircom, and Sandra J. Walde. Invertebrate mesopredators are larger in streams with fish. Aquatic Ecology, 45(2):243–253, May 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9350-3.

Prins:1979:PBU

[PSZ79] H. B. A. Prins, J. F. H. Snel, and P. E. Zanstra. Photosynthetic bicarbonate utilization in the aquatic angiosperms Potamogeton and Elodea. Aquatic Ecology, 13(2-3):106–111, December 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02284741.

Ptatscheck:2014:MAW

[PT14] Chistoph Ptatscheck and Walter Traunspurger. The meiofauna of artificial water-filled tree holes: colonization and bottom-up effects. Aquatic Ecology, 48(3):285–295, September 2014. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-014-9483-2.

Petkovska:2015:LBR

[PU15] Vesna Petkovska and Gorazd Urbanič. The links between river morphological variables and benthic invertebrate assemblages: comparison among three European ecoregions. Aquatic Ecology, 49(2):159–173, June 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9513-8.

Pustelnikovas:1994:TAS

[Pus94] Olegas Pustelnikovas. Transport and accumulation of sediment and contaminants in the Lagoon of Kuršiu̧marios (Lithuania) and Baltic Sea. Aquatic Ecology, 28(3-4):405– 411, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334210.

Parma:1968:VDP

[PV68] S. Parma and K. F. Vaas. Voorwoord. (Dutch) [Preface]. Aquatic Ecology, 2(3):87–91, September 1968. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185817.

Panis:1995:USM

[PV95] Luc Int Panis and Rudolf F. Verheyen. On the use of split moving window analysis for boundary detection in ordered dataseries from benthic communities. Aquatic Ecology, 29(1):49–53, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02061788.

Pinceel:2020:ETI

[PVB20] Tom Pinceel, Bram Vanschoenwinkel, and Luc Brendonck. An empirical test of the impact of drying events and physical disturbance on wind erosion of zooplankton egg banks in temporary ponds. Aquatic Ecology, 54(1):137–144, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09731-2.

PrudHommeVanReine:1996:SEC

[PVC96] W. F. Prud’Homme Van Reine, E. Verheij, and E. Coppejans. Species and ecads of Caulerpa (Ulvophyceae, Chlorophyta) in Malesia (South-East Asia): Taxonomy, biogeography and biodiversity. Aquatic Ecology, 30(2-3):83–98, October 1996. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02272230.

Pentek:2017:MDA

[PVH17] Attila L. Péntek, Csaba F. Vad, and Zsófia Horváth. Metacommunity dynamics of amphibians in years with differing rainfall. Aquatic Ecology, 51(1):45–57, March 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9597-9.

Piirsoo:2007:OPE

[PVV07] Kai Piirsoo, Sirje Vilbaste, and Malle Viik. Origin of phytoplankton and the environmental factors governing the structure of microalgal communities in lowland streams. Aquatic Ecology, 41(2):183–194, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9077-3.

Peeters:1973:MFR

[PW73] J. C. H. Peeters and W. J. Wolff. Macrobenthos and fishes of the rivers Meuse and Rhine, The Netherlands. Aquatic Ecology, 7(3):121–127, September 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02275613.

Pichlova:2005:LKSa

[PWG05a] Radka Pichlová, Anke Weber, and Brigitte Gosser. Leptodora kindtii survival in the laboratory. Aquatic Ecology, 38 (4):537–546, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-0328-x.

Pichlova:2005:LKSb

[PWG05b] Radka Pichlová, Anke Weber, and Brigitte Gosser. Leptodora kindtii survival in the laboratory. Aquatic Ecology, 38 (4):537–546, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-004-0328-2.

Quirino:2015:SVU

[QCF15] Bárbara Angélio Quirino, Natália Carniatto, and Rosemara Fugi. Seasonal variation in the use of food resources by small fishes inhabiting the littoral zone in a neotropical floodplain lake. Aquatic Ecology, 49(4):431–440, December 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9535-2.

Quirino:2019:AMP

[QCF19] Bárbara Angélio Quirino, Natália Carniatto, and Rosemara Fugi. The amphibian macrophyte Polygonum punctatum as a temporary habitat and feeding ground for fish.

Aquatic Ecology, 53(3):441–452, September 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09700-9.

Quesada:2002:HCM

[QJD02] Antonio Quesada, Friedrich Jüttner, and Andrei G. Degermendzhy. Heterotrophic capability of a metalimnetic plankton population in saline Lake Shira (Siberia, Khakasia). Aquatic Ecology, 36(2):219–227, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1015672800327.

Lyu:2016:GPR

[qLlZhY16] Xiao qian Lyu, Yin long Zhang, and Wen hua You. Growth and physiological responses of Eichhornia crassipes to clonal integration under experimental defoliation. Aquatic Ecology, 50(2):153–162, June 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9557-9.

Qian:2023:RRP

[QLW^{+}23] Jin Qian, Yin Liu, Peifang Wang, Jing Hu, Bianhe Lu, Sijing Tang, Chao Zhang, and Pan Gao. Responses of riparian plant roots to acute combined nitrogen and phosphorus stress: changes in root morphology and antioxidant system. Aquatic Ecology, 57(3):715–732, September 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10041-x.

Qin:2020:CFT

[QLZ20] Ying-Ying Qin, Mei-Rong Luo, and Hai-Lei Zheng. Changes in functional traits and stoichiometry of Aegiceras corniculatum propagule in three shrimp aquaculture effluent regions. Aquatic Ecology, 54(4):927–940, December 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09784-8.

Quintino:1995:SCI

[QPP95] V. Quintino, A. M. Picado, and T. H. Pearson. Sediment chemistry — infaunal community structure in a southern European estuary related to solid-phase Microtox(R) toxicity testing. Aquatic Ecology, 29(3-4):427–436, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02084241.

Quadros:2009:ICE

[QSA09] G. Quadros, Soniya Sukumaran, and R. P. Athalye. Impact of the changing ecology on intertidal polychaetes in an anthropogenically stressed tropical creek, India. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9229-8.

Qin:2024:RSL

[QZL^{+}24] Ying-Ying Qin, Ge-Ge Zhu, Ming-Zhong Liang, Hai-Lei Zheng, and Jun-Wei Li. Response strategies of leaf vein traits and stomatal traits of Aegiceras corniculatum to shrimp aquaculture wastewater discharge. Aquatic Ecology, 58(4): 1033–1048, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10122-5.

Rountree:2007:STH

[RA07] Rodney A. Rountree and Kenneth W. Able. Spatial and temporal habitat use patterns for salt marsh nekton: implications for ecological functions. Aquatic Ecology, 41(1):25– 45, March 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-006-9052-4.

Raat:1989:GPB

[Raa89] A. J. P. Raat. Growth and production of O^{+} bream (Abramis brama), O^{+} roach (Rutilus rutilus) and O^{+} carp (Cyprinus carpio) in 10 drainable 0.1 Ha ponds. Aquatic Ecology, 23(1):67–72, March 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02286428.

Rundle:1998:SMC

[RAA98] S. D. Rundle, M. J. Attrill, and A. Arshad. Seasonality in macroinvertebrate community composition across a neglected ecological boundary, the freshwater-estuarine transition zone. Aquatic Ecology, 32(3):211–216, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009934828611.

Reale:2021:DEC

[RAD21] Justin K. Reale, Thomas P. Archdeacon, and Clifford N. Dahm. Differential effects of a catastrophic wildfire on downstream fish assemblages in an aridland river. Aquatic Ecology, 55(2):483–500, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09839-4.

Ranwell:1974:SMT

[Ran74] D. S. Ranwell. The salt marsh to tidal woodland transition. Aquatic Ecology, 8(1-2):139–151, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254914.

Roy:2024:DDS

[RAN24] Mousumi Roy, Gautam Aditya, and Nepal Chandra Nandi. Drivers of the diversity and spatial heterogeneity of aquatic snails in estuarine habitats: evidence from West Bengal in India. Aquatic Ecology, 58(4):1201–1216, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10135-0.

Rao:2010:CET

[Rao10] T. S. Rao. Comparative effect of temperature on biofilm formation in natural and modified marine environment. Aquatic Ecology, 44(2):463–478, June 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9304-1.

Rocha:2024:WAC

[RAO^{+}24] R. M. Rocha, F. Azevedo, U. Oliveira, M. N. M. Cardoso, P. H. B. Clerier, R. R. Fortes, E. A. P. Lopes-Filho, M. L. Lorini, L. S. Miranda, R. B. Moura, A. R. Senna, F. M. Silva, S. N. Stampar, and V. Venekey. West Atlantic coastal marine biodiversity: the contribution of the platform iNaturalist. Aquatic Ecology, 58(1):57–71, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10062-6.

Roveta:2022:SCE

[RAP22] Camilla Roveta, Anna Annibaldi, and Stefania Puce. Single and combined effects of two trace elements (Cd and Cu) on the asexual reproduction of Aurelia sp. polyps. Aquatic Ecology, 56(3):631–637, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09940-8.

Rawi:2013:LEF

[RASA13] Che Salmah Md Rawi, Salman Abdo Al-Shami, and Abu Hassan Ahmad. Local effects of forest fragmentation on diversity of aquatic insects in tropical forest streams: implications for biological conservation. Aquatic Ecology, 47(1):75–85, March 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9426-8.

Rattan:2014:IBC

[Rat14] K. J. Rattan. An inter-basin comparison of nutrient limitation and the irradiance response of pulse-amplitude modulated (PAM) fluorescence in Lake Erie phytoplankton. Aquatic Ecology, 48(1):107–125, March 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9470-7.

Rattan:2017:CAP

[Rat17] Kim J. Rattan. Comparative analyses of physiological assays and chlorophyll-a variable fluorescence parameters: investigating the importance of phosphorus availability in oligotrophic and eutrophic freshwater systems. Aquatic Ecology, 51(3):359–375, September 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9622-7.

Ravera:2001:CBD

[Rav01] Oscar Ravera. A comparison between diversity, similarity and biotic indices applied to the macroinvertebrate community of a small stream: the Ravella river (Como Province, Northern Italy). Aquatic Ecology, 35(2):97–107, June 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011433813309.

Ruxton:1978:EAE

[RB78] T. D. Ruxton and A. C. J. Baker. Engineering aspects of estuarial water storage schemes in the U.K. Aquatic Ecology, 12(3-4):277–290, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02259189.

Rietveld:1982:QAM

[RB82] W. Rietveld and B. Beltman. A qualitative analysis of macrofaunasampling in ditches. Aquatic Ecology, 16(2-3):149–157, December 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255368.

Romeyn:1983:FSC

[RB83] K. Romeyn and L. A. Bouwman. Food selection and consumption by estuarine nematodes. Aquatic Ecology, 17(2): 103–109, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02280819.

Radwell:2008:BMA

[RB08] Andrea J. Radwell and Arthur V. Brown. Benthic meiofauna assemblage structure of headwater streams: density and distribution of taxa relative to substrate size. Aquatic Ecology, 42(3):405–414, September 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9108-0.

Rato:2024:FRE

[RBAB24] João Rato, Pedro Brandão, Pedro M. Anastácio, and Filipe Banha. First records of Mauremys sinensis in Portugal: a consequence of inadequate policies applied to the exotic pet market. Aquatic Ecology, 58(4):1091–1096, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10125-2.

Rani-Borges:2021:MFM

[RBMCP21] Bárbara Rani-Borges, Viviane Moschini-Carlos, and Marcelo Pompêo. Microplastics and freshwater microalgae: what do we know so far? Aquatic Ecology, 55(2):363–377, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09834-9.

Rodibaugh:2020:PCQ

[RBN20] Kelly J. Rodibaugh, Jesse C. Becker, and Weston H. Nowlin. Physicochemical and carbon quantity–quality gradients equally influence bacterial carbon metabolism across an arid riverscape. Aquatic Ecology, 54(3):677–696, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09767-9.

Rasmussen:2011:SEP

[RBPF11] Jes Jessen Rasmussen, Annette Baattrup-Pedersen, and Nikolai Friberg. Stream ecosystem properties and processes along a temperature gradient. Aquatic Ecology, 45(2):231– 242, May 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-010-9349-1.

Ricci:1998:SDD

[RC98] Claudia Ricci and Manuela Caprioli. Stress during dormancy: effect on recovery rates and life-history traits of anhydrobiotic animals. Aquatic Ecology, 32(4):353–359, December 1998. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009905404868.

Rocha:1995:ISW

[RCB95] C. Rocha, G. Cabeçadas, and M. J. Brogueira. On the importance of sediment-water exchange processes of ammonia to primary production in shallow areas of the Sado estuary (Portugal). Aquatic Ecology, 29(3-4):265–273, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02084224.

Roget:1994:CFL

[RCL94] E. Roget, X. Casamitjana, and J. E. Llebot. Calculation of the flow into a lake from underground springs using sedimentation rates. Aquatic Ecology, 28(2):135–141, June 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02333983.

Rieger:1993:CIA

[RCSF93] E.-M. Rieger, B. Claus, and I. Schmitz-Feuerhake. Cesium isotope-activity in the Weser ecosystem three years after the disaster at Chernobyl. Aquatic Ecology, 27(2-4):197–203, June 1993. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334783.

Roos:1991:OSD

[RCV91] Pieter J. Roos, Trudie Crommentuijn, and Renate Velzeboer. Occurrence, succession and distribution of chroöcoccal cyanobacteria in Lake Maarsseveen I. Aquatic Ecology, 24 (2):119–123, April 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260428.

Ruitenburg:1977:TSS

[RD77] G. J. Ruitenburg and C. Davids. Thecamoeba in succession series of some peat-bog vegetations. Aquatic Ecology, 11(1):22–24, April 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02282212.

Rybarczyk:1993:DNF

[RDE93] H. Rybarczyk, M. Desprez, and B. Elkaim. Dynamics of nutrients and faecal bacteria in a macrotidal estuary, the Bay of Somme (France). Aquatic Ecology, 27(2-4):395–404, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334801.

Rijkeboer:1986:CPP

[RDG86] Machteld Rijkeboer, Wim A. De Kloet, and Herman J. Gons. A comparison of primary production measurements using two laboratory systems with differences in light quality. Aquatic Ecology, 20(1-2):93–99, November 1986. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291154.

Rijkeboer:1997:SIR

[RDG97] Machteld Rijkeboer, Arnold G. Dekker, and Herman J. Gons. Subsurface irradiance reflectance spectra of inland waters differing in morphometry and hydrology. Aquatic Ecology, 31(3):313–323, September 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009916501492.

Ramos:2022:ISE

[RdMJN22] Eliêta A. Ramos, Cláudio Simões de Morais-Junior, and José L. C. Novaes. Influence of spatial and environmental factors on the structure of a zooplankton metacommunity in an intermittent river. Aquatic Ecology, 56(1):239–249, March 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-021-09912-y.

Roijackers:1992:R

[RDS92] R. Roijackers, C. Davids, and P. Schroevers. Reviews. Aquatic Ecology, 25(3):295–309, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270815.

Rip:1992:RBN

[REH92] W. J. Rip, K. Everards, and A. Houwers. Restoration of botshol (The Netherlands) by reduction of external nutrient load: The effects on physico-chemical conditions, plankton and sessile diatoms. Aquatic Ecology, 25(3):275–286, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02270813.

Reijnders:1992:HPP

[Rei92] Peter J. H. Reijnders. Harbour porpoises Phocoena phocoena in the North Sea: Numerical responses to changes in environmental conditions. Aquatic Ecology, 26(1):75–85, November 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02298029.

Reimer:1999:IGR

[Rei99] Olof Reimer. Increased gonad ratio in the blue mussel, Mytilus edulis, exposed to starfish predators. Aquatic Ecology, 33(2):185–192, September 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1023/A:1009946526175.

Repko:1979:P

[Rep79] F. F. Repko. Publications. Aquatic Ecology, 13(1):69–70, April 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260312.

Ramdani:2001:OWZ

[REP01] Mohammed Ramdani, Najat Elkhiati, and Simon T. Patrick. Open water zooplankton communities in North African wetland lakes: the CASSARINA Project. Aquatic Ecology, 35 (3-4):319–333, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1011926310469.

Ramdani:2001:ZCO

[RFF01] Mohammed Ramdani, Roger J. Flower, and Adel A. Fathi. Zooplankton (Cladocera, Ostracoda), Chironomidae and other benthic faunal remains in sediment cores from nine North African wetland lakes: the CASSARINA Project. Aquatic Ecology, 35(3-4):389–403, October 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011965226399.

Rodrigues-Filho:2020:MOI

[RFGLSB20] Carlos A. S. Rodrigues-Filho, Ronaldo C. Gurgel-Lourenço, and Jorge I. Sánchez-Botero. Metacommunity organization in an intermittent river in Brazil: the importance of riverine networks for regional biodiversity. Aquatic Ecology, 54 (1):145–161, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09732-1.

Ruiz:1994:SPM

[RFO94] A. Ruiz, J. Franco, and E. Orive. Suspended particulate matter dynamics in the shallow mesotidal urdaibai estuary (Bay of Biscay, Spain). Aquatic Ecology, 28(3-4):309–316, September 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334199.

Ramdani:2001:NAW

[RFP01] Mohammed Ramdani, Roger J. Flower, and Simon T. Patrick. North African wetland lakes: characterization of nine sites included in the CASSARINA Project. Aquatic Ecology, 35(3-4):281–302, October 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/A:
1011957324901.

Ringelberg:1991:DVM

[RFR91] J. Ringelberg, B. J. G. Flik, and K. Royackers. Diel vertical migration of Eudiaptomus gracilis during a short summer period. Aquatic Ecology, 25(1):77–84, October 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02259593.

Rossini:2017:SVE

[RFW17] Renee A. Rossini, Rod J. Fensham, and Gimme H. Walter. Spatiotemporal variance of environmental conditions in Australian artesian springs affects the distribution and abundance of six endemic snail species. Aquatic Ecology, 51(4): 511–529, December 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9633-4.

Riccardi:2007:RLH

[RG07] Nicoletta Riccardi and Gianluigi Giussani. The relevance of life-history traits in the establishment of the invader Eudiaptomus gracilis and the extinction of Eudiaptomus padanus in Lake Candia (Northern Italy): evidence for competitive exclusion? Aquatic Ecology, 41(2):243–254, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9061-3.

Rogozin:2010:SGB

[RGD10] D. Y. Rogozin, S. N. Genova, and A. G. Degermendzhy. Some generalizations based on stratification and vertical mixing in meromictic Lake Shira, Russia, in the period 2002– 2009. Aquatic Ecology, 44(3):485–496, September 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9328-6.

Ramos:2021:CDR

[RGF21] Sandra M. Ramos, Manuel A. S. Graça, and Verónica Ferreira. A comparison of decomposition rates and biological colonization of leaf litter from tropical and temperate origins. Aquatic Ecology, 55(3):925–940, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09872-3.

Rejmankova:2000:USP

[RHA00] Eliska Rejmankova, Richard M. Higashi, and Richard G. Andre. The use of Solid Phase MicroExtraction (SPME) devices in analysis for potential mosquito oviposition attractant chemicals from cyanobacterial mats. Aquatic Ecology, 34(4):413–420, December 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1011416006562.

Ruehle:2017:HPA

[RHH17] Brandon P. Ruehle, Kristin K. Herrmann, and Christopher L. Higgins. Helminth parasite assemblages in two cyprinids with different life history strategies. Aquatic Ecology, 51(2):247–256, June 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
017-9614-7.

Richardson:2002:EIW

[RHL02] Stephen M. Richardson, J. Mark Hanson, and Andrea Locke. Effects of impoundment and water-level fluctuations on macrophyte and macroinvertebrate communities of a dammed tidal river. Aquatic Ecology, 36(4):493–510, December 2002. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1021137630654.

Reinhold:1999:TMS

[RHO99] J. O. Reinhold, A. J. Hendriks, and M. Ohm. Transfer of microcontaminants from sediment to chironomids, and the risk for the pond bat Myotis dasycneme (Chiroptera) preying on them. Aquatic Ecology, 33(4):363–376, December 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009958028204.

Rey:2009:ESM

[RHO09] Jorge R. Rey, Paul E. Hargraves, and Sheila M. O’Connell. Effect of selected marine and freshwater microalgae on development and survival of the mosquito Aedes aegypti. Aquatic Ecology, 43(4):987–997, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9232-0.

Robertson:2018:SSB

[RHS18] M. D. Robertson, M. F. Hernandez, and C. D. Suski. Shelterseeking behavior of crayfish, Procambarus clarkii, in elevated carbon dioxide. Aquatic Ecology, 52(2-3):225–233, September 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-018-9657-4.

Richter:1986:BFW

[Ric86] A. F. Richter. Biomanipulation and its feasibility for water quality management in shallow eutrophic water bodies in The Netherlands. Aquatic Ecology, 20(1-2):165–172, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291160.

Riisgaard:1994:FFP

[Rii94] Hans Ulrik Riisgård. Filter-feeding in the polychaete Nereis diversicolor: a review. Aquatic Ecology, 28(3-4):453–458, September 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334216.

Ringelberg:1991:P

[RIMG91] J. Ringelberg, A. Infante, and S. Montenegro-Guillen. Preface. Aquatic Ecology, 25(2):102, June 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291237.

Ringelberg:1970:GVD

[Rin70a] J. Ringelberg. De giftigheid van diuron voor Daphnia magna. (Dutch) [The toxicity of diuron to Daphnia magna]. Aquatic Ecology, 4(1):29–33, March 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185863.

Ringelberg:1970:TVD

[Rin70b] J. Ringelberg. Het temperatuuroptimum van Daphnia magna. (Dutch) [The temperature optimum of Daphnia magna]. Aquatic Ecology, 4(3):153–156, November 1970. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185879.

Ringelberg:1970:IDI

[Rin70c] J. Ringelberg. Inleiding. (Dutch) [Introduction]. Aquatic Ecology, 4(3):144–145, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF03185875.

Ringelberg:1973:PDT

[Rin73] J. Ringelberg. Parameter dependant (temperature) tolerance levels and the influence of the complexity of the biological system. Aquatic Ecology, 7(3):106–114, September 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02275611.

Ringelberg:1976:PNK

[Rin76] J. Ringelberg. The possibilities of a new kind of microecosystem in aquatic ecosystem research. Aquatic Ecology, 10(1):17–18, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02308494.

Ringelberg:1980:EIP

[Rin80] J. Ringelberg. Eutrophication: Introduction to the process and some ecological implications. Aquatic Ecology, 14 (1-2):30–35, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260270.

Ringelberg:1981:DSW

[Rin81a] J. Ringelberg. A diel study in a water column of Lake Maarsseveen I. Aquatic Ecology, 15(1-2):60–71, October 1981. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260259.

Ringelberg:1981:GIR

[Rin81b] J. Ringelberg. General introduction to the research in the Maarsseveen Lakes. Aquatic Ecology, 15(1-2):3–4, October 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260253.

Ringelberg:1981:IRA

[Rin81c] J. Ringelberg. Introduction to the research area. Aquatic Ecology, 15(1-2):5–9, October 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260254.

Ringelberg:1993:GDB

[Rin93] J. Ringelberg. The growing difference between limnology and aquatic ecology. Aquatic Ecology, 27(1):11–19, July 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02336925.

Ringelberg:1997:TEI

[Rin97] J. Ringelberg. Two examples of the interplay between field observations and laboratory experiments from 35 years of research with planktonic organisms. Aquatic Ecology, 31(1):9– 17, March 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009984116733.

Rantala:2004:MSA

[RIT04] Markus J. Rantala, Jari Ilmonen, and Katja Tynkkynen. The macrophyte, Stratiotes aloides, protects larvae of dragonfly Aeshna viridis against fish predation. Aquatic Ecology, 38(1):77–82, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000021005.22624.16.

Roijackers:1996:TSS

[RJ96] R. M. M. Roijackers and A. M. T. Joosten. The trophic state of shallow lakes in The Netherlands. Aquatic Ecology, 30(2-3):219–226, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02272241.

Reavie:2011:ESS

[RJ11] Euan D. Reavie and Steve Juggins. Exploration of sample size and diatom-based indicator performance in three North American phosphorus training sets. Aquatic Ecology, 45(4): 529–538, November 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9373-9.

Rodriguez:2007:MMV

[RJD07] Rocío Ramírez Rodríguez, Javier Carmona Jiménez, and Carlos Martorell Delgado. Microhabitat and morphometric variation in two species of Prasiola (Prasiolales, Chlorophyta) from streams in central Mexico. Aquatic Ecology, 41(2):161–168, June 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-006-9068-9.

Rao:2002:PPS

[RK02] T. Ramakrishna Rao and Ram Kumar. Patterns of prey selectivity in the cyclopoid copepod Mesocyclops thermocyclopoides. Aquatic Ecology, 36(3):411–424, August 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1016509016852.

Ruiz:2022:ECF

[RKLB22] Thomas Ruiz, Apostolos-Manuel Koussoroplis, Delphine Latour, and Alexandre Bec. The energetic cost of facing cyanotoxins: a case study on Daphnia magna. Aquatic Ecology, 56(4):??, November 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09990-6.

Ruiz:2023:ECF

[RKLB23] Thomas Ruiz, Apostolos-Manuel Koussoroplis, Delphine Latour, and Alexandre Bec. The energetic cost of facing cyanotoxins: a case study on Daphnia magna. Aquatic Ecology, 57(1):15–20, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09990-6.

Rasanen:2014:STE

[RKS14] N. Räsänen, P. Kankaala, and S. Saarnio. Short-term effects of phosphorus addition and pH rise on bacterial utilization and biodegradation of dissolved organic carbon (DOC) from boreal mires. Aquatic Ecology, 48(4):435–446, December 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-014-9496-x.

Riisgaard:2005:WPA

[RL05] Hans Ulrik Riisgård and Poul S. Larsen. Water pumping and analysis of flow in burrowing zoobenthos: an overview. Aquatic Ecology, 39(2):237–258, June 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-1916-x.

Rijstenbil:1981:IMA

[RM81] J. W. Rijstenbil and A. G. A. Merks. The influence of microalgae on the oxygen dynamics in a brackish ditch. Aquatic Ecology, 15(3):123–135, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255171.

Ricci:1998:DMM

[RM98] Claudia Ricci and Giulio Melone. Dwarf males in monogonont rotifers. Aquatic Ecology, 32(4):361–365, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009986106280.

Riccardi:2000:SVC

[RM00] Nicoletta Riccardi and Luca Mariotto. Seasonal variations in copepod body length: a comparison between different species in the Lagoon of Venice. Aquatic Ecology, 34(3):243–252, September 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009971507797.

Rocha:1995:RMM

[RMC95] C. Rocha, M. J. Madureira, and M. Caetano. Role of microorganisms in mineralization processes in intertidal surface sediments subject to high temperatures: an incubation experiment. Aquatic Ecology, 29(3-4):257–263, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084223.

Rider:2021:MYM

[RMH21] Mitchell J. Rider, Laura H. McDonnell, and Neil Hammerschlag. Multi-year movements of adult and subadult bull sharks (Carcharhinus leucas): philopatry, connectivity, and environmental influences. Aquatic Ecology, 55(2):559–577, June 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09845-6.

Richardson:2022:RCP

[RMH22] Bryce Richardson, Holly Martin, and Jörg Detlef Hardege. The role of changing pH on olfactory success of predator–prey interactions in green shore crabs, Carcinus maenas. Aquatic Ecology, 56(2):409–418, June 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09913-x.

Ramesh:2021:BAT

[RMR21] Chatragadda Ramesh and V. Benno Meyer-Rochow. Bioluminescence in aquatic and terrestrial organisms elicited through various kinds of stimulation. Aquatic Ecology, 55(3): 737–764, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09875-0.

Riera:2011:SYS

[RMS11] Rodrigo Riera, Óscar Monterroso, and Alicia Sacramento. Six-year study of meiofaunal dynamics in fish farms in Tenerife (Canary Islands, NE Atlantic Ocean). Aquatic Ecology, 45(2):221–229, May 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-010-9348-2.

Rijstenbil:1991:PCS

[RMW91] J. W. Rijstenbil, A. G. A. Merks, and J. A. Wijnholds. Phytoplankton composition and spatial distribution of copper and zinc in the Fal Estuary (Cornwall, UK). Aquatic Ecology, 25(1):37–43, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02259587.

Rajagopal:1997:SSS

[RNJ97] S. Rajagopal, K. V. K. Nair, and H. A. Jenner. Seasonal settlement and succession of fouling communities in Kalpakkam, east coast of India. Aquatic Ecology, 30(4):309–325, May 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02085874.

Ruiz-Navarro:2020:INE

[RNJB20] Ana Ruiz-Navarro, Michelle C. Jackson, and J. Robert Britton. Influence of nutrient enrichment on the growth, recruitment and trophic ecology of a highly invasive freshwater fish. Aquatic Ecology, 54(4):1029–1039, December 2020.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09791-9.

Rodusky:2010:ILW

[Rod10] Andrew J. Rodusky. The influence of large water level fluctuations and hurricanes on periphyton and associated nutrient storage in subtropical Lake Okeechobee, USA. Aquatic Ecology, 44(4):797–815, December 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9317-9.

Roelofs:1996:RES

[Roe96] Jan G. M. Roelofs. Restoration of eutrophied shallow softwater lakes based upon carbon and phosphorus limitation. Aquatic Ecology, 30(2-3):197–202, October 1996. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02272239.

Roijackers:1981:CBT

[Roi81a] R. M. M. Roijackers. A comparison between two methods of extracting chlorophyll-a from different phytoplankton samples. Aquatic Ecology, 15(3):179–183, December 1981. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255177.

Roijackers:1981:SCN

[Roi81b] R. M. M. Roijackers. Synuraceae (Chrysophyceae) in the Netherlands. Aquatic Ecology, 15(3):195–196, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255181.

Roijackers:1988:EMT

[Roi88a] R. M. M. Roijackers. External morphology as taxonomic characteristic in planktonic scale-bearing Chrysophyceae and scaled Heliozoa. Aquatic Ecology, 22(1):69–73, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256785.

Roijakers:1988:HRN

[Roi88b] R. M. M. Roijakers. Hydrobiological research in The Netherlands fundamental and applied aspects. Aquatic Ecology, 22(1):65, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256783.

Roos:1979:ADP

[Roo79] P. J. Roos. Architecture and development of periphyton on reed-stems in Lake Maarsseveen. Aquatic Ecology, 13 (2-3):117, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02284745.

Roos:1982:FFD

[Roo82] P. J. Roos. Fate and final distribution of dead reed stems in Lake Maarsseveen. Aquatic Ecology, 16(1):113–114, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255419.

Roskam:1969:IWD

[Ros69a] R. Th. Roskam. Ijsselmeer en waterverontreiniging. (Dutch) [Lake IJssel and water pollution]. Aquatic Ecology, 3(1):9, March 1969. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185841.

Roskam:1969:SAW

[Ros69b] R. Th. Roskam. Soorten afval, die waterverontreiniging teweeg kunnen brengen. (Dutch) [Types of waste that can cause water pollution]. Aquatic Ecology, 3(1):10–11, March 1969. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185842.

Roskam:1971:M

[Ros71] Robert Thierry Roskam. In memoriam. Aquatic Ecology, 5 (4):163–171, December 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185957.

Rosenberg:1992:FBC

[Ros92a] David M. Rosenberg. Freshwater biomonitoring and Chironomidae. Aquatic Ecology, 26(2-4):101–122, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255231.

Rossaro:1992:OMC

[Ros92b] Bruno Rossaro. Ordination methods and chironomid species in stony bottom streams. Aquatic Ecology, 26(2-4):447–456, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255274.

Real:1992:FID

[RP92] Montserrat Real and Narcis Prat. Factors influencing the distribution of chironomids and oligochaetes in profundal areas of Spanish reservoirs. Aquatic Ecology, 26(2-4):405–410, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255269.

Reinart:2008:RTS

[RP08] Anu Reinart and Tiia Pedusaar. Reconstruction of the time series of the underwater light climate in a shallow turbid lake. Aquatic Ecology, 42(1):5–15, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9056-0.

Rodrigues:2022:RBD

[RP22] R. V. Rodrigues and J. S. Patil. Response of benthic dinoflagellates Amphidinium carterae and Bysmatrum gregarium to salinity changes and prolonged darkness: elucidation through laboratory experiments. Aquatic Ecology, 56(4):1113–1126, December 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-022-09960-y.

Rechulicz:2015:ODH

[RPN15] Jacek Rechulicz, Wojciech Płaska, and Dorota Nawrot. Occurrence, dispersion and habitat preferences of Amur sleeper (Perccottus glenii) in oxbow lakes of a large river and its tributary. Aquatic Ecology, 49(3):389–399, September 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9532-5.

Reitsema:2020:ICC

[RPS20] Rosanne E. Reitsema, Stefan Preiner, and Jonas Schoelynck. Implications of climate change for submerged macrophytes: effects of CO_{2}, flow velocity and nutrient concentration on Berula erecta. Aquatic Ecology, 54(3):775–793, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09776-8.

Rivera-Perez:2023:EGA

[RPSSDS23] Juan Mateo Rivera-Pérez, Luis Fernando Salazar-Salinas, and Caroline De-Souza. Effects of green areas and the urbanization process in the natural populations of the endemic freshwater crab Strengeriana fuhrmanni (Decapoda: Pseudothelphusidae) in Colombia. Aquatic Ecology, 57(3): 733–745, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-023-10042-w.

Rodrigues:1993:HBG

[RQ93] Ana Maria J. Rodrigues and Victor M. S. Quintino. Horizontal biosedimentary gradients across the Sado estuary, W. Portugal. Aquatic Ecology, 27(2-4):449–464, June 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334806.

Reede:1998:DLH

[RR98] Tineke Reede and Joop Ringelberg. Differential life history responses of several pelagic Daphnia clones differing in migratory behaviour. Aquatic Ecology, 32(3):245–253, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009968311604.

Reinikainen:2003:PPL

[RR03] Marko Reinikainen and Sari Repka. Phenotypic plasticity in life-history traits and feeding appendages in two species of Daphnia fed a natural phytoplankton assemblage. Aquatic Ecology, 37(4):409–415, October 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000007048.87897.36.

Ribeiro:2015:PCC

[RRK15] Filipe Ribeiro, Kateřina Rylková, and Lukáš Kalous. Prussian carp Carassius gibelio: a silent invader arriving to the Iberian Peninsula. Aquatic Ecology, 49(1):99–104, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9508-5.

Rodrigo:2009:MMS

[RRL09] María A. Rodrigo, Carmen Rojo, and José Larrosa. Mechanisms of microalgae selection during the assembly of a planktonic community. Aquatic Ecology, 43(1):61–72, March 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9143-x.

Rocha-Ramirez:2007:IAA

[RRRRA07] Arturo Rocha-Ramírez, Alejandro Ramírez-Rojas, and Javier Alcocer. Invertebrate assemblages associated with root masses of Eichhornia crassipes (Mart.) Solms-Laubach 1883 in the Alvarado Lagoonal System, Veracruz, Mexico. Aquatic Ecology, 41(2):319–333, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9054-2.

Reshma:2022:TRC

[RRS^{+}22] B. Reshma, B. Rahul, K. R. Sreenath, K. K. Joshi, and George Grinson. Taxonomic resolution of coral image classification with Convolutional Neural Network. Aquatic Ecology, 56(4):??, November 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09988-0.

Reshma:2023:TRC

[RRS^{+}23] B. Reshma, B. Rahul, K. R. Sreenath, K. K. Joshi, and George Grinson. Taxonomic resolution of coral image classification with Convolutional Neural Network. Aquatic Ecology, 57(4):845–861, December 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
022-09988-0.

Ringelberg:1985:DVE

[RS85] J. Ringelberg and J. Steenvoorden. Diel variation in the egg ratio of rotifers throughout the season (preliminary report). Aquatic Ecology, 19(2):153–158, December 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270761.

Rehbehn:1993:DAN

[RSK93] R. Rehbehn, B. Schuchardt, and G. O. Kirst. The distribution of Actinocyclus normanii (Bacillariophyceae) in estuaries: Field observations and laboratory investigations. Aquatic Ecology, 27(2-4):205–214, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02334784.

Riis:2009:TMR

[RSK09] Tenna Riis, Rikke Schultz, and Casper K. Katborg. Transplanting macrophytes to rehabilitate streams: experience and recommendations. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9213-8.

Rathod:2024:ASV

[RSM24] Sachin G. Rathod, Anshika Singh, and Vaibhav A. Mantri. Ascertaining the spatiotemporal variations in seaweed assemblages associated with the endemic seaweed Ulva ovata along selected locations of the Gulf of Khambhat, India. Aquatic Ecology, 58(4):1107–1127, December 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10127-0.

Rohr:2011:EHI

[RTJ11] Nicole E. Rohr, Carol S. Thornber, and Emily Jones. Epiphyte and herbivore interactions impact recruitment in a marine subtidal system. Aquatic Ecology, 45(2):213–219, May 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-010-9347-3.

Ryan:2017:EPT

[RTL17] Caitlin N. Ryan, Mridul K. Thomas, and Elena Litchman. The effects of phosphorus and temperature on the competitive success of an invasive cyanobacterium. Aquatic Ecology, 51(3):463–472, September 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9629-0.

Rossini:2017:CET

[RTW17] Renee A. Rossini, Hannah L. Tibbetts, and Gimme H. Walter. Can environmental tolerances explain convergent patterns of distribution in endemic spring snails from opposite sides of the Australian arid zone? Aquatic Ecology, 51(4): 605–624, December 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9639-y.

Roos:1987:ESS

[RU87] C. Roos and E. J. B. Uunk. Effects of stormwater sewer discharges on the aquatic community in urban canals in Lelystad. Aquatic Ecology, 21(2):207–212, December 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255446.

Ruse:1992:CBC

[Rus92] L. P. Ruse. Correlations between chironomid pupal skin collections and habitats recorded from a chalk river. Aquatic Ecology, 26(2-4):411–417, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255270.

Ryman:2008:MLP

[RVB08] Jennie E. Ryman, Jillian L. A. Van Walleghem, and Paul J. Blanchfield. Methylmercury levels in a parasite (Apophallus brevis metacercariae) and its host, yellow perch (Perca flavescens). Aquatic Ecology, 42(3):495–501, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9104-4.

Ringelberg:2005:DDVa

[RvGB05a] Joop Ringelberg, Erik van Gool, and Michaela Brehm. Density and depth variations of Daphnia multilocus genotypes during a summer period in Lake Maarsseveen. Aquatic Ecology, 38(4):525–535, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
005-5667-0.

Ringelberg:2005:DDVb

[RvGB05b] Joop Ringelberg, Erik van Gool, and Michaela Brehm. Density and depth variations of Daphnia multilocus genotypes during a summer period in Lake Maarsseveen. Aquatic Ecology, 38(4):525–535, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
004-5667-5.

Rajagopal:1998:SGG

[RVJ98] S. Rajagopal, V. P. Venugopalan, and H. A. Jenner. Settlement and growth of the green mussel Perna viridis (L.) in coastal waters: influence of water velocity. Aquatic Ecology, 32(4):313–322, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009941714978.

Rajagopal:2006:GCR

[RVJ06] S. Rajagopal, V. P. Venugopalan, and H. A. Jenner. Greening of the coasts: a review of the Perna viridis success story. Aquatic Ecology, 40(3):273–297, September 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9032-8.

Rajagopal:1996:ETS

[RVK96] S. Rajagopal, G. Van Der Velde, and A. J. Kempers. Effects of temperature, salinity and agitation on byssus thread formation of zebra mussel Dreissena polymorpha. Aquatic Ecology, 30(2-3):187–195, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02272238.

Repka:1998:VNS

[RW98] Sari Repka and Mari Walls. Variation in the neonate size of Daphnia pulex: the effects of predator exposure and clonal origin. Aquatic Ecology, 32(3):203–209, October 1998. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009978527633.

Rogozin:2009:EWC

[RZG09] D. Y. Rogozin, V. V. Zykov, and R. D. Gulati. Effect of winter conditions on distributions of anoxic phototrophic bacteria in two meromictic lakes in Siberia, Russia. Aquatic Ecology, 43(3):661–672, September 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9270-7.

Sprung:1995:DEE

[SA95] Martin Sprung and Harald Asmus. Does the energy equivalence rule apply to intertidal macrobenthic communities? Aquatic Ecology, 29(3-4):369–376, December 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084236.

Sanvicente-Anorve:2009:SRS

[SAACR09] Laura Sanvicente-Añorve, Claudia Alba, and Manuel Castillo-Rivera. Siphonophores off a riverine system in the southern Gulf of Mexico: factors affecting their distribution and spatial niche breadth and overlap. Aquatic Ecology, 43 (2):423–435, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9172-0.

Sa-Ardrit:2005:CDA

[SAB05] Phannee Sa-Ardrit and Frederick William H. Beamish. Cladocera diversity, abundance and habitat in a Western Thailand stream. Aquatic Ecology, 39(3):353–365, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-0783-4.

Schoo:2010:DNS

[SAB10] Katherina L. Schoo, Nicole Aberle, and Maarten Boersma. Does the nutrient stoichiometry of primary producers affect the secondary consumer Pleurobrachia pileus? Aquatic Ecology, 44(1):233–242, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9265-4.

Saether:1992:HBS

[Sæt92] O. A. Sæther. Heterotrissocladius boltoni sp.n., a new orthoclad from vernal pools and streams in Ohio, U.S.A. (Diptera: Chironomidae). Aquatic Ecology, 26(2-4):191–196, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255241.

Sultana:2010:MRS

[SAF10] Munira Sultana, Takashi Asaeda, and Takeshi Fujino. Morphological responses of a submerged macrophyte to epiphyton. Aquatic Ecology, 44(1):73–81, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9291-2.

Sarma:2007:CEZ

[SAGN07] S. S. S. Sarma, Rubén Azuara-García, and S. Nandini. Combined effects of zinc and algal food on the competition between planktonic rotifers, Anuraeopsis fissa and Brachionus rubens (Rotifera). Aquatic Ecology, 41(4):631–638, December 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9120-4.

Saha:2010:OFH

[SAH10] Nabaneeta Saha, Gautam Aditya, and Stephanie E. Hampton. Opportunistic foraging by heteropteran mosquito predators. Aquatic Ecology, 44(1):167–176, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9250-y.

Silkin:2016:MRI

[SAL16] V. A. Silkin, A. I. Abakumov, and A. V. Lifanchuk. Mechanisms of regulation of invasive processes in phytoplankton on the example of the north-eastern part of the Black Sea. Aquatic Ecology, 50(2):221–234, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9570-7.

Sunardi:2007:PRT

[SAM07] Sunardi, Takashi Asaeda, and Jagath Manatunge. Physiological responses of topmouth gudgeon, Pseudorasbora parva, to predator cues and variation of current velocity. Aquatic Ecology, 41(1):111–118, March 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9048-0.

Santema:1976:I

[San76] P. Santema. Introduction. Aquatic Ecology, 10(3):127–128, December 1976. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02263360.

Sarma:1999:CBB

[SAN99] S. S. S. Sarma, Mario Alfredo Fernández Araiza, and S. Nandini. Competition between Brachionus calyciflorus Pallas and Brachionus patulus (Müller) (Rotifera) in relation to algal food concentration and initial population density. Aquatic Ecology, 33(4):339–345, December 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009912816400.

Sarma:2005:BRa

[Sar05a] S. S. S. Sarma. Book review. Aquatic Ecology, 38(4):613, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-1173-7.

Sarma:2005:BRb

[Sar05b] S. S. S. Sarma. Book review. Aquatic Ecology, 38(4):614, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-1172-8.

Sarma:2005:BRc

[Sar05c] S. S. S. Sarma. Book review. Aquatic Ecology, 38(4):615, January 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-1171-9.

Sarma:2005:BRG

[Sar05d] S. S. S. Sarma. Book review: a guide to tropical freshwater zooplankton. identification, ecology and impact on fisheries. Aquatic Ecology, 38(4):613, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-1173-z.

Sarma:2005:BRSa

[Sar05e] S. S. S. Sarma. Book review: Sustainable aquaculture. Aquatic Ecology, 38(4):614, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-1172-0.

Sarma:2005:BRSb

[Sar05f] S. S. S. Sarma. Book review: Sustainable aquaculture. food for the future? Aquatic Ecology, 38(4):615, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-1171-1.

Skein:2021:CNP

[SAR21] Lisa Skein, Mhairi E. Alexander, and Tamara B. Robinson. Characteristics of native predators are more important than those of alien prey in determining the success of biotic resistance in marine systems. Aquatic Ecology, 55(1):97–106, March 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09814-5.

SampaioFranco:2024:TVH

[SASA^{+}24] Ana Clara Sampaio Franco, Valter M. Azevedo-Santos, Marluce Aparecida Mattos de Paula Nogueira, Tommaso Giarrizzo, Rachel Ann Hauser-Davis, Erick Cristofore Guimarães, Roger Henrique Dalcin, Marcelo Soeth, Matheus Oliveira Freitas, Áthila A. Bertoncini, Vinícius Abilhoa, Almir Manoel Cunico, Johnatas Adelir-Alves, Bianca Bentes, José Luís Costa Novaes, Mauricio Hostim-Silva, Jonas Rodrigues Leite, Vagner Leonardo Macêdo dos Santos, and Jean Ricardo Simões Vitule. Tilapia venturing into high-salinity environments: a cause for concern? Aquatic Ecology, 58(1):47–55, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s10452-023-10069-z. See correction [FASdPN^{+}25].

Saeijs:1978:ECC

[SB78] H. L. F. Saeijs and B. A. Bannink. Environmental consideration in a coastal engineering project the Delta Project in the south-western Netherlands. Aquatic Ecology, 12(3-4): 180–202, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259182.

Seminara:1988:TLA

[SB88] Marco Seminara and Marcello Bazzanti. Trophic level assessment of profundal sediments of the artificial Lake Campotosto (Central Italy), using midge larval community (Diptera: Chironomidae). Aquatic Ecology, 22(2):183–193, December 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02256822.

Spaak:1997:TSL

[SB97] Piet Spaak and Maarten Boersma. Tail spine length in the Daphnia galeata complex: costs and benefits of induction by fish. Aquatic Ecology, 31(1):89–98, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009935100804.

Simcic:2002:IMP

[SB02] Tatjana Simčič and Anton Brancelj. Intensity of mineralization processes in mountain lakes in NW Slovenia. Aquatic Ecology, 36(3):345–354, August 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1016557714341.

Shostell:2004:SIV

[SB04] Joseph Shostell and Paul A. Bukaveckas. Seasonal and interannual variation in nutrient fluxes from tributary inputs, consumer recycling and algal growth in a eutrophic river impoundment. Aquatic Ecology, 38(3):359–373, September 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/B:AECO.0000035167.67399.63.

Salmaso:2008:FCS

[SB08] Nico Salmaso and Maria G. Braioni. Factors controlling the seasonal development and distribution of the phytoplankton community in the lowland course of a large river in Northern Italy (River Adige). Aquatic Ecology, 42(4):533–545, December 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9135-x.

Soana:2014:SRN

[SB14] Elisa Soana and Marco Bartoli. Seasonal regulation of nitrification in a rooted macrophyte (Vallisneria spiralis L.) meadow under eutrophic conditions. Aquatic Ecology, 48 (1):11–21, March 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-013-9462-z.

Suominen:2017:CBT

[SBH17] Saara Suominen, Verena S. Brauer, and Teppo Hiltunen. Competition between a toxic and a non-toxic Microcystis strain under constant and pulsed nitrogen and phosphorus supply. Aquatic Ecology, 51(1):117–130, March 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9603-2.

Soares:2023:PAR

[SBLdAM23] Bruno Eleres Soares, Naraiana Loureiro Benone, Cleonice Maria Cardoso Lobato, and Luciano Fogaça de Assis Montag. Pasture areas reduce the abundance and trophic niche width, but not the body condition of the Amazonian whale catfish. Aquatic Ecology, 57(1):187–198, March 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10006-0.

Sueiro:2012:RPS

[SBS12] María Cruz Sueiro, Alejandro Bortolus, and Evangelina Schwindt. The role of the physical structure of Spartina densiflora Brong. in structuring macroinvertebrate assemblages. Aquatic Ecology, 46(1):25–36, March 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9379-3.

Society:1983:C

[SC83] Hydrobiological Society and J. Coosen. Congresses. Aquatic Ecology, 17(2):179, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02280826.

Serodio:1999:FLT

[SC99] João Serôdio and Fernando Catarino. Fortnightly light and temperature variability in estuarine intertidal sediments and implications for microphytobenthos primary productivity. Aquatic Ecology, 33(3):235–241, October 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009989229098.

Sharma:2000:SOF

[SC00] Jai Gopal Sharma and Rina Chakrabarti. Seasonal occurrence of freshwater medusa Limnocnida indica annandale (Cnidaria: Limnomedusae) in a lake associated with the river Yamuna, India. Aquatic Ecology, 34(2):205–207, June 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009986516652.

Spaccesi:2009:BIA

[SC09] Fernando Spaccesi and Alberto Rodrigues Capítulo. Benthic invertebrate assemblage in Samborombón River (Argentina, S. America), a brackish plain river. Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-008-9212-9.

Silva-Costa:2013:FEK

[SCB13] Augusto Silva-Costa and Leandro Bugoni. Feeding ecology of Kelp Gulls (Larus dominicanus) in marine and limnetic environments. Aquatic Ecology, 47(2):211–224, June 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9436-1.

Sreekanth:2019:ADS

[SCC19] G. B. Sreekanth, S. K. Chakraborty, and E. B. Chakurkar. Application of deterministic and stochastic geo-statistical tools for analysing spatial patterns of fish density in a tropical monsoonal estuary. Aquatic Ecology, 53(1):49–60, March 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09672-w.

Sanchez-Carmona:2012:FWS

[SCEGL12] R. Sánchez-Carmona, L. Encina, and C. Granado-Lorencio. Food web structure in Mediterranean streams: exploring stabilizing forces in these ecosystems. Aquatic Ecology, 46(3): 311–324, September 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9400-5.

Schlacher:2009:CEO

[SCG09] Thomas A. Schlacher, Rod M. Connolly, and Troy F. Gaston. Can export of organic matter from estuaries support zooplankton in nearshore, marine plumes? Aquatic Ecology, 43(2):383–393, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9196-5.

Schroevers:1968:IVA

[Sch68a] P. J. Schroevers. De invloed van de Arembergergracht op het water van de Dirkswijde en de Vossenbelt. (Dutch) [The influence of the Arembergergracht on the water of the Dirkswijde and the Vossenbelt]. Aquatic Ecology, 2(2):67–79, June 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185812.

Schroevers:1968:IDI

[Sch68b] P. J. Schroevers. Inleiding. (Dutch) [Introduction]. Aquatic Ecology, 2(2):42–44, June 1968. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185808.

Schroevers:1972:IDI

[Sch72a] P. J. Schroevers. Inleiding. (Dutch) [Introduction]. Aquatic Ecology, 6(2):48–50, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02304093.

Schroevers:1972:WVW

[Sch72b] P. J. Schroevers. Waardering van wateren in de Ooypolder aan fytoplanktonwaarnemingen. (Dutch) [Valuation of waters in the Ooypolder of phytoplankton observations]. Aquatic Ecology, 6(2):69–88, June 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02304097.

Schroevers:1974:BAW

[Sch74] P. J. Schroevers. Biological assessment of water as a fundamental problem. Aquatic Ecology, 8(3):270–273, December 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257502.

Schroder:1977:DSC

[Sch77] H. G. J. Schröder. Distribution of sulphur-cycle bacteria in the Ems–Dollard estuary. Aquatic Ecology, 11(1):14, April 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02282207.

Schreurs:1978:ACM

[Sch78] W. Schreurs. An automated colorimetric method for the determination of dissolved organic carbon in sea water by U.V. destruction. Aquatic Ecology, 12(2):137–142, July 1978. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260715.

Schroevers:1979:M

[Sch79] P. J. Schroevers. In memoriam. Aquatic Ecology, 13(1):71– 72, April 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260314.

Schroevers:1983:BAF

[Sch83] P. J. Schroevers. Biological assessment — the fairy-tale of the magic salve. Aquatic Ecology, 17(1):71–75, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255194.

Schroevers:1985:M

[Sch85] P. Schroevers. In memoriam. Aquatic Ecology, 19(2):226–230, December 1985. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02270771.

Schroevers:1988:BBW

[Sch88] P. J. Schroevers. The baby and the bath-water. Thoughts about ‘saprobity’. Aquatic Ecology, 22(1):79–80, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02256787.

Scheffer:1989:ASS

[Sch89] M. Scheffer. Alternative stable states in eutrophic, shallow freshwater systems: a minimal model. Aquatic Ecology, 23(1):73–83, March 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02286429.

Schmid:1992:HPP

[Sch92] P. E. Schmid. Habitat preferences as patch selection of larval and emerging chironomids (Diptera) in a gravel brook. Aquatic Ecology, 26(2-4):419–429, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255271.

Schories:1995:SES

[Sch95] Dirk Schories. Sporulation of Enteromorpha spp. (Chlorophyta) and overwintering of spores in sediments of the Wadden Sea, Island Sylt, North Sea. Aquatic Ecology, 29(3-4): 341–347, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02084233.

Scheffer:1997:IPA

[Sch97a] Marten Scheffer. On the implications of predator avoidance. Aquatic Ecology, 31(1):99–107, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009925025185.

Schiedek:1997:MCWa

[Sch97b] Doris Schiedek. Marenzelleria cf. wireni (Polychaeta: Spionidae) — ecophysiological adaptations to a life in the coastal waters of the Baltic Sea. Aquatic Ecology, 31(2):199–210, June 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009907606161.

Santini:2024:DST

[SCN^{+}24] Rachel Santini, Mirela Vantini Checchio, Laís Samira Correia Nunes, Priscila Lupino Gratão, and Antonio Fernando Monteiro Camargo. Do salinity, total nitrogen and phosphorus variation induce oxidative stress in emergent macrophytes along a tropical estuary? Aquatic Ecology, 58(2):399–409, June 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10079-x.

Steinman:2009:STV

[SCO09] Alan Steinman, Xuefeng Chu, and Mary Ogdahl. Spatial and temporal variability of internal and external phosphorus loads in Mona Lake, Michigan. Aquatic Ecology, 43(1):1– 18, March 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9147-6.

Schindler:2012:MOL

[SCS12] Daniel E. Schindler, Jackie L. Carter, and Dale C. Sebastian. Mysis in the Okanagan Lake food web: a time-series analysis of interaction strengths in an invaded plankton community. Aquatic Ecology, 46(2):215–227, June 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9393-0.

Stock:1983:P

[SD83] J. H. Stock and Jaap Dorgelo. Publications. Aquatic Ecology, 17(2):179–181, December 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02280827.

Seyhan:1986:ARS

[SD86] Ersin Seyhan and Arnold Dekker. Application of remote sensing techniques for water quality monitoring. Aquatic Ecology, 20(1-2):41–50, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02291149.

Saikia:2009:FEC

[SD09] S. K. Saikia and D. N. Das. Feeding ecology of common carp (Cyprinus carpio L.) in a rice–fish culture system of the Apatani plateau (Arunachal Pradesh, India). Aquatic Ecology, 43(2):559–568, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9174-y.

Storck:2025:AAA

[SdAdC^{+}25] Tamiris Rosso Storck, Aline Monique Blank do Amaral, Taisson Kroth Thomé da Cruz, Silvana Isabel Schneider, Felipe Osmari Cerezer, Júlia Antunes de Oliveira, Renato Zanella, Osmar Damian Prestes, Sandro José Giacomini, Vania Lucia Loro, Jaqueline Ineu Golombieski, and Barbara Clasen. Anthropogenic actions alter the aquatic environment quality: biomonitoring study of a river in Southern Brazil. Aquatic Ecology, 59(1):185–202, March 2025. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10156-9.

SampaioDeSouza:2023:IGB

[SDD^{+}23] Christiane Sampaio De Souza, Paulo De Oliveira Mafalda Junior, Laura Rodrigues Da Conceição, Ruy Kenji Papa De Kikuchi, and José Maria Landim Dominguez. The impact of Brazilian coast oil spill on the mesozooplankton of the tropical narrowest continental shelf. Aquatic Ecology, 57(2):481– 498, June 2023. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10024-y.

SitSit:2025:WQB

[SDN25] Fernie D. SitSit, Romeo M. Dizon, and Jones T. Napaldet. Water quality of Balili River before, during and after COVID: lessons learned and ways forwards for tropical river remediation. Aquatic Ecology, 59(1):107–124, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10151-0.

Schmidt-Drewello:2016:RBI

[SDRM16] Alexander Schmidt-Drewello, H. Wolfgang Riss, and Elisabeth I. Meyer. Relative benefit of the invasive Echinogammarus berilloni (Catta, 1878) over native gammarids under fish predation (Gasterosteus aculeatus Linnaeus, 1758). Aquatic Ecology, 50(1):75–85, March 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9555-y.

Smejkal:2024:CFR

[ŠDT^{+}24a] Marek Šmejkal, Ondřej Dočkal, Kiran Thomas, Chandani R. Verma, Pradeep Kumkar, and Lukáš Kalous. Correction: First record of highly invasive Chinese sleeper Perccottus glenii Dybowski, 1877 (Perciformes: Odontobutidae) in the Elbe River Basin, Czechia. Aquatic Ecology, 58(2):531, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10043-9. See [ŠDT^{+}24b].

Smejkal:2024:FRH

[ŠDT^{+}24b] Marek Šmejkal, Ondřej Dočkal, Kiran Thomas, Chandani R. Verma, Pradeep Kumkar, and Lukáš Kalous. First record of highly invasive Chinese sleeper Perccottus glenii Dybowski, 1877 (Perciformes: Odontobutidae) in the Elbe River Basin, Czechia. Aquatic Ecology, 58(1):125–130, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s10452-023-10036-8. See correction [ŠDT^{+}24a].

Sunardi:2024:PAR

[SDW^{+}24] Sunardi Sunardi, Moh. Dede, Sandy Budi Wibowo, Yudo Prasetyo, Annisa Joviani Astari, Lukman Lukman, Franck Lavigne, Christopher Gomez, Idea Wening Nurani, Yuichi Sakai, and Mohd Khairul Amri Kamarudin. Preliminary assessment of river ecosystem services in the volcanic area of Mount Merapi, Indonesia. Aquatic Ecology, 58(3):819–832, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10107-4.

Schneider:2011:RBV

[SEB11] Christine Schneider, Klemens Ekschmitt, and Klaus Birkhofer. Ring-based versus disc-based separation of spatial scales: a case study on the impact of arable land proportions on invertebrates in freshwater streams. Aquatic Ecology, 45(3): 351–356, September 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9358-8.

Sevindik:2022:EIA

[SEE22] Tuğba Ongun Sevindik, Mehmet Erdoğan, and Mehmet Korhan Erturaç. The effects of inter-annual fluctuations in precipitation, lake surface area, and wind speed on phytoplankton structure in three shallow Mediterranean lakes (Sakarya, Turkey). Aquatic Ecology, 56(3):697–718, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09929-3.

Sepers:1979:AMO

[Sep79] A. B. J. Sepers. The aerobic mineralization of organic compounds in the saline Lake Grevelingen (The Netherlands) (VI). Aquatic Ecology, 13(2-3):81–82, December 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284733.

Schmera:2009:MAF

[SEP09] Dénes Schmera, Tibor Erős, and János Podani. A measure for assessing functional diversity in ecological communities. Aquatic Ecology, 43(1):157–167, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9152-9.

Scherwass:2005:DPF

[SFA05] Anja Scherwass, Yvonne Fischer, and Hartmut Arndt. Detritus as a potential food source for protozoans: utilization of fine particulate plant detritus by a heterotrophic flagellate, Chilomonas paramecium, and a ciliate, Tetrahymena pyriformis. Aquatic Ecology, 39(4):439–445, December 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9012-4.

Schlechtriem:2004:SIT

[SFB04] Christian Schlechtriem, Ulfert Focken, and Klaus Becker. Stable isotopes as a tool for nutrient assimilation studies in larval fish feeding on live food. Aquatic Ecology, 38(1):93–100, March 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000020951.76155.3e.

Somogyi:2009:WBP

[SFV09] Boglárka Somogyi, Tamás Felföldi, and Lajos Vörös. Winter bloom of picoeukaryotes in Hungarian shallow turbid soda pans and the role of light and temperature. Aquatic Ecology, 43(3):735–744, September 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9269-0.

Solheim:2008:PQE

[SG08] Anne Lyche Solheim and Ramesh D. Gulati. Preface: ‘quantitative ecological responses for the water framework directive related to eutrophication and acidification of European lakes’. Aquatic Ecology, 42(2):179–181, June 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9187-6.

Supekar:2020:DAR

[SG20] Swapnil C. Supekar and Narahari P. Gramapurohit. Do antipredator responses of Euphlyctis cyanophlyctis tadpoles depend on the intensity of predation risk? Aquatic Ecology, 54(3):823–837, September 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
020-09780-y.

Szarek-Gwiazda:2006:BTE

[SGA06] Ewa Szarek-Gwiazda and Antoni Amirowicz. Bioaccumulation of trace elements in roach, silver bream, rudd, and perch living in an inundated opencast sulphur mine. Aquatic Ecology, 40(2):221–236, June 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
004-7341-3.

Siegfried:2022:NNN

[SGD22] Tabitha R. Siegfried, Silvia M. M. Gutierre, and Philip C. Darby. Native and non-native fish predators differ in their consumptive and non-consumptive impacts on a native freshwater snail. Aquatic Ecology, 56(3):865–876, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09944-y.

Szarek-Gwiazda:2006:SSC

[SGGO06] Ewa Szarek-Gwiazda, Joanna Galas, and Marcin Ollik. Surface sediment composition in an inundated opencast sulphur mine (Piaseczno reservoir, Southern Poland). Aquatic Ecology, 40(2):155–164, June 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
004-1914-z.

Simoes:2024:RSR

[SGJ^{+}24] S. Simões, A. L. Gonçalves, T. Hefin Jones, J. P. Sousa, and C. Canhoto. Reciprocal stream–riparian fluxes: effects of distinct exposure patterns on litter decomposition. Aquatic Ecology, 58(2):313–322, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10067-1.

Sushchik:2007:SDL

[SGK07] Nadezhda N. Sushchik, Michail I. Gladyshev, and Galina S. Kalachova. Seasonal dynamics of long-chain polyunsaturated fatty acids in littoral benthos in the upper Yenisei river. Aquatic Ecology, 41(2):349–365, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9065-z.

Saravia:1999:PME

[SGM99] Leonardo A. Saravia, Adonis Giorgi, and Fernando R. Momo. A photographic method for estimating chlorophyll in periphyton on artificial substrata. Aquatic Ecology, 33(4): 325–330, December 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009934626188.

Soudani:2022:HMP

[SGM22] Alireza Soudani, Ali Gholami, and Amin Mojiri. Heavy metal phytoremediation of aqueous solution by Typha domingensis. Aquatic Ecology, 56(2):513–523, June 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09945-x.

Sweerts:1986:PUC

[SGR86] Jean-Pierre R. A. Sweerts, Herman J. Gons, and Machteld Rijkeboer. Phosphate uptake capacity of summer phytoplankton of the Loosdrecht Lakes in relation to phosphorus loading and irradiance. Aquatic Ecology, 20(1-2):101–107, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291155.

Saez-Gomez:2020:MEF

[SGRMP20] P. Sáez-Gómez, A. Ramos-Merchante, and J. Prenda. Multiscale effects on freshwater fish distribution in a highly disturbed Mediterranean-type basin: community-level and species-level responses. Aquatic Ecology, 54(3):869–887, September 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09783-9.

Stamenkovic:2011:GPC

[SH11] Marija Stamenković and Dieter Hanelt. Growth and photosynthetic characteristics of several Cosmarium strains (Zygnematophyceae, Streptophyta) isolated from various geographic regions under a constant light-temperature regime. Aquatic Ecology, 45(4):455–472, November 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9367-7.

Sanchez-Hernandez:2016:DAR

[SH16] Javier Sánchez-Hernández. Do age-related changes in feeding habits of brown trout alter structural properties of food webs? Aquatic Ecology, 50(4):685–695, December 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9586-z.

Shaw:1978:EAT

[Sha78] T. L. Shaw. Engineering aspects of tidal power barrages: Their influence on the environment. Aquatic Ecology, 12(3-4):299–306, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259191.

Shiel:1998:BRL

[Shi98] R. J. Shiel. Book review: Live Food in Aquaculture. Edited by A. Hagiwara, T. W. Snell. E. Lubzens and C. S. Tamaru. Aquatic Ecology, 32(4):368–369, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009911519991.

Smit:1992:LAC

[SHK92] Henk Smit, Floor Heinis, and Frans Kerkum. Lipiniella arenicola (Chironomidae) compared with Chironomus muratensis and Ch. nudiventris: Distribution patterns related to depth and sediment characteristics, diet, and behavioural response to reduced oxygen concentrations. Aquatic Ecology, 26(2-4):431–440, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255272.

Spiers:2021:RAC

[SHP21] Lindsay J. Spiers, Sarah J. Harrison, and Valerie J. Paul. The role of algal chemical defenses in the feeding preferences of the long-spined sea urchin Diadema antillarum. Aquatic Ecology, 55(3):941–953, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09873-2.

Schuchardt:1993:TFR

[SHS93] B. Schuchardt, U. Haesloop, and M. Schirmer. The tidal freshwater reach of the Weser estuary: Riverine or estuarine? Aquatic Ecology, 27(2-4):215–226, June 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334785.

Slim:1996:TEM

[SHV96] F. J. Slim, M. A. Hemminga, and G. Van Der Velde. Tidal exchange of macrolitter between a mangrove forest and adjacent seagrass beds (Gazi Bay, Kenya). Aquatic Ecology, 30(2-3):119–128, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02272233.

Simons:1994:FEF

[Sim94] J. Simons. Field ecology of freshwater macroalgae in pools and ditches, with special attention to eutrophication. Aquatic Ecology, 28(1):25–33, March 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334242.

Sirianni:2017:DWD

[Sir17] Katherine M. Sirianni. Differential wind dispersal of cladoceran ephippia in a rock pool metacommunity. Aquatic Ecology, 51(2):203–218, June 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
016-9611-2.

Schwartz:2000:TAH

[SJ00a] Steven S. Schwartz and David G. Jenkins. Temporary aquatic habitats: constraints and opportunities. Aquatic Ecology, 34(1):3–8, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009944918152.

Stevens:2000:ASD

[SJ00b] Patrick H. Stevens and David G. Jenkins. Analyzing species distributions among temporary ponds with a permutation test approach to the join-count statistic. Aquatic Ecology, 34(1):91–99, March 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009951815795.

Sun:2022:IAP

[SJD22] Jianfan Sun, Qaiser Javed, and Daolin Du. Invasive Alternanthera philoxeroides has performance advantages over natives under flooding with high amount of nitrogen. Aquatic Ecology, 56(3):891–903, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09951-z.

Saleh:2023:MNL

[SJJA23] Alzayat Saleh, David Jones, Dean Jerry, and Mostafa Rahimi Azghadi. MFLD-net: a lightweight deep learning network for fish morphometry using landmark detection. Aquatic Ecology, 57(4):913–931, December 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
023-10044-8.

Sinhababu:1984:ELL

[SK84] Abhijit Sinhababu and Debiprasad Kushari. Effect of leaf leachates of Polyalthia longifolia on the growth and nitrogen fixation of Azolla pinnata. Aquatic Ecology, 18(2):103–108, December 1984. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257049.

Soininen:2004:CSM

[SK04] Janne Soininen and Katriina Könönen. Comparative study of monitoring South-Finnish rivers and streams using macroinvertebrate and benthic diatom community structure. Aquatic Ecology, 38(1):63–75, March 2004. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000021004.06965.bd.

Schagerl:2009:AMS

[SK09] Michael Schagerl and Maria Kerschbaumer. Autecology and morphology of selected Vaucheria species (Xanthophyceae). Aquatic Ecology, 43(2):295–303, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9163-6.

Stroom:2016:HCC

[SK16] Jasper M. Stroom and W. Edwin A. Kardinaal. How to combat cyanobacterial blooms: strategy toward preventive lake restoration and reactive control measures. Aquatic Ecology, 50(3):541–576, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9593-0.

Shiffman:2019:IDR

[SKH19] David Samuel Shiffman, Les Kaufman, and Neil Hammerschlag. Intraspecific differences in relative isotopic niche area and overlap of co-occurring sharks. Aquatic Ecology, 53(2):233–250, June 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09685-5.

Shilova:1992:LPS

[SKK92] A. I. Shilova, I. E. Kerkis, and I. I. Kiknadze. Lipiniella prima sp.nov. (Diptera, Chironomidae), larva and karyotype. Aquatic Ecology, 26(2-4):197–201, June 1992. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255242.

Spanhoff:2005:LICa

[SKM05a] Bernd Spänhoff, Norbert Kaschek, and Elisabeth Irmgard Meyer. Laboratory investigation on community composition, emergence patterns and biomass of wood-inhabiting Chironomidae (Diptera) from a sandy lowland stream in Central Europe (Germany). Aquatic Ecology, 38(4):547–560, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-4160-0.

Spanhoff:2005:LICb

[SKM05b] Bernd Spänhoff, Norbert Kaschek, and Elisabeth Irmgard Meyer. Laboratory investigation on community composition, emergence patterns and biomass of wood-inhabiting Chironomidae (Diptera) from a sandy lowland stream in Central Europe (Germany). Aquatic Ecology, 38(4):547–560, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-004-4160-5.

Schmidt:2009:FAR

[SKRB09] Susanne I. Schmidt, Marie König-Rinke, and Jürgen Benndorf. Finding appropriate reference sites in large-scale aquatic field experiments. Aquatic Ecology, 43(1):169–179, March 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9155-6.

Shevyrnogov:2002:ASD

[SKV02] Anatoly P. Shevyrnogov, Alexei V. Kartushinsky, and Galina S. Vysotskaya. Application of satellite data for investigation of dynamic processes in inland water bodies: Lake Shira (Khakasia, Siberia), a case study. Aquatic Ecology, 36(2):153–164, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1015658927683.

Saintilan:2021:ILD

[SKW21] Neil Saintilan, Jeffrey J. Kelleway, and Li Wen. Incorporation of local dissolved organic carbon into floodplain aquatic ecosystems. Aquatic Ecology, 55(3):779–790, September 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09860-7.

Santmire:2007:ISS

[SL07] J. A. Santmire and L. G. Leff. The influence of stream sediment particle size on bacterial abundance and community composition. Aquatic Ecology, 41(2):153–160, June 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9060-4.

Shi:2024:IOA

[SL24] Yuntian Shi and Yaowu Li. Impacts of ocean acidification on physiology and ecology of marine invertebrates: a comprehensive review. Aquatic Ecology, 58(2):207–226, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10058-2.

Song:2010:ERE

[SLC10] Xiaolan Song, Zhengwen Liu, and Yuwei Chen. Effects of resuspension and eutrophication level on summer phytoplankton dynamics in two hypertrophic areas of Lake Taihu, China. Aquatic Ecology, 44(1):41–54, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9258-3.

Salonen:2009:PWL

[SLG09a] K. Salonen, M. Leppäranta, and R. D. Gulati. Perspectives in winter limnology: closing the annual cycle of freezing lakes. Aquatic Ecology, 43(3):609–616, September 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9278-z.

Salonen:2009:P

[SLG09b] Kalevi Salonen, Matti Leppäranta, and Ramesh D. Gulati. Preface. Aquatic Ecology, 43(3):607–608, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9287-y.

Swain:1987:CDM

[SLH87] W. R. Swain, R. Lingeman, and F. Heinis. A characterization and description of the Maarsseveen Lake system. Aquatic Ecology, 21(1):5–16, July 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255450.

Song:2009:IAL

[SLL09] Mi-Young Song, Fabien Leprieur, and Sovan Lek. Impact of agricultural land use on aquatic insect assemblages in the Garonne river catchment (SW France). Aquatic Ecology, 43(4):999–1009, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
008-9218-3.

Schneck:2017:ENF

[SLM17] Fabiana Schneck, Katharina Lange, and Christoph D. Matthaei. Effects of a natural flood disturbance on species richness and beta diversity of stream benthic diatom communities. Aquatic Ecology, 51(4):557–569, December 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9636-1.

Savvides:2015:CSO

[SLS15] Pantelis Savvides, Vasilis Louca, and Spyros Sfenthourakis. Competition for shelter occupancy between a native freshwater crab and an invasive crayfish. Aquatic Ecology, 49(3): 273–278, September 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9522-7.

Sluys:1981:OOC

[Slu81] R. Sluys. Ostracods (Ostracoda, Crustacea) from two freshwater lakes near Maarssen, The Netherlands. Aquatic Ecology, 15(1-2):92–93, October 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260263.

Seuffert:2012:LDL

[SM12] María E. Seuffert and Pablo R. Martín. A lentic dweller in lotic habitats: the behavior of the invasive South American apple snail Pomacea canaliculata in flowing water. Aquatic Ecology, 46(1):129–142, March 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9386-4.

Smaal:1997:P

[Sma97] Aad C. Smaal. Preface. Aquatic Ecology, 30(4):I, May 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02085868.

Sharma:2008:DSM

[SMC08] Jai Gopal Sharma, Prabhat Mittal, and Rina Chakrabarti. Development of survivorship model for UV-b irradiated Catla catla larvae. Aquatic Ecology, 42(1):17–23, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9069-8.

Simoes:2021:APL

[SMC21] Sara Simões, Aingeru Martínez, and Cristina Canhoto. Annual patterns of litter decomposition in the channel and riparian areas of an intermittent stream. Aquatic Ecology, 55 (2):519–526, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09841-w.

Sanchez-Moyano:2000:EEF

[SMGAGG00] J. E. Sánchez-Moyano, E. M. García-Adiego, and J. C. García-Gómez. Effect of environmental factors on the spatial distribution of the epifauna of the alga Halopteris scoparia in Algeciras Bay, Southern Spain. Aquatic Ecology, 34(4): 355–367, December 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1011411414342.

Smith:1995:CBC

[SMJ95] J. A. Smith, G. E. Millward, and M. B. Jones. Changes in benthic community structure following construction of a harbour impoundment scheme. Aquatic Ecology, 29(3-4): 449–457, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02084243.

Sierszen:2003:BBA

[SMJ03] Michael E. Sierszen, Michael E. McDonald, and Douglas A. Jensen. Benthos as the basis for Arctic lake food webs. Aquatic Ecology, 37(4):437–445, October 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000007042.09767.dd.

Saveanu:2023:SII

[SMM23] Lucía Saveanu, Enzo Manara, and Pablo Rafael Martín. Sediment ingestion in the invasive apple snail Pomacea canaliculata. Aquatic Ecology, 57(2):433–442, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10020-2.

Sagova-Mareckova:2009:BPN

[SMPK09] Markéta Ságová-Marečková, Adam Petrusek, and Jan Květ. Biomass production and nutrient accumulation in Sparganium emersum Rehm. after sediment treatment with mineral and organic fertilisers in three mesocosm experiments. Aquatic Ecology, 43(4):??, December 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9214-7.

Schartau:2008:MIL

[SMR08] Ann Kristin Schartau, S. Jannicke Moe, and Gunnar G. Raddum. Macroinvertebrate indicators of lake acidification: analysis of monitoring data from UK, Norway and Sweden. Aquatic Ecology, 42(2):293–305, June 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9186-7.

SilvadeMelo:2024:ECC

[SMS^{+}24] Henrique Paulo Silva de Melo, Ricardo Massato Takemoto, Gabriela Silva Ribeiro Gonçalves, Renata Guimaraes Frederico, and Lucena Rocha Virgilio. Effects of climate change on the distribution of Hoplias malabaricus and its ecto- and endoparasite species in South America. Aquatic Ecology, 58(3):999–1011, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10120-7.

Sandstrom:2003:LTD

[SN03] Olof Sandström and Erik Neuman. Long-term development in a Baltic fish community exposed to bleached pulp mill effluent. Aquatic Ecology, 37(3):267–276, July 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1025855731450.

Sarma:2021:PCR

[SNG21] S. S. S. Sarma, S. Nandini, and Brij Gopal. Preface: the central role of zooplankton in freshwaters, a special issue in honour of late Ramesh D. Gulati. Aquatic Ecology, 55(4): 1121–1125, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09906-w.

Sarma:2006:ENS

[SNGV06] S. S. S. Sarma, S. Nandini, and Leticia González-Valverde. Effects of NaCl salinity on the population dynamics of freshwater zooplankton (rotifers and cladocerans). Aquatic Ecology, 40(3):349–360, September 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
006-9039-1.

Shafaghat:2024:GDL

[SNKN^{+}24] Zahra Shafaghat, Farzaneh Najafi, Ramazan-Ali Khavari-Nejad, Mehdi Mohammadi, and Sattar Tahmasebi Enferadi. Growth dynamics and lipid metabolism of Aurantiochytrium sp.: insights into its potential applications. Aquatic Ecology, 58(3):789–799, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10105-6.

Sibbing:1998:SEL

[SNO98] Ferdinand A. Sibbing, Leo A. J. Nagelkerke, and Jan W. M. Osse. Speciation of endemic Lake Tana barbs (Cyprinidae, Ethiopia) driven by trophic resource partitioning; a molecular and ecomorphological approach. Aquatic Ecology, 32 (3):217–227, October 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009920522235.

Spoljaric:2021:IEC

[ŠNZ21] Ivna Vrana Špoljarić, Tihana Novak, and Zhuoyi Zhu. Impact of environmental conditions on phospholipid fatty acid composition: implications from two contrasting estuaries. Aquatic Ecology, 55(1):1–20, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09805-6.

Slusarczyk:2012:PDR

[SOC12] M. Slusarczyk, A. Ochocka, and D. Cichocka. The prevalence of diapause response to risk of size-selective predation in small- and large-bodied prey species. Aquatic Ecology, 46(1):1–8, March 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9376-6.

Simons:1992:RBN

[SOD92] Jan Simons, Marieke Ohm, and Remco Daalder. Restoration of Botshol (The Netherlands) by reduction of external nutrient load: Recovery of the Characean community. Aquatic Ecology, 25(3):287–294, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02270814.

Soininen:2005:ACRa

[Soi05a] Janne Soininen. Assessing the current related heterogeneity and diversity patterns of benthic diatom communities in a turbid and a clear water river. Aquatic Ecology, 38(4): 495–501, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-4089-3.

Soininen:2005:ACRb

[Soi05b] Janne Soininen. Assessing the current related heterogeneity and diversity patterns of benthic diatom communities in a turbid and a clear water river. Aquatic Ecology, 38(4): 495–501, January 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-004-4089-8.

Sola:1997:RPD

[Sol97] J. Carlos Sola. Reproduction, population dynamics, growth and production of Scrobicularia plana da costa (Pelecypoda) in the Bidasoa estuary, Spain. Aquatic Ecology, 30(4):283– 296, May 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02085872.

Sommer:1983:ANC

[Som83] Ulrich Sommer. Algal nutrient competition in continuous culture. Aquatic Ecology, 17(1):21–27, June 1983. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255189.

Steinman:2011:PRS

[SON11] Alan D. Steinman, Mary E. Ogdahl, and Stephen Nold. Periphyton response to simulated nonpoint source pollution: local over regional control. Aquatic Ecology, 45(4):439–454, November 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-011-9366-8.

Sorlin:1982:GBN

[Sör82] T. Sörlin. The Gulf of Bothnia: The northernmost part of the Baltic Sea. Aquatic Ecology, 16(2-3):287–288, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255384.

Steinman:2014:IWL

[SOU14] Alan D. Steinman, Mary E. Ogdahl, and Donald G. Uzarski. Influence of water-level fluctuation duration and magnitude on sediment–water nutrient exchange in coastal wetlands. Aquatic Ecology, 48(2):143–159, June 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9472-5.

Schleiter:2001:BCH

[SOW01] Ingrid M. Schleiter, M. Obach, and H. Werner. Bioindication of chemical and hydromorphological habitat characteristics with benthic macro-invertebrates based on Artificial Neural Networks. Aquatic Ecology, 35(2):147–158, June 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011433529239.

Spaargaren:1971:MOB

[Spa71] D. H. Spaargaren. Over mechanismen en oecologische betekenis van de osmoregulatie van de garnalen Crangon crangon (L.) en Crangon allmanni kin.. (Dutch) [On mechanisms and ecological significance of the osmoregulation of the shrimp Crangon crangon (L.) and Crangon allmanni kin.]. Aquatic Ecology, 5(2):108–109, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185955.

Stasko:2012:PCD

[SPA12] Ashley D. Stasko, Theresa Patenaude, and Shelley E. Arnott. Portage connectivity does not predict establishment success of canoe-mediated dispersal for crustacean zooplankton. Aquatic Ecology, 46(1):9–24, March 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9378-4.

Smaal:1997:MRM

[SPB97] A. C. Smaal, T. C. Prins, and B. Ball. Minimum requirements for modelling bivalve carrying capacity. Aquatic Ecology, 31(4):423–428, December 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/A:
1009947627828.

Santos:2024:UJR

[SPB^{+}24] Thaís Barbosa Santos, Angelo Poliseno, Amanda G. Bendia, Vivian H. Pellizari, James D. Reimer, and Sérgio N. Stampar. Unlocking the jar: revealing gastric content in Ceriantharia (Cnidaria, Anthozoa) through whole-genome shotgun sequencing. Aquatic Ecology, 58(2):375–385, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10076-0.

Srifa:2016:PDS

[SPH16] Akeapot Srifa, Edward J. Phlips, and John C. Hendrickson. Phytoplankton dynamics in a subtropical lake dominated by cyanobacteria: cyanobacteria ‘like it hot’ and sometimes dry. Aquatic Ecology, 50(2):163–174, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9565-4.

Suhonen:2022:MSO

[SPS22] Jukka Suhonen, Lauri Paasivirta, and Erna Suutari. Macroinvertebrate species occupancy frequency distribution patterns in eutrophic lakes. Aquatic Ecology, 56(1):201–212, March 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09909-7.

Swain:1985:FZS

[SR85] W. R. Swain and K. M. Roijackers. Freshwater zooplankton sampling reconsidered: Preliminary results of a highspeed sampling device for small lakes. Aquatic Ecology, 19 (2):139–152, December 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02270760.

Suykerbuyk:1988:PBT

[SR88] R. E. M. Suykerbuyk and R. M. M. Roijackers. A phytoplankton-based typology of brackish and freshwaters in The Netherlands. Aquatic Ecology, 22(1):95–98, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02256791.

Sherk:1992:ECF

[SR92] Truman Sherk and Greg Rau. Emergence of Chironomidae from Findley Lake and two ponds in the Cascade Mountains, U.S.A. Aquatic Ecology, 26(2-4):321–330, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255258.

Santos:2021:RPA

[SRF21] Mateus Pereira Santos, Gabriel Fellipe Barros Rodrigues, and Vivian Fransozo. Resource partitioning and adequacy among ontogenetic groups in a hermit crab and gastropod shell network. Aquatic Ecology, 55(1):253–264, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09827-0.

Steele:2018:HPC

[SRG18] LaTina Steele, Courtney Ray, and Michele Guidone. High phenolic content fails to deter mesograzer consumption of Myriophyllum spicatum (Eurasian watermilfoil) in New England. Aquatic Ecology, 52(4):255–267, December 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9661-8.

Solheim:2008:ETR

[SRH08] Anne Lyche Solheim, Seppo Rekolainen, and Trygve Hesthagen. Ecological threshold responses in European lakes and their applicability for the Water Framework Directive (WFD) implementation: synthesis of lakes results from the REBECCA project. Aquatic Ecology, 42(2):317–334, June 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9188-5.

Storero:2020:PIS

[SRI20] Lorena P. Storero, Matías Ocampo Reinaldo, and Oscar O. Iribarne. Predation and interaction strength of octopuses and sea stars on different functional groups of the rocky intertidal shores of the Patagonian coast. Aquatic Ecology, 54 (1):193–203, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09736-x.

Soponis:1992:PDT

[SS92] Annelle R. Soponis and Karl W. Simpson. Polypedilum digitifer Townes and Polypedilum griseopunctatum (Malloch) (Diptera, Chironomidae): Redescription of adult males with a description and separation of the immature stages. Aquatic Ecology, 26(2-4):203–213, June 1992. CODEN AQECF9.

ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255243.

Sterner:1998:ZNR

[SS98] Robert W. Sterner and Kimberly L. Schulz. Zooplankton nutrition: recent progress and a reality check. Aquatic Ecology, 32(4):261–279, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009949400573.

Scholten:1999:ERM

[SS99] Huub Scholten and Aad C. Smaal. The ecophysiological response of mussels (Mytilus edulis) in mesocosms to a range of inorganic nutrient loads: simulations with the model EMMY. Aquatic Ecology, 33(1):83–100, March 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009995823741.

Sadegh:2024:FOE

[SS24] Naghme Sadegh and Mohammad Reza Shokri. First observation of Acropora downingi spawning in Kish Island, northern Persian Gulf. Aquatic Ecology, 58(4):1193–1199, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10134-1.

Stamenkovic:2021:DWC

[SSB21] Olivera Stamenković, Vladica Simić, and Oksana Y. Buzhdygan. Direct, water-chemistry mediated, and cascading effects of human-impact intensification on multitrophic biodiversity in ponds. Aquatic Ecology, 55(1):187–214, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09822-5.

Sanchez:2021:ELU

[SSI21] María Laura Sánchez, María Romina Schiaffino, and Irina Izaguirre. Effect of land use on the phytoplankton community of Pampean shallow lakes of the Salado River basin (Buenos Aires Province, Argentina). Aquatic Ecology, 55 (2):417–435, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09835-8.

Siikavuopio:2014:TPJ

[SSK14] Sten Ivar Siikavuopio, Bjørn-Steinar Sæther, and Rune Knudsen. Temperature preference of juvenile Arctic charr originating from different thermal environments. Aquatic Ecology, 48(3):313–320, September 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
014-9485-0.

SilvaNascimento:2024:SMS

[SSN24] Lorena Silva Nascimento, Carmem Satie Hara, and Miodeli Nogueira Júnior. Social media in service of aquatic ecology. Aquatic Ecology, 58(1):1–2, March 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10088-4.

Schwenk:2013:LCB

[SSR13] Dagmar Schwenk, Jukka Seppälä, and Heiko Rischer. Lipid content in 19 brackish and marine microalgae: influence of growth phase, salinity and temperature. Aquatic Ecology, 47 (4):415–424, December 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-013-9454-z.

Schwenk:1998:GMG

[SSS98] Klaus Schwenk, Anna Sand, and Piet Spaak. Genetic markers, genealogies and biogeographic patterns in the Cladocera. Aquatic Ecology, 32(1):37–51, March 1998. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009939901198.

Santori:2021:HSN

[SSV21] Claudia Santori, Ricky-John Spencer, and James U. Van Dyke. Hatchling short-necked turtles (Emydura macquarii) select aquatic vegetation habitats, but not after one month in captivity. Aquatic Ecology, 55(1):85–96, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09813-6.

Sakdullah:2009:OPO

[ST09] Abdulkadir Sakdullah and Makoto Tsuchiya. The origin of particulate organic matter and the diet of tilapia from an estuarine ecosystem subjected to domestic wastewater discharge: fatty acid analysis approach. Aquatic Ecology, 43 (2):577–589, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9195-6.

Silveira:2020:EDN

[ST20] Márcio José Silveira and Gabrielle Thiébaut. Effect of density and neighbours on interactions between invasive plants of similar growth form. Aquatic Ecology, 54(2):463–474, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09753-1.

Stamhuis:2006:BPP

[Sta06] Eize J. Stamhuis. Basics and principles of particle image velocimetry (PIV) for mapping biogenic and biologically relevant flows. Aquatic Ecology, 40(4):463–479, December 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-6567-z.

Steenbergen:1976:PSQ

[Ste76] C. L. M. Steenbergen. Pleomorphism in Scenedesmus quadricauda (Turp) Bréb. (Chlorophyceae). Aquatic Ecology, 10(1): 14, April 1976. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02308492.

Steinmetz:1983:DFM

[Ste83] B. Steinmetz. Developments in fishery management in The Netherlands. Aquatic Ecology, 17(1):67–69, June 1983. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255193.

Sterrenburg:1994:TME

[Ste94] F. A. S. Sterrenburg. Terpsinoe musica Ehrenberg (Bacillariophyceae, centrales), with emphasis on protoplast and cell division. Aquatic Ecology, 28(1):63–69, March 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334246.

Subina:2018:NIS

[STG18] N. S. Subina, B. R. Thorat, and Maria-Judith Gonsalves. Nitrification in intertidal sponge Cinachyrella cavernosa. Aquatic Ecology, 52(2-3):155–164, September 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9651-x.

Sun:2021:CVP

[STH21] Bin Sun, Chunyu Tang, and Peimin He. Composition and variation of phytoplankton communities during Microcystis bloom in an artificial lagoon of Hangzhou Bay, China. Aquatic Ecology, 55(2):467–481, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09838-5.

Sarma:2003:PGH

[STHN03] S. S. S. Sarma, Hugo Enrique Trujillo-Hernández, and S. Nandini. Population growth of herbivorous rotifers and their predator (Asplanchna) on urban wastewaters. Aquatic Ecology, 37(3):243–250, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1025896703470.

Stoppelaar:1976:NID

[Sto76] C. F. Den Stoppelaar. De Nederlandse invertebraten data bank. (Dutch) [The Dutch Invertebrates Database]. Aquatic Ecology, 10(1):80–81, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02308507.

Stronkhorst:1988:ISC

[Str88] J. Stronkhorst. The influence of the spectral composition of irradiance on primary production in the Eastern Scheldt (The Netherlands). Aquatic Ecology, 22(2):127–134, December 1988. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02256816.

Stronkhorst:1993:ERP

[Str93] J. Stronkhorst. The environmental risks of pollution in the Scheldt estuary. Aquatic Ecology, 27(2-4):383–393, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334800.

Sudaryanti:2001:ABH

[STT01] Sri Sudaryanti, Yulinah Trihadiningrum, and Lisa Thurtell. Assessment of the biological health of the Brantas River, East Java, Indonesia using the Australian River Assessment System (AUSRIVAS) methodology. Aquatic Ecology, 35(2):135– 146, June 2001. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1011458520966.

Saulino:2018:CHF

[STTS18a] Hugo Henrique L. Saulino, Ross M. Thompson, and Susana Trivinho-Strixino. Correction to: Herbivore functional traits and macroinvertebrate food webs have different responses to leaf chemical compounds of two macrophyte species in a tropical lake’s littoral zone. Aquatic Ecology, 52(2-3):177–178, September 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-018-9655-6.

Saulino:2018:HFT

[STTS18b] Hugo Henrique L. Saulino, Ross M. Thompson, and Susana Trivinho-Strxino. Herbivore functional traits and macroinvertebrate food webs have different responses to leaf chemical compounds of two macrophyte species in a tropical lake’s littoral zone. Aquatic Ecology, 52(2-3):165–176, September 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-018-9652-9.

Smith:2006:ENI

[STV06] Geoffrey R. Smith, Kathleen G. Temple, and David A. Vaala. Effects of nitrate on the interactions of the tadpoles of two ranids (Rana clamitans and R. catesbeiana). Aquatic Ecology, 40(1):125–130, March 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9015-1.

Sures:2001:UFP

[Sur01] Bernd Sures. The use of fish parasites as bioindicators of heavy metals in aquatic ecosystems: a review. Aquatic Ecology, 35(2):245–255, June 2001. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1011422310314.

Schot:1992:RPN

[SV92] Joke A. Schot and Piet F. M. Verdonschot. A representative pond net sample: Fiction or reality? Aquatic Ecology, 26 (2-4):457–460, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255275.

Strojsova:2009:STV

[ŠV09] Alena Štrojsová and Jaroslav Vrba. Short-term variation in extracellular phosphatase activity: possible limitations for diagnosis of nutrient status in particular algal populations. Aquatic Ecology, 43(1):19–25, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9154-7.

Schiedek:1997:MCWb

[SVB97] D. Schiedek, C. Vogan, and M. Bentley. Marenzelleria cf. wireni (Polychaeta: Spionidae) from the Tay estuary. Metabolic response to severe hypoxia and hydrogen sulphide. Aquatic Ecology, 31(2):211–222, June 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009911723907.

Snel:1998:IPI

[SVB98] Jan F. H. Snel, José H. Vos, and Theo C. M. Brock. Inhibition of photosystem II (PSII) electron transport as a convenient endpoint to assess stress of the herbicide linuron on freshwater plants. Aquatic Ecology, 32(2):113–123, September 1998. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009971930626.

Sokolova:1992:M

[SvdG92] N. Yu. Sokolova and Cees van de Guchte. In memoriam. Aquatic Ecology, 26(2-4):99, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255230.

Schmidlin:2015:TEF

[SvFN15] Lara Schmidlin, Stefanie von Fumetti, and Peter Nagel. Temperature effects on the feeding and electron transport system (ETS) activity of Gammarus fossarum. Aquatic Ecology, 49(1):71–80, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9505-8.

Smit:1994:MAL

[SVK94] Henk Smit, Joan A. Van Der Velden, and Alexander Klink. Macrozoobenthic assemblages in littoral sediments in the enclosed Rhine–Meuse Delta. Aquatic Ecology, 28(2):199–212, June 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02333991.

Seminara:2008:DRZ

[SVM08] Marco Seminara, Daria Vagaggini, and Fiorenza G. Margaritora. Differential responses of zooplankton assemblages to environmental variation in temporary and permanent ponds. Aquatic Ecology, 42(1):129–140, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9088-0.

Smit:1995:SMA

[SVW95] H. Smit, H. C. Reinhold-Dudok Van Heel, and S. M. Wiersma. Sublittoral macrozoobenthic assemblages in the enclosed sediment-polluted Rhine–Meuse Delta; their relationship to environmental conditions. Aquatic Ecology, 29 (1):31–47, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02061787.

Schaffer:2013:RDA

[SWB13] Michael Schäffer, Carola Winkelmann, and Jürgen Benndorf. Reduced drift activity of two benthic invertebrate species is mediated by infochemicals of benthic fish. Aquatic Ecology, 47(1):99–107, March 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-013-9428-1.

Swennen:1971:PBD

[Swe71] C. Swennen. Een parasitaire besmetting als doodsoorzaak van zilvermeeuwen. (Dutch) [A parasitic infestation as the cause of death of herring gulls]. Aquatic Ecology, 5(2):103– 104, June 1971. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185951.

Santana:2017:SSF

[SWF17] Lucineide Maria Santana, Guntram Weithoff, and Carla Ferragut. Seasonal and spatial functional shifts in phytoplankton communities of five tropical reservoirs. Aquatic Ecology, 51(4):531–543, December 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
017-9634-3.

Salomon:2002:MTE

[SWW02] Anne K. Salomon, Nigel P. Waller, and Carl Walters. Modeling the trophic effects of marine protected area zoning policies: a case study. Aquatic Ecology, 36(1):85–95, January 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1013346622536.

Sun:2025:DTR

[SXC^{+}25] Yan Sun, Dayong Xu, Xi Chen, Junming Zhou, Changwei Jiang, Zhiqiu Huang, and Dongming Qi. Deciphering the transcriptional regulatory networks of FOX genes in nitriteinduced spleen injury in largemouth bass. Aquatic Ecology, 59(1):1–20, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10132-3.

Si:2020:NEG

[SXY20] Chao Si, Wei Xue, and Fei-Hai Yu. No evidence of greater biomass allocation to stolons at moderate resource levels in a floating plant. Aquatic Ecology, 54(1):421–429, March 2020.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09751-3.

Skriptsova:2002:IVI

[SY02] A. V. Skriptsova and I. M. Yakovleva. The influence of variations in irradiance upon morphology in an unattached form of Gracilaria gracilis (Stackhouse) Steentoft during field cultivation, South Primorye, Russia. Aquatic Ecology, 36(4): 511–518, December 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1021189822333.

Soejima:2009:GVS

[SYW09] Akiko Soejima, Natsuko Yamazaki, and Isamu Wakana. Genetic variation and structure of the endangered freshwater benthic alga Marimo, Aegagropila linnaei (Ulvophyceae) in Japanese lakes. Aquatic Ecology, 43(2):359–370, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9204-9.

Syrovatka:2020:NGF

[SZH20] Vít Syrovátka, Marie Zhai, and Michal Horsák. Native Gammarus fossarum affects species composition of macroinvertebrate communities: evidence from laboratory, field enclosures, and natural habitat. Aquatic Ecology, 54(2):505–518, June 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09756-y.

Sha:2021:DVM

[SZH21] Yongcui Sha, Huan Zhang, and Lars-Anders Hansson. Diel vertical migration of copepods and its environmental drivers in subtropical Bahamian blue holes. Aquatic Ecology, 55(4): 1157–1169, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09807-4.

Si:2019:EPS

[SZY19] Chao Si, Li-Min Zhang, and Fei-Hai Yu. Effects of physical space and nutrients on the growth and intraspecific competition of a floating fern. Aquatic Ecology, 53(2):295–302, June 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09689-1.

Tchakonte:2014:DDE

[TAN14] Siméon Tchakonté, Gideon A. Ajeagah, and Pierre Ngassam. Diversity, dynamic and ecology of freshwater snails related to environmental factors in urban and suburban streams in Douala–Cameroon (Central Africa). Aquatic Ecology, 48(4): 379–395, December 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-014-9491-2.

Taylor:1993:ZFE

[Tay93] C. J. L. Taylor. The zooplankton of the forth estuary. Aquatic Ecology, 27(2-4):87–99, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334772.

Tasker:1992:IHA

[TB92] Mark L. Tasker and Peter H. Becker. Influences of human activities on seabird populations in the North Sea. Aquatic Ecology, 26(1):59–73, November 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02298028.

Thomas:2022:PAC

[TB22] Olivier P. Thomas and Anne-Geneviève Bagnères. Preface: Aquatic chemical ecology special issue. Aquatic Ecology, 56 (2):337–338, June 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09968-4.

Tadesse:2003:EAD

[TBA03] Zenebe Tadesse, Merike Boberg, and Gunnel Ahlgren. Effects of algal diets and temperature on the growth and fatty acid content of the cichlid fish Oreochromis niloticus L. — a laboratory study. Aquatic Ecology, 37(2):169–182, April 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1023942711822.

Thiebaut:2017:TOB

[TBBS17] Gabrielle Thiébaut, Anatole Boiché, and Marie-Hélène Barrat-Segretain. Trade-offs between growth and defence in two phylogenetically close invasive species. Aquatic Ecology, 51(3):405–415, September 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
017-9625-4.

Tomasky:1999:NLP

[TBF99] Gabrielle Tomasky, Jeri Barak, and Kenneth Foreman. Nutrient limitation of phytoplankton growth in Waquoit Bay, Massachusetts, USA: a nutrient enrichment study. Aquatic Ecology, 33(2):147–155, September 1999. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009912430833.

Testi:2009:DRE

[TBF09] Anna Testi, Sara Bisceglie, and Giuliano Fanelli. Detecting river environmental quality through plant and macroinvertebrate bioindicators in the Aniene River (Central Italy). Aquatic Ecology, 43(2):477–486, June 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-008-9205-8.

Thom:2018:PIM

[TBS18] Alexandra E. Thom, Meenakshi Bhattacharjee, and Evan Siemann. Predaceous insects may limit algal grazers in highnutrient environments: implications for wastewater remediation in open bioreactors. Aquatic Ecology, 52(1):107–118, March 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-018-9648-5.

Torres-Bejarano:2022:RVS

[TBSZ22] Angélica M. Torres-Bejarano, S. Mažeika Patricio Sulliván, and Gabriel J. Colorado Z. Riparian vegetation structure and seasonality influence functional diversity more than taxonomic diversity of stream fish assemblages in the Colombian Amazon. Aquatic Ecology, 56(1):153–172, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09904-y.

Teoh:2015:VHC

[TC15] Hong Wooi Teoh and Ving Ching Chong. Versatile hermit crabs harness multiple-source energy from coastal mudflats: implications for fish production. Aquatic Ecology, 49(1):43– 55, March 2015. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-015-9503-x.

Tolonen:2018:EFS

[TCH18] Kimmo T. Tolonen, Yongjiu Cai, and Jani Heino. Environmental filtering and spatial effects on metacommunity organisation differ among littoral macroinvertebrate groups deconstructed by biological traits. Aquatic Ecology, 52(1):119–131, March 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-018-9649-4.

Tang:2022:RER

[TCL22] Haoran Tang, Fangyuan Chen, and Yanjing Lou. Responses of early recruitment processes with rhizome to flooding depth and salinity in Manchurian wild rice (Zizania latifolia). Aquatic Ecology, 56(3):619–629, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09927-5.

TenWinkel:1987:PDA

[TD87a] Erik H. Ten Winkel and C. Davids. Population dynamic aspects of chironomid larvae of the littoral zone of Lake Maarsseveen I. Aquatic Ecology, 21(1):81–94, July 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255458.

Tundisi:1987:M

[TD87b] J. G. Tundisi and H. J. Dumont. Miscellaneous. Aquatic Ecology, 21(1):121, July 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02255462.

Tonkin:2013:PDR

[TDB13] Jonathan D. Tonkin, Russell G. Death, and José Barquín. Productivity–diversity relationships for stream invertebrates differ geographically. Aquatic Ecology, 47(1):109–121, March 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-013-9429-0.

Tubbing:1995:PDR

[TDBW95] D. M. J. Tubbing, D. De Zwart, and T. Burger-Wiersma. Phytoplankton dynamics in the River Meuse as affected by pollution. Aquatic Ecology, 29(1):103–116, April 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02061793.

Tkachenko:2023:CRC

[TDHH23] Konstantin S. Tkachenko, Vu V. Dung, Vo T. Ha, and Nguyen H. Huan. Coral reef collapse in South-Central Vietnam: a consequence of multiple negative effects. Aquatic Ecology, 57(1):65–83, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
022-09994-2.

Tian:2023:RCC

[TDL23] Yaqi Tian, Jianming Deng, and Ming Li. Response of the cellular components to environmental factors indicates limiting factors of Microcystis in Lake Taihu. Aquatic Ecology, 57(1):99–113, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09997-z.

Tolhurst:2006:ETD

[TDP06] T. J. Tolhurst, E. C. Defew, and D. M. Paterson. The effects of tidally-driven temporal variation on measuring intertidal cohesive sediment erosion threshold. Aquatic Ecology, 40(4): 521–531, December 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-005-9001-7.

Thorarinsson:2004:DHG

[TE04] Thorkell Lindberg Thórarinsson and Árni Einarsson. Dispersion of the horned grebe Podiceps auritus (L.) (Aves) on Lake Myvatn, Iceland, in late summer. Aquatic Ecology, 38(2):309–315, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.
0000032056.43542.c1.

Teesalu:2023:BRI

[TET23] Paul Teesalu, Fabio Ercoli, and Arvo Tuvikene. Behavioural responses of invasive (Gmelinoides fasciatus) and native (Gammarus lacustris) amphipods to predators on different bottom substrates. Aquatic Ecology, 57(1):139–147, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-022-09999-x.

Taheri:2024:EDG

[TFH^{+}24] Mehrshad Taheri, Maryam Yazdani Foshtomi, Mohammad Ali Hamzeh, Ahmad Manbohi, and Reza Rahnama Haratbar. Effects of discarded garbage bags on intertidal free living nematode community. Aquatic Ecology, 58(3):853–863, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10109-2.

Tolhurst:2006:ERE

[TFP06] T. J. Tolhurst, P. L. Friend, and D. M. Paterson. The effects of rain on the erosion threshold of intertidal cohesive sediments. Aquatic Ecology, 40(4):533–541, December 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-8058-z.

Tackx:1991:SDT

[TFR91] M. L. M. Tackx, J. W. Francke, and J. Rijk. Size distributions and trophic relationship of the pelagic ecosystem in the Oosterschelde (S.W. Netherlands). Aquatic Ecology, 25 (1):9–14, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259583.

Tada:2010:SFC

[TFT10] Yohei Tada, Katsunori Fujikura, and Kazushige Tanabe. In situ fluorochrome calcein marking of deep-sea molluscs using a new growth chamber. Aquatic Ecology, 44(1):217–222, March 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9290-3.

Tolkamp:1988:DBW

[TG88] Harry H. Tolkamp and Jean J. P. Gardeniers. The development of biological water quality assessment in The Netherlands. Aquatic Ecology, 22(1):87–91, August 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256789.

Thorbergsdottir:2004:ILN

[TG04] Ingunn María Thorbergsdóttir and Sigurdur Reynir Gíslason. Internal loading of nutrients and certain metals in the shallow eutrophic Lake Myvatn, Iceland. Aquatic Ecology, 38(2):191– 208, June 2004. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/B:AECO.0000032066.65901.1f.

Tootoonchi:2019:TSS

[TG19] Mohsen Tootoonchi and Lyn A. Gettys. Testing salt stress on aquatic plants: effect of salt source and substrate. Aquatic Ecology, 53(3):325–334, September 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09692-6.

Thiebaut:2016:GRC

[TGD16] G. Thiébaut, M. Gillard, and C. Deleu. Growth, regeneration and colonisation of Egeria densa fragments: the effect of autumn temperature increases. Aquatic Ecology, 50(2):175– 185, June 2016. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-016-9566-3.

Thorbergsdottir:2004:BOF

[TGE04] Ingunn María Thorbergsdóttir, Sigurdur Reynir Gíslason, and Árni Einarsson. Benthic oxygen flux in the highly productive subarctic Lake Myvatn, Iceland: In situ benthic flux chamber study. Aquatic Ecology, 38(2):177–189, June 2004. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000032057.95464.ad.

Tober:2000:GAF

[TGV00] Joanna D. Tober, Martin P. A. Griffin, and Ivan Valiela. Growth and abundance of Fundulus heteroclitus and Menidia menidia in estuaries of Waquoit Bay, Massachusetts exposed to different rates of nitrogen loading. Aquatic Ecology, 34(3): 299–306, September 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009982208382.

Thackeray:2006:SQE

[TGW06] Stephen J. Thackeray, D. Glen George, and Ian J. Winfield. Statistical quantification of the effect of thermal stratification on patterns of dispersion in a freshwater zooplankton community. Aquatic Ecology, 40(1):23–32, March 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9021-3.

Terzhevik:2009:SFT

[TGZ09] A. Terzhevik, S. Golosov, and I. Zverev. Some features of the thermal and dissolved oxygen structure in boreal, shallow icecovered Lake Vendyurskoe, Russia. Aquatic Ecology, 43(3): 617–627, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9288-x.

tenHoor:1971:IVT

[tH71] Antje ten Hoor. Isoleren van Thiobacillus. (Dutch) [Isolate from Thiobacillus]. Aquatic Ecology, 5(2):107–108, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF03185954.

Torres:2016:TFO

[TH16] Kryztal M. Medina Torres and Christopher L. Higgins. Taxonomic and functional organization in metacommunity structure of stream-fish assemblages among and within river basins in Texas. Aquatic Ecology, 50(2):247–259, June 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9572-5.

Therriault:2002:TPD

[The02] Thomas W. Therriault. Temporal patterns of diversity, abundance and evenness for invertebrate communities from coastal freshwater and brackish water rock pools. Aquatic Ecology, 36(4):529–540, December 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1021178629707.

Tix:2017:ECD

[THS17a] John A. Tix, Caleb T. Hasler, and Cory D. Suski. Elevated carbon dioxide has the potential to impact alarm cue responses in some freshwater fishes. Aquatic Ecology, 51(1):59–72, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9598-8. See

erratum [THS17b].

Tix:2017:EEC

[THS17b] John A. Tix, Caleb T. Hasler, and Cory D. Suski. Erratum to: Elevated carbon dioxide has the potential to impact alarm cue responses in some freshwater fishes. Aquatic Ecology, 51(1):73, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-016-9608-x. See

[THS17a].

Touchette:2010:ERF

[TIF10] B. W. Touchette, L. R. Iannacone, and A. Frank. Ecophysiological responses of five emergent-wetland plants to diminished water supply: an experimental microcosm study. Aquatic Ecology, 44(1):101–112, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9246-7.

Traas:1998:FWM

[TJB98] Theo P. Traas, Jan H. Janse, and Theo C. M. Brock. A food web model for fate and direct and indirect effects of Dursban(R) 4E (active ingredient chlorpyrifos) in freshwater microcosms. Aquatic Ecology, 32(2):179–190, September 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009920226083.

Tang:2024:APE

[TJZ^{+}24] Liang Tang, Haifeng Jiang, Zhuofan Zhou, Zhenling Zhang, Shanshan Hou, and Bo Liu. The adaptation path of environmental governance in the Yellow River Basin for regional ecological improvement. Aquatic Ecology, 58(3):535–549, September 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10089-3.

Taheruzzaman:1988:ESE

[TK88a] Quazi Taheruzzaman and Debi Prosad Kushari. The effect of sewage-enriched river Ganga water on the biomass production of the Azolla-Anabaena complex. Aquatic Ecology, 22(2):173–181, December 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256821.

Taheruzzaman:1988:PRE

[TK88b] Quazi Taheruzzaman and Debi Prosad Kushari. Production rate of Eichhornia crassipes (Mart.) Solms in river Ganga water. Aquatic Ecology, 22(2):163–171, December 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256820.

Taheruzzaman:1989:ESC

[TK89] Quazi Taheruzzaman and D. P. Kushari. Evaluation of some common aquatic macrophytes cultivated in enriched water as possible source of protein and biogas. Aquatic Ecology, 23(2):207–212, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256739.

Taheruzzaman:1995:BCM

[TK95] Q. Taheruzzaman and D. P. Kushari. Biomass and concentrations of macronutrients and mercury in Azolla pinnata R.Br. in Indian ponds enriched by anthropogenic effluents. Aquatic Ecology, 29(2):157–160, June 1995. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02336045.

Tilahun:2019:INN

[TK19a] Samson Tilahun and Demeke Kifle. The influence of El Niño-induced drought on cyanobacterial community structure in a shallow tropical reservoir (Koka Reservoir, Ethiopia). Aquatic Ecology, 53(1):61–77, March 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09673-9.

Tomatsuri:2019:CIO

[TK19b] Morihiko Tomatsuri and Koetsu Kon. Correction to: Impacts of ocean acidification on hermit crab communities through contrasting responses of Pagurus filholi (de Man, 1887) and Clibanarius virescens (Krauss, 1843). Aquatic Ecology, 53(4):745–746, December 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09716-1. See [TK19c].

Tomatsuri:2019:IOA

[TK19c] Morihiko Tomatsuri and Koetsu Kon. Impacts of ocean acidification on hermit crab communities through contrasting responses of Pagurus filholi (de Man, 1887) and Clibanarius virescens (Krauss, 1843). Aquatic Ecology, 53(4):569–580, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.springer.

com/article/10.1007/s10452-019-09709-0. See correc-

tion [TK19b].

Triest:2001:CMD

[TKD01] Ludwig Triest, Parminder Kaur, and Niels De Pauw. Comparative monitoring of diatoms, macroinvertebrates and macrophytes in the Woluwe River (Brussels, Belgium). Aquatic Ecology, 35(2):183–194, June 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011468615246.

Tolonen:2020:DED

[TKH20] Kimmo T. Tolonen, Juha Karjalainen, and Jani Heino. Do the ecological drivers of lake littoral communities match and lead to congruence between organism groups? Aquatic Ecology, 54(3):839–854, September 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09781-x.

Triebskorn:2003:CWA

[TKM03] R. Triebskorn, Heinz-R. Köhler, and E. Müller. Comments on Wijnhoven et al. (2003). Aquatic Ecology, 37(4):447–448, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.springer.

com/article/10.1023/B:AECO.0000007073.86590.dd. See

[WvRvdV03a] and reply [WvRvdV03b].

Titlyanov:2010:TCP

[TKR10] Eduard A. Titlyanov, Serguei I. Kiyashko, and John A. Raven. δ^{13}C and δ^{15}N in tissue of coral polyps and epilithic algae inhabiting damaged coral colonies under the influence of different light intensities. Aquatic Ecology, 44(1):13–21, March 2010. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9248-5.

Texier:1993:OMT

[TLC93] Hervé Texier, Robert Lafite, and Bruno Charriere. Organic matter transport across a macrotidal estuary grandient: the Seine estuary, France. Aquatic Ecology, 27(2-4):405–413, June 1993. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334802.

Tavares:2019:ERP

[TLF19] Diego Alberto Tavares, Richard Wilander Lambrecht, and Carla Ferragut. Epipelon responses to N and P enrichment and the relationships with phytoplankton and zooplankton in a mesotrophic reservoir. Aquatic Ecology, 53(2):303–314, June 2019. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09690-8.

Tolonen:2018:EUM

[TLH18] Katri E. Tolonen, Kirsti Leinonen, and Jani Heino. Ecological uniqueness of macroinvertebrate communities in highlatitude streams is a consequence of deterministic environmental filtering processes. Aquatic Ecology, 52(1):17–33, March 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-017-9642-3.

TenWinkel:1984:EFU

[TM84] Erik H. Ten Winkel and John T. Meulemans. Effects of fish upon submerged vegetation. Aquatic Ecology, 18(2):157–158, December 1984. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257054.

Taybi:2021:AHG

[TMM21] Abdelkhaleq Fouzi Taybi, Youness Mabrouki, and Andrés Millán. Are aquatic Hemiptera good indicators of environmental river conditions? Aquatic Ecology, 55(3):791–806, September 2021. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09861-6.

Togashi:2008:EGB

[TNC08] Tatsuya Togashi, Masaru Nagisa, and Paul Alan Cox. Effects of gamete behavior and density on fertilization success in marine green algae: insights from three-dimensional numerical simulations. Aquatic Ecology, 42(3):355–362, September 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9105-3.

Twisk:2000:EDM

[TNtK00] W. Twisk, M. A. W. Noordervliet, and W. J. ter Keurs. Effects of ditch management on caddisfly, dragonfly and amphibian larvae in intensively farmed peat areas. Aquatic Ecology, 34(4):397–411, December 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011430831180.

Twisk:2003:NVD

[TNtK03] W. Twisk, M. A. W. Noordervliet, and W. J. ter Keurs. The nature value of the ditch vegetation in peat areas in relation to farm management. Aquatic Ecology, 37(2):191–209, April 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1023944028022.

Tokeshi:1992:SDE

[Tok92] Mutsunori Tokeshi. The structure of diversity in an epiphytic chironomid community. Aquatic Ecology, 26(2-4):461– 470, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255276.

Tolkamp:1982:MML

[Tol82] Harry H. Tolkamp. Microdistribution of macroinvertebrates in lowland streams. Aquatic Ecology, 16(2-3):133–148, December 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255367.

Tolkamp:1984:BPI

[Tol84] Harry H. Tolkamp. A bibliography of publications for identification of freshwater organisms in the Netherlands. Aquatic Ecology, 18(2):141–155, December 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02257053.

Tolomeyev:2002:PDA

[Tol02] Alexander P. Tolomeyev. Phytoplankton diet of Arctodiaptomus salinus (Copepoda, Calanoida) in Lake Shira (Khakasia). Aquatic Ecology, 36(2):229–234, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1015618104520.

Timmermann:2009:PRL

[TP09] Mirna Timmermann and Martin Plath. Phototactic response and light sensitivity in an epigean and a hypogean population of a barb (Garra barreimiae, cyprinidae). Aquatic Ecology, 43(2):539–547, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9173-z.

Timofeyev:2009:CAA

[TPS09] Maxim A. Timofeyev, Marina Protopopova, and Christian E. W. Steinberg. Can acclimation of amphipods change their antioxidative response? Aquatic Ecology, 43(4):??, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-008-9217-4.

Ternon:2022:AAA

[TPT22] Eva Ternon, Anne-Sophie Pavaux, and Olivier P. Thomas. Assessment of the allelochemical activity of Ostreopsis cf. ovata and the ovatoxins towards competitive benthic microalgae. Aquatic Ecology, 56(2):475–491, June 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09953-x.

Torres-Rojas:2010:SCA

[TRHHAR10] Yassir Edén Torres-Rojas, Agustín Hernández-Herrera, and Vanessa Guadalupe Alatorre-Ramírez. Stomach content analysis of juvenile, scalloped hammerhead shark Sphyrna lewini captured off the coast of Mazatlán, Mexico. Aquatic Ecology, 44(1):301–308, March 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9245-8.

Theis:2022:MBD

[TRP22] Sebastian Theis, Jonathan L. W. Ruppert, and Mark S. Poesch. Measuring beta diversity components and beneficial effects of coarse woody habitat introduction on invertebrate and macrophyte communities in a shallow northern boreal lake; implications for offsetting. Aquatic Ecology, 56(3): 793–814, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09949-7.

Takahashi:2017:SVP

[TS17] Kae Takahashi and Takuya Sato. Spatial variability in prey phenology determines predator movement patterns and prey survival. Aquatic Ecology, 51(3):377–388, September 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9623-6.

Tavsanoglu:2017:ENW

[TŠB17] Ülkü Nihan Tavşanoğlu, Michal Šorf, and Meryem Beklioğlu. Effects of nutrient and water level changes on the composition and size structure of zooplankton communities in shallow lakes under different climatic conditions: a pan-European mesocosm experiment. Aquatic Ecology, 51 (2):257–273, June 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9615-6.

Triest:2016:BNB

[TSV16] Ludwig Triest, Iris Stiers, and Stijn Van Onsem. Biomanipulation as a nature-based solution to reduce cyanobacterial blooms. Aquatic Ecology, 50(3):461–483, September 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9548-x.

Tolomeev:2010:FSA

[TSZ10] A. P. Tolomeev, N. N. Sushchik, and T. A. Zotina. Feeding spectra of Arctodiaptomus salinus (Calanoida, Copepoda) using fatty acid trophic markers in seston food in two salt lakes in South Siberia (Khakasia, Russia). Aquatic Ecology, 44(3):513–530, September 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9331-y.

Temerova:2002:GDZ

[TTD02] Tatiana A. Temerova, Alexander P. Tolomeyev, and Andrei G. Degermendzhy. Growth of dominant zooplankton species feeding on plankton microflora in Lake Shira. Aquatic Ecology, 36(2):235–243, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1015607304508.

Tackx:1985:CRG

[TV85] M. L. M. Tackx and E. M. Van De Vrie. Calculations of results in grazing experiments using the counting method. Aquatic Ecology, 19(1):29–36, November 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255090.

Timmermans:1991:TMB

[TVD91] Klaas R. Timmermans, Bert Van Hattum, and Kees Davids. Trace metals in the benthic habitat of the Maarsseveen Lakes system, The Netherlands. Aquatic Ecology, 24(2):153–164, April 1991. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260433.

Tietjen:2003:EEC

[TW03] Todd Tietjen and Robert G. Wetzel. Extracellular enzymeclay mineral complexes: Enzyme adsorption, alteration of enzyme activity, and protection from photodegradation. Aquatic Ecology, 37(4):331–339, October 2003. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/B:AECO.0000007044.52801.6b.

Tang:2013:DRS

[TWC13] Tao Tang, Naicheng Wu, and Qinghua Cai. Disentangling the roles of spatial and environmental variables in shaping benthic algal assemblages in rivers of central and northern China. Aquatic Ecology, 47(4):453–466, December 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9458-8.

Tang:2024:PLR

[TXW^{+}24] Quehui Tang, Lei Xu, Lianggen Wang, Jiajia Ning, Delian Huang, Yafang Li, Shuangshuang Liu, and Feiyan Du. Positive linear relationship between phytoplankton diversity and productivity in an artificial reef ecosystem. Aquatic Ecology, 58(4):1267–1279, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
024-10140-3.

Taguchi:2002:MPD

[TYN02] Koichi Taguchi, Susumu Yamochi, and Yoshiharu Nakamura. Modelling population dynamics of the pelagic larval shrimp Metapenaeus ensis in the Osaka Bay estuary. Aquatic Ecology, 36(1):21–40, January 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1013342504789.

Tolomeyev:2005:SMI

[TZ05] Alexander P. Tolomeyev and Yegor S. Zadereev. An in situ method for the investigation of vertical distributions of zooplankton in lakes: test of a two-compartment enclosure. Aquatic Ecology, 39(2):181–188, June 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-5732-0.

Tang:2021:VND

[TZD21] Yali Tang, Daiying Zhou, and Henri J. Dumont. Vallisneria natans detritus supports Daphnia magna somatic growth and reproduction under addition of periphyton. Aquatic Ecology, 55(2):579–588, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
021-09846-5.

Umetsu:2012:CRG

[UET12] Cristiane Akemi Umetsu, Heloísa Beatriz Antoniazi Evangelista, and Sidinei Magela Thomaz. The colonization, regeneration, and growth rates of macrophytes from fragments: a comparison between exotic and native submerged aquatic species. Aquatic Ecology, 46(4):443–449, December 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9413-0.

Ueno:2005:ACS

[UKS05] Hiroyuki Ueno, Toshiya Katano, and Masahito Sugiyama. Abundance and community structure of picoplankton and protists in the microbial food web of Barguzin Bay, Lake Baikal. Aquatic Ecology, 39(3):263–270, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-6057-3.

Urrutia-Olvera:2021:PCD

[UOJGRL21] Ashley Urrutia-Olvera, Adán Guillermo Jordán-Garza, and Rigoberto Rosas-Luis. Prey contribution to the diet of Octopus insularis (Leite and Haimovici, 2008) using stable isotopes and stomach content analysis in the Western Gulf of Mexico. Aquatic Ecology, 55(3):765–777, September 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09859-0.

Utevsky:2016:PSM

[UT16] Serge Utevsky and Peter Trontelj. Phylogeography of the southern medicinal leech, Hirudo verbana: a response to živić et al. (2015). Aquatic Ecology, 50(1):97–100, March 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-015-9553-0.

Vaas:1975:IAA

[Vaa75] K. F. Vaas. Immigrants among the animals of the Deltaarea of the SW. Netherlands. Aquatic Ecology, 9(3):114–119, December 1975. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02263329.

Vaas:1979:SFF

[Vaa79] K. F. Vaas. Studies on the fish fauna of an estuary in the S. W. Netherlands, before and after its change into the stagnant, saline, Lake Grevelingen. Aquatic Ecology, 13(2-3): 177–188, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02284753.

Villar-Argaiz:2020:BPC

[VALRdF20] Manuel Villar-Argaiz, Manuel Jesús López-Rodríguez, and José Manuel Tierno de Figueroa. Body P content increases over ontogeny in hemimetabolous macroinvertebrates in a Mediterranean high mountain stream. Aquatic Ecology, 54(4):1185–1200, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
020-09802-9.

VanDam:1974:SDB

[Van74] Herman Van Dam. The suitability of diatoms for biological water assessment. Aquatic Ecology, 8(3):274–284, December 1974. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02257503.

VanDerVelde:1975:ITF

[Van75a] G. Van Der Velde. The immigrant triclad flatworm Dugesia tigrina (Girard) (Plathelminthes, Turbellaria). Range-extension and ecological position in The Netherlands. Aquatic Ecology, 9(3):123–130, December 1975. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02263331.

VanZon:1975:GCW

[Van75b] J. C. J. Van Zon. The grass carp for weed control. Aquatic Ecology, 9(2):88–89, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02257642.

VanDoorn:1976:RFR

[Van76] N. C. Van Doorn. Report of the fish research in a ditch at Tienhoven on 25th August, 1975. Aquatic Ecology, 10(1):74– 77, April 1976. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02308504.

VanArkel:1977:SRS

[Van77a] M. A. Van Arkel. Some results of a survey of the macrobenthic fauna of the Ems–Dollard estuary (Eastern Dutch Wadden Sea). Aquatic Ecology, 11(1):15–16, April 1977. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02282208.

VanEs:1977:MED

[Van77b] F. B. Van Es. Mineralization in the Ems–Dollard estuary effects on carbon and oxygen budgets. Aquatic Ecology, 11(1):

12, April 1977. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02282204.

VanBennekom:1978:BDW

[Van78a] A. J. Van Bennekom. The Balgzand (Dutch Wadden Sea) tidal flats as a source of soluble silicate. Aquatic Ecology, 12(2):143–144, July 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260716.

VanDerHeide:1978:SFP

[Van78b] J. Van Der Heide. Stratification in the filling phase of the man-made lake Brokopondo in Surinam (S. America). Aquatic Ecology, 12(2):85–98, July 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260709.

VanUrk:1978:MRI

[Van78c] G. Van Urk. The macrobenthos of the river IJssel. Aquatic Ecology, 12(1):21–29, April 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02260790.

VanDam:1979:DWQ

[Van79a] Herman Van Dam. Diatoms and water quality in lowland streams in the province of Northern Brabant (The Netherlands). Aquatic Ecology, 13(1):13–21, April 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260303.

VanUrk:1979:ETS

[Van79b] G. Van Urk. The effects of a temperature shock on zooplankton. Aquatic Ecology, 13(2-3):101–105, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284740.

VanDam:1982:UMS

[Van82a] Herman Van Dam. On the use of measures of structure and diversity in applied diatom ecology. Aquatic Ecology, 16(2-3):288, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255385.

VanDenHeiligenberg:1982:EMM

[Van82b] Taco Van Den Heiligenberg. The effects of mechanical and manual digging for lugworms (Arenicola marina L.) on the benthic fauna of the tidal flats in the Dutch Wadden Sea. Aquatic Ecology, 16(2-3):291–292, December 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255388.

VanDam:1984:GLI

[Van84a] Herman Van Dam. A guide to the literature for the identification of freshwater diatoms in the Netherlands. Aquatic Ecology, 18(1):11–16, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256668.

VanDonk:1984:FIP

[Van84b] Ellen Van Donk. Factors influencing phytoplankton growth and succession in Lake Maarsseveen I. Aquatic Ecology, 18(1):69–71, June 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256678.

VanIerland:1985:UFG

[Van85] E. T. Van Ierland. Use of the flowcytometer in grazing studies. Aquatic Ecology, 19(1):37–39, November 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255091.

VanLiere:1986:LLO

[Van86] Louis Van Liere. Loosdrecht Lakes, origin, eutrophication, restoration and research programme. Aquatic Ecology, 20 (1-2):9–15, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02291146.

VanDonk:1987:WQT

[Van87] Ellen Van Donk. The water quality of the two Maarsseveen Lakes in relation to their hydrodynamics. Aquatic Ecology, 21(1):17–24, July 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255451.

VanderLinden:1989:RSN

[Van89] M. J. H. A. Van der Linden. Release of sedimentary nitrogen and phosphorus in polder ditches of a low-moor peat area. Aquatic Ecology, 23(2):125–134, December 1989. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256729.

VanDerVelde:1991:PDP

[Van91] G. Van Der Velde. Population dynamics and population structure of Ferrissia wautieri (Mirolli, 1960) (Gastropoda, Ancylidae) in a pond near Nijmegen (The Netherlands). Aquatic Ecology, 24(2):141–144, April 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260431.

VanDam:1996:PRM

[Van96] Herman Van Dam. Partial recovery of moorland pools from acidification: Indications by chemistry and diatoms. Aquatic Ecology, 30(2-3):203–218, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02272240.

VanDonk:1997:DPA

[Van97] Ellen Van Donk. Defenses in phytoplankton against grazing induced by nutrient limitation, UV-b stress and infochemicals. Aquatic Ecology, 31(1):53–58, March 1997. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009951622185.

VandenBrink:1998:MAS

[VB98] Paul J. Van den Brink and Cajo J. F. Ter Braak. Multivariate analysis of stress in experimental ecosystems by Principal Response Curves and similarity analysis. Aquatic Ecology, 32(2):163–178, September 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009944004756.

Vieite:2022:FKC

[VBA22] Beatriz Vieite, Heitor O. Braga, and Ulisses M. Azeiteiro. Fishermen’s knowledge and conservation attitudes: focus on the great cormorant Phalacrocorax carbo (Linnaeus, 1758) in the Minho River, Portugal. Aquatic Ecology, 56(3):667–684, September 2022. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-021-09928-4.

Verhoeven:1996:CPB

[VBD96] J. T. A. Verhoeven, B. Beltman, and H. De Caluwe. Changes in plant biomass in fens in the vechtplassen area, as related to nutrient enrichment. Aquatic Ecology, 30(2-3):227–237, October 1996. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02272242.

Henegouwen:1975:M

[vBHvdVH75] A. L. van Berge Henegouwen, C. van der Velde, and L. W. G. Higler. Miscellanea. Aquatic Ecology, 9(1):48–50, April 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02257519.

VanDerVelde:1996:CHO

[VBK96] G. Van Der Velde, T. C. M. Brock, and A. J. Kempers. Cornelis den hartog: an outstanding aquatic ecologist. Aquatic Ecology, 30(2-3):71–82, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02272229.

VanderHoek:1989:LCG

[VC89] W. F. Van der Hoek and J. G. M. Cuppen. Life cycle and growth of Trichostegia minor (Curtis) in temporary woodland pools (Trichoptera: Phryganeidae). Aquatic Ecology, 23(2):161–168, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256733.

Vaage:2013:SSS

[VCT13] Selina Våge, Marco Castellani, and T. Frede Thingstad. Successful strategies in size structured mixotrophic food webs. Aquatic Ecology, 47(3):329–347, September 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9447-y.

vanDensen:1983:FES

[vD83] W. L. T. van Densen. Fish ecological studies in tjeukemeer and fishery management. Aquatic Ecology, 17(1):59–65, June 1983. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255192.

vanDam:1984:P

[vD84] Herman van Dam. Publications. Aquatic Ecology, 18(2): 165–173, December 1984. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257056.

VanDiepen:1985:ZMP

[VD85] Jan Van Diepen and Cees Davids. Zebra mussels and polystyrene. Aquatic Ecology, 19(2):179–181, December 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02270764.

Vegter:1987:NPP

[VD87] F. Vegter and P. R. M. De Visscher. Nutrients and phytoplankton primary production in the marine tidal Oosterschelde estuary (The Netherlands). Aquatic Ecology, 21(2): 149–158, December 1987. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255440.

Vos:1994:PRD

[VD94] Peter C. Vos and Hein De Wolf. Palaeoenvironmental research on diatoms in early and middle Holocene deposits in central North Holland (The Netherlands). Aquatic Ecology, 28(1):97–115, March 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334250.

vanDonk:1997:BOA

[vD97] Ellen van Donk. Book review: Biogeography of Freshwater Algae, J. Kristiansen. Aquatic Ecology, 31(3):342, September 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009956010333.

vanderBurgh:1970:OHG

[vdB70] P. van der Burgh. Op hydrobiologisch gebied werkzaam zijnde instituten, deinsten en werkgroepen met korte omschrijving van het onderzoek. (Dutch) [Institutes, organizations and working groups active in the hydrobiological field with a brief description of the research]. Aquatic Ecology, 4(3): 131–137, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185872.

vanDam:1978:CMP

[vDB78] H. van Dam and H. Kooyman-Van Blokland. Changes in the moorland pools near Oisterwijk between 1840 and 1975. Aquatic Ecology, 12(1):54–55, April 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260794.

VandeLannoote:1981:EAW

[VDB81] A. Vande Lannoote, G. De Gueldre, and B. Bruylants. Ecological assessment of water quality: Comparison of biological-ecological procedures in a rain-fed lowland waterway (Kleine nete, N. Belgium). Aquatic Ecology, 15(3): 161–164, December 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02255174.

Vlasblom:1977:BR

[VDD77] A. G. Vlasblom, Jaap Dorgelo, and C. Davids. Book review. Aquatic Ecology, 11(1):25–29, April 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02282213.

vanDam:1988:P

[vDDD88] H. van Dam, Jaap Dorgelo, and C. Davids. Publications. Aquatic Ecology, 22(1):99–106, August 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256792.

vanDam:1987:P

[vDDS87] H. van Dam, Niels Daan, and P. Schroevers. Publications. Aquatic Ecology, 21(1):112–120, July 1987. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255461.

vanderEyk:1970:SGB

[vdE70] R. H. van der Eyk. Sociale groepsvorming bij Gyrinus marinus gill. (Dutch) [Social group formation at Gyrinus marinus gill]. Aquatic Ecology, 4(3):157–169, November 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF03185880.

vanDam:1994:MPR

[vDG94] G. C. van Dam and M. J. J. M. Geurtz. Model particles as representatives of particles in nature. Aquatic Ecology, 28 (3-4):317–328, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334200.

vanderHeiden:2017:BAH

[vdHD17] Craig A. van der Heiden and Nathan J. Dorn. Benefits of adjacent habitat patches to the distribution of a crayfish population in a hydro-dynamic wetland landscape. Aquatic Ecology, 51(2):219–233, June 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9612-1.

VanBeusekom:1994:RSM

[VdJ94] J. E. E. Van Beusekom and V. N. de Jonge. The role of suspended matter in the distribution of dissolved inorganic phosphate, iron and aluminium in the Ems estuary. Aquatic Ecology, 28(3-4):383–395, September 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334208.

vanDuren:2006:P

[vDJ06] Luca A. van Duren and Per R. Jonsson. Preface. Aquatic Ecology, 40(4):407–408, December 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
006-9067-x.

vanDam:1980:IAD

[vDSB80] Herman van Dam, Gillis Suurmond, and Cajo Ter Braak. The impact of acidification on diatoms and chemistry of Dutch moorland pools. Aquatic Ecology, 14(3):219, December 1980. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260125.

vanderTol:1997:MAE

[vdTS97] M. W. M. van der Tol and H. Scholten. A model analysis on the effect of decreasing nutrient loads on the biomass of benthic suspension feeders in the Oosterschelde ecosystem (SW Netherlands). Aquatic Ecology, 31(4):395–408, December 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009956812400.

vdVlugt:1976:CLR

[vdV76] J. C. v. d. Vlugt. Comparative limnological research in “De Grote Rug” and model reservoirs. Aquatic Ecology, 10(3): 136–144, December 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02263362.

vanderVelde:1981:PND

[vdV81] G. van der Velde. A project on nymphaeid-dominated systems. Aquatic Ecology, 15(3):185–189, December 1981. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255178.

VanDensen:1986:FCL

[VDV86] W. L. T. Van Densen, C. Dijkers, and R. Veerman. The fish community of the Loosdrecht Lakes and the perspective for biomanipulation. Aquatic Ecology, 20(1-2):147–163, November 1986. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02291159.

vanderVelde:1999:BRF

[vdV99] Gerard van der Velde. Book review: Fish biology in Japan: an Anthology in Honour of Hiroya Kawanabe. Edited by M. Yuma, I. Nakamura and K. D. Fausch. 1998. Aquatic Ecology, 33(4):399, December 1999. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009958712820.

vanderVelde:2001:BRB

[vdV01] Gerard van der Velde. Book review: The Biology of Mangroves. By Peter J. Hogarth. Aquatic Ecology, 35(1):88, March 2001. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1011471526905.

vanderVelde:1996:P

[vdVB96] Gerard van der Velde and Theo C. M. Brock. Preface. Aquatic Ecology, 30(2-3):i, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02272228.

vanderVelde:1976:ABR

[vdVP76] G. van der Velde and P. J. G. Polderman. Atherina boyeri Risso, a genuine immigrant in the Delta Area. Aquatic Ecology, 10(2):96–97, September 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02282216.

vanderWeij:1970:MVD

[vdW70] H. G. van der Weij. De mogelijkheden van diuron voor het gezond houden van stilstaand water. (Dutch) [The possibilities of diuron for keeping stagnant water healthy]. Aquatic Ecology, 4(1):3–9, March 1970. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF03185860.

Vegter:1979:PPP

[Veg79] F. Vegter. Phytoplankton primary production in Lake Grevelingen (II). Aquatic Ecology, 13(2-3):77–78, December 1979.

CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284729.

Vesely:2020:CDF

[VEK20] Lukáš Veselý, Fabio Ercoli, and Antonín Kouba. The crayfish distribution, feeding plasticity, seasonal isotopic variation and trophic role across ontogeny and habitat in a canyon-shaped reservoir. Aquatic Ecology, 54(4):1169–1183, December 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09801-w.

Veldkamp:1972:IDI

[Vel72] H. Veldkamp. Inleiding. (Dutch) [Introduction]. Aquatic Ecology, 6(1):8–13, March 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02336205.

Vela:1991:NDF

[Vel91] Ligia Vela. Natural diet of fish from Lake Xolotlán (Managua). Aquatic Ecology, 25(2):169–172, June 1991. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291250.

Verheul:1972:TDN

[Ver72] Ir. T. Verheul. De toekomstige drinkwatervoorziening in Nederland. (Dutch) [The future drinking water supply in The Netherlands]. Aquatic Ecology, 6(3):117–124, September 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02314880.

Verwey:1973:ECW

[Ver73] J. Verwey. Environment and cooling water. Aquatic Ecology, 7(3):115–120, September 1973. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02275612.

Verdouw:1976:CCS

[Ver76a] H. Verdouw. Concentration changes of some elements during a year cycle in Lake Vechten. Aquatic Ecology, 10(1):13, April 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02308491.

Verhagen:1976:CMS

[Ver76b] J. H. G. Verhagen. The construction of models or simulation: a guideline for investigations into and control of water quality in reservoirs. Aquatic Ecology, 10(3):172–177, December 1976. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02263366.

Verhoeven:1979:BPR

[Ver79] J. T. A. Verhoeven. Biomass and productivity of Ruppia taxa in Western Europe. Aquatic Ecology, 13(2-3):113–115, December 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02284743.

Verdonschot:1980:DOH

[Ver80a] Piet F. M. Verdonschot. The distribution of oligochaetes in the Haringvliet–Hollands Diep (Delta region of the Netherlands) before and after the closure in 1970. Aquatic Ecology, 14(3):221, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260127.

Verhagen:1980:RHE

[Ver80b] J. H. G. Verhagen. The role of hydrodynamics in eutrophication. Aquatic Ecology, 14(1-2):47–54, August 1980. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260272.

Vermaat:1997:BRM

[Ver97] Jan E. Vermaat. Book review: Management and Ecology of Freshwater Plants, J. M. Caffrey, P. R. F. Barrett, K. J. Murphy and P. M. Wade. Aquatic Ecology, 31(3):341–342, September 1997. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009950826263.

Vanharanta:2020:VPN

[VET20] Mari Vanharanta, Samu Elovaara, and Tobias Tamelander. Viability of pico- and nanophytoplankton in the Baltic Sea during spring. Aquatic Ecology, 54(1):119–135, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09730-3.

Vadadi-Fulop:2012:CCF

[VFSH12] Csaba Vadadi-Fülöp, Csaba Sipkay, and Levente Hufnagel. Climate change and freshwater zooplankton: what does it boil down to? Aquatic Ecology, 46(4):501–519, December 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-012-9418-8.

vanGemerden:1972:FB

[vG72] H. van Gemerden. Fototrofe bacterien. (Dutch) [Phototrophic bacteria]. Aquatic Ecology, 6(1):13–14, March 1972. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02336206.

vanGeldermalsen:1984:MIE

[vG84] La. van Geldermalsen. Measurements of import and export in the Krabbenkreek, a tidal area in the Oosterschelde (S.W. Netherlands). Aquatic Ecology, 18(1):76, June 1984. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256684.

vanGemerden:1985:P

[vGDD85] H. van Gemerden, C. Davids, and Jaap Dorgelo. Publications. Aquatic Ecology, 19(2):217–225, December 1985. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02270769.

vanGool:1996:R

[vGDR96] Erik van Gool, Jaap Dorgelo, and J. Ringelberg. Reviews. Aquatic Ecology, 30(1):61–67, July 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02092148.

VanDonk:1989:FWM

[VGG89] E. Van Donk, R. D. Gulati, and M. P. Grimm. Food web manipulation in Lake Zwemlust: Positive and negative effects during the first two years. Aquatic Ecology, 23(1):19– 34, March 1989. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02286424.

Vivian:2020:RGE

[VGJ20] Lyndsey M. Vivian, Joe Greet, and Christopher S. Jones. Responses of grasses to experimental submergence in summer: implications for the management of unseasonal flows in regulated rivers. Aquatic Ecology, 54(4):985–999, December 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09788-4.

Vijverberg:1993:FWS

[VGM93] J. Vijverberg, R. D. Gulati, and W. M. Mooij. Food-web studies in shallow eutrophic lakes by the Netherlands Institute of Ecology: Main results, knowledge gaps and new perspectives. Aquatic Ecology, 27(1):35–49, July 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02336927.

Verma:2023:EDP

[VGP^{+}23] Chandani R. Verma, Sachin M. Gosavi, Manoj Pise, Lukáš Kalous, and Pradeep Kumkar. Effect of diethyl phthalate on predator–prey chemo-ecology in Lepidocephalichthys thermalis. Aquatic Ecology, 57(3):585–596, September 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10031-z.

vanGemerden:1978:P

[vGPD78] H. van Gemerden, S. Parma, and Jaap Dorgelo. Publications. Aquatic Ecology, 12(1):56–60, April 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260795.

vanGool:1998:QEF

[vGR98] Erik van Gool and Joop Ringelberg. Quantitative effects of fish kairomones and successive light stimuli on downward swimming responses of Daphnia. Aquatic Ecology, 32(4): 291–296, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009917929959.

VanPagee:1982:SEC

[VGV82] J. A. Van Pagee, S. Groot, and J. H. G. Verhagen. Some ecological consequences of a projected deep reservoir in the Kabalebo River in Suriname. Aquatic Ecology, 16(2-3):241–254, December 1982. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255378.

Vrieling:1994:DTP

[VGV94] Engel G. Vrieling, Winfried W. C. Gieskes, and Marten Veenhuis. Detection of toxic phytoplankton species by immunochemical particle analysis based on flow cytometry. Aquatic Ecology, 28(3-4):249–254, September 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334192.

vanGremberghe:2011:GDM

[vGVV11] Ineke van Gremberghe, Katleen Van der Gucht, and Wim Vyverman. Genetic diversity of Microcystis blooms (Cyanobacteria) in recently constructed reservoirs in Tigray (Northern Ethiopia) assessed by rDNA ITS. Aquatic Ecology, 45(2):289–306, May 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9354-z.

VanDerVelde:1988:NDE

[VH88] G. Van Der Velde and L. W. G. Higler. Notes on the distribution and ecology of Hirudinea in The Netherlands. Aquatic Ecology, 22(1):81–85, August 1988. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02256788.

Verdonschot:1989:MDD

[VH89] P. F. M. Verdonschot and L. W. G. Higler. Macroinvertebrates in Dutch ditches: a typological characterization and the status of the Demmerik ditches. Aquatic Ecology, 23(2): 135–142, December 1989. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02256730.

Voutier:2008:DAF

[VH08] Jennifer L. Voutier and John Mark Hanson. Distribution, abundance, and feeding of a disjunct population of lady crab in the southern Gulf of St. Lawrence, Canada. Aquatic Ecology, 42(1):43–60, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9078-2.

Virro:2009:DSW

[VHB09] Taavi Virro, Juta Haberman, and Kätlin Blank. Diversity and structure of the winter rotifer assemblage in a shallow eutrophic northern temperate Lake Võrtsjärv. Aquatic Ecology, 43(3):755–764, September 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9276-1.

Verdonschot:1992:LMD

[VHC92] P. F. M. Verdonschot, L. W. G. Higler, and J. G. M. Cuppen. A list of macroinvertebrates in Dutch water types: a first step towards an ecological classification of surface waters based on key factors. Aquatic Ecology, 25(3):241–259, August 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02270810.

Vinni:2000:FGA

[VHN00] Mika Vinni, Jukka Horppila, and Kari Nyberg. The food, growth and abundance of five co-existing cyprinids in lake basins of different morphometry and water quality. Aquatic Ecology, 34(4):421–431, December 2000. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011404721775.

Vendrig:2017:RVS

[VHtB17] Nadia J. Vendrig, Lia Hemerik, and Cajo J. F. ter Braak. Response variable selection in principal response curves using permutation testing. Aquatic Ecology, 51(1):131–143, March 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-016-9604-1.

Vieira:2020:LCI

[Vie20] Christophe Vieira. Lobophora–coral interactions and phase shifts: summary of current knowledge and future directions. Aquatic Ecology, 54(1):1–20, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09723-2.

Visser:2016:AMC

[VIH16] Petra M. Visser, Bas W. Ibelings, and Jef Huisman. Artificial mixing to control cyanobacterial blooms: a review. Aquatic Ecology, 50(3):423–441, September 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9537-0.

Vijverberg:1991:VPA

[Vij91] J. Vijverberg. Variability and possible adaptive significance of day-time vertical distribution of Leptodora kindtii (Focke) (Cladocera) in a shallow eutrophic lake. Aquatic Ecology, 25 (1):85–91, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02259594.

Vilenica:2017:MEA

[VIM17] Marina Vilenica, Marija Ivković, and Zlatko Mihaljević. Mayfly emergence along an oligotrophic Dinaric karst hydrosystem: spatial and temporal patterns, and species– environment relationship. Aquatic Ecology, 51(3):417–433, September 2017. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-017-9626-3.

vanIersel:1992:RBN

[vIR92] P. B. W. van Iersel and W. J. Rip. Restoration of botshol (The Netherlands) by reduction of external load: Problem analyses and restoration methods. Aquatic Ecology, 25 (3):265–273, August 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02270812.

Visser:1988:BLL

[Vis88] P. Visser. Bythotrephes longimanus Leydig in Nederland. (Dutch) [Bythotrephes longimanus Leydig in The Netherlands]. Aquatic Ecology, 22(2):211, December 1988. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256827.

VanBreemen:1980:TSD

[VK80] A. N. Van Breemen and G. J. G. Kok. Thermal stratification and deep man-made lakes: an evaluation of a predictive model. Aquatic Ecology, 14(3):158–168, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260117.

Vosjan:1982:P

[VK82] J. H. Vosjan and Frans Kouwets. Publications. Aquatic Ecology, 16(2-3):292–293, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02255389.

VanDerVlugt:1990:RLF

[VK90] J. C. Van Der Vlugt and S. P. Klapwijk. The Reeuwijk lakes: a five years water quality study in an eutrophic ecosystem. Aquatic Ecology, 24(1):57–67, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02256749.

Volk:2012:CFA

[VK12] Carol Volk and Peter Kiffney. Comparison of fatty acids and elemental nutrients in periphyton, invertebrates, and cutthroat trout (Oncorhynchus clarki) in conifer and alder streams of western Washington state. Aquatic Ecology, 46 (1):85–99, March 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9383-7.

Vos-Kelk:1977:SNL

[VKD77] Pauline Vos-Kelk and C. Davids. A short note on Lecane species (Rotifera) found in Utricularia minor vegetation. Aquatic Ecology, 11(2):53–55, September 1977. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02265709.

VanBerkum:1995:BDF

[VKG95] J. A. Van Berkum, M. Klinge, and M. P. Grimm. Biomanipulation in the Duinigermeer; first results. Aquatic Ecology, 29(1):81–90, April 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02061791.

VanDerVeer:1973:EMI

[VKK73] J. Van Der Veer and M. Koeman-Kwak. Electron microscopical identification of minute algae. Aquatic Ecology, 7(1):36– 43, March 1973. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02279919.

Vetrova:2002:BCL

[VKK02] E. Vetrova, V. Kratasyuk, and N. Kudryasheva. Bioluminescence characteristics of Lake Shira water. Aquatic Ecology, 36(2):309–315, April 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1015638909971.

vanLiere:1986:CRE

[vL86] Louis van Liere. Concluding remarks, with emphasis on remarks. Aquatic Ecology, 20(1-2):257–259, November 1986. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291167.

VanGray:2018:IVD

[VL18] Jonathon B. Van Gray and Laura G. Leff. Impacts of varying durations of passive oxygen exposure on freshwater denitrifier community structure and function. Aquatic Ecology, 52(1):35–49, March 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9643-2.

vanLiere:2007:SCN

[vLJA07] Lowie van Liere, Jan H. Janse, and Gertie H. P. Arts. Setting critical nutrient values for ditches using the eutrophication model PCDitch. Aquatic Ecology, 41(3):443–449, September 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-2835-1.

VanLiere:1979:DOA

[VM79] Louis Van Liere and Luuc R. Mur. Decay of Oscillatoria agardhi Gomont. Aquatic Ecology, 13(1):56–60, April 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260309.

Vrankovic:2025:CAA

[VM25] Jelena S. Vranković and Vanja M. Marković. Comparative analysis of antioxidant enzyme activities in Lithoglyphus naticoides and Theodoxus fluviatilis: insights into redox status and environmental adaptation. Aquatic Ecology, 59(1):277–294, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-024-10162-x.

VanMaldegem:1993:CSB

[VML93] D. C. Van Maldegem, H. P. J. Mulder, and A. Langerak. A cohesive sediment balance for the Scheldt estuary. Aquatic Ecology, 27(2-4):247–256, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02334788.

VanDam:1994:CCE

[VMS94] Herman Van Dam, Adrienne Mertens, and Jos Sinkeldam. A coded checklist and ecological indicator values of freshwater diatoms from The Netherlands. Aquatic Ecology, 28 (1):117–133, March 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02334251.

Voros:2009:FYS

[VMS09] Lajos Vörös, Andrea Mózes, and Boglárka Somogyi. A fiveyear study of autotrophic winter picoplankton in Lake Balaton, Hungary. Aquatic Ecology, 43(3):727–734, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-009-9272-5.

Vineetha:2022:TBP

[VN22] S. Vineetha and S. Bijoy Nandan. A trait-based perspective toward the mechanistic understanding of benthic macroinvertebrate assemblage in a tropical rice field. Aquatic Ecology, 56(4):1069–1097, November 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09984-4.

Valente-Neto:2015:ERD

[VNKdOR15] Francisco Valente-Neto, Ricardo Koroiva, and Fabio de Oliveira Roque. The effect of riparian deforestation on macroinvertebrates associated with submerged woody debris. Aquatic Ecology, 49 (1):115–125, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9510-y.

Valente-Neto:2016:ESS

[VNSFG16] Francisco Valente-Neto, Victor S. Saito, and Alaide Aparecida Fonseca-Gessner. Evidence of species sorting driving aquatic beetles associated with woody debris in a transitional region between Cerrado and Atlantic Forest biomes. Aquatic Ecology, 50(2):209–220, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9569-0.

Oosterom:2019:RNR

[vOCRM19] M. V. López van Oosterom, J. P. Casas-Ruiz, and I. Muñoz. Responses of a native and a recent invader snail to warming and dry conditions: the case of the lower Ebro River. Aquatic Ecology, 53(3):497–508, September 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09704-5.

VanGils:1993:MWS

[VOD93] J. A. G. Van Gils, M. R. L. Ouboter, and N. M. De Rooij. Modelling of water and sediment quality in the Scheldt estuary. Aquatic Ecology, 27(2-4):257–265, June 1993. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334789.

Vosjan:1971:S

[Vos71] J. H. Vosjan. Sulfidevorming. Aquatic Ecology, 5(2):106, June 1971. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185953.

Vosjan:1982:RET

[Vos82] J. H. Vosjan. Respiratory electron transport system activities in marine environments. Aquatic Ecology, 16(1):61–68, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255413.

Vosjan:1987:SOF

[Vos87] J. H. Vosjan. A sketchy outline of the fate of organic matter in the Dutch Wadden Sea. Aquatic Ecology, 21(2):127–132, December 1987. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255437.

VanWijk:1977:PQS

[VP77] R. J. Van Wijk and P. J. G. Polderman. A permanent quadrat study of saltmarsh algal vegetation by means of mapping. Aquatic Ecology, 11(2):46–52, September 1977. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02265708.

Valladares:2012:NLD

[VP12] Sonia Valladares and Miquel Planas. Non-lethal dorsal fin sampling for stable isotope analysis in seahorses. Aquatic Ecology, 46(3):363–370, September 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9407-y.

Vishal:2021:SMB

[VPF21] C. R. Vishal, A. Parvathi, and Camey M. Furtado. In situ measurements of bioluminescence response of Gonyaulax spinifera to various mechanical stimuli. Aquatic Ecology, 55 (2):437–451, June 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-021-09836-7.

VanLiere:1982:RSP

[VPM82] Louis Van Liere, Josje Peters, and Luuc R. Mur. Release of sediment-phosphorus and the influence of algal growth on this process. Aquatic Ecology, 16(2-3):191–200, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255372.

VanDolah:2016:ESE

[VPP16] Elizabeth R. Van Dolah, Michael Paolisso, and Allen Place. Employing a socio-ecological systems approach to engage harmful algal bloom stakeholders. Aquatic Ecology, 50(3): 577–594, September 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9562-z.

Vayghan:2013:SIH

[VPS13] Ali Haghi Vayghan, Hadi Poorbagher, and Hasan Nasrollahzadeh Saravi. Suitability indices and habitat suitability index model of Caspian kutum (Rutilus frisii kutum) in the southern Caspian Sea. Aquatic Ecology, 47(4):441–451, December 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9457-9.

Vendramin:2021:CUZ

[VPS21] Daiane Vendramin, Allana Gonçalves Piu, and Cristina Stenert. Can the use of zooplankton dormant stages from natural wetlands contribute to restoration of mined wetlands? Aquatic Ecology, 55(2):681–693, June 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09854-5.

Vieira:2024:IDE

[VRH^{+}24] Bruna Horvath Vieira, Suzelei Rodgher, Renata Natsumi Haneda, Ana Teresa Lombardi, Maria da Graça Gama Melão, Michiel Adriaan Daam, and Evaldo Luiz Gaeta Espíndola. Importance of different exposure routes on the toxicity of chromium to planktonic organisms. Aquatic Ecology, 58(2):175–189, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
023-10054-6.

Vrijhof:1980:VFM

[Vri80] H. Vrijhof. In vivo fluorescence measurements of phytoplankton. Aquatic Ecology, 14(3):213–215, December 1980. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02260123.

Veldhuis:1994:R

[VRP94] Marcel Veldhuis, Piet Roos, and J. R. Parsons. Reviews. Aquatic Ecology, 28(2):221–232, June 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02333993.

Vitt:2017:EAU

[VRR17] Simon Vitt, Anna K. Rahn, and Ingolf P. Rick. Enhanced ambient UVB light affects growth, body condition and the investment in innate and adaptive immunity in three-spined sticklebacks (Gasterosteus aculeatus). Aquatic Ecology, 51 (4):499–509, December 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-017-9632-5.

VanDeBund:1996:BCE

[VS96a] Wouter J. Van De Bund and Stefan J. H. Spaas. Benthic communities of exposed littoral sand-flats in eighteen Dutch lakes. Aquatic Ecology, 30(1):15–20, July 1996. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02092143.

VanDerHammen:1996:WMA

[VS96b] H. Van Der Hammen and H. Smit. The water mites (Acari: Hydrachnidia) of streams in The Netherlands: Distribution and ecological aspects on a regional scale. Aquatic Ecology, 30(2-3):175–185, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02272237.

Vehmaa:2009:DPL

[VS09] Anu Vehmaa and Kalevi Salonen. Development of phytoplankton in Lake Pääjärvi (Finland) during under-ice convective mixing period. Aquatic Ecology, 43(3):693–705, September 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-009-9273-4.

VanDerGeest:1997:IDC

[VSA97] H. G. Van Der Geest, S. C. Stuijfzand, and W. Admiraal. Impact of a diazinon calamity in 1996 on the aquatic macroinvertebrates in the River Meuse, The Netherlands. Aquatic Ecology, 30(4):327–330, May 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02085875.

Vanschoenwinkel:2010:HPL

[VSB10] Bram Vanschoenwinkel, Maitland Seaman, and Luc Brendonck. Hatching phenology, life history and egg bank size of fairy shrimp Branchipodopsis spp. (Branchiopoda, Crustacea) in relation to the ephemerality of their rock pool habitat. Aquatic Ecology, 44(4):771–780, December 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-010-9315-y.

Vaz:2019:WTG

[VSM19a] Ana Catarina Vaz, Giuseppe Scarcella, and Filipe Martinho. Water temperature gradients drive early life-history patterns of the common sole (Solea solea L.) in the northeast Atlantic and Mediterranean. Aquatic Ecology, 53(2):281–294, June 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09688-2.

Vijapure:2019:MCP

[VSM19b] Tejal Vijapure, Soniya Sukumaran, and Cathrine S. Manohar. Molecular characterization and phylogenetics of Indian polychaete fauna: scope for implementation in ecological monitoring. Aquatic Ecology, 53(4):665–677, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09717-0.

Voelz:2000:LTM

[VSW00] Neal J. Voelz, Sen-Her Shieh, and J. V. Ward. Long-term monitoring of benthic macroinvertebrate community structure: a perspective from a Colorado river*. Aquatic Ecology, 34(3):261–278, September 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009989510721.

Villeneuve:2019:UMA

[VTE19] Andrew Robert Villeneuve, Ian Thornhill, and Jacqualyn Eales. Upstream migration and altitudinal distribution patterns of Nereina punctulata (Gastropoda: Neritidae) in Dominica, West Indies. Aquatic Ecology, 53(2):205–215, June 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09683-7.

VanDenBrink:1985:OSY

[VV85] F. W. B. Van Den Brink and G. Van Den Velde. Observations on the seasonal and yearly occurrence and the distribution of Atyaephyra desmaresti (Millet, 1831) (Crustacea, Decapoda, Natantia) in The Netherlands. Aquatic Ecology, 19(2):193–198, December 1985. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02270766.

VanBeelen:1990:KDC

[VV90] P. Van Beelen and F. Van Keulen. The kinetics of the degradation of chloroform and benzene in anaerobic sediment from the River Rhine. Aquatic Ecology, 24(1):13–21, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256745.

VanDijk:1991:PSM

[VV91] Gerda M. Van Dijk and Ellen Van Donk. Perspectives for submerged macrophytes in shallow lake restoration projects in The Netherlands. Aquatic Ecology, 24(2):125–131, April 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02260429.

VanBeelen:1993:MCR

[VV93a] P. Van Beelen and P. L. A. Van Vlaardingen. The mineralization of chloroform in river sediments. Aquatic Ecology, 27(1):51–58, July 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02336928.

VanZoest:1993:BSP

[VV93b] R. Van Zoest and G. T. M. Van Eck. Behaviour of selected PCBs, PAHs and γ-HCH in the Scheldt estuary, S.W. Netherlands. Aquatic Ecology, 27(2-4):301–308, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334793.

Vrede:2005:CTE

[VV05] Tobias Vrede and Katarina Vrede. Contrasting ‘top-down’ effects of crustacean zooplankton grazing on bacteria and phytoflagellates. Aquatic Ecology, 39(3):283–293, September 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-1326-8.

Vesela:2007:EBS

[VV07] Sarka Vesela and Jacobus Vijverberg. Effect of body size on toxicity of zinc in neonates of four differently sized Daphnia species. Aquatic Ecology, 41(1):67–73, March 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9050-6.

VanLiere:1986:PPV

[VVA86] Louis Van Liere, Leo Van Ballegooijen, and Tom Aldenberg. Primary production in the various parts of Loosdrecht Lakes. Aquatic Ecology, 20(1-2):77–85, November 1986. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02291152.

VanDerVelde:1982:IDN

[VVB82] G. Van Der Velde, L. A. Van Der Heijden, and P. M. M. Bexkens. Initial decomposition of Nymphoides peltata (Gmel.) O. Kuntze (Menyanthaceae), as studied by the leafmarking method. Aquatic Ecology, 16(1):51–60, June 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255412.

Verberk:2012:NST

[VvdMP12] Wilco C. E. P. Verberk, Piet J. J. van den Munckhof, and Bart J. A. Pollux. Niche segregation in two closely related species of stickleback along a physiological axis: explaining multidecadal changes in fish distribution from ironinduced respiratory impairment. Aquatic Ecology, 46(2):241– 248, June 2012. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-012-9395-y.

VanWichelen:2012:IMV

[VvGV12] Jeroen Van Wichelen, Ineke van Gremberghe, and Wim Vyverman. The importance of morphological versus chemical defences for the bloom-forming cyanobacterium Microcystis against amoebae grazing. Aquatic Ecology, 46(1):73– 84, March 2012. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-011-9382-8.

VanDenBrink:1996:BLR

[VVK96] F. W. B. Van Den Brink, G. Van Der Velde, and A. G. Klink. Biodiversity in the Lower Rhine and Meuse river-floodplains: Its significance for ecological river management. Aquatic Ecology, 30(2-3):129–149, October 1996. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02272234.

Voutilainen:2009:RBP

[VvOT09] Ari Voutilainen, Tapio van Ooik, and Jouni Taskinen. Relationship between prevalence of trematode parasite Diplostomum sp. and population density of its snail host Lymnaea stagnalis in lakes and ponds in Finland. Aquatic Ecology, 43(2):351–357, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9203-x.

VanDijk:1995:UXR

[VVW95] Gerda M. Van Dijk, Bob Van Zanten, and Frederick G. Wortelboer. Using XAD resins to study the effects of reduced organic micropollutant concentrations in River Rhine and Meuse water on phytoplankton growth. Aquatic Ecology, 29(2):151–156, June 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02336044.

Vaas:1978:LHE

[VW78] K. F. Vaas and W. J. Wolff. Large hydraulic enginering projects in the Netherlands and studies of their environmental impact. Aquatic Ecology, 12(3-4):176–179, December 1978. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02259181.

VanLiere:1975:GOA

[VZM75] Louis Van Liere, Wanda Zevenboom, and Luuc R. Mur. Growth of Oscillatoria agardhii Gom. Aquatic Ecology, 9 (2):62–70, September 1975. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02257638.

Wong:1984:CCF

[WA84] M. H. Wong and K. C. Au. Contents of Cd, Fe, Mn and Zn in the tissue of Katelysia hiantina collected from Tolo Harbour, an almost land-locked sea. Aquatic Ecology, 18(2):95–101, December 1984. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257048.

Waardenburg:1980:EIS

[Waa80] H. W. Waardenburg. Ecological investigations in sublittoral biocenoses on hard substrates. Aquatic Ecology, 14(3): 221–222, December 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260128.

Wanders:1979:ETE

[Wan79] J. B. W. Wanders. Effects of thermal effluents from the Bergum power station on the phytoplankton in the Bergumermeer. Aquatic Ecology, 13(2-3):98–99, December 1979. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02284738.

Warras:1997:BRU

[War97] Carl J. Warras. Book review: UV-B and Biosphere. Edited by J. Rozema, W. W. C. Gieskes, S. C. van de Gejin, C. Nolan and H. de Boois. Aquatic Ecology, 31(1):109, March 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009996024936.

Wotton:1992:CEM

[WAW92] R. S. Wotton, P. D. Armitage, and C. A. Woodward. Colonization and emergence of midges (Chironomidae: Diptera) in slow sand filter beds. Aquatic Ecology, 26(2-4):331–339, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255259.

Wright:2010:DVC

[WB10] Ian A. Wright and Shelley Burgin. Diel variation in chironomid (Diptera: Insecta) exuviae abundance and taxonomic richness in near-pristine upland streams of the Greater Blue Mountains World Heritage Area, South-Eastern Australia. Aquatic Ecology, 44(1):131–141, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9257-4.

Wilson:1993:HVG

[WBB93] J. G. Wilson, M. Brennan, and B. Brennan. Horizontal and vertical gradients in sediment nutrients on mudflats in the Shannon estuary, Ireland. Aquatic Ecology, 27(2-4):173–180, June 1993. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334780.

Wiltshire:2000:EPF

[WBB00] Karen H. Wiltshire, Maarten Boersma, and Heinke Buhtz. Extraction of pigments and fatty acids from the green alga Scenedesmus obliquus (Chlorophyceae). Aquatic Ecology, 34 (2):119–126, June 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009911418606.

Wan:2017:TSV

[WBR17] Yu Wan, Yang Bai, and Xiaohong Ruan. Temporal and spatial variations of aquatic environmental characteristics and sediment bacterial community in five regions of Lake Taihu. Aquatic Ecology, 51(3):343–358, September 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9621-8.

Westrelin:2022:ECE

[WBS22] Samuel Westrelin, Stéphanie Boulêtreau, and Frédéric Santoul. European catfish Silurus glanis behaviour in response to a strong summer hypoxic event in a shallow lake. Aquatic Ecology, 56(4):1127–1142, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09952-y. See correction [WBS24].

Westrelin:2024:CEC

[WBS24] Samuel Westrelin, Stéphanie Boulêtreau, and Frédéric Santoul. Correction: European catfish Silurus glanis behaviour in response to a strong summer hypoxic event in a shallow lake. Aquatic Ecology, 58(2):533, June 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10068-0. See [WBS22].

Wejnerowski:2017:STD

[WCD17] Lukasz Wejnerowski, Slawek Cerbin, and Marcin Krzysztof Dziuba. Setae thickening in Daphnia magna alleviates the food stress caused by the filamentous cyanobacteria. Aquatic Ecology, 51(3):485–498, September 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9631-6.

Wedekind:2000:RHS

[WCT00] Claus Wedekind, Mira Christen, and Nathalie Treichel. Relative helminth size in crustacean hosts: in vivo determination, and effects of host gender and within-host competition in a copepod infected by a cestode. Aquatic Ecology, 34(3): 279–285, September 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1009976420423.

Wang:2011:WDP

[WCZ11] Lan Wang, Qinghua Cai, and Min Zhang. Weekly dynamics of phytoplankton functional groups under high water level fluctuations in a subtropical reservoir-bay. Aquatic Ecology, 45(2):197–212, May 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-010-9346-4.

Winter:2000:SEE

[WD00] Jennifer G. Winter and Hamish C. Duthie. Stream epilithic, epipelic and epiphytic diatoms: habitat fidelity and use in biomonitoring. Aquatic Ecology, 34(4):345–353, December 2000. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1011461727835.

Wronski:2015:BAW

[WDA15] Torsten Wronski, Marie Claire Dusabe, and Christian Albrecht. Biological assessment of water quality and biodiversity in Rwandan rivers draining into Lake Kivu. Aquatic Ecology, 49(3):309–320, September 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9525-4.

Wilkerson:2015:NUP

[WDP15] Frances P. Wilkerson, Richard C. Dugdale, and Adam Pimenta. Nutrient uptake and primary productivity in an urban estuary: using rate measurements to evaluate phytoplankton response to different hydrological and nutrient conditions. Aquatic Ecology, 49(2):211–233, June 2015. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9516-5.

Weber:1971:AEB

[Web71a] R. E. Weber. De afdeling experimentele biologie. Aquatic Ecology, 5(2):105, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185952.

Weber:1971:IVE

[Web71b] R. E. Weber. Over de inwerking van enkele milieufactoren op de functie van haemoglobine bij wadpieren. (Dutch) [On the effect of some environmental factors on the function of hemoglobin in Wadden Sea worms]. Aquatic Ecology, 5 (2):109–111, June 1971. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF03185956.

Wellershaus:1974:SCZ

[Wel74] Stefan Wellershaus. Seasonal changes in the zooplankton population in the Cochin Backwater (a South Indian estuary). Aquatic Ecology, 8(1-2):213–223, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254921.

Wetzel:1981:C

[Wet81] Robert G. Wetzel. Congresses. Aquatic Ecology, 15(1-2):94–95, October 1981. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02260265.

Wetzel:1993:MML

[Wet93] Robert G. Wetzel. Microcommunities and microgradients: Linking nutrient regeneration, microbial mutualism, and high sustained aquatic primary production. Aquatic Ecology, 27(1):3–9, July 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/BF02336924.

Weyhenmeyer:2009:DWW

[Wey09] Gesa A. Weyhenmeyer. Do warmer winters change variability patterns of physical and chemical lake conditions in Sweden? Aquatic Ecology, 43(3):653–659, September 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9284-1.

Wilson:2005:PDBa

[WF05a] James G. Wilson and Niamh Forrest. Population dynamics, biomass and productivity of Limapontia depressa (Gastropoda, Opisthobranchia) at Bull Island, Dublin, Ireland. Aquatic Ecology, 38(4):575–585, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-0310-7.

Wilson:2005:PDBb

[WF05b] James G. Wilson and Niamh Forrest. Population dynamics, biomass and productivity of Limapontia depressa (Gastropoda, Opisthobranchia) at Bull Island, Dublin, Ireland. Aquatic Ecology, 38(4):575–585, January 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-0310-z.

Wojtal-Frankiewicz:2012:EGW

[WF12] Adrianna Wojtal-Frankiewicz. The effects of global warming on Daphnia spp. population dynamics: a review. Aquatic Ecology, 46(1):37–53, March 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
011-9380-x.

Wojtal-Frankiewicz:2017:REF

[WFBJ17] Adrianna Wojtal-Frankiewicz, Joanna Bernasińska, and Tomasz Jurczak. The role of environmental factors in the induction of oxidative stress in zebra mussel (Dreissena polymorpha). Aquatic Ecology, 51(2):289–306, June 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9617-4.

Wojtal-Frankiewicz:2010:CFP

[WFFM10] A. Wojtal-Frankiewicz, P. Frankiewicz, and T. K. Mc-Carthy. Comparison of fish and phantom midge influence on cladocerans diel vertical migration in a dual basin lake. Aquatic Ecology, 44(1):243–254, March 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9280-5.

Watson:1993:BFN

[WFH93] P. G. Watson, P. E. Frickers, and R. J. M. Howland. Benthic fluxes of nutrients and some trace metals in the Tamar estuary, SW England. Aquatic Ecology, 27(2-4):135–146, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02334776.

Wiltshire:1995:PCS

[WGK95] Karen Helen Wiltshire, Claus Geisler, and Doris Krogmann. Pigment characterisation of suspended matter in association with particulate heavy metal loadings in the German Bight and Elbe Estuary. Aquatic Ecology, 29(3-4):303–314, December 1995. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02084229.

Weinmann:2021:CRI

[WGL21] Anna E. Weinmann, Susan T. Goldstein, and Martin R. Langer. Community responses of intertidal foraminifera to pH variations: a culture experiment with propagules. Aquatic Ecology, 55(1):309–325, March 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09833-w.

Watson:1992:PSC

[WH92] Charles N. Watson, Jr. and Michael W. Heyn. A preliminary survey of the Chironomidae (Diptera) of Costa Rica, with emphasis on the lotic fauna. Aquatic Ecology, 26(2-4):257– 262, June 1992. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02255249.

Weand:1976:TED

[WHP76] Barron L. Weand, Robert C. Hoehn, and Bruce C. Parker. Trace element distributions in an Antarctic meromictic lake.

Aquatic Ecology, 10(2):104–114, September 1976. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02282218.

Wedderburn:2013:ZRF

[WHS13] Scotte D. Wedderburn, Karl A. Hillyard, and Russell J. Shiel. Zooplankton response to flooding of a drought refuge and implications for the endangered fish species Craterocephalus fluviatilis cohabiting with alien Gambusia holbrooki. Aquatic Ecology, 47(3):263–275, September 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-013-9442-3.

Worischka:2014:FPC

[WHW14] Susanne Worischka, Claudia Hellmann, and Carola Winkelmann. Fish predation can induce mesohabitat-specific differences in food web structures in small stream ecosystems. Aquatic Ecology, 48(4):367–378, December 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9490-3.

Willemsen:1970:GVD

[Wil70] J. Willemsen. De giftigheid van diuron voor vis. (Dutch) [The toxicity of diuron to fish]. Aquatic Ecology, 4(1):34– 37, March 1970. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF03185864.

Willemsen:1979:ICW

[Wil79] J. Willemsen. The influence of cooling water discharge upon fish. Aquatic Ecology, 13(2-3):94–95, December 1979. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02284735.

Willemsen:1980:FAE

[Wil80] J. Willemsen. Fishery-aspects of eutrophication. Aquatic Ecology, 14(1-2):12–21, August 1980. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02260268.

Williams:2007:SAP

[Wil07] D. Dudley Williams. Size-assortative pairing in the lotic amphipod Gammarus zaddachi, an examination of hypotheses and the influence of current speed. Aquatic Ecology, 41 (2):309–317, June 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-006-9075-x.

Weilhoefer:2022:NLP

[WiNDF22] Christine L. Weilhoefer, Shin ichi Nakano, Souyma Deb, and Keitaro Fukushima. Nutrient limitation of primary production in rivers along a land use gradient in the Lake Biwa Basin, Shiga Prefecture, Japan. Aquatic Ecology, 56(4): 1177–1203, December 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09971-9.

Wong:1982:HMU

[WKH82] M. H. Wong, T. T. Kwok, and K. C. Ho. Heavy metals in Ulva lactuca collected within Tolo Harbour, an almost landlocked sea. Aquatic Ecology, 16(2-3):223–230, December 1982. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02255376.

Wrobel:2024:TDR

[WKK24] Michał Wróbel and Anna Krysztofiak-Kaniewska. Time for dam rebuilding by the Eurasian beaver. Aquatic Ecology, 58(4):1157–1160, December 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
024-10130-5.

Wernberg:2005:MPE

[WKT05] Thomas Wernberg, Gary A. Kendrick, and Benjamin D. Toohey. Modification of the physical environment by an Ecklonia radiata (Laminariales) canopy and implications for associated foliose algae. Aquatic Ecology, 39(4):419–430, December 2005. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-005-9009-z.

Wang:2024:EFG

[WL24] Xiuli Wang and Changyi Lu. Exploring the fast-growing mechanism of Laguncularia racemosa from the perspective of leaf traits and ultrastructure. Aquatic Ecology, 58(2):387– 398, June 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-023-10077-z.

Wang:2022:CIP

[WLB22] Xing Wang, Qiuhua Li, and Anton Brancelj. Composition and indication of plankton fatty acids under the influence of environmental factors in the Hongfeng Reservoir, Southwest China. Aquatic Ecology, 56(3):653–666, September 2022. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09942-0.

Wild:2012:EWC

[WLN12] Christian Wild, Christian Laforsch, and Wolfgang Niggl. Effect of water currents on organic matter release by two scleractinian corals. Aquatic Ecology, 46(3):335–341, September 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-012-9404-1.

Wondie:2007:SVP

[WMD07] Ayalew Wondie, Seyoum Mengistu, and Eshete Dejen. Seasonal variation in primary production of a large high altitude tropical lake (Lake Tana, Ethiopia): effects of nutrient availability and water transparency. Aquatic Ecology, 41(2):195– 207, June 2007. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9080-8.

Williams:1994:IRA

[WMM94] M. R. Williams, G. E. Millward, and A. W. Morris. ^{206/207}Pb isotopic ratios applied to particulate transport processes — North eastern Irish Sea. Aquatic Ecology, 28(3-4):353–358, September 1994. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02334204.

Wintzer:2011:EIP

[WMM11] Alpa P. Wintzer, Mariah H. Meek, and Bernie May. Ecological insights into the polyp stage of non-native hydrozoans in the San Francisco Estuary. Aquatic Ecology, 45(2):151– 161, May 2011. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-010-9343-7.

Witthoft-Muhlmann:2007:CIR

[WMTR07] A. Witthöft-Mühlmann, W. Traunspurger, and K. O. Rothhaupt. Combined influence of river discharge and wind on littoral nematode communities of a river mouth area of Lake Constance. Aquatic Ecology, 41(2):231–242, June 2007. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9076-9.

Williams:2009:FAS

[WMV09] Christopher D. Williams, James Moran, and Jean-Claude Vala. Factors affecting Sciomyzidae (Diptera) across a transect at Skealoghan Turlough (Co. Mayo, Ireland). Aquatic Ecology, 43(1):117–133, March 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
007-9149-4.

Wallenstein:2008:ABB

[WNS08] Francisco M. Wallenstein, Ana I. Neto, and Catarina I. Santos. Algae-based biotopes of the Azores (Portugal): spatial and seasonal variation. Aquatic Ecology, 42(4):547–559, December 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9134-y.

Wang:2024:EIH

[WNZ24] Bixue Wang, Jing Ning, and Guoping Zhu. Ectoparasite infestation and host–parasite trophic relationship for Champsocephalus gunnari (Lonnberg, 1905) at South Orkney Islands, Antarctica. Aquatic Ecology, 58(2):335–348, June 2024. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10072-4.

Wolff:1974:BDR

[Wol74] W. J. Wolff. Benthic diversity in the Rhine–Meuse estuary. Aquatic Ecology, 8(1-2):242–252, July 1974. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02254924.

Wolff:1979:TPO

[Wol79] W. J. Wolff. Transport of particulate organic material between salt marsh and estuary. Aquatic Ecology, 13(1):68, April 1979. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260311.

Wolff:1997:DAE

[Wol97] Wim J. Wolff. The development of aquatic ecology in the Netherlands. Aquatic Ecology, 30(4):241–253, May 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02085869.

Wolff:2000:RHI

[Wol00] Wim J. Wolff. Recent human-induced invasions of fresh waters by saltwater animals? Aquatic Ecology, 34(3):319–321, September 2000. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1023/A:1009908010959.

Wortelboer:1990:MCB

[Wor90] F. G. Wortelboer. A model on the competition between two macrophyte species in acidifying shallow soft-water lakes in the Netherlands. Aquatic Ecology, 24(1):91–107, July 1990. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02256751.

Wetsteyn:1991:FOO

[WP91] L. P. M. J. Wetsteyn and L. Peperzak. Field observations in the Oosterschelde (The Netherlands) on Coscinodiscus concinnus and Coscinodiscus granii (Bacillariophyceae) infected by the marine fungus Lagenisma coscinodisci (Oomycetes). Aquatic Ecology, 25(1):15–21, October 1991. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/BF02259584.

Wellnitz:2012:CMP

[WP12] Todd Wellnitz and N. LeRoy Poff. Current-mediated periphytic structure modifies grazer interactions and algal removal. Aquatic Ecology, 46(4):521–530, December 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9419-7.

Winkelmann:2008:BFR

[WPB08] Carola Winkelmann, Thomas Petzoldt, and Jürgen Benndorf. Benthivorous fish reduce stream invertebrate drift in a large-scale field experiment. Aquatic Ecology, 42(3):483–493, September 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9101-7.

Werdel:2021:CBE

[WPB21] Gina M. Werdel, Lalit K. Pandey, and Elizabeth A. Bergey. Cigarette butt effects on diatom health in a stream ecosystem. Aquatic Ecology, 55(3):999–1010, September 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09876-z.

Walter:2017:ECD

[WPvB17] Bettina Walter, Janna Peters, and Justus E. E. van Beusekom. The effect of constant darkness and short light periods on the survival and physiological fitness of two phytoplankton species and their growth potential after reillumination. Aquatic Ecology, 51(4):591–603, December 2017. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-017-9638-z.

Wellmann:1998:PSE

[WRH98] Patrick Wellmann, Hans-Toni Ratte, and Fred Heimbach. Primary and secondary effects of methabenzthiazuron on plankton communities in aquatic outdoor microcosms. Aquatic Ecology, 32(2):125–134, September 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic).

URL https://link.springer.com/article/10.1023/A:
1009992431458.

Wilson:2005:SIA

[WRS05] A. L. Wilson, D. S. Ryder, and M. M. Stevens. Stable isotope analysis of aquatic invertebrate communities in irrigated rice fields cultivated under different management regimes. Aquatic Ecology, 39(2):189–200, June 2005. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-004-7085-0.

Wiltshire:1994:PPS

[WS94] K. H. Wiltshire and F. Schroeder. Pigment patterns in suspended matter from the Elbe estuary, Northern Germany. Aquatic Ecology, 28(3-4):255–265, September 1994. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334193.

Wernand:1998:ESS

[WSvA98] M. R. Wernand, S. J. Shimwell, and H. M. van Aken. Evaluation of specific semi-empirical coastal colour algorithms using historic data sets. Aquatic Ecology, 32(1):73–91, March 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009946501534.

Weber:2014:TCM

[WT14] Sebastian Weber and Walter Traunspurger. Top-down control of a meiobenthic community by two juvenile freshwater fish species. Aquatic Ecology, 48(4):465–480, December 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-014-9498-8.

Wu:2009:SDB

[WTC09] Naicheng Wu, Tao Tang, and Qinghua Cai. Spatial distribution of benthic algae in the Gangqu River, Shangrila, China. Aquatic Ecology, 43(1):37–49, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9153-8.

Wilkinson:1995:GVD

[WTG95] Martin Wilkinson, Trevor Telfer, and Sarah Grundy. Geographical variation in the distributions of macroalgae in estuaries. Aquatic Ecology, 29(3-4):359–368, December 1995. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02084235.

Weber:2002:OSU

[WV02] Anke Weber and Sarka Vesela. Optimising survival under predation: chemical cues modify curvature in Daphnia galeata. Aquatic Ecology, 36(4):519–527, December 2002. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1021123519857.

Whatley:2014:REV

[WvLA14] Merrin H. Whatley, E. Emiel van Loon, and Wim Admiraal. The role of emergent vegetation in structuring aquatic insect communities in peatland drainage ditches. Aquatic Ecology, 48(3):267–283, September 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9482-3.

Wolters:2018:SIM

[WVM18] Jan-Willem Wolters, Ralf C. M. Verdonschot, and Patrick Meire. Stable isotope measurements confirm consumption of submerged macrophytes by macroinvertebrate and fish taxa. Aquatic Ecology, 52(4):269–280, December 2018. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-018-9662-7.

Wijnhoven:2003:EIF

[WvRvdV03a] S. Wijnhoven, M. C. van Riel, and G. van der Velde. Exotic and indigenous freshwater gammarid species: physiological tolerance to water temperature in relation to ionic content of the water. Aquatic Ecology, 37(2):151–158, April 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/

10.1023/A:1023982200529. See comments [TKM03] and reply [WvRvdV03b].

Wijnhoven:2003:RCW

[WvRvdV03b] S. Wijnhoven, M. C. van Riel, and G. van der Velde. Reply to comments on Wijnhoven et al. (2003). Aquatic Ecology, 37(4):449–451, October 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1023/B:AECO.

0000007074.31852.40. See [WvRvdV03a, TKM03].

Wilson:1984:SDC

[WW84] R. S. Wilson and S. E. Wilson. A survey of the distribution of Chironomidae (Diptera, Insecta) of the River Rhine by sampling pupal exuviae. Aquatic Ecology, 18(2):119–132, December 1984. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02257051.

Wang:2025:AHM

[WWG^{+}25] Sai Wang, Jing Wen, Tuan-Qi Guo, Hong-Jin Zhang, Yong-Duo Song, Guo-Ping Fu, and Tuan-Tuan Wang. Assessment of heavy metal ecological risk in typical seagrass beds of the China Yellow Sea. Aquatic Ecology, 59(1):83–94, March 2025. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-024-10149-8.

Wedderburn:2020:EAR

[WWS20] Scotte D. Wedderburn, Nick S. Whiterod, and Russell J. Shiel. Ecological aspects related to reintroductions to avert the extirpation of a freshwater fish from a large floodplain river. Aquatic Ecology, 54(1):281–294, March 2020. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09742-z.

Wallace:2021:MAB

[WWS21] Robert Lee Wallace, Elizabeth J. Walsh, and S. S. S. Sarma. A meta-analysis of benthic rotifer community structure as a function of lake trophic state. Aquatic Ecology, 55(4): 1297–1304, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-020-09825-2.

Wilk-Wozniak:2006:PRC

[WWŻ06] E. Wilk-Woźniak and R. Żurek. Phytoplankton and its relationships with chemical parameters and zooplankton in meromictic Piaseczno reservoir, Southern Poland. Aquatic Ecology, 40(2):165–176, June 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
005-0781-6.

Wang:2016:RAD

[WXL16] Yuyu Wang, Xiewen Xiao, and Guangchun Lei. Resource availability determines food chain length in Chinese subtropical rivers. Aquatic Ecology, 50(2):187–195, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9567-2.

Wang:2011:AES

[WXW11] Yuankun Wang, Ziqiang Xia, and Dong Wang. Assessing the effect of Separation Levee Project on Chinese sturgeon (Acipensor sinensis) spawning habitat suitability in Yangtze River, China. Aquatic Ecology, 45(2):255–266, May 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9351-2.

Wang:2012:IFA

[WXZ12] Xuefang Wang, Liuxiong Xu, and Jiangfeng Zhu. Impacts of fish aggregation devices on size structures of skipjack tuna Katsuwonus pelamis. Aquatic Ecology, 46(3):343–352, September 2012. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-012-9405-0.

Wu:2024:ILP

[WYW^{+}24] Shijie Wu, Kaiqiang Yu, Fuwen Wang, Tiezhu Mi, and Yu Zhen. Impact of low pH/high CO_{2} on the physiological response and exopolysaccharide content in cyanobacteria Trichodesmium erythraeum. Aquatic Ecology, 58(4):1225–1236, December 2024. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-024-10137-y.

Wang:2013:GTF

[WYZ13] Lei Wang, Tewu Yang, and Congming Zhu. Growth and turion formation of Potamogeton crispus in response to different phosphorus concentrations in water. Aquatic Ecology, 47(1):87–97, March 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-012-9427-7.

Wang:1992:CCR

[WZ92] X. Wang and L. Zheng. Checklist of Chironomidae records from China. Aquatic Ecology, 26(2-4):247–255, June 1992. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255248.

Wang:2018:DEF

[WZL18] Xiuli Wang, Liang Zhou, and Changyi Lu. Do environmental factors affect the male frequency of exotic mangrove species Laguncularia racemosa (Combretaceae) along the southeast coast of China? Aquatic Ecology, 52(2-3):235–244, September 2018. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-018-9658-3.

Wu:2019:CCC

[WZL19] Qiang Wu, Yapeng Zhang, and Pengfu Li. Comparison of community composition between Microcystis colonyattached and free-living bacteria, and among bacteria attached with Microcystis colonies of various sizes in culture. Aquatic Ecology, 53(3):465–481, September 2019. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09702-7.

Wang:2023:RPC

[WZM^{+}23] Liya Wang, Min Zhang, Yangyang Meng, Zhen Yang, Xiaoli Shi, Yang Yu, and Limei Shi. Responses of phytoplankton community dynamics to reduced underwater light in spring. Aquatic Ecology, 57(3):797–812, September 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10050-w.

Xu:2010:DVM

[XCL10] Yaoyang Xu, Qinghua Cai, and Daofeng Li. Diel vertical migration of Peridiniopsis niei, Liu et al., a new species of dinoflagellates in an eutrophic bay of Three-Gorge Reservoir, China. Aquatic Ecology, 44(2):387–395, June 2010. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-009-9298-8.

Huang:2023:DRR

[xHCjX^{+}23] Hui xiong Huang, Yun Cao, Kai jing Xin, Rong hua Liang, Yi ting Chen, and Jia jun Qi. Dynamic responses of root vigor, lipid peroxidation and antioxidant enzymes in Artemisia selengensis to long-term drought and rewatering. Aquatic Ecology, 57(2):321–335, June 2023. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-023-10012-2.

Xu:2015:RLE

[XWC15] Runbing Xu, Feng Wu, and Xuexiu Chang. Recovery limitation of endangered Ottelia acuminata by allelopathic interaction with cyanobacteria. Aquatic Ecology, 49(3):333–342, September 2015. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-015-9528-1.

Xu:2022:CAL

[XWJ22] Chao Xu, Hai-Jun Wang, and Erik Jeppesen. Can artificial light promote submerged macrophyte growth in summer? Aquatic Ecology, 56(1):89–98, March 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09899-6.

Xu:2011:PES

[XYH11] Jun Xu, Qiang Yang, and Lars-Anders Hansson. Preservation effects on stable isotope ratios and consequences for the reconstruction of energetic pathways. Aquatic Ecology, 45(4): 483–492, November 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-011-9369-5.

Yagci:2021:EIT

[YAC21] Meral Apaydın Yağcı, Ahmet Alp, and Mehmet Cesur. The effects and interactions of three invasive fish species introduced to the aquatic ecosystem of a Turkish lake (Eğirdir Lake). Aquatic Ecology, 55(4):1273–1283, December 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09842-9.

Youens:2008:CAO

[YB08] Abigail K. Youens and Romi L. Burks. Comparing applesnails with oranges: the need to standardize measuring techniques when studying Pomacea. Aquatic Ecology, 42(4): 679–684, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-007-9140-0.

Yuan:2020:TSV

[YBZ20] Jie Yuan, Hongmei Bu, and Quanfa Zhang. Temporal–spatial variations and source identification of dissolved nitrate in the upper Han River basin, China. Aquatic Ecology, 54 (1):89–101, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09728-x.

Yu:2018:DPT

[YF18] Haihao Yu and Shufeng Fan. Differences in physiological traits and resistances of Alternanthera philoxeroides after herbivory by generalists and specialists. Aquatic Ecology, 52 (4):323–332, December 2018. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-018-9666-3.

Yeager:2005:PPB

[YHM05] Christina L. J. Yeager, Lawrence W. Harding, Jr., and Michael E. Mallonee. Phytoplankton production, biomass and community structure following a summer nutrient pulse in Chesapeake Bay. Aquatic Ecology, 39(2):135–149, June 2005. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-004-4767-6.

Yoon:2012:PPM

[YKJ12] Ju-Duk Yoon, Jeong-Hui Kim, and Min-Ho Jang. Postpassage movement of the fluvial fish Zacco temminckii following upstream transportation by a fishway operation in dam. Aquatic Ecology, 46(4):421–430, December 2012. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9411-2.

Yoon:2015:EMI

[YKJ15] Ju-Duk Yoon, Jeong-Hui Kim, and Min-Ho Jang. Efficiency of a modified ice harbor-type fishway for Korean freshwater fishes passing a weir in South Korea. Aquatic Ecology, 49(4): 417–429, December 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9534-3.

Yoshida:2001:SSZ

[YKU01] Takehito Yoshida, Maiko Kagami, and Jotaro Urabe. Seasonal succession of zooplankton in the north basin of Lake Biwa. Aquatic Ecology, 35(1):19–29, March 2001. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1011498202050.

Yan:2008:TDP

[YL08] Yunjun Yan and Xiaoyu Li. Temporal dynamics of production and ingestion of the dominant mayflies in a subtropical stream in China. Aquatic Ecology, 42(4):657–667, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-007-9130-2.

Yeager:2011:EFT

[YL11] Lauren A. Yeager and Craig A. Layman. Energy flow to two abundant consumers in a subtropical oyster reef food web. Aquatic Ecology, 45(2):267–277, May 2011. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-011-9352-1.

Ysebaert:1993:BMA

[YMB93] T. Ysebaert, P. Meire, and J. Buijs. The benthic macrofauna along the estuarine gradient of the Schelde estuary. Aquatic Ecology, 27(2-4):327–341, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https:

//link.springer.com/article/10.1007/BF02334796.

Ysebaert:1998:ZIM

[YME98] T. Ysebaert, P. Meire, and K. Essink. Zonation of intertidal macrobenthos in the estuaries of Schelde and Ems. Aquatic Ecology, 32(1):53–71, March 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009912103505.

Yin:2008:EPS

[YN08a] Xu Wang Yin and Cui Juan Niu. Effect of pH on survival, reproduction, egg viability and growth rate of five closely related rotifer species. Aquatic Ecology, 42(4):607–616, December 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9136-9.

Yin:2008:PSF

[YN08b] Xu Wang Yin and Cui Juan Niu. Polymorphism in stem females and successive parthenogenetic generations in Brachionus calyciflorus Pallas. Aquatic Ecology, 42(3):415–420, September 2008. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9107-1.

Yamochi:2002:ARS

[YO02] Susumu Yamochi and Kazuki Oda. An attempt to restore suitable conditions for demersal fishes and crustaceans in the Port of Sakai-Semboku, north Osaka Bay, Japan. Aquatic Ecology, 36(1):67–83, January 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1013350420719.

Yang:2017:FAB

[YP17] Dongwoo Yang and Sangkyu Park. Freshwater anostracan, Branchinella kugenumaensis, as a potential controlling consumer species on toxic cyanobacteria Microcystis aeruginosa. Aquatic Ecology, 51(3):449–461, September 2017. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-017-9628-1.

Yurchenko:2009:UCS

[YRR09] Olga V. Yurchenko, Vasily I. Radashevsky, and Arkadiy A. Reunov. Ultrastructural comparison of the spermatozoa of the Pacific oyster Crassostrea gigas inhabiting polluted and relatively clean areas in Taiwan. Aquatic Ecology, 43(2):513– 519, June 2009. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-007-9161-8.

Yeager:1998:MDA

[YS98] P. E. Yeager and R. L. Sinsabaugh. Microbial diversity along a sediment detrital particle size gradient. Aquatic Ecology, 32(4):281–289, December 1998. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1023/A:1009944328443.

Yamamoto:2022:CPG

[YSY^{+}22] Takeo Yamamoto, Kay Sakuma, Akane Yoshikawa, Yosuke Igeta, and Katsuyuki Hamasaki. Changes in phototaxis, geotaxis, and sinking velocity in relation to larval development of snow crab (Chionoecetes opilio) and red snow crab (Chionoecetes japonicus) under laboratory conditions. Aquatic Ecology, 56(4):1099–1111, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09979-1.

Yang:2023:TSV

[YSYZ23] Zhen Yang, Xiaoli Shi, Yang Yu, and Min Zhang. The temporal and spatial variation in morphospecies, size and viability of Microcystis colonies in Lake Taihu. Aquatic Ecology, 57(1):53–63, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-022-09992-4.

Yemelyanova:2002:DGL

[YTD02] Anna Yu Yemelyanova, Tatiana A. Temerova, and Andrei G. Degermendzhy. Distribution of Gammarus lacustris Sars (Amphipoda, Gammaridae) in Lake Shira (Khakasia, Siberia) and laboratory study of its growth characteristics. Aquatic Ecology, 36(2):245–256, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1015624205389.

Yin:2020:JEP

[YYR20] Xuwang Yin, Haoyu Yin, and Yunjie Ruan. Joint effects of predation risk and food nutrient on sexual and asexual reproductions, and morphological defenses of freshwater rotifer Brachionus calyciflorus. Aquatic Ecology, 54(1):35–44, March 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-019-09724-1.

Yang:2023:ERC

[YYS^{+}23] Zhen Yang, Yanhong Yao, Meng Sun, Gu Li, and Jianqiang Zhu. Effects of rice–crayfish co-culture on ammonia-oxidizing microbial abundance and community structure. Aquatic Ecology, 57(1):21–33, March 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
022-09989-z.

Yin:2008:SLH

[YZ08] Xu Wang Yin and Wen Zhao. Studies on life history characteristics of Brachionus plicatilis O. F. Müller (Rotifera) in relation to temperature, salinity and food algae. Aquatic Ecology, 42(1):165–176, March 2008. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9092-4.

Yuan:2009:PCB

[YZY09] Lina Yuan, Wei Zhu, and Liuyan Yang. Phosphorus cycling between the colonial cyanobacterium Microcystis aeruginosa and attached bacteria, Pseudomonas. Aquatic Ecology, 43 (4):859–866, December 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9227-2.

Zadereev:2003:MEC

[Zad03] Yegor S. Zadereev. Maternal effects, conspecific chemical cues, and switching from parthenogenesis to gametogenesis in the cladoceran Moina macrocopa. Aquatic Ecology, 37(3):251–255, July 2003. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1023/A:1025850417717.

Zganec:2024:ISB

[ŽAL24] Krešimir Žganec, C. Brannon Andersen, and Jasna Lajtner. The impact of salinization on benthic macrocrustacean assemblages in a Mediterranean shallow lake. Aquatic Ecology, 58(3):689–706, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10099-1.

Zawierucha:2019:WBD

[ZBA19] Krzysztof Zawierucha, Jakub Buda, and Roberto Ambrosini. Water bears dominated cryoconite hole ecosystems: densities, habitat preferences and physiological adaptations of Tardigrada on an alpine glacier. Aquatic Ecology, 53(4): 543–556, December 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-019-09707-2.

Zawierucha:2019:FSS

[ZBB19] Krzysztof Zawierucha, Jakub Buda, and Michał Bogdziewicz. Fine-scale spatial heterogeneity of invertebrates within cryoconite holes. Aquatic Ecology, 53(2):179–190, June 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-019-09681-9.

Zamora-Camacho:2020:LNA

[ZCA20] Francisco Javier Zamora-Camacho and Pedro Aragón. Larval newts adjust foraging rate to perceived predator and competitor proximity. Aquatic Ecology, 54(1):271–280, March 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-019-09741-0.

Zhao:2014:CBT

[ZCZ14] Jing Zhao, Jie Cao, and Xijie Zhou. A comparison between two GAM models in quantifying relationships of environmental variables with fish richness and diversity indices. Aquatic Ecology, 48(3):297–312, September 2014. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-014-9484-1.

Zurek:2006:AES

[ŻD06] Roman Żurek and Niels De Pauw. Aquatic ecology: Special issue ‘ecology of a flooded opencast sulphur mine’. Aquatic Ecology, 40(2):133–134, June 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
006-9038-2.

Zhou:2021:CAS

[ZD21] Libin Zhou and Steven A. J. Declerck. A critical assessment of the stoichiometric knife-edge: no evidence for artifacts caused by the experimental P-supplementation of algae. Aquatic Ecology, 55(4):1317–1325, December 2021. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09867-0.

Zdorovennova:2009:STV

[Zdo09] G. E. Zdorovennova. Spatial and temporal variations of the water–sediment thermal structure in shallow ice-covered Lake Vendyurskoe (Northwestern Russia). Aquatic Ecology, 43(3):629–639, September 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
009-9277-0.

zuErmgassen:2013:QHC

[zEGB13] P. S. E. zu Ermgassen, M. W. Gray, and R. D. Brumbaugh. Quantifying the historic contribution of Olympia oysters to filtration in Pacific Coast (USA) estuaries and the implications for restoration objectives. Aquatic Ecology, 47(2):149– 161, June 2013. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-013-9431-6.

Zettler:1997:BGM

[Zet97a] Michael L. Zettler. Bibliography on the genus Marenzelleria and its geographical distribution, principal topics and nomenclature. Aquatic Ecology, 31(2):233–258, June 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1009967807978.

Zettler:1997:PDG

[Zet97b] Michael L. Zettler. Population dynamics, growth and production of the neozoon Marenzelleria cf. wireni (Verrill, 1873) (Polychaeta: Spionidae) in a coastal water of the southern Baltic Sea. Aquatic Ecology, 31(2):177–186, June 1997. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1023/A:1009903521182.

Zevenboom:1982:FCW

[Zev82] Wanda Zevenboom. N_{2}-fixing cyanobacteria: Why they do not become dominant in shallow hypertrophic lakes. Aquatic Ecology, 16(2-3):289–290, December 1982. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02255386.

Zadereev:2002:ECR

[ZG02] Yegor S. Zadereev and Michael V. Gubanov. The effect of chemicals released by Gammarus lacustris on the depth distribution of Arctodiaptomus salinus in laboratory conditions. Aquatic Ecology, 36(2):257–260, April 2002. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1023/A:1015635903281.

Zganec:2009:RBD

[ŽG09] Krešimir Žganec and Sanja Gottstein. The river before damming: distribution and ecological notes on the endemic species Echinogammarus cari (Amphipoda: Gammaridae) in the Dobra River and its tributaries, Croatia. Aquatic Ecology, 43(1):105–115, March 2009. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9157-4.

Zhang:2020:PCC

[ZGZ20] Weizhen Zhang, Peng Gu, and Zheng Zheng. Phenotype changes of cyanobacterial and microbial distribution characteristics of surface sediments in different periods of cyanobacterial blooms in Taihu Lake. Aquatic Ecology, 54(2):591– 607, June 2020. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-020-09761-1.

Zimmerman:1978:TFB

[Zim78] J. T. F. Zimmerman. Tidal flushing of the Balgzand tidal flat area, Dutch Wadden Sea. Aquatic Ecology, 12(2):151– 152, July 1978. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02260718.

Zingel:2023:CNC

[ZJN^{+}23] Priit Zingel, Erik Jeppesen, Tiina Nõges, Josef Hejzlar, Ülkü Nihan Tavşanoğlu, Eva Papastergiadou, Ulrike Scharfenberger, and Helen Agasild. Changes in nutrient concentration and water level affect the microbial loop: a 6-month mesocosm experiment. Aquatic Ecology, 57(2):369–381, June 2023. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-023-10015-z.

Zalizniak:2009:EDI

[ZKN09a] Liliana Zalizniak, Ben J. Kefford, and Dayanthi Nugegoda. Effects of different ionic compositions on survival and growth of Physa acuta. Aquatic Ecology, 43(1):145–156, March 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9144-9.

Zalizniak:2009:EPS

[ZKN09b] Liliana Zalizniak, Ben J. Kefford, and Dayanthi Nugegoda. Effects of pH on salinity tolerance of selected freshwater invertebrates. Aquatic Ecology, 43(1):135–144, March 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-007-9148-5.

Zbikowski:2010:NLS

[ŻKŻ10] Janusz Żbikowski, Jarosław Kobak, and Elżbieta Żbikowska. Is Nuphar lutea (L.) sm. a structuring factor for macrozoobenthos and selected abiotic parameters of water and bottom sediments throughout the year? Aquatic Ecology, 44(4): 709–721, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9309-9.

Zadereev:2007:IDM

[ZL07] Egor Zadereev and Tatiana Lopatina. The induction of diapause in Moina by species-specific chemical cues. Aquatic Ecology, 41(2):255–261, June 2007. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
006-9062-2.

Zheng:2010:GSF

[ZL10] Rong-Quan Zheng and Chun-Tao Liu. Giant spiny-frog (Paa spinosa) from different populations differ in thermal preference but not in thermal tolerance. Aquatic Ecology, 44(4): 723–729, December 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-009-9310-3.

Zhang:2012:MDE

[ZLC12] Yuying Zhang, Yunkai Li, and Yong Chen. Modeling the dynamics of ecosystem for the American lobster in the Gulf of Maine. Aquatic Ecology, 46(4):451–464, December 2012. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9414-z.

Zadereev:2022:ESG

[ZLT22] E. S. Zadereev, T. S. Lopatina, and A. P. Tolomeev. The effect of salinity on the grazing rate and survival of Daphnia magna females adapted to different salinities. Aquatic Ecology, 56(3):639–652, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09941-7.

Zhang:2013:EDN

[ZNP13] Jianying Y. Zhang, Wanmin M. Ni, and Yangdong D. Pan. Effects of different nitrogen species on sensitivity and photosynthetic stress of three common freshwater diatoms. Aquatic Ecology, 47(1):25–35, March 2013. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-012-9422-z.

Zoeteman:1972:OVG

[Zoe72] Ir. B. C. J. Zoeteman. Opslag van gezuiverd Rijnwater in het Veulume-Massief. (Dutch) [storage of purified Rhine water in the Veulume Massif]. Aquatic Ecology, 6(3):125–140, September 1972. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/BF02314881.

Zambrano:2006:RBF

[ZPD06] Luis Zambrano, Martin R. Perrow, and Thomas A. Davidson. Relationships between fish feeding guild and trophic structure in English lowland shallow lakes subject to anthropogenic influence: implications for lake restoration. Aquatic Ecology, 40(3):391–405, September 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-006-9037-3.

Zhu:2013:PCS

[ZPW13] Weiju Zhu, Yangdong Pan, and Quanxi Wang. Phytoplankton community and succession in a newly man-made shallow lake, Shanghai, China. Aquatic Ecology, 47(2):137–147, June 2013. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-013-9430-7.

Zhu:2015:PAN

[ZPW15] Weiju Zhu, Yangdong Pan, and Quanxi Wang. Phytoplankton assemblages in a newly man-made shallow lake and surrounding canals, Shanghai, China. Aquatic Ecology, 49 (2):147–157, June 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-015-9511-x.

Zhou:2016:EMP

[ZQH16] Jian Zhou, Boqiang Qin, and Xiaoxia Han. Effects of the magnitude and persistence of turbulence on phytoplankton in Lake Taihu during a summer cyanobacterial bloom. Aquatic Ecology, 50(2):197–208, June 2016. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-016-9568-1.

Zivic:2015:FMC

[ŽRP15] Ivana Živić, Tijana Radosavljević, and Andjeljko Petrović. The first molecular characterization of the genus Hirudo on the territory of Serbia: estimation of endangerment. Aquatic Ecology, 49(1):81–90, March 2015. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
015-9506-7.

Zivic:2016:PSM

[ŽRP16] Ivana Živić, Tijana Radosavljević, and Andjeljko Petrović. Phylogeography of the southern medicinal leech, Hirudo verbana: a reply to Utevsky and Trontelj (2015). Aquatic Ecology, 50(1):101–102, March 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
015-9560-1.

Zurburg:1994:SDB

[ZSD94] Wouter Zurburg, Aad Smaal, and Norbert Dankers. Seston dynamics and bivalve feeding in the Bay of Marennes–Oléron (France). Aquatic Ecology, 28(3-4):459–466, September 1994. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/BF02334217.

Zarkami:2014:MOR

[ZSG14] R. Zarkami, R. Sadeghi, and P. Goethals. Modelling occurrence of roach “ Rutilus rutilus ” in streams. Aquatic Ecology, 48(2):161–177, June 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-014-9473-4.

Zadereev:2010:VDA

[ZTG10] Egor S. Zadereev, Alexander P. Tolomeyev, and Mikhail V. Gubanov. The vertical distribution and abundance of Gammarus lacustris in the pelagic zone of the meromictic lakes Shira and Shunet (Khakassia, Russia). Aquatic Ecology, 44(3):531–539, September 2010. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
010-9329-5.

Zamora-Terol:2021:MEH

[ZTNW21] S. Zamora-Terol, A. Novotny, and M. Winder. Molecular evidence of host-parasite interactions between zooplankton and syndiniales. Aquatic Ecology, 55(1):125–134, March 2021. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-020-09816-3.

Zurek:2006:CPW

[Żur06a] Roman Żurek. Chemical properties of water in a flooded opencast sulphur mine. Aquatic Ecology, 40(2):135–153, June 2006. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-005-9024-0.

Zurek:2006:ZFO

[Żur06b] Roman Żurek. Zooplankton of a flooded opencast sulphur mine. Aquatic Ecology, 40(2):177–202, June 2006. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-005-9002-6.

Zwolsman:1993:DPT

[ZV93] J. J. G. Zwolsman and G. T. M. Van Eck. Dissolved and particulate trace metal geochemistry in the Scheldt estuary, S. W. Netherlands (water column and sediments). Aquatic Ecology, 27(2-4):287–300, June 1993. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://

link.springer.com/article/10.1007/BF02334792.

Zalewski:2011:AAG

[ZWF11] Alexandra Zalewski, Nicole D. Wagner, and Paul C. Frost. Antibiotics affect the growth responses of Daphnia magna to poor food quality. Aquatic Ecology, 45(4):493–504, November 2011. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-011-9370-z.

Zhao:2019:EUB

[ZWJ19] Tian Zhao, Xiaoyi Wang, and Jianping Jiang. Effects of urea on behavior and functional traits of Asiatic toad (Bufo gargarizans) tadpoles. Aquatic Ecology, 53(1):9–19, March 2019. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-018-9669-0.

Zheng:2020:SFA

[ZWY20] Xiafei Zheng, Yan Wang, and Qingyun Yan. Size-fractioned aggregates within phycosphere define functional bacterial communities related to Microcystis aeruginosa and Euglena sanguinea blooms. Aquatic Ecology, 54(2):609–623, June 2020. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125

(electronic). URL https://link.springer.com/article/
10.1007/s10452-020-09762-0.

Zeng:2022:EDC

[ZWY^{+}22] Qinghui Zeng, Zhengling Wei, Chai Yi, Yongfeng He, and Mingzhong Luo. The effect of different coverage of aquatic plants on the phytoplankton and zooplankton community structures: a study based on a shallow macrophytic lake. Aquatic Ecology, 56(4):1347–1358, December 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-022-09970-w.

Zhou:2024:RPF

[ZWY^{+}24] Bowei Zhou, Yu Wang, Jiaqi You, Feilong Li, Wei Gao, Qingping Du, Yuan Zhang, and Fen Guo. Response of periphyton fatty acids to experimental changes in aqueous phosphorus. Aquatic Ecology, 58(3):917–929, September 2024. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-024-10115-4.

Zhang:2015:LEO

[ZXH15] Min Zhang, Jun Xu, and Lars-Anders Hansson. Local environment overrides regional climate influence on regime shift in a north temperate lake. Aquatic Ecology, 49(1):105–113, March 2015. CODEN AQECF9. ISSN 1386-2588 (print),

1573-5125 (electronic). URL https://link.springer.com/
article/10.1007/s10452-015-9509-4.

Zhao:2009:ETP

[ZXZ09] Lan-Lan Zhao, Yi-Long Xi, and Chun-Wang Zha. Effects of three phthalate esters on the life-table demography of freshwater rotifer Brachionus calyciflorus Pallas. Aquatic Ecology, 43(2):395–402, June 2009. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL https://link.

springer.com/article/10.1007/s10452-008-9179-6.

Zhang:2022:RDP

[ZYL22] Yingqiu Zhang, Zonghe Yu, and Jie Li. Regionally divergent patterns of grass carp relative abundance, feeding habits and trophic niches in the subtropical Pearl River basin. Aquatic Ecology, 56(3):525–541, September 2022. CO-DEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-021-09923-9.

Zhao:2014:AWI

[ZYS14] Jiyan Zhao, Yangheshan Yang, and Shucun Sun. Artificial warming increases size at metamorphosis in plateau frogs (Rana kukunoris) in the presence of predators. Aquatic Ecology, 48(4):423–434, December 2014. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-
014-9495-y.

Zhou:2016:FDI

[ZZL16] Lei Zhou, Lei Zeng, and Guifeng Li. Fish density increases from the upper to lower parts of the Pearl River Delta, China, and is influenced by tide, chlorophyll-a, water transparency, and water depth. Aquatic Ecology, 50(1):59–74, March 2016. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10452-015-9549-9.